A Complete Bibliography of Publications in Science of Computer Programming

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

12 April 2024 Version 2.99

Title word cross-reference

1 + 118 [2359]. 10 [91]. $102.50 [267, 268]. 11 [95]. 12 [101]. 13 [103]. 14 [107]. 15 [112]. 16 [117]. 18 [125]. 19 [131]. 2 [43]. 20 [147]. 21 [160]. 22 [178]. 23 [190]. 24 [242]. 3 [47]. 3/2 [693]. $31.25 [320]. 4 [54]. 5 [61]. 6 [66]. 7 [72]. $74.25 [318]. 8 [77]. ^{+} [2065, 2428]. ^{2} [466]. ^{3} [1492]. ^{cloud} [2810]. ^{th} [2800]. _{1} [944]. fun [1352]. _{SCR} [1723]. B [2642]. M [1894]. CSP ||B [1280]. ΔIC [879]. E3 [523]. j [1050, 1534]. k [1890, 2529]. λ [146]. μ [2831, 2579, 585]. n [1354]. π [716, 574]. πZ [1467]. Z [530].

-calculus [2579, 716, 574, 146]. -driven [1507]. -functors [1894]. -induction [2529]. -induction-based [1890]. -protocol [1354].

-selected [669].

.NET [1085, 1074].

0 [195, 291, 202, 572, 183, 515, 265, 155, 192, 231, 243, 514, 245, 1980, 295, 156, 151, 206, 146, 203, 213, 289, 285, 286, 287, 219, 244, 207, 246, 182, 229, 153, 321, 296, 217, 208, 297, 193, 180, 298, 67, 204, 267, 268, 194, 216, 320, 181, 317, 145, 214, 266, 269, 397, 169, 574, 390, 215, 173, 1976, 227, 185, 575, 226, 318, 270, 201, 288, 166, 205, 196, 284, 230, 247, 248, 249, 250, 524]. 0-06-042225-4 [180]. 0-12-064460-6 [214]. 0-12-274060-2 [193]. 0-13-021999-1 [201]. 0-13-167080-8 [185]. 0-13-199936-2 [265]. 0-13-204108-1 [315]. 0-13-284027-8 [219]. 0-13-444910-X [284]. 0-13-484197-2 [202]. 0-13-536047-1 [514]. 0-13-629031-0 [195]. 0-13-629122-8 [246]. 0-13-640491-X [285]. 0-13-640509-6 [286]. 0-13-640525-8 [287]. 0-13-7248-0301-7 [194]. 0-13-728866-2 [297]. 0-13-729773-4 [155]. 0-13-729905-2 [67]. 0-13-730334-3 [192]. 0-13-730409 [227]. 0-13-820259-1 [244]. 0-13-823279-2 [217]. 0-13-850587-X [270]. 0-13-853581-7 [229]. 0-19-503691-3 [169]. 0-19-851973-7 [269]. 0-19-853757-1 [209]. 0-19-853757-3 [209]. 0-19-853814-7 [298]. 0-19-859659-6 [213]. 0-19-859664-2 [228]. 0-19-859667-7 [228]. 0-201-17448-0 [250]. 0-201-17536-3 [204]. 0-201-17812-5 [296]. 0-201-17825-7 [247]. 0-201-18053-7 [248]. 0-201-18483-4 [208]. 0-201-19240-3 [166]. 0-201-41633-6 [230]. 0-201-41661-1 [288]. 0-201-54364-8 [320]. 0-205-10450-9 [183]. 0-262-02278-8 [196]. 0-262-08171-7 [205]. 0-262-15038-7 [295]. 0-262-54058-4 [318]. 0-262-57074-2 [266]. 0-262-6-1058-2 [243]. 0-273-02673-9 [206]. 0-3-540-18657-3 [216]. 0-387-97064-9 [291]. 0-444-87132-2 [268]. 0-444-87267-1 [267]. 0-471-84899-9 [153]. 0-471-91201-8 [145]. 0-471-91282-4 [151]. 0-471-920274 [203]. 0-471-92139-4 [231]. 0-471-92741-4 [317]. 0-471-97646-6 [572]. 0-521-26896-6 [146]. 0-521-31839-4 [146]. 0-521-33429-2 [182]. 0-521-34306-2 [173]. 0-521-34632-0 [181]. 0-521-39196-2 [321]. 0-521-42036-9 [390]. 0-521-43457-2 [397]. 0-521-57651-2 [510]. 0-521-58388-8 [524]. 0-521-63107-6 [575]. 0-521-63124-6 [521]. 0-521-64320-1 [574]. 0-56067-5 [515]. 0-674-15626-9 [316]. 0-697-06841-2 [226]. 0-7167-8198-0 [245]. 0-7167-8243-X [396]. 0-7248-0274-6 [215]. 0-8176-3530-0 [319]. 0-86341-095-2 [156]. 0-86341-096-0 [156]. 0-86341-097-9 [156]. 0-86341-098-7 [156]. 0-86380-086-6 [289]. 0-932633-10-2 [207]. 0-937073-05-9 [191]. 0-937073-06-7 [191]. 0-937073-18-0 [249]. 0.17 [1081]. 0.y.z [2600]. ’09 [1480, 1355]. 0x [1466]. 1 [2398, 2393, 2438, 2472, 2522, 2416, 2409, 2448, 2466, 2433, 2443, 2516, 2495, 2477, 2557, 2594, 2621, 2538, 2531, 2574, 2588, 2548, 2565, 2615, 2610, 2605, 2662, 2695, 2717, 2636, 2629, 2677, 2689, 2649, 2670, 2706, 2700, 19, 514, 244, 218, 172, 267, 574, 201, 288, 573]. 1-57586-098-8 [573]. 1-85058-077-4 [218]. 1-85728-362-7 [516]. 10th [20, 2256]. 11 [250]. 11th [2031]. ’12 [1759]. 12.95 [142]. 128.00 [179]. 12th [2682, 1809, 2426]. 13 [141]. 13-729104-3 [141]. 13th [2687]. 14th [2694, 2686]. 15 [2421, 2460, 2508, 2486]. 15.95/$27.95 [201]. 15th [2727]. 17th [2875]. 18th [2383]. 19.50/$32.50 [214]. 19.50/$34.50 [321]. 19.95 [516]. 1981/82 [19]. 1983 [48, 49]. 1984 [79]. 1995 [2886]. 1st [21].

2 [202, 265, 243, 207, 182, 321, 217, 297, 193, 67, 171, 268, 266, 397, 173, 226, 2057, 516, 192]. 2.0 [1237, 2733, 2271]. 2003 [1025, 837]. 2004/2005 [1254]. 2006 [1058]. 2007 [1259]. 2008 [1286]. 2009 [1373, 1346]. 201 [143]. 201-14239-9 [143]. 2010 [1702, 1858, 1392]. 2011 [1504, 1941]. 2012 [1863, 1687, 1593, 1666, 2063, 1746, 1692]. 2013 [1716, 2068, 1914, 1919, 1726, 1892]. 2014 [2058, 2024, 1954, 1960, 2018]. 2015 [2143, 2239, 2156]. 2016 [2226, 2219, 2367, 2272]. 2017 [2283, 2391]. 2019 [2398, 2393, 2535, 2582, 2562, 2688]. 2020 [2438, 2472, 2522, 2416, 2421, 2409, 2448, 2460, 2466, 2433, 2443, 2516, 2495, 2508, 2477, 2486, 2697, 2665]. 2021 [2557, 2594, 2621, 2538, 2531, 2574, 2588, 2548, 2565, 2615, 2610, 2605, 2730, 2737, 2693]. 2022 [2867, 2662, 2695, 2717, 2636, 2629, 2677, 2689, 2649, 2670, 2706, 2700, 2875]. 2023 [2751, 2777, 2724, 2760, 2770, 2740]. 2024 [2801, 2821, 2868, 2840, 2855]. 21st [2562]. 22nd [2685, 2535]. 23rd [1204, 2272, 2730, 2697, 1215]. 24 [471]. 24.95 [152]. 24th [1436, 1360]. 25 [2800]. 25- [2800]. 25010 [2642]. 25th [2851, 1400].

26.95 [170]. 26th [1416]. 29th [2775]. 2nd [516, 151, 145, 230, 35].

3 [141, 184, 192, 154, 299, 110, 172, 168, 204, 171, 216, 1967, 271, 169, 730, 166]. 3-540-13721-1 [172]. 3-540-17030-8 [154]. 3-540-18008-7 [168]. 3-540-18540-2 [171]. 3-540-19577-7 [299]. 3-540-43447-X [730]. 3-540-51799-5 [271]. 3-540-96539-4 [184]. 3-540-97061-4 [290]. 3-7643-3530-0 [319]. 32.50 [144]. 35th [2625]. 39 [587]. 3D [1139]. 3rd [2880, 2885, 1173, 2881].

4 [184, 155, 231, 151, 218, 208, 180, 317, 318]. 40 [600]. 41 [610]. 42 [624]. 43 [633]. 4th [1525, 56, 2882].

5 [515, 1287, 296, 271]. 52.00/DM [319]. 5th [2884, 1143].

6 [572, 146, 213, 289, 286, 214, 215, 575, 230]. 60.75 [389]. 61131 [1967]. 61131-3 [2584, 1967]. 6th [2097, 2276].

7 [299, 295, 156, 229, 168, 298, 194, 269, 205, 247, 248]. 730409 [227]. 75 [1292]. 7th [2665, 9].

8 [154, 287, 219, 246, 2481, 320, 145, 185, 196, 524, 573]. 83g [19]. 8th [2178].

9 [291, 183, 206, 153, 143, 390]. 9.95 [167]. ’90 [2880, 2881]. 920274 [203]. ’95 [2886]. 96d [471].

AADL [1907, 1906]. Aaron [320]. AATT [2265]. AbC [2544]. abduction [1296]. ABS [2714]. absences [2361]. absolute [693, 2729]. Abstract [2404, 294, 1434, 2497, 703, 1392, 1685, 1457, 139, 487, 844, 684, 954, 725, 2407, 952, 569, 116, 2060, 571, 2632, 535, 2401, 446, 1718, 506, 379, 548, 486, 369, 376, 842, 370, 581, 1208, 1686, 1944, 1945, 1478, 862, 949, 1483, 654, 1673, 489, 1240, 784, 1694, 2062, 868, 1692]. Abstracting [685, 2502]. Abstraction [84, 1439, 352, 2604, 2163, 2306, 1255, 1651, 2073, 1910, 1568, 194, 162, 2355, 1769, 1281, 840, 144]. abstractions [1077, 1648, 1725, 2235, 2131, 2295]. abstractor [409]. ABZ [2497, 2665, 1692, 2058, 2226, 1392]. Academic [214, 1173]. Academics [1814]. Accelerating [2280]. acceleration [1685]. acceptors [223, 276]. Access [1319, 1478, 2864, 2370, 1565, 992, 1458]. Access-based [1478]. accesses [1208, 1851]. accessibility [1604, 2807]. accessible [2328]. accessing [1713]. accommodates [789]. accumulations [372, 557, 1449]. achieve [959]. Achieving [1915, 803, 1361]. Achilles [1059]. ACM [1204, 1436, 1687, 1215, 1360, 1400, 1416, 320, 2018]. ACME2012 [1814]. ACP [1273]. acquisition [336, 1012]. across [1170, 1847]. ACSmt [2809]. ACSR [615]. ACSR-VP [615]. action [677, 1497, 1596, 1032, 1183, 2352, 903]. action-semantics [677]. actions [339, 1587]. activated [863]. active [799, 461, 2379, 1352, 2290, 2748, 1491, 1247, 1720, 2564]. activities [1616, 2373]. activity [1937, 1517, 2581, 2500]. Actor [1210, 2308, 2418, 1987, 1822, 1545, 2194, 1547]. Actor-based [1210, 2308, 2418]. actors [1823, 1821, 152]. actual [1310]. acyclic [200, 890, 695]. Ad [1190, 1986, 772]. Ada [168, 53, 2592]. Adaptability [1984, 959, 1440, 1293]. adaptable [1442, 2115, 1748]. Adaptation [2034, 1562, 1706, 570, 928, 897, 2297, 1381, 2254, 1798, 803, 1304]. adapters [877]. adapting [2405, 1850]. Adaptive [1388, 2452, 2132, 2570, 2524, 2462, 1471, 1857, 1704, 2034, 2236, 2310, 1011, 748, 2639, 2651, 1379, 2827, 469, 1012, 1774, 1879]. Adaptiveness [2563]. add [773]. Adding [1826, 896]. Addison [296, 208, 204, 320, 143, 288, 166, 230, 247, 248, 250]. Addison-Wesley [296, 208, 204, 320, 143, 288, 166, 230, 247, 248, 250]. Addison-Wesley/ACM [320]. adequacy [2506, 2604]. Adequate [59, 969, 1757]. adhesive [1895]. Adjoint [1500]. administration [701, 753]. admitted [2633]. Adore [1444]. ADS [2797]. AdSiF [2298]. Advanced [575, 1071, 774]. Advances [2283, 829, 2246, 2723, 2520, 1830, 1735, 2106]. adventure [1010]. advice [941]. AErlang [2305]. affect [2843, 549]. affects [2807, 2012]. affine [1907]. after [1038]. against [953, 2735, 1189, 2676]. Agda [2475]. age [316]. age-new [316]. agenda [590]. agent [1426, 864, 1420, 2526, 1479, 2425, 645, 1425, 2721, 1427, 1822, 1188, 2160, 2066, 826, 1234, 1309, 2766, 1006, 2298, 1423, 930]. agent-based [2526, 1309]. agent-oriented [1420, 645, 1234, 1423]. agents [2653, 1415, 2223, 773, 710, 586, 1007, 1821]. Agg [1894]. agglomerations [2390]. Aggregate [2553, 2820, 2300]. Agha [152]. Agile [2259, 2541, 1775, 2365, 1507, 1791]. agnostic [2733]. agreement [1307]. AI [152, 2873, 2849, 1012]. AI-enabled [2849]. aim [1103]. air [1756]. aircraft [2213, 1431]. airplane [1888, 2577]. ALA [1471]. Alan [168]. ALDEBARAN [467]. Alexander [244]. Algebra [70, 927, 1273, 466, 1228, 869, 999, 513, 622, 982, 1195, 1621, 1501, 348, 2684, 1675, 1247, 1141, 730, 781, 1096, 2427, 721, 1485, 374]. Algebraic [734, 98, 359, 62, 1910, 307, 1220, 241, 274, 164, 188, 2471, 1596, 583, 926, 1843, 563, 187, 702, 271, 338, 179, 205]. algebraically [324, 1498]. Algebras [1599, 382, 2198]. ALGOL [224]. ALGOL-like [224]. Algorithm [419, 427, 39, 26, 4, 82, 69, 58, 260, 121, 632, 225, 1243, 1376, 901, 2411, 693, 197, 108, 1617, 1217, 2492, 235, 200, 951, 1644, 1474, 1690, 2194, 1109, 198, 2744, 405, 1266, 163, 479, 690, 1966, 1206, 771, 2400, 1963, 2023, 1267, 2513, 1622, 696, 695, 177, 1998, 529, 99, 470, 196]. Algorithmic [2035, 2159, 49, 173]. algorithmics [1496, 166]. Algorithms [124, 388, 13, 118, 92, 97, 421, 2571, 1138, 2289, 233, 358, 2407, 770, 2471, 2091, 2596, 517, 2469, 1223, 704, 446, 464, 372, 1441, 902, 1701, 2243, 1359, 1347, 237, 403, 364, 1365, 454, 1931, 424, 2336, 2709, 63, 351, 148, 436, 1105, 265, 315, 396, 145, 201]. Alias [1793, 579]. aliasing [537]. allocated [1667]. allocation [2397, 714, 2279, 712]. Alloy [2259, 1692, 2619, 2497, 2537, 1700, 1392, 2437, 2698, 1699]. AllSynth [2796]. almost [1571]. almost-certain [1571]. along [1618]. AltaRica [1841]. Alternating [1188]. Alternating-time [1188]. alternation [672, 2179]. alternation-free [672]. Alternative [12, 39, 19]. Alvey [295]. Alvey-Britain [295]. always [698]. Alzheimer [2823]. AmbieGen [2798]. Ambient [1487, 1492, 1309]. ambients [684, 868]. Ambiguity [2022, 901, 1169, 2480, 1165]. American [318, 2851]. Amir [1172]. among [1121, 2729, 1823, 1017]. amortised [2413]. Amstel [265]. Amsterdam [388, 267, 268, 1645]. analyser [2790]. analysers [2288]. analyses [2001, 1224, 2235, 579, 1765, 2349, 1903, 2763, 1239, 948]. Analysing [97, 1519, 127, 2016, 1168]. Analysis [1896, 135, 104, 681, 837, 1641, 585, 1904, 229, 2775, 1941, 532, 1465, 2732, 947, 919, 2854, 484, 2440, 662, 476, 1077, 1541, 875, 2232, 2201, 1971, 2036, 2212, 2091, 965, 1216, 2086, 2236, 2787, 2190, 1725, 2729, 2578, 821, 1804, 892, 1479, 482, 2712, 599, 2632, 2231, 2185, 1308, 2302, 2397, 1856, 1858, 982, 686, 2450, 1943, 974, 1514, 2151, 2379, 1685, 709, 1476, 2170, 841, 411, 1523, 2111, 2353, 544, 2263, 2675, 1553, 1887, 2040, 2446, 2835, 2290, 2878, 2736, 2092, 2162, 1825, 1358, 321, 2647, 2850, 615, 488, 1522, 2382, 1552, 1636, 1569, 2836]. analysis [1819, 1894, 708, 1261, 1680, 1928, 1944, 1945, 281, 2413, 860, 1344, 149, 371, 779, 563, 1203, 1478, 1018, 1851, 2773, 2570, 2251, 1401, 771, 1103, 845, 2514, 1211, 1508, 374, 1664, 578, 1965, 1580, 1727, 1608, 1728, 440, 1060, 1938, 1483, 1665, 351, 2205, 537, 1884, 301, 2221, 697, 2758, 1668, 2202, 2863, 2261, 2676, 1998, 2542, 2755, 1240, 2722, 1794, 396, 1178, 246, 249]. analytical [830]. analytics [2711, 1789, 2011, 1540]. analyze [2862]. analyzer [1049]. Analyzing [1169, 1878, 2094, 2112, 2322, 1637, 2359, 1857, 2843, 1542]. anchored [1610]. ancilla [2628]. Andrei [290]. Andrew [168, 524]. Android [2540, 2263, 2587, 2265]. Andrzej [216]. Anemone [2545]. angelic [980, 981]. animated [747]. animation [459, 1393, 2835, 196]. AnimUML [2835]. Anne [315]. annotated [1161]. annotation [2612, 2783, 1799]. annotation-based [2612]. annotation-driven [1799]. annotations [1707, 2451]. Announcement [21, 35, 73, 292]. Announcements [50, 56]. annual [232]. anomalies [888]. Anomaly [140, 870]. Anonymous [1938, 1661]. Ansart [268]. ANSI [2063]. ant [1427]. anti [1640]. anti-patterns [1640]. Antony [193]. any [762]. any-time [762]. AOP [1526, 1078]. AOSE [1421]. API [2083, 1114, 2540, 2266, 1237, 2744, 2756]. API-call [2266]. APIs [2116, 2510, 2769, 2251]. App [2540, 2744]. Appel [524]. applicability [2008]. applicable [1015]. Application [2702, 1239, 947, 1282, 457, 1216, 1056, 1262, 583, 1425, 1986, 1985, 453, 603, 1894, 591, 845, 2514, 1508, 1764, 1788, 1059, 2836]. application-level [1425]. Application-oriented [2702]. Applications [1607, 1854, 2580, 2563, 118, 1480, 1994, 1088, 1946, 1570, 1828, 1442, 570, 1528, 2262, 1688, 2115, 2280, 2291, 2248, 1489, 2469, 1089, 2138, 786, 929, 2793, 1693, 2257, 1691, 2263, 1986, 1291, 1044, 2882, 2320, 1320, 1972, 889, 2886, 1027, 2826, 1515, 2476, 1562, 1148, 723, 1062, 1125, 1234, 1309, 1802, 1491, 1664, 1879, 775, 846, 993, 1847, 1560, 1424, 2221, 1270, 1180, 142, 2536, 2613, 1794, 2572, 193]. applicative [2188]. Applied [1204, 1436, 1215, 1360, 1400, 1416, 541, 2017, 216, 1967, 2643]. Applying [864, 1631, 2828, 2577, 1305, 2485, 2829, 2103, 2652, 2768, 1104, 974]. Approach [51, 2223, 1640, 96, 52, 2592, 973, 1570, 2015, 2234, 525, 1442, 42, 2241, 2006, 1202, 2115, 864, 2787, 2248, 2405, 1297, 1505, 821, 2285, 1069, 2463, 1935, 2397, 1561, 583, 2128, 2108, 2274, 2793, 1327, 1575, 748, 2330, 2111, 1772, 2651, 1427, 559, 2266, 300, 879, 198, 615, 2796, 1153, 1310, 1501, 2858, 1709, 2826, 1893, 2655, 1848, 1988, 2476, 1449, 1604, 2033, 1148, 2381, 1010, 1062, 469, 2214, 994, 2555, 1864, 552, 1234, 1132, 2830, 1878, 2819, 761, 2783, 1705, 1732, 2489, 2756, 755, 641, 1332, 1240, 560, 226]. Approaches [588, 2838, 998, 1591, 1754, 1512, 179]. appropriate [1476]. Approximate [369, 1842, 2280, 376]. approximated [1921]. Approximately [1689]. Approximating [533]. approximation [693, 2078, 1842, 2112]. approximations [950, 1686]. apps [2247, 1772, 2587, 2265]. April [2438, 2557, 2662, 669, 442, 1172, 2751]. APSEC [2272]. Arbib [184, 179]. Arbiters [322]. arc [479, 1232]. arc-consistency [479]. Architecting [1513, 1295]. Architectural [799, 2046, 777, 2578, 511, 867, 2666, 636, 1651, 551, 1183, 1864, 1211, 1852, 2465, 2016, 2304, 2012]. architecturally [2390]. Architecture [2241, 1648, 1795, 2049, 1088, 1828, 1905, 274, 492, 2034, 2578, 2781, 1840, 1308, 1807, 1614, 887, 1603, 637, 1091, 1310, 1201, 1539, 822, 2033, 937, 1249, 1492, 2050, 1307, 1309, 1508, 1652, 817, 888, 993, 2047, 992, 1424, 2430, 641, 2144, 915, 2012, 1336, 1647, 1904]. architecture-based [2034, 1201]. Architecture-Driven [1904]. Architecture-level [2241, 1648]. Architectures [661, 1989, 385, 1143, 1349, 2044, 2303, 1907, 2370, 663, 1797, 604, 816, 815, 886, 989, 1200, 1306, 490, 821, 995, 765, 1827, 2861, 1198, 1522, 896, 2388, 1543, 818, 891, 1072, 2643, 621, 762, 871]. Arcot [389]. areas [1761]. arena [1404]. argument [1709]. argument-based [1709]. Arithmetic [31, 2407, 1942, 362, 1699, 2173]. ARM [859]. ARMv8 [2406]. Array [1120, 446, 1208]. Arrays [425, 2070, 437, 1927, 2458, 2489]. arrows [2134, 558]. Arsac [214]. art [1340]. artefact [1294]. Article [2183]. artifact [1782]. artifacts [2174, 804]. Artificial [2776, 266]. ARV [1471]. As-if-serial [1122]. Asf [1768]. Asia [2272]. Asia-Pacific [2272]. Asl [106]. Asl/F [106]. ASM [2234, 1695, 1276]. ASMs [2059, 1693, 2228, 2614]. ASP [1352]. Aspect [943, 1202, 2115, 944, 821, 2060, 1071, 939, 678, 1123, 1883, 1304, 1384, 1450, 1451, 1560, 1028, 1289, 1153, 2004]. Aspect-oriented [943, 1202, 2115, 944, 821, 2060, 939, 678, 1123, 1384, 1450, 1451, 1560, 1028, 1153, 2004]. AspectJ [1475]. AspectMaps [1526]. Aspects [2883, 2694, 1702, 2256, 1328, 1969, 1194, 2031, 1487, 2515, 1378, 2687, 1653, 2686, 1941, 2426, 2727, 2530, 2010, 1221, 1224, 892, 940, 513, 942, 2702, 1229, 810]. aspectual [1750, 1382]. assembler [1061, 738, 1168]. assembly [1366, 1857]. Assertion [1999, 450, 789, 1450, 2205]. Assertion-based [1999]. Assertional [26, 1469]. Assertions [96, 1456]. Assessing [1309, 2873]. assessment [1841, 655, 1640, 1750, 2099, 2109, 1300, 889, 2619, 986, 2757]. assessments [2642]. assign [2214]. assistance [2377]. Assistant [2865, 2845]. assistants [2750]. assisted [838, 2010, 2229]. assume [1658, 2468]. assume-guarantee [1658, 2468]. assumption [378, 2304]. assumption-commitment [378]. assurance [2811, 2667]. assured [854]. Asynchronous [2147, 2818, 2376, 792, 1377, 422, 1918, 1987, 1368, 2134, 1108, 2040, 2215, 1623, 432, 954, 2731]. asynchronously [322, 2569]. ATAM [822]. ATL [1079, 1037, 2114]. atomic [332, 474, 934, 851]. atomicity [1067, 1064]. Atomizer [1067]. ATS [1452]. attack [2300, 2157]. attack-resistant [2300]. attacks [1192, 2773, 1549, 2676]. attention [1493, 2743]. attention-based [2743]. attestation [1101, 1938]. attractors [2375]. Attribute [1749, 1162, 2462, 2343, 65, 1977, 1765, 237, 2069, 2305, 1481, 1764, 1163]. attribute-based [2462, 2305]. attributed [1027]. attributes [2121, 2807]. audio [1013]. Augmentation [2825]. Augmenting [2784]. August [2472, 2594, 2695, 2687, 2777]. Australia [21]. Authentication [350]. Author [81, 220, 345, 402, 414, 433, 587, 600, 610, 624, 633, 649, 659, 674, 688, 699, 705, 718]. authoring [1627]. authority [1630]. authorization [1102]. authorship [2359]. AutoCRUD [2309]. Automata [2704, 671, 20, 2854, 895, 1282, 1369, 1949, 2735, 2146, 2242, 2395, 685, 1353, 2479, 1380]. Automated [2447, 2350, 1161, 1363, 2712, 1603, 2584, 545, 2413, 1556, 2165, 2271, 2638, 2454, 923, 619, 2564, 2738, 2591, 2389, 2642, 446, 2507, 1651, 2481, 2619, 2348, 1660, 1307, 2819, 1713, 1756, 1567, 2781, 2824, 2038, 1572, 1752, 1934]. Automatic [2240, 303, 294, 1461, 620, 717, 1312, 1033, 1972, 76, 149, 2345, 949, 1313, 1811, 537, 1453, 2431, 739, 1220, 1770, 2116, 1710, 462, 2844, 1855, 644, 800, 2806, 2832, 2309]. Automatically [670, 1576, 41, 2768, 1686]. Automating [1250]. Automation [2440, 616, 611]. automotive [2201, 1889, 1513, 1999]. autonomic [2147]. autonomous [751, 2128, 2788, 2721, 2798, 2161]. Autoprojection [303, 294]. auxiliary [283]. availability [1194, 960]. available [1991, 2613]. avatars [2555]. avionic [541]. AVoCS [1567, 1934, 2156]. AVoCS’11 [1572]. AVoCS’12 [1752]. AVoCS’14 [2038]. Avoid [140]. Avoiding [1655, 2004, 1640]. aware [2519, 1688, 2115, 1489, 1986, 1985, 2541, 1515, 2273, 1903, 2056, 1922, 1879, 1535, 2617, 915, 2045]. awareness [1506, 1494]. Axiomatization [1105]. Axiomatizing [355]. Axioms [1803, 1950, 1244, 1861, 2684]. Azema [268]. Azure [2620].

B [2127, 2803, 219, 208, 388, 271, 397, 1692, 1939, 1901, 2127, 2213, 2447, 2519, 2497, 2159, 1395, 2231, 2666, 1392, 2490, 2064, 2065, 1155, 1571, 2063, 1435, 2437, 2655, 673, 1698, 2505, 2728]. Bach [2545]. Back [365, 120]. backbone [2432]. backbones [990]. Background [316]. Backtracking [282, 444]. backup [965]. Backward [1686, 434, 1673]. Baeten [730]. baffler [132]. Bag [361, 1363]. Bakery [1474]. Balanced [314]. balancing [1807]. Banana [1485]. band [1646]. bands [1596, 1365]. Bangalore [56]. Banzai [1362]. Barbara [144]. Barcelona [20]. bargaining [1307]. bargaining-specific [1307]. Barker [1391]. barter [826]. base [1789, 2871, 1664]. Based [2183, 2852, 2212, 1647, 1340, 17, 124, 2665, 45, 94, 498, 1137, 1810, 532, 2462, 2839, 1220, 796, 2277, 2745, 662, 2289, 2553, 1442, 2418, 2742, 476, 2006, 1202, 465, 1374, 2222, 1897, 746, 2747, 1906, 2262, 747, 478, 2115, 518, 1669, 2785, 1674, 864, 1912, 2034, 1797, 2236, 2648, 2526, 2787, 1723, 1490, 2248, 1890, 1113, 2753, 1889, 1212, 1805, 1272, 1297, 959, 703, 1615, 1224, 821, 1751, 1731, 787, 2456, 1628, 1083, 2153, 1298, 2811, 2872, 2870, 1935, 1632, 1308, 751, 1807, 2759, 2343, 686, 2450, 2108, 1425, 1322, 2540, 991, 2579, 1046, 2786, 1603]. based [788, 1970, 748, 2170, 2612, 1761, 1773, 2330, 2111, 1214, 544, 1264, 878, 2804, 2798, 2331, 1592, 2853, 2260, 2073, 1350, 2266, 2797, 1145, 453, 1532, 1037, 2850, 615, 1735, 1522, 2796, 1972, 2464, 2744, 2500, 2172, 750, 803, 798, 1201, 611, 822, 1709, 1124, 1302, 1893, 1894, 1568, 1182, 1271, 2198, 2627, 642, 2794, 2033, 1734, 1785, 2784, 890, 860, 2699, 1446, 1999, 2305, 2849, 1203, 1001, 1478, 592, 871, 2214, 2335, 1183, 2049, 2358, 501, 1812, 2076, 552, 2587, 1534, 1321, 2514, 1821, 1013, 1309, 374, 666, 1652, 1210, 2830, 1618, 2764, 175, 1963, 1555]. based [2385, 2810, 2140, 2048, 2766, 1072, 2110, 1855, 2007, 992, 619, 755, 1673, 2863, 2676, 2659, 2757, 2813, 2375, 2738, 2613, 2743, 2045, 2542, 2556, 2728, 2308, 1240, 2722, 1794, 621, 319]. Basel [319]. bases [1683]. Basic [2020, 115, 2311, 598, 1533, 1701]. basic-block [1533]. Basili [226]. basis [936, 1992, 1843, 136, 807, 1884]. Batcher [132]. Bayesian [2583]. BBS [1366]. BCS [299]. BDD [2796]. BDD-based [2796]. BDI [2653]. be [2873]. bearing [1063]. beginning [745]. behalf [156]. Behavior [784, 1831, 511, 2254, 1131, 535, 2362, 1753, 1103, 2318, 621, 1283]. Behavior-propagating [1831]. Behavioral [2363, 1828, 2020, 1381, 1284, 1827, 2141, 2311]. behaviors [2270]. behaviour [1564]. Behavioural [1390, 409, 767, 2709, 239, 1601, 1907, 2415, 2041, 2348, 1923]. Being [2506]. benchmark [2570]. Benchmarking [2110, 2749]. benefit [2559]. Bengt [298]. Berlin [184, 375, 154, 669, 172, 168, 171, 216, 179, 730]. Berlin/Heidelberg/New [184, 179]. Bertrand [195]. beside [1129]. best [1104, 1919, 1878]. better [2504]. between [2679, 1601, 1455, 921, 2103, 816, 989, 995, 1718, 1639, 1310, 1634, 857, 1508]. Beyond [332, 2453, 611]. BEZ235 [2823]. BGP [754]. biconnected [405]. bidirectional [578]. Bidirectionality [2446]. Bidirectionalization [1929]. big [2503, 2140, 2476, 2661]. bigraphs [2653, 1559, 2703]. bin [693]. binaries [2349]. binary [596, 123, 1551, 1239, 2356, 1684, 880, 2676]. binary-search [123]. binding [978, 149]. bio [2571]. bio-inspired [2571]. biometrics [1104]. BIP [1915]. bipartite [890]. biproduct [1501]. biproduct-oriented [1501]. Bird [202, 2332, 1]. Birkhäuser [319]. bisection [1644]. Bishop [230]. bisimilar [2547]. Bisimilarity [2325, 1757, 1867, 2493]. bisimulation [235]. BITAM [816]. bits [1881]. black [755]. black-box [755]. Blackboard [1986]. Blackwell [142]. blaming [1970]. Blikle [216]. bloat [1248]. block [1533, 691, 301, 619, 508]. block-based [619]. block-diagram [691]. block-structured [301]. blocking [1230, 2224]. Blockly [1828]. blocks [2408, 1870, 851, 1967]. Blueprint [1637]. Board [2245, 2255, 2264, 651, 660, 668, 675, 680, 689, 700, 726, 731, 694, 706, 743, 752, 759, 768, 776, 782, 793, 808, 814, 823, 828, 836, 847, 855, 866, 874, 882, 893, 904, 911, 918, 924, 931, 938, 945, 955, 961, 968, 975, 984, 988, 996, 1002, 1008, 1016, 1024, 1029, 1034, 1042, 1052, 1053, 1057, 1066, 1070, 1075, 1082, 1087, 1093, 1098, 1106, 1111, 1117, 1126, 1133, 1142, 1149, 1154, 1159, 1171, 1179, 1185, 1235, 1311, 1808, 1813, 1820, 1838, 1839, 1853, 1862, 1868, 1885, 1891, 2021]. Board [2030, 2037, 2043, 2051, 2067, 2081, 2082, 2087, 2095, 2096, 2105, 2113, 2117, 2123, 2126, 2129, 2133, 2137, 2139, 2142, 2148, 2155, 2166, 2171, 2177, 2182, 2192, 2193, 2196, 2203, 2218, 2225, 2237, 2238, 2281, 2282, 2293, 2301, 2312, 2315, 2319, 2324, 2337, 2346, 2351, 2357, 2360, 2364, 2368, 2372, 2378, 2387, 2394, 2399, 2410, 2417, 2422, 2434, 2439, 2444, 2449, 2461, 2467, 2473, 2478, 2487, 2496, 2509, 2517, 2523, 2532, 2539, 2549, 2558, 2566, 2575, 2589, 2595, 2606, 2611, 2616, 2622, 2630, 2637, 2650, 2663, 2671, 2678]. Board [2690, 2696, 2701, 2707, 2718, 2725, 2741, 2752, 2761, 2771, 2778, 2802, 2822, 2841, 2856, 2869]. Bognor [218]. Boiler [670]. Bonding [2180]. Bonn [35]. Book [48, 195, 141, 49, 126, 184, 291, 538, 516, 202, 375, 183, 515, 265, 155, 192, 231, 243, 514, 245, 79, 510, 154, 299, 295, 156, 151, 34, 228, 391, 206, 146, 319, 203, 521, 213, 289, 285, 286, 287, 219, 244, 78, 207, 246, 182, 229, 218, 153, 321, 172, 296, 217, 389, 208, 297, 193, 388, 180, 168, 298, 67, 204, 171, 267, 268, 194, 398, 216, 315, 179, 320, 181, 317, 143, 145, 214, 167, 266, 269, 271, 397, 169, 390, 215, 173, 499, 227, 144, 185, 226, 318, 270, 209, 290, 781, 201]. Book [288, 166, 205, 170, 196, 284, 230, 152, 191, 247, 248, 249, 250, 142, 316, 524, 396]. Books [18, 55]. boolean [2179, 484, 2435, 1777, 2375]. Bootstrapping [1744]. Bornat [141]. boroque [209]. borrow [2628]. both [1104, 789]. Bottom [223, 334, 264]. Bottom-up [223, 334]. Bound [282]. bounded [332, 425, 1368, 1314, 1951, 1264, 1825, 2229, 2187, 2124, 949, 2479, 2355, 1953]. bounding [1943]. bounds [1669, 1120]. BoundWarden [2511]. box [1051, 755]. boxes [1389, 936]. Boyer [696]. BPEL [1320, 1850, 1018, 1006]. BPMN [2302, 1270]. Brace [193]. Brains [184]. Branch [282, 1741]. Branch-and-Bound [282]. Branching [120, 29, 1757, 2493]. Brazilian [2685, 2697, 2383, 2535, 2276, 1546, 1677, 2799, 2338, 2457, 1909, 1746, 1329, 2002, 1448, 1670]. break [2607]. Brian [295]. bridge [995]. Bridging [1601, 989, 1310, 1508, 529, 556, 1761]. Britain [295]. broadcasting [413]. Broadcasts [44]. broken [2175]. Brown [226, 196]. browser [1174]. BSP [2401, 564].

Büchi [291]. bug [1216, 2870, 2570, 2762]. bugs [1401, 2265, 2261]. Building [1939, 1489, 1050, 1870, 712, 747, 2115, 1201, 2076, 1534]. built [110]. built-in [110]. bulk [2401, 2489]. Bulterman [8]. BUnity [1459]. Burns [168]. BURST [2749]. bus [466, 1555]. Bush [288]. business [1065, 816, 2285, 2397, 1325, 970, 1772, 1887, 2316, 2858, 1444, 1181, 971, 972]. Bustard [185]. Byte [1125]. Bytecode [1396, 773, 1533, 2450, 1733, 832, 1266, 1668, 1726, 1666].

C [49, 126, 375, 299, 228, 219, 267, 181, 1660, 227, 730, 247, 248, 249, 250, 142, 1882, 2001, 466, 1736, 742, 2136, 1327, 1196, 1844, 2080, 1205, 1197, 2817, 1044, 1330, 1199, 2042, 1473, 502, 1251, 1174, 1466, 1803, 623, 1128, 1031, 738, 2564, 2144]. C# [1880, 2451]. C-bus [466]. C/C [1128, 1031]. CA [1183]. Cache [535, 2252, 2504]. caches [2370]. CADP [1327, 1108, 1552, 1433]. CAESAR [467]. CAESAR/ALDEBARAN [467]. CafeOBJ [2723]. Calculating [358, 211, 1]. Calculational [517, 560]. Calculi [2138, 608, 1594]. Calculus [5, 1068, 1459, 570, 485, 423, 375, 509, 2499, 1995, 2544, 556, 1019, 431, 2579, 2333, 964, 1352, 825, 1793, 716, 564, 2388, 672, 259, 136, 819, 149, 133, 413, 780, 948, 522, 1190, 1281, 2202, 2602, 2180, 574, 146, 296]. calibration [2135]. Call [9, 20, 80, 232, 272, 485, 2429, 2266, 1588, 1707, 2150, 2404, 1535]. call-by-need [2150]. call-by-value [485, 2429, 1588, 2404]. Call-site [1535]. calling [1535]. Caltech [422]. Cambridge [515, 510, 146, 521, 182, 321, 388, 398, 181, 167, 397, 574, 390, 173, 575, 318, 205, 170, 316, 524]. camera [2135]. Caml [1979]. can [970]. cancelable [1104]. CancerGrid [1628]. Candidate [2639]. canonical [1453]. CAOVerif [1663]. capabilities [2003, 2720]. Capability [1004]. capacity [1882]. capacity-C [1882]. capture [2401, 1425]. capturing [2362]. Card [673, 1335]. cards [350, 785]. Carl [573]. Carling [218]. Carlos [171]. Carol [67]. carrying [831]. cars [2747, 2453]. Case [45, 31, 1366, 2459, 1828, 2561, 1888, 2232, 383, 2006, 2249, 963, 1714, 906, 1723, 1554, 596, 1786, 867, 2185, 1432, 833, 255, 1464, 2109, 2664, 795, 1383, 1753, 2267, 2837, 1351, 1423, 1148, 2805, 1873, 1556, 2384, 187, 1577, 1744, 2251, 1534, 2514, 2385, 2101, 301, 455, 2863, 2659, 1238, 2807, 1422, 1237, 962]. case-based [1534]. cases [2440, 765, 2111]. Castor [2470]. catches [2332]. Categorical [449, 116]. categories [1895]. categorising [2103]. causality [1470]. causally [392]. causally-ordered [392]. cause [2790, 2355]. cause-guided [2355]. caution [2518]. Cave [2152]. CBR [1050]. CC [474, 2768]. CCP [1867]. CCS [355]. CCSL [2286, 1908, 2199, 2542, 2556]. CD [1731]. CD-Form [1731]. cells [2146]. cellular [2479]. centered [2858]. centrality [2553]. Centralized [2242]. Centred [1618]. centres [1413, 609]. centric [1198, 1246]. century [2176]. CenUMs [2077]. Cerro [391]. CERT [1660]. certain [1571]. Certification [1659, 2456, 802, 2846, 1660]. certified [838, 2490, 712]. Certifying [1366, 1947]. cf4ocl [2136]. chain [2546, 545, 1044, 2715]. chains [1511, 1951, 1657]. Chaitin [173]. Chalet [21]. challenge [1112, 1064, 1974]. Challenges [2598, 1789, 1794, 1665, 891]. Championships [1858]. change [655, 2843, 1479, 1880, 473, 2111, 1793, 2320, 755, 217]. changeability [655]. changers [979]. changes [1348, 2781, 2652, 1741, 1834, 1734, 2710]. Changing [1280, 1956]. Channel [1272, 2322, 1430, 1145]. Channel-based [1272, 1145]. channels [1865, 1467]. chaos [2107]. characterisation [2660]. Characterization [2774, 2436, 1462]. characterizations [812]. Characterizing [1062, 2601, 1834]. Chari [268]. Charles [316]. charts [778, 1370, 646]. chatbots [2828]. ChatGPT [2873]. check [1120, 920]. checked [2230]. Checker [93, 2705, 1067, 1549, 2874]. checkers [1937, 2750]. Checking [2084, 427, 120, 2316, 1999, 2295, 1810, 1866, 1860, 1376, 2583, 530, 518, 1841, 1704, 1889, 1226, 1368, 1354, 1241, 380, 717, 1575, 2720, 809, 841, 2194, 2266, 1145, 2229, 2581, 1921, 789, 1859, 581, 2284, 672, 1433, 565, 1984, 2187, 1936, 591, 2348, 2569, 1372, 1851, 1450, 2042, 780, 2513, 2029, 2355, 2529, 455, 971, 2493, 1240, 2722, 1610, 1238]. checkpointing [2243]. checks [2205, 824]. chemotaxis [1415]. Chicago [191, 249]. Chichester [231, 151, 203, 289, 207, 153, 317, 145]. Chin [153]. Chin-Kuei [153]. China [2687]. chip [2340, 1998]. Cho [153]. choreographic [2546]. Choreographies [1916, 1324]. Choreography [2240, 2498, 1212]. CHOReVOLUTION [2498]. chrestomathies [1784]. Chris [521]. Christoph [572]. Chronos [1049]. Churrasco [1178]. Ciao [842]. CIbSE [2851]. CICS [1114, 657]. CIGALE [100]. CIL [2450, 1085]. circles [1917]. circuit [2627, 1927]. circuits [2322, 1108, 362, 1271, 1353, 2179, 954, 2643, 171]. Circular [1027, 11, 2753, 1720]. Circus [2386]. Cjj [623]. Claim [772]. Clarendon [228, 391, 213]. Clark [208]. clash [1464, 158, 2228]. clash-free [2228]. Class [97, 1831, 328, 360, 2371, 1399, 2132, 424, 2865, 1673, 2767]. classes [484, 2079, 1533]. classic [865]. classical [2437]. Classification [927, 704, 237, 1250, 1724, 2363, 354, 2756]. classifiers [2570]. classroom [749]. clause [2122]. clauses [319]. CLEARSY [2521]. CLG [2743]. CLG-Trans [2743]. clicker [1408]. Client [1124, 2306, 841]. Client-based [1124]. client-driven [841].

Cliffs [78, 67, 514]. clinical [1628]. clock [1374, 1561, 2542, 2556]. clock-based [2542, 2556]. clone [1731, 1733, 1132]. cloned [2698, 1734]. clones [1348, 1730]. Closed [25]. closing [458]. Closure [13, 2628, 1895]. closures [1205]. cloth [249]. Cloud [1590, 1994, 2147, 2262, 2135, 2456, 2151, 1650, 2826, 2344]. clouds [1991]. club [2447]. cluster [1897, 2268, 1305]. Clustering [654]. CM [1919]. CM2KG [2810]. CNOT [2643]. Co [1893, 2731, 2667, 2758]. Co-evolving [1893]. co-labeling [2758]. co-logic [2731]. co-verification [2667]. Coad [246]. coalgebraically [1502]. coalgebras [1461]. COBOL [657, 597]. COBOL/CICS [657]. Coccinelle [1660]. Code [1896, 1074, 262, 2784, 578, 1585, 1849, 598, 919, 2359, 1541, 1202, 164, 1348, 1248, 1434, 1255, 1296, 2598, 1486, 596, 1640, 831, 1314, 1834, 1476, 2576, 2675, 2861, 2624, 1955, 1743, 1463, 687, 1900, 1899, 2518, 1734, 2488, 1317, 2451, 1125, 1450, 1551, 666, 1132, 1451, 1247, 2004, 1239, 2783, 2035, 2710, 876, 1788, 2756, 2390, 1031, 1290, 2564, 1407, 2592, 2743, 739]. code-first [1640]. code-improving [739]. coded [1898]. CodeSparks [2784]. Coding [1660]. coercion [1207]. Coevolution [1871]. cognition [2365]. cognitive [1603]. coherence [2506]. cohesion [1124]. coinduction [2026]. cold [1290]. colibri [1050]. colibri2 [1534]. collaboration [1308, 2680]. collaborations [2285]. Collaborative [1961, 1539, 822, 1850, 1148, 1616, 1203, 2830, 992, 1178]. collect [497]. collected [1465, 291]. Collection [6, 1541, 917, 967, 916, 1405, 1059]. collections [2249, 1963]. collective [2462, 1146, 599, 2876]. Collider [1519]. Colloquium [2883, 20]. colony [1427]. colored [2379, 1324, 804]. colorful [2698]. colouring [1005]. combinability [1855]. combinational [954]. Combinations [569, 917]. combinator [737]. Combinators [1979, 1955, 1942, 1195, 146, 2250]. combined [2165]. Combining [1389, 908, 2191, 1128, 1168, 2613, 1334, 1725, 677, 1911, 1238]. ComMA [2850]. command [863, 451]. commands [1283, 843]. comment [962]. commercial [1880]. commitment [378]. Committee [402]. Common [2150, 835]. commonalities [2679]. commonality [2175]. Communicating [70, 33, 5, 254, 574, 90, 438, 322, 1108, 1359, 2569, 1345, 447, 126]. Communication [25, 84, 327, 2462, 465, 786, 599, 646, 2098, 2379, 1431, 62, 463, 819, 2305, 432, 515]. Communication-Closed [25]. communications [1798]. community [1617, 2644, 1048]. commutativity [2805]. Commuting [2190]. Compact [1222, 1292, 280, 2435, 1519]. companion [2442]. Comparative [236, 1681, 2335, 208]. comparator [1946]. Compared [396]. Comparing [663, 1937, 240, 2342, 1232]. Comparison [1646, 1132, 1521, 1442, 2600, 2189, 410, 548, 1441, 1592]. Compatibility [604, 2569, 1361, 1354, 1673]. competitive [2400, 1412]. compilation [1417, 1314, 2210, 2395, 2768, 1125, 176, 2643]. compile [2511, 846]. compile-time [2511, 846]. compiled [2314]. Compiler [524, 1162, 113, 356, 859, 1045, 678, 1032, 1532, 1968, 1978, 834]. compiler-compiler [1532]. compilers [1361, 869, 41]. Compiling [106, 469, 203]. Complete [2471, 1469, 1077, 1695, 1859, 1936, 2369, 1721, 1703]. Completeness [1276]. completion [110, 2764]. Complex [119, 1295, 1420, 1527, 2510, 2330, 1523, 2077, 2092, 1984, 2047, 2833]. Complexity [388, 23, 840, 1781, 1881, 1129, 2816, 1358, 2634, 2172, 1579, 2755, 515, 145]. compliant [1991, 623]. Component [1702, 928, 1647, 622, 1969, 1254, 1214, 2031, 1191, 1378, 1653, 2686, 621, 1649, 796, 662, 2530, 895, 1656, 1282, 1374, 1906, 2519, 1343, 1806, 1113, 1658, 1805, 511, 2456, 1298, 1807, 1322, 788, 1970, 802, 2321, 1198, 2073, 2395, 1350, 1145, 879, 2850, 1522, 1972, 1973, 2464, 803, 2033, 1785, 1446, 1492, 1812, 2363, 2617, 1072, 800, 1380]. Component-Based [1647, 621, 796, 662, 1374, 1906, 1113, 1805, 2456, 1298, 1807, 1322, 788, 1970, 2073, 1522, 1972, 2464, 2033, 1785, 1446, 1812, 1072]. component-interaction [1282]. component-oriented [1198]. components [796, 784, 1583, 2456, 322, 590, 2353, 1511, 606, 795, 1532, 1213, 405, 877, 1184, 801, 817, 2344, 1847, 1872, 806, 550]. Composing [883, 677, 934]. composite [852]. Composition [378, 788, 804, 1430, 784, 1288, 2415, 638, 797, 1214, 2146, 1796, 733, 1301, 803, 798, 779, 807, 794, 801, 1837, 1738, 1412, 805, 2395]. Compositional [1485, 2551, 1658, 2672, 1524, 1431, 607, 1315, 1978, 2493, 1377, 1971, 2350, 1804, 2579, 709, 545, 2326, 565, 1544, 1966]. Compositionality [1672, 780, 648, 1915]. compositionally [1843]. compositions [1704, 2032]. compound [2259]. comprehensibility [876]. comprehension [1490, 1540]. Comprehensive [682, 1289, 1640, 2060]. compressed [832]. compression [2721]. Comput [471, 19, 781, 648]. Computability [139, 2026, 172, 169]. computable [1225, 2658, 2479]. Computation [556, 2178, 2682, 31, 573, 632, 2181, 186, 393, 471, 1147, 2765, 2429, 909, 605, 1367, 127, 181, 152, 1458, 991]. computation-Interactive [181]. Computational [1833, 1995, 2634]. computations [2503, 2401, 1599, 1951, 256, 1252]. Compute [1528]. Computer [838, 48, 49, 997, 869, 514, 154, 1292, 172, 388, 168, 171, 216, 271, 226, 56, 232, 272, 727, 870, 611, 180, 884, 1086, 398, 1014, 404, 397, 215, 756, 2048, 418, 284, 142, 316, 34].

Computer-assisted [838]. computer-based [611]. Computing [2703, 2875, 2276, 473, 1902, 1487, 2763, 13, 35, 1905, 568, 1420, 907, 929, 957, 79, 295, 219, 2073, 2070, 2634, 1548, 1309, 2270, 2341, 2719, 1204, 1436, 2300, 1215, 1360, 1400, 1416, 1404, 166]. concept [2537, 2049, 2604]. concept-based [2049]. Concepts [642, 1849, 1197, 297, 1183, 1206, 794, 1060, 1466, 270, 1803]. conceptual [1712, 1761, 2810]. concern [1308, 1761, 1611]. concern-based [1308]. concern-specific [1611]. Concerns [1612, 1056, 162]. Concerto [2554]. Concolic [2154]. concrete [1417]. Concurrency [1574, 1863, 1914, 1598, 347, 14, 15, 848, 620, 786, 1935, 255, 932, 1275, 2077, 852, 2265]. Concurrent [51, 53, 606, 59, 254, 2206, 477, 393, 471, 1987, 614, 1327, 967, 854, 2708, 2077, 2576, 2878, 326, 2253, 1237, 2464, 1460, 2342, 819, 890, 1278, 1851, 1812, 1931, 1234, 913, 817, 283, 2406, 152, 2660, 185]. condition [2275, 1457, 790]. Conditional [1667, 2419, 1563, 2001, 1950]. Conditions [161, 2153, 1894, 1686, 2514]. Conference [9, 2730, 1373, 2875, 1341, 2625, 2562, 2178, 2682, 2800, 2775, 2665, 1809, 2272, 2686, 2693, 2851, 35, 56, 2885, 318]. conferences [1635]. confidentiality [1189]. Configurable [727]. Configuration [985, 2241, 692, 758, 1182, 987, 754, 755]. configurations [662]. Confining [929]. Conflict [1961, 2389, 1217, 1365, 1622]. conflicts [2415]. confluence [2429, 1367, 1894]. confluent [1931]. Conformance [1184, 2258, 2326, 2124, 2842]. congruence [241]. congruences [1434]. connected [606, 2453]. connection [1843]. connections [725]. Connector [1005, 2145]. Connectors [1351, 895, 1343, 1350, 1145, 1973, 2362, 1544, 888, 2317]. Conquer [118, 421, 2815, 2651, 63]. consensus [1992]. consideration [1796]. considerations [2525]. considered [428]. considering [2762]. Consistency [1671, 2631, 1273, 2440, 2877, 530, 518, 1806, 2669, 2647, 479, 2646, 2295]. consistently [1280]. ConSpec [1100]. Constants [129]. Constrained [1950, 2321, 2022, 2287, 2208, 1446]. Constraint [478, 1771, 2146, 2122, 532, 408, 425, 476, 895, 1932, 1889, 1272, 717, 2242, 1510, 2266, 1414, 454, 2683, 1931, 1353, 552, 2529, 475]. Constraint-based [478, 532, 1889, 552]. Constraints [503, 2583, 1039, 477, 2551, 1840, 2168, 543, 1908, 2186, 536, 1324, 475, 2501, 2313, 271]. construct [1109, 671, 602]. Constructing [1195, 71, 311]. Construction [46, 417, 1495, 359, 1593, 569, 1045, 431, 678, 76, 2885, 1462, 1978, 100, 2032, 695, 1776, 480, 195, 78]. constructions [410, 305]. Constructive [530, 1758]. Constructs [51, 599, 1028]. consumption [2249, 2263, 1944, 1945, 2132]. containment [2348]. content [1294]. contest [1592]. Context [2273, 1512, 2440, 1343, 2115, 1169, 1912, 2025, 1005, 1986, 458, 1301, 1515, 1634, 2187, 1879, 1535, 1120, 2045, 480, 550]. Context-aware [2273, 2115, 1986, 1515, 1879, 2045]. context-bounded [2187]. Context-dependent [1512, 1343]. context-free [1169, 1634, 550]. context-oriented [1301]. ContextErlang [1879]. ContextJS [1301]. Contexts [1277, 468, 930]. Contextual [2073, 474, 301]. contextualization [2794]. continuation [1688, 281]. continuations [885, 353]. continue [2607]. Continuous [2099, 252, 2774, 473]. contract [1993, 1865, 1864, 1705, 2859, 2704]. contract-obedient [1865]. Contracts [1805, 1390, 2296, 1917, 2423, 1656, 1671, 2726, 2824, 2846, 2618, 1900, 1833].

Contracts-refinement [1805]. contrastive [2870, 2743]. contribution [1665]. contributions [1659]. Control [1114, 30, 1706, 1319, 1574, 863, 1888, 1770, 2447, 2241, 2090, 2061, 2159, 2586, 461, 1662, 1700, 2330, 2791, 1519, 1470, 2053, 2667, 1018, 2131, 1821, 2270, 992, 619, 1458, 1756, 1528, 153]. controlled [656, 2534, 1743, 2385, 1181, 1336]. Controller [2423, 1770, 1695, 603]. controllers [2680, 2427]. Controlling [2735, 1959]. Conventions [33]. Convergence [561, 2064, 1571]. conversion [626]. conversions [2103]. convex [839]. convolution [2762, 277]. CoolTeD [2758]. cooperating [2127]. Cooperative [1123, 1959]. cooperatively [2526]. Coordinated [79]. Coordinating [1007, 667, 2147, 1986]. Coordination [2730, 2240, 2310, 490, 493, 661, 1385, 1919, 1341, 496, 2562, 2374, 2452, 1989, 2403, 930, 1143, 2563, 1349, 498, 958, 1180, 2553, 492, 663, 664, 1272, 959, 1187, 925, 926, 495, 1990, 2545, 1992, 1985, 2242, 1921, 898, 1988, 1182, 665, 1543, 871, 1183, 1184, 666, 1232, 1920, 1756, 1410, 1385, 2374]. coordinators [2552, 1072]. cope [1701]. Copenhagen [2880]. copy [497]. copy-collect [497]. Copying [92]. copyless [1586]. Coq [1138, 2534, 1973, 2845, 2430, 2317]. Corby [298]. Core [1901, 944, 2329, 2092, 2150, 2592, 2127, 2803, 2091]. CoreASM [1527]. cores [1745, 1357]. corpus [1816]. Correct [2415, 262, 1321, 2544, 2826, 1903, 1365, 690, 1774, 2032]. correcting [1296]. Correctness [98, 2721, 5, 330, 26, 1908, 520, 281, 713, 239, 52, 632, 2230, 42, 1672, 134, 1599, 2064, 526, 1999, 2420, 1776, 1105]. correctness-by-construction [1776]. Correctness-guaranteed [2721]. correspondence [1718]. COSMIC [2111, 2112]. Cost [2732, 2747, 342, 872, 2101, 1946, 1840, 1580, 1267, 2110]. Cost-effective [2747, 2101]. Costing [562]. costs [1845]. Cottam [288]. counter [2527, 1568]. counter-factual [2527]. Counterexample [2277, 2163, 1657]. counterexample-guided [2163]. counterexamples [790, 161]. countermeasures [1430]. counters [440]. coupling [2526, 2107, 857, 2292, 697]. course [142]. Courtiat [268]. Covalent [2180]. cover [2591, 154, 172, 1460, 168, 169]. coverability [1359]. Coverage [806, 1531, 857, 2384, 2400, 2863]. CPN [1203, 2385]. crash [2059, 2158]. crash-safe [2059]. crash-tolerant [2158]. Crawl [1794]. Crawl-based [1794]. Creating [1038]. creation [2488]. Creek [21]. CrimeSPOT [1491]. Crippled [57]. criteria [2189, 2027]. criterion [2483]. Critical [1428, 2088, 2197, 2354, 2603, 1567, 1236, 669, 2038, 1572, 1996, 1752, 2386, 1934, 2669, 786, 2329, 543, 1558, 671, 2251, 578, 1550, 2390]. criticality [2526]. criticality-based [2526]. critics [1394]. CRL [585]. CRM [872]. cross [1828, 1772, 333, 2746, 1125, 2733, 2278, 2859]. cross-compilation [1125]. Cross-contract [2859]. cross-cutting [1828]. cross-device [2733]. cross-layer [2278]. cross-platform [1772]. cross-project [2746]. cross-roads [333]. CrossFuzz [2859]. Crowbar [2748]. CRUD [2309]. cryptographic [1663, 709, 708, 685, 1105]. Cryptography [215]. CScout [1174]. CSLI [191, 573]. CSP [137, 423, 1283, 38, 2257, 2500, 1568, 76, 591, 791, 1371, 2369, 971]. CSP-based [1568]. CSP-OZ [791]. CSP-Z [591]. CTL [2411, 1859]. cubic [1164]. cubing [177]. CUDA [2187]. Cultural [1711]. cumulative [1633]. Current [2267, 2107]. custom [1338, 2268]. custom-tailored [2268]. customizable [761, 1738]. Customizing [978]. cut [2513, 236]. cut-off [2513]. cutting [1828, 1362]. cutting-edge [1362]. Cyan [2528]. Cyber [1520, 2258, 1888, 2157, 2286].

Cyber-Physical [1520, 2258, 1888]. cycle [1618, 1267]. cycles [2222]. cyclic [2329, 590, 1445, 177]. cyclicity [1728]. Cyclone [914].

D [183, 243, 228, 207, 143, 185, 226, 288, 142]. DAGs [2120]. Daikon [1047]. Dalvik [1668]. Daniel [79, 167]. Daniele [317]. Darrel [209]. Dart [2085]. Dashboard [2541]. Data [294, 140, 2062, 78, 139, 258, 150, 30, 239, 157, 2862, 1667, 978, 2174, 476, 1583, 1962, 2370, 1806, 589, 265, 1630, 2534, 1083, 929, 1935, 2408, 982, 926, 2274, 2299, 728, 548, 1911, 809, 521, 724, 626, 1985, 394, 1383, 1482, 687, 309, 2858, 158, 2252, 2435, 563, 2488, 2763, 713, 960, 1189, 578, 2400, 1547, 1811, 1542, 2140, 1665, 351, 224, 697, 201, 737, 1799, 2482, 1621, 2476, 2661, 194]. data-aware [1985]. data-flow [476, 2534]. data-intensive [1799]. Data-race [2482]. data-space [926]. database [2862, 1786, 2169, 722, 653, 1899, 654]. databases [1788, 48]. databases-with [1788]. dataflow [1490, 845, 2592]. dataflow-based [1490]. dataset [2512, 2814, 2812, 2758]. dataspaces [664]. datatypes [1245]. David [48, 297, 166, 170]. deactivation [2215]. dead [687]. Deadlock [1238, 482, 1825, 1522, 2382, 2500, 374]. deadlocks [2853]. Dealing [1812, 32]. deals [810]. debate [266]. Debray [318]. Debt [2267, 2633]. debugger [1924]. Debugging [1193, 2314, 2527, 2098, 595, 842, 1532, 1745, 2253, 2476, 2585]. Debugopt [2314]. decay [917]. December [56, 2522, 2621, 2717]. decentralised [1797]. decentralized [1821]. decidability [1855]. decidable [1558]. Deciding [2124]. Decision [1182, 2297, 1524, 2541, 487, 844, 1334, 1336]. Decision-making [1182, 2297]. declarations [1610]. Declarative [2149, 1584, 1925, 2204, 1924, 2128, 1986, 2381, 2380]. Decoding [2643]. decomposed [1697]. decomposing [322]. Decomposition [25, 2618, 2375]. decomposition-based [2375]. decompositions [344]. Decoupling [1056]. dedicated [1387]. Deductive [2748, 4, 1663]. deduplication [2827]. deep [803, 2691]. defect [2742]. Defending [2676]. defined [1363, 1086, 2165, 2479, 2427]. Defining [1517, 1218, 2292]. Definition [2125, 991, 401, 245, 2046]. Definitional [157]. definitions [241, 175]. deforestation [2072, 503]. degradation [2201]. degree [2175, 949, 2638]. degrees [1807]. del [391]. delayed [1247]. delaying [2019]. delays [1902]. Delegates [2240]. Delegation [1288, 350]. Delgado [171]. delimited [885]. delta [2284]. delta-oriented [2284]. demand [2446, 1848]. demand-driven [2446, 1848]. demonic [980, 981]. demonstration [2167]. Demonstrations [2866]. denial [1194, 2773]. Denmark [2880]. Denotational [212, 2836, 216, 175, 183]. density [1990]. deoptimization [1120]. Dependability [1662]. dependable [2525, 2580]. Dependence [1344, 1897, 1457, 1015]. dependencies [1637, 1445]. Dependency [2402, 2483, 484, 1005, 1224, 779]. Dependent [427, 2206, 1343, 2098, 1588, 1757, 1103, 1512, 972]. dependently [2602]. dephasing [2415]. deployed [777]. deployment [1972, 2521, 2826, 658, 1529]. dereference [1848]. Derivation [108, 2159, 326, 82, 30, 2356, 92, 197, 517, 186, 278, 256, 384, 1903, 315, 454, 148, 99]. derive [1652]. derived [331, 1998]. Deriving [568, 1497, 1596, 138, 1842, 424, 1721, 342, 121]. DescribeML [2814]. Describing [982]. Description [1600, 385, 74, 1601, 1602, 245, 636, 2814, 2817, 267, 268, 301]. Descriptions [1607, 1480, 887, 1219]. Design [598, 70, 419, 1614, 1641, 1095, 1199, 2278, 440, 762, 1994, 973, 796, 1905, 400, 1065, 423, 329, 1374, 746, 399, 770, 1836, 1806, 589, 1771, 265, 1420, 1296, 1113, 2578, 1615, 865, 571, 449, 347, 2546, 1606, 2233, 450, 244, 276, 1679, 321, 2882, 789, 2071, 405, 1556, 320, 1304, 327, 1864, 771, 1871, 613, 1652, 593, 2762, 1246, 2465, 63, 260, 277, 2048, 1315, 1732, 170, 1763, 1560, 397]. design-by-contract [1864]. Design-code [598]. Design-time [762]. Designed [1618]. Designing [1888, 1200, 1493, 765, 362, 1968, 994, 118, 644, 1287]. designs [1817, 1498, 2848, 1351]. DESOLA [1247]. destructive [191]. Destructively [150]. desync [2768]. desync-cc [2768]. desynchronisation [1655]. desynchronization [2768]. detailed [2830]. detect [1898]. Detecting [2175, 467, 175, 2274, 1017, 1732, 2261, 2375]. Detection [7, 888, 2389, 870, 2753, 599, 1047, 2482, 1711, 2675, 2624, 1961, 1132, 2819, 1165, 1267, 148, 2268, 2859, 1305]. detector [1731]. determination [606]. determinism [1356, 1277, 2188]. Deterministic [572, 1461, 1561, 2564]. develop [1528, 1320]. developer [2365, 1727]. developers [2533, 1665]. Developing [27, 1155, 1435, 153, 337, 1774, 177, 1422, 1092, 1968, 1125, 1309, 1491]. Development [2541, 30, 102, 1612, 1708, 1173, 1504, 2013, 2524, 2303, 9, 1316, 2234, 1770, 541, 113, 2815, 2247, 2223, 1090, 1395, 1632, 2062, 2793, 2540, 991, 507, 312, 1772, 1545, 2873, 2850, 2521, 2599, 2066, 2744, 1507, 279, 2655, 671, 2559, 1604, 194, 1148, 1638, 1010, 1446, 2754, 2849, 1876, 2217, 1508, 1878, 1530, 209, 1818, 2304, 1538, 1960, 144]. Developments [45, 1645]. device [2340, 2587, 2733]. DevReplay [2710]. DFAs [695]. diagnose [754]. Diagnosing [1296]. Diagnosis [2871, 1719]. diagram [691, 2786, 642, 2865]. diagram-based [2786]. diagrams [1038, 2669, 639, 1937, 1517, 2500, 1103, 2767, 971]. dialect [502]. dialogue [774]. dialysis [2353]. diamonds [1655]. DiaSuite [1528]. Diaz [267, 268]. DiCoMEF [1961]. didactic [2358]. difference [2806]. differences [1455]. different [2640, 1761, 2172, 434]. Differential [443, 2436]. differentiation [2806]. Diggers [21]. digital [750, 171, 2555]. Dijkstra [375, 204]. Diller [203]. Dilworth [124]. dimensional [2173]. Dinamica [2339]. direct [365, 1938]. Directed [82, 1890, 1221, 2703, 356, 336, 2492, 890, 2400]. directions [2598]. directory [586]. Dirichlet [277]. Dirk [291]. disambiguation [2191]. disassembler [1061]. disassembly [2768]. disc [280]. discipline [263, 728]. Disciplined [1798]. DisCoTec [2875]. Discourses [290]. discovery [1069, 2011, 1155, 1219, 1326, 2427]. Discrete [228, 2163, 1574, 1375, 492, 466, 786, 1369, 1935, 1657, 2295, 2396]. discrete-event [1574, 1375, 2396]. discrete-time [466, 786, 1657]. disease [2823]. Disentangling [1547]. disjunction [1589]. disjunctive [506, 389]. disk [965]. DiSL [1819]. Disnix [1529]. dispatch [1207]. dispatcher [1552]. dispatching [1248]. dispersion [1734]. Display [10]. Display-Oriented [10]. Dissolving [2176]. distance [2222]. distances [2054]. Distributed [491, 2240, 478, 692, 1178, 2875, 25, 495, 273, 96, 109, 26, 75, 82, 586, 1414, 1706, 1654, 1888, 1905, 1829, 870, 1420, 1146, 1856, 1527, 79, 464, 868, 1092, 1827, 289, 1427, 207, 2040, 1992, 2242, 1520, 2017, 716, 2464, 1201, 1367, 1579, 479, 1548, 1774, 1494, 1879, 817, 1229, 148, 737, 152, 1458, 99, 1529, 1411]. distributed-memory [1548]. Distributing [360, 2173]. distribution [1826, 2109]. distributions [2600]. Distributive [97]. divergence [1757, 2493]. divergence-sensitive [1757]. diverse [2656]. diversity [1639]. Divide [118, 421, 2815, 2651, 63]. Divide-And-Conquer [118, 421, 2815, 2651]. Diving [2152]. division [158]. DL [1203]. DM [184, 154, 299, 319, 172, 171, 216, 179, 271, 290]. do [2403, 1684, 549]. Docker [2291, 2593]. documental [1010]. documentation [1743, 2012]. Documenting [881]. documents [799, 1310, 320, 404]. Domain [102, 1764, 1849, 1419, 2013, 2567, 2152, 952, 1339, 2828, 1244, 1710, 2769, 2269, 2590, 305, 908, 2848, 1679, 1208, 671, 1515, 2559, 2165, 1608, 2501, 1612]. domain-independent [1710, 2269]. Domain-specific [1764, 1419, 2590, 908, 1679, 671, 1608, 2501]. Domains [1823, 533, 2829, 569, 2632, 618, 506, 1686]. domestic [2791]. Dorset [207]. Down [70, 1295, 444, 279, 440]. downwards [557]. Dowsing [142]. Draft [402]. DREMS [1905]. Driven [1624, 2866, 1960, 2524, 1088, 2303, 1202, 2115, 1979, 2328, 2247, 2829, 1630, 1912, 2025, 2248, 1090, 1489, 1113, 2656, 1505, 1631, 2828, 1628, 1626, 1322, 998, 1092, 1761, 841, 1772, 2298, 1629, 2446, 2844, 1091, 1745, 2599, 1310, 1507, 1848, 1638, 2847, 2754, 1203, 1151, 1664, 2048, 990, 1799, 1625, 1422, 1904]. driving [2747]. Dromy [34]. drug [2823]. DSL [1981, 729]. DSLs [1287]. Dual [124, 1401, 1675]. duration [1558]. during [1741, 2768]. Dutch [2055]. Dynamic [1273, 1706, 491, 2090, 692, 2246, 1429, 366, 2488, 1206, 853, 1727, 1060, 880, 1994, 1574, 1119, 1437, 885, 1420, 957, 859, 1047, 1067, 618, 728, 2584, 312, 1983, 2260, 720, 1233, 2320, 2017, 1552, 1819, 2388, 2252, 857, 2352, 1608, 1705, 1168, 1407, 712, 2542, 2556, 1305, 1516, 1830]. dynamically [1831, 1533]. dynamically-typed [1831].

e-barter [826]. e-Learning [1612, 1618]. E/E [1652]. Eames [316]. Earley [198, 45]. Early [2010, 1789, 2692, 2248, 1224, 639, 1640, 2332, 613, 1652]. Easy [1069, 1957]. EASYMAP [1083]. EATCS [154, 172, 271]. EBNF [2288]. Eclipse [1814, 2809, 2377]. Ecore [1961]. ecosystem [2645]. Ecosystems [2681, 2046, 1177, 1920, 2635]. Ed [700, 731, 299, 151, 219, 193, 388, 145, 167, 266, 230]. EDA [1799]. edge [1362, 2189, 1644]. edge-pair [2189]. edges [578]. editing [1691, 1961]. edition [516, 316]. Editor [156, 10, 976, 1602, 642, 925, 1043, 1076]. Editorial [528, 531, 555, 651, 660, 668, 675, 680, 689, 726, 743, 752, 759, 768, 776, 782, 793, 808, 814, 823, 828, 836, 847, 855, 866, 874, 882, 893, 904, 911, 918, 924, 931, 938, 945, 955, 968, 975, 984, 988, 996, 1002, 1008, 1016, 1024, 1029, 1034, 1042, 1052, 1053, 1057, 1066, 1070, 1075, 1082, 1087, 1093, 1098, 1106, 1111, 1117, 1126, 1133, 1142, 1149, 1154, 1159, 1171, 1179, 1185, 1235, 1311, 2245, 2255, 2264, 417, 1263, 1454, 962, 997, 2515, 769, 2, 415, 1647, 2383, 1624, 694, 706]. Editorial [961, 1808, 1813, 1820, 1838, 1839, 1853, 1862, 1868, 1885, 1891, 2021, 2030, 2037, 2043, 2051, 2067, 2081, 2082, 2087, 2095, 2096, 2105, 2113, 2117, 2123, 2126, 2129, 2133, 2137, 2139, 2142, 2148, 2155, 2166, 2171, 2177, 2182, 2192, 2193, 2196, 2203, 2218, 2225, 2237, 2238, 2281, 2282, 2293, 2301, 2312, 2315, 2319, 2324, 2337, 2346, 2351, 2357, 2360, 2364, 2368, 2372, 2378, 2387, 2394, 2399, 2410, 2417, 2422, 2434, 2439, 2444, 2449, 2461, 2467, 2473, 2478, 2487, 2496, 2509, 2517, 2523, 2532, 2539, 2549, 2558, 2566, 2575]. Editorial [2589, 2595, 2606, 2611, 2616, 2622, 2630, 2637, 2650, 2663, 2671, 2678, 2690, 2696, 2701, 2707, 2718, 2725, 2741, 2752, 2761, 2771, 2778, 2802, 2822, 2841, 2856, 2869, 2623]. Editors [143, 2097, 335, 1127, 1186, 1525, 1590, 1173, 1814]. Eds [730, 573, 291, 243, 391, 267, 268, 317, 318, 284]. Edsger [204].

Edu [2786]. education [2599, 750, 744, 2048]. Educational [746, 2049, 1618]. Edward [246]. Effect [2206, 2003, 1880, 2873, 697, 1336]. Effect-dependent [2206]. Effectful [2674]. effective [2747, 2746, 843, 1219, 2101]. Effectively [2265]. effectiveness [579]. effectivity [2407]. Effects [105, 851, 1084, 1724, 1021, 2860]. efficiency [616, 1388, 145, 2568]. Efficient [1248, 2527, 1889, 1147, 372, 1292, 2250, 1522, 672, 1152, 30, 1931, 1549, 913, 1386, 157, 2536, 2241, 113, 399, 1806, 906, 1228, 2554, 1368, 1241, 235, 446, 1222, 2194, 2017, 605, 1208, 1182, 1965, 1210, 1462, 2489]. Efficiently [2005]. effort [2571, 2640, 2561, 2692, 2480, 1181]. EGO [2339]. Ehrig [271]. eigensolvers [1646]. Eijk [267]. EIS [969]. elasticity [2151]. Elder [78, 185]. eLearning [2046, 2044]. election [464, 625]. Electrical [156]. electricity [448]. electronic [320]. element [2232]. Elementary [224, 1243, 1595]. Elements [24]. Eleventh [1341]. elicitation [2560]. Eliminating [687, 1255]. Elimination [257, 11, 682, 1890, 690, 2150, 1096, 508, 1120, 906]. Elsevier [388, 1088]. email [838]. Embedded [1312, 1904, 448, 1907, 2212, 1805, 1131, 859, 1417, 1092, 590, 1287, 1091, 1745, 1049, 665, 1446, 658, 1314]. Embedding [340, 2069, 1846, 2490, 2762, 1481, 2586]. embeddings [663, 1192, 2703]. Emerging [1714, 2578]. EMF [1961, 1963, 1818]. EMF-IncQuery [1818]. EMF/Ecore [1961]. EMIP [2512]. emperor [269]. Empirical [1897, 2571, 2780, 1348, 2843, 1505, 2600, 2560, 1734, 2335, 2570, 1875, 2451, 1384, 1450, 2819, 1665, 2644, 2373, 2757, 597]. Empirically [2172]. employees [2780]. Employing [1652]. Empowering [2305]. Emulating [1466]. emulators [829]. enabled [2015, 2849]. Enabledness [1141].

Enabling [2231, 2100, 2587]. Enactment [1537, 930]. encapsulated [1439]. Encoding [2227, 1918]. encryption [2657]. End [2524, 2424, 2131, 512, 1902, 1544, 2101, 2722]. end-points [512]. End-to-end [2424, 1902, 1544]. End-user [2524, 2131, 2101]. energy [2455, 1388, 2263, 2568]. EnergyPLAN [2455]. enforced [2511]. Enforcement [2240, 2019]. Enforcing [1175, 1194]. engine [2550, 1006]. Engineering [2883, 2734, 1518, 588, 1023, 2256, 2052, 2212, 1647, 1373, 2239, 2866, 1254, 2276, 1948, 2838, 2580, 1191, 1487, 2515, 2687, 835, 2272, 2007, 1590, 1006, 2426, 2727, 2851, 2839, 1355, 547, 212, 2328, 2785, 2829, 1912, 2025, 2248, 1489, 816, 2283, 820, 1631, 1759, 2828, 1628, 1626, 2185, 2652, 1870, 2774, 1773, 206, 2100, 1629, 1887, 2365, 1869, 1033, 1176, 1044, 217, 611, 2847, 1873, 1875, 1508, 2830, 1852, 1555, 654, 2617, 1762, 1872, 1874, 1512, 1635, 1877, 1346, 1382, 1780, 2694, 1625, 2391, 1624, 998, 2582, 2737, 1423, 1488]. engineering-principled [816]. Engineers [156, 156]. Englewood [514, 78, 67]. enhanced [1388, 2077]. enhancement [2380]. Enhancing [2827, 2655, 1601]. Enric [573]. Enriching [1506, 1691]. enrichments [1861]. ensemble [2624]. ensembles [2474]. Ensuring [1664, 1843, 673]. enterprise [989, 995, 2652, 887, 992]. enterprises [1631]. entertainment [1013]. entities [772]. entity [2054]. Environment [692, 643, 2034, 1889, 1022, 1046, 745, 1176, 1627, 2358, 1402, 1818, 2813, 903]. Environments [11, 1714, 1420, 1490, 1615, 1751, 1395, 2070, 749, 1650, 2008, 1835, 2050, 175, 1790]. equalities [2729]. equality [1861, 1334]. equation [2435]. Equational [199, 1950]. equationally [523]. Equations [303, 294, 121, 328, 534, 1001, 813]. equilibrium [2766]. equivalence [1461, 235, 2361, 1931]. equivalent [1724]. ERCIM [669]. Erlang [1924, 2305]. Ernest [390]. Errata [19]. Erratic [665]. Erratum [566, 471, 376, 1292, 781]. Error [135, 841, 2074, 90, 2280, 2527, 715, 1639, 790, 1719, 2431]. error-tolerant [2280]. errors [1482, 1812, 754]. Ershov [290]. ESBMC [2187]. ESBMC-GPU [2187]. ESEiC [1590]. ESOP [2881, 2880]. Essays [219]. essence [778, 1873]. essential [1455, 2186, 156]. EST [1525, 1043, 1076, 1173, 1814, 2097]. EST-4 [1525]. EST-6 [2097]. established [998]. Establishing [846]. Estelle [268]. ESTEREL [329, 541]. estimate [1476, 2597]. estimating [489]. estimation [2571, 2640, 2692, 1156, 775]. eternity [722, 2664]. eTextbook [1614]. eTutoring [2808]. Euclid [1243]. EUD [2524]. EUD-MARS [2524]. EUROMICRO [1804]. Europe [318, 191]. European [2881, 2884, 2880, 2882, 2693, 2200]. evaluate [2780, 1516, 2116, 2102]. evaluated [2405]. Evaluating [2631, 1490, 579, 902, 1616, 818, 1505, 2168, 2816, 2172]. Evaluation [1716, 2118, 1954, 129, 1591, 1259, 2304, 2757, 1642, 313, 2121, 1260, 2823, 2786, 1603, 387, 908, 2070, 1347, 2858, 1027, 822, 237, 1433, 520, 2335, 2451, 1252, 1199, 1132, 1652, 1247, 960, 2819, 2764, 445, 170]. evaluator [65]. even [2874]. Event [2127, 2803, 1212, 1574, 1375, 1282, 1374, 1228, 2510, 2064, 2396, 2536, 2613, 2505, 1939, 1901, 2127, 2213, 2447, 2519, 2159, 1395, 2231, 2666, 1155, 1571, 1435, 2655, 1698, 2728]. Event-B [2127, 2803, 2064, 2505, 1939, 1901, 2127, 2213, 2447, 2519, 2159, 1395, 2231, 2666, 1155, 1571, 1435, 2655, 1698, 2728]. event-based [2613, 1240]. event-clock [1374]. EventML [2158]. events [293, 2019, 2827, 1966]. everywhere [732]. Eves [1116, 1114]. evidence [2244, 1785]. evidence-based [1785]. EVL [1844]. Evolution [1178, 2625, 2775, 92, 1294, 875, 1786, 1339, 1840, 1298, 854, 1651, 2316, 1443, 1300, 1303, 2033, 1785, 1444, 1446, 2514, 1323, 1440, 2101, 2638, 1566, 1788, 986, 1407, 1293, 736]. Evolutionary [1606, 2107, 1539, 2781, 2011, 2744, 2381, 2719, 2759]. evolvability [1302]. evolvable [748, 2396]. Evolving [385, 727, 1681, 1750, 2792, 1379, 1893, 2273, 994, 1580, 993, 2468]. EvoMBT [2759]. Exact [1369, 2407, 2036, 1252]. Examining [2742]. example [225, 1339, 1069, 1091]. Examples [6, 326, 161]. Exception [7, 1586, 1636, 382, 1250, 1183, 1122, 1610]. exception-safety [1250]. Exceptions [38, 851, 428, 338, 935, 2042, 238, 489]. exchange [1598, 878]. Exclusion [75, 852, 1264, 1881, 1433, 470]. Executable [1527, 830, 1565]. Executing [933, 1038, 2053]. Execution [2392, 1560, 1295, 691, 1425, 2584, 2070, 300, 1406, 1899, 1850, 2132, 1447, 1353, 1947, 2228, 2585, 1884]. executives [2329]. Exercise [30, 86, 13, 322, 282, 314]. Exercises [40, 43, 47, 54, 66, 72, 77, 91, 95, 101, 103, 107, 112, 117, 125, 131, 160, 2786, 1619, 275, 61, 147, 178, 190, 242, 2047]. Exhaustive [2804]. expandable [2249]. expanders [1411]. Expansion [121, 1721]. Experience [2152, 1863, 1914, 1201, 1776, 1295, 448, 2745, 1975, 1625, 785, 1536, 2828, 1628, 544, 2861, 744, 2849, 2830, 1404]. Experiences [2619, 910, 1555, 1382]. Experiment [462, 188, 2560, 1116, 1114, 1181, 1336, 1519]. experimental [2189, 960, 1165, 1816, 736, 1525, 1590, 1043, 1076, 1173, 1814, 2097]. experimentation [2774]. experimenting [1150]. experiments [1616, 552, 729]. explainable [2790]. Explaining [435]. explanation [132]. explicit [2013, 1962, 580]. Exploiting [2120, 1714, 773, 1368, 2874]. Exploration [2210, 2656, 1083, 1429, 1176, 1562, 891, 2373]. Exploratory [2877, 2533, 653]. Exploring [2320, 2833, 1875]. Expressing [1382, 1019, 582]. Expression [2074, 1747, 1748, 2872, 1634, 1678, 100, 696, 2431]. Expressions [11, 83, 2469, 1147, 584, 1205, 1218, 2773, 1811, 2710, 2774]. Expressive [352, 2211, 2578, 304, 2080, 1883]. Expressiveness [2020, 1915, 664, 898, 2313]. extend [2455]. Extended [33, 1498, 1960, 1716, 2149, 2118, 1702, 2068, 2024, 2143, 2685, 1954, 2283, 2239, 2391, 1687, 1854, 2506, 1726, 1500, 2582, 2737, 2688, 1746, 1510, 2367, 1997, 2018, 1692]. Extending [796, 638, 1526, 1772, 1103, 1618, 1251]. Extensible [1927, 1483, 1026, 1045, 2797, 1709, 1530, 1163, 2470]. extension [1485, 60, 2060, 2286, 2343, 2504]. extensions [1932, 2437, 2069, 1249, 1028]. extents [885]. External [240, 1429]. extract [2485]. Extracting [507, 2428, 658]. Extraction [1849, 2283, 2862, 1327, 551, 1333, 2488, 2458, 2016, 800, 2564]. Extremo [2377]. extrinsic [2506]. Eyal [515]. eye [2512, 1542]. eye-tracking [1542].

F [79, 2259, 194, 398, 249, 106]. F-Alloy [2259]. F.R.G [216]. Face [2862, 1644]. face-to-face [1644]. facet [1689]. Facilitating [2045]. FACS [2530, 2143, 1969, 2686, 1702, 2688, 2367]. FACS’09 [1378]. FACS’12 [1653]. factor [693]. factories [550]. factors [2843, 1629]. factory [656]. facts [2104]. factual [2527]. failboxes [2402]. failed [1394]. Failure [1072, 2055, 1629, 2790]. Failure-free [1072]. failures [1358]. fair [1276]. fairness [32, 1860, 1469, 2728]. fairness-based [2728]. false [266]. FAMe [2091]. FAMe-TM [2091]. FAMILIAR [1419]. families [2632, 1785, 2145, 1580, 1331]. Family [122, 2537, 1616, 2027, 2336, 2722]. family-based [2722]. farewell [1683]. Fariñas [391]. FASE [2838]. Fashioning [209]. Fast [2648, 2080, 481, 1921, 832, 403]. faster [534]. Fault [109, 44, 35, 2201, 812, 1672, 1826, 1880, 1639, 2746, 2491, 666, 756, 1998]. fault-proneness [1880]. Fault-Tolerant [44, 35, 2201, 1998]. FBD [2271]. FBDTester [2271]. FCPP [2820]. FDAF [995]. FDIR [2491]. feasible [2763]. Featherweight [1119]. Feature [602, 1419, 2631, 1220, 1319, 1316, 2201, 764, 2248, 1297, 1681, 1750, 2624, 1822, 1510, 467, 2320, 2618, 1182, 1317, 2336, 2638, 1530, 1318]. feature-based [1220, 2248, 1182]. feature-oriented [1319, 1316, 2248, 1681, 1530, 1318]. Featured [2716, 2323]. FeatureIDE [1530]. features [2571, 934, 2540, 2861]. featuring [2089, 1346]. February [21, 2416, 2421, 2538, 2636, 2882]. Fed [9, 35]. Federated [2875]. Feedback [1619, 1900, 2754, 2047]. feedback-driven [2754]. feedbacks [880]. Feijen [204]. Feijs [397]. Felleisen [167]. FermaT [738]. FIACRE [1906]. fidelity [1031]. Field [1728, 1921, 1851]. Field-sensitive [1728]. fields [1942, 1995]. fifth [232, 1814, 669]. File [1114, 1116]. Filman [79]. filtering [1850, 2432]. finalization [2489]. find [2480, 966]. finders [2656]. Finding [1991, 82, 24, 2435, 1216, 1401]. fine [2254]. Finite [580, 2369, 2503, 2232, 1942, 1561, 255, 2804, 536, 2179]. finite-state [255]. First [1597, 2332, 141, 1831, 2119, 1640, 2210, 2371, 2227, 1802, 1590, 582, 142, 1684]. first-class [2371]. first-order [2119, 2210, 2227, 582]. First-past-the-post [1597]. Fitting [169]. Five [912]. Fixed [23, 186, 369, 376, 2763]. Fixed-Space [23]. fixing [2853]. fixpoint [2818, 608]. Flash [1241, 1115]. flattening [1915]. flaws [2762]. flexibility [1388]. Flexible [1745, 1613, 1078, 1437, 928, 1342, 1080, 1287, 1470]. flight [1514]. floating [512, 1825, 1753]. floating-point [512, 1753]. Flow [140, 947, 476, 2307, 2424, 1962, 2534, 982, 2446, 559, 1679, 2053, 281, 1018, 2763, 2309, 578, 1483, 351, 1187]. flow-network [1679]. flow-sensitive [2446]. fly [1376, 1921, 672, 1367, 2032]. FMICS [1428, 1236]. FMICS’11 [1550]. FMICS’2014 [1996]. FMTA [2886]. Foclasa [2452]. FOCLASA’06 [1143]. FOCLASA’09 [1349]. FOCLASA’10 [1543]. focused [1835]. fold [2818]. folds [1500, 723]. FoLLI [573]. FOPARA [1941]. forecast [1157, 1325]. forecasting [2858]. Foreword [856, 1403, 349, 719, 472, 483, 540, 640, 377, 946, 456, 207, 302, 1982, 357, 1620, 500, 406, 251, 1058, 1186, 1118]. form [798, 1731]. form-based [798]. Formal [1994, 1430, 2883, 2734, 1654, 2530, 2232, 2201, 612, 197, 785, 2091, 2799, 2697, 1373, 2457, 2578, 1564, 2383, 645, 2535, 1858, 966, 669, 1909, 228, 2353, 1887, 2260, 1545, 2161, 2290, 1379, 2160, 2031, 1552, 1279, 1237, 2066, 1997, 385, 1329, 2886, 388, 603, 826, 1018, 1967, 617, 2686, 1965, 1670, 1938, 10, 2474, 2089, 2144, 2202, 1756, 1366, 2013, 973, 1100, 1570, 1817, 2234, 448, 2418, 2213, 1601, 383, 541, 459, 2370, 113, 1906, 589, 1115, 2236, 2285, 1731, 2334, 2463, 1856, 1527, 2823, 1337, 134, 2525, 1911, 16, 1393, 1553, 2063, 2848, 584, 1983]. formal [182, 1992, 1320, 789, 2362, 2826, 1893, 708, 1579, 267, 268, 2033, 890, 2661, 930, 1550, 2094, 1017, 1555, 1705, 1512, 1713, 142, 1998, 1428, 2088, 2197, 2354, 2603, 1702, 1236, 1919, 1969, 1996, 2572, 1378, 1653, 299, 397]. Formalisation [1266, 1668, 921]. Formalisations [1270]. Formalised [2490]. Formalising [1519]. Formalism [1, 119, 15, 1083, 37, 2396, 1347]. formalisms [37, 1310]. formality [613]. Formalization [1183, 2519, 556, 1748, 2479]. Formalizing [2483, 312, 1698, 2430, 2441]. Formally [460, 2669, 2666, 2322]. format [878]. formats [1356]. formatting [529]. forms [921]. formulas [1860]. Forward [2806, 947, 547, 434]. Forward- [2806].

FOSD [1316]. FOSS [1298, 1665]. Foster [270]. Foundation [286, 287, 1220, 784, 285, 1742]. Foundational [408, 1761, 1941]. Foundations [2241, 2310, 2452, 1989, 2572, 1143, 575, 1088, 56, 272, 1912, 2507, 939, 151, 2475, 2886, 1543, 1349, 266, 231, 389, 214]. Four [156, 1881, 1940]. four-variable [1940]. FP [328, 83]. frame [1793]. Framed [1055]. framework [943, 635, 2711, 2803, 2591, 2340, 2519, 1971, 2389, 2829, 1797, 1890, 2578, 1615, 865, 2169, 2456, 2811, 1632, 1354, 2136, 2128, 1844, 709, 1970, 2612, 2490, 2298, 1690, 2798, 1198, 2577, 2797, 321, 2017, 1095, 1993, 849, 337, 2362, 2500, 1850, 1061, 1303, 1472, 1446, 835, 871, 1875, 2587, 1534, 2617, 2772, 2454, 2292, 1530, 2468, 1041, 2757, 1538, 2635, 2860, 621, 496, 1537, 2784]. frameworks [1521, 2223, 1878]. France [2882]. Fred [154]. Free [58, 332, 725, 441, 1169, 1898, 967, 1634, 672, 1929, 403, 2228, 1072, 1238, 550]. freedom [1110, 1807, 1825]. Freedoms [22]. Freeman [245, 396]. French [60]. Frenetic [2797]. Frenetic-lib [2797]. frequency [1971, 2320]. fresh [2713]. Friedman [79, 167]. friend [1172]. front [2722]. front-end [2722]. FSEN [1518, 1948, 1355, 2239, 2391, 2582, 2737]. FSM [2361]. FTS4VMC [2722]. FTSCS [2088, 2197, 2354, 2603, 2608]. FTSs [2722]. Fudgets [665]. fuel [2213]. fulfillment [1633]. Full [2306, 1915, 401, 761]. Full-abstraction [2306]. Fully [1936, 1369, 1842, 1673, 2314]. fun [261]. Function [2627, 1333, 2280, 189, 562, 1967, 619]. Functional [419, 240, 93, 2111, 159, 106, 105, 404, 94, 121, 301, 632, 2679, 2303, 2119, 202, 2212, 533, 1227, 1221, 2781, 867, 735, 444, 1718, 2154, 410, 203, 521, 1352, 1822, 333, 296, 564, 1933, 2597, 1753, 2069, 398, 1928, 563, 1927, 504, 2150, 176, 381, 238, 2112, 2738, 529, 386].

Functions [104, 1946, 484, 1642, 313, 477, 1486, 1262, 1645, 2376, 1701, 309, 1225, 1158, 1944, 627, 1456, 723, 2805, 2789, 2345, 234]. functors [1255, 1894]. Fundamental [588, 2838]. Fundamentals [1355, 2803, 2239, 1518, 2391, 1948, 2582, 2737, 271]. fusion [2135, 1449, 1245, 2072]. Future [111, 1649, 2418, 2598, 891]. futures [1344, 1122]. fuzzing [2859]. fuzzy [2668, 2787]. fuzzy-logic [2787].

G [156, 34, 208, 179, 397, 390, 173, 226, 247, 248, 250]. Gaiwan [2782]. Galletly [516]. Gallier [180]. Galois [725]. Galton [193]. game [1011, 1010, 1012]. games [1496, 1597, 2716, 1015, 2581, 1014, 1013, 2660]. Gamification [2813]. Gamification-based [2813]. gamified [2808]. Gamifying [2785]. GAMMA [263, 494]. GaMoVR [2813]. Gannon [226]. gap [1601, 989, 995, 556, 1310, 1508, 529]. Garbage [1059, 1465, 917, 967, 916, 1405]. Gate [379, 2322, 2628]. gate-level [2322]. Gaussian [906]. Gazdar [247, 248, 250]. GCL [2257]. Gehani [143]. gender [2560]. genealogy [1734]. general [1110, 534, 1970, 65, 1869, 110, 1832, 1742, 1872]. general-purpose [1832, 1742]. Generalising [2120, 558]. Generalized [1020, 363, 2053]. generalizing [1877]. generate [2492, 2769]. generated [2247, 41, 1450]. generates [2710]. Generating [1890, 189, 2817, 790, 351, 1217, 1532, 2186]. Generation [1682, 1312, 832, 550, 1694, 1574, 2591, 1202, 164, 323, 542, 2078, 1161, 2000, 1723, 1889, 356, 1421, 2153, 2824, 2713, 462, 2786, 2584, 545, 1256, 1657, 1815, 1955, 2797, 803, 2837, 1351, 2805, 779, 2384, 949, 2309, 1247, 1313, 2400, 1484, 1811, 2385, 2271, 891, 2047, 2529, 625, 990, 2863, 2592]. generational [917]. Generative [1781, 1011, 571, 2048, 2470, 1254, 1191]. generative/adaptive [1011]. generator [2829, 907, 1201, 642]. Generators [262, 2750, 1257]. Generic [614, 557, 1139, 1137, 1248, 1255, 907, 569, 1354, 1253, 1249, 1252, 1451, 1258, 1803, 858, 978]. genericity [906]. genome [2220]. GenVoca [1201]. geography [750]. geometric [1252, 243]. Geometry [749]. Geraint [155, 192]. Gerald [207]. Germany [299, 2665, 9, 35, 2885]. gestUI [2335]. gesture [1824, 2335]. gesture-based [2335]. getting [983]. Ghost [1832]. GI [35]. GI/NTG/GMR [35]. Gibbins [213]. Ginzburg [573]. Git [2638]. Gitana [2195]. GitHub [2573, 2807]. GLL [1484, 2023, 2288]. Global [294, 957, 1992, 2854, 2731, 2084, 929, 751, 951, 454]. glossaries [2877]. GLR [2641, 902]. glue [1915]. Glyph [2784]. Glyph-based [2784]. GMR [35]. Go [2482, 2403, 2482]. Goal [336, 2787, 2160, 2405]. Goal-directed [336]. goal-oriented [2160]. Goble [229]. GOFFIN [477]. Goldsmith [192]. Google [1054]. Gordon [227]. Goto [1096, 873]. governance [1021]. GPCE [1254, 1191]. GPGPU [1729]. GPU [2162, 2187, 2782]. GPUs [2739]. gradient [2289]. gradient-based [2289]. grading [2047]. Gradual [1739, 2502]. graduated [2647]. grained [2254]. grammar [1051, 1609, 1977, 2488, 1760, 754, 1608, 2617, 1163, 1749, 100, 1816]. grammar-aware [2617]. grammars [1162, 1169, 1161, 1217, 2343, 1792, 65, 1765, 1027, 2069, 1634, 1678, 1321, 2288, 1481, 1764, 1777, 1332, 2431, 550, 2074, 1747, 1748, 1763]. grammarware [1816]. grammatical [1762]. grand [1112]. granularity [1533, 340]. Graph [519, 1892, 1202, 2054, 323, 2648, 2668, 692, 645, 2870, 200, 1792, 878, 2702, 2647, 1910, 1893, 364, 890, 2646, 2186, 1321, 2514, 424, 1963, 2270, 1732, 148, 1788, 641, 2743, 1703, 1332, 2052, 319]. graph-based [1202, 878, 1963, 319]. graph-oriented [692]. Graphic [320]. graphical [1601, 2328, 1602, 1616, 2684]. Graphics [1406]. GRAPHITE [2052]. graphs [429, 1715, 1457, 2702, 2053, 487, 1261, 2810, 1097]. Graubard [266]. great [1172, 1089]. greedy [704, 2172]. greedy-based [2172]. Green [2779]. Greg [296]. Gregory [396]. Grid [465]. grids [1714, 2479]. Gries [312]. Gröbner [1683]. grounded [2185]. Group [669, 299]. grouping [2757]. groups [1983]. groupware [1308, 991, 1638]. Grover [2683]. GSOS [2325]. GSPNs [2396]. guarantee [1658, 2468]. guaranteed [2721]. guarantees [1557, 1824]. guarded [2366, 451]. guards [1086]. Guest [1647, 2383, 1624, 925, 335, 1127, 1186, 1525, 1590, 1043, 1076, 1173, 1814, 2097]. GUI [2827, 2536]. guidance [1394]. Guide [169, 1507, 1785, 286, 285, 287]. Guided [2412, 2163, 2277, 1492, 2355]. guidelines [2389]. Guilin [2687]. Gul [152]. Guttag [144]. GXL [878].

H [244, 321, 180, 204, 267, 398, 271, 397, 290, 196]. Hadron [1519]. haemodialysis [2233]. half [2176]. Hall [195, 141, 126, 202, 183, 265, 155, 192, 514, 34, 285, 286, 287, 219, 244, 78, 246, 229, 217, 297, 67, 194, 315, 215, 227, 185, 270, 201]. Hall/Open [284]. Hamlet [226]. Handbook [88, 388]. Handel [1330]. Handel-C [1330]. handled [2084]. handlers [2860]. Handles [1831]. Handling [2490, 157, 382, 1586, 1183, 766, 1122]. Hapao [1531]. happened [1858]. Harcourt [193]. hard [154, 172, 169, 2279]. hardback [291, 245, 510, 295, 228, 219, 246, 298, 267, 268, 269, 271, 397, 209, 290, 196, 191, 247, 248, 250, 524, 521]. hardbound [146]. hardcover [193, 227]. hardware [796, 1077, 2672, 1662, 1108, 912, 1552, 565, 2667, 1313]. hardware-in-the-loop [1313]. hardware/software [2667]. Harel [166]. Harlan [207, 226]. harm [1684]. harmless [428]. Harper [180]. Harvard [316]. hash [2342]. Hashing [2140]. Haskell [978, 2003, 2533, 2342, 501]. HAVi [625]. HBSniff [2675]. HDFS [2408]. healing [817]. Health [2681]. healthy [1635]. Heap [1465, 1667, 1866, 1669, 1576, 713, 846]. heap-allocated [1667]. heap-manipulating [1576]. heaps [1953]. heat [2791]. heat-pump [2791]. heel [1059]. Heidelberg [184, 299, 179, 271, 290]. helpdesk [2780]. Hemel [195, 202, 265, 155, 192, 285, 286, 287, 219, 244, 246, 229, 217, 297, 194, 315, 215, 227, 270, 201, 284]. Hempstead [195, 202, 265, 155, 192, 285, 286, 287, 219, 244, 246, 229, 217, 297, 194, 315, 215, 227, 270, 201, 284]. Hennessy [205]. Hermenegildo [318]. Hermes [1690, 2657]. Heterogeneous [2025, 2340, 1342, 1511, 1198, 1993, 2279, 2817]. heterotypic [1415]. heuristic [2526, 132, 2878, 2336]. Heuristics [311, 1476]. HFCommunity [2862]. HI [1858, 1492]. HI-Maude [1858]. Hibernate [2675]. Hidden [1898]. Hierarchical [2323, 801, 1581, 1375, 1971, 2036, 1083, 1425, 1314, 2330, 2441]. hierarchies [317, 504]. Hierarchy [361, 476, 646]. High [1014, 2684, 1031, 1362, 1200, 907, 2073, 607, 1233, 592, 454, 2173, 170, 1404]. high-dimensional [2173]. High-fidelity [1031]. High-level [2684, 607, 1233, 592, 454]. high-performance [1362, 907, 170]. Higher [303, 104, 477, 2211, 36, 2121, 940, 2343, 715, 2795, 1765, 309, 819, 601, 1719, 737, 2860]. Higher-Order [104, 36, 477, 2211, 2121, 940, 715, 2795, 1765, 819, 601, 1719, 737, 2860]. highlight [1486]. highly [2613]. Hill [79, 144]. Hille [194]. Hindley [146, 1721]. Hippocratic [1684]. historical [1383]. histories [2253, 2638]. history [1742, 972]. history-dependent [972]. HML [2831]. Hoare [126, 219, 127, 1105]. Hoare-like [127]. Hoc [1190, 1986, 772]. HOL [2864, 2795]. HOL4 [1672]. holistic [1605]. Holland [267, 268]. home [1559]. homes [1017]. homogeneous [1846]. Homomorphisms [17, 525, 507]. hookable [1589]. horizon [611]. Horn [319, 2122]. Horner [387]. hosts [1189]. hot [2779]. House [207]. HPC [2456]. Hub [2862, 2593]. Hubs [2552]. Huffman [832]. Hugging [2862]. human [727, 2680, 1639, 2581, 2214, 284]. human-robot [2680]. humble [1268]. humorous [1268]. hurricane [2232]. Hybrid [1901, 2127, 2803, 2163, 1442, 2447, 2834, 2411, 1369, 2666, 1858, 1718, 2792, 585, 1523, 2624, 2618, 2028, 1698, 2489, 2127, 2213]. HyDRa [2792]. hygienic [1193, 1955]. hypergraph [1894, 642].

Ian [270, 288, 191]. Ibero [2851]. ICFEM [1373]. ICGT [2702]. ICO [1601]. ICSME [2625]. IDE [2104]. idea [612]. idempotence [1356]. identical [1732]. identification [1445, 2363, 2609]. Identifier [1130]. identifiers [1568]. identify [2872]. identifying [2691, 2185]. identity [1441]. IEC [402, 2405, 2642, 2584, 1967]. IEEE [232, 2775, 2405]. IELR [1217]. if [1122]. IFC [731]. iFM09 [1279]. ignorance [2559]. II [1375, 2127]. III [2803]. illustrated [1037, 1433]. IMA [2604]. images [2593]. immediate [1900]. immunity [727]. impact [2631, 2780, 1836, 2034, 655, 1479, 2111, 1680]. Impacts [2644]. Imperative [647, 947, 525, 533, 1674, 1683, 944, 950, 2333, 410, 2708, 601, 813]. imperialist [2400]. implement [907]. Implementability [139, 1940]. Implementation [1162, 128, 1615, 39, 111, 36, 1871, 524, 2258, 1975, 1829, 329, 399, 770, 1836, 2083, 2544, 1856, 235, 1980, 2299, 2046, 711, 868, 908, 110, 1439, 1301, 1266, 871, 1832, 2158, 1748, 2278, 1555, 2484, 99, 2176, 1763, 227]. Implementations [98, 1663, 332, 2815, 446, 1317, 1560]. implemented [2342]. Implementing [1801, 2098, 22, 774, 1473, 1028, 408, 1362, 507, 312, 2402]. Implications [1639, 2125]. Implicit [1073, 2376, 2634, 308, 920]. implicit-invocation [920]. implicitly [715]. importance [1065, 2350]. improve [2046, 1848, 1305]. Improved [2086, 1383, 481, 1317, 861, 1998]. Improvement [74, 2674, 2746, 2008]. Improvements [2150]. Improving [2485, 2571, 1471, 2001, 313, 1951, 2624, 1815, 1611, 2253, 1219, 1740, 2268, 1799, 1563, 2647, 2132, 780, 876, 739]. ImpUnity [496]. impure [1955]. in-situ [394]. Ince [228, 209]. include [2335]. including [1919]. Inclusion [2469]. inclusive [133]. inCode [2099]. incompatibility [2871]. incomplete [2291, 1936, 2756]. Incompleteness [1230]. inconsistencies [2719, 1296]. inconsistency [509, 627, 1472]. inconsistent [1348]. incorporated [1033]. incorporating [1383]. IncQuery [1818]. Increasing [1531, 41]. Incremental [2121, 1417, 1284, 1462, 593, 639, 65, 1109, 384, 161, 2842, 736]. incrementality [1775]. incrementalization [2596]. indentation [1129]. independence [222, 2083]. independency [41]. Independent [991, 1710, 2269, 2287, 480]. Index [539, 554, 567, 576, 688, 699, 705, 718, 81, 174, 220, 345, 346, 402, 414, 433, 587, 600, 610, 624, 633, 649, 659, 674]. indexed [724]. India [56]. indicators [2745, 2858]. induction [1890, 608, 2208, 133, 2529]. Inductive [1715, 2277, 2064]. Industrial [1428, 460, 1236, 1513, 669, 1996, 1577, 2221, 448, 2440, 2241, 457, 2861, 1753, 591, 2667, 2849, 1550, 801, 619, 2592, 2738]. industry [1629]. inequalities [2208]. inequality [2191]. Infeasible [1783]. infer [1686]. Inference [1683, 1944, 1451, 978, 334, 616, 1669, 1226, 435, 580, 953, 411, 1793, 716, 2165, 853, 1721, 1762, 1740]. inferencing [597]. inferred [2252]. Inferring [2164, 2141]. Infinitary [683]. infinite [2503, 1460, 2435]. infinite-data [2435]. influence [857]. influenced [2560]. Informa [1408]. Informal [14, 187]. Information [342, 573, 947, 1294, 1694, 2289, 2307, 2424, 1898, 1626, 1606, 559, 795, 1463, 2191, 2049, 1508, 173, 2016, 1779]. infrastructure [2104, 1541, 1693, 987, 2380, 834]. inheritance [2705, 604, 1284, 1265, 2371, 1910, 791, 767, 621, 648, 317]. initial [2198]. initiative [295]. injection [1192]. inline [2252]. inlining [740]. Innermost [2739, 2483]. Input [2027, 1658, 2326, 2474]. Input-Output [2027, 2474]. input/output [1658, 2326]. inputs [2609]. insertion [843]. Insider [2577]. insight [1734]. Inspection [2052]. inspector [2195]. inspired [2571]. installation [2871]. instance [1893]. instances [2320, 2186]. instantiation [858, 1697]. Institution [1912, 156]. Institution-based [1912]. institutional [1615, 2050]. instruction [2430, 2676]. instructions [884, 1086]. instrumentation [1684, 2314]. instrumenting [1533]. integer [508]. integer-preserving [508]. integrate [1615]. Integrated [240, 842, 96, 2322, 2209, 2490, 2445, 1850, 987, 1627, 1818, 1279]. Integrating [2234, 1696, 1207, 1873, 434, 2108, 2244, 877]. Integration [1395, 1308, 1153, 1638, 827, 974, 1741, 1511, 1737, 1604, 602, 2050, 307]. Intelligence [2776, 1487, 2858, 1492, 266]. intelligent [1479, 2481, 889, 2754, 2863]. intense [1802]. intensional [1245, 391]. intensive [1295, 2626, 2070, 2465, 1799, 1175]. inter [2103]. inter-language [2103]. Interacción [1504]. interacting [1866, 1858]. Interaction [1380, 1282, 1625, 599, 1602, 1308, 2836, 1271, 2335, 1007, 1184, 2587, 2733, 2309, 284, 1863, 1914]. interaction-based [2587]. interactions [2462, 1918, 1829, 1987, 1337, 467, 1017, 1382]. Interactive [1632, 181, 1504, 1828, 746, 1614, 745, 2168, 1951, 1291, 795, 1507, 1638, 2076, 1508, 1424, 100]. interchange [1630, 1225]. interdependent [1796]. interest [2175, 2638]. Interface [887, 1600, 1601, 245, 1697, 722, 2326, 1604, 1148, 1999, 1636]. interfaces [1656, 1962, 2328, 1606, 2335, 702, 1280, 1738, 320]. interference [2003, 2274, 1959]. interlocking [2039, 543, 2089]. interlockings [1755, 463]. intermediate [835, 2072]. Internal [1244, 450, 2271]. International [141, 20, 126, 2730, 202, 2256, 155, 2875, 514, 1341, 669, 34, 2625, 219, 2562, 229, 2031, 2800, 217, 2775, 2687, 2665, 1809, 1143, 2686, 215, 227, 201, 2426, 272, 2694, 2885, 2727, 1373]. interoperability [2463, 1036]. interpretation [1434, 703, 535, 1943, 506, 379, 486, 369, 376, 842, 908, 941, 487, 844, 1944, 1945, 684, 949, 954, 489, 2654]. Interpretations [128, 725, 370]. interpreted [137]. interpreter [741]. interpreters [829, 832]. Interprocedural [368, 845]. Interrupts [38]. Interscience [572]. intersections [1109]. intertwining [2005]. Interval [1935, 2597, 2352]. Interval-based [1935]. Intervals [1434, 1643]. intracellular [2220]. intraprocedural [1483]. Introducing [2352]. Introduction [532, 2058, 2867, 875, 2226, 2668, 2776, 228, 1312, 956, 2219, 1819, 1974, 514, 925, 1990, 1127, 296, 1525, 215, 1590, 288, 396, 1043, 1076, 1173, 1814, 2097, 202, 146]. intuition [1464]. invalid [1811]. Invariance [59, 2607, 2420]. Invariant [1456, 1151, 2720, 437, 1682]. Invariant-driven [1151]. Invariants [27, 2164, 1683, 1890, 883, 1047, 605, 259, 922, 2186, 949, 2767]. invasively [1691].

Inverse [2683, 632]. inversion [343, 261, 508]. invertible [2683]. inverting [313, 723]. Investigating [1181]. investigation [1897]. investments [811]. invocation [920]. involutions [2658]. involving [713]. IOCL [2168]. IoT [2553, 2278]. IoT-based [2553]. IR [1326]. IR-style [1326]. Irish [1898]. Irsee [2885]. Isabelle [2864, 2507, 2795, 2577, 199]. Isabelle/HOL [2864, 2795]. Isabelle/UTP [2507]. ISBN [195, 141, 184, 291, 516, 202, 572, 183, 515, 265, 155, 192, 231, 243, 514, 245, 154, 299, 295, 156, 151, 228, 206, 146, 319, 203, 521, 213, 289, 285, 286, 287, 219, 244, 207, 246, 182, 229, 218, 153, 321, 172, 296, 217, 208, 297, 193, 180, 168, 298, 67, 204, 171, 267, 268, 194, 216, 315, 320, 181, 317, 143, 145, 214, 266, 269, 271, 397, 169, 574, 390, 215, 173, 227, 185, 575, 226, 318, 270, 209, 290, 730, 201, 288, 166, 205, 196, 284, 230, 191, 247, 248, 249, 250, 316, 524, 573]. ISBN [396]. ised [2028]. island [2174]. islands [1898]. IsNot [962, 963]. ISO [402, 2642, 2405]. ISO/IEC [402, 2642]. ISO/IEC/IEEE [2405]. isochronous [1315]. isolation [2260, 1650]. isomorphism [978]. Issue [1204, 2392, 2852, 1647, 2681, 2866, 1254, 1312, 2625, 2452, 2688, 2682, 2580, 1653, 2563, 2272, 1892, 2007, 2058, 2867, 1088, 1058, 2734, 2839, 919, 1316, 2730, 875, 2256, 2024, 2143, 2226, 2685, 2310, 1863, 1914, 1373, 1385, 1236, 2383, 2875, 1624, 1112, 2586, 2776, 925, 939, 669, 1641, 2276, 2156, 1960, 932, 1286, 1904, 1666, 2038, 1730, 2562, 956, 2374, 2838, 1572, 2573, 2702, 1584, 2219, 1735, 1795, 2031, 2178, 1996, 1191, 2800, 2775, 1600, 1752, 2572, 1525, 1215, 1360, 1400, 1809, 2847, 1035, 1143, 2686, 1259, 1821, 1349, 2075, 1934, 1925, 1941, 1246, 1612]. issue [2044, 2106, 1708, 1590, 2204, 1293, 1346, 2426, 1076, 1173, 1814, 2097, 1118]. Issues [661, 139, 408, 925, 1214, 1908]. ISTA [2863]. IT-governance [1021].

IT-investments [811]. IT-outsourcing [2055]. IT-portfolios [1063]. items [1690]. iterated [1717]. iteration [425, 161]. iterations [653]. Iterative [74, 71, 2830]. iteratively [1652]. itinerary [1613]. IWSC’12 [1730].

J [291, 265, 231, 243, 146, 182, 180, 168, 204, 267, 268, 214, 173, 227, 185, 288, 142, 396]. J. [730]. J2ME [749]. JAAS [864]. JaCaMo [1426]. Jaccie [1532]. Jack [389]. Jackson [341]. Jacques [151, 573]. Jan [388, 298]. January [21, 2409, 2531, 2629, 2724, 2801, 2821]. Jason [1826]. JastAdd [1162, 1045, 1980, 678]. Java [1292, 682, 858, 1442, 2705, 1119, 1389, 1437, 746, 770, 747, 1121, 773, 933, 1533, 2103, 1094, 865, 859, 827, 751, 1468, 583, 2492, 2450, 745, 748, 854, 2612, 1222, 1408, 1737, 1286, 2675, 1291, 1287, 2481, 749, 2207, 1532, 1153, 1406, 1636, 750, 744, 1124, 2386, 848, 860, 774, 1317, 2451, 771, 658, 673, 1396, 1234, 1535, 775, 2585, 1060, 1404, 1078, 2173, 623, 1289, 806, 1290, 2380, 861, 1407, 1059, 1959, 1122, 2402, 1538, 1335, 524]. Java-based [747, 751, 748, 1532, 750]. Java-like [1389, 1437]. java.util.concurrent [849]. JavaScript [1828, 2134, 1837, 2380, 2247, 2593]. Java^{T M} [1061]. JBInsTrace [1533]. JCilk [935]. JCML [1335]. Jean [573]. Jean-Jacques [573]. Jennifer [215]. Jenny [284]. JetsonLEAP [2340]. Jim [206]. JJForester [679]. JML [1911, 789, 1450]. Joeq [834]. John [516, 78, 144, 226, 288]. join [2140]. Joint [30, 339]. Jolie [2056]. Jonathan [146, 289, 573]. Jones [155, 192, 219, 244]. Jonkers [397]. Jorge [389]. Josef [215]. Jovanovich [193]. JP2 [1535]. JPC [2103]. JRE [1533]. JTC1 [402]. Judy [230]. July [20, 2466, 2588, 2689, 2687, 2770, 2868, 2885]. jump [596]. June [232, 2448, 2460, 2574, 2677, 2760]. just [2412, 2132]. just-in-time [2412, 2132]. Justifications [28]. JVM [1022, 830, 2779]. JVM-like [1022].

Kahn [2765]. KALA [1071]. Kaldewaij [315]. Kapur [243]. Keller [284]. Kenneth [295]. Kent [298]. kernel [2640, 2742, 1071]. kernels [568, 2162]. key [2745, 1569, 1780]. keyword [1223, 436]. Khasidashvili [573]. Kiama [1976]. Kin2 [2057]. Kinect [2135, 2057, 2620]. Kingdom [195, 202, 265, 155, 192, 243, 245, 206, 203, 213, 289, 285, 286, 287, 219, 244, 207, 246, 218, 321, 296, 217, 208, 297, 193, 298, 204, 194, 315, 317, 143, 214, 266, 269, 215, 270, 209, 201, 288, 284, 247, 248, 249, 250]. KinZ [2620]. Kit [772]. KLAIM [1856]. Klaus [172]. Kleene [982, 983]. Kloos [171]. Kloster [2885]. Knowledge [2283, 2011, 1012, 2007, 2016, 2013, 1712, 2010, 2871, 2540, 618, 998, 2109, 2008, 317, 1664, 2810, 2009, 2012]. Knowledge-based [2007]. known [2015]. Knows [1140]. Konstancin [2886]. Kosciusco [21]. Kraken [2587, 2733]. Kripke [1859]. Kroger [154]. Kronsjö [145]. Kruskal [39]. Kuei [153]. Kurt [151]. Kushilevitz [515].

L [243, 299, 391, 208, 181, 145, 397]. Label [382]. labeled [2609]. labeling [2758]. labelled [2812, 2547]. LabVIEW [2755]. lacking [1462]. Lake [21]. Lamb [217]. lambda [485, 1205, 296, 149]. lambda-calculus [149]. Lamport [1474]. Lane [291]. Languag [2002]. Language [2839, 6, 1607, 51, 1090, 93, 2209, 1773, 88, 87, 106, 227, 498, 573, 632, 1419, 1100, 2732, 2567, 1485, 2440, 2174, 241, 1905, 1601, 1975, 293, 329, 1409, 188, 1361, 1906, 589, 1081, 1932, 2083, 1723, 2534, 2103, 2211, 340, 1846, 944, 742, 900, 1759, 1602, 691, 1071, 2482, 245, 1506, 71, 1195, 797, 444, 41, 2579, 1557, 2210, 2442, 2590, 451, 2708, 1690, 2545, 182, 1679, 607, 452, 2475, 1743, 1883, 2464, 297, 2445, 1933, 2287, 520, 2657, 1928, 665, 601, 910, 2384, 1278, 2014, 1387, 1548, 807, 2309, 1491, 504, 1879, 1608]. language [2150, 1720, 1258, 2385, 2772, 64, 2504, 1788, 1028, 1041, 1128, 2501, 582, 247, 248, 249, 250, 1335, 2341, 1480]. Language-based [2839, 1773]. language-independency [41]. language-independent [2287]. Language-integrated [2209, 2445]. Languages [1546, 1204, 2730, 1677, 2685, 2310, 2338, 661, 1385, 2246, 2562, 2374, 2452, 1746, 1989, 1600, 1215, 1416, 352, 1143, 1349, 157, 1465, 9, 20, 1831, 525, 476, 643, 901, 1437, 1895, 1674, 1221, 490, 1187, 1417, 677, 940, 304, 925, 926, 2154, 715, 932, 908, 1990, 1822, 1036, 276, 1037, 368, 2836, 1303, 2252, 1928, 1768, 1543, 2568, 666, 176, 1847, 224, 644, 301, 1719, 1180, 386, 1341, 1830, 1360, 1400, 1448, 231, 203, 208, 398, 317]. Larch [88, 87]. Large [1625, 2062, 1419, 2163, 1295, 1541, 1369, 2169, 1657, 1198, 2267, 2570, 673, 2830, 1258, 2344, 2373, 1519]. Large-scale [1625, 2062, 1541, 1198, 2570, 2344, 2373]. Larissa [943]. latency [1291]. latent [1476]. Latham [288]. lattice [1109]. lattices [2352]. Laura [498]. Laurie [284]. Lawful [234]. Laws [361, 1594]. layer [1611, 2215, 1301, 2278, 424, 1445]. layer-oriented [424]. layered [1718, 922]. Layers [25, 2371]. layouts [2504]. laziness [1799]. Lazy [411, 1284, 983, 1252, 445, 386]. LCF [69, 181]. LCSD [1246]. LDTA [1607, 1025, 1974]. LDTA’08 [1480]. lead [2190, 1629]. leader [464, 625]. Leaf [426]. leaks [1787]. Lean [2479]. learn [970]. learned [2829, 1968, 729]. Learning [1649, 1612, 2679, 2583, 2808, 1615, 2870, 2814, 2624, 2844, 2812, 1613, 2691, 1616, 2050, 1618, 2048, 2757, 2813, 2743, 2045]. learning-based [2870]. learning-driven [2844]. Lecture [168, 171, 216, 191]. Leeuwen [388]. Left [1747, 264, 1223]. Left-bottom [264]. legacy [1443, 877, 1101, 657]. Legal [2726]. LEMMA2Jolie [2769]. Lending [1952]. length [1130]. lengths [1755]. lenses [2331]. Lenzerini [317]. Lessons [2829, 2055, 1968, 729]. let [787]. Letter [976]. level [2322, 2241, 1348, 2061, 1648, 1425, 797, 1557, 607, 1233, 463, 2661, 592, 2684, 454, 1125, 2604, 1547, 1483, 2112, 739, 398]. levels [2172, 1399, 2638, 1560]. leveraging [2450, 1815, 2858, 2380]. Lévi [573]. Lexical [135]. lib [2797]. libraries [2321, 1720, 1673, 712]. library [2104, 2103, 2408, 1253, 2331, 750, 658, 1247, 1313, 1246, 1251, 2173, 1077, 2704]. library-centric [1246]. LICS [232]. life [1618]. lifecycle [761]. lift [1878]. lifted [2235, 2632]. Light [1942]. Lightweight [616, 2263, 770, 1554, 2853, 2382, 2744, 2657, 1345, 1494, 1512, 824]. like [927, 1496, 1389, 1437, 1022, 1745, 843, 520, 127, 224]. likely [1047]. limited [1130, 1414, 2585]. Linda [927, 497, 861]. Linda-like [927]. Lindemann [572]. Line [2800, 2190, 1564, 1514, 765, 1443, 889, 520, 994, 1512, 762, 820, 1515]. Linear [917, 111, 1588, 122, 23, 504, 2831, 2163, 2164, 1860, 693, 2729, 1369, 1314, 1685, 1501, 1613, 353, 2684, 1247, 1452, 177, 529]. linear-bounded [1314]. Linear-time [504, 2831, 529]. Linearisation [159]. linearizability [1754]. linearly [1264]. lines [2020, 1801, 1509, 1513, 1681, 1750, 1129, 2284, 2273, 767, 2311, 763, 766, 2719, 1516, 2334]. linguistic [1769]. linkage [598, 1665]. Linked [2489, 1911, 1621]. Linking [1499, 2054, 2257]. links [1639]. Linter [2710]. Linux [1664]. Liskov [144]. Lisp [138, 247, 191]. LISPer [167]. List [361, 92, 507]. List-Copying [92]. Listener [1291]. Lister [168]. lists [1725, 371]. little [167]. Live [787, 1669, 686, 2253, 2476, 1818, 2083]. lived [403, 481]. Livelock [58].

Livelock-Free [58]. lively [2493]. Liveness [59, 2529, 2728]. load [1807]. loading [1290]. Lobo [389]. Local [2180, 1894, 2084, 1598, 1992, 846]. locality [1317, 780, 2504, 2140]. Locality-Sensitive [2140]. localization [551, 1478, 756]. localized [1860]. Locally [2640, 474]. Lock [967, 1110, 853]. Lock-free [967]. lock-freedom [1110]. locks [2382]. Loecks [231]. Loeckx [151]. Löf [298, 108]. Log [2736]. logging [2638]. Logic [493, 1854, 120, 130, 573, 1570, 2818, 947, 408, 232, 425, 1860, 476, 334, 2152, 1917, 2731, 443, 457, 2787, 2211, 1363, 1135, 703, 569, 1497, 1055, 2302, 2397, 2713, 29, 514, 154, 717, 2210, 1464, 312, 2664, 2795, 289, 1701, 333, 2395, 1510, 1520, 1188, 452, 1558, 453, 389, 2227, 523, 353, 698, 2683, 1582, 169, 1676, 2352, 2614, 1742, 582, 1281, 142, 2542, 2556, 1105, 1913, 972, 1187, 510, 213, 180, 181, 318]. logic-based [476]. logical [725, 505, 2327, 2436, 1296, 340, 1444, 813, 2845, 948, 1332]. Logics [240, 193, 2028, 391]. LogLogics [972]. London [141, 126, 184, 516, 183, 243, 295, 156, 34, 389, 180, 179, 214, 266, 144, 185, 196, 152, 191, 247, 248, 249, 250]. long [1095, 403, 481]. long-lived [403, 481]. long-running [1095]. Longest [310, 1644]. longest-edge [1644]. Loomes [206]. Loop [2607, 27, 1020, 1158, 1313]. Loops [27, 1667, 1406, 311, 1453]. Lost [2600]. LOTOS [1108, 379, 267]. Louvain [1711]. low [1557, 2279, 739]. low-level [1557, 739]. lowering [2132]. LR [2641, 1217, 331, 2764]. LR-parsing [331]. LR/GLR [2641]. LRU [279]. LSC [787]. LSI [1305]. Ltd [156]. LTL [1376, 1282, 1241, 1357, 2874]. Lua [2442]. Lub [2254]. LY294002 [2823].

M [184, 265, 391, 207, 182, 168, 298, 267, 179, 397, 730, 205, 249]. MA [388, 320, 167, 318, 166, 205, 170, 316, 271]. Mac [291]. MACH [2101]. machine [2679, 2230, 2742, 116, 2339, 831, 868, 2814, 2353, 2233, 2624, 2844, 2812, 1152, 2158, 2789, 673, 504, 1290, 834, 2757, 1314, 1535]. machine-checked [2230]. Machines [2497, 1392, 96, 1692, 2404, 184, 1901, 2127, 2407, 2090, 2060, 859, 571, 1561, 829, 2401, 1425, 2323, 1718, 833, 2804, 2658, 2842, 1694]. macros [1193, 1749]. macroscopic [2308]. made [1958]. Maier [48]. maintainability [1299]. Maintenance [2625, 965, 1440, 824, 923]. make [188]. makers [2541]. makes [2843]. Making [2013, 1962, 459, 2107, 1680, 2297, 1182]. malicious [1189]. Management [2200, 1340, 2344, 1419, 1712, 2711, 2877, 764, 1688, 760, 692, 758, 2586, 912, 458, 1463, 1845, 1569, 2008, 987, 774, 913, 761, 846, 914, 2454, 992, 650, 755, 1181, 2304, 1073, 2635, 985]. managerial [1629]. managers [2147]. Managing [2104, 1840, 816, 1214, 2033, 994, 960]. Manes [179, 390]. manifesting [2265]. Manipulate [150]. manipulating [2792, 1576]. Manipulation [1896, 1716, 2118, 1954, 919, 773, 1259]. manual [2783, 914]. Manuel [318]. Many [2307, 2227, 2739, 2091]. Many-Core [2091]. many-sorted [2227, 2739]. Many-to-many [2307]. Map [1245]. Maple [1260]. mapping [2825, 2675, 2861, 1443, 2287]. mappings [2679, 2474]. MapReduce [1054, 2141]. maps [1062]. Marc [196]. March [9, 2433, 2548, 2649, 442, 2740, 2840]. Marcus [320]. Maria [317]. mark&sweep [967]. market [2738]. marking [148]. Markov [1898, 1524, 1951, 545, 1657, 844]. markup [1601]. MARS [2524]. MARTE [2286, 1908, 2199]. MARTE/CCSL [2286, 1908, 2199]. Martin [206, 298, 108]. Martin-Löf [298, 108]. MAS [1422, 994]. MASCOT [818]. Mason [191]. mass [49, 824, 923]. masses [1338]. Massive [421]. Master [174, 346]. Masterpiece [114]. Matching [1438, 197, 2262, 1223, 2827, 1709, 552, 176, 1338, 1503, 436, 696].

Mathematical [46, 146, 226, 272, 1642, 1455, 401, 869, 513, 1218]. Mathematics [417, 228, 1158, 2885, 575, 344, 156, 1495, 1593, 575, 184, 206]. MATLAB [2455, 2057, 2620]. matrices [1646]. Matrix [1585, 394, 2173, 508]. Matthias [167]. Maude [1563, 2854, 1375, 2200, 1857, 999, 1856, 1858, 1151, 1965]. Maurizio [317]. Max [2766]. Max-SAT-based [2766]. maximality [2628]. maximum [2763]. May [2443, 2565, 2670, 2880, 2886, 2665, 2855]. MAZE [2066]. McGettrick [143]. McGraw [79, 144]. McGraw-Hill [79, 144]. Mclaren [114]. mCRL2 [2832]. MD [49]. MDA [991, 1511]. MDA-based [991]. MDD [1605]. MDE [1632]. MDE-based [1632]. me [2645, 244]. Mealy [2789]. Mealy-machine [2789]. mean [1921]. mean-field [1921]. meaning [1218, 338]. Means [1708, 512, 134, 2011]. measure [2591, 2340]. Measurement [2745, 2108, 2111, 2109, 2107, 2106]. measures [2553]. Measuring [2359, 1472, 916, 876, 1291]. Mechanical [2596, 1474, 1331, 1267]. Mechanised [1330]. Mechanising [1115]. Mechanism [84, 2215, 2427, 766]. Mechanisms [583, 2300, 1308]. Mechanization [2506]. mechanized [1105]. mechanizing [420]. media [2009]. Mediator [2464]. medical [2234, 2389, 2525, 2581]. Medicine [2525]. Medicine-by-wire [2525]. Meertens [1]. meet [1434, 477]. meets [2837]. Mehlhorn [231]. Mellish [247, 248, 250]. Melvin [169]. Memoized [2343]. Memoizing [444]. Memory [1340, 2864, 1471, 1130, 1115, 936, 933, 850, 2672, 2511, 1241, 1429, 2080, 2260, 912, 2146, 1565, 843, 1433, 937, 1478, 2132, 1548, 913, 846, 2271, 914, 2406, 915, 1073]. menu [708]. Mercury [2102]. Merge [140, 1441]. Merging [2698, 1961, 1963]. Message [2085, 94, 778, 1135, 608, 2857, 586, 913, 392, 646]. Message-Based [94]. message-passing [1135, 608, 2857, 392]. messages [2376]. messaging [1068, 1586]. MESSENGERS [495]. meta [1656, 2327, 1836, 1806, 1490, 1961, 798, 1893, 1444, 2131, 2186, 2377, 1720]. meta- [2327]. meta-circular [1720]. meta-control [2131]. meta-data [1806]. meta-model [1490, 798, 1444]. meta-modelling [2377]. meta-models [1961, 1893, 2186]. meta-theory [1656]. meta-tracing [1836]. metacompilation [1166]. Metadata [1628, 1847, 2380]. Metadata-based [1628]. metafront [1026]. MeTAGeM [1815]. MeTAGeM-Trace [1815]. metaheuristic [1807]. Metal [37]. metalanguage [216]. Metamodeling [1265, 635]. metamodelling [1760]. Metamorphisms [979]. MetaOCaml [906, 909]. metaprogramming [2817, 1078, 2528]. metaprograms [1473]. Metasoft [216]. meter [448]. MethApp4Mob [2793]. Method [2215, 341, 2485, 221, 2645, 541, 816, 2872, 2185, 71, 1493, 740, 347, 1573, 312, 450, 1383, 2744, 2071, 204, 1604, 2805, 1707, 1605, 1488, 2764, 2375, 736]. methodological [1504, 973, 778, 2793, 1148]. methodologies [443, 67]. Methodology [14, 15, 2091, 639, 1421, 571, 618, 1507, 565, 290]. Methods [1428, 2799, 2697, 1373, 2457, 1236, 2383, 2535, 669, 1909, 106, 2882, 1996, 1279, 1329, 2572, 2665, 1670, 2234, 448, 2742, 2232, 383, 612, 785, 1420, 1731, 299, 1844, 2848, 1320, 2172, 2779, 2886, 1199, 617, 1550, 1017, 1326, 2734]. Metrô [1432]. metric [2109, 561, 2405]. Metrics [1650, 1124, 857, 2488, 1384, 1540, 2292, 2755]. Meyer [195]. MFCS’91 [272]. Michael [192, 514, 227, 170]. Michaelson [296]. Michel [1268]. microcontroller [1434, 1551, 1239]. microengineering [1778]. microprocessor [422, 2000]. microservice [2769]. Middelburg [730]. middleware [1089, 1342, 1092, 1494, 990]. migration [2091, 1451, 1313]. military [1339]. Mills [207, 226]. Milner [574, 1721]. mind [269]. mindfulness [2780]. mined [2540]. Minimal [323, 39, 1248, 1840, 1217, 695]. MiniMAO [944]. Minimizing [2601]. Mining [2174, 1224, 1347, 2252, 2633, 2590, 2014, 1732, 2268]. MiniZinc [1932]. Minker [389]. MIPS [1845]. Miranda [234]. misalignments [816]. misbehaviour [1864]. mismatch [511]. mission [2465]. mission-related [2465]. missions [891]. Mitigating [1724]. Mixed [1632]. ML [1905, 1032, 843, 489]. ML-like [843]. MMINT [2811]. MMINT-A [2811]. MMU [2260]. Mobile [574, 751, 2793, 1986, 647, 1515, 586, 684, 1007, 1467, 1013, 1494, 666, 2045, 1190]. MobiLine [1515]. mobility [896, 1492]. MOBS [2579]. Modal [1374, 2660, 1521, 1583, 2078, 457, 2326, 351, 2313]. Modalities [120]. Mode [671, 1435, 2806]. Mode-Automata [671]. mode-rich [1435]. Model [1810, 1866, 1860, 2006, 2328, 2212, 1704, 1630, 493, 1751, 1626, 1170, 1624, 2866, 120, 380, 1092, 809, 1629, 1091, 2850, 1745, 2599, 1433, 1638, 591, 1851, 2076, 2042, 1323, 1555, 2048, 2529, 1040, 90, 2524, 1088, 1038, 1712, 2303, 784, 2418, 2711, 263, 1376, 2583, 1202, 2370, 2519, 1343, 2115, 1841, 1671, 2785, 2829, 1115, 2642, 1912, 2025, 2000, 2526, 2787, 1490, 2248, 1090, 1489, 655, 1113, 2656, 638, 820, 1505, 1631, 2334, 2169, 2828, 1628, 1937, 2811, 1368, 1241, 2759, 2823, 1337, 717, 1322, 1327, 2259, 1575, 2765, 998, 2269, 2809, 1772, 2804]. model [1815, 1983, 1822, 1036, 1079, 2835, 2194, 2746, 1145, 1591, 720, 2229, 1037, 2581, 1054, 1921, 1859, 581, 830, 798, 2102, 2114, 672, 565, 2008, 2187, 1544, 1936, 1444, 2847, 890, 2348, 1627, 2661, 2849, 1940, 1183, 1035, 1103, 673, 504, 767, 1652, 2363, 2830, 2336, 780, 1547, 872, 2874, 2029, 2355, 1818, 1041, 1920, 2564, 763, 455, 152, 971, 920, 1334, 2493, 1240, 2722, 1238, 1625, 1898, 1422, 1960, 2715, 937]. model- [1334]. Model-Based [2212, 1810, 2006, 1751, 2076, 1555, 2785, 2811, 2804, 2849, 1652, 2830]. model-checker [2874]. Model-checking [809, 591, 1921, 672, 1240]. Model-Driven [2866, 1630, 1626, 1629, 1091, 1745, 2599, 1638, 2048, 2524, 2303, 1202, 2115, 2829, 1912, 2248, 1090, 1489, 1113, 1505, 1631, 2828, 1628, 1322, 998, 1772, 1625, 1422]. model-extraction [1327]. model-integrated [1627]. model-to-model [1036]. modelers [970]. Modelica [1162]. Modeling [895, 572, 2152, 1804, 892, 461, 1693, 2792, 2040, 96, 1233, 1973, 1001, 1467, 1396, 1892, 763, 2427, 2015, 1065, 1455, 1227, 1932, 2405, 2551, 821, 2285, 645, 2286, 1858, 636, 788, 2170, 2612, 1761, 2290, 1379, 339, 1993, 2464, 1616, 1791, 993, 2396, 2009, 2144, 2308, 1351, 2865]. modelled [804]. Modelling [2653, 2213, 1857, 980, 2555, 1623, 341, 2406, 971, 1814, 2732, 1817, 2090, 1907, 2808, 1297, 1394, 1755, 1303, 1605, 2309, 2377, 2221, 2089, 1285]. Models [2730, 1385, 1341, 2562, 2374, 388, 1419, 973, 1939, 1375, 2020, 764, 663, 1841, 2034, 490, 917, 2672, 787, 2169, 2397, 1856, 646, 2774, 925, 2769, 2490, 1393, 1697, 451, 2077, 1815, 1510, 2194, 1961, 2229, 1537, 2618, 898, 2698, 2812, 2141, 1893, 1370, 2027, 2348, 987, 2488, 1304, 2186, 2420, 1313, 2604, 1210, 1963, 2810, 1232, 1732, 1566, 2311, 992, 1713, 2406, 1180, 2833, 2755, 2505, 2728, 991, 321]. MODELS’2016 [2167]. ModelSet [2812]. MODELSWARD [1960]. modern [912, 524]. modernization [1322]. modifications [777]. Modular [1918, 1975, 2059, 637, 1609, 1470, 922, 1589, 1610, 1846, 1564, 1045, 1977, 1765, 2047, 2484, 1028]. modularity [41, 1611]. modularization [1037]. module [999, 439, 271]. Modules [1870, 677, 1750, 450, 1101]. modulo [2277, 2366, 1950, 1861]. MOGA [2597]. MOGA-NN [2597]. monadic [2003, 2188, 2341]. Monadification [735]. monads [383, 1255, 2533, 558, 1473]. monitor [2332]. Monitoring [2858, 1707, 2831, 2080, 2736, 1999, 2788]. monitors [2494, 1547]. Monographs [154, 172, 271]. Monolithic [1765]. monomorphic [1929]. MontiGem [2829]. Moore [696]. Moore-style [696]. Moose [1526]. Morris [163]. Most [85, 459, 1802]. Motif [285, 286, 287]. motion [578]. MoULDyS [2788]. movements [2512]. Moving [1358]. MPC [1495]. MPI [1237, 1402]. MR [471, 19]. MT [1040]. MTL [2735]. mTOR [2823]. mu [672, 780]. mu-calculus [672, 780]. Multi [1426, 1255, 2425, 827, 2329, 1427, 2592, 2061, 2328, 1554, 1987, 2537, 2169, 2274, 1844, 2721, 2092, 1188, 2160, 2066, 2744, 1399, 2008, 910, 826, 1199, 2295, 2587, 2363, 2766, 1006, 2334, 2545]. multi-actor [1987]. Multi-agent [1426, 2425, 1427, 2721, 1188, 2160, 2066, 826, 2766, 1006]. Multi-Bach [2545]. multi-class [1399]. multi-component [2363]. multi-concept [2537]. Multi-core [2329, 2092, 2592]. multi-database [2169]. multi-device [2587]. multi-level [2061]. multi-methods [1844, 1199]. multi-model [2008]. multi-objective [2744]. Multi-paradigm [827]. multi-platform [2328]. Multi-stage [1255, 910]. Multi-task [2592]. multi-tenancy [2274]. multi-threaded [1554]. multi-view [2295]. multiagent [1424]. multicapabilities [958]. multicellular [2220]. Multicore [1345, 2370, 1452]. multicores [1947]. multidimensional [2070]. Multifunctional [1227]. multilevel [2370]. multilingual [1790]. multimethods [1439]. multimodal [1602]. Multiparty [491, 1829]. Multiple [2127, 493, 2072, 2135, 1596, 1284, 1265, 1207, 2267, 497, 2279, 436]. multiplier [278]. multirate [1654, 1888]. multirelations [980]. multiset [1754]. multistage [2314]. multithreaded [1067, 1557]. München [9]. Mundy [243]. Mungo [2207]. Munich [9]. Muon [1519]. music [2076]. Mutable [770, 2243, 2458]. mutant [1724]. mutants [1724]. Mutating [2451]. Mutation [1399, 1397, 1810, 2006, 1398, 1475, 2269, 1397]. Mutual [1881, 75, 1264, 1433, 470]. MVL [2627]. MyType [1438].

N [299, 143, 201, 249]. N.92 [402]. N.S.W [21]. name [2713]. named [2054]. named-entity [2054]. Naming [33, 583]. Nardi [317]. narrative [1013]. narrative-based [1013]. Narrowing [2878, 2005, 1950]. NASA [891]. Nash [2766]. NAT2TEST [1723]. Native [1636]. natively [2314]. Natural [2839, 2440, 2174, 1723, 2534, 1743, 2384, 2014, 2385, 249, 247, 248, 250]. natural-language [2534, 2385, 249]. naturalistic [2287]. navigation [1062, 1014]. NCL [1913]. Near [2585, 2886]. Near-omniscient [2585]. nearest [2651]. nearest-neighbor [2651]. need [2150]. needed [2480]. needs [636, 544]. negative [2514, 1267]. Negotiation [491, 1709, 1307]. neighbor [2651]. Neighborhood [2870]. NeoEMF [2169]. Neopost [1555]. Nested [937, 1245, 2186, 1515]. net [2079, 474, 380, 2579, 607, 374]. Netherlands [388, 267, 268]. Netherlands/The [388]. Nets [2219, 1694, 1918, 1952, 2181, 2222, 2223, 355, 2379, 2224, 1517, 2220, 592, 1324, 804, 572]. network [1661, 758, 1131, 957, 2870, 1679, 1358, 2796, 1387, 1491, 2762, 1998]. Network-on-Chip [1998]. Networked [1520]. Networks [82, 2854, 870, 1559, 757, 1342, 1388, 464, 2765, 1986, 1545, 2220, 1569, 1414, 1267, 756, 805, 2427, 2375, 1190]. neural [2870, 2762]. neuron [2659]. newbies’ [2559]. news [1690]. Newton [475]. NewWave [2550]. Next [891, 1421, 2750, 2403, 990]. next-generation [1421, 990]. nfer [2816, 2736]. NiCad [1536]. Nielson [398]. Nisan [515]. NISQ [2643]. NJ [514, 78, 67]. NLBSE’22 [2839]. NMM [1062]. NN [2597]. no [471, 19, 1684]. Noam [515]. Nobuo [442]. NOD4J [2585]. node [1568]. nomadic [1387]. Non [2303, 2679, 1461, 2212, 482, 1691, 1225, 1734, 2188, 2547, 1728, 993, 2564, 2144, 1217]. non-bisimilar [2547]. non-cloned [1734]. non-computable [1225]. non-cyclicity [1728]. non-determinism [2188]. non-deterministic [1461, 2564]. Non-functional [2303, 2679, 2212]. non-invasively [1691]. non-LR [1217]. non-recursive [2144]. non-secure [993]. non-standard [482]. nonblocking [1966]. nondeterminacy [981]. nondeterminism [2500, 980]. Nondeterministic [140, 85, 1460, 76, 627]. noninterference [2672]. Nonlinear [159, 121, 475]. NonNull [1266]. nonterminating [631]. nontrivial [332]. Nordström [298]. normalisers [1718]. North [267, 268, 318]. North-Holland [267, 268]. notation [1616, 1595]. notations [636, 1338]. Note [2184, 2183, 27, 1097]. Notes [2628, 168, 171, 16, 216, 191]. notion [1357]. notions [1184]. novel [1216]. November [2516, 2615, 2706, 1172]. NTG [35]. null [1848]. null-dereference [1848]. Number [31, 122, 1243, 83, 1476, 1460, 2837]. Number-Theoretic [31, 2837]. numbering [951]. numbers [512]. numeric [1686]. numerical [1077, 1725, 2632, 446, 1644, 1641]. NVP [2823]. NVP-BEZ235 [2823].

Oakley [295]. obedient [1865]. Oberon [1980, 1976]. Oberon-0 [1980, 1976]. Oberon0 [1977, 1978]. Obituary [442]. Object [491, 1074, 679, 1835, 771, 2693, 1288, 734, 670, 655, 636, 1464, 932, 1352, 2675, 825, 1983, 1822, 2878, 453, 368, 922, 910, 373, 2358, 1548, 1321, 1732, 1783, 800, 915, 1538, 2341, 1204, 1436, 1800, 2130, 1153, 1763, 195, 246]. object-aware [915]. object-based [453, 1321]. Object-focused [1835]. object-modeling [636]. Object-Oriented [2693, 1074, 679, 771, 734, 670, 655, 1464, 932, 2878, 368, 910, 373, 2358, 1548, 1732, 1783, 800, 1538, 1204, 1436, 1800, 2130, 1153, 195, 246]. object-relational [2675]. objective [2744, 1979]. Objectives [1507, 1796]. Objects [2392, 254, 407, 1377, 332, 1669, 60, 548, 2379, 2290, 2748, 647, 852, 283, 2048, 2564, 2406]. objets [60]. observation [1988]. Observations [854, 1553]. Observatory [1177]. observed [758]. observer [2579]. observing [2361]. ocaml [2331]. Occam [192, 155, 218, 516]. occurrence [2673]. Occurrences [7]. OCL [1674, 1039, 1046, 2168, 637, 1766, 2186, 2767]. OCL-based [1674]. octahedron [952]. October [2398, 2495, 2508, 2610]. ODEs [2471]. Odyssey [987]. Odyssey-SCM [987]. off [2513]. Offer [1915]. Offering [1900]. office [316]. offs [2205]. O’Haskell [629]. Ohloh [1789]. Okasaki [521]. old [545, 1595, 2176]. omniscient [2585]. on-line [889, 520]. On-the-fly [1376, 1921, 2032, 672, 1367]. onboard [1575]. one [2469]. one-unambiguous [2469]. online [1648, 2552]. only [1013]. ontological [2298]. ontologies [889]. Ontology [1203, 990, 2877, 2116, 1710, 2108, 1761, 1310, 1709, 2214]. ontology-based [2108, 2214]. Ontology-driven [1203, 990, 1761, 1310]. OnTrack [2848]. OOPS [1204, 1436, 1800, 2130]. Open [2294, 1661, 1617, 285, 286, 287, 284, 1663, 2359, 1541, 1918, 2325, 569, 1662, 1422, 2599, 1301, 1199, 2810, 2373, 2404, 1659]. open-source [1663, 1541]. OpenCert [1659]. OpenCL [2136]. OpenDSA [1614]. OpenFlow [2427]. operated [1431].

Operating [94, 1889, 1101]. Operational [363, 1418, 1845, 2836, 2381, 445]. Operationally [2864]. operations [1390, 517, 1056, 2290, 2173, 2110]. operator [1675]. operators [839, 1915, 1843]. optics [2445]. Optimal [2289, 506, 1376, 2703, 2680, 1689, 2513, 2766, 2279]. Optimisation [159, 1414]. optimiser [501]. optimistic [1290]. Optimization [572, 2277, 1807, 1711, 842, 1427, 2132, 2336, 1608, 2863]. optimizations [2053]. Optimized [1829, 2080, 2314]. Optimizing [1957, 1643, 106, 1450, 2432, 869, 2491]. Option [1997]. oracles [2683]. Orchestrating [2291, 1387]. orchestration [1144, 2424, 2344, 1006]. orchestrator [2635]. orchestrators [2645]. Order [303, 104, 1861, 36, 2119, 1282, 313, 2121, 474, 477, 2211, 940, 2343, 2210, 715, 2795, 1765, 309, 2227, 2252, 819, 601, 527, 582, 1719, 737, 2860, 407]. Order-sorted [1861]. ordered [1963, 392]. Ordering [708, 1930]. orders [2020]. organisation [2012, 1919]. organising [1921]. organization [2010, 992]. organizational [1712, 1629]. organizations [1083, 2267]. organized [1415]. Organizing [1410]. Oriented [2693, 10, 943, 2303, 1319, 1316, 1202, 1426, 734, 2115, 1420, 1797, 692, 670, 2248, 655, 1074, 1306, 944, 821, 2060, 645, 1681, 1337, 939, 1318, 1464, 932, 678, 1123, 2804, 1983, 1198, 2878, 2702, 679, 2160, 368, 2618, 2284, 1301, 1501, 1562, 1370, 910, 373, 2358, 1548, 1384, 771, 1450, 1234, 1451, 424, 1874, 1560, 1732, 1530, 1783, 1028, 800, 1538, 2341, 1204, 1436, 1800, 2130, 1795, 1153, 1423, 2004, 2676, 195, 246]. origin [123]. original [1402]. Origins [290]. OSF [285, 286, 287]. OSF/Motif [285, 286, 287]. Ostroff [289]. other [2807]. our [611]. outlier [2268]. Output [2027, 23, 1658, 2326, 2474]. outputs [2609]. outsourcing [2055, 810]. over- [948]. overfitting [2872]. Overflow [2518]. overflows [1699]. overhead [1255, 2205, 2654]. overheads [2639]. overloading [1119, 381]. overtaking [1264]. Ovm [835]. Owen [295]. OWL [1203]. OWL-DL [1203]. ownership [1389, 872, 1073]. Oxford [245, 228, 391, 213, 298, 269, 169, 226, 209, 142]. OZ [791]. oZone [1445].

P [202, 146, 267, 268, 167, 290]. PA [232]. Pacific [2272]. pack [156]. package [2871, 2600, 1637]. packages [1637, 2593]. Packet [58]. packing [693]. Pacta [2726]. pages [572, 155, 154, 146, 521, 909, 193, 171, 179, 320, 317, 143, 574, 144, 575, 316, 524, 573, 126, 396]. pain [2120]. pair [2483, 2189]. pairwise [1153]. Pallene [2442]. PALS [1888]. PAMELA [2612]. PAMOJA [2617, 2772]. paperback [265, 155, 192, 231, 243, 299, 228, 206, 146, 319, 203, 213, 289, 285, 286, 287, 244, 207, 229, 296, 217, 208, 297, 204, 216, 315, 320, 266, 215, 575, 226, 270, 209, 201, 288, 205, 284, 230, 191, 249, 316, 510]. Papers [1341, 1254, 1653, 1716, 2149, 9, 20, 80, 232, 272, 2694, 1355, 1518, 2118, 2730, 1702, 2068, 1274, 2024, 2143, 1025, 2685, 2799, 2697, 2338, 2457, 1954, 2283, 1804, 2875, 2239, 2391, 1495, 1687, 1854, 1919, 1392, 1909, 1593, 1726, 1960, 1948, 2582, 2737, 2562, 2688, 1746, 2367, 2882, 2800, 1279, 1329, 2687, 2665, 2885, 1543, 1809, 2018, 2002, 2686, 1259, 1692, 1349, 1550, 1670, 2693, 1346, 2727, 2851, 573, 669]. PapyGame [2785]. Papyrus [2809]. paradigm [1415, 1230, 1011, 827, 1092, 1464, 2298, 2597, 123, 871, 327, 1013, 1273, 1706]. paradigms [2217, 297]. paragraph [89, 529]. Parallel [1958, 2411, 599, 1406, 309, 1824, 1547, 342, 2140, 225, 233, 1376, 278, 2401, 449, 967, 372, 507, 347, 562, 607, 275, 327, 1548, 2344, 277, 270, 218]. Parallelism [2092, 1121, 1927, 1547]. parallelization [421, 2481, 2345, 1453]. parallelotopes [2164]. parameter [2854, 2411]. Parameterised [2029, 2435, 2513]. Parameterized [1987, 2328, 110]. parameters [2322, 509, 964, 162]. Parametric [615, 2854, 1801, 942]. Parglare [2641]. Paris [184, 179]. Parma [1077]. Parse [1777, 1484]. parse-tree [1484]. parser [2641, 1217, 2250]. parsers [1362, 1033]. Parsing [1791, 1748, 2174, 1958, 1026, 108, 331, 902, 198, 2250, 1634, 1678, 2764, 1164, 2023, 1587, 100, 2431, 2074, 1747]. part [513]. part-of [513]. Partial [1282, 474, 1260, 690, 52, 530, 908, 1701, 2835, 1988, 520, 527, 1576, 2420, 1290, 1716, 2118, 1954, 1259]. partial-correctness [2420]. partial-order [527]. partiality [523]. Partially [137, 1086, 1992, 2242]. partially-distributed [1992, 2242]. participant [2560]. Particular [102]. Partition [1380, 513]. Partitioning [97, 2213, 1429, 1547]. partitions [1644]. partner [2645, 2635]. parts [431]. Pascal [520, 194, 288, 230]. Pascal-like [520]. passing [1694, 1004, 1135, 608, 2857, 913, 392]. past [1597]. Pastry [2230]. patch [2676]. patches [2872]. patents [963, 962]. path [2289, 358, 274, 2189, 751, 2400]. Paths [82, 2222, 364, 2763, 1783]. pathway [2823]. Pattern [2753, 1338, 197, 1223, 569, 2254, 2046, 2266, 2320, 807, 176, 1503, 436, 696, 2659]. Pattern-based [2753, 2266]. pattern-matching [176]. patterns [1654, 2519, 1640, 599, 1672, 2666, 189, 1565, 1579, 1556, 1166, 1665, 2644, 2016, 2505, 1336, 210]. Paul [575]. Paulson [181]. PBRD [1232]. PCC [712]. pCSSL [2286]. PCTL [1921]. pedagogy [1051]. peephole [1080]. peer [1797, 772, 2715]. Peer-Model [2715]. peer-to-peer [1797]. Penrose [269]. Pentagons [1208]. Penttonen [391]. PEPM [1259]. PEPM’15 [2118]. perceptible [1291]. perceptions [2873]. Peregrinus [156]. Pereira [249]. PERF [2490]. Performance [572, 821, 2070, 889, 1404, 2459, 1471, 2745, 2742, 1648, 1490, 1362, 2843, 907, 1536, 639, 1840, 892, 2116, 2624, 1291, 2073, 1650, 1433, 2342, 2132, 818, 1727, 2023, 2112, 170, 1799]. Performance-related [889]. performances [2570]. Periodic [2199, 2295]. Perisher [21]. permutation [437, 177]. permutation-invariant [437]. persistence [2169]. persistent [2116, 2480]. personal [1615, 1391]. personalised [1618]. personalized [2808]. personas [2185]. perspective [1144, 1662, 2453, 1852, 2202, 316]. perspectives [912, 1923]. pervasive [1412, 1920]. Peter [213, 246]. Peterson [470]. Petersson [298]. Petri [572, 1918, 1952, 2181, 2222, 2079, 2223, 2579, 2379, 2224, 1517, 2220, 607, 2219, 374, 1324, 804]. Petri-net-based [374]. PetriDotNet [2221]. Pharo [2459, 2249, 1744, 2251]. Phase [860]. Phase-based [860]. phases [1447]. Philadelphia [232]. PHP [2100, 2320]. Physical [2286, 1520, 2258, 1888]. PI3K [2823]. PI3K/mTOR [2823]. Pica [2104]. picto [2833]. PiDuce [1150]. Pieprzyk [215]. Pigs [738]. pipeline [387]. pipelines [1703]. Pitfalls [613, 2004]. Pitman [206]. PiXL [974]. Placement [57]. plain [545]. plain-old [545]. plan [552, 553]. planar [1523]. plane [1000]. Planning [217, 2152, 2455, 2781, 751, 611, 2308]. platform [1663, 2328, 1772, 2521, 2619, 2014, 2733, 1101, 2810, 2738, 1698]. platform-agnostic [2733]. platforms [1905, 2073, 1548]. platooning [2161]. player [280]. Playing [210]. PLC [2655]. plugin [2809, 2377]. PLUSS [188]. Pnueli [1172]. Point [441, 512, 186, 2135, 599, 1753, 2763, 763, 1953]. Point-free [441]. point-to-point [599]. pointcuts [2175, 941]. pointer [1216, 517, 841, 1792, 579, 2191, 348, 537]. pointers [211]. pointless [1953]. points [512, 2561, 2843, 765, 369, 376]. Poirters [265]. Poland [2886, 272]. polarities [2331]. policies [2832, 1471, 2307, 2583, 2084, 864, 1627, 930, 1372, 1017]. policy [1100]. polychronous [2765, 1758]. Polychrony [1907]. PolyGloT [2808]. Polyhedra [1077, 839]. polyhedral [2086]. polykinded [630]. Polymorphic [115, 2429, 629, 334, 2499, 853, 2782, 502, 1719, 2602]. polymorphically [1221]. polymorphism [2022, 580, 942]. Polynomial [128, 950, 2471, 1683, 951, 387, 1842, 949]. polynomial-time [951, 1842]. polystores [2834, 2792]. Polytypic [630, 626]. PomCho [2546]. POP [250]. POP-11 [250]. population [1710]. populations [2220]. portability [1664]. portable [2247, 831]. portal [750]. portfolio [923, 650]. portfolios [1063]. porting [1089]. ports [1656]. POSIX [1116]. possess [630]. post [1597]. postcondition [343]. postconditions [2471]. potential [1121, 1890, 852]. Power [305, 352, 2340, 1971, 304, 603, 2279]. powerlists [421]. pp [516, 168, 730]. pp. [34, 78]. PPDP’15 [2149]. PPPJ [1058, 1118]. Practical [1941, 575, 1946, 1958, 1361, 1561, 1475, 2274, 253, 2835, 1237, 2683, 2525]. Practice [2149, 1286, 1584, 744, 1925, 2204, 2567, 2498, 1513, 2429, 1553, 2267, 1873, 1999, 617, 1166, 1404, 2199]. practices [820, 1629, 1876, 1878]. pragmatic [690]. Pre [2222, 2020]. pre-orders [2020]. Pre-synthesis [2222]. Precise [839, 2092]. Precisely [230]. precision [2086, 1848]. precompiler [1083]. precondition [1459]. preconditions [2471, 1633, 363]. Predicate [393, 471, 359, 601, 1281, 375, 390]. predicates [2001, 1363, 2165]. Predicting [2561, 2573, 1299]. prediction [2631, 2742, 1648, 535, 2263, 2746, 2597, 2570, 1812]. Predictive [1458, 2102]. Preece [284].

Preface [2149, 1428, 1204, 1436, 1800, 2130, 2392, 2839, 2852, 1316, 1578, 1886, 1964, 1242, 2730, 628, 546, 1567, 2256, 1607, 2024, 2143, 894, 1420, 1003, 1209, 1231, 1269, 2310, 1863, 1914, 2697, 1410, 652, 1385, 1236, 2875, 707, 2866, 1854, 2535, 634, 1919, 2543, 1641, 2276, 1682, 2156, 1960, 2608, 899, 1030, 1904, 1666, 2038, 2562, 2374, 2452, 2838, 1160, 2688, 1584, 977, 1735, 2031, 2178, 2682, 1996, 2580, 1191, 2800, 905, 577, 1600, 1487, 1329, 1099, 2572, 416, 1543, 2847, 2563, 2272, 1892, 1692, 1364, 1349, 1550, 1925, 1440, 1941, 1612, 594, 2106, 1134, 2204, 1480, 1767, 1293, 1346]. Preface [2426, 783, 1009, 2851, 676, 1112]. preferences [2211]. Prentice [195, 141, 126, 202, 183, 265, 155, 192, 514, 34, 285, 286, 287, 219, 244, 78, 246, 229, 217, 297, 67, 194, 315, 215, 227, 185, 270, 201, 284]. Prentice-Hall [195, 141, 126, 202, 183, 155, 192, 34, 285, 286, 287, 219, 244, 246, 229, 217, 67, 194, 315, 215, 227, 185, 201, 284]. Prentice-Hall/Open [284]. preorder [1622]. preorders [2306]. prepare [1537]. preprocessor [842, 1128]. presence [2415, 2788, 1445, 523, 1494, 578, 2035]. Presentation [14]. presentations [747]. presented [2685]. preservation [2646, 1705]. Preserve [98]. Preserving [1402, 1328, 1956, 592, 2150, 2318, 508, 2493]. Press [48, 49, 516, 515, 243, 510, 295, 228, 391, 146, 521, 213, 289, 182, 321, 389, 388, 298, 398, 320, 181, 214, 167, 266, 269, 397, 169, 574, 390, 173, 144, 575, 318, 209, 205, 170, 196, 152, 191, 249, 316, 524]. pretty [1080]. pretty-printing [1080]. prevalence [2480]. prevent [1194]. Preventing [1192, 1699]. prevention [1365]. Price [141, 184, 291, 202, 183, 265, 155, 231, 243, 245, 154, 299, 295, 156, 151, 228, 206, 146, 319, 203, 213, 289, 285, 286, 287, 219, 244, 207, 246, 182, 229, 218, 153, 321, 172, 296, 217, 389, 208, 297, 180, 168, 298, 204, 171, 267, 216, 179, 320, 181, 317, 143, 145, 214, 167, 266, 269, 271, 169, 215, 173, 227, 144, 185, 226, 318, 270, 209, 290, 201, 288, 166, 205, 170, 284, 230, 152, 247, 248, 249, 250, 142, 195, 192, 193, 194, 575, 196, 191]. PRIMA [639]. PRIMA-UML [639]. Prime [122, 2222, 2189]. Prime-Number [122]. primer [216]. Primes [31]. primitive [407, 497]. primitives [1302]. Prince [268]. Principal [381]. Principled [1778, 816]. Principles [1286, 59, 748, 432, 2149, 1584, 1925, 2204, 141, 226]. printing [1080]. priorities [1324]. prioritization [1383, 2659]. prioritized [1324]. prioritized-time [1324]. Prioritizing [2459]. priority [2551]. privacy [2436, 2185, 1627, 1372, 1938]. Probabilistic [870, 451, 2581, 2414, 2583, 2795, 2699, 1787, 2660, 1105]. probability [1595]. Problem [478, 8, 102, 23, 57, 1882, 2447, 274, 89, 399, 1146, 158, 497, 737, 2176]. problems [221, 2152, 865, 2683, 354, 1382, 310]. procedural [1822, 2035]. procedure [1985, 458, 110]. Procedures [964, 162, 509, 134, 319, 1015, 607, 1867, 1334, 2176]. Proceedings [2883, 318, 2884, 2880]. Process [1688, 2056, 730, 781, 1068, 2862, 927, 1273, 1065, 2039, 466, 457, 2285, 2397, 1710, 926, 2108, 1147, 1322, 991, 2765, 970, 608, 1594, 2316, 1537, 1988, 2008, 1638, 1444, 1190, 2644, 2112, 288, 1181, 2427, 2304, 1537, 374]. Process-aware [1688, 2056]. process-theoretic [1988]. Processes [70, 33, 5, 75, 254, 90, 1004, 355, 929, 1524, 1108, 809, 653, 1985, 647, 844, 1203, 1345, 447, 2048, 667, 971, 972, 2839, 126, 205]. Processing [78, 1406, 2510, 1690, 1902, 218, 2661, 1502, 2014, 2344, 1251, 2772, 247, 248, 250]. processor [446, 279, 1402]. processors [859, 1195, 444, 2279]. procurement [802]. Produce [262]. Product [1516, 2334, 2800, 1515, 2719, 2020, 1801, 1509, 2190, 1564, 569, 1513, 1681, 1750, 1514, 765, 1443, 2284, 2273, 1785, 371, 994, 1540, 767, 1580, 1608, 2644, 1512, 2311, 763, 766, 762, 820]. product-line [2190, 1564]. productive [16]. productivity [2692, 1727, 207]. professional [2045]. Professor [442]. Profiles [635]. profiling [1291, 1535, 1060, 1764]. Program [46, 1716, 2118, 417, 98, 186, 261, 1954, 1717, 1495, 1593, 1312, 1499, 2795, 138, 258, 2837, 30, 102, 552, 1259, 23, 2458, 13, 74, 480, 553, 92, 532, 9, 1220, 570, 921, 375, 1081, 1161, 2729, 907, 742, 2537, 2153, 599, 2632, 2872, 2713, 1019, 449, 966, 431, 507, 151, 347, 1457, 411, 964, 312, 842, 1256, 1257, 1753, 1819, 158, 1903, 1686, 194, 2885, 1261, 1922, 468, 779, 187, 133, 327, 1931, 1450, 1540, 307, 2400, 1727, 420, 1060, 1764, 561, 144, 185, 549, 277, 2270, 234, 1096, 1290, 1281, 529, 739]. programmer [1130, 287, 2873]. programmers [745, 970, 1061, 549]. Programming [2149, 1546, 2462, 1204, 1436, 1800, 2130, 2881, 6, 443, 1677, 2685, 692, 2338, 471, 231, 935, 2879, 19, 1254, 1292, 361, 1286, 1197, 2880, 105, 1746, 2475, 130, 2882, 1482, 1795, 1191, 14, 15, 352, 2002, 40, 43, 47, 54, 66, 72, 77, 86, 91, 95, 101, 103, 107, 112, 117, 125, 131, 160, 226, 318, 781, 341, 648, 943, 408, 519, 9, 20, 21, 1319, 425, 1828, 241, 263, 334, 2152, 137, 2512, 901, 293, 329, 1455, 1409, 2731, 746, 89, 1426, 734, 1248, 2808, 1420, 1255, 2211, 1074, 569, 1751, 1011, 677, 1055, 514, 304, 1681]. programming [2128, 2003, 282, 939, 797, 728, 1318, 486, 1827, 1408, 2257, 253, 1139, 1594, 2708, 1123, 1822, 333, 1510, 280, 153, 607, 1054, 296, 1237, 208, 275, 309, 2287, 2056, 136, 1768, 527, 317, 373, 187, 2568, 2358, 2014, 1548, 1931, 61, 147, 178, 190, 242, 1234, 1821, 1608, 1452, 1258, 1560, 290, 1788, 475, 1028, 301, 288, 1703, 386, 1790, 2884, 997, 2083, 1584, 1215, 1360, 1400, 1416, 1448, 2004, 1925, 2693, 2204, 2676, 2470, 202, 389, 204, 214, 169, 270, 141, 155, 192, 297, 298, 67, 315, 227].

Programs

[51, 140, 240, 25, 53, 159, 109, 59, 150, 68, 28, 353, 342, 85, 1137, 1563, 947, 2119, 1866, 2206, 568, 2001, 1729, 443, 1121, 533, 474, 1924, 1683, 2093, 2648, 933, 950, 703, 1262, 2125, 735, 2492, 1475, 71, 1067, 1327, 1314, 854, 2584, 2109, 526, 1792, 2795, 360, 2857, 626, 562, 1565, 326, 1482, 2382, 843, 1153, 1460, 384, 564, 1124, 2655, 603, 1899, 1020, 2386, 848, 76, 844, 2187, 1140, 860, 2699, 2661, 127, 434, 563, 1317, 469, 1851, 698, 2451, 813, 1576, 2042, 374, 1451, 913, 2782, 381, 2035, 2271, 392, 549, 64, 2504, 1783]. programs [537, 806, 582, 2380, 800, 2144, 2592, 489, 2314, 1477, 1335, 154]. progress [1469]. Progression [31]. Project [1177, 448, 2692, 2877, 1150, 2195, 1645, 2746]. Projected [2270]. Projecting [2125]. projection [648]. projects [2359, 2559, 2214, 2830, 2373]. PROLOG [236, 505, 401, 2103, 356, 1000, 827, 229, 248, 514, 213, 249]. PROMELA [1321, 1226]. proneness [1880, 2320]. Proof [32, 53, 109, 2229, 26, 59, 2028, 2230, 568, 459, 1805, 2666, 1662, 966, 134, 831, 2723, 2750, 709, 319, 366, 1469, 2260, 127, 181, 714, 314, 2845]. proof-carrying [831]. Proofs [1138, 429, 1051, 1366, 1461, 1394, 2435, 2032, 1097, 1105]. propagating [1831, 1812]. Propagation [90, 2289, 368, 454, 1926, 1334]. Properties [59, 927, 598, 2119, 1376, 1897, 2212, 1895, 1363, 710, 1328, 2019, 1022, 1194, 1571, 1766, 1843, 1225, 1945, 2157, 592, 307, 2420, 395, 846, 2465, 582, 2505, 2728, 1240]. Property [1757, 1285, 1890, 1889, 2579, 2804, 791]. property-based [1889]. Property-dependent [1757]. property-directed [1890]. proposal [943, 2849, 1605]. Prospects [1794]. proto [1418]. protocol [838, 1994, 542, 1354, 1431, 709, 862, 2278, 625, 1756]. protocols [465, 478, 462, 974, 711, 62, 2207, 1569, 1433, 708, 685, 148]. prototype [2358, 1540]. prototype-based [2358]. prototypes [2101]. Prototyping [1516, 130, 244, 2844, 1745]. Provably [2544, 1365, 690, 2032]. prove [722]. provenance [2209, 1782]. provide [2877, 1494, 666]. Providing [1189]. Proving [262, 2420, 395, 2864, 1110, 1085, 1701, 1401, 1754, 1930, 1238]. provisioning [1342, 1388]. proxies [1468]. proxy [1832]. pseudonym [1102]. Ptolemy [1375]. PTRebeca [2040]. public [1630]. Publications [573]. Publisher [2184, 2183]. Publishers [226, 142]. Publishing [2183, 206, 909]. pump [2213, 2791]. pure [505, 2333, 1481]. purely [444, 410, 521]. purpose [1832, 1742]. pushdown [845]. putting [16]. PVS [420]. Python [2753, 2641, 2128, 2620, 2654, 2593].

QaSD [2541]. QMaxUSE [2767]. QoC [1489]. QoC-aware [1489]. QoS [1214, 1850, 1544, 1705]. QSIC [1809]. quadratic [328, 1881]. Qualification [2750]. qualified [367]. qualitative [2211, 1750, 1132]. Quality [886, 2776, 2541, 1809, 2642, 1200, 1157, 1325, 2099, 2109, 2746, 153, 2518, 1014, 2202, 2807, 153]. Quality-aware [2541]. quantification [425]. quantified [1347]. quantifier [1890]. Quantifying [1157, 1325, 1084, 1156, 1063, 2547, 811, 1021, 2654, 1637, 1650]. Quantitative [650, 810, 1669, 1750, 1949]. quantum [2683, 1477, 2341]. quasi [2361, 2352]. quasi-action [2352]. quasi-equivalence [2361]. Queen [57]. queries [1262, 1196, 1818, 1788]. query [1069, 1690, 1902, 2445]. query-by-example [1069]. QueryAGT [2731]. querying [2488, 1742]. Question [2405]. questionnaires [1779]. Queue [1264]. queues [165]. QuPreSS [2102].

R [126, 202, 231, 34, 219, 208, 194, 398, 266, 226, 142]. Raatz [319]. race [2482]. races [953]. radioactive [917]. RAIDE [2819]. Rail [2200]. railway [2090, 2039, 1662, 543, 1755, 2848, 2290, 463]. Rajasekar [389]. random [2749, 2408, 2492, 2041, 2639, 2651, 1358]. Randomised [2000]. randomized [1138]. range [2191]. rank [1129]. Rankers [354]. Ranking [2568, 2390, 1690]. rapamycin [2823]. Rapid [130, 2652]. Rascal [1975, 2100]. rate [2485]. rates [1789]. rational [512, 189]. rationale [1310]. Rawlins [396]. Ray [316]. Rayward [142]. Rayward-Smith [142]. RBAC [2063]. RBF [2597]. RBF-NN [2597]. rbFeatures [1318]. RC [2178, 2682]. rd [497]. RDF [1715]. Re [2283, 1852]. Re-engineering [2283, 1852]. Reachability [1523, 2642, 1951, 2878]. Reactive [2114, 943, 2534, 2170, 337, 560]. read [1471, 1621]. read-write [1621]. readers [1238]. Reading [1129, 320, 166, 1986]. readings [284]. Ready [2482, 399]. Real [2200, 1559, 339, 512, 973, 662, 1654, 2407, 589, 2212, 1797, 2236, 2843, 340, 2552, 1497, 1596, 1935, 2579, 1314, 1092, 709, 543, 631, 964, 289, 615, 1859, 2362, 1824, 1936, 1001, 266, 1623, 395, 2279, 729, 2441, 2505, 1375, 1965]. Real-Time [2200, 1559, 339, 973, 662, 1654, 589, 2212, 1797, 2236, 340, 2552, 1497, 1596, 1935, 2579, 1314, 1092, 709, 543, 631, 964, 289, 615, 1859, 2362, 1824, 1936, 1001, 1623, 395, 2279, 2441, 2505, 1375, 1965]. reality [2813]. Realizability [2240, 2224]. realization [642]. realtime [2076]. reason [1544]. Reasoned [1394]. Reasoning [2297, 252, 1022, 324, 631, 1571, 1923, 2845, 2317, 2803, 2327, 734, 393, 471, 1658, 944, 1598, 1524, 1596, 1284, 2876, 2160, 830, 1182, 1158, 523, 199, 1534, 2454, 1673, 1331, 243]. Rebeca [1825, 1623, 1965, 1210]. recognisers [1164]. Recognition [1164, 863, 437, 2581, 1824, 772, 553]. recognizable [1895]. Recognizer [45]. recommendation [2598, 2540, 2744, 2518]. recommendations [2877, 1785, 1618]. recommender [2865, 2865]. reconciliation [1961]. Reconciling [1107, 636, 1446, 1727, 1038]. Reconfigurable [2396, 1971, 1745, 2071]. reconfiguration [1994, 1381, 2554, 2273, 1705, 641]. reconfigurations [1516, 1233, 1923]. reconsidered [163]. reconstructing [568]. record [2098, 601]. records [1409]. recover [1864]. Recovering [1565, 1463]. Recovery [135, 596, 598, 548, 879, 1539, 2491, 2431, 1336]. Recurrences [1]. Recursion [17, 324, 1033, 1747]. Recursive [303, 294, 2879, 129, 1866, 313, 1425, 687, 1944, 2345, 2144]. redocumentation [1176]. reduce [83, 2805]. Reducing [2249, 1845, 1580, 2205]. reduction [2366, 1282, 2079, 2297, 482, 686, 200, 2361, 2639, 1033, 2194, 2172, 2789, 1210, 1239, 2536]. reductions [1757, 2874]. Reductivity [430, 1945]. Redundancy [1780, 690]. Reengineering [2775, 875, 878, 653, 1443, 738]. reentrant [2382, 1238]. Refactoring [1640, 1503, 2485, 1880, 2481, 1591, 1443, 1680, 2819, 1174, 1740, 2373, 800]. refactorings [1167, 733, 2698]. reference [2645, 2753, 2861, 1027, 2102, 2489]. references [1831, 2502]. Refined [176]. Refinement [512, 1113, 1697, 1370, 1372, 1675, 432, 1810, 2163, 2119, 233, 1695, 2213, 438, 509, 1805, 2159, 1083, 1935, 2666, 2059, 2098, 2062, 255, 1393, 439, 964, 1275, 1435, 362, 1380, 1933, 2655, 1351, 2386, 259, 136, 468, 1278, 592, 1371, 2369, 1277, 1323, 1276, 447, 1280, 522, 2513, 1141, 2355, 2529, 2441, 2505, 2728, 721, 1274, 1499, 1140]. refinement-based [1083]. Refinement-oriented [1370]. refinements [2138, 2318, 480]. refining [1576]. Reflecting [2848]. reflection [1152]. reflective [1257, 1744, 2614, 2709, 2709]. reflectogram [2131]. Refunctionalization [1136]. regexes [1678]. Region [1340]. Region-Based [1340]. regions [1557]. Regis [218]. Register [714, 833, 330]. registers [332]. regression [2459, 2640, 2468]. regular [2469, 2019, 276, 672, 1928, 2773, 1811, 2710, 696]. regularities [1175, 1900]. regularity [400, 1644]. regulated [1422]. Regulations [1339]. reification [1094, 1926]. Reifying [2131]. reinforcement [2583]. Reinforcing [2008]. related [2243, 889, 2465]. relating [867]. Relation [513, 13, 1685, 1634]. Relational [812, 361, 1899, 48, 2862, 2119, 568, 1230, 950, 431, 2675, 1208, 1843]. relations [725, 293, 513, 683, 1456, 327, 1676, 948, 654, 2660]. relationship [921, 857]. relationships [2679]. release [1348, 2089]. releases [2600]. relevant [2571, 2270]. Reliability [1812, 2190, 1840]. reliable [2076]. relying [2462]. remarks [335]. Remembering [1172]. Remodularizing [1317]. Remote [1101]. remotely [1431]. Removal [17, 1290]. removals [314]. renaming [403, 481]. Rennes [2882]. renovation [656, 550]. Reo [895, 2242, 1233, 1271, 1544, 1353, 1232]. Rep [9, 35]. Repair [135, 2642, 2872]. Repeated [24]. repeating [2710]. repetitions [631]. replaced [2873]. replacement [1437, 1233]. replicated [2299]. replication [2712, 2158, 2528]. Report [1025, 87, 1295, 1975, 785, 299, 544, 2849, 2830]. reporting [2074, 2431]. repositories [1649, 2015, 1441]. Repository [1806, 965, 1116, 1114]. representation [1643, 979, 683, 317, 835, 2356, 2016, 2009]. representation-changers [979]. Representations [239, 1715, 2070, 713, 739]. representing [2211, 2408, 618]. reproducibility [758]. repurposing [2823]. request [1390]. request-response [1390]. required [1962]. requirement [2185, 2501]. requirements [2679, 1939, 2303, 2745, 1227, 1723, 2534, 336, 2560, 1506, 1573, 544, 2233, 2365, 452, 1084, 1302, 708, 2287, 1604, 187, 1940, 2480, 2830, 852, 993, 2385, 1382, 1488]. Research [289, 1404, 1536, 2100, 2768, 2521, 2217, 1540, 2101]. researcher [1268]. researchers [1760]. Reshape [2504]. Reshuffling [8]. resilient [2160]. resistant [2300]. Resolution [306, 2389, 1217, 2756]. resolving [632]. Resource [1941, 2732, 1442, 2249, 1228, 2397, 2413, 1414, 1446, 2604, 2150, 2261]. resource-based [1442]. resource-constrained [1446]. resource-efficient [1228]. resources [2104, 2084, 2214, 2049]. respect [2295]. response [1390]. responsiveness [2124, 916]. REST4CEP [2510]. RESTful [2510]. Restriction [1829]. Restructuring [657, 599]. result [2179]. results [1789, 840, 720, 1181]. retractable [2296]. retrenchment [1023]. retrieval [2049, 2012]. return [2676, 2676]. Return-Oriented [2676]. REUBI [1488]. Reuse [2597, 2280, 439, 843, 2827, 1738]. Reusing [1847]. reveal [2591]. reversal [2668]. Reverse [547, 1295, 1598, 1711, 1176, 1044, 654, 1346, 2806]. reverse-mode [2806]. reversibility [2181, 2180]. Reversible [2181, 2634, 2178, 2682, 2628, 2627, 2657, 2683, 2179]. Reversing [2708, 370]. REVEUR [110]. REVEUR-3 [110]. Review [48, 195, 141, 49, 126, 184, 291, 516, 202, 375, 183, 515, 265, 155, 192, 231, 243, 514, 245, 79, 510, 154, 299, 295, 156, 151, 34, 228, 391, 206, 146, 319, 203, 521, 213, 289, 285, 286, 287, 219, 244, 78, 207, 246, 182, 229, 218, 153, 321, 172, 296, 217, 389, 208, 297, 193, 388, 180, 168, 298, 67, 204, 171, 267, 268, 194, 398, 216, 315, 179, 320, 181, 317, 143, 145, 214, 167, 266, 269, 271, 397, 169, 390, 215, 173, 227, 144, 185, 226, 318, 270, 209, 290, 201, 288, 166, 205]. Review [170, 196, 284, 230, 152, 191, 247, 248, 249, 250, 142, 316, 524, 396, 538, 2561, 2598, 499, 781]. reviewer [2598]. revised [2685]. Revising [1701]. revisions [1129]. Revisited [2311, 1835, 470, 1054]. Revisiting [2673, 879, 2494]. rewrite [1859, 2413, 1803]. Rewriting [83, 128, 262, 2208, 36, 813, 1570, 2414, 1860, 2366, 1771, 2302, 2397, 453, 1956, 314, 1676, 2345, 1764, 1913, 2739, 1854]. RFI [2849]. RFP [2849]. RFP/RFI [2849]. rich [1435, 1278]. Richard [141, 291, 245, 226]. Richardson [568]. right [1223, 264]. right-to-left [1223]. right-top [264]. Rigi [1176]. rigorous [2071, 2665]. Ring [75, 464]. Rio [1432]. risk [965, 1922, 1063, 1103]. risk-aware [1922]. risk-bearing [1063]. risks [1156]. RML [2567]. Road [21, 2797]. roadmap [2257]. roads [2190, 333]. Robert [79, 290]. Robin [574]. Robot [2083, 2680, 2048]. robot-based [2048]. robotic [2667]. robotics [2524, 2347]. robots [751, 2128]. Robust [1837, 817]. Robustness [1895, 1213]. Rockville [49]. Rodin [1696, 2233, 1698]. Roger [146, 269]. role [1770, 959, 1815]. role-based [959]. Roles [757, 1714, 773, 864]. Rome [2190]. root [2790]. Rotations [1000]. routine [1493]. routines [1811]. routing [586, 1998]. Row [180]. RPC [2499]. RRPlib [2408]. RT [973, 2579]. RT-MOBS [2579]. RTSJ [1340]. Ruby [1318, 362, 2593]. rugby [2447]. Ruggles [299]. Rule [1356, 1797, 1037, 592, 2648, 387, 1855]. Rule-based [1797, 1037, 592, 2648, 1855]. Rules [109, 32, 1490, 709, 1469, 1986, 1347, 2345, 1451, 1803, 1021]. run [1248, 1516, 1446, 1864, 2205, 880]. run-time [1248, 1516, 1446, 1864, 2205, 880]. running [1095, 187, 2406]. Runtime [2392, 2019, 1551, 2567, 773, 2254, 1131, 2720, 1798, 765, 2017, 789, 1152, 1304, 1206, 1491, 1247, 736, 1335].

S [375, 289, 249, 538, 531, 555]. SAC [1687, 2075]. SAC-SVT’15 [2075]. Safe [858, 1823, 1468, 868, 2481, 2017, 914, 2583, 1801, 2554, 2059, 1019, 1196, 740, 1256, 1287, 843, 671, 1344, 1251, 2504, 1407, 1026]. SafeGen [1256]. safer [1680]. Safety [2088, 2197, 2354, 2603, 725, 1841, 2386, 463, 2667, 2669, 1889, 786, 2329, 2511, 2811, 2085, 1022, 1085, 1557, 2680, 2330, 2353, 543, 1553, 2215, 1558, 2521, 1843, 2655, 1250, 592, 1401, 1652, 846, 2529, 2402]. Safety-Critical [2088, 2197, 2354, 2603, 2386, 786, 2329, 543, 1558]. Safety-level [463]. Salespoint [1538]. SALSA [1309]. SAMOS [2711]. sample [2408]. sampling [2749, 1497, 1447, 2295]. SANER [2775]. Santander [1645]. SApHESIA [2526]. SAT [1350, 1371, 2766]. SAT-based [1350]. SAT-solving [1371]. Satan [708]. SATDBailiff [2633]. SATDBailiff-mining [2633]. satellite [1575, 1435]. satisfaction [1272, 2683]. satisfiability [2277, 2198]. Saumya [318]. Sauna [1858]. Saunders [291]. sausages [738]. SBESC [2276]. SBLP [1546, 1677, 2685, 2068]. SBLP’14 [2002]. SBLP’15 [2338]. SBMF [2799, 2697, 2457, 2383, 2535, 1909, 1329, 1670, 2024]. SBST’22 [2852]. SC22 [402]. scalability [1536, 2116]. scalable [2463, 1424, 1733]. Scale [1997, 1419, 1541, 1625, 2062, 1198, 2220, 2570, 2344, 2373]. scaling [1971, 308]. SCAM [1896]. scanning [1223]. scenario [2656, 2170, 1588, 803, 1850, 1271]. scenario-based [2170, 803, 1271]. scenarios [1828]. ScenarioTools [2170]. scene [1711]. schedulability [1971, 2036, 1825, 2542]. schedules [2224]. Scheduling [1802, 1907, 1971, 2036, 615, 1345, 2199, 2441]. schema [1786]. schemas [137, 654]. Scheme [135, 1252]. schemes [1842, 2646, 381, 1938]. Schmidt [183]. Scholten [375]. school [21]. Sci [471, 19, 781, 648]. Science [48, 49, 514, 154, 1292, 172, 388, 168, 171, 216, 271, 21, 56, 232, 272, 1074, 1051, 398, 397, 2048, 997, 1074, 219, 180, 142]. ScienceDirect [2216]. Scientific [142, 2344]. scientist [1172]. Sciit [2586]. SCJ [2386]. SCJ-Circus [2386]. SCM [987]. SCOOP [2093]. Scoping [1229, 940]. scores [2723]. SCP [2294, 2024, 1112, 1666, 1584, 1925, 2204]. SCR [1723]. Screen [68]. Script [84]. ScriptEase [1011]. scripting [1361, 1011]. Sdf [1768]. SEA [1897]. SEA-based [1897]. SEAA’12 [1804]. Seamless [1738, 803]. Search [2852, 770, 907, 2798, 2878, 1733, 2797, 123, 2627, 454, 2336, 1589]. Search-Based [2852, 2798, 2797, 2627]. searches [1811]. Searching [257, 31, 210]. Seberry [215]. SeByte [1733]. Second [407, 1076]. Second-order [407]. Section [1896, 1607, 1504, 1428, 1436, 1800, 2130, 2052, 1420, 1410, 1919, 1989, 1487, 1225, 1416, 1543, 1722, 1550, 1448, 1440, 1480, 1397]. Secure [926, 1660, 1104, 711, 559, 2754, 993, 667]. Security [710, 1719, 1366, 927, 2015, 2424, 864, 2578, 892, 995, 2185, 925, 711, 2577, 1796, 1140, 930, 215]. seeking [1639]. segment [310]. segments [437, 264]. Seldin [146]. select [2571]. Selected [1716, 2694, 2118, 1659, 2068, 1274, 2799, 2338, 2457, 1954, 1687, 1341, 1392, 1909, 1593, 1254, 1726, 1746, 2367, 2882, 1279, 1329, 2687, 2665, 2018, 1653, 2002, 1692, 1550, 1670, 2693, 284, 2727, 2851, 573, 2149, 1355, 1518, 1702, 2024, 2143, 1025, 2283, 2239, 2391, 1495, 1854, 1919, 669, 1960, 1948, 2582, 2737, 2688, 1543, 2686, 1259, 1349, 1346, 2697, 2885, 1809]. selecting [598]. selection [2631, 2645, 2747, 2624, 1689, 2027, 2384, 2570, 1206]. selective [859, 780]. Self [1415, 2310, 1410, 802, 2452, 2563, 817, 2633, 1430, 2747, 1857, 2034, 2297, 2236, 1379, 1921, 960, 1879, 1412, 1919]. self-adaptation [2297]. Self-Adaptive [2452, 2310, 2034, 2236, 1379, 1879]. Self-Adaptiveness [2563]. self-admitted [2633]. self-assembly [1857]. Self-certification [802]. self-composition [1430, 1412]. self-driving [2747]. Self-healing [817]. self-managing [960]. Self-organisation [1919]. self-organising [1921]. Self-organized [1415]. Self-Organizing [1410]. Semantic [222, 1398, 1904, 1413, 1708, 2877, 1775, 1714, 778, 2463, 2507, 559, 300, 2827, 1569, 665, 2009, 1713, 2015, 889]. Semantics [356, 2153, 940, 1856, 412, 111, 1768, 1915, 42, 505, 293, 329, 443, 897, 474, 638, 1297, 1564, 950, 1283, 786, 482, 677, 1368, 1265, 691, 1107, 449, 2134, 2299, 1108, 2579, 2224, 584, 1679, 607, 1956, 581, 363, 2836, 281, 601, 1018, 2145, 169, 1402, 1965, 445, 1705, 1560, 1587, 1418, 1673, 191, 2613, 1122, 2505, 2341, 236, 1913, 2450, 388, 375, 183, 182, 216, 390, 171]. Semantics-based [2153, 2450]. Semantics-directed [356]. semantics-preserving [1956]. semaphores [395]. Semi [1651, 2116, 2819, 1017]. Semi-automated [1651, 2819]. semi-automatic [2116]. semi-formal [1017]. semilattices [536]. semirings [1244]. sense [2107, 1528]. Sense/Compute/Control [1528]. Sensemaking [1871]. sensing [2297]. sensitive [2732, 2446, 1757, 1728, 2874, 2140]. sensor [1342, 1388, 1545, 1569, 1414, 1999, 1491, 805]. sensor-actor [1545]. Separate [1314, 1417, 162]. separation [1363]. September [35, 272, 2393, 2477, 2486, 2605, 2700]. Sequence [646, 778, 2669, 728, 1103, 2271, 2536]. SequenceL [620]. Sequences [807, 83, 1486, 1460, 1707]. Sequential [126, 1367, 1694, 1924, 344, 1599, 360, 76, 1140, 2709, 2089, 2144]. sequentialization [225]. serial [278, 1122]. serial-parallel [278]. Serialisable [2061]. serializable [722]. Series [572, 514, 152, 573]. serious [2581]. servanda [2726]. server [1554, 909, 1059]. servers [1362]. Service [1796, 2618, 498, 2303, 1144, 2423, 2115, 1704, 1797, 1615, 1069, 1337, 1342, 1388, 1194, 1983, 1562, 586, 2773, 754, 1307, 1219, 1920, 2032, 1703, 2498, 1795, 1219]. Service-Based [498, 2115, 1615]. Service-oriented [2618, 2303, 1337, 1983, 1562, 1795]. Services [491, 1324, 2424, 1991, 1150, 2456, 1619, 2102, 877, 1387, 2344, 1412, 1326, 1640]. Servilla [1342]. serving [1447]. session [1987, 2331, 2207, 2331]. session-based [2331]. Session-ocaml [2331]. sessions [2560, 2376, 1386, 1703]. Set [361, 532, 1077, 2223, 2408, 2482, 2639, 1109, 2805, 1676, 2140, 2430]. sets [2628, 2356, 2601]. SETTA [2580]. setting [1389]. Several [2632]. severity [2480]. SFr [319]. shadow [2080, 1140]. Shape [412, 1928, 1725, 1363]. Shapeliness [563]. shapes [562]. Shared [88, 87, 1865, 1294, 664, 926, 1691, 1881, 1433, 960, 1189]. shared-memory [1433]. sharing [2120, 1559, 1823, 2333, 445, 582]. Shieber [249]. shortcut [1449, 2072]. Shorter [364]. Showing [68]. shy [1262]. Sibilla [2876]. Sidcup [193]. Side [105, 2322, 1430, 851]. side-channel [1430]. Side-Effects [105, 851]. Sieber [151]. Siefkes [291]. Sieves [122]. signal [603, 293]. Signature [1962]. significance [1472]. Silver [1977, 1163]. Silverman [290]. Simi [317]. similarities [1486]. similarity [1455, 1733, 2547, 2140]. simpA [1234]. Simpl [1981]. Simple [1196, 1225, 58, 947, 709, 690, 133]. simplified [481]. simply [485]. simply-typed [485]. Simulation [2262, 1275, 2220, 1423, 2747, 399, 2347, 1214, 2298, 1623, 2094, 2385, 2660, 2308]. Simulation-based [2262, 2747, 1214, 2385]. simulations [1230]. simulator [2714].

Simulink [1962]. simultaneous [1300]. Single [1901, 2289, 1561, 1691, 697]. single-path [2289]. single-user [1691]. Sintzoff [1268]. site [1535]. Situ [12, 19, 394]. Six [1877]. sixteenth [272]. size [1363, 2111, 2585, 2782]. size-limited [2585]. size-polymorphic [2782]. sizes [1721]. sizing [2112]. skeleton [1202]. skeletons [29]. sketches [1104, 937]. Skew [1405]. Skew-space [1405]. skills [2808]. Slater [156]. SLE [1759]. slices [434]. Slicing [2699, 1211, 1563, 947, 921, 1889, 595, 715, 1457, 1210, 1128, 1168]. Slimming [1290]. Small [40, 43, 47, 54, 61, 66, 72, 77, 86, 91, 95, 101, 103, 107, 112, 117, 125, 131, 147, 160, 178, 190, 242, 2213, 1631, 275, 1177]. smallest [325]. smalltalk [170, 1739, 1736, 1737, 1735, 2520, 1740]. smart [350, 1714, 785, 1490, 2824, 2846, 1017, 2859, 2308]. smart-cards [350]. SMArtOp [2173]. smell [2675, 2624]. smells [2819, 2783, 2390]. Smith [298, 142]. smoothers [2640]. Smoothsort [19, 12]. SMORE [2009]. SMT [2854, 2366, 2000, 1395, 1696]. SMT-solver [2000]. SMT-solvers [1395]. SMUIML [1602]. Snapshots [850]. snippets [2756]. Snodgrass [245]. SOAP [1795]. social [2590, 2050, 2009, 1629]. Society [146]. Socio [2403]. Socio-technical [2403]. SODA [1423]. Soft [2395, 1824, 168]. Software [2883, 2839, 2852, 2694, 1518, 588, 1659, 2256, 760, 1509, 1647, 2223, 1516, 2781, 2283, 2681, 1759, 1178, 2334, 2239, 2391, 2626, 1969, 1641, 2625, 1948, 2582, 2737, 286, 287, 2838, 1989, 1784, 2580, 2800, 2775, 2541, 385, 611, 2515, 1515, 1378, 1525, 2687, 1423, 1809, 1653, 1577, 1143, 2563, 2272, 2686, 2050, 1349, 993, 2044, 876, 2007, 1590, 805, 2426, 2727, 2851, 1043, 1076, 1173, 1814, 1780, 2097, 2524, 2571, 2631, 1663, 2015, 2640, 1817, 195, 56, 1316, 1355, 1294, 2234, 1295, 448, 1471, 2745, 2561, 2742, 2530, 1770, 1077, 875, 400, 1065, 2232, 2645]. software [2020, 492, 963, 541, 1907, 1801, 2747, 1175, 897, 1648, 962, 1227, 656, 1789, 604, 1090, 1489, 655, 2843, 2578, 815, 886, 989, 1200, 1306, 1339, 850, 959, 821, 1617, 1804, 1840, 2195, 1631, 995, 2600, 1298, 1870, 2774, 1526, 1681, 1662, 717, 2108, 1116, 1114, 797, 1750, 1575, 1514, 1880, 156, 1645, 1782, 1441, 2011, 748, 590, 1773, 1651, 2353, 1511, 1781, 544, 2100, 285, 1639, 1519, 1435, 1730, 2365, 1869, 1778, 1091, 2746, 1176, 153, 777, 2882, 2229, 2873, 1537, 1745, 1443, 1522, 1300, 1213, 1049, 896, 1310, 2812, 1507, 798, 1177, 1539, 2141, 822, 1302]. software [2273, 2559, 2008, 1638, 1543, 2348, 2667, 987, 2754, 1873, 1923, 2569, 1876, 2214, 1304, 1875, 1812, 1774, 794, 1871, 2217, 1540, 2480, 1508, 767, 1440, 817, 888, 1246, 1555, 2465, 1787, 2106, 2644, 891, 1762, 1872, 1874, 1566, 1530, 2468, 2311, 1635, 2373, 923, 2501, 1769, 986, 1713, 763, 641, 1877, 455, 2427, 2304, 1293, 736, 766, 2807, 1538, 804, 881, 2635, 550, 621, 2719, 1702, 2776, 2625, 985, 2031, 1600, 1487, 1722, 871, 2075, 1612, 2734, 78, 79, 206, 244, 207, 217, 143, 209]. software-defined [2427]. Software-intensive [2626, 1295, 2465]. software-like [1745]. Softwarenaut [1539]. SolAR [2824]. Solenoid [1519]. Solid* [1540]. solidity [2824]. solitaire [1496]. solitaire-like [1496]. solution [1361, 399, 1306]. solution-oriented [1306]. solutions [2763, 2279, 1382]. solve [2719, 34]. solver [2000, 534]. solvers [1395, 1696]. Solving [478, 1015, 497, 221, 2854, 328, 2152, 1371]. Some [16, 124, 720, 2205, 121, 612]. Sort [97, 1146]. sorted [1861, 2227, 2739]. Sorting [12, 1415, 19, 134, 273]. SoSs [2526]. Sound [2258, 1167, 1859, 2773, 1995, 502, 1703]. soundness [673]. Source [1659, 2784, 1788, 1849, 1663, 2359, 1541, 1661, 1296, 1486, 1617, 900, 2586, 1834, 1476, 1743, 1463, 2599, 1900, 627, 2488, 1317, 876, 2373, 920, 1896, 919]. Source-code [1788, 1900]. Sourcerer [1541]. Southport [206]. Space [1944, 1945, 23, 1465, 2366, 1216, 482, 2600, 1429, 686, 926, 1943, 1990, 1562, 2555, 666, 1239, 561, 1405, 625]. Spaces [493, 2163, 1146, 1369, 957, 2712, 609, 1007, 960, 522, 667, 1189]. Spain [20]. Spanders [1411]. Spanning [39, 1411]. SPARCv8 [2430]. Spark [2408]. sparse [1927, 2173]. sparse-matrix [2173]. SPARTA [1102]. spatial [2511]. Spec [1738]. Special [1504, 1896, 1204, 2392, 2734, 2852, 919, 2052, 2226, 2685, 2310, 1373, 2681, 2866, 1919, 939, 669, 1254, 1312, 932, 2625, 1286, 2038, 1730, 1572, 2688, 2702, 1989, 1795, 2682, 2580, 2775, 1752, 1215, 1360, 1400, 1416, 1722, 1809, 1653, 1035, 1143, 2563, 2686, 1259, 1821, 1349, 1550, 2075, 1448, 1934, 1246, 2044, 1708, 2007, 1397, 2058, 1088, 1058, 1118, 1428, 1436, 1800, 2130, 2839, 1316, 2730, 1642, 875, 2256, 2024, 1420, 1863, 1914, 1385, 1236, 2383, 2875, 1624, 1112, 2776, 1641, 1645, 2276, 2156, 1960, 1904, 1666, 2562, 956, 2374, 2838, 1584, 2219, 1735, 2031]. special [2178, 1996, 1191, 2800, 1600, 1487, 2572, 2847, 1966, 2480, 1925, 1440, 1941, 1612, 2106, 2204, 1480, 1293, 1346, 2426, 1173, 1814, 2867, 1607, 2143, 1647, 1410, 925, 2452, 1543, 2272, 1892]. specialisation [741, 2122]. specialization [1717, 2153, 2412]. specializer [742]. specific [1419, 2567, 2590, 908, 2848, 1611, 1679, 671, 1307, 1608, 1764, 2501]. Specification [1729, 2138, 464, 2275, 1766, 138, 22, 2386, 791, 2158, 283, 10, 1521, 973, 1100, 1570, 519, 1583, 164, 459, 188, 530, 2061, 1227, 2236, 670, 340, 1564, 1283, 2060, 2579, 1046, 1609, 182, 1977, 62, 453, 339, 615, 1237, 1997, 708, 826, 930, 143, 187, 2165, 397, 1555, 144, 2738, 1756, 1335, 228]. Specifications [526, 342, 2832, 1374, 188, 409, 2815, 1723, 378, 867, 2125, 1527, 446, 2259, 379, 2210, 16, 1651, 1573, 2735, 2576, 452, 2428, 2886, 2287, 1271, 2381, 2699, 1278, 1576, 271, 2480, 2309, 1151, 338, 593, 351, 1285, 2542, 2556]. specified [450, 603, 64]. Specify [2591, 2185, 722, 37]. Specifying [422, 1911, 590, 1558, 2053, 254, 322, 2168, 1198, 1237]. spectral [775]. Spectrum [6, 1905]. speculative [2296]. speech [863]. Spill [452]. spirit [166]. Spivey [514, 182]. SPLC [1509, 2800]. splitting [2350, 379, 1985]. spoken [774]. Sponar [21]. Spot [2789]. SPPF [1164]. SPPF-style [1164]. spreadsheets [1751]. Springer [184, 291, 375, 154, 299, 172, 168, 171, 216, 179, 271, 290, 730]. SQA4AI [2776]. SQL [1196, 1899]. SSCOP [542]. SSDTutor [2754]. ST [2584]. stability [1302]. stabilization [2244]. Stable [1080]. stack [175, 2518]. stack-based [175]. Stackability [485]. stage [1255, 910]. staged [908, 909]. STAMP [2330]. Stan [319]. Standard [27, 489, 638, 482, 878, 2063, 2405, 1032, 1660]. standardization [1630, 1664]. standards [299, 974]. Stanford [573]. starts [266]. starvation [1238]. State [1694, 2497, 482, 686, 1392, 1340, 96, 1460, 2665, 1692, 625, 2261, 2609, 2163, 2366, 2407, 2090, 1282, 323, 2079, 2000, 1369, 2060, 1561, 1429, 2401, 1425, 2323, 255, 544, 2804, 2267, 1278, 2158, 1239, 2842, 755, 2613, 1240]. state- [2613]. State-Based [2665, 544, 755]. State-cover [1460]. state-rich [1278]. state-space [1239]. State-taint [2261]. state/event [1282, 1240]. state/event-based [1240]. statechart [1107, 593, 971]. statecharts [590, 2284, 2318, 119]. Stateflow [1370]. stateful [1883, 1467, 2484]. Statement [343, 2638]. statements [596]. states [2271]. Static [662, 1725, 2729, 681, 2151, 1557, 733, 843, 373, 11, 42, 1119, 885, 2001, 1187, 2712, 2632, 1226, 2450, 2675, 2647, 857, 281, 1478, 1851, 2773, 1549, 1580, 1239, 948, 2205, 1740, 1168, 2676, 2722, 837]. Statically [1256, 2691]. station [603]. Statistical [2036, 2077, 153]. Steam [670]. Steam-Boiler [670]. Step [1359, 2503, 1368, 2266]. Stephen [266, 270]. Steps [420]. Stepwise [233, 438, 255, 1619, 136]. Steve [1391]. Steven [510]. stigmergy [2425]. Stipula [2726]. StMungo [2207]. Stochastic [572, 2302, 2823, 2286, 2224, 2362, 1544]. stock [2738]. stop [503]. storage [49, 712]. store [1116]. Stores [525, 2540]. Stories [1624, 2847]. storm [2232]. story [2829, 2757]. Strand [270]. strata [1447]. Strategic [2541, 1262, 295]. Strategies [518, 2213, 1857, 2190, 2774, 110, 2491, 1724, 1229, 2766, 1877, 737]. Stratego [1081, 2791]. Stratego/XT [1081]. Strategy [27, 118, 1295, 2721, 591, 1280, 2728]. Stratified [1447]. Stream [1502, 2325, 1902, 1188, 779]. streaming [2418, 819, 979]. streams [1943, 1741, 2077, 2481]. Strengthening [605]. stressing [1267]. strict [635, 2720]. Strictness [104, 488, 189]. Strider [755]. Strikes [120]. String [17, 2251, 1811]. Strings [1292, 770, 1222, 210]. strong [1226, 1469, 1867]. strongest [343, 2471]. Structural [1381, 129, 28, 1175, 461, 2175, 1153, 1900, 1923, 1152, 2874]. Structure [30, 2318, 1262, 2041, 394, 1439, 158, 2072, 702, 1540]. structure-shy [1262]. Structured [2815, 494, 525, 2174, 42, 1798, 321, 1482, 1020, 301, 1227, 229]. Structures [258, 157, 265, 2159, 1911, 2797, 1859, 309, 158, 922, 713, 1732, 224, 521, 185, 201]. Structuring [45, 2023]. Student [146]. Studies [289, 575, 573, 49, 2006, 1714, 2455, 2598, 1577, 2757]. Study [45, 31, 2825, 2571, 1888, 2232, 2877, 383, 2780, 1348, 1121, 906, 1554, 2843, 1786, 867, 2185, 1432, 1681, 2533, 156, 2109, 1500, 2243, 1443, 1237, 830, 1753, 2267, 2837, 857, 1423, 1148, 1734, 1922, 1873, 1556, 187, 2570, 1384, 1450, 1665, 2644, 2101, 301, 455, 2592, 2807, 597, 1238]. Studying [2679, 1294, 1566, 1789, 944, 1302]. stutter [2874]. style [378, 365, 285, 458, 1864, 1164, 696, 1326, 1105]. styles [799, 2578]. sub [1462]. sub-system [1462]. subexpression [2150]. subgoal [1633]. sublinear [1223, 436]. submachines [2059]. subservicing [928]. subset [673, 2228, 623]. subsets [658]. Substitution [383, 441, 509]. substitutivity [1214]. subtree [2356]. subtype [580]. subtyping [1284, 2506, 1207, 2371, 601, 629, 1438, 381, 2602]. succeed [698]. success [2485, 2829, 2055, 1629, 2847, 1624]. successes [2327]. sufficient [710, 1686]. suffix [410]. suite [1528, 2824, 2170, 2172, 2755]. suites [462, 2369]. summarization [2743]. Summary [1669]. Summary-based [1669]. summer [21]. sunt [2726]. supercombinator [113]. Supervisory [1700, 1770, 1988]. supply [1511]. Support [1600, 1712, 2839, 2742, 2834, 1090, 1417, 1632, 1506, 974, 1691, 591, 1660, 1152, 1103, 1618, 755]. Supporting [1479, 1298, 797, 1741, 439, 653, 1300, 822, 2348, 468, 736, 2705, 2108, 1493, 1651, 637, 2541, 1444, 1307, 432]. supportive [1336]. suppressing [2019]. surge [2232]. surprising [1802]. survey [926, 1592, 2267, 708, 756, 1762]. suspendible [971]. Sussex [218]. Sustaining [2647]. SVT [2018]. SVT’15 [2075]. swapping [2371]. switch [2668]. Switching [58]. Switzerland [319]. SWQL [2590]. SYB [1957]. Symbolic [2854, 2366, 460, 129, 1657, 1145, 300, 2027, 1353, 2228, 1884, 1810, 1841, 1564, 1369, 1147, 1575, 2584, 2804, 2878, 2229, 2428, 581, 1899, 1936]. symmetric [1646, 2658]. symmetrically [1439]. Symmetry [2079]. Symposium [1546, 1204, 1436, 2881, 2884, 2256, 1677, 2685, 2799, 2697, 2338, 2457, 2383, 837, 2535, 2276, 2880, 1746, 2882, 2031, 1329, 1215, 1360, 1400, 1416, 2687, 2002, 1670, 2426, 573, 232, 272, 2694, 2727, 1448]. Symposiums [1909]. Synchronisation [1005]. synchronization [682, 221, 29, 2275, 932, 1992, 1460, 848, 354]. synchronizer [849]. Synchronous [293, 1676, 943, 1918, 465, 329, 340, 691, 2401, 486, 2576, 1933, 432, 2592, 2375, 2556]. synopsis [1091]. Syntactic [1775, 135, 1485, 2877, 1718]. Syntactic-semantic [1775]. syntactical [1224]. syntax [1192, 1417, 366, 1679, 2764, 2288, 1483, 2431, 1297]. Synthesis [1581, 2240, 274, 4, 1271, 28, 13, 2147, 1220, 2854, 1770, 1695, 2423, 2411, 2222, 2680, 2721, 2321, 2796, 2627, 1330, 792, 2766, 1072]. synthesisers [2537]. synthesize [29]. Synthesizing [2583, 877, 2491]. System [2200, 815, 1306, 53, 2276, 228, 67, 1872, 1377, 2013, 1295, 1888, 2213, 2340, 1102, 747, 1026, 2808, 701, 753, 670, 886, 1200, 1805, 812, 1045, 2116, 751, 1047, 1614, 2062, 543, 1408, 678, 545, 842, 2371, 2073, 333, 1977, 2736, 1825, 2817, 2229, 1443, 2618, 1257, 1399, 1579, 819, 2667, 2754, 127, 2049, 792, 613, 2420, 1462, 2865, 1280, 2094, 1452, 502, 872, 891, 2454, 308, 1855, 1163, 1424, 1006, 2045, 1756, 1240, 2027, 2344, 170]. system-on-a-chip [2340]. system/software [1200]. Systematic [7, 51, 384, 405, 1903, 2825, 2561, 2212, 347, 1443, 2783, 277]. Systems [1428, 1204, 1436, 1800, 2130, 2088, 2197, 2354, 2603, 572, 83, 128, 1567, 460, 2310, 1236, 669, 1312, 119, 2038, 2452, 1572, 96, 229, 1996, 2800, 385, 1752, 2403, 574, 1934, 94, 90, 407, 1504, 1521, 973, 2524, 2462, 2163, 943, 1581, 408, 2258, 662, 1574, 1654, 1888, 476, 1077, 1065, 2201, 2241, 2090, 2419, 2039, 1971, 2036, 589, 2212, 1227, 2091, 727, 438, 2034, 2236, 1381, 2534, 252, 655, 2669, 1889, 340, 1805, 614, 869, 1131, 1479, 1626, 1187, 2425, 2329, 1298, 1050, 1497, 1596, 1935, 1368, 2666, 1308, 461, 1807, 1858, 1527, 534, 1662, 925, 2525]. systems [1322, 2788, 2876, 1092, 1970, 2170, 590, 585, 2041, 2721, 1606, 2429, 2330, 1523, 2809, 2077, 2798, 2040, 1051, 2878, 795, 2092, 1379, 1359, 1188, 1558, 1233, 1735, 2017, 2160, 1993, 337, 889, 2066, 2464, 1201, 611, 2071, 1302, 671, 1894, 1568, 1984, 1638, 1936, 2413, 890, 1446, 2435, 826, 1001, 404, 2453, 994, 1812, 2295, 1744, 2076, 413, 2158, 658, 1534, 845, 1307, 1623, 2845, 1494, 1488, 2604, 2363, 1550, 2547, 1462, 1618, 1101, 657, 817, 2465, 1251, 1698, 1315, 1758, 2454, 2396, 2766, 992, 619, 2089, 455, 2863, 920, 2556, 2755, 2441, 2313, 881, 2609, 560]. Systems [2405, 2626, 2286, 1632, 1904, 912, 1520, 1708, 49, 289, 152]. systems-of-systems [2465, 2405, 2626]. Systolic [437, 274, 256, 276, 279].

T [321, 288]. TABASCO [2794]. table [596]. tabled [2152]. tables [1217, 2342]. tabular [584, 1218]. Tackling [2447, 2652, 2109]. TACoS [2735]. tactic [1278]. tactics [260]. tag [1993]. tailored [2268]. taint [2261]. tainted [1549]. Taming [2003, 1579]. tamper [2244]. target [831]. targeting [859, 1032]. TASE [2256, 2426, 2727, 2867]. task [2091, 1554, 1552, 1547, 2279, 2592]. task-level [1547]. tasking [168]. tasks [1493]. Tatoo [1362]. tau [1367]. tau-confluence [1367]. Taupe [1542]. taxonomy [1223, 436]. Taylor [575, 270]. Tbilisi [573]. TCP [1997]. TCTL [2194, 1936]. TDFix} [2853]. teach [865]. Teaching [1408, 2772, 748, 749, 2619, 1538]. Team [2623]. teams [2011]. Technical [2267, 2633, 746, 2403]. Technique [17, 328, 1601, 1110, 1425, 2639, 2194, 267, 268, 800, 386]. Techniques [2088, 2197, 2354, 2603, 460, 2875, 262, 1892, 1173, 2147, 2749, 1039, 1420, 79, 982, 840, 2525, 548, 2353, 2624, 1650, 2882, 1850, 1848, 1182, 1132, 1210, 1580, 756, 2112, 2676, 2643, 143]. Technische [9]. technological [1712].

Technologies [1687, 1714, 1150, 2520, 1708]. Technology [1887, 56, 2652, 462, 1181]. telephone [545]. template [1473]. templates [1196, 2853, 2032]. Tempo [742]. temporal [1860, 1055, 29, 2664, 1766, 1347, 846, 154, 289, 193]. tenancy [2274]. tensor [371]. Term [52, 2414, 2413, 1764, 1326, 2739]. Term-Wise [52]. terminating [631, 1460]. Termination [128, 2162, 1137, 1667, 2483, 1765, 1141, 148, 1884, 1930]. Terms [981, 2366, 2877, 2325]. Terry [229]. Test [1694, 2247, 1723, 2384, 2400, 2659, 2440, 2591, 1531, 2747, 542, 2078, 2824, 462, 2584, 2639, 1383, 2172, 2827, 1351, 2027, 2805, 2819, 1811, 2385, 2271, 697, 1622, 920, 2863]. Test-data [2400]. Test-driven [2247]. test-suite [2824]. testable [452]. Testing [2852, 2075, 2842, 1810, 2459, 2258, 2006, 2020, 2306, 459, 2000, 2534, 1398, 1561, 2759, 1475, 2154, 1911, 2361, 2077, 2639, 2651, 2804, 2798, 153, 2797, 2017, 1213, 1153, 1460, 2326, 1899, 2369, 2076, 2587, 2733, 1555, 806, 2536, 2738, 2402, 1397, 1722]. testing-approach [2006]. tests [1514]. tetrahedral [1644]. Text [10, 770, 1898, 1297, 1710, 2014, 2764, 2756]. text-based [1297, 2764]. text-search [770]. TextFlows [2014]. Texts [146]. textual [2758]. Their [145, 784, 525, 2469, 599, 2343, 1108, 2146, 193, 1027, 1893, 2069, 683, 2145, 845]. them [2820, 262, 2480]. Theorem [124, 568, 1701, 1238]. Theorems [121]. Theoretic [31, 941, 1302, 1988, 2837]. Theoretical [2694, 2256, 154, 172, 388, 2687, 271, 2426, 56, 1023, 136, 397, 2515, 2727]. Theories [1499, 2580, 2277, 1658, 1950, 2652, 2257, 1869, 110, 1859, 2326, 627, 1875, 260, 1877, 1331]. Theorizing [1876]. Theory [2258, 306, 723, 527, 1166, 1926, 227, 1874, 2199, 407, 48, 1243, 2296, 1583, 1656, 701, 2674, 108, 1870, 1700, 2244, 253, 366, 367, 912, 720, 1245, 884, 665, 1595, 2028, 1873, 169, 2709, 238, 1872, 205, 1703, 2567, 2429, 1871, 298, 173]. Theory-oriented [1874]. thereover [2479]. thinning [1842]. thoughts [612]. thread [2253, 1959, 2511]. Thread-enforced [2511]. threaded [1554]. threads [1574, 2382, 2253]. threats [2109]. Three [23, 2355, 1521, 765, 585, 1963, 1232, 2029]. Three-valued [2355, 2029]. three-way [1963]. threshold [1305]. Time [2200, 120, 1933, 2658, 2831, 973, 662, 1654, 492, 466, 1248, 2079, 589, 2212, 1559, 1797, 2236, 1516, 340, 2552, 786, 2511, 1497, 1596, 1935, 2302, 29, 2579, 1314, 1092, 709, 951, 543, 1951, 631, 964, 2224, 289, 1657, 1825, 1188, 339, 615, 1859, 2362, 2412, 1824, 1936, 1842, 1446, 127, 149, 1001, 2132, 1864, 1623, 1947, 504, 395, 846, 2205, 2110, 880, 2279, 1324, 2441, 2505, 529, 762, 1375, 1965]. Time-symmetric [2658]. Timed [1907, 1825, 453, 2854, 1374, 2534, 1937, 1561, 2019, 1194, 585, 2735, 1350, 2194, 1095, 1859, 898, 2842, 2027, 1623, 1965]. Timewise [447]. timing [2236, 2098, 1049, 730, 781]. timing-dependent [2098]. TinyOS [1131]. TLA [2497, 2065, 458, 2428, 2227]. TLC [2065]. TM [2091]. toil [347]. token [1486]. tokenization [1326]. Tokyo [184, 179]. tolerance [812, 1826, 666]. Tolerant [109, 44, 35, 2201, 2280, 1672, 2158, 1998]. too [244]. Tool [2167, 1537, 1660, 772, 2831, 1817, 2730, 1528, 2785, 2223, 2455, 1089, 2135, 2875, 2185, 1422, 2769, 2814, 2876, 2269, 2786, 2170, 2546, 1573, 2168, 2675, 2562, 1592, 2853, 1051, 1079, 2835, 2844, 2768, 1044, 2715, 2800, 822, 2187, 591, 2733, 1727, 1165, 2585, 1742, 100, 2221, 1713, 2758, 2767, 2722, 1974, 2834, 2787]. tool-chain [2715]. Toolbox [2455, 2057]. ToolBus [492]. toolchain [2791, 2207, 2714]. toolkit [1131, 1981, 2794, 775, 2620]. Toolkits [1525, 1043, 1076, 1173, 1814, 2097]. Tools [1607, 2866, 79, 2625, 2580, 1480, 1173, 1161, 614, 2138, 1505, 1536, 2108, 2507, 974, 1592, 2873, 2886, 1132, 549, 2101, 2048, 1740, 553, 2852, 2775]. toolset [1081, 2233, 637, 1745, 1540, 1529]. Top [70, 1295, 444, 1611, 279, 1234, 264]. Top-Down [70, 1295, 444, 279]. topic [1566]. Topics [753, 1476]. Topological [2646]. topology [1807, 1253, 1155, 2427, 510]. Tor [1589]. torch [1882]. TOSCA [1991, 2291]. TOSCA-compliant [1991]. Total [5, 526]. totality [488]. tour [742]. TrABin [2349]. Trace [2505, 1563, 450, 1815, 2836, 1371, 2585, 1673, 1815]. trace-based [1673]. traceability [598, 2454]. traceable [1939]. tracer [1533]. traces [1121, 790, 1447]. Tracing [2015, 2333, 1836, 1573]. Track [1204, 1436, 2167, 1360, 1400, 1416, 1215]. Tracking [1755, 2633, 2586, 2267, 1542]. Tractable [536]. tracts [398, 397]. trade [167, 2205]. trade-offs [2205]. trading [2738]. Traffic [2200, 1756]. train [2447, 1755, 990]. trait [1437, 1837]. TraitRecordJ [1409]. traits [1389, 1409, 1801, 2484]. tranducers [301]. Trans [2743]. transaction [2061]. transactional [1471, 933, 850, 825, 937]. transactions [936, 1870, 1071, 1095]. transfer [1887]. transfers [330]. transform [2837, 2810, 2728]. Transformation [1039, 1262, 387, 741, 1537, 258, 1892, 973, 2818, 519, 2832, 1202, 1906, 1026, 1671, 1081, 1089, 900, 645, 599, 446, 1592, 1815, 1036, 1079, 2702, 1591, 1037, 1956, 1302, 1893, 2114, 1894, 158, 2627, 642, 1922, 714, 2646, 133, 501, 1035, 1321, 2514, 644, 1041, 1031, 641, 920]. transformation-based [501]. Transformational [89, 399, 30, 45, 102, 282, 187]. Transformations [732, 74, 92, 1038, 32, 525, 2120, 383, 2206, 2328, 907, 1170, 2511, 2259, 330, 1299, 2647, 1910, 1929, 2348, 1450, 420, 1040, 824, 738, 736, 1703, 739]. transformer [603, 2794, 601, 779, 390, 2743]. Transformers [359, 1979, 393, 471, 561]. Transforming [1324]. transient [737]. Transition [2027, 2419, 2041, 1825, 2420, 2547, 432, 2313, 2609]. transition-system [2420]. transitions [2428]. Transitive [13, 200]. Translated [290]. Translating [2379, 2065, 2437, 2186, 2381]. Translation [2846, 1161, 2450, 1108, 880]. translations [2627, 2150]. Transparent [2528]. transputer [218]. Traversal [2052, 679, 1482, 163]. traversals [1957, 2120]. Tree [361, 39, 122, 2120, 517, 2872, 372, 410, 223, 2080, 679, 163, 685, 1484, 1483, 2474, 301]. trees [426, 1020, 2491, 314, 561, 1283]. Tri [2078]. Tri-modal [2078]. triagement [2843]. triaging [2870, 2762]. trials [1628]. tribute [1391, 1268]. trivial [963, 962]. trope [1761]. trope-based [1761]. true [1935]. trust [2300, 802, 1103]. trust-dependent [1103]. trusted [1101]. Trustworthy [2349]. Trylock [2664]. Tse [321]. TSO [2406]. Tuning [1536]. tuple [1146, 957, 2712, 1990, 1413, 609, 2555, 666, 522]. tuple-space [1990]. tuProlog [827]. turbine [2094]. Turing [2026, 2658]. Turku [1274]. Turning [1394, 1888]. tutorial [746, 1760]. tutoring [2754]. TVL [1297]. Two [484, 1932, 275, 1848, 1754, 85, 241, 710, 2266, 326, 2661, 2217, 2112, 398]. two-level [2661, 398]. two-step [2266]. TXL [900]. Type [1137, 2407, 1221, 427, 1226, 93, 953, 740, 715, 306, 724, 1956, 716, 978, 334, 1801, 1979, 1932, 2527, 1995, 1187, 435, 1019, 2492, 1085, 411, 2371, 1198, 2073, 2207, 368, 941, 1257, 1245, 298, 1452, 1721, 381, 502, 1251, 308, 1740, 2756, 582, 1719, 597, 1610, 1251]. Type-2 [2407]. Type-based [1137]. type-centric [1198]. Type-changing [1956]. Type-Checker [93]. Type-Checking [427]. Type-directed [1221, 2492]. type-driven [1979].

Type-indexed [724]. Type-safe [740, 1801, 1019, 1251]. type-sound [1995]. type-theoretic [941]. Typechecking [115, 2207]. Typed [942, 2145, 1068, 1831, 485, 1221, 2492, 715, 1352, 487, 149, 2602]. Types [294, 427, 139, 259, 525, 1389, 2731, 1987, 108, 2299, 548, 630, 724, 367, 2573, 1588, 1266, 2252, 563, 853, 1959, 1073, 881, 784, 2062]. Typestate [840, 2705]. typesystem [2782]. TyphonML [2834]. Typing [2134, 1501, 1739, 2596, 2673, 366, 373].

Ubiquitous [1487, 1489, 1494, 1488]. ubiquity [1492]. UCAmI [1487]. UCL [516]. UI [1603]. UIDL [1600]. UK [231, 229, 227, 230]. Ulm [2665]. UML [973, 1089, 638, 821, 639, 786, 787, 1937, 645, 1046, 2500, 987, 1203, 1103, 767, 2865, 2767, 2813]. UML-based [1046, 1203]. UML-RT [973]. Umple [2599]. unambiguous [2469]. unbounded [2229, 981, 432]. uncaught [489]. uncertainties [2788]. uncertainty [2297, 2286]. Uncovering [511]. under-approximation [2078]. under-approximations [1686]. underapproximating [948]. underlying [1083]. underpinnings [1023]. understand [549]. understandable [188]. Understanding [1786, 1782, 2365, 2518, 1787, 1769, 1779, 1781, 549, 182]. Undulant [508]. Undulant-block [508]. unfold [2818]. unfold/fold [2818]. unfoldings [380]. unfolds [1500]. Ungar [170]. unidirectional [464]. Unification [4, 69, 518]. unified [2591, 2188, 1930]. uniform [2749, 525, 1935, 1468, 1832]. unify [1594, 1611]. Unifying [2507, 2326, 1304, 321, 1875, 1499]. uninterpreted [2857]. Unique [608]. unit [1514, 1899, 2110]. United [195, 202, 265, 155, 192, 243, 245, 206, 203, 213, 289, 285, 286, 287, 219, 244, 207, 246, 218, 321, 296, 217, 208, 297, 193, 298, 204, 194, 315, 317, 143, 214, 266, 269, 215, 270, 209, 201, 288, 284, 247, 248, 249, 250]. Units [1406]. UNITY [378]. Universal [2327, 632]. Universität [9]. University [515, 510, 146, 521, 182, 321, 298, 398, 181, 269, 397, 169, 574, 390, 173, 575, 209, 284, 191, 249, 316, 524]. unobtrusive [1493]. unparsed [1166]. unreachability [1728]. unrealizable [1357]. Unrestricted [1407]. unsatisfiable [1357]. unsorted [2227]. untyped [537]. up-down [440]. Update [2623, 1019, 2796, 1779]. updates [1019]. upgrades [2419]. Uppaal [2791]. upravel [325]. urban [2447, 2308]. USA [49, 232]. usability [1505, 2858, 1605]. Usage [1900, 2084, 1943, 2109, 2077, 1788, 2593]. Use [795, 2514, 1828, 2561, 1681, 2533, 991, 16, 2111, 2077, 913, 392, 1046, 245]. Use-case [795, 1828]. User [2656, 286, 1600, 1507, 1492, 2524, 2745, 1601, 2328, 1625, 656, 1490, 1363, 1602, 1493, 1606, 1691, 2858, 1604, 1148, 320, 2131, 2335, 2165, 2101, 1738, 2757, 1618]. user-centered [2858]. user-controlled [656]. user-defined [1363, 2165]. User-driven [2656]. User-guided [1492]. users [2185]. uses [513]. Using [2322, 1563, 2054, 1, 959, 865, 995, 262, 29, 2720, 998, 1476, 1761, 1511, 722, 551, 1320, 2102, 1768, 754, 666, 2762, 1017, 342, 920, 1162, 2679, 2571, 1430, 2163, 1294, 2854, 448, 1574, 1375, 1888, 1918, 383, 1907, 1801, 663, 1175, 1559, 1161, 2000, 2300, 1889, 1283, 1596, 2666, 2185, 380, 2128, 2062, 1943, 2259, 379, 1711, 2791, 1792, 842, 1571, 2233, 2063, 2861, 244, 1611, 1977, 1825, 2175, 749, 1233, 2066, 2618, 1201, 2597, 1850, 2069, 487, 2691, 565, 2008, 194, 1928, 877, 2348, 2667, 1873, 1707, 1203, 1304, 994, 1387, 1966, 602, 1353, 1450]. using [1623, 949, 497, 1247, 1580, 2764, 2356, 1811, 420, 2585, 1721, 283, 2638, 1566, 2528, 880, 1740, 1338, 288, 1181, 2268, 2427, 2676, 2314, 2317, 1332, 804, 881, 1362, 1129]. UsiXML [1506, 1603]. usual [2070]. UTP [2507, 1331].

V [244, 268, 142]. V2 [2135]. valid [1811]. Validated [1642]. Validating [2534, 2233, 1046, 2168]. Validation [2039, 1393, 2248, 639, 2259, 2065, 2063, 1208, 1811, 739]. Vallduví [573]. Valley [21]. valuation [2659]. value [2692, 485, 2645, 2787, 2729, 1325, 595, 951, 2429, 1887, 1588, 1689, 2132, 811, 2404]. value-based [2787]. valued [2029, 2355]. values [630, 2092]. Variability [764, 2235, 2423, 2519, 760, 1517, 1903, 761, 2035, 763, 2268, 1293, 2313, 762]. variability-aware [2519, 1903]. variable [1940, 1549]. Variables [294, 2854, 2729, 686, 2275, 722, 2664, 283]. Variant [2198]. Variant-based [2198]. variants [766]. variation [765, 763]. various [464]. VDM [1692, 288, 2497, 337]. vector [2742, 1721]. VeGAn [2787]. VeGAn-Tool [2787]. vehicle [2161]. VeriFast [1577]. Verifiable [1743]. Verification [2163, 2392, 1567, 2200, 542, 84, 543, 2857, 2576, 2038, 1572, 96, 458, 1752, 1753, 150, 1722, 2075, 1934, 1477, 1332, 838, 1377, 1430, 1663, 1570, 2567, 1077, 1376, 2232, 1775, 1729, 2061, 1806, 1661, 1559, 1912, 2025, 2093, 1161, 1554, 1113, 1363, 1564, 2552, 703, 2285, 787, 1369, 1717, 2153, 462, 840, 2507, 464, 1108, 1911, 2275, 1431, 2490, 2330, 842, 1064, 1275, 1474, 2122, 2161, 2748, 2395, 1350, 2194, 62, 1095, 789, 790, 1973, 2618, 1997, 1266, 603, 1848, 2837, 565, 2667, 1556, 1001, 1967, 1330, 1577, 1966, 2158, 1551, 1580, 2094, 1267, 2465, 2035]. verification [234, 2474, 2468, 805, 619, 2089, 2556, 1756, 1335, 1641, 151]. verification-condition [790]. Verified [2347, 2680, 1906, 2666, 1116, 1114, 1621]. Verifier [1396]. verify [2332, 2330, 2187, 2613]. Verifying [2119, 1375, 460, 1114, 1792, 2157, 69, 392, 2653, 1888, 670, 2534, 2669, 614, 461, 1327, 1575, 1422, 544, 2233, 1755, 2500, 1968, 2604, 2501, 2767, 2427]. vers [60]. Versioned [936]. versions [2459, 2685, 2239, 2391, 1626, 189, 1960, 2582, 2737, 1300, 434]. very [612, 2169]. Vessels [23]. via [1810, 1866, 725, 1415, 1841, 664, 2871, 1363, 2026, 1272, 1717, 2153, 2712, 510, 717, 387, 2584, 2361, 2165, 2205, 2529, 738, 971, 2743]. VIATRA2 [1041]. Vicious [1917]. Vickers [510]. Vicky [288]. Victor [226]. video [156]. view [2298, 2295, 2188]. Viewing [1486]. Viewpoints [1472]. Views [165, 1295, 1170, 1777]. Vigorous [2294]. viii [126, 154]. Vincent [2779]. violations [2864, 1900]. Virtual [1535, 2425, 859, 829, 2339, 831, 833, 2070, 1152, 673, 1290, 834, 2813]. virtualization [1662]. virtualized [1402]. VirtuosoNext [2552]. visitors [2470]. Vissers [267]. Visual [119, 1892, 986, 1817, 643, 2011, 2358, 2014, 1540, 644, 2758]. visualization [1617, 551, 1176, 860, 1060]. Visualizations [2784, 876]. visualize [1526]. Visualizing [1131, 1542, 1532, 1177]. Visually [1637, 1834]. VLISP [60]. VLSI [280]. VM [1836]. VMC [2722]. Vogel [573]. Voice [863]. Vol [575, 398]. volatility [1084]. Volume [345, 388, 81, 402, 414, 433, 587, 600, 610, 624, 633]. Volumes [174, 346, 156]. volunteer [1661]. VP [615]. vs [2181, 2242, 1613]. vulnerabilities [2322, 2015, 2773]. vulnerability [2859].

W [375, 204, 524]. Wadler [202, 371]. Wait [403, 332]. Wait-free [403, 332]. WAKERS [2283]. Walter [142]. Warsaw [272, 2886]. Warshall [99]. was [418]. WASDeTT [1173]. Watt [297]. way [1963, 1290]. WCET [2092]. WCRE [1346]. Weak [1867, 1376, 2672, 1469, 1367, 2489]. Weakest [1633, 1459, 2471, 363]. weakly [1208]. weaving [1221, 1056, 1289]. web [2424, 1688, 1693, 1691, 2056, 1811, 2810, 2833, 1794, 2825, 1570, 2015, 746, 747, 1625, 1150, 1056, 1640, 1069, 1687, 998, 2590, 909, 1470, 889, 822, 1062, 1125, 1219, 1508, 1878, 1618, 1324, 1326]. web-based [2810, 746, 822, 1618]. WebDraw} [2844]. WebML [1625]. WebRatio [1625]. website [2844]. Weighted [845, 2640, 1524]. Weihrauch [172]. Weinberg [207]. Welcome [3]. well [2492]. well-typed [2492]. Wellings [168]. Welsh [185]. Wesley [296, 208, 204, 143, 288, 166, 230, 247, 248, 250]. Wesley/ACM [320]. West [218]. WG17 [402]. Where [2403, 2750]. Which [2533, 2214]. while [595, 311]. Whiley [1968]. Wide [6, 1905]. widening [2005, 839]. WiFi [863]. wild [1916]. wildcards [1094]. Wiley [572, 231, 151, 203, 289, 207, 218, 153, 317, 145]. Wiley-Interscience [572]. Wilhelm [231]. Will [2873]. Wilson [208]. wind [2094]. Window [1997]. windowing [862]. wire [2525, 1330]. wire-wise [1330]. wireless [1559, 1388, 1545, 749, 1569, 1491]. Wirth [201]. Wise [52, 1330]. without [2120, 347, 578]. witness [2529]. Wodel [2786]. Wodel-Edu [2786]. Wokingham [296, 208, 204, 143, 288, 230]. wollongong [21]. Wong [49]. Woodcock [206]. Word [3, 2762]. work [612, 1136, 2045]. work-based [2045]. workbench [2545, 1048]. Workflow [2344, 2732, 2551, 1606, 1611, 1387, 1285, 2550]. Working [669, 299, 2110]. works [291]. Workshop [1274, 2283, 669, 1143]. workshops [1659]. World [1858, 1736]. worlds [1104]. worm [2332]. wrapping [2086]. write [1471, 1621]. writers [1238]. writing [452]. written [2440]. WRLA [1854]. WS [1320, 1850, 1018]. WS-BPEL [1320, 1850, 1018]. WSDL [1640, 1326]. WSNs [2202]. WT [1687].

X [396, 521, 524, 285, 270, 730, 284]. XACML [2832, 1582]. XACML2mCRL2 [2832]. Xatkit [2828]. XGLR [901]. xi [575]. xii [193, 320]. xiv [78]. XML [1068, 978, 974, 1261, 779, 1251]. XML11 [862]. XPath [1262]. XPi [1068]. XPL [1846]. XSL [1299]. XSS [2691]. XT [1081]. XTP [458]. xv [317]. XVCL [766]. xx [730]. xxvi [396]. xxxi [573].

year [553]. years [997, 1626, 2521]. Yield [1608, 1063]. yielded [83]. Yoneda [442]. York [184, 291, 572, 79, 207, 179, 396]. York/London/Paris/Tokyo [184, 179]. Yourdon [246]. yRCA [2790].

Z [182, 1692, 2497, 1392, 1116, 1114, 590, 591]. Z/Eves [1116, 1114]. Z3 [2317]. zero [2600]. Ziegler [67]. zipper [2343]. zipper-based [2343]. zippers [2069]. Zone [2236]. Zonebased [2236]. zoo [1949, 1816]. zoom [1924]. zoom-declarative [1924]. Zurab [573].

References

Cai:19xx:CRU

[1] W. Cai and D. B. Skillicorn. Calculating recurrences using the Bird– Meertens formalism. Science of Computer Programming, ??(??):??, 19xx. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sintzoff:1981:E

[2] Michel Sintzoff. Editorial. Science of Computer Programming, 1 (1–2):1–2, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dijkstra:1981:WW

[3] Edsger W. Dijkstra. A word of welcome. Science of Computer Programming, 1(1–2):3–4, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Manna:1981:DSU

[4] Zohar Manna and Richard Waldinger. Deductive synthesis of the unification algorithm. Science of Computer Programming, 1(1–2):5–48, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hoare:1981:CTC

[5] C. A. R. Hoare. A calculus of total correctness for communicating processes. Science of Computer Programming, 1 (1–2):49–72, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bauer:1981:PWS

[6] F. L. Bauer, M. Broy, W. Dosch, R. Gnatz, B. Krieg-Brückner, A. Laut, M. Luckmann, T. Matzner, B. Möller, H. Partsch, P. Pepper, K. Samelson, R. Steinbrüggen, M. Wirsing, and H. Wössner. Programming in a wide spectrum language: a collection of examples. Science of Computer Programming, 1(1–2):73–114, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Best:1981:SDE

[7] E. Best and F. Cristian. Systematic detection of exception occurrences. Science of Computer Programming, 1(1– 2):115–144, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Feijen:1981:BRP

[8] W. H. J. Feijen. Bulterman’s reshuffling problem. Science of Computer Programming, 1(1–2):145–147, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1981:CPC

[9] Anonymous. Call for papers: 7th Conference on programming languages and program development: March 9 and 10, 1982, Technische Universität München, Munich, Fed. Rep. Germany. Science of Computer Programming, 1(1– 2):149–155, October 1981. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642381900095.

Sufrin:1982:FSD

[10] Bernard Sufrin. Formal specification of a display-oriented text editor. Science of Computer Programming, 1 (3):157–202, May 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sethi:1982:CEE

[11] Ravi Sethi. Circular expressions — elimination of static environments. Science of Computer Programming, 1 (3):203–222, May 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Also see Automata, Languages and Programming, Eighth Colloquium, Acre Israel, Lecture Notes in Computer Science 115, pp. 378-392, Springer-Verlag, Berlin, 1981.

Dijkstra:1982:SAS

[12] Edsger W. Dijkstra. Smoothsort, an alternative for sorting in situ. Science of Computer Programming, 1(3):223–233, May 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See errata [19].

Schmitz:1982:EPS

[13] Lothar Schmitz. An exercise in program synthesis: Algorithms for computing the transitive closure of a relation. Science of Computer Programming, 1(3):235–254, May 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lengauer:1982:MPCa

[14] Christian Lengauer and Eric C. R. Hehner. A methodology for programming with concurrency: An informal presentation. Science of Computer Programming, 2(1):1–18, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lengauer:1982:MPCb

[15] Christian Lengauer. A methodology for programming with concurrency: The formalism. Science of Computer Programming, 2(1):19–52, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Guttag:1982:SNP

[16] John V. Guttag, Jim J. Horning, and Jeannette W. Wing. Some notes on putting formal specifications to productive use. Science of Computer Programming, 2(1):53–68, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Heilbrunner:1982:TRR

[17] Stephan Heilbrunner and Werner Pohlmann. A technique for recursion removal based on string homomorphisms. Science of Computer Programming, 2(1):69–81, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1982:B

[18] Anonymous. Books. Science of Computer Programming, 2(1):83, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dijkstra:1982:ESA

[19] Edsger W. Dijkstra. Errata: “Smoothsort, an alternative for sorting in situ” [Sci. Comput. Programming 1 (1981/ 82), no. 3, 223–233; MR 83g:68090]. Science of Computer Programming, 2(1):85, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See [12].

Anonymous:1982:CPI

[20] Anonymous. Call for papers: 10th International colloquium on automata, languages, and programming: 18–22 July 1983, Barcelona, Spain. Science of Computer Programming, 2 (1):87–88, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642382900089.

Anonymous:1982:WSS

[21] Anonymous. Announcement: The 1st wollongong summer school on the science of programming: 31 January–9 February 1983, Sponar’s Chalet, Diggers Creek on Lake Kosciusco, Kosciusco Road, Perisher Valley, N.S.W., Australia. Science of Computer Programming, 2 (1):89, October 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642382900090.

London:1982:ISF

[22] Philip E. London and Martin S. Feather. Implementing specification freedoms. Science of Computer Programming, 2(2):91–131, November 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1982:FSP

[23] Martin Rem and Young Il Choo. A fixed-space program of linear output complexity for the problem of the three vessels. Science of Computer Programming, 2(2):133–141, November 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Misra:1982:FRE

[24] J. Misra and David Gries. Finding repeated elements. Science of Computer Programming, 2(2):143–152, November 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Elrad:1982:DDP

[25] Tzilla E. Elrad and Nissim Francez. Decomposition of distributed programs into communication-closed layers. Science of Computer Programming, 2(3): 155–173, December 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lamport:1982:ACP

[26] Leslie Lamport. An assertional correctness proof of a distributed algorithm. Science of Computer Programming, 2(3):175–206, December 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gries:1982:NSS

[27] David Gries. A note on a standard strategy for developing loop invariants and loops. Science of Computer Programming, 2(3):207–214, December 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mints:1982:JSS

[28] Grigori Mints and Enn Tyugu. Justifications of the structural synthesis of programs. Science of Computer Programming, 2(3):215–240, December 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Emerson:1982:UBT

[29] E. Allen Emerson and Edmund M. Clarke. Using branching time temporal logic to synthesize synchronization skeletons. Science of Computer Programming, 2(3):241–266, December 1982. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Partsch:1983:ETD

[30] H. Partsch. An exercise in the transformational derivation of an efficient program by joint development of control and data structure. Science of Computer Programming, 3(1):1–35, April 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pritchard:1983:CSN

[31] Paul A. Pritchard. A case study of number-theoretic computation: Searching for primes in arithmetic progression. Science of Computer Programming, 3(1):37–63, April 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Apt:1983:PRT

[32] K. R. Apt and E.-R. Olderog. Proof rules and transformations dealing with fairness. Science of Computer Programming, 3(1):65–100, April 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Francez:1983:ENC

[33] Nissim Francez. Extended naming conventions for communicating processes. Science of Computer Programming, 3 (1):101–114, April 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gries:1983:BRB

[34] D. Gries. Book review: How to solve it by computer, R. G. Dromy. Prentice-Hall International Inc., London, 1982, 464 pp. Science of Computer Programming, 3 (1):115–116, April 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642383900060.

Anonymous:1983:AGN

[35] Anonymous. Announcement: 2nd GI/NTG/GMR conference on faulttolerant computing: September 19– 21, 1984, Bonn, Fed. Rep. Germany. Science of Computer Programming, 3(1):117, April 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642383900072.

Paulson:1983:HOI

[36] Lawrence C. Paulson. A higher-order implementation of rewriting. Science of Computer Programming, 3(2):119–149, August 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

                                        tronic). URL http://www.cl.cam.ac.
                                        uk/Research/Reports/TR035-lcp-rewriting.
                                        dvi.gz.

Kahn:1983:MFS

[37] G. Kahn, B. Lang, B. Mélèse, and E. Morcos. Metal: a formalism to specify formalisms. Science of Computer Programming, 3(2):151–188, August 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dix:1983:EIC

[38] Trevor I. Dix. Exceptions and interrupts in CSP. Science of Computer Programming, 3(2):189–204, August 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Katajainen:1983:AIK

[39] Jyrki Katajainen and Olli Nevalainen. An alternative for the implementation of Kruskal’s minimal spanning tree algorithm. Science of Computer Programming, 3(2):205–216, August 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1983:SPEa

[40] M. Rem. Small programming exercises. Science of Computer Programming, 3 (2):217–222, August 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ganzinger:1983:IML

[41] Harald Ganzinger. Increasing modularity and language-independency in automatically generated compilers. Science of Computer Programming, 3(3):

223–278, December 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See also: Bericht TUM-I8306, Institut für Informatik, Tech. University München (July 1983).

Barbuti:1983:SAS

[42] R. Barbuti and A. Martelli. A structured approach to static semantics correctness. Science of Computer Programming, 3(3):279–312 (or 279–311??), December 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1983:SPEb

[43] M. Rem. Small programming exercises 2. Science of Computer Programming, 3(3):313–319, December 1983. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schneider:1984:FTB

[44] Fred B. Schneider, David Gries, and Richard D. Schlichting. Fault-tolerant broadcasts. Science of Computer Programming, 4(1):1–15, April 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Partsch:1984:STD

[45] H. Partsch. Structuring transformational developments: a case study based on Earley’s recognizer. Science of Computer Programming, 4(1):17–44, April 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Abrial:1984:MCP

[46] Jean-Raymond R. Abrial. The mathematical construction of a program. Science of Computer Programming, 4(1): 45–86, April 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1984:SPEa

[47] M. Rem. Small programming exercises 3. Science of Computer Programming, 4(1):87–94, April 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Adiba:1984:BRB

[48] Michel Adiba. Book review: The theory of relational databases, by David Maier. Computer Science Press, 1983. Science of Computer Programming, 4(1):95–96, April 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642384900133.

Anonymous:1984:BRB

[49] Anonymous. Book review: Algorithmic studies in mass storage systems, by C. K. Wong. Computer Science Press, Rockville, MD, USA, 1983. Science of Computer Programming, 4(1):96–97, April 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642384900145.

Anonymous:1984:A

[50] Anonymous. Announcements. Science of Computer Programming, 4 (1):99–101, April 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642384900157.

Broy:1984:SAL

[51] Manfred Broy and Friedrich L. Bauer. A systematic approach to language constructs for concurrent programs. Science of Computer Programming, 4 (2):103–139, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sokolowski:1984:PCT

[52] Stefan Sokołowski. Partial correctness: The term-wise approach. Science of Computer Programming, 4 (2):141–157, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gerth:1984:PSC

[53] Rob Gerth and Willem P. De Roever. A proof system for concurrent ADA programs. Science of Computer Programming, 4(2):159–204, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1984:SPEb

[54] M. Rem. Small programming exercises 4. Science of Computer Programming, 4(2):205–210, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Apt:1984:B

[55] Krzysztof R. Apt and Robert Paige. Books. Science of Computer Programming, 4(2):211–218, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167642384900200.

Anonymous:1984:CFS

[56] Anonymous. Announcements: 4th conference on foundations of software technology & theoretical computer science: 13–15 December 1984, Bangalore, India. Science of Computer Programming, 4(2):219, August 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642384900212.

Wagner:1984:CQP

[57] Robert Wagner and Robert Geist. The crippled queen placement problem. Science of Computer Programming, 4(3): 221–248, December 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shyamasundar:1984:SLF

[58] R. K. Shyamasundar. A simple livelock-free algorithm for packet switching. Science of Computer Programming, 4(3):249–256, December 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Manna:1984:APP

[59] Zohar Manna and Amir Pnueli. Adequate proof principles for invariance and liveness properties of concurrent programs. Science of Computer Programming, 4(3):257–290 (or 257–289??), December 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cointe:1984:EVV

[60] Pierre Cointe. Une extension de VLISP vers les objets. (French) [An extension of VLISP to objects]. Science of Computer Programming, 4(3): 291–322, December 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1984:SPEc

[61] M. Rem. Small programming exercises 5. Science of Computer Programming, 4(3):323–333, December 1984. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Koomen:1985:ASV

[62] C. J. Koomen. Algebraic specification and verification of communication protocols. Science of Computer Programming, 5(1):1–36, February 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Smith:1985:DDC

[63] Douglas R. Smith. The design of divide and conquer algorithms. Science of Computer Programming, 5 (1):37–58, February 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tarlecki:1985:LSP

[64] Andrzej Tarlecki. A language of specified programs. Science of Computer Programming, 5(1):59–81, February 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jalili:1985:GIE

[65] Fahimeh Jalili. A general incremental evaluator for attribute grammars. Science of Computer Programming, 5 (1):83–96, February 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1985:SPEa

[66] M. Rem. Small programming exercises 6. Science of Computer Programming, 5(1):97–106, February 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Meyer:1985:BRB

[67] Bertrand Meyer. Book review: Programming System methodologies, by Carol A. Ziegler. Prentice-Hall, Englewood Cliffs, NJ, 1983; ISBN 0-13-729905-2. Science of Computer Programming, 5(1):107–109, February 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167642385900073.

Meyer:1985:SPS

[68] Bertrand Meyer, Jean-Marc Nerson, and Soon Hae Ko. Showing programs on a screen. Science of Computer Programming, 5(2):111–142, June 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Paulson:1985:VUA

[69] Lawrence C. Paulson. Verifying the unification algorithm in LCF. Science of Computer Programming, 5 (2):143–169, June 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bergstra:1985:TDA

[70] J. A. Bergstra and J. V. Tucker. Top-down design and the algebra of communicating processes. Science of Computer Programming, 5(2):171–199, June 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Finance:1985:MLC

[71] Jean-Pierre Finance and Jeanine Souquières. A method and a language for constructing iterative programs. Science of Computer Programming, 5(2):201–218, June 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1985:SPEb

[72] M. Rem. Small programming exercises 7. Science of Computer Programming, 5(2):219–229, June 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1985:A

[73] Anonymous. Announcement. Science of Computer Programming, 5 (2):231, June 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642385900139.

Souquieres:1985:DII

[74] Jeanine Souquières and Jean-Pierre Finance. Description and improvement of iterative program transformations. Science of Computer Programming, 5 (3):233–264, October 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Martin:1985:DME

[75] Alain J. Martin. Distributed mutual exclusion on a ring of processes. Science of Computer Programming, 5 (3):265–276, October 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moitra:1985:ACC

[76] Abha Moitra. Automatic construction of CSP programs from sequential nondeterministic programs. Science of Computer Programming, 5 (3):277–307, October 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1985:SPEc

[77] M. Rem. Small programming exercises 8. Science of Computer Programming, 5(3):309–316, October 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jackson:1985:BRB

[78] Michael Jackson. Book review: Construction of Data Processing Software, by John Elder. Prentice Hall, Englewood Cliffs, NJ, 1984; xiv + 432 pp. Science of Computer Programming, 5 (3):317–318, ???? 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642385900188.

Feldman:1985:BRB

[79] Jerome A. Feldman and Thomas LeBlanc. Book review: Coordinated computing: Tools and techniques for distributed software, by Robert E. Filman and Daniel F. Friedman. McGraw-Hill, New York, 1984. Science of Computer Programming, 5 (3):318–320, ???? 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764238590019X.

Anonymous:1985:CP

[80] Anonymous. Call for papers. Science of Computer Programming, 5 (3):321, ???? 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642385900206.

Anonymous:1985:AIV

[81] Anonymous. Author index volume 5 (1985). Science of Computer Programming, 5(3):323, ???? 1985. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642385900218.

McCurley:1986:DDA

[82] Robert McCurley and Fred B. Schneider. Derivation of a distributed algorithm for finding paths in directed networks. Science of Computer Programming, 6(1):1–9, January 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bellegarde:1986:RSF

[83] Françoise Bellegarde. Rewriting systems on FP expressions to reduce the number of sequences yielded. Science of Computer Programming, 6 (1):11–34, January 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Francez:1986:SCA

[84] Nissim Francez, Brent Hailpern, and Gadi Taubenfeld. Script: a communication abstraction mechanism and its verification. Science of Computer Programming, 6(1):35–88, January 1986.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wiltink:1986:TMN

[85] J. G. Wiltink. Two most nondeterministic programs. Science of Computer Programming, 6(1):89–94, January 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1986:SPEa

[86] M. Rem. Small programming exercise 9. Science of Computer Programming, 6(1):95–102, January 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Guttag:1986:RLS

[87] J. V. Guttag and J. J. Horning. Report on the Larch shared language. Science of Computer Programming, 6 (2):103–134, March 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Guttag:1986:LSL

[88] J. V. Guttag and J. J. Horning. A Larch shared language handbook. Science of Computer Programming, 6(2): 135–158 (or 135–157??), March 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bird:1986:TPP

[89] R. S. Bird. Transformational programming and the paragraph problem. Science of Computer Programming, 6(2): 159–190 (or 159–189??), March 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zielinski:1986:MEP

[90] Krzysztof Zielinski. Model of error propagation in systems of communicating processes. Science of Computer Programming, 6(2):191–205, March 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1986:SPEb

[91] M. Rem. Small programming exercises 10. Science of Computer Programming, 6(2):207–211, March 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanDiepen:1986:PDT

[92] N. W. P. van Diepen and W. P. de Roever. Program derivation through transformations: the evolution of listcopying algorithms. Science of Computer Programming, 6(3):213–272, May 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fairbairn:1986:NTC

[93] Jon Fairbairn. New type-checker for a functional language. Science of Computer Programming, 6(3):273–290, May 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Stoye:1986:MBF

[94] William Stoye. Message-based functional operating systems. Science of Computer Programming, 6(3):291–311, May 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1986:SPEc

[95] M. Rem. Small programming exercises 11. Science of Computer Programming, 6(3):313–318, May 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Joseph:1986:SMA

[96] Thomas A. Joseph, Thomas Räuchle, and Sam Toueg. State machines and assertions: an integrated approach to modeling and verification of distributed systems. Science of Computer Programming, 7(1):1–22, July 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanderNat:1986:ACD

[97] M. van der Nat and A. A. Verrijn-Stuart. Analysing a class of distributive partitioning sort algorithms. Science of Computer Programming, 7(1):23–33, July 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Broy:1986:AIP

[98] Manfred Broy, Bernhard Möller, Peter Pepper, and Martin Wirsing. Algebraic implementations preserve program correctness. Science of Computer Programming, 7(1):35–53, July 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vandeSnepscheut:1986:DDI

[99] Jan L. A. van de Snepscheut. A derivation of a distributed implementation of Warshall’s algorithm. Science of Computer Programming, 7(1):55–60, July 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Voisin:1986:CTI

[100] Frederic Voisin. CIGALE: a tool for interactive grammar construction and expression parsing. Science of Computer Programming, 7(1):61–86, July 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1986:SPEd

[101] M. Rem. Small programming exercises 12. Science of Computer Programming, 7(1):87–97, July 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Partsch:1986:TPD

[102] H. Partsch. Transformational program development in a particular problem domain. Science of Computer Programming, 7(2):99–241, September 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1986:SPEe

[103] M. Rem. Small programming exercises 13. Science of Computer Programming, 7(2):243–248, September 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Burn:1986:SAH

[104] Geoffrey L. Burn, Chris Hankin, and Samson Abramsky. Strictness analysis for higher-order functions. Science of Computer Programming, 7(3): 249–278, November 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Josephs:1986:FPS

[105] Mark B. Josephs. Functional programming with side-effects. Science of Computer Programming, 7(3):279–296, November 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Inoue:1986:COM

[106] Katsuro Inoue, Hiroyuki Seki, Kenichi Taniguchi, and Tadao Kasami. Compiling and optimizing methods for the functional language Asl/F. Science of Computer Programming, 7(3): 297–312, November 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1986:SPEf

[107] M. Rem. Small programming exercises 14. Science of Computer Programming, 7(3):313–319, November 1986. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chisholm:1987:DPA

[108] Paul Chisholm. Derivation of a parsing algorithm in Martin-Löf’s theory of types. Science of Computer Programming, 8(1):1–42, February 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Joseph:1987:PRF

[109] Mathai Joseph, Abha Moitra, and Neelam Soundararajan. Proof rules for fault tolerant distributed programs. Science of Computer Programming, 8 (1):43–67, February 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kirchner:1987:RIG

[110] Claude Kirchner and Hélène Kirchner. REVEUR-3: The implementation of a general completion procedure parameterized by built-in theories and strategies. Science of Computer Programming, 8(1):69–86, February 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kolbl:1987:LFS

[111] Stefan Kolbl and Mitchell Wand. Linear future semantics and its implementation. Science of Computer Programming, 8(1):87–103, February 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1987:SPEa

[112] M. Rem. Small programming exercises 15. Science of Computer Programming, 8(1):105–112, February 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bird:1987:FDE

[113] R. S. Bird. A formal development of an efficient supercombinator compiler. Science of Computer Programming, 8 (2):113–137, April 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gries:1987:MM

[114] D. Gries and J. F. Prins. Mclaren’s masterpiece. Science of Computer Programming, 8(2):139–145, April 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cardelli:1987:BPT

[115] Luca Cardelli. Basic polymorphic typechecking. Science of Computer Programming, 8(2):147–172, April 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cousineau:1987:CAM

[116] G. Cousineau, P.-L. Curien, and M. Mauny. The categorical abstract machine. Science of Computer Programming, 8(2):173–202, April 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1987:SPEb

[117] M. Rem. Small programming exercises 16. Science of Computer Programming, 8(2):203–211, April 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Smith:1987:ASD

[118] Douglas R. Smith. Applications of a strategy for designing divideand-conquer algorithms. Science of Computer Programming, 8(3):213–229, June 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Harel:1987:SVF

[119] David Harel. Statecharts: a visual formalism for complex systems. Science of Computer Programming, 8(3):231–274, June 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Emerson:1987:MMC

[120] E. Allen Emerson and Chin-Laung Lei. Modalities for model checking: Branching time logic strikes back. Science of Computer Programming, 8(3):275–306, June 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sun:1987:DSE

[121] Yong-Qiang Sun and Ren-Ling Huang. Deriving some expansion theorems for nonlinear functional equations by an algorithm. Science of Computer Programming, 9(1):1–15, August 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pritchard:1987:LPN

[122] Paul Pritchard. Linear prime-number sieves: a family tree. Science of Computer Programming, 9(1):17–35, August 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Manna:1987:OBS

[123] Zohar Manna and Richard Waldinger. The origin of a binary-search paradigm. Science of Computer Programming, 9 (1):37–83, August 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kaldewaij:1987:SAB

[124] Anne Kaldewaij. Some algorithms based on the dual of Dilworth’s theorem. Science of Computer Programming, 9(1):85–89, August 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1987:SPEc

[125] M. Rem. Small programming exercises 18. Science of Computer Programming, 9(1):91–100, August 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1987:BRB

[126] Anonymous. Book review: Communicating Sequential Processes, By C. A. R. Hoare. Prentice-Hall International, London, 1985, viii + 256 pages. Science of Computer Programming, 9 (1):101–105, August 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642387900281.

Nielson:1987:HLP

[127] Hanne Riis Nielson. A Hoare-like proof system for analysing the computation time of programs. Science of Computer Programming, 9(2):107–136, October 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

BenCherifa:1987:TRS

[128] Ahlem Ben Cherifa and Pierre Lescanne. Termination of rewriting systems by polynomial interpretations and its implementation. Science of Computer Programming, 9(2):137–159, October 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Giannotti:1987:SES

[129] Fosca Giannotti, Attilio Matteucci, Dino Pedreschi, and Franco Turini. Symbolic evaluation with structural recursive symbolic constants. Science of Computer Programming, 9 (2):161–177, October 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Komorowski:1987:LPR

[130] Henryk Jan Komorowski and Jan Maluszynski. Logic programming and rapid prototyping. Science of Computer Programming, 9(2):179–205, October 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1987:SPEd

[131] M. Rem. Small programming exercises 19. Science of Computer Programming, 9(2):207–211, October 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dijkstra:1987:HEB

[132] Edsger W. Dijkstra. A heuristic explanation of Batcher’s baffler. Science of Computer Programming, 9(3): 213–220, December 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pepper:1987:SCP

[133] Peter Pepper. A simple calculus for program transformation (inclusive of induction). Science of Computer Programming, 9(3):221–262, December 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fokkinga:1987:CPS

[134] Maarten M. Fokkinga. A correctness proof of sorting by means of formal procedures. Science of Computer Programming, 9(3):263–270 (or 263–269??), December 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Boullier:1987:NER

[135] Pierre Boullier and Martin Jourdan. New error repair and recovery scheme for lexical and syntactic analysis. Science of Computer Programming, 9(3):

271–286, December 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morris:1987:TBS

[136] Joseph M. Morris. A theoretical basis for stepwise refinement and the programming calculus. Science of Computer Programming, 9(3):287–306, December 1987. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Baruch:1988:PIS

[137] Orit Baruch and Shmuel Katz. Partially interpreted schemas for CSP programming. Science of Computer Programming, 10(1):1–18, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

KrusemanAretz:1988:DLP

[138] F. E. J. Kruseman Aretz. On deriving a Lisp program from its specification. Science of Computer Programming, 10(1):19–32, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kapur:1988:CII

[139] Deepak Kapur and Mandayam Srivas. Computability and implementability issues in abstract data types. Science of Computer Programming, 10 (1):33–63, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Broy:1988:NDF

[140] Manfred Broy. Nondeterministic data flow programs: How to avoid the merge anomaly. Science of Computer Programming, 10(1):65–85, February 1988.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Andrews:1988:BRB

[141] D. J. Andrews. Book review: Programming from first principles, by Richard Bornat. Prentice-Hall International, London, 1986, Price £10.95, ISBN 13-729104-3. Science of Computer Programming, 10 (1):87–88, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900172.

Wordsworth:1988:BRB

[142] J. B. Wordsworth. Book review: A first course in formal logic and its applications in computer science, by R. D. Dowsing, V. J. Rayward-Smith and C. D. Walter. Blackwell Scientific Publishers, Oxford, 1986, Price £12.95. Science of Computer Programming, 10 (1):88–89, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900184.

Parke:1988:BRB

[143] Tom Parke. Book review: Software specification techniques, by N. Gehani and A. D. McGettrick, Editors. Addison-Wesley, Wokingham, United Kingdom, 1987, 477 pages, Price £23.95, ISBN 201-14239-9. Science of Computer Programming, 10 (1):90–92, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900196.

Stavridou:1988:BRB

[144] V. Stavridou. Book review: Abstraction and specification in program development, by Barbara Liskov and John Guttag. MIT Press and McGraw-Hill, London, 1986, 469 pages, Price £32.50. Science of Computer Programming, 10 (1):92–93, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642388900202.

Paterson:1988:BRB

[145] Mike Paterson. Book review: Algorithms: Their efficiency and complexity, by L. Kronsjö. Wiley, Chichester, 2nd ed., 1987, Price £24.95, ISBN 0-471-91201-8. Science of Computer Programming, 10 (1):93–95, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642388900214.

Hankin:1988:BRB

[146] Chris Hankin. Book review: Introduction to combinators and λ-calculus, by J. Roger Hindley and Jonathan P. Seldin. Mathematical Society Student Texts 1, Cambridge University Press, Cambridge, 1986, 360 pages, Price £8.95 paperback, ISBN 0-521-31839-4, £27.50 hardbound, ISBN 0-521-26896-6. Science of Computer Programming, 10(1):95–97, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642388900226.

Rem:1988:SPEa

[147] M. Rem. Small programming exercises 20. Science of Computer Programming, 10(1):99–105, February 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900238.

Tel:1988:DGM

[148] Gerard Tel, Richard B. Tan, and Jan van Leeuwen. The derivation of graph marking algorithms from distributed termination detection protocols. Science of Computer Programming, 10 (2):107–137, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nielson:1988:ABT

[149] Hanne R. Nielson and Flemming Nielson. Automatic binding time analysis for a typed lambda-calculus. Science of Computer Programming, 10 (2):139–176, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mason:1988:VPD

[150] Ian A. Mason. Verification of programs that destructively manipulate data. Science of Computer Programming, 10 (2):177–210, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Grabowski:1988:BRB

[151] Michal Grabowski. Book review: The foundations of program verification, by Jacques Loeckx and Kurt Sieber. Wiley, Chichester, 2nd ed., 1987, Price £22.50, ISBN 0-471-91282-4. Science of Computer Programming, 10 (2):211–212, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900275.

Wolczko:1988:BRB

[152] Mario Wolczko. Book review: Actors: a model of concurrent computation in distributed systems, by Gul Agha. MIT Press Series in AI, MIT Press, London, 1987, Price £24.95. Science of Computer Programming, 10(2):212–213, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900287.

Kitchenham:1988:BRB

[153] Barbara Kitchenham. Book review: Quality programming: Developing and testing software with statistical quality control, by Chin-Kuei Cho. Wiley, Chichester, 1987, Price £40.35, ISBN 0-471-84899-9. Science of Computer Programming, 10 (2):213–215, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900299.

Fisher:1988:BRB

[154] Michael Fisher. Book review: Temporal logic of programs, by Fred Kroger. EATCS Monographs on Theoretical Computer Science, Springer, Berlin, 1987, viii + 148 pages, Price DM 68.00, (hard cover). ISBN 3-540-17030-8. Science of Computer Programming, 10(2):215–216, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900305.

Capon:1988:BRB

[155] Peter C. Capon. Book review: Programming in Occam, by Geraint Jones. Prentice-Hall International, Hemel Hempstead, United Kingdom, 1987, 192 pages, Price £13.95 (paperback), ISBN 0-13-729773-4. Science of Computer Programming, 10 (2):217, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900317.

Gibbins:1988:BRB

[156] Peter Gibbins. Book review: Essential mathematics for software engineers, by G. Slater, Editor. Peter Peregrinus Ltd. on behalf of the Institution of Electrical Engineers, London, 1987, Price £75.00 (Four study pack volumes + video), ISBN 0-86341-095-2, 0-86341-096-0, 0-86341-097-9, 0-86341-098-7. Science of Computer Programming, 10(2):218–219, April 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900329.

Szymanski:1988:EHD

[157] Boleslaw K. Szymanski and Noah S. Prywes. Efficient handling of data structures in definitional languages. Science of Computer Programming, 10 (3):221–245, June 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Michel:1988:DTP

[158] Beat Michel. A “division” transformation for program and data structures and the structure clash problem. Science of Computer Programming, 10 (3):247–280, June 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Harrison:1988:LON

[159] Peter G. Harrison. Linearisation: an optimisation for nonlinear functional programs. Science of Computer Programming, 10(3):281–318, June 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1988:SPEb

[160] M. Rem. Small programming exercises 21. Science of Computer Programming, 10(3):319–325, June 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ryder:1988:CII

[161] B. G. Ryder, T. J. Marlowe, and M. C. Paull. Conditions for incremental iteration: examples and counterexamples. Science of Computer Programming, 11(1):1–15, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morgan:1988:PPA

[162] Carroll Morgan. Procedures, parameters, and abstraction: separate concerns. Science of Computer Programming, 11(1):17–28 (or 17–27??), October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mateti:1988:MTT

[163] Prabhaker Mateti and Ravi Manghirmalani. Morris’ tree traversal algorithm reconsidered. Science of Computer Programming, 11(1):29–43, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Berghammer:1988:TAS

[164] Rudolf Berghammer, Herbert Ehler, and Hans Zierer. Towards an algebraic specification of code generation. Science of Computer Programming, 11 (1):45–63, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Broy:1988:VQ

[165] Manfred Broy. Views of queues. Science of Computer Programming, 11 (1):65–86, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Walter:1988:BRB

[166] C. D. Walter. Book review: Algorithmics — The spirit of computing, by David Harel. Addison-Wesley, Reading, MA, 1987, Price £14.95, ISBN 0-201-19240-3. Science of Computer Programming, 11 (1):87–88, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900664.

Pratt:1988:BRB

[167] Ian Pratt. Book review: The little LISPer, by Daniel P. Friedman and Matthias Felleisen. MIT Press, Cambridge, MA, trade ed., 1987, Price £9.95. Science of Computer Programming, 11(1):88–90, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167642388900676.

Mearns:1988:BRB

[168] Ian Mearns. Book review: A review of Ada tasking, by Alan Burns, Andrew M. Lister and Andrew J. Wellings. Lecture Notes in Computer Science 262, Springer, Berlin, 1987, 141 pp., Price £10.00 (soft cover), ISBN 3-540-18008-

7. Science of Computer Programming, 11(1):90–91, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900688.

Ringwood:1988:BRB

[169] Graem Ringwood. Book review: Computability theory, semantics and logic programming, by Melvin Fitting. Oxford Logic Guide 13, Oxford University Press, Oxford, 1987, Price £25.00 (hard cover), ISBN 0-19-503691-3. Science of Computer Programming, 11 (1):91–94, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764238890069X.

Williams:1988:BRB

[170] Ifor Wyn Williams. Book review: The design and evaluation of a highperformance smalltalk system, by David Michael Ungar. MIT Press, Cambridge, MA, 1987, Price £26.95. Science of Computer Programming, 11 (1):94–95, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900706.

Milne:1988:BRB

[171] Robert Milne. Book review: Semantics of digital circuits, by Carlos Delgado Kloos. Lecture Notes in Computer Science 285, Springer, Berlin, 1987, 119 pages, Price DM 27.00, ISBN 3-540-18540-2. Science of Computer Programming, 11 (1):95–97, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900718.

Kuich:1988:BRB

[172] W. Kuich. Book review: Computability, by Klaus Weihrauch. EATCS Monographs on Theoretical Computer Science 9, Springer, Berlin, 1987, Price DM 98.00 (hard cover), ISBN 3-540-13721-1. Science of Computer Programming, 11 (1):97–98, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764238890072X.

Shepherdson:1988:BRB

[173] J. C. Shepherdson. Book review: Algorithmic information theory, by G. J. Chaitin. Cambridge University Press, Cambridge, 1987, Price £20.00, ISBN 0-521-34306-2. Science of Computer Programming, 11 (1):98–99, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642388900731.

Anonymous:1988:MIV

[174] Anonymous. Master index of volumes 1–10. Science of Computer Programming, 11(1):101–106, October 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schmidt:1988:DSB

[175] David A. Schmidt. Detecting stackbased environments in denotational definitions. Science of Computer Programming, 11(2):107–131, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schnoebelen:1988:RCP

[176] Ph. Schnoebelen. Refined compilation of pattern-matching for functional languages. Science of Computer Programming, 11(2):133–159, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Xue:1988:DLA

[177] Jin Yun Xue and David Gries. Developing a linear algorithm for cubing a cyclic permutation. Science of Computer Programming, 11(2): 161–165, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1988:SPEc

[178] M. Rem. Small programming exercises 22. Science of Computer Programming, 11(2):167–173, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mosses:1988:BRB

[179] Peter D. Mosses. Book review: Algebraic approaches, by E. G. Manes and M. A. Arbib. Springer, Berlin/ Heidelberg/New York/London/Paris/ Tokyo, 1986, 351 pages, Price DM 128.00. Science of Computer Programming, 11(2):175–176, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167642388900056.

Martin:1988:BRB

[180] Ursula Martin. Book review: Logic for computer science, by J. H. Gallier. Harper & Row, London, 1986, Price £17.95, ISBN 0-06-042225-4. Science of Computer Programming, 11(2): 176–178, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900068.

Nipkow:1988:BRB

[181] Tobias Nipkow. Book review: Logic and computation-Interactive proof with Cambridge LCF, by L. C. Paulson. Cambridge University Press, Cambridge, 1987, Price £27.50, ISBN 0-521-34632-0. Science of Computer Programming, 11(2):178– 180, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764238890007X.

Jones:1988:BRB

[182] Roger Bishop Jones. Book review: Understanding Z — a specification language and its formal semantics, by J. M. Spivey. Cambridge University Press, Cambridge, 1988, Price £17.50, ISBN 0-521-33429-2. Science of Computer Programming, 11 (2):180–184, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642388900081.

Brookes:1988:BRB

[183] Steve Brookes. Book review: Denotational semantics, by D. A. Schmidt. Prentice-Hall, London, 1986, Price £15.95, ISBN 0-205-10450-9. Science of Computer Programming, 11 (2):185–191, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642388900093.

Anonymous:1988:BRB

[184] Anonymous. Book review: Brains, machines and mathematics, by M. A. Arbib. Springer, Berlin/Heidelberg/ New York/London/Paris/Tokyo, 1987, Price DM 55.00, ISBN 3-540-96539-4. Science of Computer Programming, 11 (2):191–195, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764238890010X.

Stolen:1988:BRB

[185] Ketil Stølen. Book review: Concurrent program structures, by D.

Bustard, J. Elder and J. Welsh. Prentice-Hall, London, 1988, Price £15.95, ISBN 0-13-167080-8. Science of Computer Programming, 11 (2):195–196, December 1988. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642388900111.

Cai:1989:PDF

[186] Jiazhen Cai and Robert Paige. Program derivation by fixed point computation. Science of Computer Programming, 11(3):197–261, April 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Partsch:1989:IRR

[187] H. Partsch. From informal requirements to a running program: a case study in algebraic specification and transformational programming. Science of Computer Programming, 11 (3):263–297, April 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bidoit:1989:HMA

[188] M. Bidoit, M.-C. Gaudel, and A. Mauboussin. How to make algebraic specifications more understandable: an experiment with the PLUSS specification language. Science of Computer Programming, 12(1):1–38, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hall:1989:GFV

[189] Cordelia V. Hall and David S. Wise. Generating function versions with rational strictness patterns. Science of Computer Programming, 12(1):39–74, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1989:SPE

[190] M. Rem. Small programming exercises 23. Science of Computer Programming, 12(1):75–81, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wolczko:1989:BRB

[191] Mario Wolczko. Book review: The semantics of destructive Lisp, CSLI lecture notes 5. By Ian A. Mason. The University of Chicago Press, London, 1986, price £23.95 (£35.95 in Europe), ISBN 0-937073-05-9 (hardback); price £11.95 (£17.95 in Europe), ISBN 0-937073-06-7 (paperback). Science of Computer Programming, 12(1):83–84, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900294.

Capon:1989:BRB

[192] Peter C. Capon. Book review: Programming in Occam 2, by Geraint Jones and Michael Goldsmith. Prentice-Hall, Hemel Hempstead, United Kingdom, price £15.95 (paperback), ISBN 0-13-730334-3. Science of Computer Programming, 12(1):84, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167642389900312.

Lovengreen:1989:BRB

[193] Hans Henrik Løvengreen. Book review: Temporal logics and their applications, by Antony Galton, ed. Harcourt Brace Jovanovich, Sidcup, United Kingdom, 1987, xii + 244 pages, price £22.00 (hardcover), ISBN 0-12-274060-2. Science of Computer Programming, 12(1):85–87, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900324.

Mitchell:1989:BRB

[194] R. J. Mitchell. Book review: Data abstraction and program development using Pascal, by R. F. Hille. Prentice-Hall, Hemel Hempstead, United Kingdom, 1988, price £14.95, ISBN 0-13-7248-0301-7. Science of Computer Programming, 12 (1):87–88, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900336.

America:1989:BRB

[195] Pierre America. Book review: Objectoriented software construction, by Bertrand Meyer. Prentice-Hall, Hemel Hempstead, United Kingdom, 1988, price £19.95, ISBN 0-13-629031-0. Science of Computer Programming, 12(1):88–90, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900348.

Wills:1989:BRB

[196] Alan Wills. Book review: Algorithm animation, by Marc H. Brown. MIT Press, London, 1988, price £26.95 (hardback), ISBN 0-262-02278-8. Science of Computer Programming, 12(1):90–92, June 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764238990035X.

Bird:1989:FDP

[197] R. S. Bird, J. Gibbons, and G. Jones. Formal derivation of a pattern matching algorithm. Science of Computer Programming, 12(2):93–104, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

KrusemanAretz:1989:NAE

[198] F. E. J. Kruseman Aretz. A new approach to Earley’s parsing algorithm. Science of Computer Programming, 12 (2):105–121, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nipkow:1989:ERI

[199] Tobias Nipkow. Equational reasoning in Isabelle. Science of Computer Programming, 12(2):123–149, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gries:1989:ATR

[200] David Gries, Alain J. Martin, Jan L. A. van de Snepscheut, and Jan Tijmen Udding. An algorithm for transitive reduction of an acyclic graph. Science of Computer Programming, 12 (2):151–155, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wallis:1989:BRB

[201] Peter Wallis. Book review: Algorithms and data structures, by N.

Wirth. Prentice-Hall International, Hemel Hempstead, United Kingdom, 1986, Price £15.95/$27.95 (paperback), ISBN 0-13-021999-1. Science of Computer Programming, 12 (2):157–158, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900403.

Bailes:1989:BRB

[202] Paul A. Bailes. Book review: Introduction to functional programming, by R. Bird and P. Wadler. Prentice-Hall International, Hemel Hempstead, United Kingdom, 1988, Price £15.95, ISBN 0-13-484197-2. Science of Computer Programming, 12 (2):158–164, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900415.

Harrison:1989:BRB

[203] Rachel Harrison. Book review: Compiling functional languages, by A. Diller. Wiley, Chichester, United Kingdom, 1988, Price £15.95 (paperback). ISBN 0-471-920274. Science of Computer Programming, 12 (2):164–165, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900427.

Mili:1989:BRB

[204] Ali Mili. Book review: A method of programming, by Edsger W. Dijkstra and W. H. J. Feijen. Addison-Wesley, Wokingham, United Kingdom, 1988, Price £15.95 (paperback), ISBN 0-201-17536-3. Science of Computer Programming, 12 (2):165–167, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900439.

Weijland:1989:BRB

[205] W. Peter Weijland. Book review: Algebraic theory of processes, by M. Hennessy. MIT Press, Cambridge, MA, 1988, Price £16.95 (paperback), ISBN 0-262-08171-7. Science of Computer Programming, 12 (2):167–168, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900440.

Hall:1989:BRB

[206] Anthony Hall. Book review: Software engineering mathematics, by Jim Woodcock and Martin Loomes. Pitman Publishing, Southport, United Kingdom, 1988, Price £10.95 (paperback), ISBN 0-273-02673-9. Science of Computer Programming, 12 (2):168–170, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900452.

Jackson:1989:BRB

[207] Michael Jackson. Book review: Software productivity, by Harlan D. Mills; Foreword by Gerald M. Weinberg. Dorset House, New York, 1988, distributed by Wiley, Chichester, United Kingdom, Price £19.85 (paperback), ISBN 0-932633-10-2. Science of Computer Programming, 12 (2):170–172, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900464.

Lindsey:1989:BRB

[208] C. H. Lindsey. Book review: Comparative programming languages, by L. B. Wilson and R. G. Clark. Addison-Wesley, Wokingham, United Kingdom, 1988, Price £16.95 (paperback), ISBN 0-201-18483-4. Science of Computer Programming, 12 (2):172–173, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900476.

Turski:1989:BRB

[209] W. M. Turski. Book review: Software development: Fashioning the boroque, by Darrel Ince. Oxford University Press, Oxford, United Kingdom, 1988, Price £7.95, ISBN 0-19-853757-1 (paperback); Price £17.50, ISBN 0-19-853757-3 (hardback). Science of Computer Programming, 12 (2):174–175, July 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900488.

vanderWoude:1989:PPS

[210] Jaap van der Woude. Playing with patterns, searching for strings. Science of Computer Programming, 12(3): 177–190, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bijlsma:1989:CP

[211] A. Bijlsma. Calculating with pointers. Science of Computer Programming, 12 (3):191–206 (or 191–205??), September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Blikle:1989:DE

[212] Andrzej Blikle. Denotational engineering. Science of Computer Programming, 12(3):207–253, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hogger:1989:BRB

[213] Christopher Hogger. Book review: Logic with prolog, by Peter Gibbins. Clarendon Press, Oxford, United Kingdom, 1988, Price £15.00 (paperback), ISBN 0-19-859659-6. Science of Computer Programming, 12(3): 255–257, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764238990004X.

Pepper:1989:BRB

[214] P. Pepper. Book review: Foundations of programming, by J. Arsac. Academic Press, London, United Kingdom, 1985, Price £19.50/$32.50, ISBN 0-12-064460-6. Science of Computer Programming, 12(3):257– 259, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900051.

Shawe-Taylor:1989:BRB

[215] John Shawe-Taylor. Book review: Cryptography: an introduction to computer security, by Jennifer Seberry and Josef Pieprzyk. Prentice-Hall International, Hemel Hempstead, United Kingdom, 1988, Price £17.95 (paperback), ISBN 0-7248-0274-6. Science of Computer Programming, 12(3): 259–260, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900063.

Monahan:1989:BRB

[216] Brian Monahan. Book review: Metasoft primer: Towards a metalanguage for applied denotational semantics, Lecture notes in Computer Science 288. By Andrzej Blikle. Springer, Berlin, F.R.G., 1987, Price DM 31.50 (paperback), ISBN 0-3-540-18657-3. Science of Computer Programming, 12(3): 260–264, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900075.

Lehman:1989:BRB

[217] M. M. Lehman. Book review: Software engineering: Planning for change, by A. Lamb. Prentice-Hall International, Hemel Hempstead, United Kingdom, 1988, Price £13.95 (paperback), ISBN 0-13-823279-2. Science of Computer Programming, 12(3): 264–265, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900087.

Kirkham:1989:BRB

[218] C. C. Kirkham. Book review: Parallel processing: the transputer and occam, by A. Carling. Wiley, Bognor Regis, West Sussex, United Kingdom, 1988, Price £12.95, ISBN 1-85058-077-4. Science of Computer Programming, 12(3): 265–266, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900099.

Horning:1989:BRB

[219] J. J. Horning. Book review: Essays in computing science, by C. A. R. Hoare and C. B. Jones, ed., Prentice-Hall International, Hemel Hempstead, United Kingdom, 1989, Price £32.95 (hardback), ISBN 0-13-284027-8. Science of Computer Programming, 12(3): 267–270, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900105.

Anonymous:1989:AI

[220] Anonymous. Author index. Science of Computer Programming, 12 (3):271, September 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642389900117.

Andrews:1989:MSS

[221] Gregory R. Andrews. A method for solving synchronization problems. Science of Computer Programming, 13 (1):1–21, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Best:1989:SI

[222] Eike Best and Christian Lengauer. Semantic independence. Science of Computer Programming, 13(1):23–50, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hemerik:1989:BTA

[223] C. Hemerik and J. P. Katoen. Bottomup tree acceptors. Science of Computer Programming, 13(1):51–72, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tennent:1989:EDS

[224] R. D. Tennent. Elementary data structures in ALGOL-like languages. Science of Computer Programming, 13 (1):73–110, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Arnold:1989:ESP

[225] André Arnold. An example of sequentialization of a parallel algorithm. Science of Computer Programming, 13 (1):111–116, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sufrin:1989:BRB

[226] Bernard Sufrin. Book review: Principles of Computer Programming: a Mathematical Approach, by Harlan D. Mills, Victor R. Basili, John D. Gannon and Richard G. Hamlet. Brown Publishers, Oxford, 1988, Price £20.75 (paperback), ISBN 0-697-06841-2. Science of Computer Programming, 13(1): 117–119, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764238990018X.

Sokolowski:1989:BRB

[227] Stefan Sokolowski. Book review: Programming Language Theory and Its Implementation, by Michael J. C. Gordon. Prentice-Hall International, Hemel Hempstead, UK, 1988, Price £15.95 (hardcover), ISBN 0-13-730409. Science of Computer Programming, 13 (1):119–122, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900191.

Grosvenor:1989:BRB

[228] Dave Grosvenor. Book review: An Introduction to Discrete Mathematics and Formal System Specification, by D. C. Ince. Clarendon Press, Oxford, 1988, Price £15.00, ISBN 0-19-859664-2 (paperback); Price £30.00, ISBN 0-19-859667-7 (hardback). Science of Computer Programming, 13 (1):122–125, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900208.

Kay:1989:BRB

[229] S. Kay. Book review: Structured Systems Analysis through Prolog, by Terry Goble. Prentice-Hall International, Hemel Hempstead, UK, 1989, Price £13.95 (paperback), ISBN 0-13-853581-7. Science of Computer Programming, 13(1):125– 127, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764238990021X.

Wilson:1989:BRB

[230] Ian R. Wilson. Book review: Pascal Precisely, 2nd ed. By Judy Bishop. Addison-Wesley, Wokingham, UK, 1989, Price £12.95 (paperback), ISBN 0-201-41633-6. Science of Computer Programming, 13 (1):127–128, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900221.

Clement:1989:BRB

[231] T. Clement. Book review: Foundations of Programming Languages, J. Loecks, K. Mehlhorn and R. Wilhelm. Wiley, Chichester, UK, 1988, Price £24.50 (paperback), ISBN 0-471-92139-4. Science of Computer Programming, 13(1): 128–130, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900233.

Anonymous:1989:CPL

[232] Anonymous. Call for papers: LICS fifth annual IEEE symposium on logic in computer science June 4–7, 1990, Philadelphia, PA, USA. Science of Computer Programming, 13 (1):131–132, December 1989. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642389900245.

Back:1990:SRP

[233] R. J. R. Back and K. Sere. Stepwise refinement of parallel algorithms. Science of Computer Programming, 13 (2–3):133–180, May 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Thompson:1990:LFP

[234] Simon Thompson. Lawful functions and program verification in Miranda. Science of Computer Programming, 13 (2–3):181–218, May 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fernandez:1990:IEA

[235] Jean-Claude Fernandez. An implementation of an efficient algorithm for bisimulation equivalence. Science of Computer Programming, 13 (2–3):219–236, May 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

deVink:1990:CSP

[236] E. P. de Vink. Comparative semantics for PROLOG with cut. Science of Computer Programming, 13 (2–3):237–264, May 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Marcelis:1990:CAE

[237] A. J. J. M. Marcelis. On the classification of attribute evaluation algorithms. Science of Computer Programming, 14(1):1–24, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Spivey:1990:FTE

[238] Mike Spivey. A functional theory of exceptions. Science of Computer Programming, 14(1):25–42, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schoett:1990:BCD

[239] Oliver Schoett. Behavioural correctness of data representations. Science of Computer Programming, 14(1):43–57, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dybjer:1990:CIE

[240] Peter Dybjer. Comparing integrated and external logics of functional programs. Science of Computer Programming, 14(1):59–79, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Baker-Finch:1990:ACT

[241] Clement A. Baker-Finch. The (algebraic) congruence of two programming language definitions. Science of Computer Programming, 14(1):81–96, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rem:1990:SPE

[242] M. Rem. Small programming exercises 24. Science of Computer Programming, 14(1):97–101, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Davenport:1990:BRB

[243] J. H. Davenport. Book review: Geometric reasoning, by D. Kapur and J. L. Mundy, eds. MIT Press, London, United Kingdom, 1989, Price £24.75 (paperback), ISBN 0-262-6-1058-2. Science of Computer Programming, 14 (1):103–105, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090061H.

Ince:1990:BRB

[244] D. C. Ince. Book review: Software design and prototyping using me too, by H. Alexander and V. Jones. Prentice-Hall, Hemel Hempstead, United Kingdom, 1989, Price £17.95 (paperback), ISBN 0-13-820259-1. Science of Computer Programming, 14 (1):105–106, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090062I.

Feijs:1990:BRB

[245] L. M. G. Feijs. Book review: The interface description language: definition and use, by Richard Snodgrass Freeman, Oxford, United Kingdom, 1989, Price £38.95 (hardback), ISBN 0-7167-8198-0. Science of Computer Programming, 14 (1):106–109, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090063J.

Jackson:1990:BRB

[246] M. Jackson. Book review: Objectoriented analysis, by Peter Coad and Edward Yourdon. Prentice-Hall, Hemel Hempstead, United Kingdom, 1989, Price £31.90 (hardback), ISBN 0-13-629122-8. Science of Computer Programming, 14 (1):109–110, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090064K.

Wood:1990:BRBa

[247] M. McGee Wood. Book review: Natural language processing in LISP, by G. Gazdar and C. Mellish. Addison-Wesley, London, United Kingdom, 1989, Price £17.95 (hardback). ISBN 0-201-17825-7. Science of Computer Programming, 14(1):110, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090066M.

Wood:1990:BRBb

[248] M. McGee Wood. Book review: Natural language processing in Prolog, by G. Gazdar and C. Mellish. Addison-Wesley, London, United Kingdom, 1989, Price £16.95 (hardback), ISBN 0-201-18053-7. Science of Computer Programming, 14(1):110, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090067N.

Wood:1990:BRBc

[249] M. McGee Wood. Book review: Prolog and natural-language analysis, by F. C. N. Pereira and S. M. Shieber. The University of Chicago Press, London, United Kingdom, 1987, Price £23.25 (cloth), Price £11.25 (paperback), ISBN 0-937073-18-0. Science of Computer Programming, 14(1):110, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764239090065L.

Wood:1990:BRBd

[250] M. McGee Wood. Book review: Natural language processing in POP-11, by G. Gazdar and C. Mellish. Addison-Wesley, London, United Kingdom, 1989, Price £16.95 (hardback), ISBN 0-201-17448-0. Science of Computer Programming, 14 (1):110–112, June 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764239090068O.

vandeSnepscheut:1990:F

[251] Jan L. A. van de Snepscheut. Foreword. Science of Computer Programming, 14 (2–3):113–115, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642390900167.

Chandy:1990:RAC

[252] K. Mani Chandy. Reasoning about continuous systems. Science of Computer Programming, 14(2–3):117–132, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hehner:1990:PTP

[253] Eric C. R. Hehner. A practical theory of programming. Science of Computer Programming, 14(2–3):133–158, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Misra:1990:SCO

[254] Jayadev Misra. Specifying concurrent objects as communicating processes. Science of Computer Programming, 14 (2–3):159–184, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gribomont:1990:SRC

[255] E. Pascal Gribomont. Stepwise refinement and concurrency: The finite-state case. Science of Computer Programming, 14(2–3):185–228, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kaldewaij:1990:DSC

[256] Anne Kaldewaij and Martin Rem. The derivation of systolic computations. Science of Computer Programming, 14 (2–3):229–242, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kaldewaij:1990:SE

[257] Anne Kaldewaij and Berry Schoenmakers. Searching by elimination. Science of Computer Programming, 14(2– 3):243–254, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Malcolm:1990:DSP

[258] Grant Malcolm. Data structures and program transformation. Science of Computer Programming, 14(2– 3):255–279, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morgan:1990:TIR

[259] Carroll Morgan and Trevor Vickers. Types and invariants in the refinement calculus. Science of Computer Programming, 14(2–3):281–304, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Smith:1990:ATD

[260] Douglas R. Smith and Michael R. Lowry. Algorithm theories and design tactics. Science of Computer Programming, 14(2–3):305–321, October 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chen:1990:PIM

[261] Wei Chen and Jan Tijmen Udding. Program inversion: More than fun. Science of Computer Programming, 15 (1):1–13, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Despland:1990:URT

[262] Annie Despland, Monique Mazaud, and Raymond Rakotozafy. Using rewriting techniques to produce code generators and proving them correct. Science of Computer Programming, 15 (1):15–54, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Banatre:1990:GMD

[263] Jean-Pierre Banâtre and Daniel Le Métayer. The GAMMA model and its discipline of programming. Science of Computer Programming, 15 (1):55–77, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vandenEijnde:1990:LBR

[264] J. P. H. W. van den Eijnde. Leftbottom and right-top segments. Science of Computer Programming, 15 (1):79–94, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bush:1990:BRB

[265] Vicky Bush. Book review: The design of data structures and algorithms, by J. J. van Amstel and J. A. A. M. Poirters. Prentice Hall, Hemel Hempstead, United Kingdom, 1989, Price £15.95 (paperback), ISBN 0-13-199936-2. Science of Computer Programming, 15 (1):95–97, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090046G.

Pratt:1990:BRBa

[266] Ian Pratt. Book review: The artificial intelligence debate: false starts, real foundations, by Stephen R. Graubard, ed. MIT Press, London, United Kingdom, 1988, Price £8.95 (paperback), ISBN 0-262-57074-2. Science of Computer Programming, 15(1):97, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090047H.

Milne:1990:BRBa

[267] Robert E. Milne. Book review: The formal description technique LOTOS, by P. H. J. van Eijk, C. A. Vissers and M. Diaz, eds. North-Holland, Amsterdam, Netherlands, 1989, Price $102.50 (hardback), ISBN 0-444-87267-1. Science of Computer Programming, 15(1):98, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090048I.

Milne:1990:BRBb

[268] Robert E. Milne. Book review: The formal description technique Estelle, J. P. Diaz, J. P. Ansart, J. P. Courtiat, P. Azema and V. Chari, eds. North-Holland, Amsterdam, Netherlands, 1989, Prince $102.50 (hardback), ISBN 0-444-87132-2. Science of Computer Programming, 15 (1):98–100, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090049J.

Pratt:1990:BRBb

[269] Ian Pratt. Book review: The emperor’s new mind, by Roger Penrose. Oxford University Press, Oxford, United Kingdom, 1989, Price £20.00 (hardback), ISBN 0-19-851973-7. Science of Computer Programming, 15(1): 100–111, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090050N.

Talia:1990:BRB

[270] Domenico Talia. Book review: Strand: New concepts in parallel programming, by Ian Foster and Stephen Taylor. Prentice Hall, Hemel Hempstead, United Kingdom, 1989, Price £32.10 (paperback), ISBN 0-13-850587-X. Science of Computer Programming, 15(1): 111–112, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090051E.

Reichel:1990:BRB

[271] H. Reichel. Book review: Fundamentals of algebraic specifications 2, module specifications and constraints, EATCS Monographs on Theoretical Computer Science 21. By H. Ehrig and B. Ma Springer, Heidelberg, 1990, Price £98.00 DM (hardback), ISBN 3-540-51799-5. Science of Computer Programming, 15(1): 113–114, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090052F.

Anonymous:1990:CPM

[272] Anonymous. Call for papers: MFCS’91 sixteenth international symposium on mathematical foundations of computer science September 9– 13, 1991, Warsaw, Poland. Science of Computer Programming, 15(1): 115–116, November 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239090053G.

Hofstee:1990:DS

[273] H. Peter Hofstee, Alain J. Martin, and Jan L. A. Van De Snepscheut. Distributed sorting. Science of Computer Programming, 15(2–3):119–133, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Benaini:1990:SNS

[274] A. Benaini, P. Quinton, Y. Robert, Y. Saouter, and B. Tourancheau. Synthesis of a new systolic architecture for the algebraic path problem. Science of Computer Programming, 15(2– 3):135–158, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Loyens:1990:TSP

[275] L. D. J. C. Loyens and J. G. G. Van De Vorst. Two small parallel programming exercises. Science of Computer Programming, 15(2–3):159–169, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kaldewaij:1990:SDA

[276] Anne Kaldewaij and Gerard Zwaan. A systolic design for acceptors of regular languages. Science of Computer Programming, 15(2–3):171–183, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Struik:1990:SDP

[277] P. Struik. A systematic design of a parallel program for Dirichlet convolution. Science of Computer Programming, 15(2–3):185–200, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ebergen:1990:DSP

[278] Jo. C. Ebergen and Rob R. Hoogerwoord. A derivation of a serial-parallel multiplier. Science of Computer Programming, 15(2–3):201–215, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Luk:1990:SLP

[279] Wayne Luk and Geoffrey Brown. A systolic LRU processor and its topdown development. Science of Computer Programming, 15(2–3):217–233, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kessels:1990:VPC

[280] Joep L. W. Kessels and Frits D. Schalij. VLSI programming for the compact disc player. Science of Computer Programming, 15(2–3):235–248, December 1990. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Montenyohl:1991:CSF

[281] Margaret Montenyohl and Mitchell Wand. Correctness of static flow analysis in continuation semantics. Science of Computer Programming, 16(1):1–18, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fokkinga:1991:ETP

[282] Maarten M. Fokkinga. An exercise in transformational programming: Backtracking and branch-and-bound. Science of Computer Programming, 16(1): 19–48, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Singh:1991:SCO

[283] Ambuj K. Singh. Specification of concurrent objects using auxiliary variables. Science of Computer Programming, 16(1):49–88, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wills:1991:BRB

[284] Alan Wills. Book review: Human computer interaction: Selected readings, Jenny Preece and Laurie Keller, eds., (Prentice-Hall/Open University, Hemel Hempstead, United Kingdom, 1990), Price £28.50 (paperback), ISBN 0-13-444910-X. Science of Computer Programming, 16 (1):89–90, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190024R.

Hopkins:1991:BRBa

[285] Trevor P. Hopkins. Book review: OSF/Motif style guide, Open software foundation, (Prentice-Hall, Hemel Hempstead, United Kingdom, 1990), Price £12.95 (paperback), ISBN 0-13-640491-X. Science of Computer Programming, 16(1):90, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190026T.

Hopkins:1991:BRBb

[286] Trevor P. Hopkins. Book review: OSF/Motif User’s Guide, Open Software Foundation, (Prentice-Hall, Hemel Hempstead, United Kingdom, 1990), Price £11.95 (paperback), ISBN 0-13-640509-6. Science of Computer Programming, 16(1):90, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190025S.

Hopkins:1991:BRBc

[287] Trevor P. Hopkins. Book review: OSF/Motif programmer’s guide, Open Software Foundation, (Prentice-Hall, Hemel Hempstead, United Kingdom, 1990), Price £25.50 (paperback), ISBN 0-13-640525-8. Science of Computer Programming, 16 (1):91–92, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190027U.

Walshe:1991:BRB

[288] Anne Walshe. Book review: The programming process: an introduction using VDM and Pascal: John T. Latham, Vicky J. Bush and Ian D. Cottam, (Addison-Wesley, Wokingham, United Kingdom, 1990), Price £16.95 (paperback), ISBN 0-201-41661-1. Science of Computer Programming, 16 (1):92–93, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190028V.

Hooman:1991:BRB

[289] Jozef Hooman. Book review: Temporal logic for real-time systems, Jonathan S. Ostroff, (Research Studies Press, distributed by Wiley, Chichester, United Kingdom, 1989), Price £32.25 (paperback), ISBN 0-86380-086-6. Science of Computer Programming, 16(1):93–95, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764239190029W.

Turski:1991:BRB

[290] W. M. Turski. Book review: Origins of programming: Discourses on methodology, Andrei P. Ershov, (Springer, Heidelberg, 1990), Price DM 178.00 (hardback), ISBN 3-540-97061-4 (Translated by Robert H. Silverman). Science of Computer Programming, 16(1):95–97, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642391900302.

Anonymous:1991:BRB

[291] Anonymous. Book review: The collected works of J. Richard Büchi: Saunders Mac Lane and Dirk Siefkes, eds., (Springer, New York, 1990), Price £42.00 (hardback), ISBN 0-387-97064-9. Science of Computer Programming, 16(1):97–100, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764239190031R.

Anonymous:1991:A

[292] Anonymous. Announcement. Science of Computer Programming, 16(1):101, July 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/016764239190032S.

Benveniste:1991:SPE

[293] Albert Benveniste, Paul Le Guernic, and Christian Jacquemot. Synchronous programming with events and relations. The SIGNAL language and its semantics. Science of Computer Programming, 16(2):103–149, September 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bondorf:1991:AAR

[294] Anders Bondorf and Olivier Danvy. Automatic autoprojection of recursive equations with global variables and abstract data types. Science of Computer Programming, 16(2):151–195, September 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Georghiou:1991:BRB

[295] Luke Georghiou. Book review: Alvey-Britain’s strategic computing initiative, Brian Oakley and Kenneth Owen (MIT Press, London, 1990), Price £24.95 (hardback), ISBN 0-262-15038-7. Science of Computer Programming, 16(2): 197–198, September 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190003G.

Launchbury:1991:BRB

[296] John Launchbury. Book review: An introduction to functional programming through lambda calculus, Greg Michaelson (Addison-Wesley, Wokingham, United Kingdom, 1990), Price £17.95 (paperback), ISBN 0-201-17812-5. Science of Computer Programming, 16(2): 198–200, September 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190004H.

Lindsey:1991:BRB

[297] C. H. Lindsey. Book review: Programming language concepts and paradigms, David A. Watt, (Prentice Hall, Hemel Hempstead, United Kingdom, 1990), Price £16.95 (paperback), ISBN 0-13-728866-2. Science of Computer Programming, 16(2): 200–201, September 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190005I.

Mendler:1991:BRB

[298] Nax Mendler. Book review: Programming in Martin-Löf’s type theory, Bengt Nordström, Kent Petersson and Jan M. Smith (Oxford University Press, Corby, United Kingdom, 1990), Price £25.00 (hardback), ISBN 0-19-853814-7. Science of Computer Programming, 16(2): 202–203, September 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190006J.

Froome:1991:BRB

[299] Peter Froome. Book review: Formal methods in standards: a report from the BCS working group, C. L. N. Ruggles, Ed. (Springer, Heidelberg, Germany, 1990), Price 75.00 DM (paperback), ISBN 3-540-19577-7. Science of Computer Programming, 16(2): 204–205, September 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239190007K.

Kneuper:1991:SES

[300] Ralf Kneuper. Symbolic execution. A semantic approach. Science of Computer Programming, 16(3):207– 249, October 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vogler:1991:FDC

[301] Heiko Vogler. Functional description of the contextual analysis in blockstructured programming languages: a case study of tree tranducers. Science of Computer Programming, 16 (3):251–275, October 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jones:1991:F

[302] N. D. Jones. Foreword. Science of Computer Programming, 17(1–3):1–??, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bondorf:1991:AAH

[303] Anders Bondorf. Automatic autoprojection of higher order recursive equations. Science of Computer Programming, 17(1–3):3–34, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Felleisen:1991:EPP

[304] Matthias Felleisen. On the expressive power of programming languages. Science of Computer Programming, 17 (1–3):35–75, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’90 (Copenhagen, 1990).

Heckmann:1991:PDC

[305] Reinhold Heckmann. Power domain constructions. Science of Computer Programming, 17(1–3):77–117, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’90 (Copenhagen, 1990).

Helmink:1991:RTT

[306] Leen Helmink. Resolution and type theory. Science of Computer Programming, 17(1–3):119–138, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Reps:1991:APP

[307] Thomas Reps. Algebraic properties of program integration. Science of Computer Programming, 17(1–3): 139–215, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’90 (Copenhagen, 1990).

Thatte:1991:TSI

[308] Satish Thatte. A type system for implicit scaling. Science of Computer Programming, 17(1–3):217–245, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’90 (Copenhagen, 1990).

Maassen:1992:PPD

[309] Andreas Maassen. Parallel programming with data structures and higher order functions. Science of Computer Programming, 18(1):1–38, January 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zantema:1992:LSP

[310] H. Zantema. Longest segment problems. Science of Computer Programming, 18(1):39–66, January 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mili:1992:HCW

[311] Fatma Mili and Ali Mili. Heuristics for constructing while loops. Science of Computer Programming, 18 (1):67–106, January 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Heisel:1992:FIG

[312] Maritta Heisel. Formalizing and implementing Gries’ program development method in dynamic logic. Science of Computer Programming, 18 (1):107–137, January 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Boiten:1992:IRF

[313] Eerke A. Boiten. Improving recursive functions by inverting the order of evaluation. Science of Computer Programming, 18(2):139–179, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Reade:1992:BTR

[314] C. M. P. Reade. Balanced trees with removals: an exercise in rewriting and proof. Science of Computer Programming, 18(2):181–204, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morgan:1992:BRB

[315] Carroll Morgan. Book review: Programming: the derivation of algorithms, Anne Kaldewaij, (Prentice-Hall, Hemel Hempstead, United Kingdom, 1990), ISBN 0-13-204108-1 (paperback). Science of Computer Programming, 18(2):205–207, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167642392900109.

Zemanek:1992:BRB

[316] Heinz Zemanek. Book review: A computer perspective: Background to the computer age-new edition, the office of Charles and Ray Eames, (Harvard University Press, Cambridge, MA, 1990), 175 pages, ISBN 0-674-15626-9 (paperback). Science of Computer Programming, 18 (2):207–209, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239290011Y.

Osterby:1992:BRB

[317] Tom Østerby. Book review: Inheritance hierarchies in knowledge representation and programming languages, Maurizio Lenzerini, Daniele Nardi and Maria Simi, eds., (Wiley, Chichester, United Kingdom, 1991) 310 + xv pages, Price £22.50, ISBN 0-471-92741-4. Science of Computer Programming, 18(2):209–213, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/016764239290012Z.

Szeredi:1992:BRB

[318] Péter Szeredi. Book review: Logic Programming: Proceedings of the 1990 North American Conference, Saumya Debray and Manuel Hermenegildo, eds., (MIT Press, Cambridge, MA, 1990), Price £49.50 ($74.25 in Europe), ISBN 0-262-54058-4. Science of Computer Programming, 18 (2):213–215, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642392900132.

Hansson:1992:BRB

[319] Åke Hansson. Book review: Graphbased proof procedures for Horn clauses, Stan Raatz, (Birkhäuser, Basel, Switzerland, 1990), Price SFr. 52.00/DM 62.00, ISBN 3-7643-3530-0, 0-8176-3530-0 (paperback). Science of Computer Programming, 18 (2):215–218, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642392900143.

Nielsen:1992:BRB

[320] Jacob Nielsen. Book review: Graphic design for electronic documents and user interfaces, Aaron Marcus, (Addison-Wesley/ACM Press, Reading, MA, 1992), xii + 266 pages, Price $31.25 (paperback), ISBN 0-201-54364-8. Science of Computer Programming, 18 (2):218–221, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642392900154.

Knuth:1992:BRB

[321] Elod Knuth. Book review: A unifying framework for structured analysis and design models, T. H. Tse, (Cambridge University Press, Cambridge, United Kingdom, 1991), Price £19.50/$34.50, ISBN 0-521-39196-2. Science of Computer Programming, 18(2):221–222, April 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642392900165.

Ebergen:1992:AES

[322] J. C. Ebergen. Arbiters: an exercise in specifying and decomposing asynchronously communicating components. Science of Computer Programming, 18(3):223–245, June 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bouajjani:1992:MSG

[323] A. Bouajjani, J.-C. Fernandez, N. Halbwachs, P. Raymond, and C. Ratel. Minimal state graph generation. Science of Computer Programming, 18(3):247– 269, June 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gardiner:1992:RAA

[324] P. H. B. Gardiner and P. K. Pandya. Reasoning algebraically about recursion. Science of Computer Programming, 18(3):271–280, June 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bird:1992:SU

[325] Richard S. Bird. The smallest upravel. Science of Computer Programming, 18 (3):281–292, June 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Knapp:1992:DCP

[326] Edgar Knapp. Derivation of concurrent programs: two examples. Science of Computer Programming, 19 (1):1–23, October 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Perrin:1992:CRP

[327] Guy-René Perrin and Jean-Pierre Finance. Communication relations: a paradigm for parallel program design. Science of Computer Programming, 19 (1):25–59, October 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Banerjee:1992:TSC

[328] Debasish Banerjee. A technique for solving a class of quadratic FP equations. Science of Computer Programming, 19(1):61–85, October 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Berry:1992:ESP

[329] Gerard Berry and Georges Gonthier. The ESTEREL synchronous programming language: design, semantics, implementation. Science of Computer Programming, 19(2):87–152, November 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jongejan:1992:CTR

[330] Jan H. Jongejan. Correctness of transformations on register transfers. Science of Computer Programming, 19(2): 153–169, November 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hesselink:1992:LPD

[331] Wim H. Hesselink. LR-parsing derived. Science of Computer Programming, 19(2):171–196, November 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anderson:1992:BAR

[332] James H. Anderson and Bojan Grošelj. Beyond atomic registers: bounded wait-free implementations of nontrivial objects. Science of Computer Programming, 19(3):197–237, December 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kamareddine:1992:SCR

[333] Fairouz Kamareddine. A system at the cross-roads of functional and logic programming. Science of Computer Programming, 19(3):239–279, December 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Barbuti:1992:BPT

[334] Roberto Barbuti and Roberto Giacobazzi. A bottom-up polymorphic type inference in logic programming. Science of Computer Programming, 19 (3):281–313, December 1992. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ghezzi:1993:GER

[335] C. Ghezzi and G.-C. Roman. Guest Editors’ remarks. Science of Computer Programming, 20(1–2):1–??, April 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dardenne:1993:GDR

[336] Anne Dardenne, Axel van Lamsweerde, and Stephen Fickas. Goal-directed requirements acquisition. Science of Computer Programming, 20(1–2):3–50, April 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ledru:1993:DRS

[337] Y. Ledru. Developing reactive systems in a VDM framework. Science of Computer Programming, 20 (1–2):51–71, April 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schobbens:1993:EAS

[338] Pierre-Yves Schobbens. Exceptions for algebraic specifications: on the meaning of “but”. Science of Computer Programming, 20(1–2):73–111, April 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Sixth International Workshop on Software Specification and Design (6IWSSD) (Como, 1991).

Kurki-Suonio:1993:RTS

[339] Reino Kurki-Suonio, Kari Systae, and Jueri Vain. Real-time specification and modeling with joint actions. Science of Computer Programming, 20 (1–2):113–140, April 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ciapessoni:1993:ETG

[340] E. Ciapessoni, E. Corsetti, A. Montanari, and P. San Pietro. Embedding time granularity in a logical specification language for synchronous real-time systems. Science of Computer Programming, 20(1–2):141–171, April 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Sixth International Workshop on Software Specification and Design (6IWSSD) (Como, 1991).

Vivares:1993:MJP

[341] F. Vivarès. Modelling Jackson’s programming method. Science of Computer Programming, 20(3):173– 204, June 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Skillicorn:1993:DPP

[342] D. B. Skillicorn. Deriving parallel programs from specifications using cost information. Science of Computer Programming, 20(3):205–221, June 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Back:1993:SIS

[343] R. J. R. Back and J. von Wright. Statement inversion and strongest postcondition. Science of Computer Programming, 20(3):223–251, June 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Desharnais:1993:MSD

[344] Jules Desharnais, Ali Mili, and Fatma Mili. On the mathematics of sequential decompositions. Science of Computer Programming, 20(3):253– 289, June 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1993:AIV

[345] Anonymous. Author index — volume 20. Science of Computer Programming, 20(3):291, June 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1993:MIV

[346] Anonymous. Master index — volumes 11–20. Science of Computer Programming, 20(3):293–299, June 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gribomont:1993:CTS

[347] E. Pascal Gribomont. Concurrency without toil: a systematic method for parallel program design. Science of Computer Programming, 21(1):1–56, August 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moller:1993:TPA

[348] Bernhard Möller. Towards pointer algebra. Science of Computer Programming, 21(1):57–90, August 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Apt:1993:F

[349] K. R. Apt and M. Hagiya. Foreword. Science of Computer Programming, 21 (2):91–??, October 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Abadi:1993:ADS

[350] M. Abadi, M. Burrows, C. Kaufman, and B. Lampson. Authentication and delegation with smart-cards. Science of Computer Programming, 21 (2):93–113, October 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Steffen:1993:GDF

[351] Bernhard Steffen. Generating data flow analysis algorithms from modal specifications. Science of Computer Programming, 21(2):115–139, October 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mitchell:1993:AEP

[352] John C. Mitchell. On abstraction and the expressive power of programming languages. Science of Computer Programming, 21(2):141–163, October 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Conference on Theoretical Aspects of Computer Software (TACS ’91) (Sendai, 1991).

Nishizaki:1993:PCL

[353] Shin ya Nishizaki. Programs with continuations and linear logic. Science of Computer Programming, 21 (2):165–190, October 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Conference on Theoretical Aspects of Computer Software (TACS ’91) (Sendai, 1991).

Singh:1993:RCS

[354] Ambuj K. Singh and Mohamed G. Gouda. Rankers: a classification of synchronization problems. Science of Computer Programming, 21(3): 191–223, December 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

DeFrancesco:1993:ACN

[355] Nicoletta De Francesco, Ugo Montanari, and Daniel Yankelevich. Axiomatizing CCS, nets and processes. Science of Computer Programming, 21(3): 225–261, December 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Consel:1993:SDG

[356] Charles Consel and Siau Cheng Khoo. Semantics-directed generation of a Prolog compiler. Science of Computer Programming, 21(3):263–291, December 1993. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morgan:1994:F

[357] C. Morgan. Foreword. Science of Computer Programming, 22(1–2):1–??, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Backhouse:1994:CPA

[358] Roland C. Backhouse, J. P. H. W. van den Eijnde, and A. J. M. van Gasteren. Calculating path algorithms. Science of Computer Programming, 22(1–2):3–19, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Gardiner:1994:ACP

[359] Paul H. B. Gardiner, Clare E. Martin, and Oege de Moor. An algebraic construction of predicate transformers. Science of Computer Programming, 22(1–2):21–44, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Hofstee:1994:DCS

[360] H. Peter Hofstee. Distributing a class of sequential programs. Science of Computer Programming, 22 (1–2):45–65, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Hoogendijk:1994:RPL

[361] Paul F. Hoogendijk and Roland C. Backhouse. Relational programming laws in the tree, list, bag, set hierarchy. Science of Computer Programming, 22(1–2):67–105, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Jones:1994:DAC

[362] Geraint Jones and Mary Sheeran. Designing arithmetic circuits by refinement in Ruby. Science of Computer Programming, 22(1–2):107–135, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Lukkien:1994:OSG

[363] Johan J. Lukkien. Operational semantics and generalized weakest preconditions. Science of Computer Programming, 22(1–2):137–155, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Moller:1994:SPG

[364] Bernhard Möller and Martin Russling. Shorter paths to graph algorithms. Science of Computer Programming, 22 (1–2):157–180, April 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction (Oxford, 1992).

Danvy:1994:BDS

[365] Olivier Danvy. Back to direct style. Science of Computer Programming, 22 (3):183–195, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Henglein:1994:DTS

[366] Fritz Henglein. Dynamic typing: syntax and proof theory. Science of Computer Programming, 22(3):197– 230, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Fourth European Symposium on Programming (Rennes, 1992).

Jones:1994:TQT

[367] Mark P. Jones. A theory of qualified types. Science of Computer Programming, 22(3):231–256, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Fourth European Symposium on Programming (Rennes, 1992).

Larcheveque:1994:ITP

[368] J. M. Larcheveque. Interprocedural type propagation for object-oriented languages. Science of Computer Programming, 22(3):257–282, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hankin:1994:AFP

[369] Chris Hankin and Sebastian Hunt. Approximate fixed points in abstract interpretation. Science of Computer Programming, 22(3):283–306, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See erratum [376].

Hughes:1994:RAI

[370] John Hughes and John Launchbury. Reversing abstract interpretations. Science of Computer Programming, 22 (3):307–326, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Fourth European Symposium on Programming (Rennes, 1992).

Nielson:1994:TPW

[371] Flemming Nielson and Hanne Riis Nielson. The tensor product in Wadler’s analysis of lists. Science of Computer Programming, 22(3):327– 354, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Fourth European Symposium on Programming (Rennes, 1992).

Gibbons:1994:EPA

[372] Jeremy Gibbons, Wentong Cai, and David B. Skillicorn. Efficient parallel algorithms for tree accumulations. Science of Computer Programming, 23 (1):1–18, October 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Palsberg:1994:STO

[373] Jens Palsberg and Michael I. Schwartzbach. Static typing for object-oriented programming. Science of Computer Programming, 23(1):19–53, October 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rondogiannis:1994:PNB

[374] P. Rondogiannis and M. H. M. Cheng. Petri-net-based deadlock analysis of Process Algebra programs. Science of Computer Programming, 23(1):55– 89, October 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Borger:1994:BRB

[375] Egon Börger. Book review: Predicate calculus and program semantics: E. W. Dijkstra and C. S. Scholten, (Springer, Berlin, 1989). Science of Computer Programming, 23 (1):91–101, October 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642394900027.

Hankin:1994:EAF

[376] Chris Hankin and Sebastian Hunt. Erratum to: Approximate fixed points in abstract interpretation. Science of Computer Programming, 23 (1):103, October 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642394000298. See

[369].

Gaudel:1994:F

[377] Marie-Claude Gaudel. Foreword. Science of Computer Programming, 23 (2–3):105, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642394000166.

Collette:1994:CAC

[378] Pierre Collette. Composition of assumption-commitment specifications in a UNITY style. Science of Computer Programming, 23(2–3):107–125, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Colloquium on Formal Approaches of Software Engineering (Orsay, 1993).

Giannotti:1994:GSL

[379] Fosca Giannotti and Diego Latella. Gate splitting in LOTOS specifications using abstract interpretation. Science of Computer Programming, 23(2–3): 127–149, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Colloquium on Formal Approaches of Software Engineering (Orsay, 1993).

Esparza:1994:MCU

[380] Javier Esparza. Model checking using net unfoldings. Science of Computer Programming, 23(2–3):151–195, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Colloquium on Formal Approaches of Software Engineering (Orsay, 1993).

Smith:1994:PTS

[381] Geoffrey S. Smith. Principal type schemes for functional programs with overloading and subtyping. Science of Computer Programming, 23(2–3): 197–226, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Colloquium on Formal Approaches of Software Engineering (Orsay, 1993).

Bernot:1994:LAE

[382] Gilles Bernot, Pascale Le Gall, and Marc Aiguier. Label algebras and exception handling. Science of Computer Programming, 23(2–3):227–286, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Selected papers of the Colloquium on Formal Approaches of Software Engineering (Orsay, 1993).

Bellegarde:1994:SFM

[383] Françoise Bellegarde and James Hook. Substitution: a formal methods case study using monads and transformations. Science of Computer Programming, 23(2–3):287–311, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Liu:1995:SDI

[384] Yanhong A. Liu and Tim Teitelbaum. Systematic derivation of incremental programs. Science of Computer Programming, 24(1):1–39, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lucena:1995:FDE

[385] C. J. P. Lucena and P. S. C. Alencar. A formal description of evolving software systems architectures. Science of Computer Programming, 24 (1):41–61, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanGilst:1995:NPT

[386] F. A. van Gilst and P. M. van den Broek. A new programming technique for lazy functional languages. Science of Computer Programming, 24 (1):63–81, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Harrison:1995:TPE

[387] Peter G. Harrison and R. Lyndon While. Transformation of polynomial evaluation to a pipeline via Horner’s rule. Science of Computer Programming, 24(1):83–95, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Marandjan:1995:BRB

[388] Hrant. B. Marandjan. Book review: Handbook of Theoretical Computer Science. Volume A: Algorithms and Complexity. Volume B: Formal Models and Semantics, Jan van Leeuwen, ed., (Elsevier, Amsterdam, Netherlands/The MIT Press, Cambridge, MA, 1990). Science of Computer Programming, 24 (1):97–99, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642395900098.

Levi:1995:BRB

[389] Giorgio Levi. Book review: Foundations of disjunctive logic programming, Jorge Lobo, Jack Minker and Arcot Rajasekar, (MIT Press, London, 1992), Price £60.75. Science of Computer Programming, 24 (1):99–100, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239590011X.

Schmidt:1995:BRB

[390] David Schmidt. Book review: Predicate transformer semantics, Ernest G. Manes, (Cambridge University Press, Cambridge, 1993), ISBN 0-521-42036-9. Science of Computer Programming, 24(1):100–102, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/0167642395900136.

Guessarian:1995:BRB

[391] Irene Guessarian. Book review: Intensional logics, L. Fariñas del Cerro, M. Penttonen, eds., (Clarendon Press, Oxford, 1992). Science of Computer Programming, 24 (1):102–103, February 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642395900152.

Stoller:1995:VPU

[392] Scott D. Stoller and Fred B. Schneider. Verifying programs that use causally-ordered message-passing. Science of Computer Programming, 24 (2):105–128, April 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chandy:1995:PTR

[393] K. Mani Chandy and Beverly A. Sanders. Predicate transformers for reasoning about concurrent computation. Science of Computer Programming, 24(2):129–148 (or 129–147??), April 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See erratum [471].

Kaldewaij:1995:MSD

[394] Anne Kaldewaij and Laurens de Vries. The matrix as in-situ data structure. Science of Computer Programming, 24 (2):149–158, April 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Scholefield:1995:PPR

[395] David Scholefield. Proving properties of real-time semaphores. Science of Computer Programming, 24 (2):159–181, April 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanEmdeBoas:1995:BRB

[396] Peter van Emde Boas. Book review: Compared to what? An introduction to the analysis of algorithms: Gregory J. E. Rawlins, (Freeman, New York, 1992), ISBN 0-7167-8243-X, xxvi + 536 pages. Science of Computer Programming, 24 (2):183–185, April 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642395900179.

Reichel:1995:BRB

[397] Horst Reichel. Book review: Formal specification and design, Cambridge tracts in theoretical computer science: L. M. G. Feijs and H. B. M Jonkers, (Cambridge University Press, 1992) (hardback), ISBN 0-521-43457-2. Science of Computer Programming, 24(2):185–186, April 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642395900195.

Moggi:1995:BRB

[398] E. Moggi. Book review: Two-level functional languages, Cambridge tracts in computer science, vol. 34: F. Nielson and H. R. Nielson, (Cambridge University Press, 1992). Science of Computer Programming, 24 (2):186–187, April 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0167642395900217.

Bloom:1995:TDI

[399] Bard Bloom and Robert Paige. Transformational design and implementation of a new efficient solution to the ready simulation problem. Science of Computer Programming, 24 (3):189–220, June 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Banach:1995:RSD

[400] R. Banach. On regularity in software design. Science of Computer Programming, 24(3):221–248, June 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Borger:1995:MDF

[401] Egon Börger and Dean Rosenzweig. A mathematical definition of full Prolog. Science of Computer Programming, 24 (3):249–286, June 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1995:AIVa

[402] Anonymous. Author index volume 24 (1995) Committee Draft 1.0, ISO/IEC JTC1 SC22 WG17 N.92. Science of Computer Programming, 24(3):287, June 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/0167642395900055.

Moir:1995:WFA

[403] Mark Moir and James H. Anderson. Wait-free algorithms for fast, long-lived renaming. Science of Computer Programming, 25(1):1–39, October 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Parnas:1995:FDC

[404] David Lorge Parnas and Jan Madey. Functional documents for computer systems. Science of Computer Programming, 25(1):41–61, October 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Madhukar:1995:SDA

[405] K. Madhukar, D. Pavan Kumar, C. Pandu Rangan, and R. Sundar. Systematic design of an algorithm for biconnected components. Science of Computer Programming, 25(1):63–77, October 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sannella:1995:F

[406] Donald Sannella. Foreword. Science of Computer Programming, 25(2– 3):79–80, December 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/016764239590025X.

Abadi:1995:TPO

[407] Martín Abadi and Luca Cardelli. A theory of primitive objects. Secondorder systems. Science of Computer Programming, 25(2–3):81–116, December 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’94 (Edinburgh, 1994).

Andrews:1995:FII

[408] James H. Andrews. Foundational issues in implementing constraint logic programming systems. Science of Computer Programming, 25(2–3):117–147, December 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’94 (Edinburgh, 1994).

Bidoit:1995:BAS

[409] Michel Bidoit, Rolf Hennicker, and Martin Wirsing. Behavioural and abstractor specifications. Science of Computer Programming, 25(2–3):149–186, December 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’94 (Edinburgh, 1994).

Giegerich:1995:CIP

[410] Robert Giegerich and Stefan Kurtz. A comparison of imperative and purely functional suffix tree constructions. Science of Computer Programming, 25 (2–3):187–218, December 1995. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’94 (Edinburgh, 1994).

Hankin:1995:LTI

[411] Chris Hankin and Daniel Le Métayer. Lazy type inference and program analysis. Science of Computer Programming, 25(2–3):219–249, December 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’94 (Edinburgh, 1994).

Jay:1995:SS

[412] C. Barry Jay. A semantics for shape. Science of Computer Programming, 25 (2–3):251–283, December 1995. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

                                        ftp://ftp.socs.uts.edu.au/Users/
                                        cbj/shape_semantics.ps.Z. ESOP

’94 (Edinburgh, 1994).

Prasad:1995:CBS

[413] K. V. S. Prasad. A calculus of broadcasting systems. Science of Computer Programming, 25(2–3):285–327, December 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). ESOP ’94 (Edinburgh, 1994).

Anonymous:1995:AIVb

[414] Anonymous. Author index volume 25 (1995). Science of Computer Programming, 25(2–3):329, December 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/0167642395900276.

Sintzoff:1996:E

[415] Michel Sintzoff. Editorial. Science of Computer Programming, 26 (1–3):1–2, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642396900027.

Moller:1996:P

[416] Bernhard Möller. Preface. Science of Computer Programming, 26(1–3):3, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642396900040.

Backhouse:1996:MPC

[417] Roland Backhouse. Mathematics of program construction (editorial). Science of Computer Programming, 26(1– 3):5–9, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Turski:1996:WC

[418] Władysław M. Turski. I was a computer. Science of Computer Programming, 26(1–3):11–14, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Bird:1996:FAD

[419] Richard S. Bird. Functional algorithm design. Science of Computer Programming, 26(1–3):15–31, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Shankar:1996:STM

[420] Natarajan Shankar. Steps towards mechanizing program transformations using PVS. Science of Computer Programming, 26(1–3):33–57, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Achatz:1996:MPD

[421] Klaus Achatz and Wolfram Schulte. Massive parallelization of divide-and-conquer algorithms over powerlists. Science of Computer Programming, 26 (1–3):59–78, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Back:1996:SCA

[422] R. J. R. Back, A. J. Martin, and K. Sere. Specifying the Caltech asynchronous microprocessor. Science of Computer Programming, 26(1–3):79– 97, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Berghammer:1996:TDC

[423] Rudolf Berghammer and Burghard von Karger. Towards a design calculus for CSP. Science of Computer Programming, 26(1–3):99–115, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Russling:1996:DCL

[424] Martin Russling. Deriving a class of layer-oriented graph algorithms. Science of Computer Programming, 26 (1–3):117–132, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Apt:1996:ABQ

[425] Krzysztof R. Apt. Arrays, bounded quantification and iteration in logic and constraint logic programming. Science of Computer Programming, 26 (1–3):133–148, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Kaldewaij:1996:LT

[426] Anne Kaldewaij and Victor J. Dielissen. Leaf trees. Science of Computer Programming, 26(1–3):149–165, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Coquand:1996:ATC

[427] Thierry Coquand. An algorithm for type-checking dependent types. Science of Computer Programming, 26 (1–3):167–177, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Monin:1996:ECH

[428] Jean-François Monin. Exceptions considered harmless. Science of Computer Programming, 26(1–3):179–196, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Curtis:1996:PG

[429] Sharon Curtis and Gavin Lowe. Proofs with graphs. Science of Computer Programming, 26(1–3):197–216, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mathematics of program construction, Kloster Irsee, 1995.

Doornbos:1996:R

[430] Henk Doornbos and Roland Backhouse. Reductivity. Science of Computer Programming, 26(1–3):217–236, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Frappier:1996:RCP

[431] M. Frappier, A. Mili, and J. Desharnais. A relational calculus for program construction by parts. Science of Computer Programming, 26 (1–3):237–254, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Stolen:1996:RPS

[432] Ketil Stölen. Refinement principles supporting the transition from asynchronous unbounded to synchronous communication. Science of Computer Programming, 26(1–3):255–272, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Mathematics of program construction, Kloster Irsee, 1995.

Anonymous:1996:AIV

[433] Anonymous. Author index of volume 26 (1996). Science of Computer Programming, 26(1–3):273, May 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642396900064.

Nishimura:1996:IDV

[434] Susumu Nishimura. Integrating different versions of programs with backward and forward slices. Science of Computer Programming, 27(1):1–35, July 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Duggan:1996:ETI

[435] Dominic Duggan and Frederick Bent. Explaining type inference. Science of Computer Programming, 27(1):37–83, July 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Watson:1996:TSM

[436] B. W. Watson and G. Zwaan. A taxonomy of sublinear multiple keyword pattern matching algorithms. Science of Computer Programming, 27 (2):85–118, September 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Katoen:1996:SAR

[437] Joost-Pieter Katoen and Berry Schoenmakers. Systolic arrays for the recognition of permutation-invariant segments. Science of Computer Programming, 27(2):119–137, September 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Butler:1996:SRC

[438] Michael J. Butler. Stepwise refinement of communicating systems. Science of Computer Programming, 27(2): 139–173, September 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hayes:1996:SMR

[439] I. J. Hayes. Supporting module reuse in refinement. Science of Computer Programming, 27(2):175–184, September 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Segers:1996:DAC

[440] J. Segers and J. C. Ebergen. Design and analysis of up-down counters. Science of Computer Programming, 27(2):185–204, September 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bijlsma:1996:PFS

[441] A. Bijlsma and C. S. Scholten. Pointfree substitution. Science of Computer Programming, 27(2):205–214, September 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wada:1996:OPN

[442] Eiiti Wada and Akinori Yonezawa. Obituary: Professor Nobuo Yoneda (28 March 1930–22 April 1996). Science of Computer Programming, 27(3):215– 216, 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bossi:1996:DLP

[443] A. Bossi, M. Bugliesi, M. Gabbrielli, G. Levi, and M. C. Meo. Differential logic programs: Programming methodologies and semantics. Science of Computer Programming, 27(3): 217–262, November 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Frost:1996:MPF

[444] Richard A. Frost and Barbara Szydlowski. Memoizing purely functional top-down backtracking language processors. Science of Computer Programming, 27(3):263–288, November 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Seaman:1996:OSS

[445] Jill Seaman and S. Purushothaman Iyer. An operational semantics of sharing in lazy evaluation. Science of Computer Programming, 27(3): 289–322, November 1996. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fitzpatrick:1997:ATA

[446] Stephen Fitzpatrick, Terence J. Harmer, Alan Stewart, Maurice Clint, and James M. Boyle. The automated transformation of abstract specifications of numerical algorithms into efficient array processor implementations. Science of Computer Programming, 28(1):1–41, January 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schneider:1997:TRC

[447] Steve Schneider. Timewise refinement for communicating processes. Science of Computer Programming, 28 (1):43–90, January 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Arnold:1997:ESE

[448] Andre Arnold, Didier Begay, and Jean-Pierre Radoux. The embedded software of an electricity meter: An experience in using formal methods in an industrial project. Science of Computer Programming, 28(2–3):93–110, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fiadeiro:1997:CSP

[449] Jose Luiz Fiadeiro and Tom Maibaum. Categorical semantics of parallel program design. Science of Computer Programming, 28(2–3):111–138, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Iglewski:1997:IDM

[450] Michal Iglewski and Janina Mincer-Daszkiewicz. Internal design of modules specified in the trace assertion method. Science of Computer Programming, 28(2–3):139–170, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

He:1997:PMG

[451] Jifeng He, K. Seidel, and A. McIver. Probabilistic models for the guarded command language. Science of Computer Programming, 28(2–3):171–192, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Formal specifications: foundations, methods, tools and applications (Konstancin, 1995).

Kluzniak:1997:SLL

[452] Feliks Kluzniak and Miroslawa Milkowska. Spill — a logic language for writing testable requirements specifications. Science of Computer Programming, 28 (2–3):193–223, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kosiuczenko:1997:TRL

[453] Piotr Kosiuczenko and Martin Wirsing. Timed rewriting logic with an application to object-based specification. Science of Computer Programming, 28 (2–3):225–246, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Formal specifications: foundations, methods, tools and applications (Konstancin, 1995).

Pepper:1997:HLD

[454] Peter Pepper and Douglas R. Smith. A high-level derivation of global search algorithms with constraint propagation. Science of Computer Programming, 28 (2–3):247–271, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wing:1997:CSM

[455] Jeannette M. Wing and Mandana Vaziri-Farahani. A case study in model checking software systems. Science of Computer Programming, 28 (2–3):273–299, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Groote:1997:F

[456] Jan Friso Groote and Martin Rem. Foreword. Science of Computer Programming, 29(1–2):1–2, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642396000263.

Bruns:1997:IAM

[457] Glenn Bruns. An industrial application of modal process logic. Science of Computer Programming, 29(1–2):3–22, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kapus:1997:VXC

[458] Tatjana Kapus and Zmago Brezocnik. Verification of XTP context management closing procedure in style of TLA. Science of Computer Programming, 29(1–2):23–52, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bicarregui:1997:MMF

[459] Juan Bicarregui, Jeremy Dick, Brian Matthews, and Eoin Woods. Making the most of formal specification through animation, testing and proof. Science of Computer Programming, 29 (1–2):53–78, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Campos:1997:STF

[460] Sergio Campos, Edmund M. Clarke, and Marius Minea. Symbolic techniques for formally verifying industrial systems. Science of Computer Programming, 29(1–2):79–98, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Elseaidy:1997:MVA

[461] Wael M. Elseaidy, Rance Cleaveland, and John W. Baugh, Jr. Modeling and verifying active structural control systems. Science of Computer Programming, 29(1–2):99–122, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fernandez:1997:EAG

[462] Jean-Claude Fernandez, Claude Jard, Thierry Jeron, and Cesar Viho. Experiment in automatic generation of test suites for protocols with verification technology. Science of Computer Programming, 29(1–2):123–146, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morley:1997:SLC

[463] Matthew J. Morley. Safety-level communication in railway interlockings. Science of Computer Programming, 29 (1–2):147–170, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Garavel:1997:SVV

[464] Hubert Garavel and Laurent Mounier. Specification and verification of various distributed leader election algorithms for unidirectional ring networks. Science of Computer Programming, 29 (1–2):171–197, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bergstra:1997:GPB

[465] Jan A. Bergstra, Joris A. Hillebrand, and Alban Ponse. Grid protocols based on synchronous communication. Science of Computer Programming, 29 (1–2):199–233, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bos:1997:BDT

[466] S. H. J. Bos and M. A. Reniers. I^{2}Cbus in discrete-time process algebra. Science of Computer Programming, 29 (1–2):235–258, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Korver:1997:DFI

[467] Henri Korver. Detecting feature interactions with CAESAR/ALDEBARAN. Science of Computer Programming, 29 (1–2):259–278, July 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nickson:1997:SCP

[468] Ray Nickson and Ian Hayes. Supporting contexts in program refinement. Science of Computer Programming, 29 (3):279–302, September 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Palsberg:1997:NAC

[469] Jens Palsberg, Boaz Patt-Shamir, and Karl Lieberherr. A new approach to compiling adaptive programs. Science of Computer Programming, 29(3): 303–326, September 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanderSommen:1997:PME

[470] F. W. van der Sommen, W. H. J. Feijen, and A. J. M. van Gasteren. Peterson’s mutual exclusion algorithm revisited. Science of Computer Programming, 29(3):327–334, September 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chandy:1997:EPT

[471] K. Mani Chandy and Beverly A. Sanders. Erratum: “Predicate transformers for reasoning about concurrent computation” [Sci. Comput. Programming 24 (1995), no. 2, 129–148; MR 96d:68145]. Science of Computer Programming, 29(3):335, September 1997.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See [393].

Codognet:1998:F

[472] Philippe Codognet. Foreword. Science of Computer Programming, 30 (1–2):1–2, January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642397000051.

Gupta:1998:CCC

[473] V. Gupta, R. Jagadeesan, and V. A. Saraswat. Computing with continuous change. Science of Computer Programming, 30(1–2):3–49, January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Concurrent constraint programming (Venice, 1995).

Bueno:1998:POC

[474] F. Bueno, M. Hermenegildo, U. Montanari, and F. Rossi. Partial order and contextual net semantics for atomic and locally atomic CC programs. Science of Computer Programming, 30(1–2):51–82, January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Concurrent constraint programming (Venice, 1995).

VanHentenryck:1998:NCP

[475] Pascal Van Hentenryck, Laurent Michel, and Frederic Benhamou. Newton — constraint programming over nonlinear constraints. Science of Computer Programming, 30(1–2):83–118, January 1998. CODEN SCPGD4.

ISSN 0167-6423 (print), 1872-7964 (electronic).

Bagnara:1998:HCS

[476] Roberto Bagnara. A hierarchy of constraint systems for data-flow analysis of constraint logic-based languages. Science of Computer Programming, 30(1–2):119–155, January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Concurrent constraint programming (Venice, 1995).

Chakravarty:1998:GHO

[477] Manuel M. T. Chakravarty, Yike Guo, Martin Köhler, and Hendrik C. R. Lock. GOFFIN: Higher-order functions meet concurrent constraints. Science of Computer Programming, 30(1–2):157– 199, January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Concurrent constraint programming (Venice, 1995).

Borghoff:1998:CBP

[478] Uwe M. Borghoff, Remo Pareschi, Francesca Arcelli, and Ferrante Formato. Constraint-based protocols for distributed problem solving. Science of Computer Programming, 30 (1–2):201–226 (or 201–225??), January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Concurrent constraint programming (Venice, 1995).

Nguyen:1998:DAC

[479] T. Nguyen and Y. Deville. A distributed arc-consistency algorithm. Science of Computer Programming, 30 (1–2):227–250, January 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Concurrent constraint programming (Venice, 1995).

vanAmmers:1998:PCC

[480] Eric W. van Ammers. Program construction by context independent refinements. Science of Computer Programming, 30(3):251– 286, March 2, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/30/3/527.pdf.

Moir:1998:FLL

[481] Mark Moir. Fast, long-lived renaming improved and simplified. Science of Computer Programming, 30 (3):287–308, March 2, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/30/3/525.pdf.

DeFrancesco:1998:SSR

[482] Nicoletta De Francesco, Antonella Santone, and Gigliola Vaglini. State space reduction by non-standard semantics for deadlock analysis. Science of Computer Programming, 30 (3):309–338, March 2, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/30/3/526.pdf.

Cousot:1998:F

[483] Patrick Cousot and Baudouin Le Charlier. Foreword. Science of Computer Programming, 31(1):1, May 8, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/scico/sub/1998/
31/1/504.pdf.

Armstrong:1998:TCB

[484] Tania Armstrong, Kim Marriott, Peter Schachte, and Harald Søndergaard. Two classes of Boolean functions for dependency analysis. Science of Computer Programming, 31(1): 3–45, May 8, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/1/505.pdf. SAS’94

(Namur).

Banerjee:1998:SST

[485] Anindya Banerjee and David A. Schmidt. Stackability in the simplytyped call-by-value lambda calculus. Science of Computer Programming, 31 (1):47–73, May 8, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/1/506.pdf. SAS’94

(Namur).

Halbwachs:1998:ASP

[486] Nicolas Halbwachs. About synchronous programming and abstract interpretation. Science of Computer Programming, 31(1):75–89, May 8, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/scico/sub/1998/
31/1/507.pdf.

Mauborgne:1998:AIU

[487] Laurent Mauborgne. Abstract interpretation using typed decision graphs. Science of Computer Programming, 31 (1):91–112, May 8, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/1/508.pdf. SAS’94

(Namur).

LacknerSolbergGasser:1998:STA

[488] Kirsten Lackner Solberg Gasser, Hanne Riis Nielson, and Flemming Nielson. Strictness and totality analysis. Science of Computer Programming, 31(1):113–145, May 8, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/1/509.pdf. SAS’94

(Namur).

Yi:1998:AIE

[489] Kwangkeun Yi. An abstract interpretation for estimating uncaught exceptions in standard ML programs. Science of Computer Programming, 31 (1):147–173, May 8, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/1/510.pdf. SAS’94

(Namur).

Ciancarini:1998:CML

[490] Paolo Ciancarini and Roberto Gorrieri. Coordination models, languages and architectures. Science of Computer Programming, 31(2–3): 175–177, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642397000191.

Andreoli:1998:MND

[491] Jean-Marc Andreoli, François Pacull, Daniele Pagani, and Remo Pareschi. Multiparty negotiation of dynamic distributed object services. Science of Computer Programming, 31(2–3): 179–203, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/31/2-3/529.pdf.

Bergstra:1998:DTT

[492] J. A. Bergstra and P. Klint. The discrete time ToolBus — a software coordination architecture. Science of Computer Programming, 31(2–3): 205–229, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/31/2-3/530.pdf.

Ciancarini:1998:LCM

[493] P. Ciancarini, M. Mazza, and L. Pazzaglia. A logic for a coordination model with multiple spaces. Science of Computer Programming, 31(2– 3):231–261, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/31/2-3/531.pdf.

Fradet:1998:SG

[494] Pascal Fradet and Daniel Le Métayer. Structured Gamma. Science of Computer Programming, 31(2–3): 263–289, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/2-3/532.pdf.

Fukuda:1998:DCM

[495] Munehiro Fukuda, Lubomir F. Bic, Michael B. Dillencourt, and Fehmina Merchant. Distributed coordination with MESSENGERS. Science of Computer Programming, 31(2–3): 291–311, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/2-3/533.pdf.

Goeman:1998:CIF

[496] H. J. M. Goeman, J. N. Kok, K. Sere, and R. T. Udink. Coordination in the ImpUnity Framework. Science of Computer Programming, 31(2–3): 313–334, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/2-3/534.pdf.

Rowstron:1998:SLM

[497] A. I. T. Rowstron and A. M. Wood. Solving the Linda multiple rd problem using the copy-collect primitive. Science of Computer Programming, 31 (2–3):335–358, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/31/2-3/535.pdf.

Tolksdorf:1998:LSB

[498] Robert Tolksdorf. Laura — a servicebased coordination language. Science of Computer Programming, 31(2– 3):359–381, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/31/2-3/536.pdf.

Simpson:1998:BR

[499] Alex Simpson. Book review. Science of Computer Programming, 31(2– 3):383–385, July 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/31/2-3/548.pdf.

Nielson:1998:F

[500] Hanne Riis Nielson. Foreword. Science of Computer Programming, 32(1– 3):1–2, September 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/32/1-3/537.pdf.

PeytonJones:1998:TBO

[501] Simon L. Peyton Jones and André L. M. Santos. A transformationbased optimiser for Haskell. Science of Computer Programming, 32(1–3): 3–47, September 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/32/1-3/538.pdf.

Smith:1998:SPT

[502] Geoffrey Smith and Dennis Volpano. A sound polymorphic type system for a dialect of C. Science of Computer Programming, 32(1–3):49– 72, September 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/32/1-3/539.pdf.

Seidl:1998:CSD

[503] H. Seidl and M. H. Sørensen. Constraints to stop deforestation. Science of Computer Programming, 32(1–3): 73–107, September 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/32/1-3/540.pdf. 6th

European Symposium on Programming (Linköping, 1996).

Rose:1998:LTH

[504] Eva Rose. Linear-time hierarchies for a functional language machine model. Science of Computer Programming, 32(1–3):109– 143, September 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/32/1-3/541.pdf. 6th

European Symposium on Programming (Linköping, 1996).

Barbuti:1998:TLS

[505] Roberto Barbuti, Nicoletta De Francesco, Paolo Mancarella, and Antonella Santone. Towards a logical semantics for pure Prolog. Science of Computer Programming, 32(1–3):145– 176, September 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1998/32/1-3/542.pdf. 6th

European Symposium on Programming (Linköping, 1996).

Giacobazzi:1998:ODD

[506] Roberto Giacobazzi and Francesco Ranzato. Optimal domains for disjunctive abstract interpretation. Science of Computer Programming, 32(1–3): 177–210, September 1, 1998. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1998/32/1-3/543.pdf. 6th

European Symposium on Programming (Linköping, 1996).

Gorlatch:1999:EIL

[507] Sergei Gorlatch. Extracting and implementing list homomorphisms in parallel program development. Science of Computer Programming, 33 (1):1–27, January 4, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1999/33/1/523.pdf.

Wise:1999:UBE

[508] David S. Wise. Undulant-block elimination and integer-preserving matrix inversion. Science of Computer Programming, 33(1):29–85, January 4, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

                                        tronic). URL http://www.elsevier.
                                        com/cas/tree/store/scico/sub/1999/
                                        33/1/544.pdf.

Cavalcanti:1999:IPP

[509] Ana Cavalcanti, Augusto Sampaio, and Jim Woodcock. An inconsistency in procedures, parameters, and substitution in the refinement calculus. Science of Computer Programming, 33 (1):87–96, January 4, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/33/1/524.pdf.

Fiore:1999:BRB

[510] Marcelo P. Fiore. Book review: Topology via Logic: Steven Vickers, (Cambridge University Press) 0 521 36062 5 hardback, 0-521-57651-2 paperback. Science of Computer Programming, 33 (1):97–99, January 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642398000203.

Compare:1999:UAM

[511] Daniele Compare, Paola Inverardi, and Alexander L. Wolf. Uncovering architectural mismatch in component behavior. Science of Computer Programming, 33(2):101– 131, February 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/33/2/545.pdf.

Ait-Ameur:1999:RRE

[512] Yamine Ait-Ameur. Refinement of rational end-points real numbers by means of floating-point numbers. Science of Computer Programming, 33(2): 133–162, February 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/33/2/547.pdf.

Feijs:1999:RPA

[513] L. M. G. Feijs and R. C. van Ommering. Relation partition algebra — mathematical aspects of uses and part-of relations. Science of Computer Programming, 33(2):163– 212, February 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/33/2/546.pdf.

Etalle:1999:BRB

[514] Sandro Etalle. Book review: An introduction to logic programming through prolog, Michael Spivey, Prentice Hall International Series in Computer Science, Englewood cliffs, NJ, 1996. ISBN 0-13-536047-1. Science of Computer Programming, 33(2): 213–214, February 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642398000240.

Buhrman:1999:BRB

[515] Harry Buhrman. Book review: Communication complexity, Eyal Kushilevitz, Noam Nisan, Cambridge University Press, Cambridge, 1997. ISBN 0-56067-5. Science of Computer Programming, 33(2): 215–216, February 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642398000252.

Anonymous:1999:BRB

[516] Anonymous. Book review: OC-CAM 2: John Galletly, UCL Press London, 1996. ISBN 1-85728-362-7, 2nd edition, 304 pp., £19.95. Science of Computer Programming, 33 (2):217–219, February 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642398000264.

Butler:1999:CDP

[517] Michael Butler. Calculational derivation of pointer algorithms from tree operations. Science of Computer Programming, 33(3):221–260, March 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/scico/sub/1999/
33/3/552.pdf.

Bowman:1999:SCC

[518] H. Bowman, E. A. Boiten, J. Derrick, and M. W. A. Steen. Strategies for consistency checking based on unification. Science of Computer Programming, 33 (3):261–298, March 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/33/3/553.pdf.

Andries:1999:GTS

[519] Marc Andries, Gregor Engels, Annegret Habel, Berthold Hoffmann, Hans-Jörg Kreowski, Sabine Kuske, Detlef Plump, Andy Schürr, and Gabriele Taentzer. Graph transformation for specification and programming. Science of Computer Programming, 34(1):1–54, April 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/34/1/559.pdf.

Meyer:1999:CLP

[520] Uwe Meyer. Correctness of online partial evaluation for a Pascallike language. Science of Computer Programming, 34(1):55–73, April 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (elec-

tronic). URL http://www.elsevier.
com/cas/tree/store/scico/sub/1999/
34/1/551.pdf.

Hartel:1999:BRB

[521] Pieter Hartel. Book review: Purely functional data structures: Chris Okasaki, Cambridge University Press, 1998, ISBN 0-521-63124-6, x + 220 pages, hardback. Science of Computer Programming, 34(1):75–77, April 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642399000027.

Semini:1999:RCT

[522] Laura Semini and Carlo Montangero. A refinement calculus for tuple spaces. Science of Computer Programming, 34 (2):79–140, June 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/34/2/557.pdf.

Morris:1999:LRE

[523] Joseph M. Morris and Alexander Bunkenburg. E3: a logic for reasoning equationally in the presence of partiality. Science of Computer Programming, 34(2):141–158, June 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/34/2/555.pdf.

vanDeursen:1999:BRB

[524] Arie van Deursen. Book review: Modern Compiler Implementation in Java, Andrew W. Appel, Cambridge University Press, 1998, x + 548 pages, hardback, ISBN 0-521-58388-8. Science of Computer Programming, 34 (2):159–161, June 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642399000039.

Astesiano:1999:SHT

[525] Egidio Astesiano, Gianna Reggio, and Elena Zucca. Stores as homomorphisms and their transformations: a uniform approach to structured types in imperative languages. Science of Computer Programming, 34 (3):163–190, July 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/34/3/558.pdf.

Hehner:1999:SPT

[526] Eric C. R. Hehner. Specifications, programs, and total correctness. Science of Computer Programming, 34 (3):191–205, July 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.com/cas/tree/store/
scico/sub/1999/34/3/563.pdf.

Osorio:1999:TPO

[527] Mauricio Osorio, Bharat Jayaraman, and David A. Plaisted. Theory of partial-order programming. Science of Computer Programming, 34 (3):207–238, July 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1999/34/3/554.pdf.

Anonymous:1999:Ea

[528] Anonymous.  Editorial.  Science

of Computer Programming, 35(1): 1–2, September 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642399000313.

deMoor:1999:BAG

[529] Oege de Moor and Jeremy Gibbons. Bridging the algorithm gap: a linear-time functional program for paragraph formatting. Science of Computer Programming, 35(1):3– 27, September 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1999/35/1/566.pdf.

Boiten:1999:CCC

[530] Eerke Boiten, John Derrick, Howard Bowman, and Maarten Steen. Constructive consistency checking for partial specification in Z. Science of Computer Programming, 35(1):29– 75, September 1, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.com/cas/tree/store/
                                        scico/sub/1999/35/1/567.pdf.

Anonymous:1999:Eb

[531] Anonymous. Editorial(s). Science of Computer Programming, 35(2–3): 77, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
32/18/17/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/32/18/17/article.pdf.

Aiken:1999:ISC

[532] Alexander Aiken. Introduction to set constraint-based program analysis. Science of Computer Programming, 35(2–3):79–111, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.nl/gej-ng/10/39/21/32/
18/18/abstract.html;   http://www.
elsevier.nl/gej-ng/10/39/21/32/
18/18/article.pdf.

Brauburger:1999:ADF

[533] Jürgen Brauburger and Jürgen Giesl. Approximating the domains of functional and imperative programs. Science of Computer Programming, 35 (2–3):113–136, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.elsevier.nl/gej-
ng/10/39/21/32/18/19/abstract.html;
 http://www.elsevier.nl/gej-ng/
10/39/21/32/18/19/article.pdf.

Fecht:1999:FSG

[534] Christian Fecht and Helmut Seidl. A faster solver for general systems of equations. Science of Computer Programming, 35(2–3):137– 161, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
32/18/20/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/32/18/20/article.pdf.

Ferdinand:1999:CBP

[535] Christian Ferdinand, Florian Martin, Reinhard Wilhelm, and Martin Alt. Cache behavior prediction by abstract interpretation. Science of Computer Programming, 35(2–3):163– 189, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        32/18/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/32/18/21/article.pdf.

Rehof:1999:TCF

[536] Jakob Rehof and Torben Æ. Mogensen. Tractable constraints in finite semilattices. Science of Computer Programming, 35(2–3):191– 221, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        32/18/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/32/18/22/article.pdf.

Venet:1999:AAP

[537] A. Venet. Automatic analysis of pointer aliasing for untyped programs. Science of Computer Programming, 35(2–3): 223–248, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        32/18/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/32/18/23/article.pdf.

Anonymous:1999:BR

[538] Anonymous. Book review(s). Science of Computer Programming, 35(2–3): 249–253, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
32/18/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/32/18/24/article.pdf.

Anonymous:1999:I

[539] Anonymous. Index. Science of Computer Programming, 35(2–3): 255, November ??, 1999. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
32/18/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/32/18/25/article.pdf.

Cuellar:2000:F

[540] Jorge Cuellar, Stefania Gnesi, and Diego Latella. Foreword. Science of Computer Programming, 36 (1):1–3, January ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/21/21/article.pdf.

Berry:2000:EFM

[541] Gérard Berry et al. ESTEREL: a formal method applied to avionic software development. Science of Computer Programming, 36(1):5– 25, January ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/21/22/article.pdf.

Bozga:2000:VTG

[542] Marius Bozga et al. Verification and test generation for the SS-COP protocol. Science of Computer Programming, 36(1):27–52, January ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.nl/gej-ng/10/39/21/39/
21/23/abstract.html;   http://www.
elsevier.nl/gej-ng/10/39/21/39/
21/23/article.pdf.

Hartonas-Garmhausen:2000:VSC

[543] Vicky Hartonas-Garmhausen, Sergio Campos, Alessandro Cimatti, Edmund Clarke, and Fausto Giunchiglia. Verification of a safety-critical railway interlocking system with realtime constraints. Science of Computer Programming, 36(1):53–64, January ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.nl/gej-ng/10/39/21/39/
21/24/abstract.html;   http://www.
elsevier.nl/gej-ng/10/39/21/39/
21/24/article.pdf.

Heimdahl:2000:ANW

[544] Mats P. E. Heimdahl and Barbara J. Czerny. On the analysis needs when verifying state-based software requirements: an experience report. Science of Computer Programming, 36(1): 65–96, January ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/21/25/article.pdf.

Hermanns:2000:ACM

[545] Holger Hermanns and Joost-Pieter Katoen. Automated compositional Markov chain generation for a plainold telephone system. Science of Computer Programming, 36(1):97– 127, January ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/21/26/article.pdf.

Baxter:2000:P

[546] Ira Baxter, Alex Quilici, and Chris Verhoef. Preface. Science of Computer Programming, 36(2–3):129– 130, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/23/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/23/21/article.pdf.

Baxter:2000:RER

[547] Ira D. Baxter and Michael Mehlich. Reverse engineering is reverse forward engineering. Science of Computer Programming, 36(2–3):131– 147, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/23/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/23/22/article.pdf.

Girard:2000:CAD

[548] Jean-François Girard and Rainer Koschke. A comparison of abstract data types and objects recovery techniques. Science of Computer Programming, 36(2–3):149– 181, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        39/23/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/39/23/23/article.pdf.

Storey:2000:HDP

[549] M.-A. D. Storey, K. Wong, and H. A. Müller. How do program understanding tools affect how programmers understand programs? Science of Computer Programming, 36(2–3): 183–207, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        39/23/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/39/23/24/article.pdf.

vandenBrand:2000:GCS

[550] Mark van den Brand, Alex Sellink, and Chris Verhoef. Generation of components for software renovation factories from context-free grammars. Science of Computer Programming, 36(2–3): 209–266, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        39/23/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/39/23/25/article.pdf.

Jerding:2000:UVA

[551] Dean Jerding and Spencer Rugaber. Using visualization for architectural localization and extraction. Science of Computer Programming, 36(2–3): 267–284, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/23/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/23/26/article.pdf.

Quilici:2000:PPM

[552] Alex Quilici, Steven Woods, and Yongjun Zhang. Program plan matching: experiments with a constraintbased approach. Science of Computer Programming, 36(2–3):285– 302, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/23/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/23/27/article.pdf.

vanDeursen:2000:PPR

[553] Arie van Deursen, Alex Quilici, and Steve Woods. Program plan recognition for year 2000 tools. Science of Computer Programming, 36(2–3): 303–324, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/23/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/23/28/article.pdf.

Anonymous:2000:Ia

[554] Anonymous. Index. Science of Computer Programming, 36(2–3):325– 326, March ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
39/23/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/39/23/29/article.pdf.

Anonymous:2000:Ea

[555] Anonymous.  Editorial(s).  Sci-

ence of Computer Programming, 37 (1–3):1–2, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/21/article.pdf.

Dijkstra:2000:CCB

[556] Rutger M. Dijkstra. Computation calculus bridging a formalization gap. Science of Computer Programming, 37 (1–3):3–36, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/22/article.pdf.

Gibbons:2000:GDA

[557] Jeremy Gibbons. Generic downwards accumulations. Science of Computer Programming, 37(1–3): 37–65, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/23/article.pdf.

Hughes:2000:GMA

[558] John Hughes. Generalising monads to arrows. Science of Computer Programming, 37(1–3):67–111, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.nl/gej-ng/10/39/21/41/
21/24/abstract.html;   http://www.
elsevier.nl/gej-ng/10/39/21/41/
21/24/article.pdf.

Joshi:2000:SAS

[559] Rajeev Joshi and K. Rustan M. Leino. A semantic approach to secure information flow. Science of Computer Programming, 37(1–3): 113–138, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/25/article.pdf.

vonKarger:2000:CAR

[560] Burghard von Karger. A calculational approach to reactive systems. Science of Computer Programming, 37(1– 3):139–161, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/26/article.pdf.

Sorensen:2000:CPT

[561] Morten Heine B. Sørensen. Convergence of program transformers in the metric space of trees. Science of Computer Programming, 37(1–3):

163–205, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/27/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/27/article.pdf.

Jay:2000:CPP

[562] C. Barry Jay. Costing parallel programs as a function of shapes. Science of Computer Programming, 37(1– 3):207–224, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/28/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/28/article.pdf.

Nitsche:2000:SAF

[563] Thomas Nitsche. Shapeliness analysis of functional programs with algebraic data types. Science of Computer Programming, 37(1–3):225– 252, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/29/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/29/article.pdf.

Loulergue:2000:CFB

[564] Frédéric Loulergue, Gaétan Hains, and Christian Foisy. A calculus of functional BSP programs. Science of Computer Programming, 37(1–3): 253–277, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/30/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/39/
21/41/21/30/article.pdf.

McMillan:2000:MHV

[565] K. L. McMillan. A methodology for hardware verification using compositional model checking. Science of Computer Programming, 37(1–3): 279–309, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/31/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/31/article.pdf.

Anonymous:2000:Eb

[566] Anonymous.  Erratum.  Science

of Computer Programming, 37(1– 3):311, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/32/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/32/article.pdf.

Anonymous:2000:Ib

[567] Anonymous.  Index.  Science

of Computer Programming, 37(1– 3):313, May ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
41/21/33/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/41/21/33/article.pdf.

Berghammer:2000:DRP

[568] Rudolf Berghammer and Thorsten Hoffmann. Deriving relational programs for computing kernels by reconstructing a proof of Richardson’s theorem. Science of Computer Programming, 38(1–3):1–25, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.nl/gej-ng/10/39/21/42/
21/21/abstract.html;   http://www.
elsevier.nl/gej-ng/10/39/21/42/
21/21/article.pdf.

Cortesi:2000:CAD

[569] Agostino Cortesi, Baudouin Le Charlier, and Pascal Van Hentenryck. Combinations of abstract domains for logic programming: open product and generic pattern construction. Science of Computer Programming, 38(1–3): 27–71, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
42/21/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/42/21/22/article.pdf.

Ayed:2000:CPA

[570] Rahma Ben Ayed, Jules Desharnais, Marc Frappier, and Ali Mili. A calculus of program adaptation and its applications. Science of Computer Programming, 38(1–3):73– 123, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
42/21/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/42/21/23/article.pdf.

Diehl:2000:GMD

[571] Stephan Diehl. A generative methodology for the design of abstract machines. Science of Computer Programming, 38(1–3):125–142, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.nl/gej-ng/10/39/21/42/
21/24/abstract.html;   http://www.
elsevier.nl/gej-ng/10/39/21/42/
21/24/article.pdf.

Balbo:2000:PMD

[572] Gianfranco Balbo. Performance Modeling with Deterministic and Stochastic Petri Nets, Christoph Lindemann, Wiley-Interscience Series in Systems and Optimization Wiley, New York, 1998, 405 pages, ISBN 0-471-97646-6. Science of Computer Programming, 38(1–3):143– 146, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
42/21/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/42/21/25/article.pdf.

vanEijck:2000:TSL

[573] Jan van Eijck. The Tbilisi Symposium on Logic, Language and Computation: Selected Papers, Jonathan Ginzburg, Zurab Khasidashvili, Carl Vogel, Jean-Jacques Lévi, and Enric Vallduví (Eds.), Studies in Logic, Language and Information series of CSLI Publications and FoLLI, Stanford, 1998, xxxi + 376 pages, ISBN 1-57586-098-8. Science of Computer Programming, 38(1–3): 147–149, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
42/21/26/abstract.html;   http:

                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/42/21/26/article.pdf.

Sangiorgi:2000:CMS

[574] Davide Sangiorgi. Communicating and Mobile Systems: the π-calculus, Robin Milner, Cambridge University Press, Cambridge, 1999, 174 pages, ISBN 0-521-64320-1. Science of Computer Programming, 38(1–3): 151–153, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        42/21/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/42/21/27/article.pdf.

Streicher:2000:PFM

[575] Thomas Streicher. Practical Foundations of Mathematics, Paul Taylor, Cambridge Studies in Advanced Mathematics, Vol. 59, Cambridge University Press, Cambridge, 1999. xi + 572 pages, price £50 paperback, ISBN 0-521-63107-6. Science of Computer Programming, 38(1–3): 155–157, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        42/21/28/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/42/21/28/article.pdf.

Anonymous:2000:Ic

[576] Anonymous. Index. Science of Computer Programming, 38(1–3): 159, August ??, 2000. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        42/21/29/abstract.html;   http:

//www.elsevier.nl/gej-ng/10/39/
21/42/21/29/article.pdf.

Levi:2001:P

[577] Giorgio Levi. Preface. Science of Computer Programming, 39(1): 1, January ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/20/20/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/20/20/article.pdf.

Ruthing:2001:CMP

[578] Oliver Rüthing. Code motion in the presence of critical edges without bidirectional data flow analysis. Science of Computer Programming, 39 (1):3–29, January ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/20/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/20/21/article.pdf.

Hind:2001:EEP

[579] Michael Hind and Anthony Pioli. Evaluating the effectiveness of pointer alias analyses. Science of Computer Programming, 39(1):31– 55, January ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/20/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/20/22/article.pdf.

Duggan:2001:FSI

[580] Dominic Duggan. Finite subtype inference with explicit polymorphism. Science of Computer Programming, 39(1): 57–92, January ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/20/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/20/23/article.pdf.

Levi:2001:SSA

[581] Francesca Levi. A symbolic semantics for abstract model checking. Science of Computer Programming, 39(1): 93–123, January ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/20/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/20/24/article.pdf.

Volpe:2001:FOL

[582] Paolo Volpe. A first-order language for expressing sharing and type properties of logic programs. Science of Computer Programming, 39(1):125– 148, January ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/20/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/20/25/article.pdf.

Feijs:2001:MNA

[583] Loe M. G. Feijs. Mechanisms for naming: an algebraic approach with an application to Java. Science of Computer Programming, 39(2–3): 149–188, March ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/

43/23/20/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/23/20/article.pdf.

Janicki:2001:FST

[584] Ryszard Janicki and Ridha Khedri. On a formal semantics of tabular expressions. Science of Computer Programming, 39(2–3):189– 213, March ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/23/21/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/23/21/article.pdf.

Groote:2001:ATH

[585] Jan Friso Groote and Jos van Wamel. Analysis of three hybrid systems in timed μCRL. Science of Computer Programming, 39(2–3): 215–247, March ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/23/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/23/22/article.pdf.

Moreau:2001:DDS

[586] Luc Moreau. Distributed directory service and message routing for mobile agents. Science of Computer Programming, 39(2–3):249– 272, March ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
43/23/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/43/23/23/article.pdf.

Anonymous:2001:AIVa

[587] Anonymous. Author index of Volume 39. Science of Computer Programming, 39(2–3):273, March ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        43/23/24/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/43/23/24/article.pdf.

Astesiano:2001:FAS

[588] Egidio Astesiano and Carlo Ghezzi. Fundamental approaches to software engineering. Science of Computer Programming, 40(1):1–2, May ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/39/21/47/
                                        20/20/abstract.html;   http://www.
                                        elsevier.nl/gej-ng/10/39/21/47/
                                        20/20/article.pdf.

Bradley:2001:FDL

[589] Steven Bradley, William Henderson, David Kendall, and Adrian Robson. A formal design language for real-time systems with data. Science of Computer Programming, 40 (1):3–29, May ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        47/20/21/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/47/20/21/article.pdf.

Grieskamp:2001:SES

[590] Wolfgang Grieskamp, Maritta Heisel, and Heiko Dörr. Specifying embedded systems with statecharts and Z: an agenda for cyclic software components. Science of Computer Programming, 40 (1):31–57, May ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
47/20/22/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/47/20/22/article.pdf.

Mota:2001:MCC

[591] Alexandre Mota and Augusto Sampaio. Model-checking CSP-Z: strategy, tool support and industrial application. Science of Computer Programming, 40 (1):59–96, May ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
47/20/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/47/20/23/article.pdf.

Padberg:2001:RBR

[592] J. Padberg, M. Gajewsky, and C. Ermel. Rule-based refinement of highlevel nets preserving safety properties. Science of Computer Programming, 40 (1):97–118, May ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
47/20/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/47/20/24/article.pdf.

Scholz:2001:IDS

[593] Peter Scholz. Incremental design of statechart specifications. Science of Computer Programming, 40 (1):119–145, May ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        47/20/25/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/47/20/25/article.pdf.

Smith:2001:P

[594] Dennis Smith and Steve Woods. Preface. Science of Computer Programming, 40(2–3):147–149, July ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

                                        elsevier.nl/gej-ng/10/39/21/47/
                                        25/21/abstract.html;   http://www.
                                        elsevier.nl/gej-ng/10/39/21/47/
                                        25/21/article.pdf.

Francel:2001:VSW

[595] Margaret Ann Francel and Spencer Rugaber. The value of slicing while debugging. Science of Computer Programming, 40(2–3):151– 169, July ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        47/25/22/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/47/25/22/article.pdf.

Cifuentes:2001:RJT

[596] Cristina Cifuentes and Mike Van Emmerik. Recovery of jump table case statements from binary code. Science of Computer Programming, 40(2– 3):171–188, July ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        47/25/23/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/47/25/23/article.pdf.

vanDeursen:2001:ESC

[597] Arie van Deursen and Leon Moonen. An empirical study into COBOL type inferencing. Science of Computer Programming, 40(2–3): 189–211, July ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
47/25/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/47/25/24/article.pdf.

Antoniol:2001:DCT

[598] G. Antoniol, B. Caprile, A. Potrich, and P. Tonella. Design-code traceability recovery: selecting the basic linkage properties. Science of Computer Programming, 40(2–3):213– 234, July ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
47/25/25/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/47/25/25/article.pdf.

DiMartino:2001:PPA

[599] Beniamino Di Martino, Antonino Mazzeo, Nicola Mazzocca, and Umberto Villano. Parallel program analysis and restructuring by detection of point-to-point interaction patterns and their transformation into collective communication constructs. Science of Computer Programming, 40(2– 3):235–263, July ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

www.elsevier.nl/gej-ng/10/39/21/
47/25/26/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/39/
21/47/25/26/article.pdf.

Anonymous:2001:AIVb

[600] Anonymous. Author index of Volume 40. Science of Computer Programming, 40(2–3):265, July ??, 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://

                                        www.elsevier.nl/gej-ng/10/39/21/
                                        47/25/27/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/10/39/
                                        21/47/25/27/article.pdf.

Naumann:2001:PTS

[601] David A. Naumann. Predicate transformer semantics of a higher-order imperative language with record subtyping. Science of Computer Programming, 41(1):1–51, September 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/39/21/53/21/21/abstract.html.

Plath:2001:FIU

[602] Malte Plath and Mark Ryan. Feature integration using a feature construct. Science of Computer Programming, 41 (1):53–84, September 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/53/21/22/abstract.html.

Marchand:2001:FVP

[603] H. Marchand, E. Rutten, M. Le Borgne, and M. Samaan. Formal verification of programs specified with signal: application to a power transformer station controller. Science of Computer Programming, 41(1): 85–104, September 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/53/21/23/abstract.html.

Canal:2001:CIS

[604] Carlos Canal, Ernesto Pimentel, and José M. Troya. Compatibility and inheritance in software architectures. Science of Computer Programming, 41 (2):105–138, October 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/53/31/27/abstract.html.

Liu:2001:SIE

[605] Yanhong A. Liu, Scott D. Stoller, and Tim Teitelbaum. Strengthening invariants for efficient computation. Science of Computer Programming, 41 (2):139–172, October 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/53/31/28/abstract.html.

Hesselink:2001:CDC

[606] Wim H. Hesselink, Arnold Meijster, and Coenraad Bron. Concurrent determination of connected components. Science of Computer Programming, 41 (2):173–194, October 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/53/31/29/abstract.html.

Klaudel:2001:CHL

[607] Hanna Klaudel. Compositional highlevel Petri net semantics of a parallel programming language with procedures. Science of Computer Programming, 41(3):195–240, November 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/53/32/27/abstract.html.

Hennessy:2001:UFI

[608] M. Hennessy, H. Lin, and J. Rathke. Unique fixpoint induction for messagepassing process calculi. Science of Computer Programming, 41(3): 241–275, November 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/53/32/28/abstract.html.

Omicini:2001:TST

[609] Andrea Omicini and Enrico Denti. From tuple spaces to tuple centres. Science of Computer Programming, 41 (3):277–294, November 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/53/32/29/abstract.html.

Anonymous:2001:AIVc

[610] Anonymous. Author index of Volume 41. Science of Computer Programming, 41(3):295, November 2001. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/53/32/30/abstract.html.

Luqi:2002:SEB

[611] Luqi and Manfred Broy. Software engineering beyond our planning horizon: automation for computer-based systems. Science of Computer Programming, 42(1):1–10, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/61/27/27/abstract.html.

Berry:2002:FMV

[612] Daniel M. Berry. Formal methods: the very idea — Some thoughts about why they work when they work. Science of Computer Programming, 42 (1):11–27, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/61/27/28/abstract.html.

Robertson:2002:PFE

[613] David Robertson. Pitfalls of formality in early system design. Science of Computer Programming, 42 (1):29–38, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/61/27/29/abstract.html.

Cleaveland:2002:GTV

[614] Rance Cleaveland and Steven T. Sims. Generic tools for verifying concurrent systems. Science of Computer Programming, 42(1):39–47, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/61/27/30/abstract.html.

Kwak:2002:PAS

[615] Hee-Hwan Kwak, Insup Lee, and Oleg Sokolsky. Parametric approach to the specification and analysis of real-time scheduling based on ACSR-VP. Science of Computer Programming, 42 (1):49–60, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/61/27/31/abstract.html.

Berzins:2002:LIA

[616] V. Berzins. Lightweight inference for automation efficiency. Science of Computer Programming, 42(1):61–74, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/39/21/61/
                                        27/32/abstract.html.

Polak:2002:FMP

[617] Wolfgang Polak. Formal methods in practice. Science of Computer Programming, 42(1):75–85, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/39/21/61/27/33/abstract.html.

Gelfond:2002:MRK

[618] Michael Gelfond and Richard Watson. On methodology of representing knowledge in dynamic domains. Science of Computer Programming, 42 (1):87–99, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/61/27/34/abstract.html.

Volker:2002:AVF

[619] Norbert Völker and Bernd J. Krämer. Automated verification of function block-based industrial control systems. Science of Computer Programming, 42 (1):101–113, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/61/27/35/abstract.html.

Cooke:2002:ACS

[620] Daniel E. Cooke and Vladik Kreinovich. Automatic concurrency in SequenceL. Science of Computer Programming, 42 (1):115–128, January 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/61/27/36/abstract.html.

vanderAalst:2002:CBS

[621] W. M. P. van der Aalst, K. M. van Hee, and R. A. van der Toorn. Componentbased software architectures: a framework based on inheritance of behavior. Science of Computer Programming, 42(2–3):129–171, February 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/61/31/27/abstract.html.

See remark [648].

Feijs:2002:CA

[622] Loe M. G. Feijs and Yuechen Qian. Component algebra. Science of Computer Programming, 42(2–3):173–228, February 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.com/gej-ng/10/39/21/61/
31/28/abstract.html.

Tonella:2002:CSC

[623] Paolo Tonella and Alessandra Potrich. Cjj: a subset of C++ compliant with Java. Science of Computer Programming, 42(2–3):229–271, February 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/61/31/29/abstract.html.

Anonymous:2002:AIVa

[624] Anonymous. Author index of Volume 42. Science of Computer Programming, 42(2–3):273, February 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/61/31/30/abstract.html.

Usenko:2002:SSG

[625] Yaroslav S. Usenko. State space generation for the HAVi leader election protocol. Science of Computer Programming, 43(1):1–33, April 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/62/27/27/abstract.html.

Jansson:2002:PDC

[626] Patrik Jansson and Johan Jeuring. Polytypic data conversion programs. Science of Computer Programming, 43(1):35–75, April 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/62/27/28/abstract.html.

Morris:2002:SIT

[627] J. M. Morris and A. Bunkenburg. A source of inconsistency in theories of nondeterministic functions. Science of Computer Programming, 43(1):77–89, April 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/62/27/29/abstract.html.

Backhouse:2002:P

[628] Roland Backhouse and José Oliveira. Preface. Science of Computer Programming, 43(2–3):91–92, June 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/62/36/27/abstract.html.

Nordlander:2002:PSO

[629] Johan Nordlander. Polymorphic subtyping in O’Haskell. Science of Computer Programming, 43(2–3):93– 127, June 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.com/gej-ng/10/39/21/62/
36/28/abstract.html.

Hinze:2002:PVP

[630] Ralf Hinze. Polytypic values possess polykinded types. Science of Computer Programming, 43(2–3):129– 159, June 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.com/gej-ng/10/39/21/62/
36/29/abstract.html.

Hayes:2002:RAR

[631] Ian Hayes. Reasoning about real-time repetitions: terminating and nonterminating. Science of Computer Programming, 43(2–3):161–192, June 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/62/36/30/abstract.html.

Abramov:2002:URA

[632] Sergei Abramov and Robert Glück. The universal resolving algorithm and its correctness: inverse computation in a functional language. Science of Computer Programming, 43 (2–3):193–229, June 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/62/36/31/abstract.html.

Anonymous:2002:AIVb

[633] Anonymous. Author index of Volume 43. Science of Computer Programming, 43(2–3):231, June 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/62/36/32/abstract.html.

Evans:2002:P

[634] Andy Evans, Stuart Kent, and Bran Selic. Preface. Science of Computer Programming, 44(1):1–4, July 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/86/27/27/abstract.html.

Atkinson:2002:PSM

[635] Colin Atkinson and Thomas Kühne. Profiles in a strict metamodeling framework. Science of Computer Programming, 44(1):5–22, July 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/86/27/28/abstract.html.

Garlan:2002:RNA

[636] David Garlan, Shang-Wen Cheng, and Andrew J. Kompanek. Reconciling the needs of architectural description with object-modeling notations. Science of Computer Programming, 44(1):23–49, July 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/86/27/29/abstract.html.

Hussmann:2002:MAT

[637] Heinrich Hussmann, Birgit Demuth, and Frank Finger. Modular architecture for a toolset supporting OCL. Science of Computer Programming, 44(1):51–69, July 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/86/27/30/abstract.html.

Clarke:2002:ESU

[638] Siobhán Clarke. Extending standard UML with model composition semantics. Science of Computer Programming, 44(1):71–100, July 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

http://www.elsevier.com/gej-ng/
10/39/21/86/27/31/abstract.html.

Cortellessa:2002:PUP

[639] Vittorio Cortellessa and Raffaela Mirandola. PRIMA-UML: a performance validation incremental methodology on early UML diagrams. Science of Computer Programming, 44 (1):101–129, July 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.elsevier.com/gej-ng/10/39/
21/86/27/32/abstract.html.

Ehrig:2002:F

[640] Hartmut Ehrig, Gregor Engels, Hans-Jörg Kreowski, and Gabriele Taentzer. Foreword. Science of Computer Programming, 44(2):131–132, August 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.com/gej-ng/10/39/21/86/
49/27/abstract.html.

Wermelinger:2002:GTA

[641] Michel Wermelinger and José Luiz Fiadeiro. A graph transformation approach to software architecture reconfiguration. Science of Computer Programming, 44(2):133–155, August 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

                                        elsevier.com/gej-ng/10/39/21/86/
                                        49/28/abstract.html.

Minas:2002:CRD

[642] Mark Minas. Concepts and realization of a diagram editor generator based on hypergraph transformation. Science of Computer Programming, 44 (2):157–180, August 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/86/49/29/abstract.html.

Bardohl:2002:VEV

[643] Roswitha Bardohl. A visual environment for visual languages. Science of Computer Programming, 44 (2):181–203, August 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.elsevier.com/gej-ng/10/39/
                                        21/86/49/30/abstract.html.

Varro:2002:DAT

[644] Dániel Varró, Gergely Varró, and András Pataricza. Designing the automatic transformation of visual languages. Science of Computer Programming, 44(2):205–227, August 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/39/21/86/49/31/abstract.html.

Depke:2002:FAO

[645] Ralph Depke, Reiko Heckel, and Jochen Malte Küster. Formal agentoriented modeling with UML and graph transformation. Science of Computer Programming, 44(2):229– 252, August 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

elsevier.com/gej-ng/10/39/21/86/
49/32/abstract.html.

Engels:2002:HCM

[646] A. G. Engels, S. Mauw, and M. A. Reniers. A hierarchy of communication models for Message Sequence Charts. Science of Computer Programming, 44 (3):253–292, September 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kleist:2002:IOM

[647] Josva Kleist and Davide Sangiorgi. Imperative objects as mobile processes. Science of Computer Programming, 44 (3):293–342, September 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanderAalst:2002:CPI

[648] W. M. P. van der Aalst, K. M. van Hee, and R. A. van der Toorn. “Compositionality of projection inheritance” [Sci. Comput. Programming 42 (2–3) (2002) 129–171]. Science of Computer Programming, 44(3):343–344, September 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). See [621].

Anonymous:2002:AIa

[649] Anonymous. Author index. Science of Computer Programming, 44 (3):345, September 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Verhoef:2002:QIP

[650] C. Verhoef. Quantitative IT portfolio management. Science of Computer Programming, 45(1):1–96, October 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2002:EBa

[651] Anonymous. Editorial board. Science of Computer Programming, 45(1): ii, October 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chikofsky:2002:P

[652] Elliot Chikofsky and Chris Verhoef. Preface. Science of Computer Programming, 45(2–3):97–98, November/ December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jahnke:2002:SIE

[653] Jens H. Jahnke, Wilhelm Schäfer, Jörg P. Wadsack, and Albert Zündorf. Supporting iterations in exploratory database reengineering processes. Science of Computer Programming, 45(2– 3):99–136, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sousa:2002:CRA

[654] Pedro Sousa, Lurdes Pedro de Jesus, Gonçalo Pereira, and Fernando Brito e Abreu. Clustering relations into abstract ER schemas for database reverse engineering. Science of Computer Programming, 45(2–3):137–153, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chaumun:2002:CIM

[655] M. Ajmal Chaumun, Hind Kabaili, Rudolf K. Keller, and François Lustman. A change impact model for changeability assessment in objectoriented software systems. Science of Computer Programming, 45(2–3): 155–174, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brunekreef:2002:TUC

[656] Jacob Brunekreef and Bob Diertens. Towards a user-controlled software renovation factory. Science of Computer Programming, 45(2–3):175–191, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sellink:2002:RCC

[657] Alex Sellink, Harry Sneed, and Chris Verhoef. Restructuring of COBOL/ CICS legacy systems. Science of Computer Programming, 45(2–3):193–243, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rayside:2002:EJL

[658] Derek Rayside and Kostas Kontogiannis. Extracting Java library subsets for deployment on embedded systems. Science of Computer Programming, 45(2– 3):245–270, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2002:AIb

[659] Anonymous. Author index. Science of Computer Programming, 45(2–3):271, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2002:EBb

[660] Anonymous. Editorial board. Science of Computer Programming, 45 (2–3):CO2, November/December 2002. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ciancarini:2003:ICL

[661] Paolo Ciancarini and Alexander Wolf. Issues in coordination languages and architectures. Science of Computer Programming, 46(1–2):1–3, January/ February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Attanasio:2003:SAR

[662] Candida Attanasio, Flavio Corradini, and Paola Inverardi. Static analysis of real-time component-based systems configurations. Science of Computer Programming, 46(1–2):5–30, January/ February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bonsangue:2003:CCM

[663] Marcello M. Bonsangue, Joost N. Kok, and Gianluigi Zavattaro. Comparing coordination models and architectures using embeddings. Science of Computer Programming, 46(1–2):31– 69, January/February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brogi:2003:ECS

[664] Antonio Brogi and Jean-Marie Jacquet. On the expressiveness of coordination via shared dataspaces. Science of Computer Programming, 46(1–2):71– 98, January/February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moran:2003:EFS

[665] Andrew Moran, David Sands, and Magnus Carlsson. Erratic fudgets: a semantic theory for an embedded coordination language. Science of Computer Programming, 46(1–2):99– 135, January/February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rowstron:2003:UMC

[666] Antony I. T. Rowstron. Using mobile code to provide fault tolerance in tuple space based coordination languages. Science of Computer Programming, 46(1–2):137–162, January/ February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vitek:2003:CPS

[667] Jan Vitek, Ciarán Bryce, and Manuel Oriol. Coordinating processes with secure spaces. Science of Computer Programming, 46(1–2):163–193, January/ February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:EBa

[668] Anonymous. Editorial board. Science of Computer Programming, 46(1– 2):ii, January/February 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Garavel:2003:SIF

[669] Hubert Garavel, Stefania Gnesi, and Ina Schieferdecker. Special issue on the Fifth International Workshop of the ERCIM Working Group on Formal Methods for Industrial Critical Systems, Berlin, April 3–4, 2000—- selected papers. Science of Computer Programming, 46(3):195–196, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carreira:2003:AVO

[670] Paulo J. F. Carreira and Miguel E. F. Costa. Automatically verifying an object-oriented specification of the steam-boiler system. Science of Computer Programming, 46 (3):197–217, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Maraninchi:2003:MAN

[671] Florence Maraninchi and Yann Rémond. Mode-automata: a new domainspecific construct for the development of safe critical systems. Science of Computer Programming, 46(3):219– 254, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mateescu:2003:EFM

[672] Radu Mateescu and Mihaela Sighireanu. Efficient on-the-fly modelchecking for regular alternation-free mu-calculus. Science of Computer Programming, 46(3):255–281, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Requet:2003:BME

[673] Antoine Requet. A B model for ensuring soundness of a large subset of the Java Card virtual machine. Science of Computer Programming, 46 (3):283–306, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:AIa

[674] Anonymous. Author index. Science of Computer Programming, 46(3):307, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:EBb

[675] Anonymous. Editorial board. Science of Computer Programming, 46(3):CO2, March 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vandenBrand:2003:P

[676] Mark van den Brand, Marjan Mernik, and Didier Parigot. Preface. Science of Computer Programming, 47(1):1, April 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Doh:2003:CPL

[677] Kyung-Goo Doh and Peter D. Mosses. Composing programming languages by combining action-semantics modules. Science of Computer Programming, 47(1):3–36, April 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hedin:2003:JAO

[678] Görel Hedin and Eva Magnusson. JastAdd — an aspect-oriented compiler construction system. Science of Computer Programming, 47(1):37–58, April 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kuipers:2003:OOT

[679] Tobias Kuipers and Joost Visser. Object-oriented tree traversal with JJ-Forester. Science of Computer Programming, 47(1):59–87, April 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:EBc

[680] Anonymous. Editorial board. Science of Computer Programming, 47(1): ii, April 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cortesi:2003:SA

[681] Agostino Cortesi and Gilberto Filé. Static analysis. Science of Computer Programming, 47(2–3):89–90, May/ June 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aldrich:2003:CSE

[682] Jonathan Aldrich, Emin Gün Sirer, Craig Chambers, and Susan J. Eggers. Comprehensive synchronization elimination for Java. Science of Computer Programming, 47(2–3):91–120, May/ June 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mauborgne:2003:IRT

[683] Laurent Mauborgne. Infinitary relations and their representation. Science of Computer Programming, 47(2– 3):121–144, May/June 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nielson:2003:AIM

[684] Flemming Nielson, René Rydhof Hansen, and Hanne Riis Nielson. Abstract interpretation of mobile ambients. Science of Computer Programming, 47(2–3):145–175, May/June 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Monniaux:2003:ACP

[685] David Monniaux. Abstracting cryptographic protocols with tree automata. Science of Computer Programming, 47 (2–3):177–202, May/June 2003. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fernandez:2003:SSR

[686] Jean-Claude Fernandez, Marius Bozga, and Lucian Ghirvu. State space reduction based on live variables analysis. Science of Computer Programming, 47(2–3):203–220, May/June 2003. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Liu:2003:EDC

[687] Yanhong A. Liu and Scott D. Stoller. Eliminating dead code on recursive data. Science of Computer Programming, 47(2–3):221–242, May/June 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:AIb

[688] Anonymous. Author index. Science of Computer Programming, 47 (2–3):243, May/June 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:EBd

[689] Anonymous. Editorial board. Science of Computer Programming, 47 (2–3):CO2, May/June 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Paleri:2003:PRE

[690] V. K. Paleri, Y. N. Srikant, and P. Shankar. Partial redundancy elimination: a simple, pragmatic, and provably correct algorithm. Science of Computer Programming, 48(1):1–20, July 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Edwards:2003:SES

[691] Stephen A. Edwards and Edward A. Lee. The semantics and execution of a synchronous block-diagram language. Science of Computer Programming, 48(1):21–42, July 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cao:2003:DCM

[692] Jiannong Cao, Alvin Chan, Yudong Sun, and Kang Zhang. Dynamic configuration management in a graphoriented distributed programming environment. Science of Computer Programming, 48(1):43–65, July 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Berghammer:2003:LAA

[693] Rudolf Berghammer and Florian Reuter. A linear approximation algorithm for bin packing with absolute approximation factor 3/2. Science of Computer Programming, 48(1):67–80, July 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:EBe

[694] Anonymous. Editorial Board. Science of Computer Programming, 48(1): ii, July 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Watson:2003:NAC

[695] Bruce W. Watson. A new algorithm for the construction of minimal acyclic DFAs. Science of Computer Programming, 48(2–3):81–97, August/ September 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Watson:2003:BMS

[696] Bruce W. Watson and Richard E. Watson. A Boyer–Moore-style algorithm for regular expression pattern matching. Science of Computer Programming, 48(2–3):99–117, August/ September 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wah:2003:ACE

[697] K. S. How Tai Wah. An analysis of the coupling effect I: single test data. Science of Computer Programming, 48(2– 3):119–161, August/September 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pedreschi:2003:LPA

[698] Dino Pedreschi and Salvatore Ruggieri. On logic programs that always succeed. Science of Computer Programming, 48(2–3):163–196, August/ September 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:AIc

[699] Anonymous. Author index. Science of Computer Programming, 48(2–3): 197, August/September 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:EBf

[700] Anonymous. (ed. board). Science of Computer Programming, 48(2–3):CO2, August/September 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Burgess:2003:TSA

[701] Mark Burgess. On the theory of system administration. Science of Computer Programming, 49(1–3):1–46, December 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rees:2003:ASI

[702] D. Ll. L. Rees, K. Stephenson, and J. V. Tucker. The algebraic structure of interfaces. Science of Computer Programming, 49(1–3):47–88, December 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Comini:2003:AIB

[703] Marco Comini, Roberta Gori, Giorgio Levi, and Paolo Volpe. Abstract interpretation based verification of logic programs. Science of Computer Programming, 49(1–3):89–123, December 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Curtis:2003:CGA

[704] S. A. Curtis. The classification of greedy algorithms. Science of Computer Programming, 49(1–3):125–157, December 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:AId

[705] Anonymous. Author index. Science of Computer Programming, 49 (1–3):159, December 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2003:EBg

[706] Anonymous. Editorial Board. Science of Computer Programming, 49 (1–3):ii, December 2003. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Degano:2004:P

[707] Pierpaolo Degano. Preface. Science of Computer Programming, 50(1–3):1, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Meadows:2004:OSM

[708] Catherine Meadows. Ordering from Satan’s menu: a survey of requirements specification for formal analysis of cryptographic protocols. Science of Computer Programming, 50 (1–3):3–22, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gorrieri:2004:SFR

[709] Roberto Gorrieri and Fabio Martinelli. A simple framework for real-time cryptographic protocol analysis with compositional proof rules. Science of Computer Programming, 50(1–3):23– 49, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Comon-Lundh:2004:SPT

[710] Hubert Comon-Lundh and Véronique Cortier. Security properties: two agents are sufficient. Science of Computer Programming, 50(1–3):51– 71, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Giambiagi:2004:SIS

[711] Pablo Giambiagi and Mads Dam. On the secure implementation of security protocols. Science of Computer Programming, 50(1–3):73–99, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Yu:2004:BCL

[712] Dachuan Yu, Nadeem A. Hamid, and Zhong Shao. Building certified libraries for PCC: dynamic storage allocation. Science of Computer Programming, 50 (1–3):101–127, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Reddy:2004:CDR

[713] Uday S. Reddy and Hongseok Yang. Correctness of data representations involving heap data structures. Science of Computer Programming, 50 (1–3):129–160, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ohori:2004:RAP

[714] Atsushi Ohori. Register allocation by proof transformation. Science of Computer Programming, 50(1–3):161–187, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Haack:2004:TES

[715] Christian Haack and J. B. Wells. Type error slicing in implicitly typed higherorder languages. Science of Computer Programming, 50(1–3):189–224, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lhoussaine:2004:TID

[716] Cédric Lhoussaine. Type inference for a distributed π-calculus. Science of Computer Programming, 50 (1–3):225–251, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Flanagan:2004:ASM

[717] Cormac Flanagan. Automatic software model checking via constraint logic. Science of Computer Programming, 50 (1–3):253–270, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2004:AI

[718] Anonymous. Author index. Science of Computer Programming, 50(1–3):271, March 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Boiten:2004:F

[719] Eerke Boiten and Bernhard Möller. Foreword. Science of Computer Programming, 51(1–2):1–2, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kozen:2004:SRD

[720] Dexter Kozen. Some results in dynamic model theory. Science of Computer Programming, 51(1–2):3–22, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vonWright:2004:TRA

[721] J. von Wright. Towards a refinement algebra. Science of Computer Programming, 51(1–2):23–45, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hesselink:2004:UEV

[722] Wim H. Hesselink. Using eternity variables to specify and prove a serializable database interface. Science of Computer Programming, 51(1–2):47– 85, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mu:2004:TAI

[723] Shin-Cheng Mu and Richard Bird. Theory and applications of inverting functions as folds. Science of Computer Programming, 51(1–2):87–116, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hinze:2004:TID

[724] Ralf Hinze, Johan Jeuring, and Andres Löh. Type-indexed data types. Science of Computer Programming, 51 (1–2):117–151, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Backhouse:2004:SAI

[725] Kevin Backhouse and Roland Backhouse. Safety of abstract interpretations for free, via logical relations and Galois connections. Science of Computer Programming, 51(1–2):153–196, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2004:EBa

[726] Anonymous. Editorial board. Science of Computer Programming, 51(1–2):??, May 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Burgess:2004:CIE

[727] Mark Burgess. Configurable immunity for evolving human–computer systems. Science of Computer Programming, 51 (3):197–213, June 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Giegerich:2004:DDP

[728] Robert Giegerich, Carsten Meyer, and Peter Steffen. A discipline of dynamic programming over sequence data. Science of Computer Programming, 51 (3):215–263, June 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wile:2004:LLR

[729] David Wile. Lessons learned from real DSL experiments. Science of Computer Programming, 51(3):265– 290, June 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vereijken:2004:PAT

[730] J. J. Jan Joris Vereijken. Process algebra with timing: J. C. M. Baeten and C. A. Middelburg (Eds); Springer, Berlin, 2002, pp. xx + 287, ISBN 3-540-43447-X. Science of Computer Programming, 51(3):291–292, June 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2004:IEB

[731] Anonymous. IFC (Ed. board). Science of Computer Programming, 51(3):??, June 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lammel:2004:TE

[732] Ralf Lämmel. Transformations everywhere. Science of Computer Programming, 52(1–3):1–8, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kniesel:2004:SCR

[733] Günter Kniesel and Helge Koch. Static composition of refactorings. Science of Computer Programming, 52(1–3):9–51, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Borba:2004:ARO

[734] Paulo Borba, Augusto Sampaio, Ana Cavalcanti, and Márcio Cornélio. Algebraic reasoning for object-oriented programming. Science of Computer Programming, 52(1–3):53–100, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Erwig:2004:MFP

[735] Martin Erwig and Deling Ren. Monadification of functional programs. Science of Computer Programming, 52(1– 3):101–129, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zdun:2004:SIE

[736] Uwe Zdun. Supporting incremental and experimental software evolution by runtime method transformations. Science of Computer Programming, 52(1– 3):131–163, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Winter:2004:TCH

[737] Victor L. Winter and Mahadevan Subramaniam. The transient combinator, higher-order strategies, and the distributed data problem. Science of Computer Programming, 52(1–3):165– 212, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ward:2004:PSR

[738] M. P. Ward. Pigs from sausages? Reengineering from assembler to C via FermaT transformations. Science of Computer Programming, 52(1– 3):213–255, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanEngelen:2004:AVC

[739] Robert van Engelen, David Whalley, and Xin Yuan. Automatic validation of code-improving transformations on low-level program representations. Science of Computer Programming, 52(1– 3):257–280, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Glew:2004:TSM

[740] Neal Glew and Jens Palsberg. Typesafe method inlining. Science of Computer Programming, 52(1–3):281–306, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jones:2004:TIS

[741] Neil D. Jones. Transformation by interpreter specialisation. Science of Computer Programming, 52(1–3):307–339, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Consel:2004:TTP

[742] Charles Consel, Julia L. Lawall, and Anne-Françoise Le Meur. A tour of Tempo: a program specializer for the C language. Science of Computer Programming, 52(1–3):341–370, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2004:EBb

[743] Anonymous. Editorial Board. Science of Computer Programming, 52(1–3):??, August 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mahmoud:2004:PEJ

[744] Qusay H. Mahmoud. Practice and experience with Java in education. Science of Computer Programming, 53(1):1–2, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Goldman:2004:IEB

[745] Kenneth J. Goldman. An interactive environment for beginning Java programmers. Science of Computer Programming, 53(1):3–24, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bieg:2004:ETD

[746] Claudia Bieg and Stephan Diehl. Educational and technical design of a Webbased interactive tutorial on programming in Java. Science of Computer Programming, 53(1):25–36, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bonifaci:2004:JBS

[747] Vincenzo Bonifaci, Camil Demetrescu, Irene Finocchi, and Luigi Laura. A Java-based system for building animated presentations over the Web. Science of Computer Programming, 53 (1):37–49, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gray:2004:JBA

[748] Jeff Gray. A Java-based approach for teaching principles of adaptive and evolvable software. Science of Computer Programming, 53(1):51–69, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kortenkamp:2004:GTW

[749] Ulrich Kortenkamp and Dirk Materlik. Geometry teaching in wireless classroom environments using Java and J2ME. Science of Computer Programming, 53(1):71–85, October 2004.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Liu:2004:JBD

[750] Zehua Liu, Hai Yu, Ee-Peng Lim, Ming Yin, Dion Hoe-Lian Goh, Yin-Leng Theng, and Wee-Keong Ng. A Javabased digital library portal for geography education. Science of Computer Programming, 53(1):87–105, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Elnagar:2004:GPP

[751] Ashraf Elnagar and Leena Lulu. A global path planning Java-based system for autonomous mobile robots. Science of Computer Programming, 53 (1):107–122, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2004:EBc

[752] Anonymous. Editorial Board. Science of Computer Programming, 53(1): ??, October 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Burgess:2004:TSA

[753] Mark Burgess. Topics in system administration. Science of Computer Programming, 53(2):123–124, November 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Qie:2004:USG

[754] Xiaohu Qie and Sanjai Narain. Using service grammar to diagnose BGP configuration errors. Science of Computer Programming, 53(2):125–141, November 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wang:2004:SBB

[755] Yi-Min Wang, Chad Verbowski, John Dunagan, Yu Chen, Helen J. Wang, Chun Yuan, and Zheng Zhang. Strider: a black-box, state-based approach to change and configuration management and support. Science of Computer Programming, 53(2):143–164, November 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Steinder:2004:SFL

[756] Małgorzata Steinder and Adarshpal S. Sethi. A survey of fault localization techniques in computer networks. Science of Computer Programming, 53(2): 165–194, November 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Canright:2004:RN

[757] Geoffrey Canright and Kenth Engø-Monsen. Roles in networks. Science of Computer Programming, 53(2): 195–214, November 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Couch:2004:ORN

[758] Alva Couch and Yizhan Sun. On observed reproducibility in network configuration management. Science of Computer Programming, 53(2): 215–253, November 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2004:EBd

[759] Anonymous. Editorial Board. Science of Computer Programming, 53 (2):??, November 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bosch:2004:SVM

[760] Jan Bosch. Software variability management. Science of Computer Programming, 53(3):255–258, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schmid:2004:CAF

[761] Klaus Schmid and Isabel John. A customizable approach to full lifecycle variability management. Science of Computer Programming, 53(3): 259–284, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanderHoek:2004:DTP

[762] André van der Hoek. Design-time product line architectures for any-time variability. Science of Computer Programming, 53(3):285–304, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Webber:2004:MVS

[763] Diana L. Webber and Hassan Gomaa. Modeling variability in software product lines with the variation point model. Science of Computer Programming, 53(3):305–331, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Beuche:2004:VMF

[764] Danilo Beuche, Holger Papajewski, and Wolfgang Schröder-Preikschat. Variability management with feature models. Science of Computer Programming, 53(3):333–352, December 2004.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Goedicke:2004:DRV

[765] Michael Goedicke, Carsten Köllmann, and Uwe Zdun. Designing runtime variation points in product line architectures: three cases. Science of Computer Programming, 53(3):353–380, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zhang:2004:XMH

[766] Hongyu Zhang and Stan Jarzabek. XVCL: a mechanism for handling variants in software product lines. Science of Computer Programming, 53(3): 381–407, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Roubtsova:2004:BIU

[767] Ella Roubtsova and Serguei Roubtsov. Behavioural inheritance in the UML to model software product lines. Science of Computer Programming, 53(3): 409–434, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2004:EBe

[768] Anonymous. Editorial Board. Science of Computer Programming, 53 (3):??, December 2004. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Power:2005:E

[769] James F. Power and John T. Waldron. Editorial. Science of Computer Programming, 54(1):1–2, January 2005.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Boldi:2005:MSJ

[770] Paolo Boldi and Sebastiano Vigna. Mutable strings in Java: design, implementation and lightweight text-search algorithms. Science of Computer Programming, 54(1):3–23, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rajsbaum:2005:OOA

[771] Sergio Rajsbaum and Elisa Viso. Object-oriented algorithm analysis and design with Java. Science of Computer Programming, 54(1):25–47, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Seigneur:2005:CTK

[772] Jean-Marc Seigneur and Christian Damsgaard Jensen. The Claim Tool Kit for ad hoc recognition of peer entities. Science of Computer Programming, 54 (1):49–71, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cabri:2005:ERB

[773] Giacomo Cabri, Luca Ferrari, and Letizia Leonardi. Exploiting runtime bytecode manipulation to add roles to Java agents. Science of Computer Programming, 54(1):73–98, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

ONeill:2005:IAS

[774] Ian O’Neill, Philip Hanna, Xingkun Liu, Des Greer, and Michael McTear. Implementing advanced spoken dialogue management in Java. Science of Computer Programming, 54 (1):99–124, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shah:2005:SET

[775] Vijay Shah, Nicholas Younan, Torey Alford, and Anthony Skjellum. A spectral estimation toolkit for Java applications. Science of Computer Programming, 54(1):125–142, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBa

[776] Anonymous. Editorial Board. Science of Computer Programming, 54(1): ??, January 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Klusener:2005:AMD

[777] A. S. Klusener, R. Lämmel, and C. Verhoef. Architectural modifications to deployed software. Science of Computer Programming, 54(2–3):143–211, February 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Broy:2005:SME

[778] Manfred Broy. A semantic and methodological essence of message sequence charts. Science of Computer Programming, 54(2–3):213–256, February 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nishimura:2005:XST

[779] Susumu Nishimura and Keisuke Nakano. XML stream transformer generation through program composition and dependency analysis. Science of Computer Programming, 54(2– 3):257–290, February 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Santone:2005:CLI

[780] Antonella Santone and Gigliola Vaglini. Compositionality and locality for improving model checking in the selective mu-calculus. Science of Computer Programming, 54(2–3):291–311, February 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vereijken:2005:EBR

[781] Jan Joris Vereijken. Erratum to “Book review: Process algebra with timing” [Sci. Comput. Programming 51 (2004) 291–292]. Science of Computer Programming, 54(2–3):313, February 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBb

[782] Anonymous. Editorial Board. Science of Computer Programming, 54 (2–3):??, February 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

deBoer:2005:P

[783] Frank de Boer and Marcello Bonsangue. Preface. Science of Computer Programming, 55(1–3):1–2, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Arbab:2005:ABT

[784] Farhad Arbab. Abstract Behavior Types: a foundation model for components and their composition. Science of Computer Programming, 55 (1–3):3–52, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Breunesse:2005:FMS

[785] C.-B. Breunesse, N. Cataño, M. Huisman, and B. Jacobs. Formal methods for smart cards: an experience report. Science of Computer Programming, 55 (1–3):53–80, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Damm:2005:DTU

[786] Werner Damm, Bernhard Josko, Amir Pnueli, and Angelika Votintseva. A discrete-time UML semantics for concurrency and communication in safetycritical applications. Science of Computer Programming, 55(1–3):81–115, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Damm:2005:LLL

[787] Werner Damm and Bernd Westphal. Live and let die: LSC based verification of UML models. Science of Computer Programming, 55(1–3):117–159, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gossler:2005:CCB

[788] Gregor Gössler and Joseph Sifakis. Composition for component-based modeling. Science of Computer Programming, 55(1–3):161–183, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Leavens:2005:HDJ

[789] Gary T. Leavens, Yoonsik Cheon, Curtis Clifton, Clyde Ruby, and David R.

Cok. How the design of JML accommodates both runtime assertion checking and formal verification. Science of Computer Programming, 55 (1–3):185–208, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Leino:2005:GET

[790] K. Rustan M. Leino, Todd Millstein, and James B. Saxe. Generating error traces from verification-condition counterexamples. Science of Computer Programming, 55(1–3):209–226, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Olderog:2005:SPI

[791] Ernst-Rüdiger Olderog and Heike Wehrheim. Specification and (property) inheritance in CSP-OZ. Science of Computer Programming, 55 (1–3):227–257, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Plosila:2005:ASS

[792] J. Plosila, K. Sere, and M. Waldén. Asynchronous system synthesis. Science of Computer Programming, 55 (1–3):259–288, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBc

[793] Anonymous. Editorial Board. Science of Computer Programming, 55(1–3):??, March 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pulvermuller:2005:NSC

[794] Elke Pulvermüller, Gerhard Goos, and Uwe Aßmann. New software composition concepts. Science of Computer Programming, 56(1–2):1–4, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kantorowitz:2005:UCC

[795] Eliezer Kantorowitz and Alexander Lyakas. Use-case components for interactive information systems. Science of Computer Programming, 56(1–2):5–21, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Arato:2005:ECB

[796] Péter Arató, Zoltán Ádám Mann, and András Orbán. Extending componentbased design with hardware components. Science of Computer Programming, 56(1–2):23–39, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Frohlich:2005:SSC

[797] Peter H. Fröhlich, Andreas Gal, and Michael Franz. Supporting software composition at the programming language level. Science of Computer Programming, 56(1–2):41–57, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lumpe:2005:FBM

[798] M. Lumpe and J.-G. Schneider. A form-based meta-model for software composition. Science of Computer Programming, 56(1–2):59–78, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Assmann:2005:ASA

[799] Uwe Aßmann. Architectural styles for active documents. Science of Computer Programming, 56(1–2):79– 98, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Washizaki:2005:TAC

[800] Hironori Washizaki and Yoshiaki Fukazawa. A technique for automatic component extraction from objectoriented programs by refactoring. Science of Computer Programming, 56 (1–2):99–116, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Scheben:2005:HCI

[801] Ursula Scheben. Hierarchical composition of industrial components. Science of Computer Programming, 56 (1–2):117–139, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gross:2005:SCT

[802] H.-G. Gross, M. Melideo, and A. Sillitti. Self-certification and trust in component procurement. Science of Computer Programming, 56(1–2):141–156, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Liu:2005:ASC

[803] Xiaodong Liu, Beihu Wang, and Jon Kerridge. Achieving seamless component composition through scenariobased deep adaptation and generation. Science of Computer Programming, 56 (1–2):157–170, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

daSilva:2005:CSA

[804] Leandro Dias da Silva and Angelo Perkusich. Composition of software artifacts modelled using Colored Petri nets. Science of Computer Programming, 56(1–2):171–189, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Volgyesi:2005:SCV

[805] P. Völgyesi, M. Maróti, S. Dóra, E. Osses, and Á. Lédeczi. Software composition and verification for sensor networks. Science of Computer Programming, 56(1–2):191–210, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vincenzi:2005:CTJ

[806] A. M. R. Vincenzi, J. C. Maldonado, W. E. Wong, and M. E. Delamaro. Coverage testing of Java programs and components. Science of Computer Programming, 56(1–2):211–230, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Porter:2005:SBP

[807] Ronald Porter, James O. Coplien, and Tiffany Winn. Sequences as a basis for pattern language composition. Science of Computer Programming, 56 (1–2):231–249, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBd

[808] Anonymous. Editorial Board. Science of Computer Programming, 56(1–2):??, April 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Groote:2005:MCP

[809] J. F. Groote and T. A. C. Willemse. Model-checking processes with data. Science of Computer Programming, 56 (3):251–273, May/June 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Verhoef:2005:QAO

[810] C. Verhoef. Quantitative aspects of outsourcing deals. Science of Computer Programming, 56(3):275–313, May/ June 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Verhoef:2005:QVI

[811] C. Verhoef. Quantifying the value of IT-investments. Science of Computer Programming, 56(3):315–342, May/ June 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cortellessa:2005:RCS

[812] V. Cortellessa, D. Del Gobbo, M. Shereshevsky, J. Desharnais, and A. Mili. Relational characterizations of system fault tolerance. Science of Computer Programming, 56(3): 343–361, May/June 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ponsini:2005:RIP

[813] Olivier Ponsini, Carine Fédèle, and Emmanuel Kounalis. Rewriting of imperative programs into logical equations. Science of Computer Programming, 56(3):363–401, May/June 2005.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBe

[814] Anonymous. Editorial Board. Science of Computer Programming, 56 (3):??, May/June 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chung:2005:SSA

[815] Lawrence Chung and Nary Subramanian. System and software architectures. Science of Computer Programming, 57(1):1–4, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chen:2005:BEP

[816] Hong-Mei Chen, Rick Kazman, and Aditya Garg. BITAM: an engineeringprincipled method for managing misalignments between business and IT architectures. Science of Computer Programming, 57(1):5–26, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shin:2005:SHC

[817] Michael E. Shin. Self-healing components in robust software architecture for concurrent and distributed systems. Science of Computer Programming, 57(1):27–44, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sancho:2005:EPA

[818] Pere P. Sancho, Carlos Juiz, and Ramon Puigjaner. Evaluating the performance of architectures in MAS-COT. Science of Computer Programming, 57(1):45–60, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Murakami:2005:CCS

[819] Masaki Murakami. A calculus for concurrent system with higher-order streaming communication. Science of Computer Programming, 57(1):61–72, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Coallier:2005:PLE

[820] François Coallier and Roger Champagne. A Product Line engineering practices model. Science of Computer Programming, 57(1):73–87, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cooper:2005:PMA

[821] Kendra Cooper, Lirong Dai, and Yi Deng. Performance modeling and analysis of software architectures: an aspect-oriented UML based approach. Science of Computer Programming, 57(1):89–108, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Maheshwari:2005:SAC

[822] Piyush Maheshwari and Albert Teoh. Supporting ATAM with a collaborative Web-based software architecture evaluation tool. Science of Computer Programming, 57(1):109–128, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBf

[823] Anonymous. Editorial Board. Science of Computer Programming, 57(1): ??, July 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Veerman:2005:TLC

[824] Niels Veerman. Towards lightweight checks for mass maintenance transformations. Science of Computer Programming, 57(2):129–163, August 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jagannathan:2005:TOC

[825] Suresh Jagannathan, Jan Vitek, Adam Welc, and Antony Hosking. A transactional object calculus. Science of Computer Programming, 57 (2):164–186, August 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nunez:2005:FSM

[826] Manuel Núñez, Ismael Rodríguez, and Fernando Rubio. Formal specification of multi-agent e-barter systems. Science of Computer Programming, 57 (2):187–216, August 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Denti:2005:MPJ

[827] Enrico Denti, Andrea Omicini, and Alessandro Ricci. Multi-paradigm Java–Prolog integration in tuProlog. Science of Computer Programming, 57 (2):217–250, August 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642305000158.

Anonymous:2005:EBg

[828] Anonymous. Editorial Board. Science of Computer Programming, 57(2):??, August 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ertl:2005:AIV

[829] M. Anton Ertl. Advances in interpreters, virtual machines and emulators. Science of Computer Programming, 57(3):251–252, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Liu:2005:EJM

[830] Hanbing Liu and J. Strother Moore. Executable JVM model for analytical reasoning: a study. Science of Computer Programming, 57(3):253–274, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Franz:2005:PVM

[831] Michael Franz, Deepak Chandra, Andreas Gal, Vivek Haldar, Christian W. Probst, Fermín Reig, and Ning Wang. A portable virtual machine target for proof-carrying code. Science of Computer Programming, 57(3):275–294, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Latendresse:2005:GFI

[832] Mario Latendresse and Marc Feeley. Generation of fast interpreters for Huffman compressed bytecode. Science of Computer Programming, 57(3): 295–317, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gregg:2005:CVR

[833] David Gregg, Andrew Beatty, Kevin Casey, Brian Davis, and Andy Nisbet. The case for virtual register machines. Science of Computer Programming, 57(3):319–338, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Whaley:2005:JVM

[834] John Whaley. Joeq: a virtual machine and compiler infrastructure. Science of Computer Programming, 57(3): 339–356, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Palacz:2005:ECI

[835] K. Palacz, J. Baker, C. Flack, C. Grothoff, H. Yamauchi, and J. Vitek. Engineering a common intermediate representation for the Ovm framework. Science of Computer Programming, 57(3):357–378, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBh

[836] Anonymous. Editorial Board. Science of Computer Programming, 57 (3):??, September 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cousot:2005:SAS

[837] Radhia Cousot. Static Analysis Symposium 2003. Science of Computer Programming, 58(1–2):1–2, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Abadi:2005:CAV

[838] Martín Abadi and Bruno Blanchet. Computer-assisted verification of a protocol for certified email. Science of Computer Programming, 58(1–2):3–27, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bagnara:2005:PWO

[839] Roberto Bagnara, Patricia M. Hill, Elisa Ricci, and Enea Zaffanella. Precise widening operators for convex polyhedra. Science of Computer Programming, 58(1–2):28–56, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Field:2005:TVA

[840] J. Field, D. Goyal, G. Ramalingam, and E. Yahav. Typestate verification: Abstraction techniques and complexity results. Science of Computer Programming, 58(1–2):57–82, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Guyer:2005:ECC

[841] Samuel Z. Guyer and Calvin Lin. Error checking with client-driven pointer analysis. Science of Computer Programming, 58(1–2):83–114, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hermenegildo:2005:IPD

[842] Manuel V. Hermenegildo, Germán Puebla, Francisco Bueno, and Pedro López-García. Integrated program debugging, verification, and optimization using abstract interpretation (and the Ciao system preprocessor). Science of Computer Programming, 58(1– 2):115–140, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lee:2005:SIS

[843] Oukseh Lee, Hongseok Yang, and Kwangkeun Yi. Static insertion of safe and effective memory reuse commands into ML-like programs. Science of Computer Programming, 58(1– 2):141–178, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Monniaux:2005:AIP

[844] David Monniaux. Abstract interpretation of programs as Markov decision processes. Science of Computer Programming, 58(1–2):179–205, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Reps:2005:WPS

[845] Thomas Reps, Stefan Schwoon, Somesh Jha, and David Melski. Weighted pushdown systems and their application to interprocedural dataflow analysis. Science of Computer Programming, 58(1– 2):206–263, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shaham:2005:ELT

[846] Ran Shaham, Eran Yahav, Elliot K. Kolodner, and Mooly Sagiv. Establishing local temporal heap safety properties with applications to compiletime memory management. Science of Computer Programming, 58(1–2): 264–289, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBi

[847] Anonymous. Editorial Board. Science of Computer Programming, 58 (1–2):??, October 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moir:2005:CSJ

[848] Mark Moir, Nir Shavit, and Jan Vitek. Concurrency and synchronization in Java programs. Science of Computer Programming, 58(3):291–292, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lea:2005:JUC

                                   [849] Doug Lea. The java.util.concurrent

synchronizer framework. Science of Computer Programming, 58(3): 293–309, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cole:2005:SST

[850] Christopher Cole and Maurice Herlihy. Snapshots and software transactional memory. Science of Computer Programming, 58(3):310–324, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Harris:2005:ESE

[851] Tim Harris. Exceptions and side-effects in atomic blocks. Science of Computer Programming, 58(3):325–343, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shanneb:2005:ERP

[852] Abdelsalam Shanneb, John Potter, and James Noble. Exclusion requirements and potential concurrency for composite objects. Science of Computer Programming, 58(3):344–365, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rose:2005:DIP

[853] James Rose, Nikhil Swamy, and Michael Hicks. Dynamic inference of polymorphic lock types. Science of Computer Programming, 58 (3):366–383, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Greenhouse:2005:OAE

[854] Aaron Greenhouse, T. J. Halloran, and William L. Scherlis. Observations on the assured evolution of concurrent Java programs. Science of Computer Programming, 58(3):384–411, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2005:EBj

[855] Anonymous. Editorial Board. Science of Computer Programming, 58 (3):??, December 2005. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aleksy:2006:F

[856] Markus Aleksy, Ralf Gitzel, and John Waldron. Foreword. Science of Computer Programming, 59(1–2):1–3, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mitchell:2006:SIC

[857] Áine Mitchell and James F. Power. A study of the influence of coverage on the relationship between static and dynamic coupling metrics. Science of Computer Programming, 59 (1–2):4–25, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Allen:2006:SIG

[858] Eric E. Allen and Robert Cartwright. Safe instantiation in Generic Java. Science of Computer Programming, 59 (1–2):26–37, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Debbabi:2006:SDC

[859] Mourad Debbabi, Abdelouahed Gherbi, Azzam Mourad, and Hamdi Yahyaoui. A selective dynamic compiler for embedded Java virtual machines targeting ARM processors. Science of Computer Programming, 59(1–2):38–63, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nagpurkar:2006:PBV

[860] Priya Nagpurkar and Chandra Krintz. Phase-based visualization and analysis of Java programs. Science of Computer Programming, 59(1–2):64–81, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wells:2006:NIL

[861] George C. Wells. New and improved: Linda in Java. Science of Computer Programming, 59(1–2):82–96, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Puder:2006:XAW

[862] Arno Puder. XML11 — an abstract windowing protocol. Science of Computer Programming, 59(1–2):97– 108, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ayres:2006:VAC

[863] Tony Ayres and Brian Nolan. Voice activated command and control with speech recognition over WiFi. Science of Computer Programming, 59(1– 2):109–126, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cabri:2006:ASP

[864] Giacomo Cabri, Luca Ferrari, and Letizia Leonardi. Applying security policies through agent roles: a JAAS based approach. Science of Computer Programming, 59(1–2):127–146, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cunningham:2006:UCP

[865] H. Conrad Cunningham, Yi Liu, and Cuihua Zhang. Using classic problems to teach Java framework design. Science of Computer Programming, 59(1– 2):147–169, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBa

[866] Anonymous. Editorial Board. Science of Computer Programming, 59 (1–2):??, January 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Corradini:2006:RFS

[867] Flavio Corradini, Paola Inverardi, and Alexander L. Wolf. On relating functional specifications to architectural specifications: a case study. Science of Computer Programming, 59 (3):171–208, February 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Giannini:2006:SAA

[868] Paola Giannini, Davide Sangiorgi, and Andrea Valente. Safe Ambients: Abstract machine and distributed implementation. Science of Computer Programming, 59(3):209–249, February 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

DallOsso:2006:CAS

[869] Aldo Dall’Osso. Computer algebra systems as mathematical optimizing compilers. Science of Computer Programming, 59(3):250–273, February 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Burgess:2006:PAD

[870] Mark Burgess. Probabilistic anomaly detection in distributed computer networks. Science of Computer Programming, 60(1):1–26, March 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Papadopoulos:2006:IFS

[871] George A. Papadopoulos, Aristos Stavrou, and Odysseas Papapetrou. An implementation framework for Software Architectures based on the coordination paradigm. Science of Computer Programming, 60(1):27–67, March 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sohn:2006:COM

[872] So Young Sohn and Ji Soo Lee. Cost of ownership model for a CRM system. Science of Computer Programming, 60 (1):68–81, March 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kondoh:2006:G

[873] Hidetaka Kondoh and Kokichi Futatsugi. goto. Science of Computer Programming, 60(1):82–116, March 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBb

[874] Anonymous. Editorial Board. Science of Computer Programming, 60(1):??, March 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Balmas:2006:ISI

[875] Françoise Balmas and Kostas Kontogiannis. Introduction to the special issue on software analysis, evolution and reengineering. Science of Computer Programming, 60(2):117– 120, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Umphress:2006:SVI

[876] David A. Umphress, T. Dean Hendrix, James H. Cross II, and Saeed Maghsoodloo. Software visualizations for improving and measuring the comprehensibility of source code. Science of Computer Programming, 60 (2):121–133, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mudiam:2006:SIL

[877] Sudhakiran V. Mudiam, Gerald C. Gannod, and Timothy E. Lindquist. Synthesizing and integrating legacy components as services using adapters. Science of Computer Programming, 60 (2):134–148, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Holt:2006:GGB

[878] Richard C. Holt, Andy Schürr, Susan Elliott Sim, and Andreas Winter. GXL: a graph-based standard exchange format for reengineering. Science of Computer Programming, 60 (2):149–170, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Koschke:2006:RAC

[879] Rainer Koschke, Gerardo Canfora, and Jörg Czeranski. Revisiting the ΔIC approach to component recovery. Science of Computer Programming, 60 (2):171–188, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ung:2006:DBT

[880] David Ung and Cristina Cifuentes. Dynamic binary translation using runtime feedbacks. Science of Computer Programming, 60(2):189–204, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanDeursen:2006:DSS

[881] Arie van Deursen and Leon Moonen. Documenting software systems using types. Science of Computer Programming, 60(2):205–220, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBc

[882] Anonymous. Editorial Board. Science of Computer Programming, 60(2):??, April 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Charpentier:2006:CI

[883] Michel Charpentier. Composing invariants. Science of Computer Programming, 60(3):221–243, May 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Maurer:2006:TCI

[884] Ward Douglas Maurer. A theory of computer instructions. Science of Computer Programming, 60(3):244– 273, May 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Biernacki:2006:SDE

[885] Dariusz Biernacki, Olivier Danvy, and Chung chieh Shan. On the static and dynamic extents of delimited continuations. Science of Computer Programming, 60(3):274–297, May 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chung:2006:QSS

[886] Lawrence Chung and Nary Subramanian. Quality system and software architectures. Science of Computer Programming, 61(1):1–3, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Garg:2006:IDE

[887] Aditya Garg, Rick Kazman, and Hong-Mei Chen. Interface descriptions for enterprise architecture. Science of Computer Programming, 61(1):4–15, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shin:2006:DAS

[888] Michael E. Shin, Yan Xu, Fernando Paniagua, and Jung Hoon An. Detection of anomalies in software architecture with connectors. Science of Computer Programming, 61(1):16–26, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lera:2006:PRO

[889] Isaac Lera, Carlos Juiz, and Ramon Puigjaner. Performance-related ontologies and Semantic Web applications for on-line performance assessment of intelligent systems. Science of Computer Programming, 61(1):27–37, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Murakami:2006:FMC

[890] Masaki Murakami. A formal model of concurrent systems based on bipartite directed acyclic graph. Science of Computer Programming, 61(1):38–47, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sterritt:2006:NGS

[891] Roy Sterritt, Christopher A. Rouff, Michael G. Hinchey, James L. Rash, and Walt Truszkowski. Next generation system and software architectures: Challenges from future NASA exploration missions. Science of Computer Programming, 61(1):48–57, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dai:2006:MPA

[892] Lirong Dai and Kendra Cooper. Modeling and performance analysis for security aspects. Science of Computer Programming, 61(1):58–71, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBd

[893] Anonymous. Editorial Board. Science of Computer Programming, 61(1):??, June 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brogi:2006:P

[894] Antonio Brogi, Jean-Marie Jacquet, Jeff Kramer, and Ernesto Pimentel. Preface. Science of Computer Programming, 61(2):73–74, July 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Baier:2006:MCC

[895] Christel Baier, Marjan Sirjani, Farhad Arbab, and Jan Rutten. Modeling component connectors in Reo by constraint automata. Science of Computer Programming, 61(2):75–113, July 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lopes:2006:AMS

[896] Antónia Lopes and José Luiz Fiadeiro. Adding mobility to software architectures. Science of Computer Programming, 61(2):114–135, July 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brogi:2006:SSA

[897] Antonio Brogi, Carlos Canal, and Ernesto Pimentel. On the semantics of software adaptation. Science of Computer Programming, 61 (2):136–151, July 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Linden:2006:ETC

[898] I. Linden, J.-M. Jacquet, K. De Bosschere, and A. Brogi. On the expressiveness of timed coordination models. Science of Computer Programming, 61(2):152–187, July 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hedin:2006:P

[899] Görel Hedin and Eric Van Wyk. Preface. Science of Computer Programming, 61(3):189, August 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cordy:2006:TST

[900] James R. Cordy. The TXL source transformation language. Science of Computer Programming, 61(3):190– 210, August 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Begel:2006:XAA

[901] Andrew Begel and Susan L. Graham. XGLR—an algorithm for ambiguity in programming languages. Science of Computer Programming, 61 (3):211–227, August 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Johnstone:2006:EGP

[902] Adrian Johnstone, Elizabeth Scott, and Giorgios Economopoulos. Evaluating GLR parsing algorithms. Science of Computer Programming, 61 (3):228–244, August 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vandenBrand:2006:AE

[903] Mark van den Brand, Jørgen Iversen, and Peter D. Mosses. An Action Environment. Science of Computer Programming, 61(3):245–264, August 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBe

[904] Anonymous. Editorial Board. Science of Computer Programming, 61(3):??, August 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lengauer:2006:P

[905] Christian Lengauer and Walid Taha. Preface. Science of Computer Programming, 62(1):1–2, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carette:2006:GEC

[906] Jacques Carette. Gaussian Elimination: a case study in efficient genericity with MetaOCaml. Science of Computer Programming, 62(1):3–24, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cohen:2006:SPG

[907] Albert Cohen, Sébastien Donadio, Maria-Jesus Garzaran, Christoph Herrmann, Oleg Kiselyov, and David Padua. In search of a program generator to implement generic transformations for high-performance computing. Science of Computer Programming, 62 (1):25–46, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Herrmann:2006:CPE

[908] Christoph A. Herrmann and Tobias Langhammer. Combining partial evaluation and staged interpretation in the implementation of domain-specific languages. Science of Computer Programming, 62(1):47–65, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

League:2006:MSP

[909] Christopher League. MetaOCaml server pages: Web publishing as staged computation. Science of Computer Programming, 62(1):66–84, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Neverov:2006:EOO

[910] Gregory Neverov and Paul Roe. Experiences with an object-oriented, multistage language. Science of Computer Programming, 62(1):85–94, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBf

[911] Anonymous. Editorial Board. Science of Computer Programming, 62 (1):??, September 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jones:2006:FPM

[912] Richard Jones. Five perspectives on modern memory management: Systems, hardware and theory. Science of Computer Programming, 62 (2):95–97, October 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sagonas:2006:EMM

[913] Konstantinos Sagonas and Jesper Wilhelmsson. Efficient memory management for concurrent programs that use message passing. Science of Computer Programming, 62(2):98–121, October 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Swamy:2006:SMM

[914] Nikhil Swamy, Michael Hicks, Greg Morrisett, Dan Grossman, and Trevor Jim. Safe manual memory management in Cyclone. Science of Computer Programming, 62(2):122–144, October 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wright:2006:OAM

[915] Greg Wright, Matthew L. Seidl, and Mario Wolczko. An object-aware memory architecture. Science of Computer Programming, 62(2):145–163, October 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Printezis:2006:MGC

[916] Tony Printezis. On measuring garbage collection responsiveness. Science of Computer Programming, 62(2):164– 183, October 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Clinger:2006:LCR

[917] William D. Clinger and Fabio V. Rojas. Linear combinations of radioactive decay models for generational garbage collection. Science of Computer Programming, 62(2):184–203, October 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBg

[918] Anonymous. Editorial Board. Science of Computer Programming, 62 (2):??, October 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Antoniol:2006:SIS

[919] Giuliano Antoniol, Jens Krinke, and Paolo Tonella. Special issue on Source code analysis and manipulation. Science of Computer Programming, 62(3): 205–208, October 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zhang:2006:UST

[920] Hongyu Zhang, Jeremy S. Bradbury, James R. Cordy, and Juergen Dingel. Using source transformation to test and model check implicit-invocation systems. Science of Computer Programming, 62(3):209–227, October 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Binkley:2006:FRB

[921] Dave Binkley, Sebastian Danicic, Tibor Gyimóthy, Mark Harman, Ákos Kiss, and Bogdan Korel. A formalisation of the relationship between forms of program slicing. Science of Computer Programming, 62(3):228–252, October 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Muller:2006:MIL

[922] Peter Müller, Arnd Poetzsch-Heffter, and Gary T. Leavens. Modular invariants for layered object structures. Science of Computer Programming, 62(3): 253–286, October 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Veerman:2006:AMM

[923] Niels Veerman. Automated mass maintenance of a software portfolio. Science of Computer Programming, 62(3): 287–317, October 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBh

[924] Anonymous. Editorial Board. Science of Computer Programming, 62 (3):??, October 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Focardi:2006:GEI

[925] Riccardo Focardi and Gianluigi Zavattaro. Guest Editor’s introduction: Special issue on security issues in coordination models, languages, and systems. Science of Computer Programming, 63 (1):1–2, November 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Focardi:2006:SSD

[926] Riccardo Focardi, Roberto Lucchi, and Gianluigi Zavattaro. Secure shared data-space coordination languages: a process algebraic survey. Science of Computer Programming, 63(1):3–15, November 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aldini:2006:CSP

[927] Alessandro Aldini. Classification of security properties in a Linda-like process algebra. Science of Computer Programming, 63(1):16–38, November 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brogi:2006:CAT

[928] Antonio Brogi, Carlos Canal, and Ernesto Pimentel. Component adaptation through flexible subservicing. Science of Computer Programming, 63 (1):39–56, November 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

DeNicola:2006:CDP

[929] Rocco De Nicola, Daniele Gorla, and Rosario Pugliese. Confining data and processes in global computing applications. Science of Computer Programming, 63(1):57–87, November 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Omicini:2006:ACC

[930] Andrea Omicini, Alessandro Ricci, and Mirko Viroli. Agent Coordination Contexts for the formal specification and enactment of coordination and security policies. Science of Computer Programming, 63(1):88–107, November 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBi

[931] Anonymous. Editorial Board. Science of Computer Programming, 63 (1):??, November 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Harris:2006:SIS

[932] Tim Harris and Doug Lea. Special issue on synchronization and concurrency in object-oriented languages. Science of Computer Programming, 63(2): 109–110, December 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carlstrom:2006:EJP

[933] Brian D. Carlstrom, JaeWoong Chung, Hassan Chafi, Austen McDonald, Chi Cao Minh, Lance Hammond, Christos Kozyrakis, and Kunle Olukotun. Executing Java programs with transactional memory. Science of Computer Programming, 63(2):111– 129, December 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Eugster:2006:CAF

[934] Patrick Eugster and Sebastien Vaucouleur. Composing atomic features. Science of Computer Programming, 63 (2):130–146, December 1, 2006. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Danaher:2006:PEJ

[935] John S. Danaher, I.-Ting Angelina Lee, and Charles E. Leiserson. Programming with exceptions in JCilk. Science of Computer Programming, 63(2): 147–171, December 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cachopo:2006:VBB

[936] João Cachopo and António Rito-Silva. Versioned boxes as the basis for memory transactions. Science of Computer Programming, 63(2):172–185, December 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moss:2006:NTM

[937] J. Eliot B. Moss and Antony L. Hosking. Nested transactional memory: Model and architecture sketches. Science of Computer Programming, 63(2): 186–201, December 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBj

[938] Anonymous. Editorial Board. Science of Computer Programming, 63 (2):??, December 1, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fradet:2006:SIF

[939] Pascal Fradet and Ralf Lämmel. Special issue on foundations of aspectoriented programming. Science of Computer Programming, 63(3):203– 206, December 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dutchyn:2006:SSA

[940] Christopher Dutchyn, David B. Tucker, and Shriram Krishnamurthi. Semantics and scoping of aspects in higherorder languages. Science of Computer Programming, 63(3):207–239, December 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ligatti:2006:TTI

[941] Jay Ligatti, David Walker, and Steve Zdancewic. A type-theoretic interpretation of pointcuts and advice. Science of Computer Programming, 63(3): 240–266, December 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jagadeesan:2006:TPP

[942] Radha Jagadeesan, Alan Jeffrey, and James Riely. Typed parametric polymorphism for aspects. Science of Computer Programming, 63(3):267– 296, December 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Altisen:2006:AOP

[943] K. Altisen, F. Maraninchi, and D. Stauch. Aspect-oriented programming for reactive systems: Larissa, a proposal in the synchronous framework. Science of Computer Programming, 63(3):297–320, December 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Clifton:2006:MIC

[944] Curtis Clifton and Gary T. Leavens. MiniMAO_{1}: an imperative core language for studying aspect-oriented reasoning. Science of Computer Programming, 63(3):321–374, December 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2006:EBk

[945] Anonymous. Editorial Board. Science of Computer Programming, 63 (3):??, December 15, 2006. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Giacobazzi:2007:F

[946] Roberto Giacobazzi. Foreword. Science of Computer Programming, 64 (1):1–2, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Amtoft:2007:LIF

[947] Torben Amtoft and Anindya Banerjee. A logic for information flow analysis with an application to forward slicing of simple imperative programs. Science of Computer Programming, 64 (1):3–28, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schmidt:2007:CLR

[948] David A. Schmidt. A calculus of logical relations for over- and underapproximating static analyses. Science of Computer Programming, 64 (1):29–53, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rodriguez-Carbonell:2007:AGP

[949] E. Rodríguez-Carbonell and D. Kapur. Automatic generation of polynomial invariants of bounded degree using abstract interpretation. Science of Computer Programming, 64 (1):54–75, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Colon:2007:PAR

[950] Michael A. Colón. Polynomial approximations of the relational semantics of imperative programs. Science of Computer Programming, 64 (1):76–96, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gulwani:2007:PTA

[951] Sumit Gulwani and George C. Necula. A polynomial-time algorithm for global value numbering. Science of Computer Programming, 64(1):97–114, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Clariso:2007:OAD

[952] Robert Clarisó and Jordi Cortadella. The octahedron abstract domain. Science of Computer Programming, 64(1): 115–139, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Flanagan:2007:TIA

[953] Cormac Flanagan and Stephen N. Freund. Type inference against races. Science of Computer Programming, 64(1): 140–165, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Thompson:2007:AIC

[954] Sarah Thompson and Alan Mycroft. Abstract interpretation of combinational asynchronous circuits. Science of Computer Programming, 64(1): 166–183, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBa

[955] Anonymous. Editorial Board. Science of Computer Programming, 64 (1):??, January 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jacquet:2007:ISI

[956] Jean-Marie Jacquet and Gian Pietro Picco. Introduction to the special issue. Science of Computer Programming, 64(2):185–186, January 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

DeNicola:2007:GCD

[957] Rocco De Nicola, Daniele Gorla, and Rosario Pugliese. Global computing in a dynamic network of tuple spaces. Science of Computer Programming, 64(2): 187–204, January 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Udzir:2007:CM

[958] Nur Izura Udzir, Alan M. Wood, and Jeremy L. Jacob. Coordination with multicapabilities. Science of Computer Programming, 64(2):205–222, January 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Colman:2007:URB

[959] Alan Colman and Jun Han. Using rolebased coordination to achieve software adaptability. Science of Computer Programming, 64(2):223–245, January 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Russello:2007:EES

[960] Giovanni Russello, Michel R. V. Chaudron, Maarten van Steen, and Ibrahim Bokharouss. An experimental evaluation of self-managing availability in shared data spaces. Science of Computer Programming, 64(2):246– 262, January 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBb

[961] Anonymous. Editorial Board. Science of Computer Programming, 64 (2):??, January 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Broy:2007:ECA

[962] Manfred Broy. Editorial comment to ‘About “trivial” software patents: The IsNot case’. Science of Computer Programming, 64(3):263, February 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bergstra:2007:ATS

[963] Jan A. Bergstra and Paul Klint. About “trivial” software patents: The IsNot case. Science of Computer Programming, 64(3):264–285, February 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hayes:2007:PPR

[964] Ian J. Hayes. Procedures and parameters in the real-time program refinement calculus. Science of Computer Programming, 64(3):286–311, February 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Burgess:2007:RAD

[965] Mark Burgess and Trond Reitan. A risk analysis of disk backup or repository maintenance. Science of Computer Programming, 64(3):312–331, February 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Filliatre:2007:FPP

[966] Jean-Christophe Filliâtre. Formal proof of a program: Find. Science of Computer Programming, 64(3): 332–340, February 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gao:2007:LFP

[967] H. Gao, J. F. Groote, and W. H. Hesselink. Lock-free parallel and concurrent garbage collection by mark&sweep. Science of Computer Programming, 64(3):341–374, February 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBc

[968] Anonymous. Editorial Board. Science of Computer Programming, 64 (3):??, February 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Barjis:2007:TMA

[969] Joseph Barjis, Juan Carlos Augusto, and Ulrich Ultes-Nitsche. Towards more adequate EIS. Science of Computer Programming, 65(1):1–3, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gruhn:2007:WBP

[970] Volker Gruhn and Ralf Laue. What business process modelers can learn from programmers. Science of Computer Programming, 65(1):4–13, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Yeung:2007:MMC

[971] W. L. Yeung, K. R. P. H. Leung, Ji Wang, and Wei Dong. Modelling and model checking suspendible business processes via statechart diagrams and CSP. Science of Computer Programming, 65(1):14–29, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanHee:2007:LLH

[972] Kees van Hee, Olivia Oanea, Alexander Serebrenik, Natalia Sidorova, and Marc Voorhoeve. LogLogics: a logic for history-dependent business processes. Science of Computer Programming, 65 (1):30–40, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Akhlaki:2007:MAF

[973] K. Benghazi Akhlaki, M. I. Capel Tuñón, J. A. Holgado Terriza, and L. E. Mendoza Morales. A methodological approach to the formal specification of real-time systems by transformation of UML-RT design models. Science of Computer Programming, 65 (1):41–56, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gallardo:2007:PAX

[974] María del Mar Gallardo, Jesús Martínez, Pedro Merino, Pablo Nuñez, and Ernesto Pimentel. PiXL: Applying XML standards to support the integration of analysis tools for protocols. Science of Computer Programming, 65 (1):57–69, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBd

[975] Anonymous. Editorial Board. Science of Computer Programming, 65 (1):??, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bergstra:2007:LE

[976] Jan Bergstra. Letter from the editor. Science of Computer Programming, 65(1):v, March 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kozen:2007:P

[977] Dexter Kozen. Preface. Science of Computer Programming, 65(2):71, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Atanassow:2007:CXH

[978] Frank Atanassow and Johan Jeuring. Customizing an XML–Haskell data binding with type isomorphism inference in Generic Haskell. Science of Computer Programming, 65(2): 72–107, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gibbons:2007:MSR

[979] Jeremy Gibbons. Metamorphisms: Streaming representation-changers. Science of Computer Programming, 65 (2):108–139, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Martin:2007:MAD

[980] C. E. Martin, S. A. Curtis, and I. Rewitzky. Modelling angelic and demonic nondeterminism with multirelations. Science of Computer Programming, 65(2):140–158, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morris:2007:TUD

[981] Joseph M. Morris and Malcolm Tyrrell. Terms with unbounded demonic and angelic nondeterminacy. Science of Computer Programming, 65(2):159– 172, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fernandes:2007:DDF

[982] Therrezinha Fernandes and Jules Desharnais. Describing data flow analysis techniques with Kleene algebra. Science of Computer Programming, 65(2): 173–194, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moller:2007:KGL

[983] Bernhard Möller. Kleene getting lazy. Science of Computer Programming, 65 (2):195–214, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBe

[984] Anonymous. Editorial Board. Science of Computer Programming, 65 (2):??, March 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Krikhaar:2007:SCM

[985] René Krikhaar and Ivica Crnkovic. Software Configuration Management. Science of Computer Programming, 65 (3):215–221, April 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Voinea:2007:VAS

[986] Lucian Voinea, Johan Lukkien, and Alexandru Telea. Visual assessment of software evolution. Science of Computer Programming, 65 (3):222–248, April 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Murta:2007:OSI

[987] Leonardo Murta, Hamilton Oliveira, Cristine Dantas, Luiz Gustavo Lopes, and Cláudia Werner. Odyssey-SCM: an integrated software configuration management infrastructure for UML models. Science of Computer Programming, 65(3):249–274, April 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBf

[988] Anonymous. Editorial Board. Science of Computer Programming, 65(3):??, April 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chung:2007:BGB

[989] Lawrence Chung and Nary Subramanian. Bridging the gap between enterprise architectures and software architectures. Science of Computer Programming, 66(1):1–3, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Verstichel:2007:ODM

[990] Stijn Verstichel, Sofie Van Hoecke, Matthias Strobbe, Steven Van den Berghe, Filip De Turck, Bart Dhoedt, Piet Demeester, and Frederik Vermeulen. Ontology-driven middleware for next-generation train backbones. Science of Computer Programming, 66 (1):4–24, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Garrido:2007:DUC

[991] José Luis Garrido, Manuel Noguera, Miguel González, María V. Hurtado, and María L. Rodríguez. Definition and use of Computation Independent Models in an MDA-based groupware development process. Science of Computer Programming, 66(1):25–43, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vela:2007:AAC

[992] F. L. Gutiérrez Vela, J. L. Isla Montes, P. Paderewski Rodríguez, M. Sánchez Román, and B. Jiménez Valverde. An architecture for access control management in collaborative enterprise systems based on organization models. Science of Computer Programming, 66 (1):44–59, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shin:2007:SRA

[993] Michael E. Shin and Hassan Gomaa. Software requirements and architecture modeling for evolving non-secure applications into secure applications. Science of Computer Programming, 66 (1):60–70, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pena:2007:DME

[994] Joaquin Peña, Michael G. Hinchey, Manuel Resinas, Roy Sterritt, and James L. Rash. Designing and managing evolving systems using a MAS product line approach. Science of Computer Programming, 66(1):71–86, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dai:2007:UFB

[995] Lirong Dai and Kendra Cooper. Using FDAF to bridge the gap between enterprise and software architectures for security. Science of Computer Programming, 66(1):87–102, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBg

[996] Anonymous. Editorial Board. Science of Computer Programming, 66 (1):??, April 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Broy:2007:ESC

[997] Manfred Broy. Editorial — Science of Computer Programming — 25 years. Science of Computer Programming, 66 (2):103–104, April 30, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gitzel:2007:UEW

[998] Ralf Gitzel, Axel Korthaus, and Martin Schader. Using established Web Engineering knowledge in model-driven approaches. Science of Computer Programming, 66(2):105–124, April 30, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Duran:2007:MMA

[999] Francisco Durán and José Meseguer. Maude’s module algebra. Science of Computer Programming, 66(2):125– 153, April 30, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Csenki:2007:RPP

[1000] Attila Csenki. Rotations in the plane and Prolog. Science of Computer Programming, 66(2):154–161, April 30, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ogata:2007:MVR

[1001] Kazuhiro Ogata and Kokichi Futatsugi. Modeling and verification of realtime systems based on equations. Science of Computer Programming, 66(2): 162–180, April 30, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBh

[1002] Anonymous. Editorial Board. Science of Computer Programming, 66 (2):??, April 30, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Canal:2007:P

[1003] Carlos Canal and Mirko Viroli. Preface. Science of Computer Programming, 66(3):181–183, May 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chothia:2007:CPP

[1004] Tom Chothia and Dominic Duggan. Capability passing processes. Science of Computer Programming, 66 (3):184–204, May 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Clarke:2007:CCS

[1005] Dave Clarke, David Costa, and Farhad Arbab. Connector colouring I: Synchronisation and context dependency. Science of Computer Programming, 66 (3):205–225, May 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Viroli:2007:EBO

[1006] Mirko Viroli, Enrico Denti, and Alessandro Ricci. Engineering a BPEL orchestration engine as a multi-agent system. Science of Computer Programming, 66(3):226–245, May 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Peschanski:2007:CMA

[1007] Frédéric Peschanski, Alexis Darrasse, Nataliya Guts, and Jérémy Bobbio. Coordinating mobile agents in interaction spaces. Science of Computer Programming, 66(3):246–265, May 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBi

[1008] Anonymous. Editorial Board. Science of Computer Programming, 66(3):??, May 15, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

deLeeuw:2007:P

[1009] Karl de Leeuw. Preface. Science of Computer Programming, 67(1):1–2, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moreno-Ger:2007:DAA

[1010] Pablo Moreno-Ger, José Luis Sierra, Iván Martínez-Ortiz, and Baltasar Fernández-Manjón. A documental approach to adventure game development. Science of Computer Programming, 67(1):3–31, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cutumisu:2007:SGA

[1011] Maria Cutumisu, Curtis Onuczko, Matthew McNaughton, Thomas Roy, Jonathan Schaeffer, Allan Schumacher, Jeff Siegel, Duane Szafron, Kevin Waugh, Mike Carbonaro, Harvey Duff, and Stephanie Gillis. ScriptEase: a generative/adaptive programming paradigm for game scripting. Science of Computer Programming, 67 (1):32–58, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ponsen:2007:KAA

[1012] Marc Ponsen, Pieter Spronck, Héctor Muñoz-Avila, and David W. Aha. Knowledge acquisition for adaptive game AI. Science of Computer Programming, 67(1):59–75, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Roden:2007:TME

[1013] Timothy E. Roden, Ian Parberry, and David Ducrest. Toward mobile entertainment: a paradigm for narrativebased audio only games. Science of Computer Programming, 67(1):76–90, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Nieuwenhuisen:2007:HQN

[1014] D. Nieuwenhuisen, A. Kamphuis, and M. H. Overmars. High quality navigation in computer games. Science of Computer Programming, 67 (1):91–104, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Heule:2007:SGD

[1015] M. J. H. Heule and L. J. M. Rothkrantz. Solving games: Dependence of applicable solving procedures. Science of Computer Programming, 67 (1):105–124, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBj

[1016] Anonymous. Editorial Board. Science of Computer Programming, 67(1):??, June 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shehata:2007:USF

[1017] Mohamed Shehata, Armin Eberlein, and Abraham Fapojuwo. Using semiformal methods for detecting interactions among smart homes policies. Science of Computer Programming, 67(2– 3):125–161, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ouyang:2007:FSA

[1018] Chun Ouyang, Eric Verbeek, Wil M. P.

van der Aalst, Stephan Breutel, Marlon Dumas, and Arthur H. M. ter Hofstede. Formal semantics and analysis of control flow in WS-BPEL. Science of Computer Programming, 67(2– 3):162–198, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Erwig:2007:UCE

[1019] Martin Erwig and Deling Ren. An update calculus for expressing type-safe program updates. Science of Computer Programming, 67(2–3):199–222, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Maurer:2007:GSP

[1020] Ward Douglas Maurer. Generalized structured programs and loop trees. Science of Computer Programming, 67 (2–3):223–246, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Verhoef:2007:QEI

[1021] C. Verhoef. Quantifying the effects of IT-governance rules. Science of Computer Programming, 67(2–3):247–277, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fong:2007:RAS

[1022] Philip W. L. Fong. Reasoning about safety properties in a JVM-like environment. Science of Computer Programming, 67(2–3):278–300, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Banach:2007:ETU

[1023] R. Banach, M. Poppleton, C. Jeske, and S. Stepney. Engineering and theoretical underpinnings of retrenchment. Science of Computer Programming, 67 (2–3):301–329, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBk

[1024] Anonymous. Editorial Board. Science of Computer Programming, 67 (2–3):??, July 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bryant:2007:RSP

[1025] Barrett R. Bryant and João Saraiva. Report on selected papers from LDTA 2003. Science of Computer Programming, 68(1):1, August 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brabrand:2007:MSS

[1026] Claus Brabrand and Michael I. Schwartzbach. The metafront system: Safe and extensible parsing and transformation. Science of Computer Programming, 68(1):2–20, August 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Magnusson:2007:CRA

[1027] Eva Magnusson and Görel Hedin. Circular reference attributed grammars — their evaluation and applications. Science of Computer Programming, 68 (1):21–37, August 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

VanWyk:2007:IAO

[1028] Eric Van Wyk. Implementing aspectoriented programming constructs as modular language extensions. Science of Computer Programming, 68 (1):38–61, August 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBl

[1029] Anonymous. Editorial Board. Science of Computer Programming, 68 (1):??, August 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hedin:2007:P

[1030] Görel Hedin and John Tang Boyland.

Preface. Science of Computer Programming, 68(2):63, September 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Waddington:2007:HFC

[1031] Daniel Waddington and Bin Yao. Highfidelity C/C++ code transformation. Science of Computer Programming, 68 (2):64–78, September 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Iversen:2007:ACT

[1032] Jørgen Bøndergaard Iversen. An action compiler targeting Standard ML. Science of Computer Programming, 68 (2):79–94, September 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Johnstone:2007:ARE

[1033] Adrian Johnstone and Elizabeth Scott. Automatic recursion engineering of reduction incorporated parsers. Science of Computer Programming, 68(2):95– 110, September 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBm

[1034] Anonymous. Editorial Board. Science of Computer Programming, 68 (2):??, September 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pierantonio:2007:SIM

[1035] Alfonso Pierantonio, Antonio Vallecillo, Bran Selic, and Jeff Gray. Special issue on model transformation. Science of Computer Programming, 68(3): 111–113, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jouault:2007:IMM

[1036] Frédéric Jouault and Ivan Kurtev. On the interoperability of model-to-model transformation languages. Science of Computer Programming, 68(3): 114–137, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kurtev:2007:RBM

[1037] Ivan Kurtev, Klaas van den Berg, and Frédéric Jouault. Rule-based modularization in model transformation languages illustrated with ATL. Science of Computer Programming, 68(3): 138–154, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Alanen:2007:CRD

[1038] Marcus Alanen, Torbjörn Lundkvist, and Ivan Porres. Creating and reconciling diagrams after executing model transformations. Science of Computer Programming, 68(3):155–178, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cabot:2007:TTO

[1039] J. Cabot and E. Teniente. Transformation techniques for OCL constraints. Science of Computer Programming, 68 (3):179–195, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tratt:2007:MTM

[1040] Laurence Tratt. Model transformations in MT. Science of Computer Programming, 68(3):196–213, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Varro:2007:MTL

[1041] Dániel Varró and András Balogh.

The model transformation language of the VIATRA2 framework. Science of Computer Programming, 68(3): 214–234, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBn

[1042] Anonymous. Editorial Board. Science of Computer Programming, 68 (3):??, October 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vandenBrand:2007:GEI

[1043] M. G. J. van den Brand. Guest Editor’s introduction: Experimental Software and Toolkits (EST). Science of Computer Programming, 69(1– 3):1–2, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kraft:2007:TCR

[1044] Nicholas A. Kraft, Brian A. Malloy, and James F. Power. A tool chain for reverse engineering C++ applications. Science of Computer Programming, 69(1–3):3–13, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ekman:2007:JSM

[1045] Torbjörn Ekman and Görel Hedin. The JastAdd system — modular extensible compiler construction. Science of Computer Programming, 69(1–3): 14–26, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gogolla:2007:UUB

[1046] Martin Gogolla, Fabian Büttner, and Mark Richters. USE: a UML-based specification environment for validating UML and OCL. Science of Computer Programming, 69(1–3):27– 34, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ernst:2007:DSD

[1047] Michael D. Ernst, Jeff H. Perkins, Philip J. Guo, Stephen McCamant, Carlos Pacheco, Matthew S. Tschantz, and Chen Xiao. The Daikon system for dynamic detection of likely invariants. Science of Computer Programming, 69(1–3):35–45, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Oliveira:2007:CW

[1048] Cristóvão Oliveira and Michel Wermelinger. The CommUnity Workbench. Science of Computer Programming, 69(1–3):46–55, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Li:2007:CTA

[1049] Xianfeng Li, Yun Liang, Tulika Mitra, and Abhik Roychoudhury. Chronos: a timing analyzer for embedded software. Science of Computer Programming, 69(1–3):56–67, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Diaz-Agudo:2007:BCS

[1050] Belén Díaz-Agudo, Pedro A. González-Calero, Juan A. Recio-García, and Antonio A. Sánchez-Ruiz-Granados. Building CBR systems with j colibri. Science of Computer Programming, 69 (1–3):68–75, December 1, 2007. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Johnstone:2007:PPS

[1051] Adrian Johnstone and Elizabeth Scott.

Proofs and pedagogy; science and systems: The grammar tool box. Science of Computer Programming, 69(1– 3):76–85, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBo

[1052] Anonymous. Editorial Board. Science of Computer Programming, 69(1– 3):??, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2007:EBp

[1053] Anonymous. Editorial Board. Science of Computer Programming, 69(1– 3):??, December 1, 2007. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lammel:2008:GMP

[1054] Ralf Lämmel. Google’s MapReduce programming model — Revisited. Science of Computer Programming, 70 (1):1–30, January 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Duan:2008:FTL

[1055] Zhenhua Duan, Xiaoxiao Yang, and Maciej Koutny. Framed temporal logic programming. Science of Computer Programming, 70(1):31–61, January 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cicchetti:2008:DWA

[1056] Antonio Cicchetti and Davide Di Ruscio. Decoupling Web application concerns through weaving operations. Science of Computer Programming, 70 (1):62–86, January 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBa

[1057] Anonymous. Editorial Board. Science of Computer Programming, 70 (1):??, January 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aleksy:2008:PSI

[1058] Markus Aleksy, Ralf Gitzel, John Waldron, and James F. Power. PPPJ 2006 special issue — foreword. Science of Computer Programming, 70(2– 3):87–88, February 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Xian:2008:GCJ

[1059] Feng Xian, Witawas Srisa-an, and Hong Jiang. Garbage collection: Java application servers’ Achilles heel. Science of Computer Programming, 70(2– 3):89–110, February 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Singer:2008:DAJ

[1060] Jeremy Singer and Chris Kirkham.

Dynamic analysis of Java program concepts for visualization and profiling. Science of Computer Programming, 70(2–3):111–126, February 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mathiske:2008:ADF

[1061] Bernd Mathiske, Doug Simon, and Dave Ungar. An assembler and disassembler framework for Java^{T M} programmers. Science of Computer Programming, 70(2–3):127–148, February 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Navarro:2008:CNM

[1062] A. Navarro, A. Fernández-Valmayor, B. Fernández-Manjón, and J. L. Sierra. Characterizing navigation maps for Web applications with the NMM approach. Science of Computer Programming, 71(1):1–16, March 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Peters:2008:QYR

[1063] R. J. Peters and C. Verhoef. Quantifying the yield of risk-bearing ITportfolios. Science of Computer Programming, 71(1):17–56, March 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hesselink:2008:CAV

[1064] Wim H. Hesselink. A challenge for atomicity verification. Science of Computer Programming, 71(1):57–72, March 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Barjis:2008:IBP

[1065] Joseph Barjis. The importance of business process modeling in software systems design. Science of Computer Programming, 71(1):73–87, March 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBb

[1066] Anonymous. Editorial Board. Science of Computer Programming, 71 (1):??, March 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Flanagan:2008:ADA

[1067] Cormac Flanagan and Stephen N. Freund. Atomizer: a dynamic atomicity checker for multithreaded programs. Science of Computer Programming, 71 (2):89–109, April 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Acciai:2008:XTP

[1068] Lucia Acciai and Michele Boreale. XPi:

a typed process calculus for XML messaging. Science of Computer Programming, 71(2):110–143, April 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Crasso:2008:EWS

[1069] Marco Crasso, Alejandro Zunino, and Marcelo Campo. Easy Web service discovery: a query-by-example approach. Science of Computer Programming, 71 (2):144–164, April 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBc

[1070] Anonymous. Editorial Board. Science of Computer Programming, 71(2):??, April 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fabry:2008:KKA

[1071] Johan Fabry, Éric Tanter, and Theo D’Hondt. KALA: Kernel aspect language for advanced transactions. Science of Computer Programming, 71 (3):165–180, May 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tivoli:2008:FFC

[1072] Massimo Tivoli and Paola Inverardi. Failure-free coordinators synthesis for component-based architectures. Science of Computer Programming, 71 (3):181–212, May 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zhao:2008:IOT

[1073] Tian Zhao, Jason Baker, James Hunt, James Noble, and Jan Vitek. Implicit ownership types for memory management. Science of Computer Programming, 71(3):213–241, May 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chung:2008:SCN

[1074] Myung-Hoon Chung. Science Code .Net: Object-oriented programming for science. Science of Computer Programming, 71(3):242–247, May 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBd

[1075] Anonymous. Editorial Board. Science of Computer Programming, 71(3):??, May 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vandenBrand:2008:GEI

[1076] M. G. J. van den Brand. Guest Editor’s introduction: Second issue of Experimental Software and Toolkits (EST). Science of Computer Programming, 72 (1–2):1–2, June 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bagnara:2008:PPL

[1077] Roberto Bagnara, Patricia M. Hill, and Enea Zaffanella. The Parma Polyhedra Library: Toward a complete set of numerical abstractions for the analysis and verification of hardware and software systems. Science of Computer Programming, 72(1–2):3–21, June 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tanter:2008:FMA

[1078] Éric Tanter, Rodolfo Toledo, Guillaume Pothier, and Jacques Noyé. Flexible metaprogramming and AOP in Java. Science of Computer Programming, 72(1–2):22–30, June 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jouault:2008:AMT

[1079] Frédéric Jouault, Freddy Allilaire, Jean Bézivin, and Ivan Kurtev. ATL: a model transformation tool. Science of Computer Programming, 72 (1–2):31–39, June 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jackson:2008:SFP

[1080] Stoney Jackson, Premkumar Devanbu, and Kwan-Liu Ma. Stable, flexible, peephole pretty-printing. Science of Computer Programming, 72 (1–2):40–51, June 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bravenboer:2008:SXL

[1081] Martin Bravenboer, Karl Trygve Kalleberg, Rob Vermaas, and Eelco Visser. Stratego/XT 0.17. A language and toolset for program transformation. Science of Computer Programming, 72 (1–2):52–70, June 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBe

[1082] Anonymous. Editorial Board. Science of Computer Programming, 72 (1–2):??, June 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Daylight:2008:FUE

[1083] Edgar G. Daylight, Arnout Vandecappelle, and Francky Catthoor. The formalism underlying EASYMAP: a precompiler for refinement-based exploration of hierarchical data organizations. Science of Computer Programming, 72(3):71–135, August 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kulk:2008:QRV

[1084] G. P. Kulk and C. Verhoef. Quantifying requirements volatility effects. Science of Computer Programming, 72 (3):136–175, August 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fruja:2008:TPT

[1085] Nicu G. Fruja. Towards proving type safety of .NET CIL. Science of Computer Programming, 72(3): 176–219, August 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Maurer:2008:PDC

[1086] Ward Douglas Maurer. Partially defined computer instructions and guards. Science of Computer Programming, 72(3):220–239, August 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBf

[1087] Anonymous. Editorial Board. Science of Computer Programming, 72 (3):??, August 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aksit:2008:ESI

[1088] Mehmet Aksit and Ivan Kurtev. Elsevier special issue on foundations and applications of model driven architecture. Science of Computer Programming, 73(1):1–2, September 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Christoph:2008:GUT

[1089] Alexander Christoph and Matthias M.

Müller. GREAT: UML transformation tool for porting middleware applications. Science of Computer Programming, 73(1):3–12, September 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cepa:2008:LSM

[1090] Vasian Cepa and Mira Mezini. Language support for model-driven software development. Science of Computer Programming, 73(1):13–25, September 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Karsai:2008:MDA

[1091] Gabor Karsai, Sandeep Neema, and David Sharp. Model-driven architecture for embedded software: a synopsis and an example. Science of Computer Programming, 73(1):26–38, September 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gokhale:2008:MDM

[1092] Aniruddha Gokhale, Krishnakumar Balasubramanian, Arvind S. Krishna, Jaiganesh Balasubramanian, George Edwards, Gan Deng, Emre Turkay, Jeffrey Parsons, and Douglas C. Schmidt. Model driven middleware: a new paradigm for developing distributed real-time and embedded systems. Science of Computer Programming, 73(1): 39–58, September 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBg

[1093] Anonymous. Editorial Board. Science of Computer Programming, 73 (1):??, September 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cimadamore:2008:RJW

[1094] Maurizio Cimadamore and Mirko Viroli. On the reification of Java wildcards. Science of Computer Programming, 73(2–3):59–75, October 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lanotte:2008:DVL

[1095] Ruggero Lanotte, Andrea Maggiolo-Schettini, Paolo Milazzo, and Angelo Troina. Design and verification of longrunning transactions in a timed framework. Science of Computer Programming, 73(2–3):76–94, October 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vu:2008:GEP

[1096] Thuy Duong Vu. Goto elimination in program algebra. Science of Computer Programming, 73(2–3):95–128, October 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

deFreitas:2008:NPG

[1097] Renata de Freitas and Petrucio Viana.

A note on proofs with graphs. Science of Computer Programming, 73(2– 3):129–135, October 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBh

[1098] Anonymous. Editorial Board. Science of Computer Programming, 73 (2–3):??, October 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Massacci:2008:P

[1099] Fabio Massacci, Frank Piessens, and Sjouke Mauw. Preface. Science of Computer Programming, 74(1–2):1, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aktug:2008:CFL

[1100] Irem Aktug and Katsiaryna Naliuka.

ConSpec — a formal language for policy specification. Science of Computer Programming, 74(1–2):2–12, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schellekens:2008:RAL

[1101] Dries Schellekens, Brecht Wyseur, and Bart Preneel. Remote attestation on legacy operating systems with trusted platform modules. Science of Computer Programming, 74(1–2): 13–22, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bianchi:2008:SPA

[1102] G. Bianchi, M. Bonola, V. Falletta, F. S. Proto, and S. Teofili. The SPARTA pseudonym and authorization system. Science of Computer Programming, 74(1–2):23–33, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Refsdal:2008:EUS

[1103] Atle Refsdal and Ketil Stølen. Extending UML sequence diagrams to model trust-dependent behavior with the aim to support risk analysis. Science of Computer Programming, 74(1– 2):34–42, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bringer:2008:BBW

[1104] Julien Bringer, Hervé Chabanne, and Bruno Kindarji. The best of both worlds: Applying secure sketches to cancelable biometrics. Science of Computer Programming, 74(1–2):43– 51, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

denHartog:2008:TMC

[1105] Jerry den Hartog. Towards mechanized correctness proofs for cryptographic algorithms: Axiomatization of a probabilistic Hoare style logic. Science of Computer Programming, 74(1– 2):52–63, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2008:EBi

[1106] Anonymous. Editorial Board. Science of Computer Programming, 74(1– 2):??, December 1, 2008. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Eshuis:2009:RSS

[1107] Rik Eshuis. Reconciling statechart semantics. Science of Computer Programming, 74(3):65–99, January 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Garavel:2009:SCH

[1108] Hubert Garavel, Gwen Salaün, and Wendelin Serwe. On the semantics of communicating hardware processes and their translation into LOTOS for the verification of asynchronous circuits with CADP. Science of Computer Programming, 74(3):100–127, January 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kourie:2009:IAC

[1109] Derrick G. Kourie, Sergei Obiedkov, Bruce W. Watson, and Dean van der Merwe. An incremental algorithm to construct a lattice of set intersections. Science of Computer Programming, 74 (3):128–142, January 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Colvin:2009:GTP

[1110] Robert Colvin and Brijesh Dongol.

A general technique for proving lockfreedom. Science of Computer Programming, 74(3):143–165, January 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBa

[1111] Anonymous. Editorial Board. Science of Computer Programming, 74 (3):??, January 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dong:2009:SSI

[1112] Jin Song Dong and Jing Sun. SCP special issue on the grand challenge — Preface. Science of Computer Programming, 74(4):167, February 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chen:2009:RVC

[1113] Zhenbang Chen, Zhiming Liu, Anders P. Ravn, Volker Stolz, and Naijun Zhan. Refinement and verification in component-based model-driven design. Science of Computer Programming, 74(4):168–196, February 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Freitas:2009:VCF

[1114] Leo Freitas, Jim Woodcock, and Yichi Zhang. Verifying the CICS File Control API with Z/Eves: An experiment in the verified software repository. Science of Computer Programming, 74(4): 197–218, February 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Butterfield:2009:MFM

[1115] Andrew Butterfield, Leo Freitas, and Jim Woodcock. Mechanising a formal model of flash memory. Science of Computer Programming, 74(4):

219–237, February 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Freitas:2009:PFS

[1116] Leo Freitas, Jim Woodcock, and Zheng Fu. POSIX file store in Z/Eves: an experiment in the verified software repository. Science of Computer Programming, 74(4):238–257, February 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBb

[1117] Anonymous. Editorial Board. Science of Computer Programming, 74 (4):??, February 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aleksy:2009:PSI

[1118] Markus Aleksy, Vasco Amaral, Ralf Gitzel, James Power, and John Waldron. PPPJ special issue–Foreword. Science of Computer Programming, 74 (5–6):259–260, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bettini:2009:FJD

[1119] Lorenzo Bettini, Sara Capecchi, and Betti Venneri. Featherweight Java with dynamic and static overloading. Science of Computer Programming, 74(5– 6):261–278, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wurthinger:2009:ABC

[1120] Thomas Würthinger, Christian Wimmer, and Hanspeter Mössenböck. Array bounds check elimination in the context of deoptimization. Science of Computer Programming, 74(5–6): 279–295, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bradel:2009:SPP

[1121] Borys J. Bradel and Tarek S. Abdelrahman. A study of potential parallelism among traces in Java programs. Science of Computer Programming, 74(5– 6):296–313, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zhang:2009:ISE

[1122] Lingli Zhang and Chandra Krintz. Asif-serial exception handling semantics for Java futures. Science of Computer Programming, 74(5–6):314–332, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hoffman:2009:CAO

[1123] Kevin Hoffman and Patrick Eugster. Cooperative aspect-oriented programming. Science of Computer Programming, 74(5–6):333–354, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Makela:2009:CBC

[1124] Sami Mäkelä and Ville Leppänen. Client-based cohesion metrics for Java programs. Science of Computer Programming, 74(5–6):355–378, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Puder:2009:BCL

[1125] Arno Puder and Sascha Häberling. Byte code level cross-compilation for developing Web applications. Science of Computer Programming, 74(5–6):

379–396, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBc

[1126] Anonymous. Editorial Board. Science of Computer Programming, 74 (5–6):??, March 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Krikhaar:2009:GEI

[1127] René Krikhaar and Ralf Lämmel.

Guest Editors’ introduction. Science of Computer Programming, 74 (7):397–398, May 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Vidacs:2009:CPS

[1128] László Vidács, Árpád Beszédes, and Tibor Gyimóthy. Combining preprocessor slicing with C/C++ language slicing. Science of Computer Programming, 74 (7):399–413, May 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hindle:2009:RBL

[1129] Abram Hindle, Michael W. Godfrey, and Richard C. Holt. Reading beside the lines: Using indentation to rank revisions by complexity. Science of Computer Programming, 74 (7):414–429, May 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Binkley:2009:ILL

[1130] Dave Binkley, Dawn Lawrie, Steve Maex, and Christopher Morrell. Identifier length and limited programmer memory. Science of Computer Programming, 74(7):430–445, May 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dalton:2009:VRB

[1131] Andrew R. Dalton, Sally K. Wahba, Sravanthi Dandamudi, and Jason O. Hallstrom. Visualizing the runtime behavior of embedded network systems: a toolkit for TinyOS. Science of Computer Programming, 74 (7):446–469, May 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Roy:2009:CEC

[1132] Chanchal K. Roy, James R. Cordy, and Rainer Koschke. Comparison and evaluation of code clone detection techniques and tools: a qualitative approach. Science of Computer Programming, 74(7):470–495, May 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBd

[1133] Anonymous. Editorial Board. Science of Computer Programming, 74(7):??, May 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Uustalu:2009:P

[1134] Tarmo Uustalu. Preface. Science of Computer Programming, 74(8):497, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cockett:2009:LMP

[1135] J. R. B. Cockett and Craig Pastro. The logic of message-passing. Science of Computer Programming, 74 (8):498–533, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Danvy:2009:RW

[1136] Olivier Danvy and Kevin Millikin.

Refunctionalization at work. Science of Computer Programming, 74 (8):534–549, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Abel:2009:TBT

[1137] Andreas Abel. Type-based termination of generic programs. Science of Computer Programming, 74 (8):550–567, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Audebaud:2009:PRA

[1138] Philippe Audebaud and Christine Paulin-Mohring. Proofs of randomized algorithms in Coq. Science of Computer Programming, 74(8):568– 589, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hinze:2009:GP

[1139] Ralf Hinze and Andres Löh. Generic programming in 3D. Science of Computer Programming, 74(8):590– 628, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Morgan:2009:SKR

[1140] Carroll Morgan. The Shadow knows:

Refinement and security in sequential programs. Science of Computer Programming, 74(8):629–653, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Solin:2009:ETR

[1141] Kim Solin and Joakim von Wright. Enabledness and termination in refinement algebra. Science of Computer Programming, 74(8):654–668, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBe

[1142] Anonymous. Editorial Board. Science of Computer Programming, 74(8):??, June 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pimentel:2009:SII

[1143] Ernesto Pimentel and Jean-Marie Jacquet. Special issue on the 5th International Workshop on Foundations of Coordination Languages and Software Architectures (FOCLASA’06). Science of Computer Programming, 74 (9):669–670, July 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Barbosa:2009:PSO

[1144] Marco A. Barbosa and Luis S. Barbosa. A perspective on service orchestration. Science of Computer Programming, 74(9):671–687, July 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kluppelholz:2009:SMC

[1145] Sascha Klüppelholz and Christel Baier. Symbolic model checking for channelbased component connectors. Science of Computer Programming, 74 (9):688–701, July 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Casadei:2009:CSP

[1146] Matteo Casadei, Mirko Viroli, and Luca Gardelli. On the collective sort problem for distributed tuple spaces. Science of Computer Programming, 74 (9):702–722, July 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fraikin:2009:ESC

[1147] Benoît Fraikin and Marc Frappier. Efficient symbolic computation of process expressions. Science of Computer Programming, 74(9):723–753, July 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Molina:2009:MAU

[1148] Ana Isabel Molina, Miguel Ángel Redondo, and Manuel Ortega. A methodological approach for user interface development of collaborative applications: a case study. Science of Computer Programming, 74 (9):754–776, July 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBf

[1149] Anonymous. Editorial Board. Science of Computer Programming, 74(9):??, July 1, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carpineti:2009:PPE

[1150] Samuele Carpineti, Cosimo Laneve, and Luca Padovani. PiDuce — a project for experimenting Web services technologies. Science of Computer Programming, 74(10):777–811, August 15, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Roldan:2009:IDS

[1151] Manuel Roldán, Francisco Durán, and Antonio Vallecillo. Invariant-driven specifications in Maude. Science of Computer Programming, 74(10):812– 835, August 15, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ortin:2009:EVM

[1152] Francisco Ortin, Jose Manuel Redondo, and J. Baltasar García Perez-Schofield. Efficient virtual machine support of runtime structural reflection. Science of Computer Programming, 74(10): 836–860, August 15, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lemos:2009:ITO

[1153] Otávio Augusto Lazzarini Lemos, Ivan Gustavo Franchin, and Paulo Cesar Masiero. Integration testing of Object-Oriented and Aspect-Oriented programs: a structural pairwise approach for Java. Science of Computer Programming, 74(10):861–878, August 15, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBg

[1154] Anonymous. Editorial Board. Science of Computer Programming, 74 (10):??, August 15, 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hoang:2009:DTD

[1155] Thai Son Hoang, Hironobu Kuruma, David Basin, and Jean-Raymond Abrial. Developing topology discovery in Event-B. Science of Computer Programming, 74(11–12):879–899, November 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kulk:2009:QIE

[1156] G. P. Kulk, R. J. Peters, and C. Verhoef. Quantifying IT estimation risks. Science of Computer Programming, 74 (11–12):900–933, November 2009. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Eveleens:2009:QIF

[1157] J. L. Eveleens and C. Verhoef. Quantifying IT forecast quality. Science of Computer Programming, 74(11–12): 934–988, November 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mili:2009:MRA

[1158] Ali Mili, Shir Aharon, and Chaitanya Nadkarni. Mathematics for reasoning about loop functions. Science of Computer Programming, 74(11–12): 989–1020, November 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2009:EBh

[1159] Anonymous. Editorial Board. Science of Computer Programming, 74 (11–12):??, November 2009. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Johnstone:2010:P

[1160] Adrian Johnstone, Anthony Sloane, and John Tang Boyland. Preface. Science of Computer Programming, 75(1– 2):1–2, January 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Camacho:2010:AGP

[1161] Diego Ordóñez Camacho, Kim Mens, Mark van den Brand, and Jurgen Vinju. Automated generation of program translation and verification tools using annotated grammars. Science of Computer Programming, 75(1–2): 3–20, January 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Aakesson:2010:IMC

[1162] Johan Åkesson, Torbjörn Ekman, and Görel Hedin. Implementation of a Modelica compiler using JastAdd attribute grammars. Science of Computer Programming, 75(1–2):21–38, January 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

VanWyk:2010:SEA

[1163] Eric Van Wyk, Derek Bodin, Jimin Gao, and Lijesh Krishnan. Silver: an extensible attribute grammar system. Science of Computer Programming, 75 (1–2):39–54, January 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Scott:2010:RPS

[1164] Elizabeth Scott and Adrian Johnstone. Recognition is not parsing — SPPFstyle parsing from cubic recognisers. Science of Computer Programming, 75 (1–2):55–70, January 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schmitz:2010:EAD

[1165] Sylvain Schmitz. An experimental ambiguity detection tool. Science of Computer Programming, 75(1–2): 71–84, January 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rinderknecht:2010:TPU

[1166] Christian Rinderknecht and Nic Volanschi. Theory and practice of unparsed patterns for metacompilation. Science of Computer Programming, 75 (3):85–105, March 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cornelio:2010:SR

[1167] Márcio Cornélio, Ana Cavalcanti, and Augusto Sampaio. Sound refactorings. Science of Computer Programming, 75 (3):106–133, March 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ward:2010:CDS

[1168] Martin Ward and Hussein Zedan.

Combining dynamic and static slicing for analysing assembler. Science of Computer Programming, 75 (3):134–175, March 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brabrand:2010:AAC

[1169] Claus Brabrand, Robert Giegerich, and Anders Møller. Analyzing ambiguity of context-free grammars. Science of Computer Programming, 75 (3):176–191, March 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Derrick:2010:MTA

[1170] John Derrick and Heike Wehrheim. Model transformations across views. Science of Computer Programming, 75 (3):192–210, March 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2010:EBa

[1171] Anonymous. Editorial Board. Science of Computer Programming, 75(3):??, March 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

deRoever:2010:RAP

[1172] Willem Paul de Roever. Remembering Amir Pnueli, a great scientist and a great friend (April 22, 1941–November 2, 2009). Science of Computer Programming, 75(4):211–213, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vandenBrand:2010:GEI

[1173] Mark G. J. van den Brand and Kim Mens. Guest Editors’ introduction to the 3rd issue of Experimental Software and Toolkits (EST): a special issue on Academic Software Development Tools and Techniques (WAS-DeTT 2008). Science of Computer Programming, 75(4):214–215, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Spinellis:2010:CRB

[1174] Diomidis Spinellis. CScout: a refactoring browser for C. Science of Computer Programming, 75(4):216–231, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Brichau:2010:ESR

[1175] Johan Brichau, Andy Kellens, Sergio Castro, and Theo D’Hondt. Enforcing structural regularities in software using IntensiVE. Science of Computer Programming, 75(4):232–246, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kienle:2010:RES

[1176] Holger M. Kienle and Hausi A. Müller.

Rigi—an environment for software reverse engineering, exploration, visualization, and redocumentation. Science of Computer Programming, 75 (4):247–263, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lungu:2010:SPO

[1177] Mircea Lungu, Michele Lanza, Tudor Gîrba, and Romain Robbes. The Small Project Observatory: Visualizing software ecosystems. Science of Computer Programming, 75(4):264– 275, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

DAmbros:2010:DCS

[1178] Marco D’Ambros and Michele Lanza.

Distributed and Collaborative Software Evolution Analysis with Churrasco. Science of Computer Programming, 75 (4):276–287, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2010:EBb

[1179] Anonymous. Editorial Board. Science of Computer Programming, 75(4):??, April 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wood:2010:CML

[1180] Alan Wood and Michael Schumacher. Coordination models, languages and applications. Science of Computer Programming, 75(5):289–291, May 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Weber:2010:IEU

[1181] Barbara Weber, Bela Mutschler, and Manfred Reichert. Investigating the effort of using business process management technology: Results from a controlled experiment. Science of Computer Programming, 75(5):292– 310, May 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mendonca:2010:DMC

[1182] Marcilio Mendonca and Donald Cowan. Decision-making coordination and efficient reasoning techniques for featurebased configuration. Science of Computer Programming, 75(5):311–332, May 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pereira:2010:FAM

[1183] David P. Pereira and Ana C. V. de Melo. Formalization of an architectural model for exception handling coordination based on CA action concepts. Science of Computer Programming, 75(5):333–349, May 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ramos:2010:CNC

[1184] Rodrigo Ramos, Augusto Sampaio, and Alexandre Mota. Conformance notions for the coordination of interaction components. Science of Computer Programming, 75(5):350–373, May 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2010:EBc

[1185] Anonymous. Editorial Board. Science of Computer Programming, 75(5):??, May 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lea:2010:GEF

[1186] Doug Lea and Gianluigi Zavattaro.

Guest Editors’ foreword. Science of Computer Programming, 75(6):375, June 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

DeNicola:2010:FLS

[1187] Rocco De Nicola, Daniele Gorla, René Rydhof Hansen, Flemming Nielson, Hanne Riis Nielson, Christian W. Probst, and Rosario Pugliese. From Flow Logic to static type systems for coordination languages. Science of Computer Programming, 75(6):376– 397, June 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kluppelholz:2010:ATS

[1188] Sascha Klüppelholz and Christel Baier. Alternating-time stream logic for multi-agent systems. Science of Computer Programming, 75(6):398–425, June 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Russello:2010:PDC

[1189] Giovanni Russello, Changyu Dong, Naranker Dulay, Michel Chaudron, and Maarten van Steen. Providing data confidentiality against malicious hosts in Shared Data Spaces. Science of Computer Programming, 75 (6):426–439, June 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Singh:2010:PCM

[1190] Anu Singh, C. R. Ramakrishnan, and Scott A. Smolka. A process calculus for Mobile Ad Hoc Networks. Science of Computer Programming, 75 (6):440–469, June 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lawall:2010:PSI

[1191] Julia L. Lawall. Preface to special issue on Generative Programming and Component Engineering (GPCE 2007). Science of Computer Programming, 75 (7):471–472, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bravenboer:2010:PIA

[1192] Martin Bravenboer, Eelco Dolstra, and Eelco Visser. Preventing injection attacks with syntax embeddings. Science of Computer Programming, 75 (7):473–495, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Culpepper:2010:DHM

[1193] Ryan Culpepper and Matthias Felleisen.

Debugging hygienic macros. Science of Computer Programming, 75(7):496– 515, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Fradet:2010:AAE

[1194] Pascal Fradet and Stéphane Hong Tuan Ha. Aspects of availability: Enforcing timed properties to prevent denial of service. Science of Computer Programming, 75(7):516–542, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Frisby:2010:CLP

[1195] Nicolas Frisby, Garrin Kimmell, Philip Weaver, and Perry Alexander. Constructing language processors with algebra combinators. Science of Computer Programming, 75(7):543– 572, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gil:2010:SSS

[1196] Joseph (Yossi) Gil and Keren Lenz.

Simple and safe SQL queries with C++ templates. Science of Computer Programming, 75(7):573–595, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jarvi:2010:PCC

[1197] Jaakko Järvi, Mat Marcus, and Jacob N. Smith. Programming with C++ concepts. Science of Computer Programming, 75(7):596–614, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jung:2010:TCF

[1198] Georg Jung and John Hatcliff. A type-centric framework for specifying heterogeneous, large-scale, componentoriented, architectures. Science of Computer Programming, 75(7):615– 637, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pirkelbauer:2010:DEC

[1199] Peter Pirkelbauer, Yuriy Solodkyy, and Bjarne Stroustrup. Design and evaluation of C++ open multi-methods. Science of Computer Programming, 75 (7):638–667, July 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chung:2010:DHQ

[1200] Lawrence Chung, José Luis Garrido, Nary Subramanian, Manuel Noguera, and Kawtar Benghazi. Designing high quality system/software architectures. Science of Computer Programming, 75 (8):669–671, August 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lung:2010:EBA

[1201] Chung-Horng Lung, Pragash Rajeswaran, Sathyanarayanan Sivadas, and Theleepan Sivabalasingam. Experience of building an architecturebased generator using GenVoca for distributed systems. Science of Computer Programming, 75(8):672–688, August 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bennett:2010:AOM

[1202] Jeannette Bennett, Kendra Cooper, and Lirong Dai. Aspect-oriented model-driven skeleton code generation: a graph-based transformation approach. Science of Computer Programming, 75(8):689–725, August 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Noguera:2010:ODA

[1203] Manuel Noguera, María V. Hurtado, María Luisa Rodríguez, Lawrence Chung, and José Luis Garrido. Ontology-driven analysis of UMLbased collaborative processes using OWL-DL and CPN. Science of Computer Programming, 75(8):726–760, August 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ancona:2010:PSI

[1204] Davide Ancona and Alex Buckley. Preface to the special issue on Object-Oriented Programming Languages and Systems (OOPS 2008), A Special Track at the 23rd ACM Symposium on Applied Computing. Science of Computer Programming, 75(9):761, September 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jarvi:2010:CLE

[1205] Jaakko Järvi and John Freeman. C++ lambda expressions and closures. Science of Computer Programming, 75(9): 762–772, September 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pirkelbauer:2010:DAS

[1206] Peter Pirkelbauer, Sean Parent, Mat Marcus, and Bjarne Stroustrup. Dynamic algorithm selection for runtime concepts. Science of Computer Programming, 75(9):773–786, September 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hallett:2010:ICS

[1207] J. J. Hallett, Victor Luchangco, Sukyoung Ryu, and Guy L. Steele Jr. Integrating coercion with subtyping and multiple dispatch. Science of Computer Programming, 75(9):787–795, September 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Logozzo:2010:PWR

[1208] Francesco Logozzo and Manuel Fähndrich. Pentagons: a weakly relational abstract domain for the efficient validation of array accesses. Science of Computer Programming, 75(9):796– 807, September 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Canal:2010:P

[1209] Carlos Canal and Corina S. Pasareanu. Preface. Science of Computer Programming, 75(10):809–810, October 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Sabouri:2010:ABS

[1210] Hamideh Sabouri and Marjan Sirjani. Actor-based slicing techniques for efficient reduction of Rebeca models. Science of Computer Programming, 75(10):811–827, October 1, 2010.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rodrigues:2010:SAA

[1211] Nuno F. Rodrigues and Luis S. Barbosa. Slicing for architectural analysis. Science of Computer Programming, 75(10):828–847, October 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ciancia:2010:EBC

[1212] Vincenzo Ciancia, Gianluigi Ferrari, Roberto Guanciale, and Daniele Strollo. Event based choreography. Science of Computer Programming, 75 (10):848–878, October 1, 2010. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lei:2010:RTS

[1213] Bin Lei, Xuandong Li, Zhiming Liu, Charles Morisset, and Volker Stolz. Robustness testing for software components. Science of Computer Programming, 75(10):879–897, October 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Heam:2010:CSB

[1214] Pierre-Cyrille Héam, Olga Kouchnarenko, and Jérôme Voinot. Component simulation-based substitutivity managing QoS and composition issues. Science of Computer Programming, 75 (10):898–917, October 1, 2010. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mernik:2010:SIP

[1215] Marjan Mernik, Barrett R. Bryant, and Chang-Hyun Jo. Special issue on the Programming Languages Track at the 23rd ACM Symposium on Applied Computing. Science of Computer Programming, 75(11):919–920, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Buss:2010:NAS

[1216] Marcio Buss, Daniel Brand, Vugranam Sreedhar, and Stephen A. Edwards. A novel analysis space for pointer analysis and its application for bug finding. Science of Computer Programming, 75(11):921–942, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Denny:2010:IAG

[1217] Joel E. Denny and Brian A. Malloy. The IELR(1) algorithm for generating minimal LR(1) parser tables for non-LR(1) grammars with conflict resolution. Science of Computer Programming, 75(11):943–979, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jin:2010:DMT

[1218] Ying Jin and David Lorge Parnas. Defining the meaning of tabular mathematical expressions. Science of Computer Programming, 75(11):980–1000, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Rodriguez:2010:IWS

[1219] Juan Manuel Rodriguez, Marco Crasso, Alejandro Zunino, and Marcelo Campo. Improving Web Service descriptions for effective service discovery. Science of Computer Programming, 75(11):1001–1021, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Apel:2010:AFA

[1220] Sven Apel, Christian Lengauer, Bernhard Möller, and Christian Kästner. An algebraic foundation for automatic feature-based program synthesis. Science of Computer Programming, 75 (11):1022–1047, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chen:2010:TDW

[1221] Kung Chen, Shu-Chun Weng, Meng Wang, Siau-Cheng Khoo, and Chung-Hsin Chen. Type-directed weaving of aspects for polymorphically typed functional languages. Science of Computer Programming, 75(11):1048–1076, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Haubl:2010:CES

[1222] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Compact and efficient strings for Java. Science of Computer Programming, 75(11):1077– 1094, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cleophas:2010:NTS

[1223] Loek Cleophas, Bruce W. Watson, and Gerard Zwaan. A new taxonomy of sublinear right-to-left scanning keyword pattern matching algorithms. Science of Computer Programming, 75(11):1095–1112, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Conejero:2010:MEA

[1224] José M. Conejero, Juan Hernández, Elena Jurado, and Klaas van den Berg. Mining early aspects based on syntactical and dependency analyses. Science of Computer Programming, 75(11):1113–1141, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Maurer:2010:SSI

[1225] Ward Douglas Maurer. Simple section interchange and properties of noncomputable functions. Science of Computer Programming, 75(11):1142– 1164, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Donaldson:2010:TIS

[1226] Alastair F. Donaldson and Simon J. Gay. Type inference and strong static type checking for Promela. Science of Computer Programming, 75(11):1165– 1191, November 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Broy:2010:MSS

[1227] Manfred Broy. Multifunctional software systems: Structured modeling and specification of functional requirements. Science of Computer Programming, 75(12):1193–1214, December 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carlson:2010:REE

[1228] Jan Carlson and Björn Lisper. A resource-efficient event algebra. Science of Computer Programming, 75 (12):1215–1234, December 1, 2010.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Tanter:2010:SSD

[1229] Éric Tanter, Johan Fabry, Rémi Douence, Jacques Noyé, and Mario Südholt. Scoping strategies for distributed aspects. Science of Computer Programming, 75(12):1235–1261, December 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Boiten:2010:IRS

[1230] Eerke Boiten and John Derrick. Incompleteness of relational simulations in the blocking paradigm. Science of Computer Programming, 75(12):1262– 1269, December 1, 2010. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Canal:2011:Pa

[1231] Carlos Canal, Pascal Poizat, and Mirko Viroli. Preface. Science of Computer Programming, 76(1):1–2, January 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Talcott:2011:CTC

[1232] Carolyn Talcott, Marjan Sirjani, and Shangping Ren. Comparing three coordination models: Reo, ARC, and PBRD. Science of Computer Programming, 76(1):3–22, January 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Krause:2011:MDR

[1233] Christian Krause, Ziyan Maraikar, Alexander Lazovik, and Farhad Arbab. Modeling dynamic reconfigurations in Reo using high-level replacement systems. Science of Computer Programming, 76(1):23–36, January 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ricci:2011:SAO

[1234] Alessandro Ricci, Mirko Viroli, and Giulio Piancastelli. simpA: an agentoriented approach for programming concurrent applications on top of Java. Science of Computer Programming, 76 (1):37–62, January 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2011:EB

[1235] Anonymous. Editorial Board. Science of Computer Programming, 76 (1):??, January 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cofer:2011:PSI

[1236] Darren Cofer, Alessandro Fantechi, Stefan Leue, and Pedro Merino. Preface to the special issue on Formal Methods for Industrial Critical Systems (FMICS 2007 + FMICS 2008). Science of Computer Programming, 76 (2):63–64, February 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Li:2011:FSM

[1237] Guodong Li, Robert Palmer, Michael DeLisi, Ganesh Gopalakrishnan, and Robert M. Kirby. Formal specification of MPI 2.0: Case study in specifying a practical concurrent programming API. Science of Computer Programming, 76(2):65–81, February 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanGastel:2011:DSF

[1238] Bernard van Gastel, Leonard Lensink, Sjaak Smetsers, and Marko van Eekelen. Deadlock and starvation free reentrant readers–writers: a case study combining model checking with theorem proving. Science of Computer Programming, 76(2):82–99, February 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schlich:2011:ASA

[1239] Bastian Schlich, Jörg Brauer, and Stefan Kowalewski. Application of static analyses for state-space reduction to the microcontroller binary code. Science of Computer Programming, 76(2): 100–118, February 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

terBeek:2011:SEB

[1240] Maurice H. ter Beek, Alessandro Fantechi, Stefania Gnesi, and Franco Mazzanti. A state/event-based modelchecking approach for the analysis of abstract system properties. Science of Computer Programming, 76(2):119– 135, February 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Edelkamp:2011:FME

[1241] S. Edelkamp, D. Sulewski, J. Barnat, L. Brim, and P. Šimeček. Flash memory efficient LTL model checking. Science of Computer Programming, 76(2): 136–157, February 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Audebaud:2011:P

[1242] Philippe Audebaud and Christine Paulin-Mohring. Preface. Science of Computer Programming, 76(3):159, March 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Backhouse:2011:EAE

[1243] Roland Backhouse and João F. Ferreira. On Euclid’s algorithm and elementary number theory. Science of Computer Programming, 76(3):160– 180, March 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Desharnais:2011:IAD

[1244] Jules Desharnais and Georg Struth. Internal axioms for domain semirings. Science of Computer Programming, 76 (3):181–203, March 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Matthes:2011:MFN

[1245] Ralph Matthes. Map fusion for nested datatypes in intensional type theory. Science of Computer Programming, 76 (3):204–224, March 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Siek:2011:SIL

[1246] Jeremy G. Siek. Special issue on library-centric software design (LCSD 2006). Science of Computer Programming, 76(4):225–226, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Russell:2011:DAL

[1247] Francis P. Russell, Michael R. Mellor, Paul H. J. Kelly, and Olav Beckmann. DESOLA: an active linear algebra library using delayed evaluation and runtime code generation. Science of Computer Programming, 76 (4):227–242, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bourdev:2011:ERT

[1248] Lubomir Bourdev and Jaakko Järvi.

Efficient run-time dispatching in generic programming with minimal code bloat. Science of Computer Programming, 76(4):243–257, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Oancea:2011:AGE

[1249] Cosmin E. Oancea and Stephen M.

Watt. An architecture for generic extensions. Science of Computer Programming, 76(4):258–277, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Munkby:2011:AES

[1250] Gustav Munkby and Sibylle Schupp.

Automating exception-safety classification. Science of Computer Programming, 76(4):278–289, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Solodkyy:2011:ETS

[1251] Yuriy Solodkyy and Jaakko Järvi. Extending type systems in a library: Type-safe XML processing in C++. Science of Computer Programming, 76 (4):290–306, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Pion:2011:GLE

[1252] Sylvain Pion and Andreas Fabri. A generic lazy evaluation scheme for exact geometric computations. Science of Computer Programming, 76 (4):307–323, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Heinzl:2011:GTL

[1253] René Heinzl and Philipp Schwaha. A generic topology library. Science of Computer Programming, 76 (4):324–346, April 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Gluck:2011:SIG

[1254] Robert Glück and Eelco Visser. Special issue on Generative Programming and Component Engineering (Selected Papers from GPCE 2004/2005). Science of Computer Programming, 76 (5):347–348, May 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carette:2011:MSP

[1255] Jacques Carette and Oleg Kiselyov. Multi-stage programming with functors and monads: Eliminating abstraction overhead from generic code. Science of Computer Programming, 76 (5):349–375, May 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Huang:2011:SSP

[1256] Shan Shan Huang, David Zook, and Yannis Smaragdakis. Statically safe program generation with SafeGen. Science of Computer Programming, 76 (5):376–391, May 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lutteroth:2011:TSR

[1257] Christof Lutteroth, Dirk Draheim, and Gerald Weber. A type system for reflective program generators. Science of Computer Programming, 76 (5):392–422, May 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Siek:2011:LGP

[1258] Jeremy G. Siek and Andrew Lumsdaine. A language for generic programming in the large. Science of Computer Programming, 76(5):423–465, May 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ramalingam:2011:SIP

[1259] G. Ramalingam and Eelco Visser. Special issue on Partial Evaluation and Program Manipulation (selected papers from PEPM 2007). Science of Computer Programming, 76(6):467– 468, June 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Carette:2011:PEM

[1260] Jacques Carette and Michael Kucera.

Partial evaluation of Maple. Science of Computer Programming, 76 (6):469–491, June 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Moller:2011:XGP

[1261] Anders Møller and Michael Schwartzbach. XML graphs in program analysis. Science of Computer Programming, 76 (6):492–515, June 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Cunha:2011:TSS

[1262] Alcino Cunha and Joost Visser. Transformation of structure-shy programs with application to XPath queries and strategic functions. Science of Computer Programming, 76(6):516– 539, June 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bergstra:2011:E

[1263] J. A. Bergstra. Editorial. Science of Computer Programming, 76(7):541, July 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hesselink:2011:QBM

[1264] Wim H. Hesselink and Alex A. Aravind. Queue based mutual exclusion with linearly bounded overtaking. Science of Computer Programming, 76 (7):542–554, July 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Ducournau:2011:MSM

[1265] Roland Ducournau and Jean Privat. Metamodeling semantics of multiple inheritance. Science of Computer Programming, 76(7):555–586, July 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Male:2011:FIA

[1266] Chris Male, David J. Pearce, Alex Potanin, and Constantine Dymnikov. Formalisation and implementation of an algorithm for bytecode verification of @NonNull types. Science of Computer Programming, 76(7):587– 608, July 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Shankar:2011:MVS

[1267] Natarajan Shankar and K. Subramani. A mechanical verification of the stressing algorithm for negative cost cycle detection in networks. Science of Computer Programming, 76 (7):609–626, July 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanLamsweerde:2011:HHR

[1268] Axel van Lamsweerde. The humble humorous researcher: a tribute to Michel Sintzoff. Science of Computer Programming, 76(8):627–630, August 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Canal:2011:Pb

[1269] Carlos Canal, Pascal Poizat, and Marjan Sirjani. Preface. Science of Computer Programming, 76(8):631– 632, August 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wong:2011:FAB

[1270] Peter Y. H. Wong and Jeremy Gibbons. Formalisations and applications of BPMN. Science of Computer Programming, 76(8):633–650, August 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Meng:2011:SRC

[1271] Sun Meng, Farhad Arbab, and Christel Baier. Synthesis of Reo circuits from scenario-based interaction specifications. Science of Computer Programming, 76(8):651–680, August 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Clarke:2011:CBC

[1272] Dave Clarke, José Proença, Alexander Lazovik, and Farhad Arbab. Channelbased coordination via constraint satisfaction. Science of Computer Programming, 76(8):681–710, August 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Andova:2011:DCP

[1273] S. Andova, L. P. J. Groenewegen, and E. P. de Vink. Dynamic consistency in process algebra: From Paradigm to ACP. Science of Computer Programming, 76(8):711–735, August 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Boiten:2011:SPR

[1274] Eerke Boiten, John Derrick, and Gerhard Schellhorn. Selected papers of the Refinement Workshop Turku (2008). Science of Computer Programming, 76(9):737–738, September 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Hesselink:2011:SRC

[1275] Wim H. Hesselink. Simulation refinement for concurrency verification. Science of Computer Programming, 76(9):

739–755, September 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schellhorn:2011:CFA

[1276] Gerhard Schellhorn. Completeness of fair ASM refinement. Science of Computer Programming, 76(9):756– 773, September 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Reeves:2011:CRD

[1277] Steve Reeves and David Streader. Contexts, refinement and determinism. Science of Computer Programming, 76(9): 774–791, September 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Oliveira:2011:TLR

[1278] Marcel Oliveira, Frank Zeyda, and Ana Cavalcanti. A tactic language for refinement of state-rich concurrent specifications. Science of Computer Programming, 76(9):792–833, September 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Leuschel:2011:SPI

[1279] Michael Leuschel and Heike Wehrheim.

Selected papers on Integrated Formal Methods (iFM09). Science of Computer Programming, 76(10):835– 836, October 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Schneider:2011:CSI

[1280] Steve Schneider and Helen Treharne.

Changing system interfaces consistently: a new refinement strategy for CSP ||B. Science of Computer Programming, 76(10):837–860, October 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Weiss:2011:PAP

[1281] Benjamin Weiß. Predicate abstraction in a program logic calculus. Science of Computer Programming, 76(10): 861–876, October 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Benes:2011:POR

[1282] N. Beneš, L. Brim, B. Buhnova, I. Černá, J. Sochor, and P. Vařeková. Partial order reduction for state/event LTL with application to componentinteraction automata. Science of Computer Programming, 76(10):877–890, October 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Colvin:2011:SBT

[1283] Robert J. Colvin and Ian J. Hayes. A semantics for Behavior Trees using CSP with specification commands. Science of Computer Programming, 76 (10):891–914, October 1, 2011. CO-DEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Dovland:2011:IRL

[1284] Johan Dovland, Einar Broch Johnsen, Olaf Owe, and Martin Steffen. Incremental reasoning with lazy behavioral subtyping for multiple inheritance. Science of Computer Programming, 76(10):915–941, October 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wong:2011:PSW

[1285] Peter Y. H. Wong and Jeremy Gibbons.

Property specifications for workflow modelling. Science of Computer Programming, 76(10):942–967, October 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Horspool:2011:PPP

[1286] R. Nigel Horspool, Ralf Gitzel, Giacomo Cabri, and Markus Aleksy. Principles and practice of programming in Java 2008—special issue. Science of Computer Programming, 76 (11):969, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Kabanov:2011:DSF

[1287] Jevgeni Kabanov, Michael Hunger, and Rein Raudjärv. On designing safe and flexible embedded DSLs with Java 5. Science of Computer Programming, 76(11):970–991, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Bettini:2011:DOC

[1288] Lorenzo Bettini, Viviana Bono, and Betti Venneri. Delegation by object composition. Science of Computer Programming, 76(11):992–1014, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Villazon:2011:CAW

[1289] Alex Villazón, Walter Binder, Philippe Moret, and Danilo Ansaloni. Comprehensive aspect weaving for Java. Science of Computer Programming, 76 (11):1015–1036, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Wagner:2011:SJV

[1290] Gregor Wagner, Andreas Gal, and Michael Franz. “slimming” a Java virtual machine by way of cold code removal and optimistic partial program loading. Science of Computer Programming, 76(11):1037–1053, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Jovic:2011:LLP

[1291] Milan Jovic and Matthias Hauswirth. Listener latency profiling: Measuring the perceptible performance of interactive Java applications. Science of Computer Programming, 76(11):1054– 1072, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Haubl:2011:ECE

[1292] Christian Häubl, Christian Wimmer, and Hanspeter Mössenböck. Erratum to “Compact and Efficient Strings for Java” [Science of Computer Programming 75 (2010) 1077–1094]. Science of Computer Programming, 76 (11):1073, November 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Zaidman:2011:PSI

[1293] Andy Zaidman and Johan Brichau. Preface to the special issue on software evolution, adaptability and variability. Science of Computer Programming, 76(12):1075–1077, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Arbuckle:2011:SSE

[1294] Tom Arbuckle. Studying software evolution using artefacts’ shared information content. Science of Computer Programming, 76(12):1078–1097, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Arias:2011:TSR

[1295] Trosky B. Callo Arias, Paris Avgeriou, Pierre America, Krelis Blom, and Sergiy Bachynskyy. A top-down strategy to reverse architecting execution views for a large and complex softwareintensive system: an experience report. Science of Computer Programming, 76(12):1098–1112, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Castro:2011:DCD

[1296] Sergio Castro, Coen De Roover, Andy Kellens, Angela Lozano, Kim Mens, and Theo D’Hondt. Diagnosing and correcting design inconsistencies in source code with logical abduction. Science of Computer Programming, 76 (12):1113–1129, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Classen:2011:TBA

[1297] Andreas Classen, Quentin Boucher, and Patrick Heymans. A text-based approach to feature modelling: Syntax and semantics of TVL. Science of Computer Programming, 76(12):1130– 1143, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

DiCosmo:2011:SSE

[1298] Roberto Di Cosmo, Davide Di Ruscio, Patrizio Pelliccione, Alfonso Pierantonio, and Stefano Zacchiroli. Supporting software evolution in componentbased FOSS systems. Science of Computer Programming, 76(12):1144– 1160, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Karus:2011:PMX

[1299] Siim Karus and Marlon Dumas. Predicting the maintainability of XSL transformations. Science of Computer Programming, 76(12):1161–1176, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Laval:2011:SSV

[1300] Jannik Laval, Simon Denier, Stéphane Ducasse, and Jean-Rémy Falleri. Supporting simultaneous versions for software evolution assessment. Science of Computer Programming, 76(12):1177– 1193, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Lincke:2011:OIC

[1301] Jens Lincke, Malte Appeltauer, Bastian Steinert, and Robert Hirschfeld. An open implementation for contextoriented layer composition in ContextJS. Science of Computer Programming, 76(12):1194–1209, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Mannaert:2011:TRS

[1302] Herwig Mannaert, Jan Verelst, and Kris Ven. The transformation of requirements into software primitives: Studying evolvability based on systems theoretic stability. Science of Computer Programming, 76(12):1210– 1222, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Meyers:2011:FEM

[1303] Bart Meyers and Hans Vangheluwe. A framework for evolution of modelling languages. Science of Computer Programming, 76(12):1223–1246, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Parra:2011:UDR

[1304] Carlos Parra, Xavier Blanc, Anthony Cleve, and Laurence Duchien. Unifying design and runtime software adaptation using aspect models. Science of Computer Programming, 76(12):1247– 1260, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

vanderSpek:2011:ADT

[1305] Pieter van der Spek and Steven Klusener. Applying a dynamic threshold to improve cluster detection of LSI. Science of Computer Programming, 76(12):1261–1274, December 1, 2011. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Chung:2012:SSS

[1306] Lawrence Chung, Manuel Noguera, Nary Subramanian, and José Luis Garrido. System and software solution-oriented architectures. Science of Computer Programming, 77 (1):1–3, January 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001419.

Resinas:2012:BSA

[1307] Manuel Resinas, Pablo Fernández, and Rafael Corchuelo. A bargaining-specific architecture for supporting automated service agreement negotiation systems. Science of Computer Programming, 77 (1):4–28, January 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001851.

Duque:2012:ICI

[1308] Rafael Duque, María Luisa Rodríguez, María Visitación Hurtado, Crescencio Bravo, and Carlos Rodríguez-Domínguez. Integration of collaboration and interaction analysis mechanisms in a concern-based architecture for groupware systems. Science of Computer Programming, 77 (1):29–45, January 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310000936.

Rodriguez:2012:ASA

[1309] Marcela D. Rodríguez and Jesús Favela. Assessing the SALSA architecture for developing agent-based ambient computing applications. Science of Computer Programming, 77 (1):46–65, January 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310002261.

Lopez:2012:BGB

[1310] Claudia López, Víctor Codocedo, Hernán Astudillo, and Luiz Marcio Cysneiros. Bridging the gap between software architecture rationale formalisms and actual architecture documents: an ontology-driven approach. Science of Computer Programming, 77 (1):66–80, January 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001218.

Anonymous:2012:EB

[1311] Anonymous. Editorial Board. Science of Computer Programming, 77 (1):??, January 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001614.

Hammond:2012:ISI

[1312] Kevin Hammond and Paul H. J.

Kelly. Introduction to the special issue on automatic program generation for embedded systems. Science of Computer Programming, 77(2): 81–82, February 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001924.

Ryssel:2012:ALM

[1313] Uwe Ryssel, Joern Ploennigs, and Klaus Kabitzsch. Automatic library migration for the generation of hardware-in-the-loop models. Science of Computer Programming, 77(2): 83–95, February 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001115.

Ghosal:2012:SCH

[1314] Arkadeb Ghosal, Daniel Iercan, Christoph M. Kirsch, Thomas A. Henzinger, and Alberto Sangiovanni-Vincentelli. Separate compilation of hierarchical real-time programs into linear-bounded Embedded Machine code. Science of Computer Programming, 77(2):96– 112, February 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001103.

Talpin:2012:CDI

[1315] Jean-Pierre Talpin, Julien Ouy, Thierry Gautier, Loïc Besnard, and Paul Le Guernic. Compositional design of isochronous systems. Science of Computer Programming, 77(2): 113–128, February 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001139.

Apel:2012:PSI

[1316] Sven Apel, Christian Lengauer, and Julia Lawall. Preface to the special issue on feature-oriented software development (FOSD 2009). Science of Computer Programming, 77 (3):129–130, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000487.

Olszak:2012:RJP

[1317] Andrzej Olszak and Bo Nørregaard Jørgensen. Remodularizing Java programs for improved locality of feature implementations in source code. Science of Computer Programming, 77 (3):131–151, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001917.

Gunther:2012:RFO

[1318] Sebastian Günther and Sagar Sunkle. rbFeatures: Feature-oriented programming with Ruby. Science of Computer Programming, 77(3): 152–173, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000025.

Apel:2012:ACF

[1319] Sven Apel, Sergiy Kolesnikov, Jörg Liebig, Christian Kästner, Martin Kuhlemann, and Thomas Leich. Access control in featureoriented programming. Science of Computer Programming, 77(3):174– 187, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001528.

Lapadula:2012:UFM

[1320] Alessandro Lapadula, Rosario Pugliese, and Francesco Tiezzi. Using formal methods to develop WS-BPEL applications. Science of Computer Programming, 77(3):189–213, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642311000694.

Ribeiro:2012:CTO

[1321] Leila Ribeiro, Osmar Marchi dos Santos, Fernando Luís Dotti, and Luciana Foss. Correct transformation: From object-based graph grammars to PROMELA. Science of Computer Programming, 77(3):214– 246, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311000979.

Fuentes-Fernandez:2012:MDP

[1322] Rubén Fuentes-Fernández, Juan Pavón, and Francisco Garijo. A modeldriven process for the modernization of component-based systems. Science of Computer Programming, 77 (3):247–269, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001110.

Ruhroth:2012:MER

[1323] Thomas Ruhroth and Heike Wehrheim. Model evolution and refinement. Science of Computer Programming, 77 (3):270–289, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001298.

Valero:2012:TWS

[1324] Valentín Valero, Hermenegilda Macià, Juan José Pardo, María Emilia Cambronero, and Gregorio Díaz. Transforming Web Services Choreographies with priorities and time constraints into prioritized-time colored Petri nets. Science of Computer Programming, 77 (3):290–313, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001407.

Eveleens:2012:QFQ

[1325] J. L. Eveleens, M. van der Pas, and C. Verhoef. Quantifying forecast quality of IT business value. Science of Computer Programming, 77 (3):314–354, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001560.

Wu:2012:WTT

[1326] Chen Wu. WSDL term tokenization methods for IR-style Web services discovery. Science of Computer Programming, 77(3):355– 374, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001705.

Gallardo:2012:MEA

[1327] M. M. Gallardo, C. Joubert, P. Merino, and D. Sanán. A model-extraction approach to verifying concurrent C programs with CADP. Science of Computer Programming, 77(3):375– 392, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001808.

Djoko:2012:APP

[1328] Simplice Djoko Djoko, Rémi Douence, and Pascal Fradet. Aspects preserving properties. Science of Computer Programming, 77(3):393– 422, March 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001870.

Machado:2012:PSP

[1329] Patrícia D. L. Machado. Preface: Selected papers from the Brazilian Symposium on Formal Methods (SBMF 2008). Science of Computer Programming, 77(4): 423, April 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001791.

Perna:2012:MWW

[1330] Juan Perna and Jim Woodcock. Mechanised wire-wise verification of Handel-C synthesis. Science of Computer Programming, 77(4):424– 443, April 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642310000341.

Zeyda:2012:MRA

[1331] Frank Zeyda and Ana Cavalcanti. Mechanical reasoning about families of UTP theories. Science of Computer Programming, 77(4): 444–479, April 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310000481.

daCosta:2012:VGG

[1332] Simone André da Costa and Leila Ribeiro. Verification of graph grammars using a logical approach. Science of Computer Programming, 77 (4):480–504, April 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231000033X.

Myreen:2012:FE

[1333] Magnus O. Myreen and Michael J. C.

Gordon. Function extraction. Science of Computer Programming, 77 (4):505–517, April 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231000184X.

deOliveira:2012:CDP

[1334] Diego Caminha B. de Oliveira, David Déharbe, and Pascal Fontaine. Combining decision procedures by (model- )equality propagation. Science of Computer Programming, 77(4):518– 532, April 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310000729.

daCosta:2012:JSL

[1335] Umberto Souza da Costa, Anamaria Martins Moreira, Martin A. Musicante, and Plácido A. Souza Neto. JCML: a specification language for the runtime verification of Java Card programs. Science of Computer Programming, 77(4): 533–550, April 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310000596.

vanHeesch:2012:SEP

[1336] Uwe van Heesch, Paris Avgeriou, Uwe Zdun, and Neil Harrison. The supportive effect of patterns in architecture decision recovery — a controlled experiment. Science of Computer Programming, 77(5):551– 576, May 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002085.

Fiadeiro:2012:FMS

[1337] José Fiadeiro, Antónia Lopes, and João Abreu. A formal model for service-oriented interactions. Science of Computer Programming, 77 (5):577–608, May 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002255.

Volanschi:2012:PMM

[1338] Nic Volanschi. Pattern matching for the masses using custom notations. Science of Computer Programming, 77 (5):609–635, May 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002243.

Codur:2012:RSE

[1339] K. Burak Codur and Ali H. Dogru.

Regulations and software evolution: an example from the military domain. Science of Computer Programming, 77 (5):636–643, May 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002231.

Hamza:2012:RBR

[1340] H. Hamza and S. Counsell. Regionbased RTSJ memory management: State of the art. Science of Computer Programming, 77(5):644– 659, May 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000032.

Field:2012:SPE

[1341] John Field and Vasco T. Vasconcelos. Selected papers from the Eleventh International Conference on Coordination Models and Languages. Science of Computer Programming, 77 (6):661–662, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001900.

Fok:2012:SFS

[1342] Chien-Liang Fok, Gruia-Catalin Roman, and Chenyang Lu. Servilla: a flexible service provisioning middleware for heterogeneous sensor networks. Science of Computer Programming, 77 (6):663–684, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642310002054.

Bonsangue:2012:MCD

[1343] Marcello Bonsangue, Dave Clarke, and Alexandra Silva. A model of contextdependent component connectors. Science of Computer Programming, 77 (6):685–706, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311000189.

Navabi:2012:DAS

[1344] Armand Navabi, Xiangyu Zhang, and Suresh Jagannathan. Dependence analysis for safe futures. Science of Computer Programming, 77(6): 707–726, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642310001577.

Ritson:2012:MSL

[1345] Carl G. Ritson, Adam T. Sampson, and Frederick R. M. Barnes. Multicore scheduling for lightweight communicating processes. Science of Computer Programming, 77(6): 727–740, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001286.

Zaidman:2012:PSI

[1346] Andy Zaidman and Giuliano Antoniol. Preface to the special issue on reverse engineering (featuring selected papers from WCRE 2009). Science of Computer Programming, 77 (6):741–742, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000207.

Lo:2012:MQT

[1347] David Lo, G. Ramalingam, Venkatesh-Prasad Ranganath, and Kapil Vaswani. Mining quantified temporal rules: Formalism, algorithms, and evaluation. Science of Computer Programming, 77 (6):743–759, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001875.

Bettenburg:2012:ESI

[1348] Nicolas Bettenburg, Weiyi Shang, Walid M. Ibrahim, Bram Adams, Ying Zou, and Ahmed E. Hassan. An empirical study on inconsistent changes to code clones at the release level. Science of Computer Programming, 77 (6):760–776, June 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310002091.

Salaun:2012:PSI

[1349] Gwen Salaün and Marjan Sirjani.

Preface: Special issue on Foundations of Coordination Languages and Software Architectures (selected papers from FOCLASA’09). Science of Computer Programming, 77(7–8): 777–778, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001766.

Kemper:2012:SBV

[1350] S. Kemper. SAT-based verification for timed component connectors. Science of Computer Programming, 77(7– 8):779–798, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311000499.

Meng:2012:CDM

[1351] Sun Meng, Farhad Arbab, Bernhard K. Aichernig, Lacramioara Astefanoaei, Frank S. de Boer, and Jan Rutten. Connectors as designs: Modeling, refinement and test case generation. Science of Computer Programming, 77(7– 8):799–822, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001006.

Henrio:2012:ATF

[1352] Ludovic Henrio, Florian Kammüller, and Bianca Lutz. ASP_{fun}: a typed functional active object calculus. Science of Computer Programming, 77(7– 8):823–847, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311000037.

Pourvatan:2012:SER

[1353] Bahman Pourvatan, Marjan Sirjani, Hossein Hojjat, and Farhad Arbab. Symbolic execution of Reo circuits using constraint automata. Science of Computer Programming, 77(7–8):

848–869, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000980.

Duran:2012:GFP

[1354] Francisco Durán, Meriem Ouederni, and Gwen Salaün. A generic framework for n-protocol compatibility checking. Science of Computer Programming, 77 (7–8):870–886, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000955.

Arbab:2012:FSE

[1355] Farhad Arbab and Marjan Sirjani. Fundamentals of software engineering (selected papers of FSEN ’09). Science of Computer Programming, 77(7– 8):887–888, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001304.

Aceto:2012:RFD

[1356] Luca Aceto, Arnar Birgisson, Anna Ingolfsdottir, MohammadReza Mousavi, and Michel A. Reniers. Rule formats for determinism and idempotence. Science of Computer Programming, 77(7– 8):889–907, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310000717.

Schuppan:2012:TNU

[1357] Viktor Schuppan. Towards a notion of unsatisfiable and unrealizable cores for LTL. Science of Computer Programming, 77(7–8): 908–939, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310002030.

Klein:2012:MNU

[1358] Dominik Klein, Frank G. Radmacher, and Wolfgang Thomas. Moving in a network under random failures: a complexity analysis. Science of Computer Programming, 77(7– 8):940–954, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001176.

Kleijn:2012:SCA

[1359] Jetty Kleijn and Maciej Koutny. Step coverability algorithms for communicating systems. Science of Computer Programming, 77(7–8): 955–967, July 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310002029.

Mernik:2012:SIP

[1360] Marjan Mernik and Barrett R. Bryant. Special issue on the Programming Languages track at the 24th ACM Symposium on Applied Computing. Science of Computer Programming, 77(9): 969–970, August 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000153.

Biggar:2012:PSA

[1361] Paul Biggar, Edsko de Vries, and David Gregg. A practical solution for achieving language compatibility in scripting language compilers. Science of Computer Programming, 77(9): 971–989, August 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000165.

Cervelle:2012:BUC

[1362] Julien Cervelle, Rémi Forax, Gautier Loyauté, and Gilles Roussel. Banzai+Tatoo: Using cuttingedge parsers for implementing highperformance servers. Science of Computer Programming, 77(9):990– 1005, August 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000177.

Chin:2012:AVS

[1363] Wei-Ngan Chin, Cristina David, Huu Hai Nguyen, and Shengchao Qin. Automated verification of shape, size and bag properties via user-defined predicates in separation logic. Science of Computer Programming, 77(9): 1006–1036, August 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001516.

Roggenbach:2012:P

[1364] Markus Roggenbach. Preface. Science of Computer Programming, 77 (10–11):1037–1038, September 1, 2012.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764231100147X.

Narkawicz:2012:PCC

[1365] Anthony Narkawicz, César Muñoz, and Gilles Dowek. Provably correct conflict prevention bands algorithms. Science of Computer Programming, 77(10–11):1039–1057, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001481.

Affeldt:2012:CAF

[1366] Reynald Affeldt, David Nowak, and Kiyoshi Yamada. Certifying assembly with formal security proofs: the case of BBS. Science of Computer Programming, 77(10–11):1058– 1074, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001493.

Mateescu:2012:SDF

[1367] Radu Mateescu and Anton Wijs. Sequential and distributed on-the-fly computation of weak tau-confluence. Science of Computer Programming, 77(10–11):1075–1094, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764231100150X.

Dubrovin:2012:ESS

[1368] Jori Dubrovin, Tommi Junttila, and Keijo Heljanko. Exploiting step semantics for efficient bounded model checking of asynchronous systems. Science of Computer Programming, 77(10–11):1095– 1121, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001511.

Damm:2012:EFS

[1369] Werner Damm, Henning Dierks, Stefan Disch, Willem Hagemann, Florian Pigorsch, Christoph Scholl, Uwe Waldmann, and Boris Wirtz. Exact and fully symbolic verification of linear hybrid automata with large discrete state spaces. Science of Computer Programming, 77(10–11):1122– 1150, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001523.

Miyazawa:2012:ROM

[1370] Alvaro Miyazawa and Ana Cavalcanti. Refinement-oriented models of Stateflow charts. Science of Computer Programming, 77(10–11):1151– 1177, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001535.

Palikareva:2012:SSC

[1371] Hristina Palikareva, Joël Ouaknine, and A. W. Roscoe. SAT-solving in CSP trace refinement. Science of Computer Programming, 77(10–11):1178– 1197, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001547.

Papanikolaou:2012:RCP

[1372] Nikolaos Papanikolaou, Sadie Creese, and Michael Goldsmith. Refinement checking for privacy policies. Science of Computer Programming, 77(10–11):1198–1209, September 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642311001559.

Cavalcanti:2012:SII

[1373] Ana Cavalcanti. Special issue: International Conference on Formal Engineering Methods — ICFEM 2009. Science of Computer Programming, 77 (12):1211, October 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001062.

Bertrand:2012:MEC

[1374] Nathalie Bertrand, Axel Legay, Sophie Pinchinat, and Jean-Baptiste Raclet. Modal event-clock specifications for timed component-based design. Science of Computer Programming, 77(12):1212–1234, October 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642311000190.

Bae:2012:VHP

[1375] Kyungmin Bae, Peter Csaba Ölveczky, Thomas Huining Feng, Edward A. Lee, and Stavros Tripakis. Verifying hierarchical Ptolemy II discrete-event models using Real-Time Maude. Science of Computer Programming, 77(12): 1235–1271, October 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310001863.

Barnat:2012:FPM

[1376] Jirí Barnat, Lubos Brim, and Petr Rockai. On-the-fly parallel model checking algorithm that is optimal for verification of weak LTL properties. Science of Computer Programming, 77(12):1272– 1288, October 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000682.

Ahrendt:2012:SCV

[1377] Wolfgang Ahrendt and Maximilian Dylla. A system for compositional verification of asynchronous objects. Science of Computer Programming, 77(12):1289–1309, October 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642310001553.

Meng:2012:FAC

[1378] Sun Meng and Bernhard Schätz. Formal Aspects of Component Software (FACS’09). Science of Computer Programming, 78(1):1–2, November 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312000330.

Khakpour:2012:FME

[1379] Narges Khakpour, Saeed Jalili, Carolyn Talcott, Marjan Sirjani, and MohammadReza Mousavi. Formal modeling of evolving self-adaptive systems. Science of Computer Programming, 78 (1):3–26, November 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001742.

Lumpe:2012:PRC

[1380] Markus Lumpe. Partition refinement of Component Interaction Automata. Science of Computer Programming, 78 (1):27–45, November 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001912.

Canal:2012:SRS

[1381] Carlos Canal, Javier Cámara, and Gwen Salaün. Structural reconfiguration of systems under behavioral adaptation. Science of Computer Programming, 78(1):46– 64, November 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001730.

Zambrano:2012:EAI

[1382] Arturo Zambrano, Johan Fabry, and Silvia Gordillo. Expressing aspectual interactions in requirements engineering: Experiences, problems and solutions. Science of Computer Programming, 78(1):65– 92, November 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002267.

Khalilian:2012:IMT

[1383] Alireza Khalilian, Mohammad Abdollahi Azgomi, and Yalda Fazlalizadeh. An improved method for test case prioritization by incorporating historical test case data. Science of Computer Programming, 78(1): 93–116, November 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000159.

Piveta:2012:ESA

[1384] Eduardo Kessler Piveta, Ana Moreira, Marcelo Soares Pimenta, João Araújo, Pedro Guerreiro, and R. Tom Price. An empirical study of aspect-oriented metrics. Science of Computer Programming, 78(1):117– 144, November 1, 2012. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000287.

Clarke:2013:PSI

[1385] Dave Clarke and Gul Agha. Preface to the special issue on Coordination Models and Languages (Coordination 2010). Science of Computer Programming, 78(2):145– 146, February 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000858.

Sivaramakrishnan:2013:ES

[1386] K. C. Sivaramakrishnan, Mohammad Qudeisat, Lukasz Ziarek, Karthik Nagaraj, and Patrick Eugster. Efficient sessions. Science of Computer Programming, 78(2):147–167, February 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642312000494.

Philips:2013:NOS

[1387] E. Philips, R. Van Der Straeten, and V. Jonckers. NOW: Orchestrating services in a nomadic network using a dedicated workflow language. Science of Computer Programming, 78(2): 168–194, February 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001894.

Fok:2013:ASP

[1388] Chien-Liang Fok, Gruia-Catalin Roman, and Chenyang Lu. Adaptive service provisioning for enhanced energy efficiency and flexibility in wireless sensor networks. Science of Computer Programming, 78(2):195– 217, February 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311002280.

Bettini:2013:CTB

[1389] Lorenzo Bettini, Ferruccio Damiani, Kathrin Geilmann, and Jan Schäfer. Combining traits with boxes and ownership types in a Java-like setting. Science of Computer Programming, 78(2):

218–247, February 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001833.

Acciai:2013:BCR

[1390] Lucia Acciai, Michele Boreale, and Gianluigi Zavattaro. Behavioural contracts with request-response operations. Science of Computer Programming, 78(2):248–267, February 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642311001845.

Crampton:2013:PTS

[1391] Jason Crampton. A personal tribute to Steve Barker. Science of Computer Programming, 78(3): 269, March 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001864.

Frappier:2013:ASM

[1392] Marc Frappier, Uwe Glässer, Sarfraz Khurshid, Régine Laleau, and Steve Reeves. Abstract state machines, Alloy, B and Z: Selected papers from ABZ 2010. Science of Computer Programming, 78(3): 270–271, March 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000906.

Hallerstede:2013:VFM

[1393] Stefan Hallerstede, Michael Leuschel, and Daniel Plagge. Validation of formal models by refinement animation. Science of Computer Programming, 78 (3):272–292, March 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000918.

Ireland:2013:RMC

[1394] Andrew Ireland, Gudmund Grov, Maria Teresa Llano, and Michael Butler. Reasoned modelling critics: Turning failed proofs into modelling guidance. Science of Computer Programming, 78(3):293– 309, March 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231100092X.

Deharbe:2013:ISS

[1395] David Déharbe. Integration of SMTsolvers in B and Event-B development environments. Science of Computer Programming, 78(3):310– 326, March 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000931.

Reynolds:2013:MJB

[1396] Mark C. Reynolds. Modeling the Java Bytecode Verifier. Science of Computer Programming, 78(3): 327–342, March 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311000943.

duBousquet:2013:SSM

[1397] Lydie du Bousquet, Jeremy S. Bradbury, and Gordon Fraser. Special section on Mutation testing (Mutation 2010). Science of Computer Programming, 78(4):343–344, April 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312001219.

Clark:2013:SMT

[1398] John A. Clark, Haitao Dan, and Robert M. Hierons. Semantic mutation testing. Science of Computer Programming, 78(4):345–363, April 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642311000992.

Mateo:2013:MMC

[1399] Pedro Reales Mateo, Macario Polo Usaola, and Jeff Offutt. Mutation at the multi-class and system levels. Science of Computer Programming, 78 (4):364–387, April 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000305.

Mernik:2013:SIP

[1400] Marjan Mernik and Barrett R. Bryant.

Special issue on the Programming Languages track at the 25th ACM Symposium on Applied Computing. Science of Computer Programming, 78 (4):388–389, April 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001311.

Popeea:2013:DAP

[1401] Corneliu Popeea and Wei-Ngan Chin. Dual analysis for proving safety and finding bugs. Science of Computer Programming, 78(4):390– 411, April 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001323.

Rodrigues:2013:POM

[1402] Eduardo R. Rodrigues, Philippe O. A. Navaux, Jairo Panetta, and Celso L. Mendes. Preserving the original MPI semantics in a virtualized processor environment. Science of Computer Programming, 78 (4):412–421, April 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001335.

Aleksy:2013:F

[1403] Markus Aleksy, Ralf Gitzel, Andreas Krall, Hanspeter Mössenböck, Christian W. Probst, and Ben Stephenson. Foreword. Science of Computer Programming, 78(5): 423–424, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001821.

Taboada:2013:JHP

[1404] Guillermo L. Taboada, Sabela Ramos, Roberto R. Expósito, Juan Touriño, and Ramón Doallo. Java in the High Performance Computing arena: Research, practice and experience. Science of Computer Programming, 78 (5):425–444, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001420.

Tong:2013:SSG

[1405] Liangliang Tong and Francis C. M.

Lau. Skew-space garbage collection. Science of Computer Programming, 78 (5):445–457, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001432.

Leung:2013:PEJ

[1406] Alan Leung, Ondrej Lhoták, and Ghulam Lashari. Parallel execution of Java loops on Graphics Processing Units. Science of Computer Programming, 78 (5):458–480, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001444.

Wurthinger:2013:USD

[1407] Thomas Würthinger, Christian Wimmer, and Lukas Stadler. Unrestricted and safe dynamic code evolution for Java. Science of Computer Programming, 78(5):481– 498, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001456.

Hauswirth:2013:TJP

[1408] Matthias Hauswirth and Andrea Adamoli. Teaching Java programming with the Informa clicker system. Science of Computer Programming, 78 (5):499–520, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001468.

Bettini:2013:TPL

[1409] Lorenzo Bettini, Ferruccio Damiani, Ina Schaefer, and Fabio Strocco. TraitRecordJ: a programming language with traits and records. Science of Computer Programming, 78 (5):521–541, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001572.

Casadei:2013:PSS

[1410] Matteo Casadei, Marco Mamei, and Ronaldo Menezes. Preface to the Special section on Self-Organizing Coordination. Science of Computer Programming, 78(5): 542–543, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000300.

Dolev:2013:SDS

[1411] Shlomi Dolev and Nir Tzachar. Spanders: Distributed spanning expanders. Science of Computer Programming, 78 (5):544–555, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001840.

Viroli:2013:CSC

[1412] Mirko Viroli. On competitive selfcomposition in pervasive services. Science of Computer Programming, 78 (5):556–568, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001852.

Nardini:2013:STC

[1413] Elena Nardini, Andrea Omicini, and Mirko Viroli. Semantic tuple centres. Science of Computer Programming, 78 (5):569–582, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001876.

Muldoon:2013:DCO

[1414] Conor Muldoon, Gregory M. P.

O’Hare, Michael J. O’Grady, Richard Tynan, and Niki Trigoni. Distributed constraint optimisation for resource limited sensor networks. Science of Computer Programming, 78 (5):583–593, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001888.

Bai:2013:SOS

[1415] Linge Bai, Manolya Eyiyurekli, Peter I. Lelkes, and David E. Breen. Self-organized sorting of heterotypic agents via a chemotaxis paradigm. Science of Computer Programming, 78 (5):594–611, May 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231200192X.

Mernik:2013:SSP

[1416] Marjan Mernik and Barrett R. Bryant. Special section on the Programming Languages track at the 26th ACM Symposium on Applied Computing. Science of Computer Programming, 78 (6):613–614, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312002122.

Dinkelaker:2013:ICS

[1417] Tom Dinkelaker, Michael Eichberg, and Mira Mezini. Incremental concrete syntax for embedded languages with support for separate compilation. Science of Computer Programming, 78 (6):615–632, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312002134.

Viroli:2013:OSP

[1418] Mirko Viroli, Jacob Beal, and Kyle Usbeck. Operational semantics of proto. Science of Computer Programming, 78 (6):633–656, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312002146.

Acher:2013:FDS

[1419] Mathieu Acher, Philippe Collet, Philippe Lahire, and Robert B. France. FAMILIAR: a domainspecific language for large scale management of feature models. Science of Computer Programming, 78 (6):657–681, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002158.

Cabri:2013:PSS

[1420] Giacomo Cabri and Maria Ganzha.

Preface to the special section on agent-oriented design methods and programming techniques for distributed computing in dynamic and complex environments. Science of Computer Programming, 78(6): 682–683, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000020.

Dam:2013:TNG

[1421] Hoa Khanh Dam and Michael Winikoff.

Towards a next-generation AOSE methodology. Science of Computer Programming, 78(6):684–694, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642311002279.

Garcia:2013:MDC

[1422] Emilia Garcia, Adriana Giret, and Vicente Botti. A Model-Driven CASE tool for developing and verifying regulated open MAS. Science of Computer Programming, 78 (6):695–704, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001869.

Molesini:2013:SAO

[1423] Ambra Molesini, Matteo Casadei, Andrea Omicini, and Mirko Viroli. Simulation in Agent-Oriented Software Engineering: the SODA case study. Science of Computer Programming, 78 (6):705–714, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001778.

Vigueras:2013:SMS

[1424] Guillermo Vigueras, Juan M. Orduña, Miguel Lozano, and Yvon Jégou. A scalable multiagent system architecture for interactive applications. Science of Computer Programming, 78 (6):715–724, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001729.

Fortino:2013:ALT

[1425] Giancarlo Fortino and Francesco Rango. An application-level technique based on recursive hierarchical state machines for agent execution state capture. Science of Computer Programming, 78(6): 725–746, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231100178X.

Boissier:2013:MAO

[1426] Olivier Boissier, Rafael H. Bordini, Jomi F. Hübner, Alessandro Ricci, and Andrea Santi. Multi-agent oriented programming with JaCaMo. Science of Computer Programming, 78 (6):747–761, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231100181X.

Ilie:2013:MAA

[1427] Sorin Ilie and Costin Badica. Multiagent approach to distributed ant colony optimization. Science of Computer Programming, 78(6):762– 774, June 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001717.

Alpuente:2013:PSS

[1428] María Alpuente, Christophe Joubert, Stefan Kowalewski, and Marco Roveri. Preface to the special section on Formal Methods for Industrial Critical Systems (FMICS 2009 + FMICS 2010). Science of Computer Programming, 78 (7):775–777, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001074.

Evangelista:2013:DSS

[1429] Sami Evangelista and Lars Michael Kristensen. Dynamic state space partitioning for external memory state space exploration. Science of Computer Programming, 78(7): 778–795, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311001754.

Almeida:2013:FVS

[1430] J. Bacelar Almeida, Manuel Barbosa, Jorge S. Pinto, and Bárbara Vieira. Formal verification of side-channel countermeasures using self-composition. Science of Computer Programming, 78(7):796– 812, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001857.

Goodloe:2013:CVC

[1431] Alwyn E. Goodloe and César A. Muñoz. Compositional verification of a communication protocol for a remotely operated aircraft. Science of Computer Programming, 78 (7):813–827, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311001882.

Ferrari:2013:MRC

[1432] Alessio Ferrari, Alessandro Fantechi, Gianluca Magnani, Daniele Grasso, and Matteo Tempestini. The Metrô Rio case study. Science of Computer Programming, 78(7): 828–842, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000676.

Mateescu:2013:MCP

[1433] Radu Mateescu and Wendelin Serwe. Model checking and performance evaluation with CADP illustrated on sharedmemory mutual exclusion protocols.

Science of Computer Programming, 78 (7):843–861, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000123.

Brauer:2013:AIM

[1434] Jörg Brauer, Andy King, and Stefan Kowalewski. Abstract interpretation of microcontroller code: Intervals meet congruences. Science of Computer Programming, 78(7): 862–883, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001116.

Iliasov:2013:DMR

[1435] Alexei Iliasov, Elena Troubitsyna, Linas Laibinis, Alexander Romanovsky, Kimmo Varpaaniemi, Dubravka Ilic, and Timo Latvala. Developing mode-rich satellite software by refinement in Event-B. Science of Computer Programming, 78(7):884– 905, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000883.

Ancona:2013:PSS

[1436] Davide Ancona. Preface to the special section on Object-Oriented Programming and Systems (OOPS 2009), a special track at the 24th ACM Symposium on Applied Computing. Science of Computer Programming, 78(7):906, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231300066X.

Bettini:2013:FDT

[1437] Lorenzo Bettini, Sara Capecchi, and Ferruccio Damiani. On flexible dynamic trait replacement for Java-like languages. Science of Computer Programming, 78(7):907– 932, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312002092.

Saito:2013:MMS

[1438] Chieri Saito and Atsushi Igarashi. Matching MyType to subtyping. Science of Computer Programming, 78 (7):933–952, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313000038.

Lievens:2013:AIS

[1439] David Lievens and William Harrison. Abstraction over implementation structure with symmetrically encapsulated multimethods. Science of Computer Programming, 78(7): 953–968, July 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312002225.

Seinturier:2013:PSS

[1440] Lionel Seinturier and Mark van den Brand. Preface to the special section on software evolution, adaptability, and maintenance. Science of Computer Programming, 78(8):

969–970, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000439.

Goeminne:2013:CIM

[1441] Mathieu Goeminne and Tom Mens.

A comparison of identity merge algorithms for software repositories. Science of Computer Programming, 78(8): 971–986, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002048.

Autili:2013:HAR

[1442] Marco Autili, Paolo Di Benedetto, and Paola Inverardi. A hybrid approach for resource-based comparison of adaptable Java applications. Science of Computer Programming, 78(8): 987–1009, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000147.

Laguna:2013:SMS

[1443] Miguel A. Laguna and Yania Crespo. A systematic mapping study on software product line evolution: From legacy system reengineering to product line refactoring. Science of Computer Programming, 78(8):1010– 1034, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000895.

Mosser:2013:SAS

[1444] Sébastien Mosser and Mireille Blay-Fornarino. “Adore”, a logical metamodel supporting business process evolution. Science of Computer Programming, 78(8):1035–1054, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642312001220.

Laval:2013:OLI

[1445] Jannik Laval, Nicolas Anquetil, Usman Bhatti, and Stéphane Ducasse. oZone: Layer identification in the presence of cyclic dependencies. Science of Computer Programming, 78(8): 1055–1072, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001499.

Navas:2013:RRT

[1446] Juan F. Navas, Jean-Philippe Babau, and Jacques Pulou. Reconciling run-time evolution and resourceconstrained embedded systems through a component-based development framework. Science of Computer Programming, 78(8):1073–1098, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642312001633.

Pirzadeh:2013:SSE

[1447] Heidar Pirzadeh, Sara Shanian, Abdelwahab Hamou-Lhadj, Luay Alawneh, and Arya Shafiee. Stratified sampling of execution traces: Execution phases serving as strata. Science of Computer Programming, 78(8):

1099–1118, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002080.

Santos:2013:SSB

[1448] André Santos and João Saraiva. Special section on the Brazilian Symposium on Programming Languages. Science of Computer Programming, 78(8): 1119–1120, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001669.

Martinez:2013:SFA

[1449] Mónica Martínez and Alberto Pardo.

A shortcut fusion approach to accumulations. Science of Computer Programming, 78(8):1121–1136, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312001670.

Rebelo:2013:OGA

[1450] Henrique Rebêlo, Ricardo Lima, Gary T. Leavens, Márcio Cornélio, Alexandre Mota, and César Oliveira. Optimizing generated aspect-oriented assertion checking code for JML using program transformations: an empirical study. Science of Computer Programming, 78(8):1137–1156, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312001682.

Rubbo:2013:IRG

[1451] Fernando Barden Rubbo, Eduardo Kessler Piveta, and Daltro José Nunes. Inference rules for generic code migration of aspect-oriented programs. Science of Computer Programming, 78(8): 1157–1175, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001694.

Shi:2013:LTS

[1452] Rui Shi and Hongwei Xi. A linear type system for multicore programming in ATS. Science of Computer Programming, 78(8):1176– 1192, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001700.

daMata:2013:APC

[1453] Leonardo Luiz Padovani da Mata, Fernando Magno Quintão Pereira, and Renato Ferreira. Automatic parallelization of canonical loops. Science of Computer Programming, 78(8): 1193–1206, August 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001712.

Bergstra:2013:E

[1454] Jan Bergstra. Editorial. Science of Computer Programming, 78(9): 1207, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310000997.

Berry:2013:ESD

[1455] Daniel M. Berry. The essential similarity and differences between mathematical modeling and programming. Science of Computer Programming, 78(9):1208–1211, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642310000924.

Mraihi:2013:IAI

[1456] Olfa Mraihi, Asma Louhichi, Lamia Labed Jilani, Jules Desharnais, and Ali Mili. Invariant assertions, invariant relations, and invariant functions. Science of Computer Programming, 78(9):1212–1239, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312001086.

Halder:2013:APS

[1457] Raju Halder and Agostino Cortesi.

Abstract program slicing on dependence condition graphs. Science of Computer Programming, 78(9):1240– 1263, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001098.

Yang:2013:PAC

[1458] Fan Yang, Chris Hankin, Flemming Nielson, and Hanne Riis Nielson. Predictive access control for distributed computation. Science of Computer Programming, 78(9):1264– 1277, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001104.

Astefanoaei:2013:WPC

[1459] L. Astefanoaei, F. S. de Boer, M. Dastani, and J.-J. Ch. Meyer. A weakest precondition calculus for BUnity. Science of Computer Programming, 78(9):1278–1293, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231200113X.

Lin:2013:SCT

[1460] Che-Sheng Lin and Gwan-Hwan Hwang. State-cover testing for nondeterministic terminating concurrent programs with an infinite number of synchronization sequences. Science of Computer Programming, 78(9):1294– 1323, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001165.

Bonsangue:2013:AEP

[1461] Marcello Bonsangue, Georgiana Caltais, Eugen-Ioan Goriac, Dorel Lucanu, Jan Rutten, and Alexandra Silva. Automatic equivalence proofs for nondeterministic coalgebras. Science of Computer Programming, 78(9):1324– 1345, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001207.

Santone:2013:ICS

[1462] Antonella Santone, Gigliola Vaglini, and Maria Luisa Villani. Incremental construction of systems: an efficient characterization of the lacking sub-system. Science of Computer Programming, 78(9):1346– 1367, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001438.

Kwiatkowski:2013:RMI

[1463] L. M. Kwiatkowski and C. Verhoef. Recovering management information from source code. Science of Computer Programming, 78(9):1368– 1406, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231200144X.

Hadar:2013:WIL

[1464] Irit Hadar. When intuition and logic clash: the case of the objectoriented paradigm. Science of Computer Programming, 78(9):1407– 1426, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001918.

Albert:2013:HSA

[1465] Elvira Albert, Samir Genaim, and Miguel Gómez-Zamalloa. Heap space analysis for garbage collected languages. Science of Computer Programming, 78(9):1427–1448, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001931.

Sutton:2013:ECC

[1466] Andrew Sutton and Jonathan I. Maletic. Emulating C++0x concepts. Science of Computer Programming, 78(9):1449–1469, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642312001943.

Poroor:2013:MMS

[1467] Jayaraj Poroor and Bharat Jayaraman. Modeling mobile stateful channels in πZ. Science of Computer Programming, 78(9):1470– 1489, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001955.

Eugster:2013:SUP

[1468] Patrick Eugster. Safe uniform proxies for Java. Science of Computer Programming, 78(9):1490–1520, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312002043.

Hesselink:2013:CAP

[1469] Wim H. Hesselink. Complete assertional proof rules for progress under weak and strong fairness. Science of Computer Programming, 78(9):1521– 1537, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002055.

Leger:2013:MFC

[1470] Paul Leger, Éric Tanter, and Rémi Douence. Modular and flexible causality control on the Web. Science of Computer Programming, 78(9):1538– 1558, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231200216X.

Atoofian:2013:AAI

[1471] Ehsan Atoofian, Amirali Baniasadi, and Yvonne Coady. ARV–ALA: Improving performance of software transactional memory through adaptive read and write policies. Science of Computer Programming, 78(9):1559– 1571, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002183.

Mu:2013:MSI

[1472] Kedian Mu, Zhi Jin, Weiru Liu, Didar Zowghi, and Bo Wei. Measuring the significance of inconsistency in the Viewpoints framework. Science of Computer Programming, 78(9):1572– 1599, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002213.

Sinkovics:2013:IMC

[1473] Ábel Sinkovics and Zoltán Porkoláb.

Implementing monads for C++ template metaprograms. Science of Computer Programming, 78(9):1600– 1621, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000051.

Hesselink:2013:MVL

[1474] Wim H. Hesselink. Mechanical verification of Lamport’s Bakery algorithm. Science of Computer Programming, 78(9):1622–1638, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000634.

Ferrari:2013:TPM

[1475] Fabiano Cutigi Ferrari, Awais Rashid, and José Carlos Maldonado. Towards the practical mutation testing of AspectJ programs. Science of Computer Programming, 78(9):1639– 1662, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000713.

Grant:2013:UHE

[1476] Scott Grant, James R. Cordy, and David B. Skillicorn. Using heuristics to estimate an appropriate number of latent topics in source code analysis. Science of Computer Programming, 78(9):1663– 1678, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000762.

Ying:2013:VQP

[1477] Mingsheng Ying, Nengkun Yu, Yuan Feng, and Runyao Duan. Verification of quantum programs. Science of Computer Programming, 78(9):1679– 1700, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000774.

Oh:2013:ABA

[1478] Hakjoo Oh and Kwangkeun Yi.

Access-based abstract memory localization in static analysis. Science of Computer Programming, 78(9):1701– 1727, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001007.

Dam:2013:SCI

[1479] Hoa Khanh Dam and Aditya Ghose.

Supporting change impact analysis for intelligent agent systems. Science of Computer Programming, 78(9):1728– 1750, September 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001068.

Vinju:2013:PSS

[1480] Jurgen J. Vinju. Preface to the special section on Language Descriptions Tools and Applications (LDTA’08 & ’09). Science of Computer Programming, 78 (10):1751, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002377.

Sloane:2013:PEA

[1481] Anthony M. Sloane, Lennart C. L. Kats, and Eelco Visser. A pure embedding of attribute grammars. Science of Computer Programming, 78(10): 1752–1769, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231100205X.

Lammel:2013:PET

[1482] Ralf Lämmel, Simon Thompson, and Markus Kaiser. Programming errors in traversal programs over structured data. Science of Computer Programming, 78(10):1770– 1808, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311002061.

Soderberg:2013:EIF

[1483] Emma Söderberg, Torbjörn Ekman, Görel Hedin, and Eva Magnusson. Extensible intraprocedural flow analysis at the abstract syntax tree level. Science of Computer Programming, 78(10):1809– 1827, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000172.

Scott:2013:GPT

[1484] Elizabeth Scott and Adrian Johnstone. GLL parse-tree generation. Science of Computer Programming, 78(10): 1828–1844, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000627.

Andersen:2013:BAC

[1485] Jacob Andersen, Claus Brabrand, and David Raymond Christiansen. Banana Algebra: Compositional syntactic language extension. Science of Computer Programming, 78(10):1845– 1870, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002110.

Chilowicz:2013:VFT

[1486] Michel Chilowicz, Étienne Duris, and Gilles Roussel. Viewing functions as token sequences to highlight similarities in source code. Science of Computer Programming, 78(10):1871– 1891, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002195.

Lopez-de-Ipina:2013:PSS

[1487] Diego López de Ipiña, Sergio F.

Ochoa, and José Bravo. Preface to the special section on Software Engineering Aspects of Ubiquitous Computing and Ambient Intelligence (UCAmI 2011). Science of Computer Programming, 78(10):1892– 1894, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000464.

Ruiz-Lopez:2013:RRE

[1488] Tomás Ruiz-López, Manuel Noguera, María José Rodríguez, José Luis Garrido, and Lawrence Chung. REUBI: a Requirements Engineering method for ubiquitous systems. Science of Computer Programming, 78(10):1895– 1911, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001645.

Chabridon:2013:BUQ

[1489] Sophie Chabridon, Denis Conan, Zied Abid, and Chantal Taconet. Building ubiquitous QoC-aware applications through model-driven software engineering. Science of Computer Programming, 78(10):1912– 1929, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001475.

Catala:2013:MMD

[1490] Alejandro Catala, Patricia Pons, Javier Jaen, Jose A. Mocholi, and Elena Navarro. A meta-model for dataflow-based rules in smart environments: Evaluating user comprehension and performance. Science of Computer Programming, 78(10):1930– 1950, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001232.

Roover:2013:CLR

[1491] Coen De Roover, Christophe Scholliers, Wouter Amerijckx, Theo D’Hondt, and Wolfgang De Meuter. CrimeSPOT: a language and runtime for developing active wireless sensor network applications. Science of Computer Programming, 78(10):1951– 1970, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001463.

Paz-Lopez:2013:TUA

[1492] A. Paz-Lopez, G. Varela, J. A. Becerra, S. Vazquez-Rodriguez, and R. J. Duro. Towards ubiquity in ambient intelligence: User-guided component mobility in the HI^{3} architecture. Science of Computer Programming, 78(10): 1971–1986, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002067.

Gil:2013:DUA

[1493] Miriam Gil, Estefanía Serral, Pedro Valderas, and Vicente Pelechano. Designing for user attention: a method for supporting unobtrusive routine tasks. Science of Computer Programming, 78(10):1987– 2008, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000476.

Rodriguez-Covili:2013:LDM

[1494] Juan Rodríguez-Covili and Sergio F.

Ochoa. A lightweight and distributed middleware to provide presence awareness in mobile ubiquitous systems. Science of Computer Programming, 78(10):2009– 2025, October 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313000282.

Desharnais:2013:MPC

[1495] Jules Desharnais. Mathematics of Program Construction (selected papers from MPC 2010). Science of Computer Programming, 78(11):2027– 2028, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001347.

Backhouse:2013:ASL

[1496] Roland Backhouse, Wei Chen, and João F. Ferreira. The algorithmics of solitaire-like games. Science of Computer Programming, 78(11):2029– 2046, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001359.

Dongol:2013:DRT

[1497] Brijesh Dongol and Ian J. Hayes. Deriving real-time action systems in a sampling logic. Science of Computer Programming, 78(11):2047– 2063, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001360.

Guttmann:2013:EDA

[1498] Walter Guttmann. Extended designs algebraically. Science of Computer Programming, 78(11):2064– 2085, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001372.

Hayes:2013:LUT

[1499] Ian J. Hayes, Steve E. Dunne, and Larissa A. Meinicke. Linking Unifying Theories of Program Refinement. Science of Computer Programming, 78(11):2086– 2107, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001384.

Hinze:2013:AFU

[1500] Ralf Hinze. Adjoint folds and unfolds — an extended study. Science of Computer Programming, 78(11):2108– 2159, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001396.

Macedo:2013:TLA

[1501] Hugo Daniel Macedo and José Nuno Oliveira. Typing linear algebra: a biproduct-oriented approach. Science of Computer Programming, 78(11): 2160–2191, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001402.

Niqui:2013:SPC

[1502] Milad Niqui and Jan J. M. M. Rutten. Stream processing coalgebraically.

Science of Computer Programming, 78 (11):2192–2215, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312001414.

Wang:2013:RPM

[1503] Meng Wang, Jeremy Gibbons, Kazutaka Matsuda, and Zhenjiang Hu. Refactoring pattern matching. Science of Computer Programming, 78(11): 2216–2242, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001426.

Abascal:2013:SSM

[1504] Julio Abascal and Nestor Garay-Vitoria. Special section on methodological development of interactive systems from Interacción 2011. Science of Computer Programming, 78(11):2243– 2244, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000221.

Condori-Fernandez:2013:EAE

[1505] Nelly Condori-Fernández, Jose Ignacio Panach, Arthur Iwan Baars, Tanja Vos, and Óscar Pastor. An empirical approach for evaluating the usability of model-driven tools. Science of Computer Programming, 78(11):2245– 2258, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001451.

Figueroa-Martinez:2013:EUL

[1506] Jose Figueroa-Martinez, Víctor López-Jaquero, Francisco Luis Gutiérrez Vela, and Pascual González. Enriching UsiXML language to support awareness requirements. Science of Computer Programming, 78(11):2259– 2267, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001906.

Losada:2013:GAD

[1507] Begoña Losada, Maite Urretavizcaya, and Isabel Fernández-Castro. A guide to agile development of interactive software with a “User Objectives”- driven methodology. Science of Computer Programming, 78(11):2268– 2281, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001657.

Rojas:2013:BGB

[1508] Luis A. Rojas and José A. Macías.

Bridging the gap between information architecture analysis and software engineering in interactive Web application development. Science of Computer Programming, 78(11):2282– 2291, November 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001487.

Bosch:2013:SPL

[1509] Jan Bosch and Jaejoon Lee. Software product lines (SPLC 2010). Science of Computer Programming, 78(12):2293– 2294, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000257.

Karatas:2013:EFM

[1510] Ahmet Serkan Karatas, Halit Oguztüzün, and Ali Dogru. From extended feature models to constraint logic programming. Science of Computer Programming, 78(12):2295– 2312, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001153.

Hartmann:2013:UMI

[1511] Herman Hartmann, Mila Keren, Aart Matsinger, Julia Rubin, Tim Trew, and Tali Yatzkar-Haham. Using MDA for integration of heterogeneous components in software supply chains. Science of Computer Programming, 78(12):2313– 2330, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000688.

Ubayashi:2013:CDP

[1512] Naoyasu Ubayashi, Shin Nakajima, and Masayuki Hirayama. Contextdependent product line engineering with lightweight formal approaches. Science of Computer Programming, 78(12):2331–2346, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312001177.

Eklund:2013:AAP

[1513] Ulrik Eklund and Håkan Gustavsson. Architecting automotive product lines: Industrial practice. Science of Computer Programming, 78(12):2347– 2359, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312001190.

Ganesan:2013:AUT

[1514] Dharmalingam Ganesan, Mikael Lindvall, David McComas, Maureen Bartholomew, Steve Slegel, Barbara Medina, Rene Krikhaar, Chris Verhoef, and Lisa P. Montgomery. An analysis of unit tests of a flight software product line. Science of Computer Programming, 78(12):2360– 2380, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000317.

Marinho:2013:MNS

[1515] Fabiana G. Marinho, Rossana M. C.

Andrade, Cláudia Werner, Windson Viana, Marcio E. F. Maia, Lincoln S. Rocha, Eldânae Teixeira, João B. Ferreira Filho, Valéria L. L. Dantas, Fabrício Lima, and Saulo Aguiar. MobiLine: a Nested Software Product Line for the domain of mobile and context-aware applications. Science of Computer Programming, 78(12):2381– 2398, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000871.

Cetina:2013:PDS

[1516] Carlos Cetina, Pau Giner, Joan Fons, and Vicente Pelechano. Prototyping Dynamic Software Product Lines to evaluate run-time reconfigurations. Science of Computer Programming, 78(12):2399– 2413, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001189.

Heuer:2013:DVA

[1517] André Heuer, Vanessa Stricker, Christof J. Budnik, Sascha Konrad, Kim Lauenroth, and Klaus Pohl. Defining variability in activity diagrams and Petri nets. Science of Computer Programming, 78(12):2414– 2432, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001141.

Arbab:2013:FSE

[1518] Farhad Arbab and Marjan Sirjani. Fundamentals of Software Engineering (selected papers of FSEN 2011). Science of Computer Programming, 78(12): 2433–2434, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002049.

Hwong:2013:FAC

[1519] Yi Ling Hwong, Jeroen J. A. Keiren, Vincent J. J. Kusters, Sander Leemans, and Tim A. C. Willemse. Formalising and analysing the control software of the Compact Muon Solenoid Experiment at the Large Hadron Collider. Science of Computer Programming, 78(12):2435– 2452, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002365.

Kim:2013:DLN

[1520] Minyoung Kim, Mark-Oliver Stehr, and Carolyn Talcott. A distributed logic for Networked Cyber-Physical Systems. Science of Computer Programming, 78(12):2453– 2467, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000245.

Aceto:2013:SMS

[1521] Luca Aceto, Ignacio Fábregas, David de Frutos-Escrig, Anna Ingólfsdóttir, and Miguel Palomino. On the specification of modal systems: a comparison of three frameworks. Science of Computer Programming, 78(12):2468– 2487, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000294.

Lambertz:2013:EDA

[1522] Christian Lambertz and Mila Majster-Cederbaum. Efficient deadlock analysis of component-based software architectures. Science of Computer Programming, 78(12):2488– 2510, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313000415.

Hansen:2013:RAC

[1523] Hallstein A. Hansen, Gerardo Schneider, and Martin Steffen. Reachability analysis of complex planar hybrid systems. Science of Computer Programming, 78(12):2511– 2536, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313000427.

Deng:2013:CRW

[1524] Yuxin Deng and Matthew Hennessy. Compositional reasoning for weighted Markov decision processes. Science of Computer Programming, 78(12):2537– 2579, December 1, 2013. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313000440.

Mens:2014:GEI

[1525] Kim Mens, M. G. J. van den Brand, and Holger M. Kienle. Guest Editors’ introduction to the 4th issue of Experimental Software and Toolkits (EST-4). Science of Computer Programming, 79 (??):1–5, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312002201.

Fabry:2014:AEM

[1526] Johan Fabry, Andy Kellens, Simon Denier, and Stéphane Ducasse. AspectMaps: Extending Moose to visualize AOP software. Science of Computer Programming, 79(??):6– 22, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000329.

Farahbod:2014:EFS

[1527] Roozbeh Farahbod, Vincenzo Gervasi, and Uwe Glässer. Executable formal specifications of complex distributed systems with CoreASM. Science of Computer Programming, 79 (??):23–38, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000160.

Bertran:2014:DTS

[1528] Benjamin Bertran, Julien Bruneau, Damien Cassou, Nicolas Loriant, Emilie Balland, and Charles Consel. Dia-Suite: a tool suite to develop Sense/ Compute/Control applications. Science of Computer Programming, 79 (??):39–51, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000652.

vanderBurg:2014:DTD

[1529] Sander van der Burg and Eelco Dolstra. Disnix: a toolset for distributed deployment. Science of Computer Programming, 79(??):52– 69, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000639.

Thum:2014:FEF

[1530] Thomas Thüm, Christian Kästner, Fabian Benduhn, Jens Meinicke, Gunter Saake, and Thomas Leich. FeatureIDE: an extensible framework for feature-oriented software development. Science of Computer Programming, 79 (??):70–85, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001128.

Bergel:2014:ITC

[1531] Alexandre Bergel and Vanessa Peña. Increasing test coverage with Hapao. Science of Computer Programming, 79 (??):86–100, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000706.

Krebs:2014:JJB

[1532] Nico Krebs and Lothar Schmitz. Jaccie: a Java-based compiler-compiler for generating, visualizing and debugging compiler components. Science of Computer Programming, 79(??): 101–115, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000469.

Caserta:2014:JTJ

[1533] Pierre Caserta and Olivier Zendra. JBInsTrace: a tracer of Java and JRE classes at basicblock granularity by dynamically instrumenting bytecode. Science of Computer Programming, 79(??):116– 125, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000299.

Recio-Garcia:2014:CFB

[1534] Juan A. Recio-García, Pedro A.

González-Calero, and Belén Díaz-Agudo. j colibri2: a framework for building case-based reasoning systems. Science of Computer Programming, 79(??):126–145, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642312000664.

Sarimbekov:2014:JCS

[1535] Aibek Sarimbekov, Andreas Sewe, Walter Binder, Philippe Moret, and Mira Mezini. JP2: Call-site aware calling context profiling for the Java Virtual Machine. Science of Computer Programming, 79(??):146– 157, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002036.

Cordy:2014:TRT

[1536] James R. Cordy and Chanchal K.

Roy. Tuning research tools for scalability and performance: the NiCad experience. Science of Computer Programming, 79(??):158– 171, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642311002024.

Kuhrmann:2014:PET

[1537] Marco Kuhrmann, Georg Kalus, and Manuel Then. The Process Enactment Tool Framework — transformation of software process models to prepare enactment. Science of Computer Programming, 79(??): 172–188, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000640.

Zschaler:2014:SJF

[1538] Steffen Zschaler, Birgit Demuth, and Lothar Schmitz. Salespoint: a Java framework for teaching object-oriented software development. Science of Computer Programming, 79(??):189– 203, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231200069X.

Lungu:2014:ECS

[1539] Mircea Lungu, Michele Lanza, and Oscar Nierstrasz. Evolutionary and collaborative software architecture recovery with Softwarenaut. Science of Computer Programming, 79(??): 204–223, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312000718.

Reniers:2014:STS

[1540] Dennie Reniers, Lucian Voinea, Ozan Ersoy, and Alexandru Telea. The Solid* toolset for software visual analytics of program structure and metrics comprehension: From research prototype to product. Science of Computer Programming, 79(??):224– 240, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231200086X.

Bajracharya:2014:SIL

[1541] Sushil Bajracharya, Joel Ossher, and Cristina Lopes. Sourcerer: an infrastructure for large-scale collection and analysis of open-source code. Science of Computer Programming, 79(??): 241–259, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231200072X.

Smet:2014:TVA

[1542] Benoît De Smet, Lorent Lempereur, Zohreh Sharafi, Yann-Gaël Guéhéneuc, Giuliano Antoniol, and Naji Habra. Taupe: Visualizing and analyzing eye-tracking data. Science of Computer Programming, 79(??):260– 278, January 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000135.

Mousavi:2014:PSS

[1543] Mohammad Reza Mousavi and Gwen Salaün. Preface: Special section on foundations of coordination languages and software architectures (selected papers from FOCLASA’10). Science of Computer Programming, 80 (??):1–2, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000482.

Moon:2014:CMR

[1544] Young-Joo Moon, Alexandra Silva, Christian Krause, and Farhad Arbab. A compositional model to reason about end-to-end QoS in Stochastic Reo connectors. Science of Computer Programming, 80(??):3– 24, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642311002073.

Kamali:2014:FDW

[1545] Maryam Kamali, Linas Laibinis, Luigia Petre, and Kaisa Sere. Formal development of wireless sensor-actor networks. Science of Computer Programming, 80 (??):25–49, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312000470.

Aldrich:2014:BSP

[1546] Jonathan Aldrich and Ricardo Massa Ferreira Lima. Brazilian Symposium on Programming Languages (SBLP 2010). Science of Computer Programming, 80 (??):51, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000749.

Scholliers:2014:PAM

[1547] Christophe Scholliers, Éric Tanter, and Wolfgang De Meuter. Parallel actor monitors: Disentangling tasklevel parallelism from data partitioning in the actor model. Science of Computer Programming, 80(??): 52–64, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000725.

Pinho:2014:OOP

[1548] Eduardo Gurgel Pinho and Francisco Heron de Carvalho Junior. An object-oriented parallel programming language for distributed-memory parallel computing platforms. Science of Computer Programming, 80(??): 65–90, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000750.

Rimsa:2014:ESC

[1549] Andrei Rimsa, Marcelo d’Amorim, Fernando Magno Quintão Pereira, and Roberto S. Bigonha. Efficient static checker for tainted variable attacks. Science of Computer Programming, 80 (??):91–105, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000737.

Salaun:2014:PSS

[1550] Gwen Salaün and Bernhard Schätz. Preface: Special section on formal methods for industrial critical systems (selected papers from FMICS’11). Science of Computer Programming, 80(??):107–108, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764231300021X.

Reinbacher:2014:RVM

[1551] Thomas Reinbacher, Jörg Brauer, Martin Horauer, Andreas Steininger, and Stefan Kowalewski. Runtime verification of microcontroller binary code. Science of Computer Programming, 80(??):109–129, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642312002109.

Lantreibecq:2014:FAH

[1552] Etienne Lantreibecq and Wendelin Serwe. Formal analysis of a hardware dynamic task dispatcher with CADP. Science of Computer Programming, 80(??):130–149, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313000166.

Huhn:2014:OFS

[1553] Michaela Huhn and Stefan Milius. Observations on formal safety analysis in practice. Science of Computer Programming, 80(??):150– 168, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300004X.

Catano:2014:CSL

[1554] Néstor Cataño, Ijaz Ahmed, Radu I.

Siminiceanu, and Jonathan Aldrich. A case study on the lightweight verification of a multi-threaded task server. Science of Computer Programming, 80(??):169–187, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313000178.

Sijtema:2014:EFE

[1555] M. Sijtema, A. Belinfante, M. I. A.

Stoelinga, and L. Marinelli. Experiences with formal engineering: Model-based specification, implementation and testing of a software bus at Neopost. Science of Computer Programming, 80(??):188– 209, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300107X.

Nicholson:2014:AVD

[1556] Jon Nicholson, Amnon H. Eden, Epameinondas Gasparis, and Rick Kazman. Automated verification of design patterns: a case study. Science of Computer Programming, 80(??): 211–222, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001354.

Gerakios:2014:SSG

[1557] Prodromos Gerakios, Nikolaos Papaspyrou, and Konstantinos Sagonas. Static safety guarantees for a low-level multithreaded language with regions. Science of Computer Programming, 80(??):223– 263, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001433.

Konur:2014:SSC

[1558] Savas Konur. Specifying safety-critical systems with a decidable duration logic. Science of Computer Programming, 80(??):264–287, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764231300186X.

Calder:2014:RTV

[1559] Muffy Calder, Alexandros Koliousis, Michele Sevegnani, and Joseph Sventek. Real-time verification of wireless home networks using bigraphs with sharing. Science of Computer Programming, 80(??):288– 310, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001974.

Tanter:2014:ELA

[1560] Éric Tanter, Ismael Figueroa, and Nicolas Tabareau. Execution levels for aspect-oriented programming: Design, semantics, implementations and applications. Science of Computer Programming, 80(??):311– 342, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002244.

El-Fakih:2014:PAT

[1561] Khaled El-Fakih, Nina Yevtushenko, and Adenilso Simao. A practical approach for testing timed deterministic finite state machines with single clock. Science of Computer Programming, 80(??):343– 355, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002372.

Mirandola:2014:ASE

[1562] Raffaela Mirandola, Pasqualina Potena, and Patrizia Scandurra. Adaptation space exploration for serviceoriented applications. Science of Computer Programming, 80(??):356– 384, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002475.

Alpuente:2014:UCT

[1563] María Alpuente, Demis Ballis, Francisco Frechina, and Daniel Romero. Using conditional trace slicing for improving Maude programs. Science of Computer Programming, 80(??): 385–415, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002505.

Classen:2014:FSM

[1564] Andreas Classen, Maxime Cordy, Patrick Heymans, Axel Legay, and Pierre-Yves Schobbens. Formal semantics, modular specification, and symbolic verification of product-line behaviour. Science of Computer Programming, 80(??):416–439, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313002517.

Ketterlin:2014:RMA

[1565] Alain Ketterlin and Philippe Clauss. Recovering memory access patterns of executable programs. Science of Computer Programming, 80(??):440– 456, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001505.

Thomas:2014:SSE

[1566] Stephen W. Thomas, Bram Adams, Ahmed E. Hassan, and Dorothea Blostein. Studying software evolution using topic models. Science of Computer Programming, 80(??):457– 479, February 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642312001621.

Bendisposto:2014:PAV

[1567] Jens Bendisposto, Michael Leuschel, and Markus Roggenbach. Preface of Automated Verification of Critical Systems 2010 (AVoCS 2010). Science of Computer Programming, 81 (??):1–2, February 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001986.

Mazur:2014:CBC

[1568] Tomasz Mazur and Gavin Lowe. CSPbased counter abstraction for systems with node identifiers. Science of Computer Programming, 81(??): 3–52, February 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000798.

Macedonio:2014:SAK

[1569] Damiano Macedonio and Massimo Merro. A semantic analysis of key management protocols for wireless sensor networks. Science of Computer Programming, 81(??):53– 78, February 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300018X.

Alpuente:2014:RLA

[1570] María Alpuente, Demis Ballis, and Daniel Romero. A rewriting logic approach to the formal specification and verification of Web applications. Science of Computer Programming, 81(??):79–107, February 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313001883.

Hoang:2014:RAA

[1571] Thai Son Hoang. Reasoning about almost-certain convergence properties using Event-B. Science of Computer Programming, 81(??):108– 121, February 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001998.

Jones:2014:SIA

[1572] Cliff Jones and Alexander Romanovsky. Special issue on Automated Verification of Critical Systems (AVoCS’11). Science of Computer Programming, 82(??):1, March 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313001391.

Hallerstede:2014:MTT

[1573] Stefan Hallerstede, Michael Jastram, and Lukas Ladenberger. A method and tool for tracing requirements into specifications. Science of Computer Programming, 82(??): 2–21, March 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313000683.

Auer:2014:CCG

[1574] Anthony Auer, Juergen Dingel, and Karen Rudie. Concurrency control generation for dynamic threads using discrete-event systems. Science of Computer Programming, 82(??): 22–43, March 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000208.

Gan:2014:SMC

[1575] Xiang Gan, Jori Dubrovin, and Keijo Heljanko. A symbolic model checking approach to verifying satellite onboard software. Science of Computer Programming, 82(??): 44–55, March 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000658.

Qin:2014:ARP

[1576] Shengchao Qin, Guanhua He, Chenguang Luo, Wei-Ngan Chin, and Hongli Yang. Automatically refining partial specifications for heapmanipulating programs. Science of Computer Programming, 82(??): 56–76, March 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000646.

Philippaerts:2014:SVV

[1577] Pieter Philippaerts, Jan Tobias Mühlberg, Willem Penninckx, Jan Smans, Bart Jacobs, and Frank Piessens. Software verification with VeriFast: Industrial case studies. Science of Computer Programming, 82(??):77– 97, March 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000191.

Arbab:2014:P

[1578] Farhad Arbab and Peter Csaba Ölveczky. Preface. Science of Computer Programming, 83(??):1–2, April 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313002864.

Meseguer:2014:TDS

[1579] José Meseguer. Taming distributed system complexity through formal patterns. Science of Computer Programming, 83(??):3–34, April 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313001615.

Sabouri:2014:RVC

[1580] Hamideh Sabouri and Ramtin Khosravi. Reducing the verification cost of evolving product families using static analysis techniques. Science of Computer Programming, 83(??): 35–55, April 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001585.

Aminof:2014:SHS

[1581] Benjamin Aminof, Fabio Mogavero, and Aniello Murano. Synthesis of hierarchical systems. Science of Computer Programming, 83(??): 56–79, April 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231300155X.

Ramli:2014:LX

[1582] Carroline Dewi Puspa Kencana Ramli, Hanne Riis Nielson, and Flemming Nielson. The logic of XACML. Science of Computer Programming, 83 (??):80–105, April 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001238.

Bauer:2014:MST

[1583] Sebastian S. Bauer, Kim G. Larsen, Axel Legay, Ulrik Nyman, and Andrzej Wasowski. A modal specification theory for components with data. Science of Computer Programming, 83 (??):106–128, April 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300141X.

King:2014:PSS

[1584] Andy King. Preface for SCP special issue on Principles and Practice of Declarative Programming. Science of Computer Programming, 84 (??):1–2, May 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001949.

vanEmden:2014:MC

[1585] M. H. van Emden. Matrix code. Science of Computer Programming, 84 (??):3–21, May 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001251.

Jaksic:2014:EHC

[1586] Svetlana Jaksić and Luca Padovani. Exception handling for copyless messaging. Science of Computer Programming, 84(??):22–51, May 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313001214.

Thielecke:2014:SPA

[1587] Hayo Thielecke. On the semantics of parsing actions. Science of Computer Programming, 84(??): 52–76, May 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001202.

Lago:2014:LDT

[1588] Ugo Dal Lago and Barbara Petit. Linear dependent types in a call-by-value scenario. Science of Computer Programming, 84(??): 77–100, May 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001676.

Schrijvers:2014:TMS

[1589] Tom Schrijvers, Bart Demoen, Markus Triska, and Benoit Desouter. Tor: Modular search with hookable disjunction. Science of Computer Programming, 84(??):101–120, May 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764231300138X.

VanGorp:2014:GEI

[1590] Pieter Van Gorp and Louis Rose. Guest Editors’ introduction to the first issue on Experimental Software Engineering in the Cloud (ESEiC). Science of Computer Programming, 85 ((part A)):1–4, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003523.

Kolahdouz-Rahimi:2014:EMT

[1591] S. Kolahdouz-Rahimi, K. Lano, S. Pillay, J. Troya, and P. Van Gorp. Evaluation of model transformation approaches for model refactoring. Science of Computer Programming, 85 ((part A)):5–40, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001871.

Jakumeit:2014:SCT

[1592] Edgar Jakumeit, Sebastian Buchwald, Dennis Wagelaar, Li Dan, Ábel Hegedüs, Markus Herrmannsdörfer, Tassilo Horn, Elina Kalnina, Christian Krause, Kevin Lano, Markus Lepper, Arend Rensink, Louis Rose, Sebastian Wätzoldt, and Steffen Mazanek. A survey and comparison of transformation tools based on the transformation tool contest. Science of Computer Programming, 85((part A)): 41–99, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002803.

Gibbons:2014:SPM

[1593] Jeremy Gibbons and Pablo Nogueira. Selected papers from Mathematics of Program Construction 2012. Science of Computer Programming, 85 ((part B)):101, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002323.

Hoare:2014:LPU

[1594] Tony Hoare and Stephan van Staden. The laws of programming unify process calculi. Science of Computer Programming, 85((part B)): 102–114, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002207.

Morgan:2014:ONN

[1595] Carroll Morgan. An old new notation for elementary probability theory. Science of Computer Programming, 85((part B)): 115–136, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002256.

Dongol:2014:DRT

[1596] Brijesh Dongol, Ian J. Hayes, and John Derrick. Deriving real-time action systems with multiple time bands using algebraic reasoning. Science of Computer Programming, 85((part B)):137–165, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002025.

Backhouse:2014:FPP

[1597] Roland Backhouse. First-past-the-post games. Science of Computer Programming, 85((part B)): 166–203, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001640.

Dang:2014:CLR

[1598] H.-H. Dang and B. Möller. Concurrency and local reasoning under reverse exchange. Science of Computer Programming, 85((part B)): 204–223, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001639.

Guttmann:2014:ACS

[1599] Walter Guttmann. Algebras for correctness of sequential computations. Science of Computer Programming, 85((part B)):224–240, June 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313002013.

Lopez-Jaquero:2014:PSI

[1600] Víctor López-Jaquero and Jean Vanderdonckt. Preface to the special issue on Software Support for User Interface Description Languages (UIDL 2011). Science of Computer Programming, 86 (??):1–2, June 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002426.

Barboni:2014:BGB

[1601] Eric Barboni, Célia Martinie, David Navarre, Philippe Palanque, and Marco Winckler. Bridging the gap between a behavioural formal description technique and a user interface description language: Enhancing ICO with a graphical user interface markup language. Science of Computer Programming, 86(??): 3–29, June 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000993.

Dumas:2014:GES

[1602] Bruno Dumas, Beat Signer, and Denis Lalanne. A graphical editor for the SMUIML multimodal user interaction description language. Science of Computer Programming, 86 (??):30–42, June 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001019.

Gonzalez-Calleros:2014:AUE

[1603] Juan Gonzalez-Calleros, Jan-Patrick Osterloh, Rene Feil, and Andreas Lüdtke. Automated UI evaluation based on a cognitive architecture and UsiXML. Science of Computer Programming, 86(??): 43–57, June 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001020.

Minon:2014:AIA

[1604] Raúl Miñón, Lourdes Moreno, Paloma Martínez, and Julio Abascal. An approach to the integration of accessibility requirements into a user interface development method. Science of Computer Programming, 86 (??):58–73, June 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001032.

Panach:2014:PMU

[1605] Jose Ignacio Panach, Nathalie Aquino, and Óscar Pastor. A proposal for modelling usability in a holistic MDD method. Science of Computer Programming, 86(??):74– 88, June 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001573.

Guerrero-Garcia:2014:EDU

[1606] Josefina Guerrero-García. Evolutionary design of user interfaces for workflow information systems. Science of Computer Programming, 86(??): 89–102, June 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001603.

Brabrand:2014:PSS

[1607] Claus Brabrand and Eric Van Wyk.

Preface to the Special Section on Language Descriptions, Tools, and Applications (LDTA 2011). Science of Computer Programming, 87 (??):1, July 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002384.

Sauthoff:2014:YGA

[1608] Georg Sauthoff and Robert Giegerich. Yield grammar analysis and product optimization in a domain-specific language for dynamic programming. Science of Computer Programming, 87 (??):2–22, July 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002402.

Johnstone:2014:MGS

[1609] Adrian Johnstone, Elizabeth Scott, and Mark van den Brand. Modular grammar specification. Science of Computer Programming, 87 (??):23–43, July 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002414.

vanDooren:2014:MTC

[1610] Marko van Dooren, Bart Jacobs, and Wouter Joosen. Modular type checking of anchored exception declarations. Science of Computer Programming, 87 (??):44–61, July 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002785.

Joncheere:2014:IWM

[1611] Niels Joncheere, Sebastian Günther, Ragnhild Van Der Straeten, and Viviane Jonckers. Improving workflow modularity using a concernspecific layer on top of Unify. Science of Computer Programming, 87 (??):62–94, July 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002876.

Sierra:2014:PSI

[1612] José-Luis Sierra and Antonio Sarasa-Cabezuelo. Preface for the special issue on Software Development Concerns in the e-Learning Domain. Science of Computer Programming, 88 (??):1–2, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001361.

Martin-SanJose:2014:FLI

[1613] Juan-Fernando Martín-SanJosé, M.- Carmen Juan, José-A. Gil-Gómez, and Noemí Rando. Flexible learning itinerary vs. linear learning itinerary. Science of Computer Programming, 88 (??):3–21, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000033.

Fouh:2014:DAI

[1614] Eric Fouh, Ville Karavirta, Daniel A.

Breakiron, Sally Hamouda, Simin Hall, Thomas L. Naps, and Clifford A. Shaffer. Design and architecture of an interactive eTextbook — the OpenDSA system. Science of Computer Programming, 88(??): 22–40, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300333X.

Conde:2014:IDS

[1615] Miguel Á. Conde, Francisco García-Peñalvo, Marc Alier, Enric Mayol, and Camino Fernández-Llamas. Implementation and design of a service-based framework to integrate personal and institutional learning environments. Science of Computer Programming, 88 (??):41–53, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002840.

Molina:2014:EGN

[1616] Ana I. Molina, Miguel A. Redondo, Manuel Ortega, and Carmen Lacave. Evaluating a graphical notation for modeling collaborative learning activities: a family of experiments. Science of Computer Programming, 88 (??):54–81, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000860.

Cooper:2014:OSS

[1617] Matthew L. Cooper, Clifford A. Shaffer, Stephen H. Edwards, and Sean P. Ponce. Open source software and the algorithm visualization community. Science of Computer Programming, 88 (??):82–91, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000021.

Santos:2014:EWB

[1618] Olga C. Santos, Jesus G. Boticario, and Diana Pérez-Marín. Extending Web-based educational systems with personalised support through User Centred Designed recommendations along the e-learning life cycle. Science of Computer Programming, 88 (??):92–109, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003407.

Heeren:2014:FSS

[1619] Bastiaan Heeren and Johan Jeuring.

Feedback services for stepwise exercises. Science of Computer Programming, 88(??):110–129, August 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314000884.

Mousavi:2014:F

[1620] Mohammad Reza Mousavi and Antonio Ravara. Foreword. Science of Computer Programming, 89 (part A) (??):1, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003328.

Horne:2014:VAR

[1621] Ross Horne and Vladimiro Sassone.

A verified algebra for read-write Linked Data. Science of Computer Programming, 89 (part A)(??):

2–22, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001627.

Ware:2014:ATC

[1622] Simon Ware and Robi Malik. An algorithm to test the conflict preorder. Science of Computer Programming, 89 (part A)(??):23–40, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002359.

Reynisson:2014:MSA

[1623] Arni Hermann Reynisson, Marjan Sirjani, Luca Aceto, Matteo Cimini, Ali Jafari, Anna Ingolfsdottir, and Steinar Hugi Sigurdarson. Modelling and simulation of asynchronous real-time systems using Timed Rebeca. Science of Computer Programming, 89 (part A)(??): 41–68, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000239.

DiRuscio:2014:GES

[1624] Davide Di Ruscio, Richard F. Paige, and Alfonso Pierantonio. Guest editorial to the special issue on Success Stories in Model Driven Engineering. Science of Computer Programming, 89 (part B)(??):69– 70, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003420.

Brambilla:2014:LSM

[1625] Marco Brambilla and Piero Fraternali. Large-scale Model-Driven Engineering of Web user interaction: the WebML and WebRatio experience. Science of Computer Programming, 89 (part B)(??):71– 87, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000701.

Davies:2014:MDE

[1626] Jim Davies, Jeremy Gibbons, James Welch, and Edward Crichton. Modeldriven engineering of information systems: 10 years and 1000 versions. Science of Computer Programming, 89 (part B)(??):88–104, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000270.

Nadas:2014:MIA

[1627] Andras Nadas, Tihamer Levendovszky, Ethan K. Jackson, Istvan Madari, and Janos Sztipanovits. A modelintegrated authoring environment for privacy policies. Science of Computer Programming, 89 (part B)(??): 105–125, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300124X.

Davies:2014:CEM

[1628] Jim Davies, Jeremy Gibbons, Steve Harris, and Charles Crichton. The CancerGrid experience: Metadatabased model-driven engineering for clinical trials. Science of Computer Programming, 89 (part B)(??):126– 143, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000452.

Hutchinson:2014:MDE

[1629] John Hutchinson, Jon Whittle, and Mark Rouncefield. Model-driven engineering practices in industry: Social, organizational and managerial factors that lead to success or failure. Science of Computer Programming, 89 (part B)(??):144– 161, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000786.

Buttner:2014:MDS

[1630] Fabian Büttner, Ullrich Bartels, Lars Hamann, Oliver Hofrichter, Mirco Kuhlmann, Martin Gogolla, Lutz Rabe, Frank Steimke, Yorck Rabenstein, and Alina Stosiek. Model-driven standardization of public authority data interchange. Science of Computer Programming, 89 (part B)(??): 162–175, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000695.

Cuadrado:2014:AMD

[1631] Jesús Sánchez Cuadrado, Javier Luis Cánovas Izquierdo, and Jesús García Molina. Applying model-driven engineering in small software enterprises. Science of Computer Programming, 89 (part B)(??):176– 198, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001056.

Dubois:2014:MBF

[1632] Emmanuel Dubois, Christophe Bortolaso, Damien Appert, and Guillaume Gauffre. An MDE-based framework to support the development of Mixed Interactive Systems. Science of Computer Programming, 89 (part B)(??): 199–221, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000671.

Braude:2014:WPC

[1633] Eric J. Braude. Weakest preconditions and cumulative subgoal fulfillment. Science of Computer Programming, 89 (part C)(??):223– 234, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000252.

Mascarenhas:2014:RBC

[1634] Fabio Mascarenhas, Sérgio Medeiros, and Roberto Ierusalimschy. On the relation between context-free grammars and parsing expression grammars. Science of Computer Programming, 89 (part C)(??):235– 250, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000276.

Vasilescu:2014:HHS

[1635] Bogdan Vasilescu, Alexander Serebrenik, Tom Mens, Mark G. J. van den Brand, and Ekaterina Pek. How healthy are software engineering conferences? Science of Computer Programming, 89 (part C)(??):251– 272, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000318.

Li:2014:EAJ

[1636] Siliang Li and Gang Tan. Exception analysis in the Java Native Interface. Science of Computer Programming, 89 (part C)(??):273– 297, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000446.

Abdeen:2014:PBV

[1637] Hani Abdeen, Stéphane Ducasse, Damien Pollet, Ilham Alloui, and Jean-Rémy Falleri. The Package Blueprint: Visually analyzing and quantifying packages dependencies. Science of Computer Programming, 89 (part C)(??):298– 319, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000835.

Molina:2014:MDD

[1638] Ana I. Molina, William J. Giraldo, Manuel Ortega, Miguel A. Redondo, and César A. Collazos. Model-driven development of interactive groupware systems: Integration into the software development process. Science of Computer Programming, 89 (part C)(??): 320–349, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400104X.

Huang:2014:LBH

[1639] Fuqun Huang, Bin Liu, You Song, and Shreya Keyal. The links between human error diversity and software diversity: Implications for fault diversity seeking. Science of Computer Programming, 89 (part C)(??): 350–373, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001312.

Coscia:2014:RCF

[1640] José Luis Ordiales Coscia, Cristian Mateos, Marco Crasso, and Alejandro Zunino. Refactoring code-first Web Services for early avoiding WSDL anti-patterns: Approach and comprehensive assessment. Science of Computer Programming, 89 (part C)(??): 374–407, September 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001646.

Gil:2014:PSI

[1641] Amparo Gil, Jean-Michel Muller, and Javier Segura. Preface to the special issue on Numerical Software: Design, Analysis and Verification. Science of Computer Programming, 90 (part A) (??):1, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314001063.

Backeljauw:2014:VES

[1642] Franky Backeljauw, Stefan Becuwe, Annie Cuyt, Joris Van Deun, and Daniel W. Lozier. Validated evaluation of special mathematical functions. Science of Computer Programming, 90 (part A)(??):2–20, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001263.

Bruguera:2014:ORI

[1643] Javier D. Bruguera. Optimizing the representation of intervals. Science of Computer Programming, 90 (part A)(??):21–33, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001408.

Hannukainen:2014:NRF

[1644] Antti Hannukainen, Sergey Korotov, and Michal Krízek. On numerical regularity of the face-to-face longestedge bisection algorithm for tetrahedral partitions. Science of Computer Programming, 90 (part A)(??):

34–41, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001226.

Gil:2014:RSD

[1645] Amparo Gil, Javier Segura, and Nico M. Temme. Recent software developments for special functions in the Santander–Amsterdam project. Science of Computer Programming, 90 (part A)(??):42–54, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002888.

Moldaschl:2014:CES

[1646] Michael Moldaschl and Wilfried N.

Gansterer. Comparison of eigensolvers for symmetric band matrices. Science of Computer Programming, 90 (part A)(??):55–66, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000112.

Buhnova:2014:GES

[1647] Barbora Buhnova, Antonio Vallecillo, Nenad Medvidovic, Magnus Larsson, Javier López, and Jorge Cuellar. Guest editorial to the Special Issue on Component-Based Software Engineering and Software Architecture. Science of Computer Programming, 90 (part B)(??):67– 70, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003213.

Brosig:2014:ALS

[1648] Fabian Brosig, Nikolaus Huber, and Samuel Kounev. Architecturelevel software performance abstractions for online performance prediction. Science of Computer Programming, 90 (part B)(??):71–92, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001421.

Abate:2014:LFC

[1649] Pietro Abate, Roberto Di Cosmo, Ralf Treinen, and Stefano Zacchiroli. Learning from the future of component repositories. Science of Computer Programming, 90 (part B)(??): 93–115, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001561.

Krebs:2014:MTQ

[1650] Rouven Krebs, Christof Momm, and Samuel Kounev. Metrics and techniques for quantifying performance isolation in cloud environments. Science of Computer Programming, 90 (part B)(??):116–134, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001962.

Haitzer:2014:SAA

[1651] Thomas Haitzer and Uwe Zdun. Semiautomated architectural abstraction specifications for supporting software evolution. Science of Computer Programming, 90 (part B)(??):135– 160, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002542.

Rupanov:2014:EEM

[1652] V. Rupanov, C. Buckl, L. Fiege, M. Armbruster, A. Knoll, and G. Spiegelberg. Employing early model-based safety evaluation to iteratively derive E/E architecture design. Science of Computer Programming, 90 (part B)(??):161–179, September 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002554.

Pasareanu:2014:SIF

[1653] Corina S. Pasareanu and Gwen Salaün.

Special issue on Formal Aspects of Component Software (Selected Papers from FACS’12). Science of Computer Programming, 91 (part A) (??):1–2, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000306.

Bae:2014:FPM

[1654] Kyungmin Bae, José Meseguer, and Peter Csaba Ölveczky. Formal patterns for multirate distributed realtime systems. Science of Computer Programming, 91 (part A)(??): 3–44, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002396.

Beohar:2014:ADD

[1655] H. Beohar and P. J. L. Cuijpers. Avoiding diamonds in desynchronisation. Science of Computer Programming, 91 (part A)(??):45– 69, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003377.

Bauer:2014:MTC

[1656] Sebastian Bauer, Rolf Hennicker, and Axel Legay. A meta-theory for component interfaces with contracts on ports. Science of Computer Programming, 91 (part A)(??): 70–89, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002438.

Jansen:2014:SCG

[1657] Nils Jansen, Ralf Wimmer, Erika Ábrahám, Barna Zajzon, Joost-Pieter Katoen, Bernd Becker, and Johann Schuster. Symbolic counterexample generation for large discrete-time Markov chains. Science of Computer Programming, 91 (part A)(??): 90–114, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000331.

Chilton:2014:CAG

[1658] Chris Chilton, Bengt Jonsson, and Marta Kwiatkowska. Compositional assume-guarantee reasoning for input/output component theories. Science of Computer Programming, 91 (part A)(??):115– 137, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000057.

Barbosa:2014:SCO

[1659] Luís Soares Barbosa and Siraj Ahmed Shaikh. Selected contributions from the Open Source Software Certification (OpenCert) workshops. Science of Computer Programming, 91 (part B)(??):139–140, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314001713.

Olesen:2014:CTS

[1660] Mads Chr. Olesen, René Rydhof Hansen, Julia L. Lawall, and Nicolas Palix. Coccinelle: Tool support for automated CERT C Secure Coding Standard certification. Science of Computer Programming, 91 (part B)(??): 141–160, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002031.

Breuer:2014:OSV

[1661] Peter T. Breuer and Simon Pickin. Open source verification in an anonymous volunteer network. Science of Computer Programming, 91 (part B)(??):161–187, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313002037.

Feuser:2014:DOP

[1662] Johannes Feuser and Jan Peleska. Dependability in open proof software with hardware virtualization

— The railway control systems perspective. Science of Computer Programming, 91 (part B)(??):188– 215, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002001.

Almeida:2014:COS

[1663] José Bacelar Almeida, Manuel Barbosa, Jean-Christophe Filliâtre, Jorge Sousa Pinto, and Bárbara Vieira. CAOVerif: an open-source deductive verification platform for cryptographic software implementations. Science of Computer Programming, 91 (part B)(??): 216–233, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231200189X.

Rubanov:2014:EPL

[1664] Vladimir Rubanov and Denis Silakov. Ensuring portability of Linux applications through standardization and knowledge base driven analysis. Science of Computer Programming, 91 (part B)(??):234– 248, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000240.

Sowe:2014:ESF

[1665] Sulayman K. Sowe, Antonio Cerone, and Dimitrios Settas. An empirical study of FOSS developers patterns of contribution: Challenges for data linkage and analysis. Science of Computer Programming, 91 (part B)(??): 249–265, October 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003262.

Huisman:2014:SSI

[1666] Marieke Huisman. SCP special issue on Bytecode 2012 — preface. Science of Computer Programming, 92 (part A)(??):1, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000355.

Albert:2014:CTL

[1667] Elvira Albert, Puri Arenas, Samir Genaim, Germán Puebla, and Guillermo Román-Díez. Conditional termination of loops over heap-allocated data. Science of Computer Programming, 92 (part A)(??):2–24, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313001044.

Wognsen:2014:FAD

[1668] Erik Ramsgaard Wognsen, Henrik Søndberg Karlsen, Mads Chr. Olesen, and René Rydhof Hansen. Formalisation and analysis of Dalvik bytecode. Science of Computer Programming, 92 (part A)(??):25– 55, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003304.

Braberman:2014:SBI

[1669] Víctor Braberman, Diego Garbervetsky, Samuel Hym, and Sergio Yovine. Summary-based inference of quantitative bounds of live heap objects. Science of Computer Programming, 92 (part A)(??):56–84, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313003298.

Simao:2014:SPB

[1670] Adenilso Simao and Carroll Morgan. Selected papers from the Brazilian Symposium on Formal Methods (SBMF 2011). Science of Computer Programming, 92 (part B)(??): 85, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000458.

Braga:2014:CMT

[1671] Christiano Braga, Cássio Santos, and Viviane Torres da Silva. Consistency of model transformation contracts. Science of Computer Programming, 92 (part B)(??):86– 104, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002219.

Dias:2014:CCF

[1672] Diego Dias and Juliano Iyoda. Compositionality and correctness of fault tolerant patterns in HOL4. Science of Computer Programming, 92 (part B)(??):105–128, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313001664.

Welsch:2014:FAT

[1673] Yannick Welsch and Arnd Poetzsch-Heffter. A fully abstract tracebased semantics for reasoning about backward compatibility of class libraries. Science of Computer Programming, 92 (part B)(??):129– 161, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002529.

Buttner:2014:OBI

[1674] Fabian Büttner and Martin Gogolla.

On OCL-based imperative languages. Science of Computer Programming, 92 (part B)(??):162–178, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313002530.

Preoteasa:2014:RAD

[1675] Viorel Preoteasa. Refinement algebra with dual operator. Science of Computer Programming, 92 (part B)(??): 179–210, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001597.

Rocha:2014:SSR

[1676] Camilo Rocha and César Muñoz. Synchronous set relations in rewriting logic. Science of Computer Programming, 92 (part B)(??):211– 228, October 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001652.

Braga:2014:BSP

[1677] Christiano Braga and José Luiz Fiadeiro. Brazilian Symposium on Programming Languages (SBLP 2011). Science of Computer Programming, 93 (part A)(??):1–2, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314000045.

Medeiros:2014:RPE

[1678] Sérgio Medeiros, Fabio Mascarenhas, and Roberto Ierusalimschy. From regexes to parsing expression grammars. Science of Computer Programming, 93 (part A)(??): 3–18, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002171.

Kfoury:2014:SSD

[1679] Assaf Kfoury. The syntax and semantics of a domain-specific language for flow-network design. Science of Computer Programming, 93 (part A)(??): 19–38, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000026.

Mongiovi:2014:MRS

[1680] Melina Mongiovi, Rohit Gheyi, Gustavo Soares, Leopoldo Teixeira, and Paulo Borba. Making refactoring safer through impact analysis. Science of Computer Programming, 93 (part A) (??):39–64, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002827.

Ferreira:2014:UFO

[1681] Gabriel Coutinho Sousa Ferreira, Felipe Nunes Gaia, Eduardo Figueiredo, and Marcelo de Almeida Maia. On the use of feature-oriented programming for evolving software product lines — a comparative study. Science of Computer Programming, 93 (part A)(??): 65–85, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002815.

Grov:2014:PIG

[1682] Gudmund Grov and Thomas Wies.

Preface — Invariant generation. Science of Computer Programming, 93 (part B)(??):87–88, November 1, 2014.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642314002500.

Cachera:2014:IPI

[1683] David Cachera, Thomas Jensen, Arnaud Jobin, and Florent Kirchner. Inference of polynomial invariants for imperative programs: a farewell to Gröbner bases. Science of Computer Programming, 93 (part B)(??): 89–109, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000951.

Sridhar:2014:HBI

[1684] Meera Sridhar, Richard Wartell, and Kevin W. Hamlen. Hippocratic binary instrumentation: First do no harm. Science of Computer Programming, 93 (part B)(??):110– 124, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000914.

Gonnord:2014:AAL

[1685] Laure Gonnord and Peter Schrammel. Abstract acceleration in linear relation analysis. Science of Computer Programming, 93 (part B)(??): 125–153, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002463.

Mine:2014:BUA

[1686] A. Miné. Backward under-approximations in numeric abstract domains to automatically infer sufficient program conditions. Science of Computer Programming, 93 (part B)(??):154– 182, November 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300244X.

DiIorio:2014:WTS

[1687] Angelo Di Iorio, Davide Rossi, and Stefano Zacchiroli. Web technologies: Selected & extended papers from WT ACM SAC 2012. Science of Computer Programming, 94 (part 1)(??): 1–2, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001051.

Book:2014:PAC

[1688] Matthias Book, Marco Buss, and Volker Gruhn. Process-aware continuation management in web applications. Science of Computer Programming, 94 (part 1)(??):3– 17, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001895.

Liberman:2014:AOF

[1689] Sonya Liberman and Ronny Lempel. Approximately optimal facet value selection. Science of Computer Programming, 94 (part 1)(??): 18–31, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001937.

Hogenboom:2014:QLR

[1690] Frederik Hogenboom, Damir Vandic, Flavius Frasincar, Arnout Verheij, and Allard Kleijn. A query language and ranking algorithm for news items in the Hermes news processing framework. Science of Computer Programming, 94 (part 1)(??): 32–52, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001925.

Heinrich:2014:ESU

[1691] Matthias Heinrich, Franz Lehmann, Franz Josef Grüneberger, Martin Gaedke, Thomas Springer, and Alexander Schill. Enriching single-user web applications non-invasively with shared editing support. Science of Computer Programming, 94 (part 1)(??): 53–66, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001913.

Riccobene:2014:PAS

[1692] Elvinia Riccobene and Steve Reeves. Preface: Abstract State Machines, Alloy, B, VDM, and Z. Selected & extended papers from ABZ 2012. Science of Computer Programming, 94 (part 2)(??):67–68, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314002809.

Gervasi:2014:MWA

[1693] Vincenzo Gervasi, Egon Börger, and Antonio Cisternino. Modeling web applications infrastructure with ASMs. Science of Computer Programming, 94 (part 2)(??):69– 92, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000926.

Arcaini:2014:TGS

[1694] Paolo Arcaini and Angelo Gargantini. Test generation for sequential nets of Abstract State Machines with information passing. Science of Computer Programming, 94 (part 2)(??): 93–108, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000483.

Banach:2014:ACS

[1695] Richard Banach, Huibiao Zhu, Wen Su, and Xiaofeng Wu. ASM, controller synthesis, and complete refinement. Science of Computer Programming, 94 (part 2)(??):109– 129, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001932.

Deharbe:2014:ISS

[1696] David Déharbe, Pascal Fontaine, Yoann Guyot, and Laurent Voisin. Integrating SMT solvers in Rodin. Science of Computer Programming, 94 (part 2)(??):130–143, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016764231400183X.

Hallerstede:2014:RDM

[1697] Stefan Hallerstede and Thai Son Hoang. Refinement of decomposed models by interface instantiation. Science of Computer Programming, 94 (part 2)(??):144–163, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002330.

Su:2014:FHS

[1698] Wen Su, Jean-Raymond Abrial, and Huibiao Zhu. Formalizing hybrid systems with Event-B and the Rodin Platform. Science of Computer Programming, 94 (part 2)(??):164– 202, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002482.

Milicevic:2014:PAO

[1699] Aleksandar Milicevic and Daniel Jackson. Preventing arithmetic overflows in Alloy. Science of Computer Programming, 94 (part 2)(??):203– 216, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002470.

Fraikin:2014:SCT

[1700] Benoît Fraikin, Marc Frappier, and Richard St-Denis. Supervisory control theory with Alloy. Science of Computer Programming, 94 (part 2)(??): 217–237, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002494.

Jones:2014:RBT

[1701] Cliff B. Jones, Matthew J. Lovert, and L. Jason Steggles. Revising basic theorem proving algorithms to cope with the logic of partial functions. Science of Computer Programming, 94 (part 2)(??):238– 252, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002360.

Barbosa:2014:FAC

[1702] Luís Soares Barbosa and Markus Lumpe. Formal Aspects of Component Software (FACS 2010 selected and extended papers). Science of Computer Programming, 94 (part 3)(??): 253–254, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001506.

Zhao:2014:SCT

[1703] Liang Zhao, Roberto Bruni, and Zhiming Liu. A sound and complete theory of graph transformations for service programming with sessions and pipelines. Science of Computer Programming, 94 (part 3)(??):255– 288, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003225.

Bugliesi:2014:MCA

[1704] M. Bugliesi, A. Marin, and S. Rossi. Model checking adaptive service compositions. Science of Computer Programming, 94 (part 3)(??):289– 306, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003249.

Tamura:2014:QCP

[1705] Gabriel Tamura, Rubby Casallas, Anthony Cleve, and Laurence Duchien. QoS contract preservation through dynamic reconfiguration: a formal semantics approach. Science of Computer Programming, 94 (part 3)(??): 307–332, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003390.

Andova:2014:DAD

[1706] S. Andova, L. P. J. Groenewegen, and E. P. de Vink. Dynamic adaptation with distributed control in Paradigm. Science of Computer Programming, 94 (part 3)(??):333– 361, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003274.

Nobakht:2014:MMC

[1707] B. Nobakht, F. S. de Boer, M. M.

Bonsangue, S. de Gouw, and M. M. Jaghoori. Monitoring method call sequences using annotations. Science of Computer Programming, 94 (part 3)(??):362–378, November 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313003237.

Valencia-Garcia:2014:SIS

[1708] Rafael Valencia-García, Alejandro Rodríguez-González, and Ricardo Colomo-Palacios. Special issue on Systems Development by Means of Semantic Technologies. Science of Computer Programming, 95 (part 1)(??): 1–2, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001749.

Maio:2014:EAB

[1709] Paulo Maio and Nuno Silva. An extensible argument-based ontology matching negotiation approach. Science of Computer Programming, 95 (part 1) (??):3–25, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000264.

Faria:2014:DIP

[1710] Carla Faria, Ivo Serra, and Rosario Girardi. A domain-independent process for automatic ontology population from text. Science of Computer Programming, 95 (part 1)(??):

26–43, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003419.

Hamdaqa:2014:CSD

[1711] Mohammad Hamdaqa, Ladan Tahvildari, Neil LaChapelle, and Brian Campbell. Cultural scene detection using reverse Louvain optimization. Science of Computer Programming, 95 (part 1)(??):44– 72, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000124.

Ale:2014:CMT

[1712] Mariel A. Ale, Carlos M. Toledo, Omar Chiotti, and María R. Galli. A conceptual model and technological support for organizational knowledge management. Science of Computer Programming, 95 (part 1)(??): 73–92, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000100.

Wang:2014:ATS

[1713] Hai H. Wang, Danica Damljanović, and Jing Sun. An automated tool for semantic accessing to formal software models. Science of Computer Programming, 95 (part 1)(??):93– 111, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400094X.

Bonino:2014:EST

[1714] Dario Bonino and Giuseppe Procaccianti. Exploiting semantic technologies in smart environments and grids: Emerging roles and case studies. Science of Computer Programming, 95 (part 1)(??):112– 134, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000859.

Gayo:2014:IRR

[1715] Jose Emilio Labra Gayo, Johan Jeuring, and Jose María Álvarez Rodríguez. Inductive representations of RDF graphs. Science of Computer Programming, 95 (part 1)(??): 135–146, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000094.

Albert:2014:SEP

[1716] Elvira Albert and Shin-Cheng Mu.

Selected and extended papers from Partial Evaluation and Program Manipulation 2013. Science of Computer Programming, 95 (part 2)(??): 147–148, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002949.

DeAngelis:2014:PVI

[1717] E. De Angelis, F. Fioravanti, A. Pettorossi, and M. Proietti. Program verification via iterated specialization. Science of Computer Programming, 95 (part 2)(??):149– 175, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002871.

Garcia-Perez:2014:SFC

[1718] A. García-Pérez and P. Nogueira. On the syntactic and functional correspondence between hybrid (or layered) normalisers and abstract machines. Science of Computer Programming, 95 (part 2)(??):176– 199, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002597.

Weijers:2014:STE

[1719] Jeroen Weijers, Jurriaan Hage, and Stefan Holdermans. Security type error diagnosis for higher-order, polymorphic languages. Science of Computer Programming, 95 (part 2)(??): 200–218, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001518.

Servetto:2014:MCL

[1720] Marco Servetto and Elena Zucca. A meta-circular language for active libraries. Science of Computer Programming, 95 (part 2)(??):219– 253, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002317.

Simon:2014:DCT

[1721] Axel Simon. Deriving a complete type inference for Hindley– Milner and vector sizes using expansion. Science of Computer Programming, 95 (part 2)(??):254– 271, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001324.

Mousavi:2014:SSS

[1722] Mohammad Reza Mousavi and Jun Pang. Special section on Software Verification and Testing. Science of Computer Programming, 95 (part 3)(??): 273–274, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002950.

Carvalho:2014:ITC

[1723] Gustavo Carvalho, Diogo Falcão, Flávia Barros, Augusto Sampaio, Alexandre Mota, Leonardo Motta, and Mark Blackburn. NAT2TEST_{SCR}: Test case generation from natural language requirements based on SCR specifications. Science of Computer Programming, 95 (part 3)(??):275– 297, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002858.

Papadakis:2014:MEE

[1724] Mike Papadakis, Marcio Delamaro, and Yves Le Traon. Mitigating the effects of equivalent mutants with mutant classification strategies. Science of Computer Programming, 95 (part 3)(??):298– 319, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002603.

Chen:2014:SAL

[1725] Liqian Chen, Renjian Li, Xueguang Wu, and Ji Wang. Static analysis of lists by combining shape and numerical abstractions. Science of Computer Programming, 95 (part 3)(??): 320–342, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002822.

Gomez-Zamalloa:2014:SEP

[1726] Miguel Gómez-Zamalloa and Germán Puebla. Selected and extended papers from Bytecode 2013. Science of Computer Programming, 95 (part 3) (??):343, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400286X.

Sarimbekov:2014:DPA

[1727] Aibek Sarimbekov, Yudi Zheng, Danilo Ansaloni, Lubomír Bulej, Lukás Marek, Walter Binder, Petr Tuma, and Zhengwei Qi. Dynamic program analysis — reconciling developer productivity and tool performance. Science of Computer Programming, 95 (part 3)(??):344– 358, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001543.

Scapin:2014:FSU

[1728] Enrico Scapin and Fausto Spoto.

Field-sensitive unreachability and noncyclicity analysis. Science of Computer Programming, 95 (part 3)(??): 359–375, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400152X.

Blom:2014:SVG

[1729] Stefan Blom, Marieke Huisman, and Matej Mihelcić. Specification and verification of GPGPU programs. Science of Computer Programming, 95 (part 3)(??):376– 388, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001531.

Inoue:2014:SIS

[1730] Katsuro Inoue, Rainer Koschke, and Jens Krinke. Special issue on software clones (IWSC’12). Science of Computer Programming, 95 (part 4) (??):389, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003372.

Cuomo:2014:CFC

[1731] Antonio Cuomo, Antonella Santone, and Umberto Villano. CD-Form: a clone detector based on formal methods. Science of Computer Programming, 95 (part 4)(??):390– 405, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003067.

Tekin:2014:GMA

[1732] Umut Tekin and Feza Buzluca. A graph mining approach for detecting identical design structures in object-oriented design models. Science of Computer Programming, 95 (part 4)(??):406– 425, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002451.

Keivanloo:2014:SSC

[1733] Iman Keivanloo, Chanchal K. Roy, and Juergen Rilling. SeByte: Scalable clone and similarity search for bytecode. Science of Computer Programming, 95 (part 4)(??):426– 444, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002773.

Mondal:2014:IDC

[1734] Manishankar Mondal, Chanchal K. Roy, and Kevin A. Schneider. An insight into the dispersion of changes in cloned and non-cloned code: a genealogy based empirical study. Science of Computer Programming, 95 (part 4)(??):445–468, December 1, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313003110.

Lagadec:2014:PSI

[1735] Loïc Lagadec and Alain Plantec. Preface to the special issue on advances in Smalltalk based systems. Science of Computer Programming, 96 (part 1) (??):1–3, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003141.

Chisnall:2014:SCW

[1736] David Chisnall. Smalltalk in a C world. Science of Computer Programming, 96 (part 1)(??):4– 16, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002852.

Hlopko:2014:ISJ

[1737] Marcel Hlopko, Jan Kurs, Jan Vraný, and Claus Gittinger. On the integration of Smalltalk and Java. Science of Computer Programming, 96 (part 1)(??):17–33, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002839.

VanRyseghem:2014:SCR

[1738] Benjamin Van Ryseghem, Stéphane Ducasse, and Johan Fabry. Seamless composition and reuse of customizable user interfaces with Spec. Science of Computer Programming, 96 (part 1)(??):34–51, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313003286.

Allende:2014:GTS

[1739] Esteban Allende, Oscar Callaú, Johan Fabry, Éric Tanter, and Marcus Denker. Gradual typing for Smalltalk. Science of Computer Programming, 96 (part 1)(??):52– 69, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313001445.

Unterholzner:2014:IRT

[1740] Martin Unterholzner. Improving refactoring tools in Smalltalk using static type inference. Science of Computer Programming, 96 (part 1)(??): 70–83, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003250.

Gomez:2014:SSC

[1741] Verónica Uquillas Gómez, Stéphane Ducasse, and Andy Kellens. Supporting streams of changes during branch integration. Science of Computer Programming, 96 (part 1)(??): 84–106, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314003384.

Stevens:2014:LFG

[1742] Reinout Stevens, Coen De Roover, Carlos Noguera, Andy Kellens, and Viviane Jonckers. A logic foundation for a general-purpose history querying tool. Science of Computer Programming, 96 (part 1)(??):107– 120, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000653.

Kuhn:2014:VSC

[1743] Tobias Kuhn and Alexandre Bergel.

Verifiable source code documentation in controlled natural language. Science of Computer Programming, 96 (part 1)(??):121–140, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000069.

Polito:2014:BRS

[1744] G. Polito, S. Ducasse, L. Fabresse, N. Bouraqadi, and B. van Ryseghem. Bootstrapping reflective systems: the case of Pharo. Science of Computer Programming, 96 (part 1)(??):141– 155, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002797.

Lagadec:2014:MDT

[1745] Loïc Lagadec, Ciprian Teodorov, Jean-Christophe Le Lann, Damien Picard, and Erwan Fabiani. Modeldriven toolset for embedded reconfigurable cores: Flexible prototyping and software-like debugging. Science of Computer Programming, 96 (part 1)(??):156–174, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000823.

Junior:2014:SEP

[1746] Francisco Heron de Carvalho Junior and Luís Soares Barbosa. Selected and extended papers of the Brazilian Symposium on Programming Languages 2012. Science of Computer Programming, 96 (part 2)(??):175– 176, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231400330X.

Medeiros:2014:LRP

[1747] Sérgio Medeiros, Fabio Mascarenhas, and Roberto Ierusalimschy. Left recursion in Parsing Expression Grammars. Science of Computer Programming, 96 (part 2)(??):177– 190, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000288.

Reis:2014:FIA

[1748] Leonardo V. S. Reis, Roberto S. Bigonha, Vladimir O. Di Iorio, and Luis Eduardo S. Amorim. The formalization and implementation of Adaptable Parsing Expression Grammars. Science of Computer Programming, 96 (part 2)(??):191–210, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000872.

Viera:2014:AGM

[1749] Marcos Viera and S. Doaitse Swierstra. Attribute grammar macros. Science of Computer Programming, 96 (part 2)(??):211–229, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016764231400029X.

Gaia:2014:QQA

[1750] Felipe Nunes Gaia, Gabriel Coutinho Sousa Ferreira, Eduardo Figueiredo, and Marcelo de Almeida Maia. A quantitative and qualitative assessment of aspectual feature modules for evolving software product lines. Science of Computer Programming, 96 (part 2)(??): 230–253, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001336.

Cunha:2014:MBP

[1751] Jácome Cunha, Jorge Mendes, João Saraiva, and Joost Visser. Modelbased programming environments for spreadsheets. Science of Computer Programming, 96 (part 2)(??):254– 275, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000343.

Luttgen:2014:SIA

[1752] Gerald Lüttgen and Stephan Merz.

Special issue on Automated Verification of Critical Systems (AVoCS’12). Science of Computer Programming, 96 (part 3)(??):277– 278, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400135X.

Marche:2014:VFB

[1753] Claude Marché. Verification of the functional behavior of a floatingpoint program: an industrial case study. Science of Computer Programming, 96 (part 3)(??):279–296, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001671.

Tofan:2014:TAP

[1754] Bogdan Tofan, Oleg Travkin, Gerhard Schellhorn, and Heike Wehrheim. Two approaches for proving linearizability of multiset. Science of Computer Programming, 96 (part 3)(??): 297–314, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001658.

James:2014:MVR

[1755] Phillip James, Faron Moller, Hoang Nga Nguyen, Markus Roggenbach, Steve Schneider, and Helen Treharne. On modelling and verifying railway interlockings: Tracking train lengths. Science of Computer Programming, 96 (part 3)(??):315–336, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001695.

Zhao:2014:FSV

[1756] Yang Zhao and Kristin Yvonne Rozier. Formal specification and verification of a coordination protocol for an automated air traffic control system. Science of Computer Programming, 96 (part 3)(??):337– 353, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400166X.

Mateescu:2014:PDR

[1757] Radu Mateescu and Anton Wijs.

Property-dependent reductions adequate with divergence-sensitive branching bisimilarity. Science of Computer Programming, 96 (part 3)(??): 354–376, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001683.

Talpin:2014:CPS

[1758] Jean-Pierre Talpin, Jens Brandt, Mike Gemünde, Klaus Schneider, and Sandeep Shukla. Constructive polychronous systems. Science of Computer Programming, 96 (part 3)(??): 377–394, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001737.

Czarnecki:2014:SLE

[1759] Krzysztof Czarnecki and Görel Hedin.

Software language engineering (SLE ’12). Science of Computer Programming, 96 (part 4)(??):395, December 15, 2014. CODEN SCPGD4.

ISSN 0167-6423 (print), 1872-7964 (electronic). URL http://www.

                                        sciencedirect.com/science/article/
                                        pii/S0167642314003803.

Paige:2014:TMG

[1760] Richard F. Paige, Dimitrios S. Kolovos, and Fiona A. C. Polack. A tutorial on metamodelling for grammar researchers. Science of Computer Programming, 96 (part 4)(??): 396–416, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002457.

Guizzardi:2014:UTB

[1761] Giancarlo Guizzardi and Veruska Zamborlini. Using a trope-based foundational ontology for bridging different areas of concern in ontology-driven conceptual modeling. Science of Computer Programming, 96 (part 4)(??): 417–443, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000896.

Stevenson:2014:SGI

[1762] Andrew Stevenson and James R. Cordy. A survey of grammatical inference in software engineering. Science of Computer Programming, 96 (part 4)(??):444–459, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002469.

vanderStorm:2014:DIO

[1763] Tijs van der Storm, William R.

Cook, and Alex Loh. The design and implementation of Object Grammars. Science of Computer Programming, 96 (part 4)(??):460– 487, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000902.

Sloane:2014:DSP

[1764] Anthony M. Sloane and Matthew Roberts. Domain-specific program profiling and its application to attribute grammars and term rewriting. Science of Computer Programming, 96 (part 4)(??):488–510, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000628.

Krishnan:2014:MMT

[1765] Lijesh Krishnan and Eric Van Wyk.

Monolithic and modular termination analyses for higher-order attribute grammars. Science of Computer Programming, 96 (part 4)(??):511– 526, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002780.

Kanso:2014:STP

[1766] Bilal Kanso and Safouan Taha.

Specification of temporal properties with OCL. Science of Computer Programming, 96 (part 4)(??):527– 551, December 15, 2014. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000963.

Vinju:2015:P

[1767] Jurgen J. Vinju. Preface. Science of Computer Programming, 97 (part 1)(??):1, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314001828.

Mosses:2015:SPL

[1768] Peter D. Mosses. Semantics of programming languages: Using Asf + Sdf. Science of Computer Programming, 97 (part 1)(??):2– 10, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003316.

Visser:2015:UST

[1769] Eelco Visser. Understanding software through linguistic abstraction. Science of Computer Programming, 97 (part 1) (??):11–16, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003365.

Baeten:2015:RSC

[1770] Jos Baeten and Jasen Markovski. The role of supervisory controller synthesis in automatic control software development. Science of Computer Programming, 97 (part 1)(??): 17–22, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003006.

Bruni:2015:CDR

[1771] Roberto Bruni, Alberto Lluch Lafuente, and Ugo Montanari. Constraint design rewriting. Science of Computer Programming, 97 (part 1) (??):23–30, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002992.

Heitkotter:2015:EMD

[1772] Henning Heitkötter, Herbert Kuchen, and Tim A. Majchrzak. Extending a model-driven cross-platform development approach for business apps. Science of Computer Programming, 97 (part 1)(??):31–36, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313002979.

Gupta:2015:LBS

[1773] Gopal Gupta. Language-based software engineering. Science of Computer Programming, 97 (part 1)(??): 37–40, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000616.

Preda:2015:DCD

[1774] Mila Dalla Preda, Maurizio Gabbrielli, Saverio Giallorenzo, Ivan Lanese, and Jacopo Mauro. Developing correct, distributed, adaptive software. Science of Computer Programming, 97 (part 1)

(??):41–46, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003031.

Bianculli:2015:SSI

[1775] Domenico Bianculli, Antonio Filieri, Carlo Ghezzi, and Dino Mandrioli. Syntactic-semantic incrementality for agile verification. Science of Computer Programming, 97 (part 1)(??): 47–54, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003109.

Watson:2015:ECC

[1776] B. W. Watson, D. G. Kourie, and L. Cleophas. Experience with correctness-by-construction. Science of Computer Programming, 97 (part 1) (??):55–58, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003080.

Stevenson:2015:PVB

[1777] Andrew Stevenson and James R. Cordy. Parse views with Boolean grammars. Science of Computer Programming, 97 (part 1)(??):59– 63, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002918.

Johnstone:2015:PSM

[1778] Adrian Johnstone and Elizabeth Scott. Principled software microengineering. Science of Computer Programming, 97 (part 1)(??):64– 68, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300302X.

vanEijck:2015:UIU

[1779] Jan van Eijck and Tijs van der Storm.

Understanding information update in questionnaires. Science of Computer Programming, 97 (part 1)(??): 69–74, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003353.

vandenBrand:2015:SER

[1780] Mark van den Brand and Jan Friso Groote. Software engineering: Redundancy is key. Science of Computer Programming, 97 (part 1)(??): 75–81, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003043.

Heering:2015:GSC

[1781] Jan Heering. Generative software complexity and software understanding. Science of Computer Programming, 97 (part 1)(??):82– 85, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003018.

Godfrey:2015:USA

[1782] Michael W. Godfrey. Understanding software artifact provenance. Science of Computer Programming, 97 (part 1) (??):86–90, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003055.

Tip:2015:IPO

[1783] Frank Tip. Infeasible paths in objectoriented programs. Science of Computer Programming, 97 (part 1)(??): 91–97, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231300289X.

Lammel:2015:SC

[1784] Ralf Lämmel. Software chrestomathies. Science of Computer Programming, 97 (part 1)(??):98–104, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642313002980.

Moonen:2015:TEB

[1785] Leon Moonen. Towards evidencebased recommendations to guide the evolution of component-based product families. Science of Computer Programming, 97 (part 1)(??):105– 112, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002931.

Cleve:2015:UDS

[1786] Anthony Cleve, Maxime Gobert, Loup Meurice, Jerome Maes, and Jens Weber. Understanding database schema evolution: a case study. Science of Computer Programming, 97 (part 1)(??):113–121, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313003092.

Snelting:2015:UPS

[1787] Gregor Snelting. Understanding probabilistic software leaks. Science of Computer Programming, 97 (part 1)(??): 122–126, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231300292X.

Urma:2015:SCQ

[1788] Raoul-Gabriel Urma and Alan Mycroft. Source-code queries with graph databases-with application to programming language usage and evolution. Science of Computer Programming, 97 (part 1)(??):127– 134, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002943.

Bruntink:2015:TBR

[1789] Magiel Bruntink. Towards base rates in software analytics: Early results and challenges from studying Ohloh. Science of Computer Programming, 97 (part 1)(??):135– 142, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003079.

vanderStorm:2015:TMP

[1790] Tijs van der Storm and Jurgen J. Vinju. Towards multilingual programming environments. Science of Computer Programming, 97 (part 1)(??): 143–149, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313003341.

Nierstrasz:2015:PAM

[1791] Oscar Nierstrasz and Jan Kurs. Parsing for agile modeling. Science of Computer Programming, 97 (part 1)(??): 150–156, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002955.

Heinen:2015:VPP

[1792] Jonathan Heinen, Christina Jansen, Joost-Pieter Katoen, and Thomas Noll. Verifying pointer programs using graph grammars. Science of Computer Programming, 97 (part 1)(??): 157–162, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642313002967.

Kogtenkov:2015:ACC

[1793] Alexander Kogtenkov, Bertrand Meyer, and Sergey Velder. Alias calculus, change calculus and frame inference. Science of Computer Programming, 97 (part 1)(??):163– 172, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002906.

vanDeursen:2015:CBA

[1794] Arie van Deursen, Ali Mesbah, and Alex Nederlof. Crawl-based analysis of web applications: Prospects and challenges. Science of Computer Programming, 97 (part 1)(??): 173–180, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003815.

Lanese:2015:SIS

[1795] Ivan Lanese, Manuel Mazzara, and Fabrizio Montesi. Special issue on Service-Oriented Architecture and Programming (SOAP 2013). Science of Computer Programming, 97 (part 2)(??):181–182, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314003293.

Karatas:2015:SCC

[1796] Fatih Karatas, Lars Fischer, and Dogan Kesdogan. Service composition with consideration of interdependent security objectives. Science of Computer Programming, 97 (part 2)(??): 183–201, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002962.

Cameron:2015:RBP

[1797] Alexander Cameron, Markus Stumptner, Nanda Nandagopal, Wolfgang Mayer, and Todd Mansell. Rulebased peer-to-peer framework for decentralised real-time service oriented architectures. Science of Computer Programming, 97 (part 2)(??):202– 234, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002834.

Giusto:2015:DSC

[1798] Cinzia Di Giusto and Jorge A. Pérez. Disciplined structured communications with disciplined runtime adaptation. Science of Computer Programming, 97 (part 2)(??):235– 265, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002512.

Zagarese:2015:IDI

[1799] Quirino Zagarese, Gerardo Canfora, Eugenio Zimeo, Iyad Alshabani, Laurent Pellegrino, Amjad Alshabani, and Françoise Baude. Improving data-intensive EDA performance with annotation-driven laziness. Science of Computer Programming, 97 (part 2)(??):266– 279, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314001348.

Ancona:2015:PSS

[1800] Davide Ancona. Preface to the special section on Object-Oriented Programming and Systems (OOPS 2010). Science of Computer Programming, 97 (part 3)(??):281, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314002937.

Bettini:2015:ITS

[1801] Lorenzo Bettini, Ferruccio Damiani, and Ina Schaefer. Implementing type-safe software product lines using parametric traits. Science of Computer Programming, 97 (part 3)(??): 282–308, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001901.

Rooney:2015:SIA

[1802] Sean Rooney. Scheduling intense applications most ‘surprising’ first. Science of Computer Programming, 97 (part 3)(??):309–319, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642313002232.

Tang:2015:AGR

[1803] Xiaolong Tang and Jaakko Järvi. Axioms as generic rewrite rules in C++ with concepts. Science of Computer Programming, 97 (part 3)(??): 320–330, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002433.

Cortellessa:2015:MAC

[1804] Vittorio Cortellessa and Henry Muccini. Modeling and analysis of compositional software (papers from EUROMI-CRO SEAA’12). Science of Computer Programming, 97 (part 3)(??): 331–332, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002883.

Cimatti:2015:CRP

[1805] Alessandro Cimatti and Stefano Tonetta. Contracts-refinement proof system for component-based embedded systems. Science of Computer Programming, 97 (part 3)(??):333– 348, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002901.

Brada:2015:RMD

[1806] Premek Brada and Kamil Jezek. Repository and meta-data design for efficient component consistency verification. Science of Computer Programming, 97 (part 3)(??):349– 365, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002925.

Etemaadi:2015:NDF

[1807] Ramin Etemaadi and Michel R. V. Chaudron. New degrees of freedom in metaheuristic optimization of component-based systems architecture: Architecture topology and load balancing. Science of Computer Programming, 97 (part 3)(??): 366–380, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002913.

Anonymous:2015:EBa

[1808] Anonymous. Editorial Board. Science of Computer Programming, 97 (part 3)(??):ifc, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004924.

Muccini:2015:SIS

[1809] Henry Muccini and Antony Tang.

Special issue: Selected papers from the 12th International Conference on Quality Software (QSIC 2012). Science of Computer Programming, 97 (part 4)(??):381–382, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314002160.

Aichernig:2015:MBM

[1810] Bernhard K. Aichernig, Elisabeth Jöbstl, and Stefan Tiran. Modelbased mutation testing via symbolic refinement checking. Science of Computer Programming, 97 (part 4)(??): 383–404, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002329.

Shahbaz:2015:AGV

[1811] Muzammil Shahbaz, Phil McMinn, and Mark Stevenson. Automatic generation of valid and invalid test data for string validation routines using web searches and regular expressions. Science of Computer Programming, 97 (part 4)(??):405– 425, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001725.

Pham:2015:RPC

[1812] Thanh-Trung Pham, Xavier Défago, and Quyet-Thang Huynh. Reliability prediction for componentbased software systems: Dealing with concurrent and propagating errors. Science of Computer Programming, 97 (part 4)(??):426– 457, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002172.

Anonymous:2015:EBb

[1813] Anonymous. Editorial Board. Science of Computer Programming, 97 (part 4)(??):ifc, January 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005024.

vandenBrand:2015:GEI

[1814] Mark van den Brand, Davide Di Ruscio, Dimitrios S. Kolovos, and Louis M. Rose. Guest Editors’ introduction to the fifth issue of Experimental Software and Toolkits (EST): a special issue on Academics Modelling with Eclipse (ACME2012). Science of Computer Programming, 98 (part 1)(??): 1–2, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005103.

Jimenez:2015:MTI

[1815] Álvaro Jiménez, Juan M. Vara, Verónica A. Bollati, and Esperanza Marcos. MeTAGeM-Trace: Improving trace generation in model transformation by leveraging the role of transformation models. Science of Computer Programming, 98 (part 1)(??): 3–27, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003700.

Zaytsev:2015:GZC

[1816] Vadim Zaytsev. Grammar Zoo:

a corpus of experimental grammarware. Science of Computer Programming, 98 (part 1)(??):28– 51, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003347.

Amalio:2015:TVF

[1817] Nuno Amálio and Christian Glodt. A tool for visual and formal modelling of software designs. Science of Computer Programming, 98 (part 1)(??): 52–79, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002305.

Ujhelyi:2015:EII

[1818] Zoltán Ujhelyi, Gábor Bergmann, Ábel Hegedüs, Ákos Horváth, Benedek Izsó, István Ráth, Zoltán Szatmári, and Dániel Varró. EMF-IncQuery: an integrated development environment for live model queries. Science of Computer Programming, 98 (part 1)(??): 80–99, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000082.

Marek:2015:IDP

[1819] Lukás Marek, Yudi Zheng, Danilo Ansaloni, Lubomír Bulej, Aibek Sarimbekov, Walter Binder, and Petr Tuma. Introduction to dynamic program analysis with DiSL. Science of Computer Programming, 98 (part 1)(??): 100–115, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000070.

Anonymous:2015:EBc

[1820] Anonymous. Editorial Board. Science of Computer Programming, 98 (part 1) (??):ifc, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005127.

Ricci:2015:SIP

[1821] Alessandro Ricci, Gul Agha, Rafael H. Bordini, and Assaf Marron. Special issue on programming based on actors, agents and decentralized control. Science of Computer Programming, 98 (part 2)(??):117– 119, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004626.

Jordan:2015:FMA

[1822] Howell Jordan, Goetz Botterweck, John Noll, Andrew Butterfield, and Rem Collier. A feature model of actor, agent, functional, object, and procedural programming languages. Science of Computer Programming, 98 (part 2)(??):120– 139, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000501.

DeKoster:2015:DSS

[1823] Joeri De Koster, Stefan Marr, Theo D’Hondt, and Tom Van Cutsem. Domains: Safe sharing among actors. Science of Computer Programming, 98 (part 2)(??):140– 158, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000495.

Marr:2015:PGR

[1824] Stefan Marr, Thierry Renaux, Lode Hoste, and Wolfgang De Meuter. Parallel gesture recognition with soft realtime guarantees. Science of Computer Programming, 98 (part 2)(??): 159–183, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231400063X.

Khamespanah:2015:TRS

[1825] Ehsan Khamespanah, Marjan Sirjani, Zeynab Sabahi Kaviani, Ramtin Khosravi, and Mohammad-Javad Izadi. Timed Rebeca schedulability and deadlock freedom analysis using bounded floating time transition system. Science of Computer Programming, 98 (part 2)(??):184– 204, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314003153.

Fernandez-Diaz:2015:ADF

[1826] Á. Fernández-Díaz, C. Benac-Earle, and L. Fredlund. Adding distribution and fault tolerance to Jason. Science of Computer Programming, 98 (part 2)(??):205– 232, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314000227.

Harel:2015:TBP

[1827] David Harel, Amir Kantor, Guy Katz, Assaf Marron, Gera Weiss, and Guy Wiener. Towards behavioral programming in distributed architectures. Science of Computer Programming, 98 (part 2)(??):233– 267, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314001075.

Ashrov:2015:UCB

[1828] Adiel Ashrov, Assaf Marron, Gera Weiss, and Guy Wiener. A use-case for behavioral programming: an architecture in JavaScript and Blockly for interactive applications with crosscutting scenarios. Science of Computer Programming, 98 (part 2)(??): 268–292, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400032X.

Bensalem:2015:ODI

[1829] Saddek Bensalem, Marius Bozga, Jean Quilbeuf, and Joseph Sifakis. Optimized distributed implementation of multiparty interactions with Restriction. Science of Computer Programming, 98 (part 2)(??):293– 316, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000641.

Laval:2015:ADL

[1830] Jannik Laval and Andy Kellens. Advances in Dynamic Languages. Science of Computer Programming, 98 (part 3) (??):317, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004869.

Arnaud:2015:HBP

[1831] Jean-Baptiste Arnaud, Stéphane Ducasse, Marcus Denker, and Camille Teruel. Handles: Behavior-propagating first class references for dynamically-typed languages. Science of Computer Programming, 98 (part 3)(??):318– 338, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003359.

Peck:2015:GUG

[1832] Mariano Martinez Peck, Noury Bouraqadi, Luc Fabresse, Marcus Denker, and Camille Teruel. Ghost: a uniform and general-purpose proxy implementation. Science of Computer Programming, 98 (part 3)(??):339– 359, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002779.

Scholliers:2015:CC

[1833] Christophe Scholliers, Éric Tanter, and Wolfgang De Meuter. Computational contracts. Science of Computer Programming, 98 (part 3)(??): 360–375, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313002347.

Gomez:2015:VCS

[1834] Verónica Uquillas Gómez, Stéphane Ducasse, and Theo D’Hondt. Visually characterizing source code changes. Science of Computer Programming, 98 (part 3)(??):376– 393, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001950.

Olivero:2015:OFE

[1835] Fernando Olivero, Michele Lanza, and Marco D’Ambros. Object-focused environments revisited. Science of Computer Programming, 98 (part 3)(??): 394–407, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313001688.

Bolz:2015:IMT

[1836] Carl Friedrich Bolz and Laurence Tratt. The impact of metatracing on VM design and implementation. Science of Computer Programming, 98 (part 3)(??):408– 421, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313000269.

VanCutsem:2015:RTC

[1837] Tom Van Cutsem and Mark S.

Miller. Robust trait composition for JavaScript. Science of Computer Programming, 98 (part 3)(??): 422–438, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642312002079.

Anonymous:2015:EBd

[1838] Anonymous. Editorial Board. Science of Computer Programming, 98 (part 2) (??):ifc, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005206.

Anonymous:2015:EBe

[1839] Anonymous. Editorial Board. Science of Computer Programming, 98 (part 3) (??):ifc, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005279.

Cortellessa:2015:MES

[1840] Vittorio Cortellessa, Raffaela Mirandola, and Pasqualina Potena. Managing the evolution of a software architecture at minimal cost under performance and reliability constraints. Science of Computer Programming, 98 (part 4)(??):439– 463, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002792.

Bozzano:2015:SAA

[1841] Marco Bozzano, Alessandro Cimatti, Oleg Lisagor, Cristian Mattarei, Sergio Mover, Marco Roveri, and Stefano Tonetta. Safety assessment of AltaRica models via symbolic model checking. Science of Computer Programming, 98 (part 4)(??): 464–483, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002810.

Mu:2015:ATD

[1842] Shin-Cheng Mu, Yu-Han Lyu, and Akimasa Morihata. Approximate by thinning: Deriving fully polynomialtime approximation schemes. Science of Computer Programming, 98 (part 4)(??):484–515, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314002974.

Majster-Cederbaum:2015:BCE

[1843] Mila Majster-Cederbaum and Nils Semmelrock. A basis for compositionally ensuring safety properties and its connection to relational algebraic operators. Science of Computer Programming, 98 (part 4)(??): 516–530, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003165.

Goc:2015:EFM

[1844] Yannick Le Goc and Alexandre Donzé.

EVL: a framework for multi-methods in C++. Science of Computer Programming, 98 (part 4)(??):531– 550, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003360.

Kwiatkowski:2015:ROC

[1845] Lukasz M. Kwiatkowski and Chris Verhoef. Reducing operational costs through MIPS management. Science of Computer Programming, 98 (part 4)(??):551–588, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314003487.

Clark:2015:XLM

[1846] Tony Clark. XPL: a language for modular homogeneous language embedding. Science of Computer Programming, 98 (part 4)(??):589– 616, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314003670.

Song:2015:RMA

[1847] Myoungkyu Song and Eli Tilevich. Reusing metadata across components, applications, and languages. Science of Computer Programming, 98 (part 4)(??):617– 644, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314003682.

Margoor:2015:TTI

[1848] Amogh Margoor and Raghavan Komondoor. Two techniques to improve the precision of a demanddriven null-dereference verification approach. Science of Computer Programming, 98 (part 4)(??):645– 679, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314003827.

Abebe:2015:EDC

[1849] Surafel Lemma Abebe and Paolo Tonella. Extraction of domain concepts from the source code. Science of Computer Programming, 98 (part 4)(??): 680–706, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004419.

Margaris:2015:IFA

[1850] D. Margaris, C. Vassilakis, and P. Georgiadis. An integrated framework for adapting WS-BPEL scenario execution using QoS and collaborative filtering techniques. Science of Computer Programming, 98 (part 4)(??): 707–734, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004778.

Parizek:2015:MCC

[1851] Pavel Parízek and Ondrej Lhoták.

Model checking of concurrent programs with static analysis of field accesses. Science of Computer Programming, 98 (part 4)(??):735– 763, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400478X.

Sanchez:2015:PAR

[1852] Alejandro Sanchez, Nuno Oliveira, Luis S. Barbosa, and Pedro Henriques. A perspective on architectural re-engineering. Science of Computer Programming, 98 (part 4)(??): 764–784, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000938.

Anonymous:2015:EBf

[1853] Anonymous. Editorial Board. Science of Computer Programming, 98 (part 4) (??):ifc, February 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005474.

Duran:2015:PRL

[1854] Francisco Durán and Narciso Martí-Oliet. Preface to Rewriting Logic and Its Applications (extended selected papers from WRLA 2012). Science of Computer Programming, 99 (??):1–2, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314003311.

Tushkanova:2015:RBS

[1855] E. Tushkanova, A. Giorgetti, C. Ringeissen, and O. Kouchnarenko. A rulebased system for automatic decidability and combinability. Science of Computer Programming, 99(??): 3–23, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231400046X.

Eckhardt:2015:SDI

[1856] Jonas Eckhardt, Tobias Mühlbauer, José Meseguer, and Martin Wirsing. Semantics, distributed implementation, and formal analysis of KLAIM models in Maude. Science of Computer Programming, 99(??): 24–74, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004420.

Bruni:2015:MAA

[1857] Roberto Bruni, Andrea Corradini, Fabio Gadducci, Alberto Lluch Lafuente, and Andrea Vandin. Modelling and analyzing adaptive selfassembly strategies with Maude. Science of Computer Programming, 99 (??):75–94, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642313003389.

Fadlisyah:2015:FMA

[1858] Muhammad Fadlisyah, Peter Csaba Ölveczky, and Erika Ábrahám. Formal modeling and analysis of interacting hybrid systems in HI-Maude: What happened at the 2010 Sauna World Championships? Science of Computer Programming, 99(??): 95–127, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002895.

Lepri:2015:SCT

[1859] Daniela Lepri, Erika Ábrahám, and Peter Csaba Ölveczky. Sound and complete timed CTL model checking of timed Kripke structures and real-time rewrite theories. Science of Computer Programming, 99(??): 128–192, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314002846.

Bae:2015:MCL

[1860] Kyungmin Bae and José Meseguer. Model checking linear temporal logic of rewriting formulas under localized fairness. Science of Computer Programming, 99(??):193–234, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642314000471.

Gutierrez:2015:OSE

[1861] Raúl Gutiérrez, José Meseguer, and Camilo Rocha. Order-sorted equality enrichments modulo axioms. Science of Computer Programming, 99 (??):235–261, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002998.

Anonymous:2015:EBg

[1862] Anonymous. Editorial Board. Science of Computer Programming, 99 (??):ifc, March 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005620.

Carbone:2015:PSIa

[1863] Marco Carbone, Ivan Lanese, Alexandra Silva, and Ana Sokolova. Preface for the special issue on Interaction and Concurrency Experience 2012. Science of Computer Programming, 100(??):1, March 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005188.

Poyias:2015:DCA

[1864] Kyriakos Poyias and Emilio Tuosto.

A design-by-contract approach to recover the architectural style from run-time misbehaviour. Science of Computer Programming, 100(??):2– 27, March 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004754.

Lozes:2015:SCO

[1865] Étienne Lozes and Jules Villard.

Shared contract-obedient channels. Science of Computer Programming, 100(??):28–60, March 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314003931.

Asavoae:2015:MCR

[1866] I. M. Asavoae, F. de Boer, M. M.

Bonsangue, D. Lucanu, and J. Rot. Model checking recursive programs interacting via the heap. Science of Computer Programming, 100(??): 61–83, March 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003943.

Pino:2015:WCB

[1867] Luis F. Pino, Andrés Aristizábal, Filippo Bonchi, and Frank Valencia. Weak CCP bisimilarity with strong procedures. Science of Computer Programming, 100(??):84– 104, March 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231400392X.

Anonymous:2015:EBh

[1868] Anonymous. Editorial Board. Science of Computer Programming, 100 (??):ifc, March 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000039.

Johnson:2015:TGT

[1869] Pontus Johnson, Mathias Ekstedt, Michael Goedicke, and Ivar Jacobson. Towards general theories of software engineering. Science of Computer Programming, 101 (??):1–5, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005371.

Erbas:2015:MTB

[1870] Cengiz Erbas and Bahar Celikkol Erbas. Modules and transactions: Building blocks for a theory of software engineering. Science of Computer Programming, 101 (??):6–20, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005383.

Ralph:2015:SCI

[1871] Paul Ralph. The Sensemaking– Coevolution–Implementation Theory of software design. Science of Computer Programming, 101(??): 21–41, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005395.

Stoica:2015:SCG

[1872] Anca-Juliana Stoica, Kristiaan Pelckmans, and William Rowe. System components of a general theory of software engineering. Science of Computer Programming, 101 (??):42–65, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005401.

Ng:2015:ISE

[1873] Pan-Wei Ng. Integrating software engineering theory and practice using essence: a case study. Science of Computer Programming, 101 (??):66–78, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005413.

Stol:2015:TOS

[1874] Klaas-Jan Stol and Brian Fitzgerald.

Theory-oriented software engineering. Science of Computer Programming, 101(??):79–98, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005425.

Perry:2015:FEU

[1875] Dewayne E. Perry. A framework for exploring unifying theories of empirical software engineering. Science of Computer Programming, 101 (??):99–123, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005437.

Paivarinta:2015:TAS

[1876] Tero Päivärinta and Kari Smolander. Theorizing about software development practices. Science of Computer Programming, 101(??):124– 135, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005449.

Wieringa:2015:SSG

[1877] Roel Wieringa and Maya Daneva. Six strategies for generalizing software engineering theories. Science of Computer Programming, 101(??): 136–152, April 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005450.

Salas-Zarate:2015:ABP

[1878] María del Pilar Salas-Zárate, Giner Alor-Hernández, Rafael Valencia-García, Lisbeth Rodríguez-Mazahua, Alejandro Rodríguez-González, and José Luis López Cuadrado. Analyzing best practices on Web development frameworks: the lift approach. Science of Computer Programming, 102 (??):1–19, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005735.

Salvaneschi:2015:CLD

[1879] Guido Salvaneschi, Carlo Ghezzi, and Matteo Pradella. ContextErlang: a language for distributed contextaware self-adaptive applications. Science of Computer Programming, 102 (??):20–43, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005577.

Gatrell:2015:ERC

[1880] M. Gatrell and S. Counsell. The effect of refactoring on change and faultproneness in commercial C# software. Science of Computer Programming, 102(??):44–56, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005711.

Hesselink:2015:MEF

[1881] Wim H. Hesselink. Mutual exclusion by four shared bits with not more than quadratic complexity. Science of Computer Programming, 102 (??):57–75, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500009X.

Backhouse:2015:CCT

[1882] Roland Backhouse and Hai Truong.

The capacity-C torch problem. Science of Computer Programming, 102 (??):76–107, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000118.

Leger:2015:ESA

[1883] Paul Leger, Éric Tanter, and Hiroaki Fukuda. An expressive stateful aspect language. Science of Computer Programming, 102(??): 108–141, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000258.

Vidal:2015:SEB

[1884] Germán Vidal. Symbolic execution as a basis for termination analysis. Science of Computer Programming, 102 (??):142–157, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000271.

Anonymous:2015:EBi

[1885] Anonymous. Editorial Board. Science of Computer Programming, 102 (??):ifc, May 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000313.

Artho:2015:Pa

[1886] Cyrille Artho and Peter Csaba Ölveczky. Preface. Science of Computer Programming, 103(??): 1–2, June 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000246.

Huuck:2015:TTF

[1887] Ralf Huuck. Technology transfer: Formal analysis, engineering, and business value. Science of Computer Programming, 103(??): 3–12, June 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005358.

Bae:2015:DVD

[1888] Kyungmin Bae, Joshua Krisiloff, José Meseguer, and Peter Csaba Ölveczky. Designing and verifying distributed cyber-physical systems using Multirate PALS: an airplane turning control system case study. Science of Computer Programming, 103(??): 13–50, June 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004109.

Choi:2015:ESC

[1889] Yunja Choi, Mingyu Park, Taejoon Byun, and Dongwoo Kim. Efficient safety checking for automotive operating systems using propertybased slicing and constraint-based environment generation. Science of Computer Programming, 103(??): 51–70, June 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004766.

Champion:2015:GPD

[1890] Adrien Champion, Rémi Delmas, and Michael Dierkes. Generating property-directed potential invariants by quantifier elimination in a kinduction-based framework. Science of Computer Programming, 103 (??):71–87, June 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004638.

Anonymous:2015:EBj

[1891] Anonymous. Editorial Board. Science of Computer Programming, 103 (??):ifc, June 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000416.

Ribeiro:2015:PSI

[1892] Leila Ribeiro and Matthias Tichy. Preface for the Special Issue on Graph Transformation and Visual Modeling Techniques 2013. Science of Computer Programming, 104 (??):1, June 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000544.

Mantz:2015:CEM

[1893] Florian Mantz, Gabriele Taentzer, Yngve Lamo, and Uwe Wolter. Coevolving meta-models and their instance models: a formal approach based on graph transformation. Science of Computer Programming, 104 (??):2–43, June 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000106.

Maximova:2015:LCA

[1894] Maria Maximova, Hartmut Ehrig, and Claudia Ermel. Local confluence analysis of hypergraph transformation systems with application conditions based on M-functors and Agg. Science of Computer Programming, 104 (??):44–70, June 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004432.

Bruggink:2015:RCP

[1895] H. J. Sander Bruggink, Barbara König, and Sebastian Küpper. Robustness and closure properties of recognizable languages in adhesive categories. Science of Computer Programming, 104 (??):71–98, June 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003694.

Adams:2015:SSS

[1896] Bram Adams and Juergen Rilling. Special section on Source Code Analysis and Manipulation (SCAM 2013). Science of Computer Programming, 105(??):1–2, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000817.

Beszedes:2015:EIS

[1897] Árpád Beszédes, Lajos Schrettner, Béla Csaba, Tamás Gergely, Judit Jász, and Tibor Gyimóthy. Empirical investigation of SEA-based dependence cluster properties. Science of Computer Programming, 105 (??):3–25, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314004092.

Cerulo:2015:IHM

[1898] Luigi Cerulo, Massimiliano Di Penta, Alberto Bacchelli, Michele Ceccarelli, and Gerardo Canfora. Irish: a Hidden Markov Model to detect coded information islands in free text. Science of Computer Programming, 105 (??):26–43, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005589.

Marcozzi:2015:RSE

[1899] Michaël Marcozzi, Wim Vanhoof, and Jean-Luc Hainaut. Relational symbolic execution of SQL code for unit testing of database programs. Science of Computer Programming, 105 (??):44–72, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000660.

Lozano:2015:UCO

[1900] Angela Lozano, Kim Mens, and Andy Kellens. Usage contracts: Offering immediate feedback on violations of structural source-code regularities. Science of Computer Programming, 105 (??):73–91, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500012X.

Banach:2015:CHE

[1901] Richard Banach, Michael Butler, Shengchao Qin, Nitika Verma, and Huibiao Zhu. Core hybrid Event-B I: Single hybrid Event-B machines. Science of Computer Programming, 105 (??):92–123, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000283.

Kakkad:2015:CEE

[1902] Vasvi Kakkad, Andrew E. Santosa, Alan Fekete, and Bernhard Scholz. Computing end-to-end delays in stream query processing. Science of Computer Programming, 105 (??):124–144, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000696.

Midtgaard:2015:SDC

[1903] Jan Midtgaard, Aleksandar S. Dimovski, Claus Brabrand, and Andrzej Wasowski. Systematic derivation of correct variability-aware program analyses. Science of Computer Programming, 105(??):145– 170, July 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000714.

Hugues:2015:PSI

[1904] Jérôme Hugues and Oleg Sokolsky. Preface to the special issue: Architecture-Driven Semantic Analysis of Embedded Systems. Science of Computer Programming, 106 (??):1–2, August 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500091X.

Balasubramanian:2015:DMW

[1905] Daniel Balasubramanian, Abhishek Dubey, William Otte, Tihamer Levendovszky, Aniruddha Gokhale, Pranav Kumar, William Emfinger, and Gabor Karsai. DREMS ML: a wide spectrum architecture design language for distributed computing platforms. Science of Computer Programming, 106 (??):3–29, August 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000684.

Bodeveix:2015:TVT

[1906] Jean-Paul Bodeveix, Mamoun Filali, Manuel Garnacho, Régis Spadotti, and Zhibin Yang. Towards a verified transformation from AADL to the formal component-based language FI-ACRE. Science of Computer Programming, 106(??):30–53, August 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642315000647.

Besnard:2015:TBM

[1907] Loïc Besnard, Adnan Bouakaz, Thierry Gautier, Paul Le Guernic, Yue Ma, Jean-Pierre Talpin, and Huafeng Yu. Timed behavioural modelling and affine scheduling of embedded software architectures in the AADL using Polychrony. Science of Computer Programming, 106(??):

54–77, August 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231400269X.

Mallet:2015:CIM

[1908] Frédéric Mallet and Robert de Simone. Correctness issues on MARTE/ CCSL constraints. Science of Computer Programming, 106(??):78– 92, August 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000519.

Gheyi:2015:SPB

[1909] Rohit Gheyi and Juliano Iyoda.

Selected papers from the Brazilian Symposiums on Formal Methods (SBMF 2012 and 2013). Science of Computer Programming, 107– 108(??):1, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001057.

Lowe:2015:AGT

[1910] Michael Löwe, Harald König, Christoph Schulz, and Marius Schultchen. Algebraic graph transformations with inheritance and abstraction. Science of Computer Programming, 107– 108:2–18, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000295.

Gladisch:2015:SLD

[1911] Christoph Gladisch and Shmuel Tyszberowicz. Specifying linked data structures in JML for combining formal verification and testing. Science of Computer Programming, 107– 108:19–40, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000398.

Calegari:2015:IBF

[1912] Daniel Calegari and Nora Szasz. Institution-based foundations for verification in the context of modeldriven engineering. Science of Computer Programming, 107–108: 41–63, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000507.

dosSantos:2015:RLS

[1913] Joel dos Santos, Christiano Braga, and Débora C. Muchaluat-Saade. A rewriting logic semantics for NCL. Science of Computer Programming, 107– 108:64–92, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000799.

Carbone:2015:PSIb

[1914] Marco Carbone, Ivan Lanese, Alberto Lluch Lafuente, and Ana Sokolova. Preface for the special issue of Interaction and Concurrency Experience 2013. Science of Computer Programming, 109 (??):1, October 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500115X.

Baranov:2015:OSA

[1915] Eduard Baranov and Simon Bliudze.

Offer semantics: Achieving compositionality, flattening and full expressiveness for the glue operators in BIP. Science of Computer Programming, 109 (??):2–35, October 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001148.

Bartoletti:2015:CW

[1916] Massimo Bartoletti, Julien Lange, Alceste Scalas, and Roberto Zunino. Choreographies in the wild. Science of Computer Programming, 109(??): 36–60, October 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005565.

Bartoletti:2015:VCC

[1917] Massimo Bartoletti, Tiziana Cimoli, Paolo Di Giamberardino, and Roberto Zunino. Vicious circles in contracts and in logic. Science of Computer Programming, 109(??):61– 95, October 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000131.

Baldan:2015:MES

[1918] Paolo Baldan, Filippo Bonchi, Fabio Gadducci, and Giacoma Valentina Monreale. Modular encoding of synchronous and asynchronous interactions using open Petri nets. Science of Computer Programming, 109(??): 96–124, October 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005607.

Fernandez-Marquez:2015:PFC

[1919] Jose Luis Fernandez-Marquez, Giovanna Di Marzo Serugendo, and Gabriella Castelli. Preface to Formal Coordination and Self-organisation: Special section including selected best papers from CM 2012–2013. Science of Computer Programming, 110 (??):1–2, October 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315001446.

Viroli:2015:CMP

[1920] Mirko Viroli, Danilo Pianini, Sara Montagna, Graeme Stevenson, and Franco Zambonelli. A coordination model of pervasive service ecosystems. Science of Computer Programming, 110(??):3–22, October 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642315001264.

Latella:2015:FPF

[1921] Diego Latella, Michele Loreti, and Mieke Massink. On-the-fly PCTL fast mean-field approximated modelchecking for self-organising coordination. Science of Computer Programming, 110(??):23–50, October 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016764231500132X.

Murta:2015:SRA

[1922] Daniel Murta and José Nuno Oliveira.

A study of risk-aware program transformation. Science of Computer Programming, 110(??):51–77, October 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642315000908.

Oliveira:2015:RAS

[1923] Nuno Oliveira and Luís S. Barbosa. Reasoning about software reconfigurations: the behavioural and structural perspectives. Science of Computer Programming, 110(??):78– 103, October 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001252.

Caballero:2015:ZDD

[1924] Rafael Caballero, Enrique Martin-Martin, Adrián Riesco, and Salvador Tamarit. A zoom-declarative debugger for sequential Erlang programs. Science of Computer Programming, 110(??): 104–118, October 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001343.

Schrijvers:2015:PSS

[1925] Tom Schrijvers. Preface for SCP special issue on Principles and Practice of Declarative Programming. Science of Computer Programming, 111 (part 1) (??):1–2, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315001355.

Robbins:2015:TPR

[1926] Ed Robbins, Jacob M. Howe, and Andy King. Theory propagation and reification. Science of Computer Programming, 111 (part 1)(??): 3–22, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002615.

ODonnell:2015:ESF

[1927] John T. O’Donnell. Extensible sparse functional arrays with circuit parallelism. Science of Computer Programming, 111 (part 1)(??):23– 50, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314005747.

Montenegro:2015:SAF

[1928] Manuel Montenegro, Ricardo Peña, and Clara Segura. Shape analysis in a functional language by using regular languages. Science of Computer Programming, 111 (part 1)(??): 51–78, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005759.

Matsuda:2015:BFM

[1929] Kazutaka Matsuda and Meng Wang.

“Bidirectionalization for free” for monomorphic transformations. Science of Computer Programming, 111 (part 1)(??):79–109, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642314003323.

Yamada:2015:UOT

[1930] Akihisa Yamada, Keiichirou Kusakari, and Toshiki Sakabe. A unified ordering for termination proving. Science of Computer Programming, 111 (part 1)(??):110– 134, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314003335.

Pino:2015:EAP

[1931] Luis F. Pino, Filippo Bonchi, and Frank Valencia. Efficient algorithms for program equivalence for confluent concurrent constraint programming. Science of Computer Programming, 111 (part 1)(??):135– 155, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005723.

Caballero:2015:TTE

[1932] Rafael Caballero, Peter J. Stuckey, and Antonio Tenorio-Fornés. Two type extensions for the constraint modeling language MiniZinc. Science of Computer Programming, 111 (part 1)(??): 156–189, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000805.

Mandel:2015:TRF

[1933] Louis Mandel, Cédric Pasteur, and Marc Pouzet. Time refinement in a functional synchronous language. Science of Computer Programming, 111 (part 1)(??):190– 211, November 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001367.

Schneider:2015:SIA

[1934] Steve Schneider and Helen Treharne. Special issue on Automated Verification of Critical Systems (AVoCS 2013). Science of Computer Programming, 111 (part 2)(??):213–, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001987. Spe-

cial Issue on Automated Verification of Critical Systems (AVoCS 2013).

Dongol:2015:IBD

[1935] Brijesh Dongol and John Derrick. Interval-based data refinement: A uniform approach to true concurrency in discrete and real-time systems. Science of Computer Programming, 111 (part 2)(??):214–247, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/

article/pii/S016764231500101X. Spe-

cial Issue on Automated Verification of Critical Systems (AVoCS 2013).

Morb:2015:FST

[1936] Georges Morbé and Christoph Scholl.

Fully symbolic TCTL model checking for complete and incomplete real-time systems. Science of Computer Programming, 111 (part 2)(??):248–276, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001458. Spe-

cial Issue on Automated Verification of Critical Systems (AVoCS 2013).

Daw:2015:CMC

[1937] Zamira Daw and Rance Cleaveland. Comparing model checkers for timed UML activity diagrams. Science of Computer Programming, 111 (part 2)(??):277–299, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001045. Spe-

cial Issue on Automated Verification of Critical Systems (AVoCS 2013).

Smyth:2015:FAP

[1938] Ben Smyth, Mark D. Ryan, and Liqun Chen. Formal analysis of privacy in Direct Anonymous Attestation schemes. Science of Computer Programming, 111 (part 2)(??):300–317, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000702. Spe-

cial Issue on Automated Verification of Critical Systems (AVoCS 2013).

Alkhammash:2015:BTE

[1939] Eman Alkhammash, Michael Butler, Asieh Salehi Fathabadi, and Corina Cîrstea. Building traceable Event-B models from requirements. Science of Computer Programming, 111 (part 2)(??):318–338, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001240. Spe-

cial Issue on Automated Verification of Critical Systems (AVoCS 2013).

Patcas:2015:IRF

[1940] Lucian M. Patcas, Mark Lawford, and Tom Maibaum. Implementability of requirements in the four-variable model. Science of Computer Programming, 111 (part 2)(??):339–362, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001033. Spe-

cial Issue on Automated Verification of Critical Systems (AVoCS 2013).

Shkaravska:2015:PSI

[1941] Olha Shkaravska, Simona Ronchi Della Rocca, and Marko van Eekelen. Preface of the special issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011. Science of Computer Programming, 111 (part 3)(??):363–364, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002506. Spe-

cial Issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011.

Canavese:2015:LCF

[1942] D. Canavese, E. Cesena, R. Ouchary, M. Pedicini, and L. Roversi. Light combinators for finite fields arithmetic. Science of Computer Programming, 111 (part 3)(??):365–394, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000672. Spe-

cial Issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011.

Gaboardi:2015:BSU

[1943] Marco Gaboardi and Romain Péchoux.

On bounding space usage of streams using interpretation analysis. Science of Computer Programming, 111 (part 3)(??):395–425, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001008. Spe-

cial Issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011.

Montenegro:2015:SCAa

[1944] Manuel Montenegro, Ricardo Peña, and Clara Segura. Space consumption analysis by abstract interpretation: Inference of recursive functions. Science of Computer Programming, 111 (part 3)(??):426–457, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314001701. Spe-

cial Issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011.

Montenegro:2015:SCAb

[1945] Manuel Montenegro, Ricardo Peña, and Clara Segura. Space consumption analysis by abstract interpretation: Reductivity properties. Science of Computer Programming, 111 (part 3)(??):458–482, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002445. Spe-

cial Issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011.

Albert:2015:PCC

[1946] Elvira Albert, Puri Arenas, Samir Genaim, and Germán Puebla. A practical comparator of cost functions and its applications. Science of Computer Programming, 111 (part 3)(??):483–504, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231400570X. Spe-

cial Issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011.

Rodrigues:2015:CET

[1947] Vítor Rodrigues, Benny Akesson, Mário Florido, Simão Melo de Sousa, João Pedro Pedroso, and Pedro Vasconcelos. Certifying execution time in multicores. Science of Computer Programming, 111 (part 3)(??):505–534, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500129X. Spe-

cial Issue on Foundational and Practical Aspects of Resource Analysis (FOPARA) 2009 & 2011.

Hojjat:2015:FSE

[1948] Hossein Hojjat, Marjan Sirjani, and Farhad Arbab. Fundamentals of Software Engineering (selected papers of FSEN 2013). Science of Computer Programming, 112 (part 1)(??): 1–2, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002671.

Hartmanns:2015:QAZ

[1949] Arnd Hartmanns and Holger Hermanns. In the quantitative automata zoo. Science of Computer Programming, 112 (part 1)(??):3– 23, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002580.

Cholewa:2015:CNC

[1950] Andrew Cholewa, Santiago Escobar, and José Meseguer. Constrained narrowing for conditional equational theories modulo axioms. Science of Computer Programming, 112 (part 1)(??): 24–57, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001239.

Hatefi:2015:ITB

[1951] Hassan Hatefi and Holger Hermanns. Improving time bounded reachability computations in interactive Markov chains. Science of Computer Programming, 112 (part 1)(??): 58–74, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315000994.

Bartoletti:2015:LPN

[1952] Massimo Bartoletti, Tiziana Cimoli, and G. Michele Pinna. Lending Petri nets. Science of Computer Programming, 112 (part 1)(??):75– 101, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315001021.

deBoer:2015:IPP

[1953] F. S. de Boer, M. M. Bonsangue, and J. C. Rot. It is pointless to point in bounded heaps. Science of Computer Programming, 112 (part 1)(??): 102–118, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315001306.

Chin:2015:SEP

[1954] Wei-Ngan Chin and Jurriaan Hage. Selected and extended papers from Partial Evaluation and Program Manipulation 2014. Science of Computer Programming, 112 (part 2)(??): 119, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500283X.

Kameyama:2015:CIY

[1955] Yukiyoshi Kameyama, Oleg Kiselyov, and Chung chieh Shan. Combinators for impure yet hygienic code generation. Science of Computer Programming, 112 (part 2)(??):120– 144, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500249X.

Leather:2015:TCR

[1956] Sean Leather, Johan Jeuring, Andres Löh, and Bram Schuur. Type-changing rewriting and semantics-preserving transformation. Science of Computer Programming, 112 (part 2)(??):145– 169, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001434.

Adams:2015:OST

[1957] Michael D. Adams, Andrew Farmer, and José Pedro Magalhães. Optimizing SYB traversals is easy! Science of Computer Programming, 112 (part 2)(??):170–193, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642315002622.

Barenghi:2015:PPM

[1958] Alessandro Barenghi, Stefano Crespi Reghizzi, Dino Mandrioli, Federica Panella, and Matteo Pradella. Parallel parsing made practical. Science of Computer Programming, 112 (part 3)(??):195–226, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642315002610.

Yi:2015:CTC

[1959] Jaeheon Yi, Tim Disney, Stephen N. Freund, and Cormac Flanagan. Cooperative types for controlling thread interference in Java. Science of Computer Programming, 112 (part 3)(??): 227–260, November 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001483.

Hammoudi:2015:PSI

[1960] Slimane Hammoudi. Preface for special issue on Model Driven Development: Extended versions of selected papers from MODELSWARD 2014. Science of Computer Programming, 113 (part 1)(??):1–2, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003007.

Koshima:2015:CEE

[1961] Amanuel Alemayehu Koshima and Vincent Englebert. Collaborative editing of EMF/Ecore meta-models and models: Conflict detection, reconciliation, and merging in Di-CoMEF. Science of Computer Programming, 113 (part 1)(??):3– 28, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001380.

Bender:2015:SRM

[1962] Marc Bender, Karen Laurin, Mark Lawford, Vera Pantelic, Alexandre Korobkine, Jeff Ong, Bennett Mackenzie, Monika Bialy, and Steven Postma. Signature required: Making Simulink data flow and interfaces explicit. Science of Computer Programming, 113 (part 1) (??):29–50, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001392.

Schwagerl:2015:GBA

[1963] Felix Schwägerl, Sabrina Uhrig, and Bernhard Westfechtel. A graphbased algorithm for three-way merging of ordered collections in EMF models. Science of Computer Programming, 113 (part 1)(??):51– 81, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000532.

Artho:2015:Pb

[1964] Cyrille Artho and Peter Csaba Ölveczky. Preface. Science of Computer Programming, 113 (part 2)(??): 83–84, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002646.

Sabahi-Kaviani:2015:FSE

[1965] Zeynab Sabahi-Kaviani, Ramtin Khosravi, Peter Csaba Ölveczky, Ehsan Khamespanah, and Marjan Sirjani. Formal semantics and efficient analysis of Timed Rebeca in Real-Time Maude. Science of Computer Programming, 113 (part 2)(??):85– 118, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001379.

Pilbrow:2015:ACN

[1966] Colin Pilbrow and Robi Malik. An algorithm for compositional nonblocking verification using special events. Science of Computer Programming, 113 (part 2)(??):119– 148, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001136.

Pang:2015:FVF

[1967] Linna Pang, Chen-Wei Wang, Mark Lawford, and Alan Wassyng. Formal verification of function blocks applied to IEC 61131-3. Science of Computer Programming, 113 (part 2)(??): 149–190, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002981.

Pearce:2015:DVC

[1968] David J. Pearce and Lindsay Groves. Designing a verifying compiler: Lessons learned from developing Whiley. Science of Computer Programming, 113 (part 2)(??):191–220, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016764231500266X.

Fiadeiro:2015:FAC

[1969] José Luiz Fiadeiro and Zhiming Liu.

Formal Aspects of Component Software (FACS 2013). Science of Computer Programming, 113 (part 3)(??): 221–222, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003366.

Gossler:2015:GFB

[1970] Gregor Gössler and Daniel Le Métayer.

A general framework for blaming in component-based systems. Science of Computer Programming, 113 (part 3)(??):223–235, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642315001331.

Boudjadar:2015:RFC

[1971] Abdeldjalil Boudjadar, Alexandre David, Jin Hyun Kim, Kim G. Larsen, Marius Mikucionis, Ulrik Nyman, and Arne Skou. A reconfigurable framework for compositional schedulability and power analysis of hierarchical scheduling systems with frequency scaling. Science of Computer Programming, 113 (part 3)(??):236– 260, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002828.

Lascu:2015:ADC

[1972] Tudor A. Lascu, Jacopo Mauro, and Gianluigi Zavattaro. Automatic deployment of component-based applications. Science of Computer Programming, 113 (part 3)(??):261– 284, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315001409.

Li:2015:MVC

[1973] Yi Li and Meng Sun. Modeling and verification of component connectors in Coq. Science of Computer Programming, 113 (part 3)(??):285– 301, December 1, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315003536.

vandenBrand:2015:ILT

[1974] Mark van den Brand. Introduction — the LDTA tool challenge. Science of Computer Programming, 114 (??):1–6, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315003184.

Basten:2015:MLI

[1975] Bas Basten, Jeroen van den Bos, Mark Hills, Paul Klint, Arnold Lankamp, Bert Lisser, Atze van der Ploeg, Tijs van der Storm, and Jurgen Vinju. Modular language implementation in Rascal — experience report. Science of Computer Programming, 114(??): 7–19, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003524.

Sloane:2015:OK

[1976] Anthony M. Sloane and Matthew Roberts. Oberon-0 in Kiama. Science of Computer Programming, 114(??): 20–32, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003032.

Kaminski:2015:MSO

[1977] Ted Kaminski and Eric Van Wyk.

A modular specification of Oberon0 using the Silver attribute grammar system. Science of Computer Programming, 114(??):33–44, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003020.

Viera:2015:CCC

[1978] Marcos Viera and S. Doaitse Swierstra. Compositional compiler construction: Oberon0. Science of Computer Programming, 114(??):45– 56, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003019.

Boulytchev:2015:CTD

[1979] Dmitry Boulytchev. Combinators and type-driven transformers in Objective Caml. Science of Computer Programming, 114(??):57–73, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315001422.

Fors:2015:JIO

[1980] Niklas Fors and Görel Hedin. A JastAdd implementation of Oberon-

0. Science of Computer Programming, 114(??):74–84, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764231500026X.

Freudenthal:2015:SDT

[1981] Margus Freudenthal. Simpl DSL toolkit. Science of Computer Programming, 114(??):85–91, December 15, 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642314005590.

Kokash:2015:F

[1982] Natallia Kokash and António Ravara. Foreword. Science of Computer Programming, 115–116(??):1–2, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642315003561.

Johnsen:2015:FMS

[1983] Einar Broch Johnsen, Olaf Owe, Dave Clarke, and Joakim Bjørk. A formal model of service-oriented dynamic object groups. Science of Computer Programming, 115–116 (??):3–22, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314005553.

Merelli:2015:ACC

[1984] Emanuela Merelli, Nicola Paoletti, and Luca Tesei. Adaptability checking in complex systems. Science of Computer Programming, 115–116 (??):23–46, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000659.

Jongmans:2015:PSD

[1985] S.-S. T. Q. Jongmans, D. Clarke, and J. Proença. A procedure for splitting data-aware processes and its application to coordination. Science of Computer Programming, 115–116 (??):47–78, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642314000847.

Jacquet:2015:BRD

[1986] Jean-Marie Jacquet, Isabelle Linden, and Mihail-Octavian Staicu. Blackboard rules: From a declarative reading to its application for coordinating context-aware applications in mobile ad hoc networks. Science of Computer Programming, 115–116 (??):79–99, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001471.

Charalambides:2015:PCS

[1987] Minas Charalambides, Peter Dinges, and Gul Agha. Parameterized, concurrent session types for asynchronous multi-actor interactions. Science of Computer Programming, 115–116(??): 100–126, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315002993.

Markovski:2015:PTA

[1988] Jasen Markovski. A process-theoretic approach to supervisory coordination under partial observation. Science of Computer Programming, 115–116 (??):127–145, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642314002986.

Kokash:2015:SSF

[1989] Natallia Kokash and Javier Cámara. Special section on foundations of coordination languages and software architectures. Science of Computer Programming, 115–116(??): 147–148, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315003597.

Jacquet:2015:IDT

[1990] Jean-Marie Jacquet, Isabelle Linden, and Denis Darquennes. On the introduction of density in tuple-space coordination languages. Science of Computer Programming, 115–116(??):

149–176, ????  2015.  CODEN

SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003147.

Brogi:2015:FAS

[1991] Antonio Brogi and Jacopo Soldani. Finding available services in TOSCA-compliant clouds. Science of Computer Programming, 115–116 (??):177–198, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002634.

Jongmans:2015:GCT

[1992] S.-S. T. Q. Jongmans and F. Arbab.

Global consensus through local synchronization: a formal basis for partially-distributed coordination. Science of Computer Programming, 115– 116(??):199–224, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002609.

Le:2015:TCF

[1993] Thi Thieu Hoa Le, Roberto Passerone, Uli Fahrenberg, and Axel Legay. A tag contract framework for modeling heterogeneous systems. Science of Computer Programming, 115–116 (??):225–246, ???? 2015. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001276.

Abid:2016:FDD

[1994] Rim Abid, Gwen Salaün, and Noel De Palma. Formal design of dynamic reconfiguration protocol for cloud applications. Science of Computer Programming, 117(??):1– 16, February 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315004128.

Damiani:2016:TSC

[1995] Ferruccio Damiani, Mirko Viroli, and Jacob Beal. A type-sound calculus of computational fields. Science of Computer Programming, 117(??): 17–44, February 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003573.

Lang:2016:PSI

[1996] Frédéric Lang and Francesco Flammini. Preface to the special issue on Formal Methods for Industrial Critical Systems (FMICS’2014). Science of Computer Programming, 118 (??):1–2, March 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000149.

Lockefeer:2016:FSV

[1997] Lars Lockefeer, David M. Williams, and Wan Fokkink. Formal specification and verification of TCP extended with the Window Scale Option. Science of Computer Programming, 118 (??):3–23, March 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001835.

Zhang:2016:IFT

[1998] Zhen Zhang, Wendelin Serwe, Jian Wu, Tomohiro Yoneda, Hao Zheng, and Chris Myers. An improved fault-tolerant routing algorithm for a Network-on-Chip derived with formal analysis. Science of Computer Programming, 118(??):24–39, March 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642316000125.

Nguyen:2016:ABM

[1999] Thang Nguyen and Dejan Nicković.

Assertion-based monitoring in practice — checking correctness of an automotive sensor interface. Science of Computer Programming, 118(??): 40–59, March 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003512.

Campbell:2016:RTM

[2000] Brian Campbell and Ian Stark. Randomised testing of a microprocessor model using SMT-solver state generation. Science of Computer Programming, 118(??):60–76, March 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642315003159.

Blazy:2016:ISA

[2001] Sandrine Blazy, David Bühler, and Boris Yakobowski. Improving static analyses of C programs with conditional predicates. Science of Computer Programming, 118(??):77– 95, March 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315004116.

Pereira:2016:SPB

[2002] Fernando Magno Quintão Pereira. Selected papers of the Brazilian Symposium on Programming Languag’s 2014 (SBLP’14). Science of Computer Programming, 119(??): 1–2, April 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316000289.

Figueroa:2016:ECH

[2003] Ismael Figueroa, Nicolas Tabareau, and Éric Tanter. Effect capabilities for Haskell: Taming effect interference in monadic programming. Science of Computer Programming, 119 (??):3–30, April 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315004062.

Santos:2016:ACP

[2004] Adriano Santos, Péricles Alves, Eduardo Figueiredo, and Fabiano Ferrari. Avoiding code pitfalls in Aspect-Oriented Programming. Science of Computer Programming, 119 (??):31–50, April 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315004141.

Amato:2016:EIW

[2005] Gianluca Amato, Francesca Scozzari, Helmut Seidl, Kalmer Apinis, and Vesal Vojdani. Efficiently intertwining widening and narrowing. Science of Computer Programming, 120 (??):1–24, ???? 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315004165.

Belli:2016:MBM

[2006] Fevzi Belli, Christof J. Budnik, Axel Hollmann, Tugkan Tuglular, and W. Eric Wong. Model-based mutation testing-approach and case studies. Science of Computer Programming, 120(??):25–48, ???? 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000137.

Valencia-Garcia:2016:SIK

[2007] Rafael Valencia-García and Giner Alor-Hernández. Special issue on knowledgebased software engineering. Science of Computer Programming, 121 (??):1–2, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000496.

Mejia:2016:RAM

[2008] Jezreel Mejia, Edrisi Muñoz, and Mirna Muñoz. Reinforcing the applicability of multi-model environments for software process improvement using knowledge management. Science of Computer Programming, 121 (??):3–15, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500413X.

Villanueva:2016:STS

[2009] Daniel Villanueva, Israel González-Carrasco, J. L. López-Cuadrado, and Nora Lado. SMORE: Towards a semantic modeling for knowledge representation on social media. Science of Computer Programming, 121 (??):16–33, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001318.

Barra:2016:EKO

[2010] Eduardo Barra and Jorge Morato. Early knowledge organization assisted by aspects. Science of Computer Programming, 121(??): 34–54, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001410.

Gonzalez-Torres:2016:KDS

[2011] Antonio González-Torres, Francisco J. García-Peñalvo, Roberto Therón-Sánchez, and Ricardo Colomo-Palacios. Knowledge discovery in software teams by means of evolutionary visual software analytics. Science of Computer Programming, 121(??): 55–74, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002658.

deGraaf:2016:HOA

[2012] K. A. de Graaf, P. Liang, A. Tang, and H. van Vliet. How organisation of architecture documentation affects architectural knowledge retrieval. Science of Computer Programming, 121 (??):75–99, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315003172.

Ait-Ameur:2016:MED

[2013] Yamine Ait-Ameur and Dominique Méry. Making explicit domain knowledge in formal system development. Science of Computer Programming, 121(??):100–127, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642315004153.

Perovsek:2016:TVP

[2014] Matic Perovsek, Janez Kranjc, Tomaz Erjavec, Bojan Cestnik, and Nada Lavrac. TextFlows: a visual programming platform for text mining and natural language processing. Science of Computer Programming, 121 (??):128–152, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000113.

Alqahtani:2016:TKS

[2015] Sultan S. Alqahtani, Ellis E. Eghan, and Juergen Rilling. Tracing known security vulnerabilities in software repositories — a Semantic Web enabled modeling approach. Science of Computer Programming, 121(??): 153–175, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000253.

Velasco-Elizondo:2016:KRI

[2016] Perla Velasco-Elizondo, Rosario Marín-Piña, Sodel Vazquez-Reyes, Arturo Mora-Soto, and Jezreel Mejia. Knowledge representation and information extraction for analysing architectural patterns. Science of Computer Programming, 121(??):176– 189, June 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000101.

Lahami:2016:SER

[2017] Mariam Lahami, Moez Krichen, and Mohamed Jmaiel. Safe and efficient runtime testing framework applied in dynamic and distributed systems. Science of Computer Programming, 122 (??):1–28, June 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000460.

Pang:2016:SEP

[2018] Jun Pang and Mariëlle Stoelinga. Selected and extended papers from ACM SVT 2014. Science of Computer Programming, 123(??):1, July 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642316000708.

Falcone:2016:RER

[2019] Yliès Falcone, Thierry Jéron, Hervé Marchand, and Srinivas Pinisetty. Runtime enforcement of regular timed properties by suppressing and delaying events. Science of Computer Programming, 123(??): 2–41, July 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231600068X.

Beohar:2016:BBM

[2020] Harsh Beohar, Mahsa Varshosaz, and Mohammad Reza Mousavi. Basic behavioral models for software product lines: Expressiveness and testing pre-orders. Science of Computer Programming, 123(??): 42–60, July 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315001288.

Anonymous:2016:EBa

[2021] Anonymous. Editorial Board. Science of Computer Programming, 123 (??):ifc, July 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000721.

Camarao:2016:ACP

[2022] Carlos Camarão, Lucília Figueiredo, and Rodrigo Ribeiro. Ambiguity and constrained polymorphism. Science of Computer Programming, 124 (??):1–19, August 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316000836.

Scott:2016:SGP

[2023] Elizabeth Scott and Adrian Johnstone. Structuring the GLL parsing algorithm for performance. Science of Computer Programming, 125(??): 1–22, September 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231630003X.

Braga:2016:PSS

[2024] Christiano Braga and Narciso Martí-Oliet. Preface to SCP special issue with extended selected papers from SBMF 2014. Science of Computer Programming, 126(??):1– 2, September 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300326.

Calegari:2016:HVC

[2025] Daniel Calegari, Till Mossakowski, and Nora Szasz. Heterogeneous verification in the context of model driven engineering. Science of Computer Programming, 126(??):3– 30, September 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316000472.

Ciaffaglione:2016:TTC

[2026] Alberto Ciaffaglione. Towards Turing computability via coinduction. Science of Computer Programming, 126(??):

31–51, September 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000484.

Moraes:2016:FTS

[2027] Alan Moraes, Wilkerson L. Andrade, and Patrícia D. L. Machado. A family of test selection criteria for Timed Input-Output Symbolic Transition System models. Science of Computer Programming, 126(??):52– 72, September 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300302.

Neves:2016:PTH

[2028] Renato Neves, Alexandre Madeira, Manuel A. Martins, and Luis S. Barbosa. Proof theory for hybrid(ised) logics. Science of Computer Programming, 126(??):73–93, September 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000691.

Timm:2016:PTV

[2029] Nils Timm and Stefan Gruner. Parameterised three-valued model checking. Science of Computer Programming, 126(??):94–110, September 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642316000265.

Anonymous:2016:EBb

[2030] Anonymous. Editorial Board. Science of Computer Programming, 126 (??):ifc, September 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300454.

Lanese:2016:PSI

[2031] Ivan Lanese and Eric Madelaine. Preface for the special issue of the 11th International Symposium on Formal Aspects of Component Software. Science of Computer Programming, 127 (??):1, October 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300156.

Walther:2016:FCP

[2032] Sven Walther and Heike Wehrheim. On-the-fly construction of provably correct service compositions — templates and proofs. Science of Computer Programming, 127(??):2– 23, October 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300028.

Mokni:2016:FAM

[2033] Abderrahman Mokni, Christelle Urtado, Sylvain Vauttier, Marianne Huchard, and Huaxi Yulin Zhang. A formal approach for managing component-based architecture evolution. Science of Computer Programming, 127(??):24–49, October 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642316000794.

Camara:2016:AIE

[2034] Javier Cámara, Antónia Lopes, David Garlan, and Bradley Schmerl. Adaptation impact and environment models for architecture-based self-adaptive systems. Science of Computer Programming, 127(??):50– 75, October 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315004177.

Soleimanifard:2016:AVP

[2035] Siavash Soleimanifard and Dilian Gurov. Algorithmic verification of procedural programs in the presence of code variability. Science of Computer Programming, 127(??): 76–102, October 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315002592.

Boudjadar:2016:SES

[2036] Abdeldjalil Boudjadar, Alexandre David, Jin Hyun Kim, Kim G. Larsen, Marius Mikucionis, Ulrik Nyman, and Arne Skou. Statistical and exact schedulability analysis of hierarchical scheduling systems. Science of Computer Programming, 127(??):103– 130, October 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300430.

Anonymous:2016:EBc

[2037] Anonymous. Editorial Board. Science of Computer Programming, 127 (??):ifc, October 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300594.

Huisman:2016:PSI

[2038] Marieke Huisman and Jaco van de Pol. Preface of special issue on Automated Verification of Critical Systems (AVoCS’14). Science of Computer Programming, 128(??): 1, October 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300223.

Bonacchi:2016:VPR

[2039] A. Bonacchi, A. Fantechi, S. Bacherini, and M. Tempestini. Validation process for railway interlocking systems. Science of Computer Programming, 128 (??):2–21, October 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300053.

Jafari:2016:PMA

[2040] Ali Jafari, Ehsan Khamespanah, Marjan Sirjani, Holger Hermanns, and Matteo Cimini. PTRebeca: Modeling and analysis of distributed and asynchronous systems. Science of Computer Programming, 128(??):22– 50, October 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000800.

Groote:2016:RSB

[2041] Jan Friso Groote, Remco van der Hofstad, and Matthias Raffelsieper. On the random structure of behavioural transition systems. Science of Computer Programming, 128(??): 51–67, October 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000599.

Rockai:2016:MCC

[2042] P. Rockai, J. Barnat, and L. Brim.

Model checking C++ programs with exceptions. Science of Computer Programming, 128(??):68–85, October 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642316300338.

Anonymous:2016:EBd

[2043] Anonymous. Editorial Board. Science of Computer Programming, 128 (??):ifc, October 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300740.

Sierra:2016:SIE

[2044] José-Luis Sierra. Special issue on eLearning software architectures. Science of Computer Programming, 129 (??):1–2, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300910.

Zhang:2016:FPW

[2045] Bingxue Zhang, Chuantao Yin, Bertrand David, Zhang Xiong, and Wei Niu. Facilitating professionals’ work-based learning with contextaware mobile system. Science of Computer Programming, 129(??):3– 19, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316000459.

Garcia-Holgado:2016:API

[2046] Alicia García-Holgado and Francisco José García-Peñalvo. Architectural pattern to improve the definition and implementation of eLearning ecosystems. Science of Computer Programming, 129(??):20– 34, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300259.

Striewe:2016:AMG

[2047] Michael Striewe. An architecture for modular grading and feedback generation for complex exercises. Science of Computer Programming, 129(??): 35–47, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300260.

Stuikys:2016:MDP

[2048] Vytautas Stuikys, Renata Burbaite, Kristina Bespalova, and Giedrius Ziberkas. Model-driven processes and tools to design robot-based generative learning objects for computer science education. Science of Computer Programming, 129(??):48– 71, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300247.

Perez-Rodriguez:2016:ACB

[2049] Roberto Pérez-Rodríguez, Luis Anido-Rifón, Miguel Gómez-Carballa, and Marcos Mouriño-García. Architecture of a concept-based information retrieval system for educational resources. Science of Computer Programming, 129(??):72– 91, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300314.

Raspopovic:2016:SAI

[2050] Miroslava Raspopović, Svetlana Cvetanović, Dusan Stanojević, and Mateja Opacić. Software architecture for integration of institutional and social learning environments. Science of Computer Programming, 129(??):92– 102, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300855.

Anonymous:2016:EBe

[2051] Anonymous. Editorial Board. Science of Computer Programming, 129 (??):ifc, November 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300934.

Bosnacki:2016:SSG

[2052] Dragan Bosnacki, Stefan Edelkamp, Alberto Lluch Lafuente, and Anton Wijs. Special section on Graph Inspection and Traversal Engineering (GRAPHITE 2014). Science of Computer Programming, 130(??): 1, November 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300843.

Mansky:2016:SEO

[2053] William Mansky, Elsa L. Gunter, Dennis Griffith, and Michael D. Adams. Specifying and executing optimizations for generalized control flow graphs. Science of Computer Programming, 130 (??):2–23, November 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300703.

Blanco:2016:UGD

[2054] Roi Blanco, Paolo Boldi, and Andrea Marino. Using graph distances for named-entity linking. Science of Computer Programming, 130(??):24– 36, November 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315003160.

Delen:2016:LDI

[2055] G. P. A. J. Delen, R. J. Peters, C. Verhoef, and S. F. M. van Vlijmen. Lessons from Dutch IToutsourcing success and failure. Science of Computer Programming, 130(??): 37–68, November 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300016.

Montesi:2016:PAW

[2056] Fabrizio Montesi. Process-aware web programming with Jolie. Science of Computer Programming, 130(??):69– 96, November 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300284.

Terven:2016:KKT

[2057] Juan R. Terven and Diana M. Córdova-Esparza. Kin2. A Kinect 2 toolbox for MATLAB. Science of Computer Programming, 130(??):97– 106, November 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300569.

Ait-Ameur:2016:IAS

[2058] Yamine Ait-Ameur and Klaus-Dieter Schewe. Introduction to the ABZ 2014 special issue. Science of Computer Programming, 131(??):1– 2, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301083.

Ernst:2016:MCS

[2059] Gidon Ernst, Jörg Pfähler, Gerhard Schellhorn, and Wolfgang Reif.

Modular, crash-safe refinement for ASMs with submachines. Science of Computer Programming, 131(??): 3–21, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300181.

Dausend:2016:TCE

[2060] Marcel Dausend and Alexander Raschke. Towards a comprehensive extension of abstract state machines for aspectoriented specification. Science of Computer Programming, 131(??):22– 41, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300168.

Borger:2016:SML

[2061] Egon Börger, Klaus-Dieter Schewe, and Qing Wang. Serialisable multilevel transaction control: a specification and verification. Science of Computer Programming, 131(??):42– 58, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300041.

Furst:2016:LSS

[2062] Andreas Fürst, Thai Son Hoang, David Basin, Naoto Sato, and Kunihiko Miyazaki. Large-scale system development using Abstract Data Types and refinement. Science of Computer Programming, 131(??):59– 75, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300193.

Huynh:2016:FVR

[2063] Nghi Huynh, Marc Frappier, Amel Mammar, Régine Laleau, and Jules Desharnais. A formal validation of the RBAC ANSI 2012 standard using B. Science of Computer Programming, 131(??):76– 93, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231630020X.

Hallerstede:2016:CEB

[2064] Stefan Hallerstede. The correctness of event-B inductive convergence. Science of Computer Programming, 131(??): 94–108, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300211.

Hansen:2016:TBT

[2065] Dominik Hansen and Michael Leuschel.

Translating B to TLA^{+} for validation with TLC. Science of Computer Programming, 131(??):109– 125, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300235.

Li:2016:FDM

[2066] Qin Li and Graeme Smith. Formal development of multi-agent systems using MAZE. Science of Computer Programming, 131(??):126– 150, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231630017X.

Anonymous:2016:EBf

[2067] Anonymous. Editorial Board. Science of Computer Programming, 131 (??):ifc, December 1, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301368.

Bois:2016:SEP

[2068] André Rauber Du Bois and Phil Trinder. Selected and extended papers from SBLP 2013. Science of Computer Programming, 132 (part 1) (??):1, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301071.

Martins:2016:EAG

[2069] Pedro Martins, João Paulo Fernandes, João Saraiva, Eric Van Wyk, and Anthony Sloane. Embedding attribute grammars and their extensions using functional zippers. Science of Computer Programming, 132 (part 1)(??): 2–28, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316000812.

Junior:2016:PEV

[2070] Francisco Heron de Carvalho Junior and Cenez Araújo Rezende. Performance evaluation of virtual execution environments for intensive computing on usual representations of multidimensional arrays. Science of Computer Programming, 132 (part 1)(??): 29–49, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300065.

Madeira:2016:MRD

[2071] Alexandre Madeira, Renato Neves, Luís S. Barbosa, and Manuel A. Martins. A method for rigorous design of reconfigurable systems. Science of Computer Programming, 132 (part 1)(??):50–76, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300272.

Pardo:2016:MIS

[2072] Alberto Pardo, João Paulo Fernandes, and João Saraiva. Multiple intermediate structure deforestation by shortcut fusion. Science of Computer Programming, 132 (part 1)(??): 77–95, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300880.

Junior:2016:CAT

[2073] Francisco Heron de Carvalho Junior, Cenez Araújo Rezende, Jefferson de Carvalho Silva, Wagner Guimarães Al-Alam, and João Marcelo Uchoa de Alencar. Contextual abstraction in a type system for componentbased high performance computing platforms. Science of Computer Programming, 132 (part 1)(??):96– 128, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300892.

Maidl:2016:ERP

[2074] André Murbach Maidl, Fabio Mascarenhas, Sérgio Medeiros, and Roberto Ierusalimschy. Error reporting in Parsing Expression Grammars. Science of Computer Programming, 132 (part 1)(??):129– 140, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301046.

Salaun:2016:SIS

[2075] Gwen Salaün and Mariëlle Stoelinga. Special issue on Software Verification and Testing (SAC-SVT’15). Science of Computer Programming, 132 (part 2)(??):141–142, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642316301484.

Poncelet:2016:MBT

[2076] Clement Poncelet and Florent Jacquemard. Model-based testing for building reliable realtime interactive music systems. Science of Computer Programming, 132 (part 2)(??):143– 172, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301022.

Homm:2016:CCE

[2077] Daniel Homm, Jürgen Eckert, and Reinhard German. CenUMs — concurrency enhanced usage models for statistical testing of complex systems with concurrent streams of use. Science of Computer Programming, 132 (part 2)(??):173– 189, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300867.

Bride:2016:TMU

[2078] Hadrien Bride, Jacques Julliand, and Pierre-Alain Masson. Tri-modal under-approximation for test generation. Science of Computer Programming, 132 (part 2)(??):190– 208, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300879.

Bourdil:2016:SRT

[2079] Pierre-Alain Bourdil, Bernard Berthomieu, Silvano Dal Zilio, and François Vernadat. Symmetry reduction for time Petri net state classes. Science of Computer Programming, 132 (part 2)(??): 209–225, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301095.

Jakobsson:2016:FSE

[2080] Arvid Jakobsson, Nikolai Kosmatov, and Julien Signoles. Fast as a shadow, expressive as a tree: Optimized memory monitoring for C. Science of Computer Programming, 132 (part 2)(??): 226–246, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301344.

Anonymous:2016:EBg

[2081] Anonymous. Editorial Board. Science of Computer Programming, 132 (part 1)(??):ifc, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301514.

Anonymous:2016:EBh

[2082] Anonymous. Editorial Board. Science of Computer Programming, 132 (part 2)(??):ifc, December 15, 2016. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301629.

Campusano:2017:LRP

[2083] Miguel Campusano and Johan Fabry. Live Robot Programming: the language, its implementation, and robot API independence. Science of Computer Programming, 133 (part 1)(??): 1–19, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316300697.

Bodei:2017:CGU

[2084] Chiara Bodei, Viet Dung Dinh, and Gian-Luigi Ferrari. Checking global usage of resources handled with local policies. Science of Computer Programming, 133 (part 1)(??):20– 50, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300727.

Ernst:2017:MSD

[2085] Erik Ernst, Anders Møller, Mathias Schwarz, and Fabio Strocco. Message safety in Dart. Science of Computer Programming, 133 (part 1)(??): 51–73, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300831.

Bygde:2017:IPP

[2086] Stefan Bygde, Björn Lisper, and Niklas Holsti. Improved precision in polyhedral analysis with wrapping. Science of Computer Programming, 133 (part 1) (??):74–87, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300909.

Anonymous:2017:EBa

[2087] Anonymous. Editorial Board. Science of Computer Programming, 133 (part 1)(??):ifc, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301848.

Artho:2017:FTS

[2088] Cyrille Artho and Peter Csaba Ölveczky. Formal Techniques for Safety-Critical Systems (FTSCS 2014). Science of Computer Programming, 133 (part 2)(??):89–90, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642316301496.

Vu:2017:FMV

[2089] Linh Hong Vu, Anne E. Haxthausen, and Jan Peleska. Formal modelling and verification of interlocking systems featuring sequential release. Science of Computer Programming, 133 (part 2) (??):91–115, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300570.

Benerecetti:2017:DSM

[2090] M. Benerecetti, R. De Guglielmo, U. Gentile, S. Marrone, N. Mazzocca, R. Nardone, A. Peron, L. Velardi, and V. Vittorini. Dynamic state machines for modelling railway control systems. Science of Computer Programming, 133 (part 2)(??): 116–153, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301332.

Bukhari:2017:FTF

[2091] Syed Ali Asadullah Bukhari, Faiq Khalid Lodhi, Osman Hasan, Muhammad Shafique, and Jörg Henkel. FAMe-TM: Formal analysis methodology for task migration algorithms in Many-Core systems. Science of Computer Programming, 133 (part 2)(??):154– 174, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300715.

Kelter:2017:PAP

[2092] Timon Kelter and Peter Marwedel.

Parallelism analysis: Precise WCET values for complex multi-core systems. Science of Computer Programming, 133 (part 2)(??):175– 193, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000277.

Caltais:2017:VSP

[2093] Georgiana Caltais and Bertrand Meyer.

On the verification of SCOOP programs. Science of Computer Programming, 133 (part 2)(??):194– 215, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301058.

Seceleanu:2017:AWT

[2094] Cristina Seceleanu, Morgan Johansson, Jagadish Suryadevara, Gaetana Sapienza, Tiberiu Seceleanu, Stein-Erik Ellevseth, and Paul Pettersson. Analyzing a wind turbine system: From simulation to formal verification. Science of Computer Programming, 133 (part 2)(??):216– 242, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301472.

Anonymous:2017:EBb

[2095] Anonymous. Editorial Board. Science of Computer Programming, 133 (part 2)(??):ifc, January 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301903.

Anonymous:2017:EBc

[2096] Anonymous. Editorial Board. Science of Computer Programming, 133 (part 2)(??):ifc, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301903.

vandenBrand:2017:GEI

[2097] Mark G. J. van den Brand, Jurgen J. Vinju, and Kim Mens. Guest Editors’ introduction to the 6th issue of Experimental Software and Toolkits (EST-6). Science of Computer Programming, 134(??):1–3, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301885.

Felgentreff:2017:IRR

[2098] T. Felgentreff, M. Perscheid, and R. Hirschfeld. Implementing record and refinement for debugging timingdependent communication. Science of Computer Programming, 134 (??):4–18, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315003585.

Ganea:2017:CQA

[2099] George Ganea, Ioana Verebi, and Radu Marinescu. Continuous quality assessment with inCode. Science of Computer Programming, 134 (??):19–36, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642315000520.

Hills:2017:EPS

[2100] Mark Hills, Paul Klint, and Jurgen J.

Vinju. Enabling PHP software engineering research in Rascal. Science of Computer Programming, 134 (??):37–46, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300296.

Storrle:2017:CEE

[2101] Harald Störrle. Cost-effective evolution of research prototypes into enduser tools: the MACH case study. Science of Computer Programming, 134(??):47–60, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231500146X.

Martinez-Fernandez:2017:MUQ

[2102] S. Martínez-Fernández, X. Franch, and J. Bisbal. Mercury: Using the QuPreSS reference model to evaluate predictive services. Science of Computer Programming, 134 (??):61–74, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315004050.

Castro:2017:JLC

[2103] Sergio Castro, Kim Mens, and Paulo Moura. JPC: a library for categorising and applying inter-language conversions between Java and Prolog. Science of Computer Programming, 134 (??):75–99, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642315004049.

Bagge:2017:MFR

[2104] Anya Helene Bagge. Managing facts and resources with the Pica IDE infrastructure library. Science of Computer Programming, 134(??): 100–111, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301447.

Anonymous:2017:EBd

[2105] Anonymous. Editorial Board. Science of Computer Programming, 134 (??):ifc, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316302118.

Staron:2017:PSI

[2106] Miroslaw Staron, Wilhelm Meding, Alain Abran, and Jan Bosch. Preface to the special issue on advances in software measurement. Science of Computer Programming, 135(??): 1–3, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316302222.

Kirbas:2017:ECM

[2107] Serkan Kirbas, Tracy Hall, and Alper Sen. Evolutionary coupling measurement: Making sense of the current chaos. Science of Computer Programming, 135(??):4– 19, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301587.

Fonseca:2017:OBA

[2108] Vinícius Soares Fonseca, Monalessa Perini Barcellos, and Ricardo de Almeida Falbo. An ontology-based approach for integrating tools supporting the software measurement process. Science of Computer Programming, 135(??):20– 44, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301599.

Hebig:2017:TQT

[2109] Regina Hebig and Haoyu Wang. On tackling quality threats for the assessment of measurement programs: a case study on the distribution of metric usage and knowledge. Science of Computer Programming, 135(??): 45–74, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301460.

Tsunoda:2017:BIO

[2110] Masateru Tsunoda, Akito Monden, Kenichi Matsumoto, Sawako Ohiwa, and Tomoki Oshino. Benchmarking IT operations cost based on working time and unit cost. Science of Computer Programming, 135(??):75– 87, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231630171X.

Haoues:2017:FCI

[2111] Mariem Haoues, Asma Sellami, and Hanêne Ben-Abdallah. Functional change impact analysis in use cases: an approach based on COSMIC functional size measurement. Science of Computer Programming, 135(??):88– 104, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316301459.

Valdes-Souto:2017:APT

[2112] Francisco Valdés-Souto. Analyzing the performance of two COSMIC approximation sizing techniques at the functional process level. Science of Computer Programming, 135(??):105– 121, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316302192.

Anonymous:2017:EBe

[2113] Anonymous. Editorial Board. Science of Computer Programming, 135 (??):ifc, February 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316302246.

Martinez:2017:RMT

[2114] Salvador Martínez, Massimo Tisi, and Rémi Douence. Reactive model transformation with ATL. Science of Computer Programming, 136 (??):1–16, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231630106X.

Boudaa:2017:AOM

[2115] Boudjemaa Boudaa, Slimane Hammoudi, Leila Amel Mebarki, Abdelkader Bouguessa, and Mohammed Amine Chikh. An aspect-oriented modeldriven approach for building adaptable context-aware service-based applications. Science of Computer Programming, 136(??):17–42, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642316301253.

Elias:2017:SAS

[2116] Endhe Elias, Jário Santos, Ig Ibert Bittencourt, Seiji Isotani, Olavo Holanda, and Patrick H. S. Brito. A semi-automatic system to evaluate the performance and scalability of ontology persistent APIs. Science of Computer Programming, 136 (??):43–59, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301605.

Anonymous:2017:EBf

[2117] Anonymous. Editorial Board. Science of Computer Programming, 136 (??):ifc, ???? 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300138.

Asai:2017:SEP

[2118] Kenichi Asai and Konstantinos Sagonas. Selected and extended papers from Partial Evaluation and Program Manipulation 2015 (PEPM’15). Science of Computer Programming, 137(??):1, April 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300230.

Asada:2017:VRP

[2119] Kazuyuki Asada, Ryosuke Sato, and Naoki Kobayashi. Verifying relational properties of functional programs by first-order refinement. Science of Computer Programming, 137 (??):2–62, April 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642316000678.

Bahr:2017:GTT

[2120] Patrick Bahr and Emil Axelsson. Generalising tree traversals and tree transformations to DAGs: Exploiting sharing without the pain. Science of Computer Programming, 137 (??):63–97, April 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316000824.

Bransen:2017:IEH

[2121] Jeroen Bransen, Atze Dijkstra, and S. Doaitse Swierstra. Incremental evaluation of higher-order attributes. Science of Computer Programming, 137 (??):98–124, April 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316300685.

Kafle:2017:CSH

[2122] Bishoksan Kafle and John P. Gallagher. Constraint specialisation in Horn clause verification. Science of Computer Programming, 137(??): 125–140, April 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300205.

Anonymous:2017:EBg

[2123] Anonymous. Editorial Board. Science of Computer Programming, 137 (??):ifc, April 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300266.

Muller:2017:DCB

[2124] Richard Müller, Christian Stahl, and Walter Vogler. Deciding conformance for bounded responsiveness. Science of Computer Programming, 138 (??):1–25, May 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316302179.

Desharnais:2017:PPS

[2125] Jules Desharnais, Nafi Diallo, Wided Ghardallou, and Ali Mili. Projecting programs on specifications: Definition and implications. Science of Computer Programming, 138 (??):26–48, May 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300217.

Anonymous:2017:EBh

[2126] Anonymous. Editorial Board. Science of Computer Programming, 138 (??):ifc, May 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300370.

Banach:2017:CHE

[2127] Richard Banach, Michael Butler, Shengchao Qin, and Huibiao Zhu. Core Hybrid Event-B II: Multiple cooperating Hybrid Event-B machines. Science of Computer Programming, 139 (??):1–35, June 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300229.

Fichera:2017:PFP

[2128] Loris Fichera, Fabrizio Messina, Giuseppe Pappalardo, and Corrado Santoro. A Python framework for programming autonomous robots using a declarative approach. Science of Computer Programming, 139 (??):36–55, June 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300242.

Anonymous:2017:EBi

[2129] Anonymous. Editorial Board. Science of Computer Programming, 139 (??):ifc, June 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300448.

Ancona:2017:PSS

[2130] Davide Ancona. Preface to the special section on Object-Oriented Programming and Systems (OOPS 2015). Science of Computer Programming, 140(??):1, June 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300539.

Papoulias:2017:EUA

[2131] N. Papoulias, M. Denker, S. Ducasse, and L. Fabresse. End-user abstractions for meta-control: Reifying the reflectogram. Science of Computer Programming, 140(??):2– 16, June 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300011.

Pape:2017:AJT

[2132] Tobias Pape, Carl Friedrich Bolz, and Robert Hirschfeld. Adaptive just-in-time value class optimization for lowering memory consumption and improving execution time performance. Science of Computer Programming, 140 (??):17–29, June 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316301034.

Anonymous:2017:EBj

[2133] Anonymous. Editorial Board. Science of Computer Programming, 140 (??):ifc, June 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300588.

Fritz:2017:TSA

[2134] Eric Fritz and Tian Zhao. Typing and semantics of asynchronous arrows in JavaScript. Science of Computer Programming, 141–142(??): 1–39, July/August 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300527.

Cordova-Esparza:2017:MCC

[2135] Diana-Margarita Córdova-Esparza, Juan R. Terven, Hugo Jiménez-Hernández, and Ana-Marcela Herrera-Navarro. A multiple camera calibration and point cloud fusion tool for Kinect V2. Science of Computer Programming, 143(??):1– 8, September 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642316302180.

Fachada:2017:CCF

[2136] Nuno Fachada, Vitor V. Lopes, Rui C. Martins, and Agostinho C. Rosa. cf4ocl: a C framework for OpenCL. Science of Computer Programming, 143(??):9– 19, September 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300540.

Anonymous:2017:EBk

[2137] Anonymous. Editorial Board. Science of Computer Programming, 143 (??):ifc, September 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300874.

Codescu:2017:SRC

[2138] Mihai Codescu, Till Mossakowski, Donald Sannella, and Andrzej Tarlecki. Specification refinements: Calculi, tools, and applications. Science of Computer Programming, 144(??): 1–49, September 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300801.

Anonymous:2017:EBl

[2139] Anonymous. Editorial Board. Science of Computer Programming, 144 (??):ifc, September 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300989.

Sohrabi:2017:PSS

[2140] Mohammad Karim Sohrabi and Hosseion Azgomi. Parallel set similarity join on big data based on Locality-Sensitive Hashing. Science of Computer Programming, 145(??): 1–12, October 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300813.

Luo:2017:ISB

[2141] Chen Luo, Fei He, and Carlo Ghezzi. Inferring software behavioral models with MapReduce. Science of Computer Programming, 145(??):13– 36, October 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300795.

Anonymous:2017:EBm

[2142] Anonymous. Editorial Board. Science of Computer Programming, 145 (??):ifc, October 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301107.

Braga:2017:PSI

[2143] Christiano Braga. Preface to Special issue with extended selected papers from FACS 2015. Science of Computer Programming, 146(??): 1, October 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301247.

Westman:2017:FAM

[2144] Jonas Westman, Mattias Nyberg, Joakim Gustavsson, and Dilian Gurov. Formal architecture modeling of sequential non-recursive C programs. Science of Computer Programming, 146(??):2–27, October 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642317300564.

Proenca:2017:TCF

[2145] José Proença and Dave Clarke. Typed connector families and their semantics. Science of Computer Programming, 146(??):28–49, October 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642317300515.

Jongmans:2017:CAM

[2146] S.-S. T. Q. Jongmans, T. Kappé, and F. Arbab. Constraint automata with memory cells and their composition. Science of Computer Programming, 146(??):50– 86, October 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300552.

Abid:2017:AST

[2147] Rim Abid, Gwen Salaün, and Noel De Palma. Asynchronous synthesis techniques for coordinating autonomic managers in the cloud. Science of Computer Programming, 146(??): 87–103, October 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301089.

Anonymous:2017:EBn

[2148] Anonymous. Editorial Board. Science of Computer Programming, 146 (??):ifc, October 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301351.

Albert:2017:PSE

[2149] Elvira Albert. Preface for selected and extended papers from Principles and Practice of Declarative Programming (PPDP’15). Science of Computer Programming, 147(??):1– 2, November 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301296.

Schmidt-Schauss:2017:ICN

[2150] Manfred Schmidt-Schauß and David Sabel. Improvements in a call-by-need functional core language: Common subexpression elimination and resource preserving translations. Science of Computer Programming, 147 (??):3–26, November 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300199.

Garcia:2017:SAC

[2151] Abel Garcia, Cosimo Laneve, and Michael Lienhardt. Static analysis of cloud elasticity. Science of Computer Programming, 147(??):27– 53, November 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300679.

Bartak:2017:MSP

[2152] Roman Barták, Lukás Chrpa, Agostino Dovier, Jindrich Vodrázka, and Neng-Fa Zhou. Modeling and solving planning problems in tabled logic programming: Experience from the Cave Diving domain. Science of Computer Programming, 147(??):54– 77, November 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300825.

DeAngelis:2017:SBG

[2153] E. De Angelis, F. Fioravanti, A. Pettorossi, and M. Proietti. Semanticsbased generation of verification conditions via program specialization. Science of Computer Programming, 147(??):78–108, November 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016764231630199X.

Giantsios:2017:CTF

[2154] Aggelos Giantsios, Nikolaos Papaspyrou, and Konstantinos Sagonas. Concolic testing for functional languages. Science of Computer Programming, 147(??):109– 134, November 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300837.

Anonymous:2017:EBo

[2155] Anonymous. Editorial Board. Science of Computer Programming, 147 (??):ifc, November 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231730148X.

Grov:2017:PSI

[2156] Gudmund Grov and Andrew Ireland. Preface of the special issue for AVoCS 2015. Science of Computer Programming, 148(??):1– 2, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301284.

OHalloran:2017:VCA

[2157] Colin O’Halloran, Tom Gibson Robinson, and Neil Brock. Verifying cyber attack properties. Science of Computer Programming, 148(??):3– 25, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301272.

Rahli:2017:ESV

[2158] Vincent Rahli, David Guaspari, Mark Bickford, and Robert L. Constable. EventML: Specification, verification, and implementation of crash-tolerant state machine replication systems. Science of Computer Programming, 148(??):26–48, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301193.

Dalvandi:2017:DAC

[2159] Mohammadsadegh Dalvandi, Michael Butler, and Abdolbaghi Rezazadeh. Derivation of algorithmic control structures in Event-B refinement. Science of Computer Programming, 148(??): 49–65, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231730120X.

Laibinis:2017:FRA

[2160] L. Laibinis, I. Pereverzeva, and E. Troubitsyna. Formal reasoning about resilient goal-oriented multiagent systems. Science of Computer Programming, 148(??):66–87, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301181.

Kamali:2017:FVA

[2161] Maryam Kamali, Louise A. Dennis, Owen McAree, Michael Fisher, and Sandor M. Veres. Formal verification of autonomous vehicle platooning. Science of Computer Programming, 148(??):88–106, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301168.

Ketema:2017:TAG

[2162] Jeroen Ketema and Alastair F. Donaldson. Termination analysis for GPU kernels. Science of Computer Programming, 148(??):107–122, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300849.

Althaus:2017:VLH

[2163] Ernst Althaus, Björn Beber, Werner Damm, Stefan Disch, Willem Hagemann, Astrid Rakow, Christoph Scholl, Uwe Waldmann, and Boris Wirtz. Verification of linear hybrid systems with large discrete state spaces using counterexample-guided abstraction refinement. Science of Computer Programming, 148(??):123– 160, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300850.

Amato:2017:ILI

[2164] Gianluca Amato, Marco Rubino, and Francesca Scozzari. Inferring linear invariants with parallelotopes. Science of Computer Programming, 148(??): 161–188, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301211.

Qin:2017:ASI

[2165] Shengchao Qin, Guanhua He, Wei-Ngan Chin, Florin Craciun, Mengda He, and Zhong Ming. Automated specification inference in a combined domain via user-defined predicates. Science of Computer Programming, 148(??):189– 212, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231730117X.

Anonymous:2017:EBp

[2166] Anonymous. Editorial Board. Science of Computer Programming, 148 (??):ifc, November 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301818.

Clarke:2017:TDT

[2167] Peter J. Clarke, Juan de Lara, and Mark van den Brand. Tool demonstration track of MODELS’2016. Science of Computer Programming, 149 (??):1–2, December 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301764.

Hammad:2017:IIT

[2168] Muhammad Hammad, Tao Yue, Shuai Wang, Shaukat Ali, and Jan F. Nygård. IOCL: an interactive tool for specifying, validating and evaluating OCL constraints. Science of Computer Programming, 149(??): 3–8, December 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301466.

Daniel:2017:NMD

[2169] Gwendal Daniel, Gerson Sunyé, Amine Benelallam, Massimo Tisi, Yoann Vernageau, Abel Gómez, and Jordi Cabot. NeoEMF: a multidatabase model persistence framework for very large models. Science of Computer Programming, 149(??): 9–14, December 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301600.

Greenyer:2017:STS

[2170] Joel Greenyer, Daniel Gritzner, Timo Gutjahr, Florian König, Nils Glade, Assaf Marron, and Guy Katz. ScenarioTools — a tool suite for the scenario-based modeling and analysis of reactive systems. Science of Computer Programming, 149(??):15– 27, December 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301442.

Anonymous:2017:EBq

[2171] Anonymous. Editorial Board. Science of Computer Programming, 149 (??):ifc, December 1, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301983.

Lin:2017:EEG

[2172] Chu-Ti Lin, Kai-Wei Tang, Jiun-Shiang Wang, and Gregory M. Kapfhammer. Empirically evaluating greedy-based test suite reduction methods at different levels of test suite complexity. Science of Computer Programming, 150(??):1– 25, December 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301077.

Tommasel:2017:SJL

[2173] Antonela Tommasel, Daniela Godoy, and Alejandro Zunino. SMArtOp: a Java library for distributing high-dimensional sparse-matrix arithmetic operations. Science of Computer Programming, 150(??):26– 30, December 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301260.

Bacchelli:2017:MSD

[2174] Alberto Bacchelli, Andrea Mocci, Anthony Cleve, and Michele Lanza. Mining structured data in natural language artifacts with island parsing. Science of Computer Programming, 150(??): 31–55, December 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301302.

Khatchadourian:2017:DBP

[2175] Raffi Khatchadourian, Awais Rashid, Hidehiko Masuhara, and Takuya Watanabe. Detecting broken pointcuts using structural commonality and degree of interest. Science of Computer Programming, 150(??):56– 74, December 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301326.

vandenHove:2017:DHC

[2176] Gauthier van den Hove. Dissolving a half century old problem about the implementation of procedures. Science of Computer Programming, 150(??): 75–86, December 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301582. See

[2879].

Anonymous:2017:EBr

[2177] Anonymous. Editorial Board. Science of Computer Programming, 150 (??):ifc, December 15, 2017. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302095.

Lanese:2018:PSI

[2178] Ivan Lanese and Simon Devitt. Preface for the special issue of the 8th Conference on Reversible Computation (RC 2016). Science of Computer Programming, 151(??): 1, January 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302277.

Selinger:2018:FAR

[2179] Peter Selinger. A finite alternation result for reversible boolean circuits. Science of Computer Programming, 151 (??):2–17, January 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301703.

Kuhn:2018:LRC

[2180] Stefan Kuhn and Irek Ulidowski.

Local reversibility in a Calculus of Covalent Bonding. Science of Computer Programming, 151(??):18– 47, January 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231730196X.

Barylska:2018:RCV

[2181] Kamila Barylska, Maciej Koutny, Lukasz Mikulski, and Marcin Piatkowski. Reversible computation vs. reversibility in Petri nets. Science of Computer Programming, 151(??): 48–60, January 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302204.

Anonymous:2018:EBa

[2182] Anonymous. Editorial Board. Science of Computer Programming, 151 (??):ifc, January 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302307.

Anonymous:2018:PNA

[2183] Anonymous. Publisher’s note: Article based publishing. Science of Computer Programming, 152(??): iii, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231730254X.

Anonymous:2018:PN

[2184] Anonymous. Publisher’s note. Science of Computer Programming, 152 (??):iv, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302575.

Dupree:2018:CSU

[2185] Janna-Lynn (Weber) Dupree, Edward Lank, and Daniel M. Berry. A case study of using grounded analysis as a requirement engineering method: Identifying personas that specify privacy and security tool users. Science of Computer Programming, 152 (??):1–37, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301697.

Radke:2018:TEO

[2186] Hendrik Radke, Thorsten Arendt, Jan Steffen Becker, Annegret Habel, and Gabriele Taentzer. Translating essential OCL invariants to nested graph constraints for generating instances of meta-models. Science of Computer Programming, 152 (??):38–62, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231730165X.

Monteiro:2018:EGC

[2187] Felipe R. Monteiro, Erickson H.

da S. Alves, Isabela S. Silva, Hussama I. Ismail, Lucas C. Cordeiro, and Eddie B. de Lima Filho. ESBMC-GPU: a context-bounded model checking tool to verify CUDA programs. Science of Computer Programming, 152 (??):63–69, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301934.

Rivas:2018:UVM

[2188] Exequiel Rivas, Mauro Jaskelioff, and Tom Schrijvers. A unified view of monadic and applicative non-determinism. Science of Computer Programming, 152(??):70– 98, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301958.

Durelli:2018:ECE

[2189] Vinicius H. S. Durelli, Marcio E. Delamaro, and Jeff Offutt. An experimental comparison of edge, edgepair, and prime path criteria. Science of Computer Programming, 152(??): 99–115, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302150.

Castro:2018:ARL

[2190] Thiago Castro, André Lanna, Vander Alves, Leopoldo Teixeira, Sven Apel, and Pierre-Yves Schobbens.

All roads lead to Rome: Commuting strategies for product-line reliability analysis. Science of Computer Programming, 152(??):116– 160, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302253.

Maalej:2018:CRI

[2191] Maroua Maalej, Vitor Paisante, Fernando Magno Quintão Pereira, and Laure Gonnord. Combining range and inequality information for pointer disambiguation. Science of Computer Programming, 152(??):161– 184, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302265.

Anonymous:2018:EBb

[2192] Anonymous. Editorial Board. Science of Computer Programming, 152 (??):ifc, January 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302472.

Anonymous:2018:EBc

[2193] Anonymous. Editorial Board. Science of Computer Programming, 153 (??):ii, February 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231730299X.

Khamespanah:2018:ETM

[2194] Ehsan Khamespanah, Ramtin Khosravi, and Marjan Sirjani. An efficient TCTL model checking algorithm and a reduction technique for verification of timed actor models. Science of Computer Programming, 153 (??):1–29, February 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302423.

Cosentino:2018:GSP

[2195] Valerio Cosentino, Javier Luis Cánovas Izquierdo, and Jordi Cabot. Gitana: a software project inspector. Science of Computer Programming, 153(??): 30–33, February 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302666.

Anonymous:2018:EBd

[2196] Anonymous. Editorial Board. Science of Computer Programming, 154 (??):ii, March 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231830008X.

Artho:2018:FTS

[2197] Cyrille Artho and Peter Csaba Ölveczky. Formal Techniques for Safety-Critical Systems (FTSCS 2015). Science of Computer Programming, 154(??):1–2, March 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302939.

Meseguer:2018:VBS

[2198] José Meseguer. Variant-based satisfiability in initial algebras. Science of Computer Programming, 154 (??):3–41, March 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301788.

Zhang:2018:PSM

[2199] Min Zhang, Feng Dai, and Frédéric Mallet. Periodic scheduling for MARTE/CCSL: Theory and practice. Science of Computer Programming, 154(??):42–60, March 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301740.

Berger:2018:VER

[2200] Ulrich Berger, Phillip James, Andrew Lawrence, Markus Roggenbach, and Monika Seisenberger. Verification of the European Rail Traffic Management System in Real-Time Maude. Science of Computer Programming, 154 (??):61–88, March 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231730223X.

Becker:2018:FAF

[2201] Klaus Becker, Sebastian Voss, and Bernhard Schätz. Formal analysis of feature degradation in faulttolerant automotive systems. Science of Computer Programming, 154 (??):89–133, March 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302198.

Wu:2018:FAC

[2202] Xi Wu and Huibiao Zhu. Formal analysis of a calculus for WSNs from quality perspective. Science of Computer Programming, 154(??): 134–153, March 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301661.

Anonymous:2018:EBe

[2203] Anonymous. Editorial Board. Science of Computer Programming, 155 (??):ii, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300315.

Vidal:2018:PSS

[2204] Germán Vidal. Preface for SCP special issue on Principles and Practice of Declarative Programming. Science of Computer Programming, 155 (??):1–2, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302903.

Stulova:2018:STO

[2205] Nataliia Stulova, José F. Morales, and Manuel V. Hermenegildo. Some tradeoffs in reducing the overhead of assertion run-time checks via static analysis. Science of Computer Programming, 155(??):3–26, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302915.

Benton:2018:EDT

[2206] Nick Benton, Martin Hofmann, and Vivek Nigam. Effect-dependent transformations for concurrent programs. Science of Computer Programming, 155(??):27–51, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231730237X.

Kouzapas:2018:TPM

[2207] Dimitrios Kouzapas, Ornela Dardha, Roly Perera, and Simon J. Gay. Typechecking protocols with Mungo and StMungo: A session type toolchain for Java. Science of Computer Programming, 155(??): 52–75, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302186.

Nagao:2018:RIC

[2208] Takahiro Nagao and Naoki Nishida. Rewriting induction for constrained inequalities. Science of Computer Programming, 155(??):76–102, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642317302241.

Fehrenbach:2018:LIP

[2209] Stefan Fehrenbach and James Cheney. Language-integrated provenance. Science of Computer Programming, 155 (??):103–145, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301685.

Grewe:2018:ELS

[2210] Sylvia Grewe, Sebastian Erdweg, André Pacak, Michael Raulf, and Mira Mezini. Exploration of language specifications by compilation to first-order logic. Science of Computer Programming, 155(??): 146–172, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301594.

Charalambidis:2018:HOL

[2211] Angelos Charalambidis, Panos Rondogiannis, and Antonis Troumpoukis. Higher-order logic programming: An expressive language for representing qualitative preferences. Science of Computer Programming, 155(??): 173–197, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301909.

Brau:2018:TSA

[2212] Guillaume Brau, Jérôme Hugues, and Nicolas Navet. Towards the systematic analysis of non-functional properties in model-based engineering for real-time embedded systems. Science of Computer Programming, 156 (??):1–20, May 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302927.

Banach:2018:MFR

[2213] Richard Banach. Modelling, formal refinement and partitioning strategies for a small aircraft fuel pump system in Hybrid Event-B. Science of Computer Programming, 156 (??):21–44, May 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300029.

Paredes-Valverde:2018:OBA

[2214] Mario Andrés Paredes-Valverde, María del Pilar Salas-Zárate, Ricardo Colomo-Palacios, Juan Miguel Gómez-Berbís, and Rafael Valencia-García. An ontology-based approach with which to assign human resources to software projects. Science of Computer Programming, 156(??):90–103, May 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642318300042.

Kamina:2018:MSM

[2215] Tetsuo Kamina, Tomoyuki Aotani, Hidehiko Masuhara, and Atsushi Igarashi. Method safety mechanism for asynchronous layer deactivation. Science of Computer Programming, 156 (??):104–120, May 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300236.

Anonymous:2018:S

[2216] Anonymous. ScienceDirect. Science of Computer Programming, 156 (??):??, May 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302952.

Ralph:2018:TPS

[2217] Paul Ralph. The two paradigms of software development research. Science of Computer Programming, 156 (??):??, May 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300030.

Anonymous:2018:EBf

[2218] Anonymous. Editorial Board. Science of Computer Programming, 157 (??):ii, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300698.

Kordon:2018:ISI

[2219] Fabrice Kordon and Daniel Moldt.

Introduction to the special issue from Petri Nets 2016. Science of Computer Programming, 157 (??):1–2, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231830025X.

Kedzia:2018:SMP

[2220] Kamil Kedzia, Wojtek Ptak, Jacek Sroka, and Andrzej M. Kierzek. Simulation of multicellular populations with Petri nets and genome scale intracellular networks. Science of Computer Programming, 157 (??):3–16, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302149.

Voros:2018:IAP

[2221] András Vörös, Dániel Darvas, Ákos Hajdu, Attila Klenik, Kristóf Marussy, Vince Molnár, Tamás Bartha, and István Majzik. Industrial applications of the PetriDotNet modelling and analysis tool. Science of Computer Programming, 157(??): 17–40, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301910.

Best:2018:PSP

[2222] Eike Best and Raymond Devillers. Pre-synthesis of Petri nets based on prime cycles and distance paths. Science of Computer Programming, 157 (??):41–55, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301454.

Cabac:2018:SDP

[2223] Lawrence Cabac, Michael Haustermann, and David Mosteller. Software development with Petri nets and agents: Approach, frameworks and tool set. Science of Computer Programming, 157(??): 56–70, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302678.

Helouet:2018:RSS

[2224] Loïc Hélouët and Karim Kecir. Realizability of schedules by stochastic time Petri nets with blocking semantics. Science of Computer Programming, 157 (??):71–102, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231730268X.

Anonymous:2018:EBg

[2225] Anonymous. Editorial Board. Science of Computer Programming, 158 (??):ii, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300881.

Butler:2018:IAS

[2226] Michael Butler and Klaus-Dieter Schewe. Introduction to the ABZ 2016 special issue. Science of Computer Programming, 158(??):1–2, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642318300169.

Merz:2018:ETU

[2227] Stephan Merz and Hernán Vanzetto.

Encoding TLA+ into unsorted and many-sorted first-order logic. Science of Computer Programming, 158 (??):3–20, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301922.

Schellhorn:2018:SEC

[2228] Gerhard Schellhorn, Gidon Ernst, Jörg Pfähler, Stefan Bodenmüller, and Wolfgang Reif. Symbolic execution for a clash-free subset of ASMs. Science of Computer Programming, 158 (??):21–40, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301739.

Krings:2018:PAB

[2229] Sebastian Krings and Michael Leuschel. Proof assisted bounded and unbounded symbolic model checking of software and system models. Science of Computer Programming, 158 (??):41–63, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301727.

Azmy:2018:MCC

[2230] Noran Azmy, Stephan Merz, and Christoph Weidenbach. A machinechecked correctness proof for Pastry. Science of Computer Programming, 158(??):64–80, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301612.

Dobrikov:2018:EAE

[2231] Ivaylo Dobrikov and Michael Leuschel. Enabling analysis for Event-B. Science of Computer Programming, 158 (??):81–99, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301624.

Baugh:2018:FMF

[2232] John Baugh and Alper Altuntas. Formal methods and finite element analysis of hurricane storm surge: A case study in software verification. Science of Computer Programming, 158 (??):100–121, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301715.

Hoang:2018:VVR

[2233] Thai Son Hoang, Colin Snook, Asieh Salehi, Michael Butler, and Lukas Ladenberger. Validating and verifying the requirements and design of a haemodialysis machine using the Rodin toolset. Science of Computer Programming, 158(??): 122–147, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302381.

Arcaini:2018:IFM

[2234] Paolo Arcaini, Silvia Bonfanti, Angelo Gargantini, Atif Mashkoor, and Elvinia Riccobene. Integrating formal methods into medical software development: The ASM approach. Science of Computer Programming, 158 (??):148–167, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301430.

Dimovski:2018:VAL

[2235] Aleksandar S. Dimovski, Claus Brabrand, and Andrzej Wasowski. Variability abstractions for lifted analyses. Science of Computer Programming, 159 (??):1–27, July 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300583.

Camilli:2018:ZBF

[2236] Matteo Camilli, Angelo Gargantini, and Patrizia Scandurra. Zone-based formal specification and timing analysis of real-time self-adaptive systems. Science of Computer Programming, 159 (??):28–57, July 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300753.

Anonymous:2018:EBh

[2237] Anonymous. Editorial Board. Science of Computer Programming, 159 (??):1–58, July 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301242.

Anonymous:2018:EBi

[2238] Anonymous. Editorial Board. Science of Computer Programming, 160 (??):ii, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301424.

Dastani:2018:FSE

[2239] Mehdi Dastani, Hossein Hojjat, and Marjan Sirjani. Fundamentals of software engineering (extended versions of selected papers of FSEN 2015). Science of Computer Programming, 160(??):1–2, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302459.

Autili:2018:CRE

[2240] Marco Autili, Paola Inverardi, and Massimo Tivoli. Choreography realizability enforcement through the automatic synthesis of distributed coordination delegates. Science of Computer Programming, 160(??): 3–29, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302228.

Behjati:2018:ALC

[2241] Razieh Behjati and Shiva Nejati.

Architecture-level configuration of industrial control systems: Foundations for an efficient approach. Science of Computer Programming, 160 (??):30–47, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302071.

Jongmans:2018:CCV

[2242] S.-S. T. Q. Jongmans and F. Arbab.

Centralized coordination vs. partiallydistributed coordination with Reo and constraint automata. Science of Computer Programming, 160 (??):48–77, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317301259.

Kiehn:2018:SMC

[2243] Astrid Kiehn and Deepanker Aggarwal. A study of mutable checkpointing and related algorithms. Science of Computer Programming, 160 (??):78–92, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300680.

Hajisheykhi:2018:TIT

[2244] Reza Hajisheykhi, Ali Ebnenasir, and Sandeep S. Kulkarni. A theory of integrating tamper evidence with stabilization. Science of Computer Programming, 160(??): 93–114, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317300503.

Anonymous:2018:EBj

[2245] Anonymous. Editorial board. Science of Computer Programming, 161 (??):ii, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301758.

Etien:2018:ADL

[2246] Anne Etien and Jannik Laval. Advances in dynamic languages. Science of Computer Programming, 161

(??):1, ????  2018.  CODEN

SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300911.

Bouraqadi:2018:TDD

[2247] Noury Bouraqadi and Dave Mason.

Test-driven development for generated portable Javascript apps. Science of Computer Programming, 161 (??):2–17, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300595.

Cavarle:2018:FOM

[2248] Glenn Cavarlé, Alain Plantec, Steven Costiou, and Vincent Ribaud. A feature-oriented model-driven engineering approach for the early validation of feature-based applications. Science of Computer Programming, 161(??):18–33, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300017.

Bergel:2018:RRC

[2249] Alexandre Bergel, Alejandro Infante, Sergio Maass, and Juan Pablo Sandoval Alcocer. Reducing resource consumption of expandable collections: The Pharo case. Science of Computer Programming, 161 (??):34–56, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302940.

Kurs:2018:EPP

[2250] Jan Kurs, Jan Vraný, Mohammad Ghafari, Mircea Lungu, and Oscar Nierstrasz. Efficient parsing with parser combinators. Science of Computer Programming, 161 (??):57–88, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302654.

Pollet:2018:CAS

[2251] Damien Pollet and Stéphane Ducasse. A critical analysis of string APIs: The case of Pharo. Science of Computer Programming, 161(??): 89–104, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302435.

Milojkovic:2018:MIC

[2252] Nevena Milojković, Clément Béra, Mohammad Ghafari, and Oscar Nierstrasz. Mining inline cache data to order inferred types in dynamic languages. Science of Computer Programming, 161(??): 105–121, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642317302411.

Leske:2018:ILD

[2253] Max Leske, Andrei Chis, and Oscar Nierstrasz. Improving live debugging of concurrent threads through thread histories. Science of Computer Programming, 161(??):

122–148, ????  2018.  CODEN

SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302174.

Costiou:2018:LPF

[2254] Steven Costiou, Mickaël Kerboeuf, Glenn Cavarlé, and Alain Plantec. Lub: A pattern for fine grained behavior adaptation at runtime. Science of Computer Programming, 161 (??):149–171, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301946.

Anonymous:2018:EBk

[2255] Anonymous. Editorial board. Science of Computer Programming, 162 (??):ii, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301990.

Bonsangue:2018:PSI

[2256] Marcello Bonsangue and Yuxin Deng.

Preface for the special issue of the 10th International Symposium on Theoretical Aspects of Software Engineering (TASE 2016). Science of Computer Programming, 162 (??):1–2, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301023.

He:2018:NRL

[2257] Jifeng He and Qin Li. A new roadmap for linking theories of programming and its applications on GCL and CSP. Science of Computer Programming, 162(??):3–34, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302216.

Araujo:2018:SCT

[2258] Hugo Araujo, Gustavo Carvalho, Morteza Mohaqeqi, Mohammad Reza Mousavi, and Augusto Sampaio. Sound conformance testing for cyber-physical systems: Theory and implementation. Science of Computer Programming, 162(??):35–54, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301429.

Gammaitoni:2018:AVM

[2259] Loïc Gammaitoni, Pierre Kelsen, and Qin Ma. Agile validation of model transformations using compound f-alloy specifications. Science of Computer Programming, 162 (??):55–75, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301417.

Jomaa:2018:FPD

[2260] Narjes Jomaa, David Nowak, Gilles Grimaud, and Samuel Hym. Formal proof of dynamic memory isolation based on MMU. Science of Computer Programming, 162(??): 76–92, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301338.

Xu:2018:STA

[2261] Zhiwu Xu, Cheng Wen, and Shengchao Qin. State-taint analysis for detecting resource bugs. Science of Computer Programming, 162(??): 93–109, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301314.

Bonchi:2018:SBM

[2262] Filippo Bonchi, Antonio Brogi, Andrea Canciani, and Jacopo Soldani. Simulation-based matching of cloud applications. Science of Computer Programming, 162(??): 110–131, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301223.

Hu:2018:LEC

[2263] Yan Hu, Jiwei Yan, Dong Yan, Qiong Lu, and Jun Yan. Lightweight energy consumption analysis and prediction for Android applications. Science of Computer Programming, 162 (??):132–147, ???? 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300953.

Anonymous:2018:EBl

[2264] Anonymous. Editorial board. Science of Computer Programming, 163 (??):ii, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318302132.

Wang:2018:AEM

[2265] Jue Wang, Yanyan Jiang, Chang Xu, Qiwei Li, Tianxiao Gu, Jun Ma, Xiaoxing Ma, and Jian Lu. AATT+: Effectively manifesting concurrency bugs in Android apps. Science of Computer Programming, 163(??): 1–18, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301047.

Kim:2018:TSA

[2266] Dongwoo Kim and Yunja Choi. A two-step approach for pattern-based API-call constraint checking. Science of Computer Programming, 163(??): 19–41, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301205.

Martini:2018:TDT

[2267] Antonio Martini, Terese Besker, and Jan Bosch. Technical debt tracking: Current state of practice: A survey and multiple case study in 15 large organizations. Science of Computer Programming, 163(??):42– 61, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301035.

Wille:2018:ICT

[2268] David Wille, Önder Babur, Loek Cleophas, Christoph Seidl, Mark van den Brand, and Ina Schaefer. Improving custom-tailored variability mining using outlier and cluster detection. Science of Computer Programming, 163(??):62–84, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642318301382.

Gomez-Abajo:2018:TDI

[2269] Pablo Gómez-Abajo, Esther Guerra, Juan de Lara, and Mercedes G. Merayo. A tool for domainindependent model mutation. Science of Computer Programming, 163 (??):85–92, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300261.

Tamrawi:2018:PCG

[2270] Ahmed Tamrawi and Suresh Kothari.

Projected control graph for computing relevant program behaviors. Science of Computer Programming, 163 (??):93–114, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301497.

Song:2018:FAT

[2271] Jiyoung Song, Eunkyoung Jee, and Doo-Hwan Bae. FBDTester 2.0: Automated test sequence generation for FBD programs with internal memory states. Science of Computer Programming, 163(??):115– 137, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301552.

Potanin:2018:PSI

[2272] Alex Potanin and Gail Murphy. Preface for the Special Issue on the 23rd Asia-Pacific Software Engineering Conference (APSEC) 2016. Science of Computer Programming, 163 (??):138, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301710.

Mauro:2018:CAR

[2273] Jacopo Mauro, Michael Nieke, Christoph Seidl, and Ingrid Chieh Yu. Contextaware reconfiguration in evolving software product lines. Science of Computer Programming, 163(??):139– 159, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301692.

Furda:2018:PAD

[2274] Andrei Furda, Colin Fidge, and Alistair Barros. A practical approach for detecting multi-tenancy data interference. Science of Computer Programming, 163(??):160– 173, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301564.

Gomes:2018:SVS

[2275] Pedro de C. Gomes, Dilian Gurov, Marieke Huisman, and Cyrille Artho. Specification and verification of synchronization with condition variables. Science of Computer Programming, 163(??):174–189, October 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642318301588.

Gracioli:2018:PSI

[2276] Giovani Gracioli and Rivalino Matias.

Preface for the special issue of the 6th Brazilian Symposium on Computing System Engineering (SBESC 2016). Science of Computer Programming, 165(??):ii, November 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300625.

Araujo:2018:CGI

[2277] Rodrigo F. Araújo, Higo F. Albuquerque, Iury V. de Bessa, Lucas C. Cordeiro, and João E. Chaves Filho. Counterexample guided inductive optimization based on satisfiability modulo theories. Science of Computer Programming, 165(??): 1–2, November 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317302162.

Resner:2018:DIC

[2278] Davi Resner, Gustavo Medeiros de Araujo, and Antônio Augusto Fröhlich. Design and implementation of a cross-layer IoT protocol. Science of Computer Programming, 165 (??):3–23, November 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301673.

Valentin:2018:TOS

[2279] Eduardo Valentin, Rosiane de Freitas, and Raimundo Barreto. Towards optimal solutions for the low power hard real-time task allocation on multiple heterogeneous processors. Science of Computer Programming, 165 (??):24–37, November 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301648.

Brandalero:2018:AET

[2280] Marcelo Brandalero, Leonardo Almeida da Silveira, Jeckson Dellagostin Souza, and Antonio Carlos Schneider Beck. Accelerating errortolerant applications with approximate function reuse. Science of Computer Programming, 165(??):38– 53, November 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317300965.

Anonymous:2018:EBm

[2281] Anonymous. Editorial Board. Science of Computer Programming, 165 (??):??, November 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303149.

Anonymous:2018:EBn

[2282] Anonymous. Editorial Board. Science of Computer Programming, 166 (??):ii, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303423.

Cleophas:2018:WAK

[2283] Loek Cleophas, Ina Schaefer, and Bruce W. Watson. Workshop on advances in knowledge extraction and re-engineering of software (selected and extended papers from WAKERS 2017). Science of Computer Programming, 166(??):1– 2, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301849.

Lienhardt:2018:CDO

[2284] Michael Lienhardt, Ferruccio Damiani, Lorenzo Testa, and Gianluca Turin. On checking delta-oriented product lines of statecharts. Science of Computer Programming, 166(??): 3–34, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301953.

Corradini:2018:FAM

[2285] Flavio Corradini, Fabrizio Fornari, Andrea Polini, Barbara Re, and Francesco Tiezzi. A formal approach to modeling and verification of business process collaborations. Science of Computer Programming, 166(??):35– 70, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318302053.

Du:2018:PSE

[2286] Dehui Du, Ping Huang, Kaiqiang Jiang, and Frédéric Mallet. pCSSL: a stochastic extension to MARTE/ CCSL for modeling uncertainty in Cyber Physical Systems. Science of Computer Programming, 166(??):71– 88, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301916.

Mefteh:2018:TNP

[2287] Mariem Mefteh, Nadia Bouassida, and Hanêne Ben-Abdallah. Towards naturalistic programming: Mapping language-independent requirements to constrained language specifications. Science of Computer Programming, 166(??):89–119, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642318301941.

Scott:2018:GSA

[2288] Elizabeth Scott and Adrian Johnstone. GLL syntax analysers for EBNF grammars. Science of Computer Programming, 166(??):120–145, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318302375.

Audrito:2018:OSP

[2289] Giorgio Audrito, Ferruccio Damiani, and Mirko Viroli. Optimal single-path information propagation in gradient-based algorithms. Science of Computer Programming, 166(??):146– 166, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302387.

Kamburjan:2018:FMA

[2290] Eduard Kamburjan, Reiner Hähnle, and Sebastian Schön. Formal modeling and analysis of railway operations with active objects. Science of Computer Programming, 166(??):167– 193, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302508.

Brogi:2018:OIT

[2291] Antonio Brogi, Davide Neri, Luca Rinaldi, and Jacopo Soldani. Orchestrating incomplete TOSCA applications with Docker. Science of Computer Programming, 166(??):194– 213, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302673.

Tempero:2018:FDC

[2292] Ewan Tempero and Paul Ralph. A framework for defining coupling metrics. Science of Computer Programming, 166(??):214–230, November 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642318300613.

Anonymous:2018:EBo

[2293] Anonymous. Editorial Board. Science of Computer Programming, 167 (??):ii, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318303691.

Anonymous:2018:VOS

[2294] Anonymous. A vigorous and open SCP. Science of Computer Programming, 167(??):iii, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318303721.

Pittou:2018:CMV

[2295] Maria Pittou, Panagiotis Manolios, Jan Reineke, and Stavros Tripakis. Checking multi-view consistency of discrete systems with respect to periodic sampling abstractions. Science of Computer Programming, 167(??): 1–24, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318302612.

Barbanera:2018:TRS

[2296] Franco Barbanera, Ivan Lanese, and Ugo de’Liguoro. A theory of retractable and speculative contracts. Science of Computer Programming, 167(??): 25–50, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302636.

Camara:2018:RAS

[2297] Javier Cámara, Wenxin Peng, David Garlan, and Bradley Schmerl. Reasoning about sensing uncertainty and its reduction in decision-making for self-adaptation. Science of Computer Programming, 167(??):51– 69, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302521.

Hocaoglu:2018:AAD

[2298] Mehmet F. Hocaoglu. AdSiF: Agent driven simulation framework paradigm and ontological view. Science of Computer Programming, 167(??):70– 90, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302648.

Gadducci:2018:SIR

[2299] Fabio Gadducci, Hernán Melgratti, and Christian Roldán. On the semantics and implementation of replicated data types. Science of Computer Programming, 167(??):91– 113, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302429.

Casadei:2018:TAR

[2300] Roberto Casadei, Alessandro Aldini, and Mirko Viroli. Towards attackresistant Aggregate Computing using trust mechanisms. Science of Computer Programming, 167(??):114– 137, December 1, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303046.

Anonymous:2018:EBp

[2301] Anonymous. Editorial Board. Science of Computer Programming, 168 (??):ii, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303939.

Duran:2018:SAB

[2302] Francisco Durán, Camilo Rocha, and Gwen Salaün. Stochastic analysis of BPMN with time in rewriting logic. Science of Computer Programming, 168(??):1–17, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303307.

Ameller:2018:NFR

[2303] David Ameller, Xavier Burgués, Dolors Costal, Carles Farré, and Xavier Franch. Non-functional requirements in model-driven development of serviceoriented architectures. Science of Computer Programming, 168(??):18– 37, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303034.

Yang:2018:EPA

[2304] Chen Yang, Peng Liang, and Paris Avgeriou. Evaluation of a process for architectural assumption management in software development. Science of Computer Programming, 168(??): 38–70, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303253.

Nicola:2018:AEE

[2305] Rocco De Nicola, Tan Duong, Omar Inverso, and Catia Trubiani. AErlang: Empowering Erlang with attributebased communication. Science of Computer Programming, 168(??):71– 93, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303290.

Bernardi:2018:FAC

[2306] Giovanni Bernardi and Adrian Francalanza. Full-abstraction for client testing preorders. Science of Computer Programming, 168(??):94–117, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303277.

Baldan:2018:MMI

[2307] Paolo Baldan and Alberto Lluch Lafuente. Many-to-many information flow policies. Science of Computer Programming, 168(??):118– 141, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303265.

deBerardinis:2018:ABM

[2308] Jacopo de Berardinis, Giorgio Forcina, Ali Jafari, and Marjan Sirjani. Actorbased macroscopic modeling and simulation for smart urban planning. Science of Computer Programming, 168(??):142–164, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642318303459.

Rodriguez-Echeverria:2018:AAG

[2309] Roberto Rodriguez-Echeverria, Juan C. Preciado, Javier Sierra, Jose M. Conejero, and Fernando Sanchez-Figueroa. AutoCRUD: Automatic generation of CRUD specifications in interaction flow modelling language. Science of Computer Programming, 168(??):165– 168, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303575.

Canal:2018:PSI

[2310] Carlos Canal and Gwen Salaün. Preface: Special issue on foundations of coordination languages and self-adaptive systems. Science of Computer Programming, 168(??):169– 170, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231830354X.

Varshosaz:2018:BBM

[2311] Mahsa Varshosaz, Harsh Beohar, and Mohammad Reza Mousavi. Basic behavioral models for software product lines: Revisited. Science of Computer Programming, 168(??):171– 185, December 15, 2018. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303381.

Anonymous:2019:EBa

[2312] Anonymous. Editorial Board. Science of Computer Programming, 169 (??):ii, January 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318304167.

terBeek:2019:EMT

[2313] Maurice H. ter Beek, Ferruccio Damiani, Stefania Gnesi, Franco Mazzanti, and Luca Paolini. On the expressiveness of modal transition systems with variability constraints. Science of Computer Programming, 169(??): 1–17, January 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303769.

Yin:2019:DDF

[2314] Jie Yin, Gang Tan, Hao Li, Xiaolong Bai, Yu-Ping Wang, and Shi-Min Hu. Debugopt: Debugging fully optimized natively compiled programs using multistage instrumentation. Science of Computer Programming, 169 (??):18–32, January 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318303629.

Anonymous:2019:EBb

[2315] Anonymous. Editorial Board. Science of Computer Programming, 170 (??):ii, January 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318304258.

Krishna:2019:CBP

[2316] Ajay Krishna, Pascal Poizat, and Gwen Salaün. Checking business process evolution. Science of Computer Programming, 170(??):1– 26, January 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318304052.

Zhang:2019:RAC

[2317] Xiyue Zhang, Weijiang Hong, Yi Li, and Meng Sun. Reasoning about connectors using Coq and Z3. Science of Computer Programming, 170(??): 27–44, January 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318304076.

Syriani:2019:SBP

[2318] Eugene Syriani, Vasco Sousa, and Levi Lúcio. Structure and behavior preserving statecharts refinements. Science of Computer Programming, 170 (??):45–79, January 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318304118.

Anonymous:2019:EBc

[2319] Anonymous. Editorial Board. Science of Computer Programming, 171 (??):ii, February 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318304398.

Kyriakakis:2019:EFC

[2320] Panos Kyriakakis, Alexander Chatzigeorgiou, Apostolos Ampatzoglou, and Stelios Xinogalos. Exploring the frequency and change proneness of dynamic feature pattern instances in PHP applications. Science of Computer Programming, 171(??):1– 20, February 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318304106.

Iannopollo:2019:CSC

[2321] Antonio Iannopollo, Stavros Tripakis, and Alberto Sangiovanni-Vincentelli. Constrained synthesis from component libraries. Science of Computer Programming, 171(??): 21–41, February 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231830409X.

Abbassi:2019:UGL

[2322] Imran Hafeez Abbassi, Faiq Khalid, Osman Hasan, and Awais Mehmood Kamboh. Using gate-level side channel parameters for formally analyzing vulnerabilities in integrated circuits. Science of Computer Programming, 171 (??):42–66, February 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318304210.

Fragal:2019:HFS

[2323] Vanderson Hafemann Fragal, Adenilso Simao, and Mohammad Reza Mousavi. Hierarchical featured state machines. Science of Computer Programming, 171(??):67–88, February 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642318304064.

Anonymous:2019:EBd

[2324] Anonymous. Editorial Board. Science of Computer Programming, 172 (??):ii, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300073.

Bonchi:2019:BOT

[2325] Filippo Bonchi, Tom van Bussel, Matias David Lee, and Jurriaan Rot. Bisimilarity of open terms in stream GSOS. Science of Computer Programming, 172(??): 1–26, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300406.

Luthmann:2019:UMI

[2326] Lars Luthmann, Stephan Mennicke, and Malte Lochau. Unifying modal interface theories and compositional input/output conformance testing. Science of Computer Programming, 172 (??):27–47, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318304088.

Benzmuller:2019:UML

[2327] Christoph Benzmüller. Universal (meta-)logical reasoning: Recent successes. Science of Computer Programming, 172(??):48–62, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642318301461.

Bouraoui:2019:MDE

[2328] Amina Bouraoui and Imen Gharbi.

Model driven engineering of accessible and multi-platform graphical user interfaces by parameterized model transformations. Science of Computer Programming, 172(??): 63–101, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318304301.

Deutschbein:2019:MCC

[2329] Calvin Deutschbein, Tom Fleming, Alan Burns, and Sanjoy Baruah. Multi-core cyclic executives for safetycritical systems. Science of Computer Programming, 172(??):102– 116, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300601.

Han:2019:HVA

[2330] Xiao Han, Tao Tang, and Jidong Lv. A hierarchical verification approach to verify complex safety control systems based on STAMP. Science of Computer Programming, 172 (??):117–134, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318304325.

Imai:2019:SOS

[2331] Keigo Imai, Nobuko Yoshida, and Shoji Yuen. Session-ocaml: a session-based library with polarities and lenses. Science of Computer Programming, 172 (??):135–159, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318303289.

Ferrando:2019:EBC

[2332] Angelo Ferrando. The early bird catches the worm: First verify, then monitor! Science of Computer Programming, 172(??):160– 179, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318304349.

Giannini:2019:TSI

[2333] Paola Giannini, Tim Richter, Marco Servetto, and Elena Zucca. Tracing sharing in an imperative pure calculus. Science of Computer Programming, 172(??):180–202, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642318304337.

Damiani:2019:FMM

[2334] Ferruccio Damiani, Michael Lienhardt, and Luca Paolini. A formal model for Multi Software Product Lines. Science of Computer Programming, 172 (??):203–231, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300364.

Parra:2019:ECE

[2335] Otto Parra, Sergio España, Jose Ignacio Panach, and Oscar Pastor. An empirical comparative evaluation of gestUI to include gesture-based interaction in user interfaces. Science of Computer Programming, 172(??): 232–263, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300741.

Sanchez:2019:FHS

[2336] Luis Emiliano Sánchez, Jorge Andrés Diaz-Pace, and Alejandro Zunino. A family of heuristic search algorithms for feature model optimization. Science of Computer Programming, 172 (??):264–293, March 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318304350.

Anonymous:2019:EBe

[2337] Anonymous. Editorial Board. Science of Computer Programming, 173 (??):ii, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300140.

Castor:2019:SPB

[2338] Fernando Castor and Yu David Liu. Selected papers of the Brazilian Symposium on Programming Languages (SBLP’15+16). Science of Computer Programming, 173(??): 1–2, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318304313.

Ferreira:2019:DEV

[2339] Bruno Morais Ferreira, Britaldo Silveira Soares-Filho, and Fernando Magno Quintão Pereira. The Dinamica EGO virtual machine. Science of Computer Programming, 173 (??):3–20, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300455.

Bessa:2019:JFM

[2340] Tarsila Bessa, Christopher Gull, Pedro Quintão, Michael Frank, José Nacif, and Fernando Magno Quintão Pereira. JetsonLEAP: a framework to measure power on a heterogeneous system-on-a-chip device. Science of Computer Programming, 173 (??):21–36, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642317301776.

daSilvaFeitosa:2019:MSQ

[2341] Samuel da Silva Feitosa, Juliana Kaizer Vizzotto, Eduardo Kessler Piveta, and Andre Rauber Du Bois. A monadic semantics for quantum computing in an object oriented language. Science of Computer Programming, 173 (??):37–55, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300844.

MedeirosDuarte:2019:CPC

[2342] Rodrigo Medeiros Duarte, André Rauber Du Bois, Maurício Lima Pilla, Gerson Geraldo H. Cavalheiro, and Renata Hax Sander Reiser. Comparing the performance of concurrent hash tables implemented in Haskell. Science of Computer Programming, 173 (??):56–70, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302491.

Fernandes:2019:MZB

[2343] João Paulo Fernandes, Pedro Martins, Alberto Pardo, João Saraiva, and Marcos Viera. Memoized zipperbased attribute grammars and their higher order extension. Science of Computer Programming, 173(??): 71–94, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231830412X.

Silva:2019:SWM

[2344] Jefferson de Carvalho Silva, Allberson Bruno de Oliveira Dantas, and Francisco Heron de Carvalho Junior. A Scientific Workflow Management System for orchestration of parallel components in a cloud of large-scale parallel processing services. Science of Computer Programming, 173(??): 95–127, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301527.

Rocha:2019:APR

[2345] Rodrigo C. O. Rocha, Luís F. W. Góes, and Fernando M. Q. Pereira. Automatic parallelization of recursive functions with rewriting rules. Science of Computer Programming, 173(??): 128–152, March 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300066.

Anonymous:2019:EBf

[2346] Anonymous. Editorial Board. Science of Computer Programming, 174 (??):ii, April 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300243.

Cavalcanti:2019:VSR

[2347] Ana Cavalcanti, Augusto Sampaio, Alvaro Miyazawa, Pedro Ribeiro, Madiel Conserva Filho, André Didier, Wei Li, and Jon Timmis. Verified simulation for robotics. Science of Computer Programming, 174 (??):1–37, April 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301655.

Muram:2019:SAC

[2348] Faiz UL Muram, Huy Tran, and Uwe Zdun. Supporting automated containment checking of software behavioural models using model transformations and model checking. Science of Computer Programming, 174 (??):38–71, April 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300036.

Lindner:2019:TTA

[2349] Andreas Lindner, Roberto Guanciale, and Roberto Metere. TrABin: Trustworthy analyses of binaries. Science of Computer Programming, 174 (??):72–89, April 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301667.

Budde:2019:ACI

[2350] Carlos E. Budde, Pedro R. D’Argenio, and Arnd Hartmanns. Automated compositional importance splitting. Science of Computer Programming, 174(??):90–108, April 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301503.

Anonymous:2019:EBg

[2351] Anonymous. Editorial Board. Science of Computer Programming, 175 (??):ii, April 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300322.

Santiago:2019:IDL

[2352] Regivan Santiago, Benjamín Bedregal, Alexandre Madeira, and Manuel A. Martins. On interval dynamic logic: Introducing quasi-action lattices. Science of Computer Programming, 175 (??):1–16, April 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300103.

Harrison:2019:FTS

[2353] Michael D. Harrison, Leo Freitas, Michael Drinnan, José C. Campos, Paolo Masci, Costanzo di Maria, and Michael Whitaker. Formal techniques in the safety analysis of software components of a new dialysis machine. Science of Computer Programming, 175 (??):17–34, April 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300819.

Artho:2019:FTS

[2354] Cyrille Artho and Peter Csaba Ölveczky. Formal Techniques for Safety-Critical Systems (FTSCS 2016). Science of Computer Programming, 175(??):35–36, April 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231930019X.

Timm:2019:TVB

[2355] Nils Timm and Stefan Gruner. Threevalued bounded model checking with cause-guided abstraction refinement. Science of Computer Programming, 175(??):37–62, April 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300206.

Scott:2019:DRU

[2356] Elizabeth Scott, Adrian Johnstone, and L. Thomas van Binsbergen. Derivation representation using binary subtree sets. Science of Computer Programming, 175(??):63– 84, April 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302302.

Anonymous:2019:EBh

[2357] Anonymous. Editorial Board. Science of Computer Programming, 176 (??):ii, May 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300401.

Perez-Schofield:2019:DOO

[2358] Baltasar García Perez-Schofield and Francisco Ortin. A didactic objectoriented, prototype-based visual programming environment. Science of Computer Programming, 176(??): 1–13, May 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318302326.

Avelino:2019:MAC

[2359] Guilherme Avelino, Leonardo Passos, Andre Hora, and Marco Tulio Valente. Measuring and analyzing code authorship in 1 + 118 open source projects. Science of Computer Programming, 176(??):14–32, May 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318300388.

Anonymous:2019:EBi

[2360] Anonymous. Editorial Board. Science of Computer Programming, 177 (??):ii, May 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300486.

Hierons:2019:FQE

[2361] Robert M. Hierons. FSM quasiequivalence testing via reduction and observing absences. Science of Computer Programming, 177(??): 1–18, May 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231830323X.

Li:2019:FFC

[2362] Yi Li, Xiyue Zhang, Yuanyi Ji, and Meng Sun. A formal framework capturing real-time and stochastic behavior in connectors. Science of Computer Programming, 177 (??):19–40, May 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301631.

Sabahi-Kaviani:2019:BMI

[2363] Zeynab Sabahi-Kaviani and Fatemeh Ghassemi. Behavioral model identification and classification of multi-component systems. Science of Computer Programming, 177(??): 41–66, May 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231830039X.

Anonymous:2019:EBj

[2364] Anonymous. Editorial Board. Science of Computer Programming, 178 (??):ii, June 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300565.

Jia:2019:USD

[2365] Jingdong Jia, Xiaoying Yang, Rong Zhang, and Xi Liu. Understanding software developers’ cognition in agile requirements engineering. Science of Computer Programming, 178 (??):1–19, June 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302764.

Bae:2019:SSS

[2366] Kyungmin Bae and Camilo Rocha. Symbolic state space reduction with guarded terms for rewriting modulo SMT. Science of Computer Programming, 178(??):20–42, June 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642318301187.

Kouchnarenko:2019:SEP

[2367] Olga Kouchnarenko and Eric Madelaine. Selected and extended papers from FACS 2016. Science of Computer Programming, 178(??): 43–44, June 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300449.

Anonymous:2019:EBk

[2368] Anonymous. Editorial Board. Science of Computer Programming, 179 (??):ii, June 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300693.

Peleska:2019:FCS

[2369] Jan Peleska, Wen ling Huang, and Ana Cavalcanti. Finite complete suites for CSP refinement testing. Science of Computer Programming, 179 (??):1–23, June 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300620.

Bijo:2019:FMD

[2370] Shiji Bijo, Einar Broch Johnsen, Ka I Pun, and S. Lizeth Tapia Tarifa. A formal model of data access for multicore architectures with multilevel caches. Science of Computer Programming, 179(??):24–53, June 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301138.

Inoue:2019:TSF

[2371] Hiroaki Inoue and Atsushi Igarashi. A type system for first-class layers with inheritance, subtyping, and swapping. Science of Computer Programming, 179(??):54–86, June 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300619.

Anonymous:2019:EBl

[2372] Anonymous. Editorial Board. Science of Computer Programming, 180 (??):ii, July 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300796.

Vassallo:2019:LSE

[2373] Carmine Vassallo, Giovanni Grano, Fabio Palomba, Harald C. Gall, and Alberto Bacchelli. A large-scale empirical exploration on refactoring activities in open source software projects. Science of Computer Programming, 180 (??):1–15, July 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318302557.

Jacquet:2019:PSI

[2374] Jean-Marie Jacquet and Mieke Massink. Preface to the special issue on Coordination Models and Languages (Coordination 2017). Science of Computer Programming, 180(??): 16–17, July 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300632.

Yuan:2019:NDB

[2375] Qixia Yuan, Andrzej Mizera, Jun Pang, and Hongyang Qu. A new decomposition-based method for detecting attractors in synchronous Boolean networks. Science of Computer Programming, 180 (??):18–35, July 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318301060.

Jeffery:2019:ASI

[2376] Alex Jeffery and Martin Berger. Asynchronous sessions with implicit functions and messages. Science of Computer Programming, 180(??): 36–70, July 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300656.

Segura:2019:EEP

[2377] Ángel Mora Segura and Juan de Lara. Extremo: an Eclipse plugin for modelling and meta-modelling assistance.

Science of Computer Programming, 180(??):71–80, July 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300644.

Anonymous:2019:EBm

[2378] Anonymous. Editorial Board. Science of Computer Programming, 181 (??):ii, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0167642319300887.

Gkolfi:2019:TAO

[2379] Anastasia Gkolfi, Crystal Chang Din, Einar Broch Johnsen, Lars Michael Kristensen, Martin Steffen, and Ingrid Chieh Yu. Translating active objects into colored Petri nets for communication analysis. Science of Computer Programming, 181 (??):1–26, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0167642318300418.

Wang:2019:DEJ

[2380] Yuchen Wang, Kwok Sun Cheng, Myoungkyu Song, and Eli Tilevich. A declarative enhancement of JavaScript programs by leveraging the Java metadata infrastructure. Science of Computer Programming, 181 (??):27–46, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0167642319300723.

Molina:2019:EAT

[2381] Facundo Molina, César Cornejo, Renzo Degiovanni, Germán Regis, Pablo F. Castro, Nazareno Aguirre, and Marcelo F. Frias. An evolutionary approach to translating operational specifications into declarative specifications. Science of Computer Programming, 181(??): 47–63, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300735.

Laneve:2019:LDA

[2382] Cosimo Laneve. A lightweight deadlock analysis for programs with threads and reentrant locks. Science of Computer Programming, 181 (??):64–81, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318303812.

Cornelio:2019:GES

[2383] M. L. Cornélio and A. W. Roscoe. Guest editorial for the special issue from the 18th Brazilian Symposium on Formal Methods (SBMF 2015). Science of Computer Programming, 181 (??):82–83, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300747.

Nogueira:2019:TCG

[2384] Sidney Nogueira, Hugo Araujo, Renata Araujo, Juliano Iyoda, and Augusto Sampaio. Test case generation, selection and coverage from natural language. Science of Computer Programming, 181(??):84–110, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL https://www.sciencedirect.
com/science/article/pii/S0167642319300024.

Silva:2019:CSB

[2385] Bruno Cesar F. Silva, Gustavo Carvalho, and Augusto Sampaio. CPN simulation-based test case generation from controlled natural-language requirements. Science of Computer Programming, 181(??):111– 139, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0167642319300516.

Miyazawa:2019:SCS

[2386] Alvaro Miyazawa, Ana Cavalcanti, and Andy Wellings. SCJ-Circus: Specification and refinement of safetycritical Java programs. Science of Computer Programming, 181(??): 140–176, July 15, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0167642319300012.

Anonymous:2019:EBn

[2387] Anonymous. Editorial Board. Science of Computer Programming, 182 (??):ii, August 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S016764231930098X.

Marmsoler:2019:CDA

[2388] Diego Marmsoler. A calculus for dynamic architectures. Science of Computer Programming, 182(??):1– 41, August 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764231830385X.

Bowles:2019:FAC

[2389] J. Bowles, M. B. Caminati, S. Cha, and J. Mendoza. A framework for automated conflict detection and resolution in medical guidelines. Science of Computer Programming, 182 (??):42–63, August 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319300759.

Vidal:2019:RAC

[2390] Santiago Vidal, Willian Oizumi, Alessandro Garcia, Andrés Díaz Pace, and Claudia Marcos. Ranking architecturally critical agglomerations of code smells. Science of Computer Programming, 182(??): 64–85, August 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318303514.

Dastani:2019:FSE

[2391] Mehdi Dastani and Marjan Sirjani. Fundamentals of Software Engineering (extended versions of selected papers of FSEN 2017). Science of Computer Programming, 182(??): 86–87, August 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0167642319300838.

Ancona:2019:PSI

[2392] Davide Ancona. Preface: Special issue on verification of objects at runtime execution. Science of Computer Programming, 182(??):88– 89, August 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S016764231930084X.

Anonymous:2019:S

[2393] Anonymous. 1 September 2019. Science of Computer Programming, 183 (??):??, September 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2019:EBo

[2394] Anonymous. Editorial Board. Science of Computer Programming, 183 (??):??, September 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301108.

Kappe:2019:SCA

[2395] Tobias Kappé, Benjamin Lion, Farhad Arbab, and Carolyn Talcott. Soft component automata: Composition, compilation, logic, and verification. Science of Computer Programming, 183 (??):??, September 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301175.

Tigane:2019:RGM

[2396] Samir Tigane, Laid Kahloul, Saber Benharzallah, Souheib Baarir, and Samir Bourekkache. Reconfigurable GSPNs: a modeling formalism of evolvable discrete-event systems. Science of Computer Programming, 183(??): ??, September 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642318302363.

Duran:2019:RLA

[2397] Francisco Durán, Camilo Rocha, and Gwen Salaün. A rewriting logic approach to resource allocation analysis in business process models. Science of Computer Programming, 183 (??):??, September 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319301017.

Anonymous:2019:O

[2398] Anonymous. 1 October 2019. Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2019:EBp

[2399] Anonymous. Editorial Board. Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319301297.

Saadatjoo:2019:TDG

[2400] Mohammad Ali Saadatjoo and Seyed Morteza Babamir. Test-data generation directed by program path coverage through imperialist competitive algorithm. Science of Computer Programming, 184(??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642319301029.

Ferrarotti:2019:BAS

[2401] Flavio Ferrarotti, Senén González, and Klaus-Dieter Schewe. BSP abstract state machines capture bulk synchronous parallel computations. Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303848.

Zhang:2019:DSJ

[2402] Dan Zhang, Dragan Bošnački, Mark van den Brand, Cornelis Huizing, Bart Jacobs, Ruurd Kuiper, and Anton Wijs. Dependency safety for Java — Implementing and testing failboxes. Science of Computer Programming, 184(??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301145.

Mariani:2019:CST

[2403] Stefano Mariani. Coordination in socio-technical systems: Where are we now? where do we go next? Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301157.

Accattoli:2019:AMO

[2404] Beniamino Accattoli and Giulio Guerrieri. Abstract machines for Open Call-by-Value. Science of Computer Programming, 184(??): ??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231830042X.

Chaabane:2019:MAS

[2405] Mariam Chaabane, Ismael Bouassida Rodriguez, Ricardo Colomo-Palacios, Walid Gaaloul, and Mohamed Jmaiel. A modeling approach for Systems-of-Systems by adapting ISO/IEC/IEEE 42010 Standard evaluated by Goal–Question–Metric. Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301030.

Winter:2019:MCO

[2406] Kirsten Winter, Graeme Smith, and John Derrick. Modelling concurrent objects running on the TSO and ARMv8 memory models. Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301066.

Beierle:2019:TEA

[2407] Christoph Beierle and Klaus-Dieter Schewe. Type-2 effectivity in abstract state machines for algorithms with exact real arithmetic. Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301054.

Emara:2019:RSL

[2408] Tamer Z. Emara and Joshua Zhexue Huang. RRPlib: a Spark library for representing HDFS blocks as a set of random sample data blocks. Science of Computer Programming, 184 (??):??, October 1, 2019. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319300942.

Anonymous:2020:Ja

[2409] Anonymous. 1 January 2020. Science of Computer Programming, 185 (??):??, January 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBa

[2410] Anonymous. Editorial Board. Science of Computer Programming, 185 (??):??, January 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301509.

Benes:2020:PPS

[2411] Nikola Benes, Lubos Brim, Samuel Pastva, and David Safránek. Parallel parameter synthesis algorithm for hybrid CTL. Science of Computer Programming, 185(??):??, January 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642319301170.

Lima:2020:GJT

[2412] Caio Lima, Junio Cezar, Guilherme Vieira Leobas, Erven Rohou, and Fernando Magno Quintão Pereira. Guided just-in-time specialization. Science of Computer Programming, 185 (??):??, January 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318303058.

Moser:2020:AAR

[2413] Georg Moser and Manuel Schneckenreither. Automated amortised resource analysis for term rewrite systems. Science of Computer Programming, 185 (??):??, January 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301042.

Avanzini:2020:PTR

[2414] Martin Avanzini, Ugo Dal Lago, and Akihisa Yamada. On probabilistic term rewriting. Science of Computer Programming, 185(??): ??, January 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301339.

Bowles:2020:CCP

[2415] Juliana Küster Filipe Bowles and Marco B. Caminati. Correct composition in the presence of behavioural conflicts and dephasing. Science of Computer Programming, 185(??): ??, January 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301072.

Anonymous:2020:Fa

[2416] Anonymous. 1 February 2020. Science of Computer Programming, 186 (??):??, February 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBb

[2417] Anonymous. Editorial Board. Science of Computer Programming, 186 (??):??, February 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301601.

Azadbakht:2020:FAB

[2418] Keyvan Azadbakht, Frank S. de Boer, Nikolaos Bezirgiannis, and Erik de Vink. A formal actor-based model for streaming the future. Science of Computer Programming, 186(??): ??, February 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301364.

Beohar:2020:CTS

[2419] Harsh Beohar, Barbara König, Sebastian Küpper, and Alexandra Silva. Conditional transition systems with upgrades. Science of Computer Programming, 186(??): ??, February 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301169.

Rusu:2020:PPC

[2420] Vlad Rusu, Gilles Grimaud, and Michaël Hauspie. Proving partialcorrectness and invariance properties of transition-system models. Science of Computer Programming, 186 (??):??, February 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301376.

Anonymous:2020:Fb

[2421] Anonymous. 15 February 2020. Science of Computer Programming, 187 (??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBc

[2422] Anonymous. Editorial Board. Science of Computer Programming, 187 (??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301741.

Basile:2020:CSS

[2423] Davide Basile, Maurice H. ter Beek, Pierpaolo Degano, Axel Legay, Gian-Luigi Ferrari, Stefania Gnesi, and Felicita Di Giandomenico. Controller synthesis of service contracts with variability. Science of Computer Programming, 187(??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302260.

BenSaid:2020:EEI

[2424] Najah Ben Said and Ioana Cristescu.

End-to-end information flow security for web services orchestration. Science of Computer Programming, 187 (??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301618.

DeNicola:2020:MAS

[2425] Rocco De Nicola, Luca Di Stefano, and Omar Inverso. Multi-agent systems with virtual stigmergy. Science of Computer Programming, 187 (??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764231930139X.

Zhang:2020:PSI

[2426] Chenyi Zhang and Jun Pang. Preface for the special issue of the 12th International Symposium on Theoretical Aspects of Software Engineering (TASE 2018). Science of Computer Programming, 187(??): ??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301686.

Xiang:2020:MVT

[2427] Shuangqing Xiang, Huibiao Zhu, Xi Wu, Lili Xiao, Marcello Bonsangue, Wanling Xie, and Lei Zhang. Modeling and verifying the topology discovery mechanism of Open-Flow controllers in software-defined networks using process algebra. Science of Computer Programming, 187 (??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319301388.

Kukovec:2020:EST

[2428] Jure Kukovec, Thanh-Hai Tran, and Igor Konnov. Extracting symbolic transitions from TLA^{+} specifications. Science of Computer Programming, 187(??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319301558.

Hamana:2020:PCS

[2429] Makoto Hamana, Tatsuya Abe, and Kentaro Kikuchi. Polymorphic computation systems: Theory and practice of confluence with call-by-value. Science of Computer Programming, 187 (??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319301182.

Wang:2020:FSI

[2430] Jiawei Wang, Ming Fu, Lei Qiao, and Xinyu Feng. Formalizing SPARCv8 instruction set architecture in Coq. Science of Computer Programming, 187 (??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318301163.

deMedeiros:2020:ASE

[2431] Sérgio Queiroz de Medeiros, Gilney de Azevedo Alvez Junior, and Fabio Mascarenhas. Automatic syntax error reporting and recovery in parsing expression grammars. Science of Computer Programming, 187(??): ??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301662.

Zhang:2020:OBF

[2432] Yueling Zhang, Min Zhang, and Geguang Pu. Optimizing backbone filtering. Science of Computer Programming, 187(??):??, February 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642319301674.

Anonymous:2020:Ma

[2433] Anonymous. 1 March 2020. Science of Computer Programming, 188 (??):??, March 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBd

[2434] Anonymous. Editorial Board. Science of Computer Programming, 188 (??):??, March 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300101.

Neele:2020:FCP

[2435] Thomas Neele, Tim A. C. Willemse, and Jan Friso Groote. Finding compact proofs for infinite-data parameterised Boolean equation systems. Science of Computer Programming, 188 (??):??, March 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301807.

Castiglioni:2020:LCD

[2436] Valentina Castiglioni, Konstantinos Chatzikokolakis, and Catuscia Palamidessi. A logical characterization of differential privacy. Science of Computer Programming, 188 (??):??, March 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301790.

Krings:2020:TAE

[2437] Sebastian Krings, Michael Leuschel, Joshua Schmidt, David Schneider, and Marc Frappier. Translating Alloy and extensions to classical B. Science of Computer Programming, 188 (??):??, March 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301704.

Anonymous:2020:Aa

[2438] Anonymous. 1 April 2020. Science of Computer Programming, 189 (??):??, April 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBe

[2439] Anonymous. Editorial Board. Science of Computer Programming, 189 (??):??, April 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300204.

Arruda:2020:ACA

[2440] Filipe Arruda, Flávia Barros, and Augusto Sampaio. Automation and consistency analysis of test cases written in natural language: an industrial context. Science of Computer Programming, 189(??): ??, April 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642319301698.

Zhu:2020:FHS

[2441] Chenyang Zhu, Michael Butler, and Corina Cirstea. Formalizing hierarchical scheduling for refinement of real-time systems. Science of Computer Programming, 189 (??):??, April 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300010.

Gualandi:2020:PCL

[2442] Hugo Musso Gualandi and Roberto Ierusalimschy. Pallene: a companion language for Lua. Science of Computer Programming, 189 (??):??, April 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320300046.

Anonymous:2020:Mb

[2443] Anonymous. 1 May 2020. Science of Computer Programming, 190(??):??, May 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBf

[2444] Anonymous. Editorial Board. Science of Computer Programming, 190 (??):??, May 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320300332.

Lopez-Gonzalez:2020:OLI

[2445] J. López-González and Juan M. Serrano. The optics of language-integrated query. Science of Computer Programming, 190(??):??, May 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764232030006X.

Jaiswal:2020:BFS

[2446] Swati Jaiswal, Uday P. Khedker, and Supratik Chakraborty. Bidirectionality in flow-sensitive demand-driven analysis. Science of Computer Programming, 190(??):??, May 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320300022.

Banach:2020:AUT

[2447] Richard Banach. Automated urban train control with hybrid Event-B: ‘tackling’ the rugby club problem. Science of Computer Programming, 190(??):??, May 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300150.

Anonymous:2020:Jb

[2448] Anonymous. 1 June 2020. Science of Computer Programming, 191(??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBg

[2449] Anonymous. Editorial Board. Science of Computer Programming, 191 (??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300447.

Ferrara:2020:CJB

[2450] Pietro Ferrara, Agostino Cortesi, and Fausto Spoto. From CIL to Java bytecode: Semantics-based translation for static analysis leveraging. Science of Computer Programming, 191(??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300034.

Pinheiro:2020:MCA

[2451] Pedro Pinheiro, José Carlos Viana, Márcio Ribeiro, Leo Fernandes, Fabiano Ferrari, Rohit Gheyi, and Baldoino Fonseca. Mutating code annotations: an empirical evaluation on Java and C# programs. Science of Computer Programming, 191 (??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300290.

Jacquet:2020:PSI

[2452] Jean-Marie Jacquet and Jacopo Soldani. Preface to the Special Issue on Foundations of Coordination Languages and Self-Adaptive Systems (Foclasa 2018). Science of Computer Programming, 191 (??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300289.

Pelliccione:2020:BCC

[2453] Patrizio Pelliccione, Eric Knauss, S. Magnus Ågren, Rogardt Heldal, Carl Bergenhem, Alexey Vinel, and Oliver Brunnegård. Beyond connected cars: a systems of systems perspective. Science of Computer Programming, 191(??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300253.

Tekinerdogan:2020:ARF

[2454] Bedir Tekinerdogan and Ferhat Erata. Automated reasoning framework for traceability management of system of systems. Science of Computer Programming, 191 (??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300277.

Cabrera:2020:MTE

[2455] Pedro Cabrera, Henrik Lund, Jakob Zinck Thellufsen, and Peter Sorknæs. The MATLAB Toolbox for EnergyPLAN: a tool to extend energy planning studies. Science of Computer Programming, 191(??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300162.

Dantas:2020:CBF

[2456] Allberson Bruno de Oliveira Dantas, Francisco Heron de Carvalho Junior, and Luis Soares Barbosa. A component-based framework for certification of components in a cloud of HPC services. Science of Computer Programming, 191 (??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300807.

Cavalheiro:2020:SPB

[2457] Simone Cavalheiro and José Luiz Fiadeiro. Selected papers from the Brazilian Symposium on Formal Methods (SBMF 2017). Science of Computer Programming, 191 (??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300381.

Sakaguchi:2020:PEM

[2458] Kazuhiko Sakaguchi. Program extraction for mutable arrays. Science of Computer Programming, 191 (??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642319301650.

Alcocer:2020:PVP

[2459] Juan Pablo Sandoval Alcocer, Alexandre Bergel, and Marco Tulio Valente. Prioritizing versions for performance regression testing: the Pharo case. Science of Computer Programming, 191(??):??, June 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320300265.

Anonymous:2020:Jc

[2460] Anonymous. 15 June 2020. Science of Computer Programming, 192 (??):??, June 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBh

[2461] Anonymous. Editorial Board. Science of Computer Programming, 192 (??):??, June 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320300563.

Alrahman:2020:PIC

[2462] Yehia Abd Alrahman, Rocco De Nicola, and Michele Loreti. Programming interactions in collective adaptive systems by relying on attribute-based communication. Science of Computer Programming, 192 (??):??, June 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300393.

Davies:2020:FSA

[2463] Jim Davies, James Welch, David Milward, and Steve Harris. A formal, scalable approach to semantic interoperability. Science of Computer Programming, 192(??): ??, June 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232030037X.

Li:2020:MCB

[2464] Yi Li, Weidi Sun, and Meng Sun. Mediator: a componentbased modeling language for concurrent and distributed systems. Science of Computer Programming, 192 (??):??, June 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300484.

Silva:2020:VMR

[2465] Eduardo Silva, Thais Batista, and Flavio Oquendo. On the verification of mission-related properties in software-intensive systems-of-systems architectural design. Science of Computer Programming, 192 (??):??, June 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320300368.

Anonymous:2020:Jd

[2466] Anonymous. 1 July 2020. Science of Computer Programming, 193(??):??, July 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBi

[2467] Anonymous. Editorial Board. Science of Computer Programming, 193 (??):??, July 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320300642.

Tran:2020:FAG

[2468] Hoang-Viet Tran, Pham Ngoc Hung, Viet-Ha Nguyen, and Toshiaki Aoki. A framework for assume-guarantee regression verification of evolving software. Science of Computer Programming, 193(??):??, July 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642320300496.

Chen:2020:IAO

[2469] Haiming Chen and Zhiwu Xu. Inclusion algorithms for one-unambiguous regular expressions and their applications. Science of Computer Programming, 193(??):??, July 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764231830296X.

Zhang:2020:CPE

[2470] Weixin Zhang and Bruno C. d. S. Oliveira.

Castor: Programming with extensible generative visitors. Science of Computer Programming, 193 (??):??, July 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300599.

Boreale:2020:CAA

[2471] Michele Boreale. Complete algorithms for algebraic strongest postconditions and weakest preconditions in polynomial ODEs. Science of Computer Programming, 193 (??):??, July 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300514.

Anonymous:2020:Ab

[2472] Anonymous. 1 August 2020. Science of Computer Programming, 194 (??):??, August 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBj

[2473] Anonymous. Editorial Board. Science of Computer Programming, 194 (??):??, August 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300745.

Tornblom:2020:FVI

[2474] John Törnblom and Simin Nadjm-Tehrani. Formal verification of inputoutput mappings of tree ensembles.

Science of Computer Programming, 194(??):??, August 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300605.

Kokke:2020:PLF

[2475] Wen Kokke, Jeremy G. Siek, and Philip Wadler. Programming language foundations in Agda. Science of Computer Programming, 194 (??):??, August 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300502.

Marra:2020:DAL

[2476] Matteo Marra, Guillermo Polito, and Elisa Gonzalez Boix. A debugging approach for live Big Data applications. Science of Computer Programming, 194(??):??, August 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300708.

Anonymous:2020:Sa

[2477] Anonymous. 1 September 2020. Science of Computer Programming, 195 (??):??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBk

[2478] Anonymous. Editorial Board. Science of Computer Programming, 195 (??):??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300964.

Silvasi:2020:LFB

[2479] Frantisek Silvási and Martin Tomásek.

Lean formalization of bounded grids and computable cellular automata defined thereover. Science of Computer Programming, 195(??): ??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300812.

Ribeiro:2020:PSP

[2480] Cristina Ribeiro and Daniel Berry.

The prevalence and severity of persistent ambiguity in software requirements specifications: Is a special effort needed to find them? Science of Computer Programming, 195 (??):??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300824.

Khatchadourian:2020:SAR

[2481] Raffi Khatchadourian, Yiming Tang, and Mehdi Bagherzadeh. Safe automated refactoring for intelligent parallelization of Java 8 streams. Science of Computer Programming, 195 (??):??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300861.

Fava:2020:RSG

[2482] Daniel Schnetzer Fava and Martin Steffen. Ready, set, go!: Data-race detection and the Go language. Science of Computer Programming, 195 (??):??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300836.

Almeida:2020:FDP

[2483] Ariane Alves Almeida and Mauricio Ayala-Rincón. Formalizing the dependency pair criterion for innermost termination. Science of Computer Programming, 195(??): ??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300848.

Tesone:2020:NMI

[2484] Pablo Tesone, Stéphane Ducasse, Guillermo Polito, Luc Fabresse, and Noury Bouraqadi. A new modular implementation for stateful traits. Science of Computer Programming, 195 (??):??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300800.

Alcocer:2020:ISR

[2485] Juan Pablo Sandoval Alcocer, Alejandra Siles Antezana, Gustavo Santos, and Alexandre Bergel. Improving the success rate of applying the extract method refactoring. Science of Computer Programming, 195(??): ??, September 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232030085X.

Anonymous:2020:Sb

[2486] Anonymous. 15 September 2020. Science of Computer Programming, 196 (??):??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBl

[2487] Anonymous. Editorial Board. Science of Computer Programming, 196 (??):??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301131.

Nunez-Varela:2020:DCS

[2488] Alberto S. Nuñez-Varela, Héctor G.

Pérez-Gonzalez, Francisco E. Martínez-Perez, Carlos Soubervielle-Montalvo, and Oscar E. Perez-Cham. Dynamic creation of source code models for the extraction of code metrics data through grammar querying. Science of Computer Programming, 196(??): ??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300903.

Valloud:2020:LWR

[2489] Andrés Valloud. Linked weak reference arrays: a hybrid approach to efficient bulk finalization. Science of Computer Programming, 196(??): ??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300897.

Halchin:2020:HBM

[2490] Alexandra Halchin, Yamine Ait-Ameur, Neeraj Kumar Singh, Julien Ordioni, and Abderrahmane Feliachi. Handling B models in the PERF integrated verification framework: Formalised and certified embedding. Science of Computer Programming, 196 (??):??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300873.

Muller:2020:SOF

[2491] Sascha Müller, Liana Mikaelyan, Andreas Gerndt, and Thomas Noll. Synthesizing and optimizing FDIR recovery strategies from fault trees. Science of Computer Programming, 196 (??):??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300885.

Feitosa:2020:TDA

[2492] Samuel Feitosa, Rodrigo Ribeiro, and Andre Du Bois. A type-directed algorithm to generate random welltyped Java 8 programs. Science of Computer Programming, 196(??): ??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301039.

dePutter:2020:CMC

[2493] Sander de Putter, Frédéric Lang, and Anton Wijs. Compositional model checking with divergence preserving branching bisimilarity is lively. Science of Computer Programming, 196 (??):??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318300789.

Santos:2020:RM

[2494] Renan Santos, Noemi Rodriguez, and Roberto Ierusalimschy. Revisiting monitors. Science of Computer Programming, 196(??):??, September 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642320301015.

Anonymous:2020:Oa

[2495] Anonymous. 1 October 2020. Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBm

[2496] Anonymous. Editorial Board. Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301362.

Butler:2020:ASM

[2497] Michael Butler and Alexander Raschke.

Abstract state machines, Alloy, B, TLA, VDM and Z (ABZ 2018). Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301222.

Autili:2020:CSC

[2498] Marco Autili, Amleto Di Salle, Francesco Gallo, Claudio Pompilio, and Massimo Tivoli. CHOReVO-LUTION: Service choreography in practice. Science of Computer Programming, 197(??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642320301076.

Choi:2020:PRC

[2499] Kwanghoon Choi, James Cheney, Simon Fowler, and Sam Lindley. A polymorphic RPC calculus. Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301088.

Lima:2020:FVD

[2500] Lucas Lima, Amaury Tavares, and Sidney C. Nogueira. A framework for verifying deadlock and nondeterminism in UML activity diagrams based on CSP. Science of Computer Programming, 197(??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301064.

Vidal:2020:DSL

[2501] Marzina Vidal, Tiago Massoni, and Franklin Ramalho. A domainspecific language for verifying software requirement constraints. Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301180.

Toro:2020:AGR

[2502] Matías Toro and Éric Tanter. Abstracting gradual references. Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301052.

Ancona:2020:BSF

[2503] Davide Ancona, Francesco Dagnino, Jurriaan Rot, and Elena Zucca. A big step from finite to infinite computations. Science of Computer Programming, 197(??): ??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301027.

Tasos:2020:RYL

[2504] Alexandros Tasos, Juliana Franco, Sophia Drossopoulou, Tobias Wrigstad, and Susan Eisenbach. Reshape your layouts, not your programs: a safe language extension for better cache locality. Science of Computer Programming, 197(??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320300915.

Zhu:2020:TSR

[2505] Chenyang Zhu, Michael Butler, and Corina Cirstea. Trace semantics and refinement patterns for real-time properties in event-B models. Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301210.

Gadea:2020:MCA

[2506] Alejandro Gadea, Emmanuel Gunther, and Miguel Pagano. Mechanization of coherence and adequacy: Being extrinsic extended to subtyping. Science of Computer Programming, 197 (??):??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642318302442.

Foster:2020:USF

[2507] Simon Foster, James Baxter, Ana Cavalcanti, Jim Woodcock, and Frank Zeyda. Unifying semantic foundations for automated verification tools in Isabelle/UTP. Science of Computer Programming, 197(??): ??, October 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301192.

Anonymous:2020:Ob

[2508] Anonymous. 15 October 2020. Science of Computer Programming, 198 (??):??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBn

[2509] Anonymous. Editorial Board. Science of Computer Programming, 198 (??):??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301490.

Gamaza:2020:RRA

[2510] Ángel Gamaza, Guadalupe Ortiz, Juan Boubeta-Puig, and Alfonso Garcia de Prado. REST4CEP: RESTful APIs for complex event processing. Science of Computer Programming, 198 (??):??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301234.

Dhumbumroong:2020:BTE

[2511] Smith Dhumbumroong and Krerk Piromsopa. BoundWarden: Threadenforced spatial memory safety through compile-time transformations. Science of Computer Programming, 198 (??):??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301271.

Bednarik:2020:EEM

[2512] Roman Bednarik, Teresa Busjahn, Agostino Gibaldi, Alireza Ahadi, Maria Bielikova, Martha Crosby, Kai Essig, Fabian Fagerholm, Ahmad Jbara, Raymond Lister, Pavel Orlov, James Paterson, Bonita Sharif, Teemu SirkiÃ, Jan Stelovsky, Jozef Tvarozek, Hana Vrzakova, and Ian van der Linde. EMIP: the eye movements in programming dataset. Science of Computer Programming, 198(??): ??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301283.

Siirtola:2020:OCA

[2513] Antti Siirtola and Keijo Heljanko. An optimal cut-off algorithm for parameterised refinement checking. Science of Computer Programming, 198 (??):??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301258.

Ribeiro:2020:UCE

[2514] Leila Ribeiro, Lucio Mauro Duarte, Rodrigo Machado, Andrei Costa, Érika Fernandes Cota, and Jonas Santos Bezerra. Use case evolution analysis based on graph transformation with negative application conditions. Science of Computer Programming, 198 (??):??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301040.

Mallet:2020:ETA

[2515] Frederic Mallet and Min Zhang. Editorial — Theoretical Aspects of Software Engineering (2017). Science of Computer Programming, 198(??): ??, October 15, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301295.

Anonymous:2020:N

[2516] Anonymous. 1 November 2020. Science of Computer Programming, 199 (??):??, November 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBo

[2517] Anonymous. Editorial Board. Science of Computer Programming, 199 (??):??, November 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301593.

Meldrum:2020:USO

[2518] Sarah Meldrum, Sherlock A. Licorish, Caitlin A. Owen, and Bastin Tony Roy Savarimuthu. Understanding Stack Overflow code quality: a recommendation of caution. Science of Computer Programming, 199(??): ??, November 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301246.

Bodeveix:2020:EBF

[2519] Jean-Paul Bodeveix, Arnaud Dieumegard, and Mamoun Filali. Event-B formalization of a variability-aware component model patterns framework. Science of Computer Programming, 199 (??):??, November 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301209.

Lagadec:2020:AST

[2520] Loïc Lagadec, Anne Etien, and Jannik Laval. Advances in Smalltalk technologies. Science of Computer Programming, 199(??):??, November 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S016764232030126X.

Lecomte:2020:CSP

[2521] Thierry Lecomte, David Deharbe, Paulin Fournier, and Marcel Oliveira. The CLEARSY safety platform: 5 years of research, development and deployment. Science of Computer Programming, 199(??):??, November 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642320301325.

Anonymous:2020:D

[2522] Anonymous. 1 December 2020. Science of Computer Programming, 200 (??):??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2020:EBp

[2523] Anonymous. Editorial Board. Science of Computer Programming, 200 (??):??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301684.

Akiki:2020:EME

[2524] Pierre A. Akiki, Paul A. Akiki, Arosha K. Bandara, and Yijun Yu. EUD-MARS: End-user development of model-driven adaptive robotics software systems. Science of Computer Programming, 200(??): ??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301428.

Freitas:2020:MWP

[2525] Leo Freitas, William E. Scott, and Patrick Degenaar. Medicineby-wire: Practical considerations on formal techniques for dependable medical systems. Science of Computer Programming, 200(??): ??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301532.

Camps:2020:SAB

[2526] Valérie Camps, Stéphanie Combettes, and Teddy Bouziat. SApH-ESIA: an agent-based model and a criticality-based heuristic for cooperatively coupling SoSs. Science of Computer Programming, 200(??): ??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301416.

Chen:2020:ECF

[2527] Sheng Chen and Baijun Wu. Efficient counter-factual type error debugging. Science of Computer Programming, 200(??):??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301520.

Ugliara:2020:TRU

[2528] Fellipe A. Ugliara, Gustavo M. D. Vieira, and José de O. Guimarães. Transparent replication using metaprogramming in Cyan. Science of Computer Programming, 200(??): ??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301398.

Timm:2020:MCS

[2529] Nils Timm, Stefan Gruner, Madoda Nxumalo, and Josua Botha. Model checking safety and liveness via kinduction and witness refinement with constraint generation. Science of Computer Programming, 200(??): ??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301404.

Bae:2020:FAC

[2530] Kyungmin Bae and Peter Csaba Ölveczky. Formal aspects of component software (FACS 2018). Science of Computer Programming, 200 (??):??, December 1, 2020. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301556.

Anonymous:2021:Ja

[2531] Anonymous. 1 January 2021. Science of Computer Programming, 201 (??):??, January 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBa

[2532] Anonymous. Editorial Board. Science of Computer Programming, 201 (??):??, January 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301829.

Figueroa:2021:WMH

[2533] Ismael Figueroa, Paul Leger, and Hiroaki Fukuda. Which monads Haskell developers use: an exploratory study. Science of Computer Programming, 201(??):??, January 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301313.

Carvalho:2021:VVT

[2534] Gustavo Carvalho and Igor Meira.

Validating, verifying and testing timed data-flow reactive systems in Coq from controlled naturallanguage requirements. Science of Computer Programming, 201(??): ??, January 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301453.

Duran:2021:PBS

[2535] Adolfo Duran and Philip Wadler. Preface — 22nd Brazilian Symposium on Formal Methods — SBMF 2019. Science of Computer Programming, 201 (??):??, January 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301738.

Yan:2021:ETG

[2536] Jiwei Yan, Hao Zhou, Xi Deng, Ping Wang, Rongjie Yan, Jun Yan, and Jian Zhang. Efficient testing of GUI applications by event sequence reduction. Science of Computer Programming, 201 (??):??, January 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301301.

Correia:2021:FMC

[2537] Alexandre Correia, Juliano Iyoda, and Alexandre Mota. A family of multiconcept program synthesisers in Alloy. Science of Computer Programming, 201(??):??, January 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301441.

Anonymous:2021:F

[2538] Anonymous. 1 February 2021. Science of Computer Programming, 202 (??):??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBb

[2539] Anonymous. Editorial Board. Science of Computer Programming, 202 (??):??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301969.

Gao:2021:ARD

[2540] Shanquan Gao, Lei Liu, Yuzhou Liu, Huaxiao Liu, and Yihui Wang. API recommendation for the development of Android App features based on the knowledge mined from App stores. Science of Computer Programming, 202 (??):??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301647.

Lopez:2021:QQA

[2541] L. López, M. Manzano, C. Gómez, M. Oriol, C. Farré, X. Franch, S. Martínez-Fernández, and A. M. Vollmer. QaSD: a quality-aware strategic dashboard for supporting decision makers in agile software development. Science of Computer Programming, 202(??):??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301763.

Zhang:2021:CBDa

[2542] Yuanrui Zhang, Frédéric Mallet, Huibiao Zhu, Yixiang Chen, Bo Liu, and Zhiming Liu. A clock-based dynamic logic for schedulability analysis of CCSL specifications. Science of Computer Programming, 202 (??):??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301544.

Gallagher:2021:P

[2543] John P. Gallagher and Martin Sulzmann. Preface. Science of Computer Programming, 202(??): ??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301854.

DeNicola:2021:PCI

[2544] Rocco De Nicola, Tan Duong, and Michele Loreti. Provably correct implementation of the abc calculus. Science of Computer Programming, 202 (??):??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301751.

Jacquet:2021:AWM

[2545] Jean-Marie Jacquet and Manel Barkallah. Anemone: a workbench for the Multi-Bach coordination language. Science of Computer Programming, 202 (??):??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642320301878.

Guanciale:2021:TCC

[2546] Roberto Guanciale and Emilio Tuosto. PomCho: a tool chain for choreographic design. Science of Computer Programming, 202(??):

??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232030143X.

Salaun:2021:QSN

[2547] Gwen Salaün. Quantifying the similarity of non-bisimilar labelled transition systems. Science of Computer Programming, 202(??): ??, February 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232030188X.

Anonymous:2021:Ma

[2548] Anonymous. 1 March 2021. Science of Computer Programming, 203 (??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBc

[2549] Anonymous. Editorial Board. Science of Computer Programming, 203 (??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320302112.

Kaplar:2021:NWE

[2550] Sebastijan Kaplar, Miroslav Zarić, and Stéphane Ducasse. NewWave: Workflow engine. Science of Computer Programming, 203(??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642320301891.

Changizi:2021:CWM

[2551] Behnaz Changizi, Natallia Kokash, Farhad Arbab, and Leonid Makhnist. Compositional workflow modeling with priority constraints. Science of Computer Programming, 203 (??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301866.

Cledou:2021:HVO

[2552] Guillermina Cledou, José Proença, Bernhard H. C. Sputh, and Eric Verhulst. Hubs for VirtuosoNext: Online verification of real-time coordinators. Science of Computer Programming, 203(??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232030174X.

Audrito:2021:ACM

[2553] Giorgio Audrito, Danilo Pianini, Ferruccio Damiani, and Mirko Viroli. Aggregate centrality measures for IoT-based coordination. Science of Computer Programming, 203 (??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301921.

Chardet:2021:TSE

[2554] Maverick Chardet, Hélène Coullon, and Simon Robillard. Toward safe and efficient reconfiguration with Concerto. Science of Computer Programming, 203(??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301908.

Perez-Vereda:2021:MDA

[2555] Alejandro Pérez-Vereda, Carlos Canal, and Ernesto Pimentel. Modelling digital avatars: a tuple space approach. Science of Computer Programming, 203(??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232030191X.

Zhang:2021:CBDb

[2556] Yuanrui Zhang, Hengyang Wu, Yixiang Chen, and Frédéric Mallet. A clock-based dynamic logic for the verification of CCSL specifications in synchronous systems. Science of Computer Programming, 203 (??):??, March 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320301994.

Anonymous:2021:Aa

[2557] Anonymous. 1 April 2021. Science of Computer Programming, 204 (??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBd

[2558] Anonymous. Editorial Board. Science of Computer Programming, 204 (??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000083.

Mehrotra:2021:HBN

[2559] Gaurav Mehrotra and Daniel M. Berry. How to benefit from newbies’ domain ignorance in software development projects. Science of Computer Programming, 204 (??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232030201X.

Diaz:2021:RES

[2560] Eduardo Díaz, José Ignacio Panach, Silvia Rueda, Marcela Ruiz, and Oscar Pastor. Are requirements elicitation sessions influenced by participants’ gender? An empirical experiment. Science of Computer Programming, 204(??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320302033.

Azzeh:2021:PSE

[2561] Mohammad Azzeh, Ali Bou Nassif, and Imtinan Basem Attili. Predicting software effort from use case points: a systematic review. Science of Computer Programming, 204 (??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320302045.

Inverso:2021:PSI

[2562] Omar Inverso and Hugo Torres Vieira. Preface for the special issue on tool papers of the 21st International Conference on Coordination Models and Languages, COORDINATION 2019. Science of Computer Programming, 204(??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320302008.

Pimentel:2021:PSI

[2563] Ernesto Pimentel and Jacopo Soldani. Preface to the special issue on coordination and self-adaptiveness of software applications. Science of Computer Programming, 204 (??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320302021.

Wasser:2021:AME

[2564] Nathan Wasser, Asmae Heydari Tabar, and Reiner Hähnle. Automated model extraction: From non-deterministic C code to active objects. Science of Computer Programming, 204 (??):??, April 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642320302057.

Anonymous:2021:Mb

[2565] Anonymous. 1 May 2021. Science of Computer Programming, 205(??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBe

[2566] Anonymous. Editorial Board. Science of Computer Programming, 205 (??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000290.

Ancona:2021:RTP

[2567] Davide Ancona, Luca Franceschini, Angelo Ferrando, and Viviana Mascardi. RML: Theory and practice of a domain specific language for runtime verification. Science of Computer Programming, 205 (??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000034.

Pereira:2021:RPL

[2568] Rui Pereira, Marco Couto, Francisco Ribeiro, Rui Rua, Jácome Cunha, João Paulo Fernandes, and João Saraiva. Ranking programming languages by energy efficiency. Science of Computer Programming, 205 (??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000022.

Ouederni:2021:CCA

[2569] Meriem Ouederni. Compatibility checking for asynchronously communicating software. Science of Computer Programming, 205(??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642320301775.

Pecorelli:2021:ASC

[2570] Fabiano Pecorelli and Dario Di Nucci.

Adaptive selection of classifiers for bug prediction: a large-scale empirical analysis of its performances and a benchmark study. Science of Computer Programming, 205 (??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000046.

Ali:2021:ISE

[2571] Asad Ali and Carmine Gravino. Improving software effort estimation using bio-inspired algorithms to select relevant features: an empirical study. Science of Computer Programming, 205(??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000149.

Massoni:2021:PSI

[2572] Tiago Massoni and Mohammad Mousavi. Preface to the special issue on Formal Methods: Foundations and Applications. Science of Computer Programming, 205 (??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000125.

Kallis:2021:PIT

[2573] Rafael Kallis, Andrea Di Sorbo, Gerardo Canfora, and Sebastiano Panichella. Predicting issue types on GitHub. Science of Computer Programming, 205(??):??, May 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642320302069.

Anonymous:2021:Jb

[2574] Anonymous. 1 June 2021. Science of Computer Programming, 206(??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBf

[2575] Anonymous. Editorial Board. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000423.

Hu:2021:VCC

[2576] Kai Hu, Teng Zhang, Yi Ding, Jian Zhu, and Jean-Pierre Talpin. Verification of concurrent code from synchronous specifications. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000186.

Kammuller:2021:AII

[2577] Florian Kammüller and Manfred Kerber. Applying the Isabelle Insider framework to airplane security. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000162.

Chondamrongkul:2021:FSA

[2578] Nacha Chondamrongkul, Jing Sun, and Ian Warren. Formal security analysis for software architecture design: an expressive framework to emerging architectural styles. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000241.

Ge:2021:RMC

[2579] Ning Ge, Silvano Dal Zilio, Hongyu Liu, Li Zhang, and Lianyi Zhang. RT-MOBS: a compositional observer semantics of time Petri net for real-time property specification language based on μ-calculus. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000174.

Larsen:2021:PSI

[2580] Kim G. Larsen, Oleg Sokolsky, and Ji Wang. Preface to the special issue on Dependable Software Engineering: Theories, Tools and Applications (SETTA 2017). Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000150.

LYvonnet:2021:PMC

[2581] Thibaud L’Yvonnet, Elisabetta De Maria, Sabine Moisan, and Jean-Paul Rigault. Probabilistic model checking for human activity recognition in medical serious games. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000228.

Hojjat:2021:FSE

[2582] Hossein Hojjat and Mieke Massink. Fundamentals of Software Engineering (extended versions of selected papers of FSEN 2019). Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000198.

Belzner:2021:SSP

[2583] Lenz Belzner and Martin Wirsing. Synthesizing safe policies under probabilistic constraints with reinforcement learning and Bayesian model checking. Science of Computer Programming, 206(??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000137.

He:2021:ATG

[2584] Weigang He, Jianqi Shi, Ting Su, Zeyu Lu, Li Hao, and Yanhong Huang. Automated test generation for IEC 61131-3 ST programs via dynamic symbolic execution. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000010.

Shimari:2021:NNO

[2585] Kazumasa Shimari, Takashi Ishio, Tetsuya Kanda, Naoto Ishida, and Katsuro Inoue. NOD4J: Nearomniscient debugging tool for Java using size-limited execution trace. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100023X.

Edwards:2021:SEI

[2586] Nystrom Edwards, Dhitiwat Jongsuebchoke, and Tim Storer. Sciit: Embedding issue tracking in source control management. Science of Computer Programming, 206(??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642321000216.

Ravelo-Mendez:2021:KFE

[2587] William Ravelo-Méndez, Camilo Escobar-Velásquez, and Mario Linares-Vásquez. Kraken: a framework for enabling multi-device interaction-based testing of Android apps. Science of Computer Programming, 206 (??):??, June 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000204.

Anonymous:2021:Jc

[2588] Anonymous. 1 July 2021. Science of Computer Programming, 207(??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBg

[2589] Anonymous. Editorial Board. Science of Computer Programming, 207 (??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000551.

Guzman-Guzman:2021:SND

[2590] Xiomarah Guzmán-Guzmán, Edward Rolando Núñez-Valdez, Raysa Vásquez-Reynoso, Angel Asencio, and Vicente García-Díaz. SWQL: a new domain-specific language for mining the social Web. Science of Computer Programming, 207 (??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000356.

Bardin:2021:SMC

[2591] Sébastien Bardin, Nikolai Kosmatov, Michaël Marcozzi, and Mickaël Delahaye. Specify and measure, cover and reveal: a unified framework for automated test generation. Science of Computer Programming, 207(??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000344.

Yang:2021:MTA

[2592] Zhibin Yang, Shenghao Yuan, Jean-Paul Bodeveix, Mamoun Filali, Tiexin Wang, and Yong Zhou. Multi-task Ada code generation from synchronous dataflow programs on multi-core: Approach and industrial study. Science of Computer Programming, 207 (??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100037X.

Zerouali:2021:UJP

[2593] Ahmed Zerouali, Tom Mens, and Coen De Roover. On the usage of javascript, Python and Ruby packages in Docker Hub images. Science of Computer Programming, 207 (??):??, July 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000460.

Anonymous:2021:Ab

[2594] Anonymous. 1 August 2021. Science of Computer Programming, 208 (??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBh

[2595] Anonymous. Editorial Board. Science of Computer Programming, 208 (??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000666.

Busi:2021:MIT

[2596] Matteo Busi, Pierpaolo Degano, and Letterio Galletta. Mechanical incrementalization of typing algorithms. Science of Computer Programming, 208(??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000502.

Manjhi:2021:REI

[2597] Devesh Manjhi and Amrita Chaturvedi. Reuse estimate and interval prediction using MOGA-NN and RBF-NN in the functional paradigm. Science of Computer Programming, 208 (??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000368.

Cetin:2021:RCR

[2598] H. Alperen Çetin, Emre Dogan, and Eray Tüzün. A review of code reviewer recommendation studies: Challenges and future directions. Science of Computer Programming, 208 (??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000459.

Lethbridge:2021:UMD

[2599] Timothy C. Lethbridge, Andrew Forward, Omar Badreddin, Dusan Brestovansky, Miguel Garzon, Hamoud Aljamaan, Sultan Eid, Ahmed Husseini Orabi, Mahmoud Husseini Orabi, Vahdat Abdelzad, Opeyemi Adesina, Aliaa Alghamdi, Abdulaziz Algablan, and Amid Zakariapour. Umple: Model-driven development for open source and education. Science of Computer Programming, 208(??): ??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000587.

Decan:2021:LZS

[2600] Alexandre Decan and Tom Mens.

Lost in zero space — an empirical comparison of 0.y.z releases in software package distributions. Science of Computer Programming, 208 (??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000496.

Turker:2021:MCS

[2601] Uraz Cengiz Türker, Robert M. Hierons, and Guy-Vincent Jourdan. Minimizing characterizing sets. Science of Computer Programming, 208 (??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000381.

Xue:2021:DTC

[2602] Mingqi Xue and Bruno C. d. S. Oliveira.

A dependently typed calculus with polymorphic subtyping. Science of Computer Programming, 208(??): ??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000484.

Artho:2021:FTS

[2603] Cyrille Artho and Peter Csaba Ölveczky. Formal Techniques for Safety-Critical Systems (FTSCS 2018). Science of Computer Programming, 208(??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000514.

SaardeMoraes:2021:AMV

[2604] Rodrigo Saar de Moraes and Simin Nadjm-Tehrani. Abstraction models for verifying resource adequacy of IMA systems at concept level. Science of Computer Programming, 208 (??):??, August 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000472.

Anonymous:2021:S

[2605] Anonymous. 1 September 2021. Science of Computer Programming, 209 (??):??, September 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBi

[2606] Anonymous. Editorial Board. Science of Computer Programming, 209 (??):??, September 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000770.

Chen:2021:LIB

[2607] Wei Chen. Loop invariance with break and continue. Science of Computer Programming, 209(??): ??, September 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000721.

Hasan:2021:PF

[2608] Osman Hasan and Frédéric Mallet.

Preface — FTSCS 2019. Science of Computer Programming, 209(??): ??, September 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000629.

vandenBos:2021:SIL

[2609] Petra van den Bos and Frits Vaandrager. State identification for labeled transition systems with inputs and outputs. Science of Computer Programming, 209(??): ??, September 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100071X.

Anonymous:2021:O

[2610] Anonymous. 1 October 2021. Science of Computer Programming, 210 (??):??, October 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBj

[2611] Anonymous. Editorial Board. Science of Computer Programming, 210 (??):??, October 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000927.

Guerin:2021:PAB

[2612] Sylvain Guérin, Guillaume Polet, Caine Silva, Joel Champeau, Jean-Christophe Bach, Salvador Martínez, Fabien Dagnat, and Antoine Beugnard. PAMELA: an annotationbased Java modeling framework. Science of Computer Programming, 210 (??):??, October 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000617.

Zeller:2021:CSE

[2613] Peter Zeller, Annette Bieniusa, and Arnd Poetzsch-Heffter. Combining state- and event-based semantics to verify highly available applications. Science of Computer Programming, 210(??):??, October 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000800.

Schewe:2021:LRA

[2614] Klaus-Dieter Schewe, Flavio Ferrarotti, and Senén González. A logic for reflective ASMs. Science of Computer Programming, 210(??): ??, October 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000848.

Anonymous:2021:N

[2615] Anonymous. 1 November 2021. Science of Computer Programming, 211 (??):??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBk

[2616] Anonymous. Editorial Board. Science of Computer Programming, 211 (??):??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001039.

Ssanyu:2021:PCF

[2617] Jackline Ssanyu, Engineer Bainomugisha, and Benjamin Kanagwa. PAMOJA: a component framework for grammar-aware engineering. Science of Computer Programming, 211(??): ??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000964.

Liebrenz:2021:SOD

[2618] Timm Liebrenz, Paula Herber, and Sabine Glesner. Service-oriented decomposition and verification of hybrid system models using feature models and contracts. Science of Computer Programming, 211(??): ??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000873.

Macedo:2021:ETA

[2619] Nuno Macedo, Alcino Cunha, José Pereira, Renato Carvalho, Ricardo Silva, Ana C. R. Paiva, Miguel Sozinho Ramalho, and Daniel Silva. Experiences on teaching alloy with an automated assessment platform. Science of Computer Programming, 211 (??):??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000836.

Terven:2021:KAK

[2620] Juan R. Terven and Diana M. Córdova-Esparza. KinZ: an Azure Kinect toolkit for Python and Matlab. Science of Computer Programming, 211 (??):??, November 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000952.

Anonymous:2021:D

[2621] Anonymous. 1 December 2021. Science of Computer Programming, 212 (??):??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2021:EBl

[2622] Anonymous. Editorial Board. Science of Computer Programming, 212 (??):??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764232100112X.

Lucia:2021:UET

[2623] Andrea De Lucia and Mohammad Reza Mousavi. Update from the Editorial Team. Science of Computer Programming, 212(??): ??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100085X.

Jain:2021:IPH

[2624] Shivani Jain and Anju Saha. Improving performance with hybrid feature selection and ensemble machine learning techniques for code smell detection. Science of Computer Programming, 212(??):??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001064.

Hayashi:2021:SIS

[2625] Shinpei Hayashi and Michael L. Collard. Special issue on Software Maintenance Tools at 35th International Conference on Software Maintenance and Evolution (ICSME 2019). Science of Computer Programming, 212 (??):??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100099X.

Drira:2021:SIS

[2626] Khalil Drira and Carlos E. Cuesta.

Software-intensive Systems-of-Systems. Science of Computer Programming, 212(??):??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000988.

Miller:2021:FTS

[2627] D. Michael Miller and Gerhard W. Dueck. Function translations and search-based transformation for MVL reversible circuit synthesis. Science of Computer Programming, 212(??): ??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321000976.

Boykett:2021:NMR

[2628] Tim Boykett. Notes on the maximality of reversible gate sets under borrow and ancilla closure. Science of Computer Programming, 212 (??):??, December 1, 2021. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001076.

Anonymous:2022:Ja

[2629] Anonymous. 1 January 2022. Science of Computer Programming, 213 (??):??, January 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBa

[2630] Anonymous. Editorial Board. Science of Computer Programming, 213 (??):??, January 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001313.

Ali:2022:EIF

[2631] Asad Ali and Carmine Gravino. Evaluating the impact of feature selection consistency in software prediction. Science of Computer Programming, 213 (??):??, January 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001088.

Dimovski:2022:SLA

[2632] Aleksandar S. Dimovski, Sven Apel, and Axel Legay. Several lifted abstract domains for static analysis of numerical program families. Science of Computer Programming, 213 (??):??, January 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001180.

AlOmar:2022:SMT

[2633] Eman Abdullah AlOmar, Ben Christians, Mihal Busho, Ahmed Hamad AlKhalid, Ali Ouni, Christian Newman, and Mohamed Wiem Mkaouer. SATDBailiff-mining and tracking selfadmitted technical debt. Science of Computer Programming, 213(??): ??, January 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321000861.

Kristiansen:2022:RCI

[2634] Lars Kristiansen. Reversible computing and implicit computational complexity. Science of Computer Programming, 213(??):??, January 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001167.

vanVulpen:2022:OPM

[2635] Paul van Vulpen, Slinger Jansen, and Sjaak Brinkkemper. The orchestrator’s partner management framework for software ecosystems. Science of Computer Programming, 213 (??):??, January 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001155.

Anonymous:2022:F

[2636] Anonymous. 1 February 2022. Science of Computer Programming, 214 (??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBb

[2637] Anonymous. Editorial Board. Science of Computer Programming, 214 (??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001465.

Tang:2022:AEF

[2638] Yiming Tang, Allan Spektor, Raffi Khatchadourian, and Mehdi Bagherzadeh. Automated evolution of feature logging statement levels using Git histories and degree of interest. Science of Computer Programming, 214 (??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001179.

Huang:2022:CTS

[2639] Rubing Huang, Haibo Chen, Weifeng Sun, and Dave Towey. Candidate test set reduction for adaptive random testing: an overheads reduction technique. Science of Computer Programming, 214(??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001234.

Alqasrawi:2022:LWR

[2640] Yousef Alqasrawi, Mohammad Azzeh, and Yousef Elsheikh. Locally weighted regression with different kernel smoothers for software effort estimation. Science of Computer Programming, 214(??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642321001374.

Dejanovic:2022:PLG

[2641] Igor Dejanović. Parglare: a LR/ GLR parser for Python. Science of Computer Programming, 214(??): ??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001271.

Cai:2022:MQA

[2642] Cheng-Hao Cai, Jing Sun, and Gillian Dobbie. B model quality assessments on automated reachability repair with ISO/IEC 25010. Science of Computer Programming, 214(??): ??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001258.

deBrugiere:2022:DTA

[2643] Timothée Goubault de Brugière, Marc Baboulin, Benoît Valiron, Simon Martiel, and Cyril Allouche. Decoding techniques applied to the compilation of CNOT circuits for NISQ architectures. Science of Computer Programming, 214(??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001192.

Stefano:2022:ISC

[2644] Manuel De Stefano, Emanuele Iannone, Fabiano Pecorelli, and Damian Andrew Tamburri. Impacts of software community patterns on process and product: an empirical study. Science of Computer Programming, 214 (??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001246.

Beelen:2022:YVM

[2645] Luc Beelen, Slinger Jansen, and Sietse Overbeek. Are you of value to me? A partner selection reference method for software ecosystem orchestrators. Science of Computer Programming, 214 (??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100126X.

Pascual:2022:TCP

[2646] Romain Pascual, Pascale Le Gall, Agnès Arnould, and Hakim Belhaouari. Topological consistency preservation with graph transformation schemes. Science of Computer Programming, 214(??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001210.

Kosiol:2022:SIG

[2647] Jens Kosiol, Daniel Strüber, Gabriele Taentzer, and Steffen Zschaler. Sustaining and improving graduated graph consistency: a static analysis of graph transformations. Science of Computer Programming, 214 (??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001222.

Campbell:2022:FRB

[2648] Graham Campbell, Brian Courtehoute, and Detlef Plump. Fast rule-based graph programs. Science of Computer Programming, 214 (??):??, February 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001209.

Anonymous:2022:Ma

[2649] Anonymous. 1 March 2022. Science of Computer Programming, 215 (??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBc

[2650] Anonymous. Editorial Board. Science of Computer Programming, 215 (??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000053.

Huang:2022:NND

[2651] Rubing Huang, Weifeng Sun, Haibo Chen, Chenhui Cui, and Ning Yang. A nearest-neighbor divide-and-conquer approach for adaptive random testing. Science of Computer Programming, 215(??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001362.

Dvorak:2022:TRT

[2652] Ondrej Dvorák and Robert Pergl. Tackling rapid technology changes by applying enterprise engineering theories. Science of Computer Programming, 215(??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642321001404.

Archibald:2022:MVB

[2653] Blair Archibald, Muffy Calder, Michele Sevegnani, and Mengwei Xu. Modelling and verifying BDI agents with bigraphs. Science of Computer Programming, 215(??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642321001532.

Zhang:2022:QIO

[2654] Qiang Zhang, Lei Xu, Xiangyu Zhang, and Baowen Xu. Quantifying the interpretation overhead of Python. Science of Computer Programming, 215 (??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001520.

Mao:2022:RDA

[2655] Xia Mao, Yueling Zhang, Jianqi Shi, Yanhong Huang, and Qin Li. A refinement development approach for enhancing the safety of PLC programs with Event-B. Science of Computer Programming, 215 (??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001568.

Clariso:2022:UDD

[2656] Robert Clarisó and Jordi Cabot.

User-driven diverse scenario exploration in model finders. Science of Computer Programming, 215(??): ??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001386.

Mogensen:2022:HRL

[2657] Torben Ægidius Mogensen. Hermes:

a reversible language for lightweight encryption. Science of Computer Programming, 215(??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642321001398.

Nakano:2022:TST

[2658] Keisuke Nakano. Time-symmetric Turing machines for computable involutions. Science of Computer Programming, 215(??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642321001416.

Yan:2022:TCP

[2659] Rongjie Yan, Yuhang Chen, Hongyu Gao, and Jun Yan. Test case prioritization with neuron valuation based pattern. Science of Computer Programming, 215(??): ??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001544.

Zhang:2022:MCS

[2660] Chenyi Zhang and Jun Pang. Modal characterisation of simulation relations in probabilistic concurrent games. Science of Computer Programming, 215 (??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642321001556.

Neto:2022:TLF

[2661] João Batista de Souza Neto, Anamaria Martins Moreira, Genoveva Vargas-Solar, and Martin A. Musicante. A two-level formal model for Big Data processing programs. Science of Computer Programming, 215 (??):??, March 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100157X.

Anonymous:2022:Aa

[2662] Anonymous. 1 April 2022. Science of Computer Programming, 216 (??):??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBd

[2663] Anonymous. Editorial Board. Science of Computer Programming, 216 (??):??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000193.

Hesselink:2022:TCT

[2664] Wim H. Hesselink. Trylock, a case for temporal logic and eternity variables. Science of Computer Programming, 216(??):??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232100160X.

Mery:2022:SPR

[2665] Dominique Méry and Alexander Raschke. Selected papers from the Rigorous State-Based Methods 7th International Conference, ABZ 2020, Ulm, Germany, May 27–29, 2020. Science of Computer Programming, 216(??):??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000132.

Dupont:2022:FVA

[2666] Guillaume Dupont, Yamine Ait-Ameur, Neeraj Kumar Singh, and Marc Pantel. Formally verified architectural patterns of hybrid systems using proof and refinement with Event-B. Science of Computer Programming, 216(??):??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001581.

Murray:2022:SAI

[2667] Yvonne Murray, Martin Sirevåg, Pedro Ribeiro, David A. Anisi, and Morten Mossige. Safety assurance of an industrial robotic control system using hardware/software co-verification. Science of Computer Programming, 216(??):??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642321001593.

Campos:2022:IRF

[2668] Suene Campos, Regivan Santiago, Manuel A. Martins, and Daniel Figueiredo. Introduction to reversal fuzzy switch graph. Science of Computer Programming, 216 (??):??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000090.

Chen:2022:FVC

[2669] Xiaohong Chen, Qianqian Liu, Frédéric Mallet, Qin Li, Shubin Cai, and Zhi Jin. Formally verifying consistency of sequence diagrams for safety critical systems. Science of Computer Programming, 216(??): ??, April 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000107.

Anonymous:2022:Mb

[2670] Anonymous. 1 May 2022. Science of Computer Programming, 217(??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBe

[2671] Anonymous. Editorial Board. Science of Computer Programming, 217 (??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000314.

Coughlin:2022:CNH

[2672] Nicholas Coughlin and Graeme Smith.

Compositional noninterference on hardware weak memory models. Science of Computer Programming, 217 (??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000120.

Castagna:2022:ROT

[2673] Giuseppe Castagna, Victor Lanvin, Mickaël Laurent, and Kim Nguyen.

Revisiting occurrence typing. Science of Computer Programming, 217 (??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000144.

Ceresa:2022:EIT

[2674] Martín A. Ceresa and Mauro J. Jaskelioff. Effectful improvement theory. Science of Computer Programming, 217(??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000259.

Huang:2022:HSA

[2675] Zijie Huang, Zhiqing Shao, Guisheng Fan, Huiqun Yu, Kang Yang, and Ziyi Zhou. HBSniff: a static analysis tool for Java Hibernate object-relational mapping code smell detection. Science of Computer Programming, 217 (??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000119.

Xu:2022:DAR

[2676] Shenglin Xu and Yongjun Wang. Defending against Return-Oriented Programming attacks based on return instruction using static analysis and binary patch techniques. Science of Computer Programming, 217 (??):??, May 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000016.

Anonymous:2022:Jb

[2677] Anonymous. 1 June 2022. Science of Computer Programming, 218(??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBf

[2678] Anonymous. Editorial Board. Science of Computer Programming, 218 (??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000478.

Alashqar:2022:SCM

[2679] Abdelkareem M. Alashqar. Studying the commonalities, mappings and relationships between non-functional requirements using machine learning. Science of Computer Programming, 218(??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000399.

Gleirscher:2022:VSO

[2680] Mario Gleirscher, Radu Calinescu, James Douthwaite, Benjamin Lesage, Colin Paterson, Jonathan Aitken, Rob Alexander, and James Law. Verified synthesis of optimal safety controllers for human-robot collaboration. Science of Computer Programming, 218(??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000429.

Constantinou:2022:SIS

[2681] Eleni Constantinou, Dario Di Nucci, Raula Gaikovina Kula, and Henrique Rocha. Special issue on software health of software ecosystems. Science of Computer Programming, 218 (??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000387.

Lanese:2022:PSI

[2682] Ivan Lanese and Mariusz Rawski. Preface for the special issue of the 12th Conference on Reversible Computation (RC 2020). Science of Computer Programming, 218 (??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000417.

Perkowski:2022:IPC

[2683] Marek Perkowski. Inverse problems, constraint satisfaction, reversible logic, invertible logic and Grover quantum oracles for practical problems. Science of Computer Programming, 218 (??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000089.

Paixao:2022:HLA

[2684] João Paixão, Lucas Rufino, and Paweł Sobociński. High-level axioms for graphical linear algebra. Science of Computer Programming, 218 (??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000247.

Camarao:2022:SIR

[2685] Carlos Camarão and Martin Sulzmann.

Special issue on revised and extended versions of papers presented at the 22nd Brazilian Symposium on Programming Languages (SBLP 2018). Science of Computer Programming, 218(??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000260.

Proenca:2022:SIS

[2686] José Proença and Markus Lumpe.

Special issue on selected papers from the 14th International Conference on Formal Aspects of Component Software (FACS 2017). Science of Computer Programming, 218 (??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000363.

Mery:2022:SPI

[2687] Dominique Méry and Shengchao Qin.

Selected papers from The 13th International Symposium on Theoretical Aspects of Software Engineering 29 July — 1 August 2019, Guilin, China. Science of Computer Programming, 218(??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000375.

Jongmans:2022:PSI

[2688] Sung-Shik Jongmans and Farhad Arbab. Preface — special issue on selected and extended papers from FACS 2019. Science of Computer Programming, 218 (??):??, June 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000405.

Anonymous:2022:Jc

[2689] Anonymous. 1 July 2022. Science of Computer Programming, 219(??):??, July 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBg

[2690] Anonymous. Editorial Board. Science of Computer Programming, 219 (??):??, July 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000594.

Maurel:2022:SIX

[2691] Heloise Maurel, Santiago Vidal, and Tamara Rezk. Statically identifying XSS using deep learning. Science of Computer Programming, 219 (??):??, July 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000430.

Azzeh:2022:VPP

[2692] Mohammad Azzeh, Ali Bou Nassif, Yousef Elsheikh, and Lefteris Angelis. On the value of project productivity for early effort estimation. Science of Computer Programming, 219(??):??, July 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000521.

Sridharan:2022:SPE

[2693] Manu Sridharan. Selected papers from the 2021 European Conference on Object-Oriented Programming. Science of Computer Programming, 219 (??):??, July 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000533.

Aoki:2022:SPI

[2694] Toshiaki Aoki and Qin Li. Selected papers from the 14th international symposium on Theoretical Aspects of Software Engineering. Science of Computer Programming, 219 (??):??, July 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000545.

Anonymous:2022:Ab

[2695] Anonymous. 1 August 2022. Science of Computer Programming, 220 (??):??, August 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBh

[2696] Anonymous. Editorial Board. Science of Computer Programming, 220 (??):??, August 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000673.

Carvalho:2022:PSP

[2697] Gustavo Carvalho and Volker Stolz. Preface — Selected papers from the 23rd Brazilian Symposium on Formal Methods — SBMF 2020. Science of Computer Programming, 220 (??):??, August 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000636.

Liu:2022:MCA

[2698] Chong Liu, Nuno Macedo, and Alcino Cunha. Merging cloned Alloy models with colorful refactorings. Science of Computer Programming, 220 (??):??, August 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000624.

Navarro:2022:SPP

[2699] Marcelo Navarro and Federico Olmedo. Slicing of probabilistic programs based on specifications. Science of Computer Programming, 220(??): ??, August 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000557.

Anonymous:2022:S

[2700] Anonymous. 1 September 2022. Science of Computer Programming, 221 (??):??, September 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBi

[2701] Anonymous. Editorial Board. Science of Computer Programming, 221 (??):??, September 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000843.

Kehrer:2022:SIA

[2702] Timo Kehrer and Fabio Gadducci.

Special issue on application-oriented aspects of graphs and graph transformation (ICGT 2020). Science of Computer Programming, 221(??): ??, September 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000788.

Chiapperini:2022:COE

[2703] Alessio Chiapperini, Marino Miculan, and Marco Peressotti. Computing (optimal) embeddings of directed bigraphs. Science of Computer Programming, 221(??): ??, September 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000752.

Basile:2022:CAL

[2704] Davide Basile and Maurice H. ter Beek. Contract Automata Library. Science of Computer Programming, 221 (??):??, September 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000740.

Bacchiani:2022:JTC

[2705] Lorenzo Bacchiani, Mario Bravetti, Marco Giunti, João Mota, and António Ravara. A Java typestate checker supporting inheritance. Science of Computer Programming, 221(??): ??, September 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000776.

Anonymous:2022:N

[2706] Anonymous. 1 November 2022. Science of Computer Programming, 223 (??):??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBj

[2707] Anonymous. Editorial Board. Science of Computer Programming, 223 (??):??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764232200123X.

Hoey:2022:RIC

[2708] James Hoey and Irek Ulidowski. Reversing an imperative concurrent programming language. Science of Computer Programming, 223(??): ??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764232200106X.

Schewe:2022:BTR

[2709] Klaus-Dieter Schewe and Flavio Ferrarotti. Behavioural theory of reflective algorithms I: Reflective sequential algorithms. Science of Computer Programming, 223(??): ??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000971.

Ueda:2022:DLG

[2710] Yuki Ueda, Takashi Ishio, and Kenichi Matsumoto. DevReplay: Linter that generates regular expressions for repeating code changes. Science of Computer Programming, 223(??): ??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322000909.

Babur:2022:SFM

[2711] Önder Babur, Loek Cleophas, and Mark van den Brand. SAMOS — a framework for model analytics and management. Science of Computer Programming, 223(??): ??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001101.

DeNicola:2022:ART

[2712] Rocco De Nicola, Luca Di Stefano, Omar Inverso, and Aline Uwimbabazi. Automated replication of tuple spaces via static analysis. Science of Computer Programming, 223 (??):??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764232200096X.

Eliott:2022:PLF

[2713] Harold Pancho Eliott and Martin Berger. A program logic for fresh name generation. Science of Computer Programming, 223(??): ??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000934.

Schlatte:2022:AST

[2714] Rudolf Schlatte, Einar Broch Johnsen, Eduard Kamburjan, and S. Lizeth Tapia Tarifa. The ABS simulator toolchain. Science of Computer Programming, 223(??):??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322000946.

Kuehn:2022:PMT

[2715] Eva Maria Kuehn. The Peer-Model tool-chain. Science of Computer Programming, 223(??):??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642322001095.

Fahrenberg:2022:FG

[2716] Uli Fahrenberg and Axel Legay. Featured games. Science of Computer Programming, 223(??):??, November 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642322001071.

Anonymous:2022:D

[2717] Anonymous. 1 December 2022. Science of Computer Programming, 224 (??):??, December 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2022:EBk

[2718] Anonymous. Editorial Board. Science of Computer Programming, 224 (??):??, December 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001368.

Afzal:2022:ECS

[2719] Uzma Afzal, Tariq Mahmood, and Shazia Usmani. Evolutionary Computing to solve product inconsistencies in Software Product Lines. Science of Computer Programming, 224 (??):??, December 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001083.

Gariano:2022:UCS

[2720] Isaac Oscar Gariano, Marco Servetto, and Alex Potanin. Using capabilities for strict runtime invariant checking. Science of Computer Programming, 224(??):??, December 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001113.

Gu:2022:CGS

[2721] Rong Gu, Peter G. Jensen, Cristina Seceleanu, Eduard Enoiu, and Kristina Lundqvist. Correctness-guaranteed strategy synthesis and compression for multi-agent autonomous systems. Science of Computer Programming, 224 (??):??, December 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001277.

terBeek:2022:FFE

[2722] Maurice H. ter Beek, Ferruccio Damiani, Michael Lienhardt, Franco Mazzanti, Luca Paolini, and Giordano Scarso. FTS4VMC: a front-end tool for static analysis and family-based model checking of FTSs with VMC. Science of Computer Programming, 224 (??):??, December 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001125.

Futatsugi:2022:APS

[2723] Kokichi Futatsugi. Advances of proof scores in CafeOBJ. Science of Computer Programming, 224(??): ??, December 1, 2022. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001265.

Anonymous:2023:Ja

[2724] Anonymous. January 2023. Science of Computer Programming, 225(??): ??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2023:EBa

[2725] Anonymous. Editorial Board. Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000047.

Crafa:2023:PSS

[2726] Silvia Crafa, Cosimo Laneve, Giovanni Sartor, and Adele Veschetti. Pacta sunt servanda: Legal contracts in Stipula. Science of Computer Programming, 225(??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001447.

Zhang:2023:SPI

[2727] Min Zhang and Kazuhiro Ogata.

Selected papers from the 15th international symposium on Theoretical Aspects of Software Engineering (TASE 2021). Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001459.

Zhu:2023:FBR

[2728] Chenyang Zhu, Michael Butler, Corina Cirstea, and Thai Son Hoang. A fairness-based refinement strategy to transform liveness properties in Event-B models. Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764232200140X.

Chen:2023:SAL

[2729] Liqian Chen, Dengping Wei, Banghu Yin, and Ji Wang. Static analysis of linear absolute value equalities among variables of a program. Science of Computer Programming, 225(??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322001393.

Audrito:2023:PSI

[2730] Giorgio Audrito, Omar Inverso, and Hugo Torres Vieira. Preface for the special issue on tool papers of the 23rd International Conference on Coordination Models and Languages, COORDINATION 2021. Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322001320.

Bianchini:2023:PQA

[2731] Riccardo Bianchini and Francesco Dagnino. QueryAGT: Asynchronous global types in co-logic programming. Science of Computer Programming, 225(??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322001289.

Ali:2023:CAR

[2732] Muhammad Rizwan Ali, Yngve Lamo, and Violet Ka I. Pun. Cost analysis for a resource sensitive workflow modelling language. Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001290.

Ravelo-Mendez:2023:KPA

[2733] William Ravelo-Méndez, Camilo Escobar-Velásquez, and Mario Linares-Vásquez. Kraken 2.0: a platform-agnostic and cross-device interaction testing tool. Science of Computer Programming, 225(??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001307.

Anonymous:2023:SIS

[2734] Anonymous. Special issue on ‘Software Engineering and Formal Methods’. Science of Computer Programming, 225(??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001484.

Hofmann:2023:CTA

[2735] Till Hofmann and Stefan Schupp.

Controlling timed automata against MTL specifications with TACoS. Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322001319.

Kauffman:2023:LAS

[2736] Sean Kauffman. Log analysis and system monitoring with nfer. Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322001423.

Hojjat:2023:FSE

[2737] Hossein Hojjat and Mieke Massink. Fundamentals of Software Engineering (extended versions of selected papers of FSEN 2021). Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322001460.

Zakeriyan:2023:ATI

[2738] Arvin Zakeriyan, Ramtin Khosravi, Hadi Safari, Ehsan Khamespanah, and Seyede Mehrnaz Shamsabadi. Automated testing of an industrial stock market trading platform based on functional specification. Science of Computer Programming, 225 (??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642322001411.

vanEerd:2023:IMS

[2739] Johri van Eerd, Jan Friso Groote, Pieter Hijma, Jan Martens, Muhammad Osama, and Anton Wijs. Innermost many-sorted term rewriting on GPUs. Science of Computer Programming, 225(??):??, January 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642322001435.

Anonymous:2023:M

[2740] Anonymous. March 2023. Science of Computer Programming, 226(??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2023:EBb

[2741] Anonymous. Editorial Board. Science of Computer Programming, 226 (??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000199.

Azzeh:2023:EPK

[2742] Mohammad Azzeh, Yousef Elsheikh, Ali Bou Nassif, and Lefteris Angelis. Examining the performance of kernel methods for software defect prediction based on support vector machine. Science of Computer Programming, 226(??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001496.

Zeng:2023:CTC

[2743] Jianwei Zeng, Yutong He, Tao Zhang, Zhou Xu, and Qiang Han. CLG-Trans: Contrastive learning for code summarization via graph attention-based transformer. Science of Computer Programming, 226 (??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000072.

Li:2023:LAR

[2744] Xun Li, Lei Liu, Yuzhou Liu, and Huaxiao Liu. A lightweight API recommendation method for App development based on multiobjective evolutionary algorithm. Science of Computer Programming, 226 (??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000096.

Atoum:2023:MKP

[2745] Issa Atoum. Measurement of key performance indicators of user experience based on software requirements. Science of Computer Programming, 226(??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000114.

Khatri:2023:ESC

[2746] Yogita Khatri and Sandeep Kumar Singh. An effective software cross-project fault prediction model for quality improvement. Science of Computer Programming, 226 (??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001514.

Birchler:2023:CES

[2747] Christian Birchler, Nicolas Ganz, Sajad Khatiri, Alessio Gambi, and Sebastiano Panichella. Cost-effective simulation-based test selection in self-driving cars software. Science of Computer Programming, 226 (??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000084.

Kamburjan:2023:DVA

[2748] Eduard Kamburjan, Marco Scaletta, and Nils Rollshausen. Deductive verification of active objects with Crowbar. Science of Computer Programming, 226(??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000102.

Acher:2023:BBU

[2749] Mathieu Acher, Gilles Perrouin, and Maxime Cordy. BURST: Benchmarking uniform random sampling techniques. Science of Computer Programming, 226(??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642322001472.

Gleirscher:2023:QPA

[2750] Mario Gleirscher, Robert Sachtleben, and Jan Peleska. Qualification of proof assistants, checkers, and generators: Where are we and what next? Science of Computer Programming, 226(??):??, March 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000126.

Anonymous:2023:Aa

[2751] Anonymous. April 2023. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2023:EBc

[2752] Anonymous. Editorial Board. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000321.

Chen:2023:PBC

[2753] Jie Chen, Tao Jiang, Dongjin Yu, and Haiyang Hu. Pattern-based circular reference detection in Python. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764232300014X.

Newar:2023:SFD

[2754] Dip Kiran Pradhan Newar, Rui Zhao, Harvey Siy, Leen-Kiat Soh, and Myoungkyu Song. SSDTutor: a feedback-driven intelligent tutoring system for secure software development. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000151.

Zhao:2023:TMS

[2755] Xin Zhao and Jeff Gray. Towards a metrics suite for the complexity analysis of LabVIEW systems models. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000138.

Velazquez-Rodriguez:2023:TCA

[2756] Camilo Velázquez-Rodríguez, Dario Di Nucci, and Coen De Roover. A text classification approach to API type resolution for incomplete code snippets. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000230.

Yang:2023:EAM

[2757] Bo Yang, Haoran Guo, and Huai Liu. Evaluation and assessment of machine learning based user story grouping: a framework and empirical studies. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000254.

Wang:2023:CTC

[2758] Chong Wang, Jingwen Jiang, Maya Daneva, and Marten van Sinderen. CoolTeD: a tool for co-labeling and visual analysis of textual dataset.

Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000229.

Ferdous:2023:EEM

[2759] Raihana Ferdous, Chia kang Hung, Fitsum Kifetew, Davide Prandi, and Angelo Susi. EvoMBT: Evolutionary model based testing. Science of Computer Programming, 227(??):??, April 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000242.

Anonymous:2023:Jb

[2760] Anonymous. June 2023. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2023:EBd

[2761] Anonymous. Editorial Board. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000485.

Sepahvand:2023:UWE

[2762] Reza Sepahvand, Reza Akbari, Behnaz Jamasb, Sattar Hashemi, and Omid Boushehrian. Using word embedding and convolution neural network for bug triaging by considering design flaws. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000278.

Pathade:2023:CMF

[2763] Komal Pathade and Uday Khedker. Computing maximum fixed point solutions over feasible paths in data flow analyses. Science of Computer Programming, 228 (??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000266.

Sasano:2023:TBS

[2764] Isao Sasano and Kwanghoon Choi. A text-based syntax completion method using LR parsing and its evaluation. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000394.

Gautier:2023:PMC

[2765] Thierry Gautier, Paul Le Guernic, Loïc Besnard, and Jean-Pierre Talpin. The polychronous model of computation and Kahn process networks. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000400.

Timm:2023:MSB

[2766] Nils Timm, Josua Botha, and Steven Jordaan. Max-SAT-based synthesis of optimal and Nash equilibrium strategies for multi-agent systems. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232300028X.

Wu:2023:QNT

[2767] Hao Wu. QMaxUSE: a new tool for verifying UML class diagrams and OCL invariants. Science of Computer Programming, 228 (??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000370.

Kargen:2023:DCR

[2768] Ulf Kargén, Ivar Härnqvist, Johannes Wilson, Gustav Eriksson, Evelina Holmgren, and Nahid Shahmehri. desync-cc: a research tool for automatically applying disassembly desynchronization during compilation. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000369.

Giallorenzo:2023:LTG

[2769] Saverio Giallorenzo, Fabrizio Montesi, Marco Peressotti, and Florian Rademacher. LEMMA2Jolie: a tool to generate microservice APIs from domain models. Science of Computer Programming, 228(??):??, June 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000382.

Anonymous:2023:Jc

[2770] Anonymous. July 2023. Science of Computer Programming, 229(??):??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2023:EBe

[2771] Anonymous. Editorial Board. Science of Computer Programming, 229(??):??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000643.

Ssanyu:2023:TLP

[2772] Jackline Ssanyu, Engineer Bainomugisha, and Benjamin Kanagwa. Teaching language processing with the PAMOJA framework. Science of Computer Programming, 229(??):??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000412.

Parolini:2023:SSA

[2773] Francesco Parolini and Antoine Miné.

Sound static analysis of regular expressions for vulnerabilities to denial of service attacks. Science of Computer Programming, 229(??):??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000424.

Erthal:2023:CCE

[2774] Vladimir M. Erthal, Bruno P. de Souza, Paulo Sérgio M. dos Santos, and Guilherme H. Travassos. Characterization of continuous experimentation in software engineering: Expressions, models, and strategies. Science of Computer Programming, 229(??):??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000436.

Liang:2023:SIT

[2775] Peng Liang and Nikolaos Tsantalis. Special issue on Tools at the 29th IEEE International Conference on Software Analysis, Evolution and Reengineering (SANER 2022). Science of Computer Programming, 229(??):??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000552.

Felderer:2023:ISQ

[2776] Michael Felderer, Valentina Lenarduzzi, Fabio Palomba, and Fabiano Pecorelli. Introduction to the Software Quality for Artificial Intelligence (SQA4AI) special issue. Science of Computer Programming, 229(??):??, July 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000576.

Anonymous:2023:Ab

[2777] Anonymous. August 2023. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2023:EBf

[2778] Anonymous. Editorial Board. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232300103X.

Liu:2023:VGH

[2779] Kenan Liu, Khaled Mahmoud, Joonhwan Yoo, and Yu David Liu. Vincent: Green hot methods in the JVM. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000448.

Bernardez:2023:ESE

[2780] Beatriz Bernárdez, José Ignacio Panach, José A. Parejo, Amador Durán, Natalia Juristo, and Antonio Ruiz-Cortés. An empirical study to evaluate the impact of mindfulness on helpdesk employees. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232300059X.

Chondamrongkul:2023:SEA

[2781] Nacha Chondamrongkul and Jing Sun. Software evolutionary architecture: Automated planning for functional changes. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000606.

Singh:2023:GSP

[2782] Robbert Gurdeep Singh and Christophe Scholliers. Gaiwan: a size-polymorphic typesystem for GPU programs. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000710.

Slivka:2023:TSA

[2783] Jelena Slivka, Nikola Luburić, Simona Prokić, Katarina-Glorija Grujić, Aleksandar Kovačević, Goran Sladić, and Dragan Vidaković. Towards a systematic approach to manual annotation of code smells. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000813.

Moseler:2023:CFA

[2784] Oliver Moseler and Stephan Diehl. The CodeSparks Framework — augmenting source code with glyph-based visualizations. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642323000801.

Bucchiarone:2023:GMB

[2785] Antonio Bucchiarone, Maxime Savary-Leblanc, Xavier Le Pallec, Jean-Michel Bruel, Antonio Cicchetti, Jordi Cabot, and Sébastien Gérard. Gamifying model-based engineering: the PapyGame tool. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000564.

Gomez-Abajo:2023:WET

[2786] Pablo Gómez-Abajo, Esther Guerra, and Juan de Lara. Wodel-Edu: a tool for the generation and evaluation of diagram-based exercises. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000825.

Cano-Genoves:2023:VTF

[2787] C. Cano-Genoves, E. Insfran, and S. Abrahão. VeGAn-Tool: a fuzzy-logic approach for value-based goal model analysis. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642323000837.

Ghosh:2023:MMA

[2788] Bineet Ghosh and Étienne André.

MoULDyS: Monitoring of autonomous systems in the presence of uncertainties. Science of Computer Programming, 230(??):??, August 2023.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642323000588.

Renkin:2023:MMR

[2789] Florian Renkin, Philipp Schlehuber-Caissier, Alexandre Duret-Lutz, and Adrien Pommellet. The Mealymachine reduction functions of Spot. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000771.

Soldani:2023:YEF

[2790] Jacopo Soldani, Stefano Forti, and Antonio Brogi. yRCA: an explainable failure root cause analyser. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000795.

Hasrat:2023:TDH

[2791] Imran Riaz Hasrat, Peter Gjøl Jensen, Kim Guldstrand Larsen, and Jirí Srba. A toolchain for domestic heatpump control using Uppaal Stratego. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000692.

Gobert:2023:MME

[2792] Maxime Gobert, Loup Meurice, and Anthony Cleve. Modeling, manipulating and evolving hybrid polystores with HyDRa. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000540.

Galeano:2023:MAM

[2793] Iris Galeano, Mauricio Merín, Magalí González, and Luca Cernuzzi. A methodological approach for mobile applications development: MethApp4Mob. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000680.

Moharil:2023:TTB

[2794] Ambarish Moharil and Arpit Sharma.

TABASCO: a transformer based contextualization toolkit. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232300076X.

Hirata:2023:PLH

[2795] Michikazu Hirata, Yasuhiko Minamide, and Tetsuya Sato. Program logic for higher-order probabilistic programs in Isabelle/HOL. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000758.

Larsen:2023:ABB

[2796] Kim G. Larsen, Anders Mariegaard, Stefan Schmid, and Jirí Srba. All-Synth: a BDD-based approach for network update synthesis. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000746.

Klikovits:2023:FLE

[2797] Stefan Klikovits, Ezequiel Castellano, Ahmet Cetinkaya, and Paolo Arcaini. Frenetic-lib: an extensible framework for search-based generation of road structures for ADS testing. Science of Computer Programming, 230(??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000783.

Humeniuk:2023:ASB

[2798] Dmytro Humeniuk, Foutse Khomh, and Giuliano Antoniol. AmbieGen: a search-based framework for autonomous systems testing. Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000722.

Campos:2023:SPB

[2799] Sérgio Campos and Marius Minea. Selected papers from the Brazilian Symposium on Formal Methods (SBMF 2021). Science of Computer Programming, 230(??):??, August 2023.

CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642323000734.

Lee:2023:PSI

[2800] Jaejoon Lee and Tomoji Kishi. Preface for the special issue on tool papers of the 25-^{th} International Systems and Software Product Line Conference (SPLC 2021). Science of Computer Programming, 230 (??):??, August 2023. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000709.

Anonymous:2024:Ja

[2801] Anonymous. January 2024. Science of Computer Programming, 231(??): ??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2024:EBa

[2802] Anonymous. Editorial Board. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001235.

Banach:2024:CHE

[2803] Richard Banach. Core Hybrid Event-B III: Fundamentals of a reasoning framework. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642323000849.

Huang:2024:EPO

[2804] Wen ling Huang, Niklas Krafczyk, and Jan Peleska. Exhaustive property oriented model-based testing with symbolic finite state machines. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000874.

Mu:2024:MTC

[2805] Xiangyu Mu, Lei Liu, Peng Zhang, Jingyao Li, and Hui Li. A method of test case set generation in the commutativity test of reduce functions. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000886.

vandenBerg:2024:FRM

[2806] Birthe van den Berg, Tom Schrijvers, James McKinna, and Alexander Vandenbroucke. Forward- or reverse-mode automatic differentiation: What’s the difference? Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000928.

Zhao:2024:HAA

[2807] Yaxin Zhao, Lina Gong, Wenhua Yang, and Yu Zhou. How accessibility affects other quality attributes of software? A case study of GitHub.

Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001090.

Bucchiarone:2024:PPG

[2808] Antonio Bucchiarone, Tommaso Martorella, Davide Frageri, and Diego Colombo. PolyGloT: a personalized and gamified eTutoring system for learning modelling and programming skills. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642323000850.

Harbo:2024:APE

[2809] Sean Kristian Remond Harbo, Emil Palmelund Voldby, Jonas Madsen, and Michele Albano. ACSmt: a plugin for Eclipse Papyrus to model systems of systems. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000904.

Smajevic:2024:COW

[2810] Muhamed Smajevic, Syed Juned Ali, and Dominik Bork. CM2KG^{cloud}

— an open web-based platform to transform conceptual models into knowledge graphs. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000898.

DiSandro:2024:MFM

[2811] Alessio Di Sandro, Logan Murphy, Torin Viger, and Marsha Chechik. MMINT-A: a framework for model-based safety assurance. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000862.

Lopez:2024:MLD

[2812] José Antonio Hernández López, Javier Luis Cánovas Izquierdo, and Jesús Sánchez Cuadrado. ModelSet: a labelled dataset of software models for machine learning. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323000916.

Yigitbas:2024:GGB

[2813] Enes Yigitbas, Maximilian Schmidt, Antonio Bucchiarone, Sebastian Gottschalk, and Gregor Engels. GaMoVR: Gamification-based UML learning environment in virtual reality. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001119.

Giner-Miguelez:2024:DDD

[2814] Joan Giner-Miguelez, Abel Gómez, and Jordi Cabot. DescribeML: a dataset description tool for machine learning. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001120.

Bortin:2024:SDI

[2815] Maksym Bortin. Structured development of implementations for divide-and-conquer specifications. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232300093X.

Kauffman:2024:CEN

[2816] Sean Kauffman and Martin Zimmermann. The complexity of evaluating nfer. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642323000941.

Kiselyov:2024:GCH

[2817] Oleg Kiselyov. Generating C: Heterogeneous metaprogramming system description. Science of Computer Programming, 231(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167642323000977.

Ameen:2024:AUF

[2818] Mahmudul Faisal Al Ameen, Naoki Kobayashi, and Ryosuke Sato. Asynchronous unfold/fold transformation for fixpoint logic. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000965.

Santana:2024:EER

[2819] Railana Santana, Luana Martins, Tássio Virgínio, Larissa Rocha, Heitor Costa, and Ivan Machado. An empirical evaluation of RAIDE: a semiautomated approach for test smells detection and refactoring. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323000953.

Audrito:2024:FAT

[2820] Giorgio Audrito and Gianluca Torta. FCPP to aggregate them all. Science of Computer Programming, 231 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001089.

Anonymous:2024:Jb

[2821] Anonymous. January 2024. Science of Computer Programming, 232(??): ??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2024:EBb

[2822] Anonymous. Editorial Board. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001466.

Fernandes:2024:SFM

[2823] Herbert Rausch Fernandes, Giovanni Freitas Gomes, Antonio Carlos Pinheiro de Oliveira, and Sérgio Vale Aguiar Campos. Stochastic formal model of PI3K/mTOR pathway in Alzheimer’s disease for drug repurposing: an evaluation of rapamycin, LY294002, and NVP-BEZ235. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001107.

Driessen:2024:SAT

[2824] S. W. Driessen, D. Di Nucci, D. A.

Tamburri, and W. J. van den Heuvel. SolAR: Automated test-suite generation for solidity smart contracts. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001181.

Aldalur:2024:WAS

[2825] Iñigo Aldalur. Web Augmentation:

a systematic mapping study. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001272.

Mammar:2024:FAC

[2826] Amel Mammar, Meriem Belguidoum, and Saddam Hocine Hiba. A formal approach for the correct deployment of cloud applications. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001302.

Liu:2024:ETR

[2827] Shuqi Liu, Yu Zhou, Longbing Ji, Tingting Han, and Taolue Chen. Enhancing test reuse with GUI events deduplication and adaptive semantic matching. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S016764232300134X.

Daniel:2024:AMD

[2828] Gwendal Daniel and Jordi Cabot. Applying model-driven engineering to the domain of chatbots: the Xatkit experience. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642323001144.

Buschhaus:2024:LLA

[2829] Constantin Buschhaus, Arkadii Gerasimov, Jörg Christian Kirchhof, Judith Michael, Lukas Netz, Bernhard Rumpe, and Sebastian Stüber. Lessons learned from applying modeldriven engineering in 5 domains: the success story of the MontiGem generator framework. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001156.

Sadovykh:2024:IAM

[2830] Andrey Sadovykh, Bilal Said, Dragos Truscan, and Hugo Bruneliere. An iterative approach for model-based requirements engineering in large collaborative projects: a detailed experience report. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001296.

Aceto:2024:MTL

[2831] Luca Aceto, Antonis Achilleos, Duncan Paul Attard, Léo Exibard, Adrian Francalanza, and Anna Ingólfsdóttir. A monitoring tool for linear-time μHML. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001132.

Arshad:2024:XAT

[2832] Hamed Arshad, Ross Horne, Christian Johansen, Olaf Owe, and Tim A. C. Willemse. XACML2mCRL2:

Automatic transformation of XACML policies into mCRL2 specifications. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001284.

Yohannis:2024:ECM

[2833] Alfa Yohannis, Dimitris Kolovos, and Antonio García-Domínguez. Exploring complex models with picto web. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001193.

Basciani:2024:TTS

[2834] Francesco Basciani, Juri Di Rocco, Ludovico Iovino, and Alfonso Pierantonio. TyphonML: Tool support for hybrid polystores. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001260.

Jouault:2024:APT

[2835] Frédéric Jouault, Valentin Besnard, Matthias Brun, Théo Le Calvar, Fabien Chhel, Mickaël Clavreul, Jérome Delatour, Maxime Méré, Matthias Pasquier, and Ciprian Teodorov. AnimUML: a practical tool for partial model animation and analysis. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001326.

Mahe:2024:DOS

[2836] Erwan Mahe, Christophe Gaston, and Pascale Le Gall. Denotational and operational semantics for interaction languages: Application to trace analysis. Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001168.

Masuda:2024:PGM

[2837] Masahiro Masuda and Yukiyoshi Kameyama. Program generation meets program verification: a case study on number-theoretic transform. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232300117X.

Jakobs:2024:PSI

[2838] Marie-Christine Jakobs, Einar Broch Johnsen, Eduard Kamburjan, and Manuel Wimmer. Preface for the special issue on “Fundamental Approaches to Software Engineering” (FASE 2022). Science of Computer Programming, 232(??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001375.

Anonymous:2024:PSIa

[2839] Anonymous. Preface for the special issue on NLBSE’22: Natural Language-based Software to support software engineering processes. Science of Computer Programming, 232 (??):??, January 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001314.

Anonymous:2024:Ma

[2840] Anonymous. March 2024. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2024:EBc

[2841] Anonymous. Editorial Board. Science of Computer Programming, 233 (??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000066.

Tvardovskii:2024:TIC

[2842] Aleksandr Tvardovskii, Khaled El-Fakih, and Nina Yevtushenko. Testing and incremental conformance testing of timed state machines. Science of Computer Programming, 233 (??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001351.

Chen:2024:WMR

[2843] Jie Chen, Ziyuan Zhang, Dongjin Yu, and Haiyang Hu. What makes a real change in software performance? An empirical study on analyzing the factors that affect the triagement of performance change points. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001508.

Kaluarachchi:2024:WML

[2844] Thisaranie Kaluarachchi and Manjusri Wickramasinghe. WebDraw: a machine learning-driven tool for automatic website prototyping. Science of Computer Programming, 233 (??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001387.

Reynolds:2024:RAL

[2845] Conor Reynolds and Rosemary Monahan. Reasoning about logical systems in the Coq proof assistant. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001363.

Krijnen:2024:TCS

[2846] Jacco O. G. Krijnen, Manuel M. T.

Chakravarty, Gabriele Keller, and Wouter Swierstra. Translation certification for smart contracts. Science of Computer Programming, 233 (??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001338.

Munoz:2024:PSI

[2847] Paula Muñoz, Steffen Zschaler, and Richard F. Paige. Preface to the special issue on success stories in model driven engineering. Science of Computer Programming, 233 (??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001545.

James:2024:ORD

[2848] Phillip James, Faron Moller, and Filippos Pantekis. OnTrack: Reflecting on domain specific formal methods for railway designs. Science of Computer Programming, 233 (??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001399.

Nistala:2024:IER

[2849] Padmalata Nistala, Asha Rajbhoj, Vinay Kulkarni, Sapphire Noronha, and Ankit Joshi. An industrial experience report on modelbased, AI-enabled proposal development for an RFP/RFI. Science of Computer Programming, 233 (??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001405.

Kurtev:2024:MBC

[2850] Ivan Kurtev, Jozef Hooman, Mathijs Schuts, and Daan van der Munnik. Model based component development and analysis with ComMA. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001491.

delaVara:2024:PSP

[2851] Jose Luis de la Vara, Monalessa Barcellos, Marcela Genero, and Renata Guizzardi. Preface for “Selected papers from the 25th Ibero–American Conference on Software Engineering (CIbSE 2022)”. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001417.

Anonymous:2024:PSIb

[2852] Anonymous. Preface for the special issue on SBST’22: Search-Based Software Testing — Tools. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001533.

Ji:2024:TLT

[2853] Wangjie Ji, Lili Bo, Yanchi Yuan, and Xiaobing Sun. TDFix: a lightweight tool for fixing deadlocks based on templates. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001557.

Arias:2024:SAP

[2854] Jaime Arias, Kyungmin Bae, Carlos Olarte, Peter Csaba Ölveczky, Laure Petrucci, and Fredrik Rømming. Symbolic analysis and parameter synthesis for networks of parametric timed automata with global variables using Maude and SMT solving. Science of Computer Programming, 233(??):??, March 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642323001569.

Anonymous:2024:Mb

[2855] Anonymous. May 2024. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2024:EBd

[2856] Anonymous. Editorial Board. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016764232400025X.

Hong:2024:VMP

[2857] Weijiang Hong, Zhenbang Chen, Yufeng Zhang, Hengbiao Yu, Yide Du, and Ji Wang. Verification of message-passing uninterpreted programs. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001570.

Macias:2024:MFU

[2858] José A. Macías and Clemente R.

Borges. Monitoring and forecasting usability indicators: a business intelligence approach for leveraging user-centered evaluation data. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001594.

Yang:2024:CCC

[2859] Huiwen Yang, Xiguo Gu, Xiang Chen, Liwei Zheng, and Zhanqi Cui. CrossFuzz: Cross-contract fuzzing for smart contract vulnerability detection. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642323001582.

vandenBerg:2024:FHO

[2860] Birthe van den Berg and Tom Schrijvers. A framework for higherorder effects and handlers. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000091.

Ignaim:2024:IEU

[2861] Karam Ignaim, João M. Fernandes, and André L. Ferreira. An industrial experience of using reference architectures for mapping features to code. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000108.

Ait:2024:HEP

[2862] Adem Ait, Javier Luis Cánovas Izquierdo, and Jordi Cabot. HF-Community: an extraction process and relational database to analyze Hugging Face Hub data. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000029.

Wu:2024:ITC

[2863] Xiaoxue Wu, Yizeng Gu, Lidan Lin, Wei Zheng, and Xiang Chen. ISTA+: Test case generation and optimization for intelligent systems based on coverage analysis. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000017.

Ahmadi:2024:OPM

[2864] Sharar Ahmadi, Brijesh Dongol, and Matt Griffin. Operationally proving memory access violations in Isabelle/ HOL. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S016764232400011X.

Savary-Leblanc:2024:MAR

[2865] Maxime Savary-Leblanc, Xavier Le Pallec, and Sébastien Gérard. The Modeling Assistant Recommender: a UML class diagram recommender system. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000133.

DiRuscio:2024:PSI

[2866] Davide Di Ruscio, Jessie Galasso, and Richard Paige. Preface for the special issue on tools and demonstrations in model-driven engineering. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000145.

Ait-Ameur:2024:ITS

[2867] Yamine Ait-Ameur and Florin Craciun.

Introduction to the TASE 2022 Special issue. Science of Computer Programming, 234(??):??, May 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000157.

Anonymous:2024:Jc

[2868] Anonymous. July 2024. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:2024:EBe

[2869] Anonymous. Editorial Board. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000480.

Dong:2024:NCL

[2870] Haozhen Dong, Hongmin Ren, Jialiang Shi, Yichen Xie, and Xudong Hu. Neighborhood contrastive learningbased graph neural network for bug triaging. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0167642324000169.

Cao:2024:DPI

[2871] Yulu Cao, Zhifei Chen, Xiaowei Zhang, Yanhui Li, Lin Chen, and Linzhang Wang. Diagnosis of package installation incompatibility via knowledge base. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000212.

Dong:2024:MIO

[2872] Yukun Dong, Xiaotong Cheng, Yufei Yang, Lulu Zhang, Shuqi Wang, and Lingjie Kong. A method to identify overfitting program repair patches based on expression tree. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000285.

Kuhail:2024:WRA

[2873] Mohammad Amin Kuhail, Sujith Samuel Mathew, Ashraf Khalil, Jose Berengueres, and Syed Jawad Hussain Shah. “Will I be replaced?” Assessing Chat-GPT’s effect on software development and programmer perceptions of AI tools. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000340.

Thierry-Mieg:2024:MCE

[2874] Yann Thierry-Mieg, Etienne Renault, Emmanuel Paviot-Adet, and Denis Poitrenaud. A model-checker exploiting structural reductions even with stutter sensitive LTL. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000121.

Damiani:2024:PSI

[2875] Ferruccio Damiani, David Eyers, and Anna Philippou. Preface for the special issue on tool papers of the 17th International Federated Conference on Distributed Computing Techniques, DisCoTec 2022. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000194.

Giudice:2024:STR

[2876] Nicola Del Giudice, Lorenzo Matteucci, Michela Quadrini, Aniqa Rehman, and Michele Loreti. Sibilla: a tool for reasoning about collective systems. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000182.

Becker:2024:ESS

[2877] Pablo Becker, María Fernanda Papa, and Luis Olsina. Exploratory study on the syntactic and semantic consistency of terms in project management glossaries to provide recommendations for a project management ontology. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167642324000170.

Kang:2024:NHS

[2878] Byeongjee Kang and Kyungmin Bae. Narrowing and heuristic search for symbolic reachability analysis of concurrent object-oriented systems. Science of Computer Programming, 235(??):??, July 2024. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167642324000200.

Dijkstra:1960:RPb

[2879] Edsger W. Dijkstra. Recursive programming. Numerische Mathematik, 2 (1):312–318, December 1960. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic). URL http:

//gdzdoc.sub.uni-goettingen.de/
sub/digbib/loader?did=D196346. See

[2176].

Jones:1990:EES

[2880] Neil D. Jones, editor. ESOP ’90:

3rd European Symposium on Programming, Copenhagen, Denmark, May 15– 18, 1990: proceedings, volume 432 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 1990. CODEN LNCSD9. ISBN 3-540-52592-0 (Berlin), 0-387-52592-0 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.6 .E976 1990.

Anonymous:1991:ESP

[2881] Anonymous, editor. 3rd European Symposium on Programming — ESOP ’90, volume 17(1–3) of Science of Computer Programming. Elsevier, Amsterdam, The Netherlands, December 1991. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic). Also published as [2880].

Krieg-Bruckner:1994:MSD

[2882] B. Krieg-Brückner, editor. Methods of software design: techniques and applications. Selected papers of the 4th European Symposium on Programming, February 26–28, 1992, Rennes, France, volume 22(3) of Science of Computer Programming. Elsevier, Amsterdam, The Netherlands, June 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1994:PCF

[2883] Anonymous, editor. Proceedings of the Colloquium on Formal Aspects of Software Engineering, volume 23(2–3) of Science of Computer Programming. Elsevier, Amsterdam, The Netherlands, December 1994. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Anonymous:1995:PES

[2884] Anonymous, editor. Proceedings of the 1994 5th European Symposium on Programming, volume 25(2–3) of Science of Computer Programming. Elsevier, Amsterdam, The Netherlands, December 1995. CODEN SCPGD4. ISBN ???? ISSN 0167-6423 (print), 1872-7964 (electronic). LCCN ????

Moller:1995:MPC

[2885] B. Möller, editor. Mathematics of program construction. Selected papers from the 3rd international conference on mathematics of program construction, July 17–21, 1995, Kloster Irsee, Germany, volume 26(1–3) of Science of Computer Programming. Elsevier, Amsterdam, The Netherlands, May 1995. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).

Madey:1997:FSF

[2886] J. Madey, A. Tarlecki, and W. M. Turski, editors. Formal specifications: foundations, methods, tools and applications. FMTA ’95. Konstancin near Warsaw, Poland, May 29–31, 1995, volume 28(2–3) of Science of Computer Programming. Elsevier, Amsterdam, The Netherlands, April 1997. CODEN SCPGD4. ISSN 0167-6423 (print), 1872-7964 (electronic).